A Complete Bibliography of Publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches

sur les pêcheries du Canada (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

−1 [1383, 1691]. −10 [1673]. −23 [2580, 2197]. −4 [1383, 1691]. −5 [1607, 1384]. ^{−} [2585]. ^{131} [2522]. ^{14} [2075, 2180, 1733, 2639, 2193, 2037, 907, 2570, 300, 812, 377]. ^{203} [2570]. ^{3} [2075, 2645]. ^{65} [1660, 6]. _{12} [890]. _{2} [2452, 2956, 1092]. _{3} [2570]. a [2472, 753, 1290, 1493, 1895, 1903, 1399]. α [1032, 2581, 336]. β [214]. : [401, 1399]. Δ [2914, 1855, 2247, 948]. F_{1} [118, 246, 2105]. F_{3} [1848]. F_{5} [1848]. K [2899]. n [1547]. R [1909, 2899]. (R) [2663, 1992, 2673]. sn [2155]. × [954, 1280, 1423, 118, 1283, 1848, 1974, 2105, 5].

-Amino [1855, 2247]. -Aminolevulinate [948]. -Butyl [1547].

-Glucuronidase [214]. -glycero-3-phosphorylcholine [2155]. -Methyltestosterone [336, 1032]. -Selection [2899]. -Strategists [2899]. -Tocopherol [2581].

1 [2948]. 1-Methyladenosine [1149]. 100-Yr-Old [2186]. 104th [1818]. 11-Ketotestosterone [336]. 11th [2865]. 12-hr [922]. 1221 [662]. 13th [2072]. 140 [1561]. 145W [614]. 160^{o} [257]. 1971 [557]. 1974 [1623, 1818, 2216]. 1976 [2654, 2865]. 1980s [2651].

2 [2666, 1327, 2922, 2795]. 2-Methylisoborneol [1823]. 200-Mile [1810]. 214 [2666]. 222 [1259, 2003, 290, 1591, 97]. 227 [1092, 1085]. 24-Hour [1972]. 25-day [922].

3-Trifluoromethyl-4-Aminophenol [931]. 3-Trifluoromethyl-4-Nitrophenol [1147, 1732, 1674, 2208]. 3-Year [2625]. 30-Minute [1793].

4 [247, 1569, 1327]. 4’-bis [247]. 4-D [1327]. 450 [2619].

5 [436, 225, 1569]. 5-Chlorouracil [1770]. 5-Dichloro-4’-Nitrosalicylanilide [1569]. 5’-Monophosphate [436]. 5’-Nucleotidase [225]. 50N [614].

6 [2568, 2777]. 6-Month [2618]. 6-Year [2839]. 6/76 [2568].

75^{o} [1340, 1336]. ’76 [2654, 2568].

= [1269, 778].

île [649]. Îles [2226]. Îles-de-la-Madeleine [2226].

A. [2503, 55]. A.D. [672, 673]. Abdominally [2223, 2488]. Abductor [1759]. Aberrant [512]. Ability [613, 1608, 5, 2887]. Abiotic [2152]. Ablation [2064, 1108, 2774]. Abnormal [559]. Above [1519, 1413, 223]. Above-ground [1519]. Abraliopsis [359]. Abramis [2885]. Abrupt [1463]. Absence [2180, 2142, 1725, 1049, 2359, 2614]. Absolute [23]. Absorption [2921, 84]. Abundance [2743, 220, 1636, 2411, 1295, 572, 2400, 2333, 323, 1360, 1751, 73, 545, 2436, 1364, 264, 60, 1368, 2526, 883, 666, 465, 2689, 254, 1664, 1144, 304, 1562, 1629, 1387, 2605, 2316, 1294, 1346, 1896, 1047]. Abyssal [488]. Academy [626]. acanthias [2633, 1589, 793, 740, 2114, 2171, 2440, 2356, 834, 1512, 2043, 2127, 2853]. Acanthocephala [2219]. Acanthocephalan [204]. Acartia [2716]. Accelerate [1108]. Accelerated [1384]. Accelerates [2844]. Acceleration [2725]. Acceptability [2658]. Access [1723, 2879]. Acclimated [1054].

Acclimation [143, 1565, 606, 2815, 1424]. Accommodated [1917]. Accommodation [1111]. Accommodative [1036]. Accumulation [955, 2168, 1725, 552, 386, 1992, 858, 938, 172, 2143, 1854, 781, 1451, 1660, 1883]. Accurate [236, 453]. Acetylcholinesterase [1759]. Acetylene [2233]. Achievements [1676]. Acid [189, 32, 353, 868, 444, 1401, 1795, 1733, 2267, 2568, 93, 860, 2299, 2362, 1556, 452, 885, 2917, 2457, 2271, 2577, 64, 9, 1855, 2247, 1860, 1581, 52, 1675, 1566, 420, 888, 1558, 1551, 541, 65, 12, 1118]. Acid-Soluble [9]. Acid-Stressed [1551]. Acidic [2632]. Acidification [731, 2771, 812]. Acidified [1251]. Acids [56, 1476, 1298, 1477, 862, 860, 1263, 1568, 2155, 848, 1536, 1410, 1367, 1986, 2449, 792, 1312, 580, 1109, 2175]. Acipenser [949]. Acknowledgments [1872]. Acoustic [2975, 1666, 2682, 2684, 1282, 2908, 1805, 1567]. Acoustical [326]. Acquired [109]. Acquisition [908]. ACTH [1388, 137]. Action [1468, 79, 2497]. Activated [2514]. Activation [1749, 2740, 1989]. Active [351, 2045]. Activities [2735, 2786, 1586]. Activity [570, 1316, 1735, 351, 1388, 27, 2952, 1826, 899, 1855, 2247, 1122, 948, 470, 1325, 609, 2746, 300, 1394, 2348, 2959, 1313, 541, 1022, 2491, 2487, 573, 589, 2370]. Actual [2699]. aculeatus [1379, 14, 96, 1049, 79, 413]. Acute [1849, 955, 1968, 2163, 787, 1099, 759, 606, 1434, 1972, 2849, 1507, 1732, 1860, 1049, 1807, 1576, 1604, 1987, 2450, 2722, 1729, 726, 1798, 1452, 1984, 2692, 2304, 745, 2034, 1742]. acutorostrata [1624]. acutus [2713, 2712]. acyl [2155, 2155]. Acyl-CoA [2155]. Acyltransferase [2155]. Adams [499]. Adaptation [2701, 2515, 195, 2079]. Adaptations [2383, 803]. Adapted [2142]. Adaptive [583, 533, 2837, 1836]. Addendum [1082]. Addition [530, 1085, 2506]. Additional [293, 310, 334, 361, 388, 407, 430, 456, 481, 543, 582, 601, 623, 644, 669, 725, 751, 776, 800, 825, 852, 934, 177]. Additions [495, 1153]. Additive [1423]. Additives [2234]. Address [1168]. Adductor [64, 9, 396, 858, 1384, 1103, 1659]. Adenosine [1826, 1583, 1095]. Adenosinetriphosphatase [589]. Adjacent [2299, 502, 728, 877, 2972, 2372, 1907]. Adjusted [1160]. Adjustment [202, 1374, 1538, 2699]. Adjustments [2877]. Adjuvant [1466]. Administered [1466]. Administration [2645, 1743]. Adopted [1239]. Adrenal [339, 612]. Adrenal-Cortical [339]. Adsorption [2740]. adspersus [2747]. Adult [461, 2437, 733, 1407, 1071, 708, 801, 785, 1148, 1388, 437, 273, 2907, 102, 710, 2064, 337, 1054, 611, 2133, 816, 1266, 2379, 1377, 2920, 2324]. Adults [1472, 1907]. Advantages [451]. Advective [1890]. Advice [1171]. aeglefinus [1414, 311]. aequispina [86]. Aerated [738, 1984]. Aeration [817]. Aerial [2680]. Aerobic [537, 617]. Aerococcus [1596]. Aeromonads [1397]. Aeromonas [164, 596, 2196, 512, 2334, 2503, 109, 598, 1437, 1465, 1466, 1909, 2913, 2236]. Aetideidae [2013]. Affairs [1703]. Affected [2142, 736, 891]. Affecting [2371, 2939, 1713, 1822, 1350, 2271, 2172, 1112, 2386, 641, 556, 1952, 2051]. Afferent [1473]. affinis [204, 1809, 379, 472]. Africa [2886, 1210, 1813, 488, 1626]. African [1186]. Afrique [1210]. After [2608, 2627, 1839, 2645, 748, 1116, 2599, 2601, 210, 2618, 2622, 2774, 1312, 1406, 258, 287, 1732, 180, 1768, 1050]. Again [1418]. Against [164, 1925, 109, 869, 1488]. Agar [1437]. Agarophyte [1949]. AGARS [2761]. Age [958, 1370, 929, 116, 954, 2435, 960, 2358, 2782, 341, 2636, 2017, 2650, 244, 2941, 1248, 473, 1045, 905, 2171, 2505, 98, 1512, 2218, 2816, 2727, 1788, 1533, 2896, 1415, 2523, 1385, 1475, 1462, 1017, 2384, 1287, 893, 2532, 45, 2925, 2926]. Age-0 [2384]. Age-Frequency [2532]. Age-Groups [2505, 2896]. Age-Specific [2636]. Aged [2255]. Agent [2696]. Agents [1851]. Agglutinating [109, 869]. Agglutination [2164]. Agglutinins [590, 2913]. aggregata [1277]. Aggregation [1320, 2971]. Aggregations [714]. Aggression [263]. Aggressive [664, 1598]. Aging [819, 638]. Aglantha [945]. Agmenellum [2268]. Agonid [113]. Agonidae [126]. Agonistic [1905]. Agricultural [2827]. AHH [2614]. Aholehole [571]. Air [16, 142, 2133, 1975, 2545, 1619]. Air-Dried [16]. Air-Supersaturated [2133, 1975, 2545]. Aishihik [1532, 2113, 1532]. al [1212]. Alaska [2874, 929, 1154, 946, 2182, 116, 1315, 71, 610, 2824, 943, 2940, 1968, 1594, 2509, 239, 355, 1470, 291, 1044, 2317, 199, 1327, 1856, 577, 2026, 2875, 2473]. Alaskan [603, 631, 2025, 1843]. Alberta [159, 31, 1479, 722, 2104, 2264, 621, 947, 1357, 719, 1281, 1311]. Albino [164]. Albumin [2641]. alcicornis [1323]. Alcohols [1477, 2154]. Aldrin [368, 2075]. Aleutian [489]. aleutianus [252]. Alevins [141, 1495, 288, 1588, 2486, 2170, 2559]. Alewife [2070, 2542, 1422, 2209]. Alewives [646, 452, 1284, 186, 2348]. Alga [2268, 2492, 781]. Algae [2683, 2734, 39, 991, 1153, 567, 2229, 2954, 1310, 1988, 2889, 2152, 1967, 880, 569, 1823, 2887, 1145, 2586]. Algal [1401, 1681, 593, 916]. Algonquin [2556]. Alien [201]. Alimentary [2713]. Alkaline [539]. Alleles [1299]. Allochthonous [1334, 2367]. Allometry [631, 1520]. Allosmerus [520]. Allosteric [1749]. Allothunnus [1964]. Allows [1088]. Alone [1434]. Along [1832, 2292, 2191, 2894, 648, 840, 637, 257, 2487]. Alosa [646, 1590, 866, 785, 1148, 1447, 452, 2042, 1431, 1284, 2070, 2542, 440, 2369, 2732, 795, 1422, 2209, 2348, 24, 999]. alosoides [1905]. Alpha [101]. Alpine [312]. alpinus [1594, 2509, 2736, 980, 2112, 291, 1044, 1413, 1856, 1404, 2745]. Alteration [1593]. Alterations [2612, 4]. Alternation [341]. Alternative [2646, 1194, 812]. Alternatives [2964, 2871, 1201, 2668, 2953, 2965]. alterniflora [1519]. altivelis [1944]. alutus [2958, 323, 714, 1248, 101, 219, 893, 1305]. Alvariño [599]. am [2483]. Ambicoloration [517]. Ambient [891, 2270, 1987, 2379]. Ambloplites [2441]. Amchitka [2026]. Ameliorate [2887]. America [1815, 2517, 1260, 2796, 14, 400, 1250, 1162, 591, 2768, 672, 1629, 932, 1113]. American [2777, 1134, 1818, 1066, 1724, 2587, 272, 900, 2823, 160, 772, 2448, 2391, 1590, 1668, 866, 185, 2272, 2479, 1012, 785, 1148, 1447, 2212, 2836, 2456, 2551, 928, 2042, 1374, 1431, 2665, 1116, 652, 186, 2893, 440, 2369, 2732, 838, 358, 2044, 2064, 2150, 1396, 161, 835, 1371, 1500, 1438, 1748, 854, 514, 2740, 1602, 1363, 1361, 1058, 1664, 2957, 966, 699, 412, 1574, 1564, 78, 2020, 1036, 1377, 998, 1753, 2774, 971, 24, 1312, 2019, 1134, 1408]. americana [2325, 2731]. americanus [353, 1066, 1724, 272, 900, 160, 2448, 1405, 1668, 398, 185, 1497, 2272, 2021, 128, 664, 105, 2836, 585, 1161, 1531, 1766, 2456, 746, 1686, 2665, 1598, 1116, 1364, 439, 2893, 2044, 2064, 161, 835, 984, 992, 898, 1371, 1500, 1438, 1748, 2740, 1363, 103, 338, 1072, 33, 427, 2020, 268, 1158, 1108, 634, 641, 1490, 998, 1918, 1753, 2774, 524, 1827, 556, 1073, 2019, 3]. Americas [2093]. Amine [734]. Amine-Citrate [734]. Amino [444, 2917, 1855, 1558, 2175, 2247]. Aminolevulinate [948]. Aminonaphthylamidase [1656]. Aminophenol [931]. Ammodytes [353, 1924, 827, 828, 927, 902]. Ammonia [1077, 2667, 2785, 2705, 1780, 2849, 2664, 2759, 1763, 798, 2860]. Ammonium [374, 384]. Amoebae [2354]. Amoebocytes [2823]. Among [2179, 1789, 170, 500, 308, 2214, 706, 833, 1109, 1184, 984]. Amorphous [2252]. Amounts [1027]. Amphibians [1679]. Amphipacific [487]. Amphipod [547, 1527, 379, 2073]. Amphipoda [73, 472, 765, 766, 1546]. Amphipods [1809, 1671, 369]. ampullatus [1627]. Amur [2479]. anadromes [2969]. Anadromous [1315, 943, 2509, 1373, 1261, 291, 2209, 2969]. Anaerobic [2928, 2230]. Analogues [1807, 2046]. Analogy [2425]. analyse [910]. Analyses [2023, 2597, 252]. Analysing [419]. Analysis [1, 1723, 2661, 184, 318, 2511, 2344, 384, 744, 2917, 1053, 2341, 2817, 2037, 2933, 1016, 1160, 389, 2078, 1870, 2009, 2508, 80, 661, 1124, 878, 1415, 1563, 1245, 2291, 2819, 760, 1561, 1534, 516, 2088, 2846, 2315, 2108, 1805, 2535, 1484, 1989, 910]. analytical [1253]. analytique [1253]. Analyzing [2896]. Anarhichas [2903]. Anatomie [2867]. Anchor [640]. Anchovy [1203, 609]. Ancyrocephalinae [1038, 1271]. and/or [483]. Androgens [337]. Anemia [2198]. Anemone [2004]. Anesthesia [363, 1122, 97]. Anesthetic [1990, 2563]. Anesthetization [1259, 362, 767]. Angeles [494]. Anguilla [772, 2212, 928, 652, 1396, 719, 966, 1377, 971, 1134, 1408]. anguillarum [1985, 2196, 1925, 342, 2792, 2802]. angulosus [765]. Aniline [2715]. Animal [1938, 1061, 913]. Animals [2629, 1322, 37, 2701, 1326, 1081, 41, 760, 1019, 1102, 768, 2932]. Anisakis [2178, 1510]. Anisogammarus [1527, 2073]. Ann [1070]. Annelida [492]. Annelids [488, 494]. Announcement [826, 1483, 2058, 2138]. Announcement/Avis [1483, 2058, 2138]. Annual [1274, 1555, 1139, 1641, 2470, 1689, 915, 2443, 1818]. Anomalous [475]. anomalum [1486, 969]. Anonchohaptor [1486]. Anoplopoma [512, 195, 112, 721]. Anotopterus [55]. Anse [649]. Antarctic [1625, 2963]. Anthopleura [2004]. Antibacterial [1851, 783]. Antibiotic [801]. Antibiotics [2711]. Antibodies [2691, 397, 109, 598, 869]. Antibody [164, 2851, 1769, 460, 1115, 2164, 1502, 598, 2635, 1465, 2972]. Antigen [2851, 2635, 1465]. Antigens [1421]. Antillean [1638]. Antilles [1391].

Antimicrobial [1516]. Antimora [1060, 2482]. Antimycin [594, 1606, 1959]. Antioxidants [1525]. Antisera [1998]. Apatite [2489]. Aphanizomenon [1824]. Aphyonidae [1365]. Apotheosis [2259]. Apparatus [1306, 1734, 1735, 899, 2460, 1102, 1106, 1726]. Appareil [2867]. Apparent [1468, 1426, 1725, 2927, 2494]. Apparently [1101]. Appearance [909, 1626]. Appearing [1521]. Appendix [350]. Appetite [322]. Applicability [2556]. Application [2330, 184, 1658, 779, 1087, 2711, 2282, 1972, 2264, 2044, 1941, 1989]. Applications [2467, 1958, 2424]. Applied [953, 1006, 2352, 2004, 1875, 2539]. Appraising [1727]. appreciation [482]. Approach [2517, 1088, 806, 1868, 676, 2315, 1118]. Approaches [247, 1460, 1084]. April [1623, 1894, 557, 2601]. Aquacultural [1326, 1201, 1948, 1359]. Aquaculture [2653, 1931, 2829, 1939, 1958, 1196, 2652, 1932, 1200, 1927, 1933, 2777, 2500, 2777]. aquae [1824]. Aquaria [384, 817]. Aquarium [2328, 1908, 1909, 536, 632, 783]. Aquatic [56, 953, 1169, 2811, 2072, 2463, 2277, 2464, 1733, 1308, 2646, 506, 2897, 1744, 1851, 1735, 2761, 1774, 1866, 1727, 1965, 2796, 2327, 1352, 2467, 2135, 2134, 2945, 1081, 2238, 469, 626, 2808, 1436, 2140, 2215, 2806, 1541, 2589, 2250]. Aqueous [2923, 1780, 2140, 2451, 798, 2986]. Arab [1758]. araneus [2695]. Archipelago [2031]. Archosargus [2583]. Arctic [2031, 2606, 2096, 2159, 605, 1755, 1809, 1040, 1524, 1635, 797, 991, 1594, 2025, 2509, 2736, 1808, 1067, 980, 2112, 2738, 2093, 832, 551, 372, 291, 1044, 1413, 1856, 2616, 2510, 2745, 1340, 2466, 2053, 2674]. Arctica [1476]. arcticus [1594, 372, 2053]. Arctogadus [605]. Area [299, 1231, 307, 297, 208, 296, 509, 27, 305, 2199, 1364, 209, 2628, 1605, 1303, 303, 2885, 254, 2202, 2203, 301, 530, 308, 298, 300, 304, 1093, 306, 302, 1366, 2307, 237, 999, 656, 1907, 1354, 2533]. Areal [2744]. Areas [2610, 488, 55, 814]. arenaria [179, 2148, 2612, 2613]. arenatus [1163]. Arenicola [2607]. argentatus [2723]. Argentine [1268]. Argopecten [2921]. armatus [1398, 577]. Aroclor [2663, 1992, 2673, 662]. Aromatic [2973, 2610, 2609, 2597]. Arrow [471, 2599]. Arsenic [2554, 1999, 1781, 1654, 897]. artedii [987, 2141, 2542, 403, 2730]. Artemia [2759]. Articles [29, 50, 68, 90, 115, 139, 158, 174, 188, 215, 241, 292, 309, 333, 360, 387, 406, 429, 455, 480, 542, 581, 600, 622, 643, 668, 724, 750, 775, 799, 824, 851, 879, 914, 936, 957, 979, 1026, 1110, 1000, 1057, 1083]. Artificial [1258, 2479, 509, 808, 384, 567, 2062, 2314, 1837, 2912, 2020, 2246, 2370]. Artificially [1906, 1920, 1090, 397, 1096]. Aryl [2614, 1561]. Arylamidase [1316]. Arylsulphatase [214]. Ascaridata [2178, 1510]. Ascorbic [2457, 2271, 2577]. Ash [2911]. Ashfalls [603]. ashore [1637]. Asia [1937, 1908, 1933]. Aspartate [1749]. Aspects [1406, 501, 1001, 1342, 467, 1592, 1954, 1570, 1132, 988, 791, 1991, 1461, 2306, 1632, 1630, 766, 2335, 217, 120, 2297, 990, 2847, 1900, 2446, 2696, 145]. aspera [2262]. Assay [593]. Assaying [2666, 377]. Assess [1021]. Assessing [1176]. Assessment [1522, 1203, 1951, 2047, 2282, 2625, 1911, 2864, 2218, 2601, 1720, 1177, 2466, 2304]. Assessments [2278]. Assigned [1125]. Assimilation [129, 385, 736, 2980]. Assiniboine [1555, 2077]. Associated [501, 1719, 227, 2363, 2695, 2233, 2978, 2112, 2889, 816, 2403, 2788, 1909, 2230, 586, 2106, 589]. Association [2478]. Associations [2219, 2525, 521, 2390]. Assumption [2491]. Asterionella [2554]. Asymptomatic [1785]. Athabasca [2264]. Athecate [602]. Atheresthes [1444]. atherinoides [1923]. Atlantic [1636, 1230, 2048, 1307, 214, 635, 217, 189, 56, 1477, 2024, 2075, 777, 2669, 1, 2983, 2984, 1997, 2591, 2452, 1212, 375, 2178, 2713, 132, 2976, 743, 1455, 2511, 1635, 2517, 339, 1495, 436, 1607, 779, 423, 1024, 1333, 1612, 2866, 2515, 599, 759, 933, 1996, 2513, 2712, 1700, 1471, 1208, 11, 1209, 874, 1751, 1961, 1585, 613, 566, 844, 1232, 612, 1507, 1806, 2841, 274, 1581, 1921, 114, 316, 2453, 1749, 891, 186, 332, 1973, 2508, 2530, 1179, 2130, 2239, 1693, 1993, 2339, 2149, 2553, 1775, 2821, 170, 1163, 2878, 2469, 1472, 1577, 2080, 1642, 212, 1034, 415]. Atlantic [1515, 1228, 182, 1661, 2523, 1004, 1325, 591, 1301, 1420, 66, 235, 347, 1427, 1361, 1324, 1031, 1588, 2038, 2486, 2170, 254, 1539, 1276, 2524, 2559, 966, 985, 1558, 151, 2746, 2802, 792, 21, 828, 1574, 2186, 855, 2518, 1721, 1722, 2963, 2961, 279, 765, 766, 998, 218, 349, 1011, 2700, 2843, 965, 2147, 2648, 2531, 2781, 55, 54, 176, 1098, 1146, 1117, 1366, 1907, 814, 408, 2491, 2766, 2340, 2539, 2475, 345, 2913, 2482, 2494, 745, 897, 45, 2041, 2223, 2488, 2487, 2697, 1159, 3, 2184, 829]. Atlántico [1212]. Atlas [1679, 2866, 2688, 2468, 2687]. Atmosphere [1297, 2860]. Atmospheric [2309, 2273, 2307]. Atomic [2921, 84]. ATP [1095, 2787]. ATPase [2952, 2746]. atrarius [563]. atratulus [242, 202, 1249]. Atresia [60]. atromaculatus [2555, 917]. atropurpurea [2492]. Attached [567]. Attachment [709]. Attraction [161, 1448, 1601]. Attributable [234]. Attributed [1420]. Attributes [1775]. Aulorhynchidae [324]. auratus [761, 1849, 1850, 1690, 764, 251, 1943, 568, 981, 2507, 1392, 1842, 125, 2928]. Australia [2331, 2376]. Australian [2876]. australis [1964]. Automated [2761, 716]. Automatic [2097, 2649, 35, 1541]. Autotrophic [2014, 2506]. Autumn [415, 1563, 1627, 347]. Autumn-Spawning [347]. aux [2226]. Availability [1248, 795, 2758, 2309, 2609, 263, 2395]. Available [2897, 2832]. Avalon [213]. Average [2650, 820, 1457]. Avis [1483, 2058, 2138]. Avoidance [2362, 2761, 2459, 440, 2256, 1106, 665]. Avoided [1565, 2201]. Axenic [2154]. Axis [2087]. ayresi [950]. Ayu [1944]. azote [857]. azteca [73, 369]. aztecus [1412]. Azygokeras [2013].

B [890]. B.C. [2598]. B2033 [1261]. Babine [587, 1464, 36, 2364, 1008, 1784]. Back [15, 2186]. Back-Calculated [2186]. Back-Calculation [15]. Backcrosses [2105]. Background [295]. Backgrounds [1151]. Bacteria [1733, 2349, 179, 1351, 178, 311, 2889, 1602, 2152, 2002, 2605, 2928, 1286, 314]. Bacterial [1308, 501, 512, 1101, 2955, 2972, 1893, 2300, 771, 2504, 1650, 1908, 1490, 1488, 2167, 443, 632, 655, 2755, 2913, 2236, 1521, 2696]. Bactericidin [2740]. Bacterins [1985, 2196, 1279]. Bacteriological [222]. Bacterium [342, 2164, 2978, 1502, 2230, 2236]. Baeoctenus [2525, 2481]. Baffin [2159, 467, 1303]. bairdi [631, 2123, 1128]. Bait [138]. Baja [2662]. Balaena [551]. Balaenoptera [790, 1624]. Balance [1495, 2803]. Balanced [859]. Balances [2546]. balanoides [2207]. Balanus [2207, 1503]. Balaton [2398]. baldwini [1271]. Baleen [2372, 2834]. Ball [2856]. Balon [1978]. balthica [1067]. Baltic [2626]. Band [2922]. Banded [240]. Banff [31]. Bangia [2492]. Bank [2693, 850, 330, 1058, 2693, 1658, 2090, 83, 1475, 1664, 418]. Bankia [130]. Banks [464]. Barathronus [1365]. barbue [1730]. Bark [1968]. Barnacle [86]. Barnacles [1503]. Barren [677]. Barriers [192, 125]. Barrow [1470]. Basal [2959]. Base [2748]. Based [1684, 2508, 1641, 1520, 1854, 2345]. Basic [2806, 1875]. Basin [2228, 915, 2116, 412, 1878, 2855, 488, 2790, 1355]. Basins [297, 1888, 1720]. Basis [653, 59, 2359, 1011, 515]. Basket [2073]. Bass [1468, 2914, 639, 1409, 2362, 544, 1302, 618, 2265, 651, 1046, 2393, 2390, 2526, 1729, 2185, 2441, 1759, 2579, 46, 2108, 2766]. Batch [880]. Batchawana [378]. Bathyl [1060]. Bathyl-Demersal [1060]. Bathymetric [323]. Bathymetry [297]. bathypelagica [599]. Bathyraja [1146, 1125]. Bathysauridae [723]. Bathysaurus [723]. Bathythermograph [285]. Bay [1265, 790, 422, 71, 572, 2598, 136, 585, 1161, 949, 2612, 2035, 2752, 2754, 2977, 2835, 253, 232, 2301, 154, 2599, 439, 378, 2678, 2756, 411, 2296, 768, 1377, 2605, 2753, 863, 2623, 1665, 1408, 2855, 2207, 1283, 572, 471, 656, 192, 2415, 401, 2101, 818, 935, 557, 642, 2763, 2417, 1364, 521, 523, 522, 2305, 2130, 1627, 859, 2626, 2609, 2624, 1030, 19, 2597, 2692, 1768, 2230, 2600, 2297, 1073, 2232, 464]. Bayer [2698, 1569]. Bays [625]. be [1789]. Beach [494, 3]. Beam [2908]. Bear [1741, 1740]. Bearing [939]. Bears [2335]. Beats [972]. Beaufort [2102, 2316, 2071, 2288]. Beaver [1314, 1548]. Bed [2789]. Bedford [2790, 915]. Beds [2561, 666, 1948]. Beef [1437]. Before [2371, 1839, 748, 280]. Behavior [1033, 756, 1757, 2636, 467, 1487, 2706, 764, 1858, 2380, 785, 1148, 1447, 2221, 664, 2721, 2767, 1905, 1471, 607, 2904, 1977, 1598, 2542, 417, 981, 440, 1014, 1448, 1112, 2213, 2280, 1151, 1626, 2379, 466, 654, 927, 1036, 1387, 2579, 2891, 965, 2147, 2648, 1587, 2381, 2702, 636, 2882, 231]. Behavioral [2378, 2279, 1495, 1068, 1531, 1249, 1256, 1039, 2531]. Behaviour [207, 227, 320, 76]. Behavioural [507]. Below [789, 223]. Beluga [1126]. Beneath [122]. Beneficial [171]. Benthic [2629, 2606, 2322, 797, 1345, 762, 2610, 1064, 2035, 907, 521, 523, 522, 1741, 763, 2628, 69, 255, 1674, 1470, 1701, 1718, 498, 30, 243, 1967, 2560, 2956, 2805, 1720, 2624, 768, 377, 2117, 2297, 2812, 1337, 2446, 2685, 2686, 1001, 2865]. benthique [1001]. Benthos [2971, 191, 205, 1814, 1334, 394, 317, 275, 747, 2967, 2855]. Benthosema [11]. Benzene [1729, 2639]. Benzocaine [97]. Bergendal [932]. Bering [2123, 1398, 2262, 1154, 577, 1224]. Berkeley [482, 483]. berkeleyi [492, 491, 496, 486]. berkeleyorum [510]. Bertalanffy [2330, 2733]. Best [2890]. Between [2512, 505, 461, 242, 2387, 2734, 490, 1636, 2643, 2976, 185, 2611, 471, 2100, 1060, 2479, 1241, 2235, 2456, 1174, 752, 1304, 1403, 1015, 1857, 2393, 2195, 2505, 2519, 1094, 2962, 2820, 666, 2382, 291, 2890, 947, 2699, 1383, 2152, 1542, 532, 823, 2715, 2480, 837, 65, 536, 462, 2318, 1794, 314, 1297, 77, 380, 283, 161, 347, 243, 998, 2491, 2297]. Beverton [1114]. Bias [2925, 2926, 1797, 1805, 2532]. Biases [2783]. bibliography [2329]. Bicarbonate [1153]. Bicinctus [2177, 2202, 2203]. Bicinctus-Group [2177]. bicolor [2525, 2481]. bicuspidatus [1793]. Bideford [747]. Bielersee [682]. Big [282, 1921, 144]. Bight [2866, 44]. Bilateral [2064, 1432]. Bile [2129]. Biliary [1062]. bilinearis [2119, 2483, 1673]. Bill [1080]. Billfishes [1237]. Binary [2785]. Binder [1326]. Binding [2734, 62, 1035, 1715]. Binnengewässer [2922]. binoculata [282]. Bioaccumulation [2176]. Bioassay [761, 1099, 647, 2244]. Bioassays [1850, 1556, 1573, 2365, 1541, 913]. Bioavailability [2617]. Biochemical [189, 527, 1951, 2517, 2263, 2550, 1958, 802, 224, 2915, 2790, 1362, 1397, 1491, 2443, 1753, 904, 2847]. Biochemistry [1681, 2906]. Bioconcentration [2905]. Biodegradation [1547, 738]. Biodeposition [1274]. Bioeconomic [2947]. Bioeconomics [2863]. Bioenergetics [592, 2424, 1991, 1854, 2538]. Bioenergetics-Based [1854]. Biogeochemistry [2546]. Biological [1342, 797, 2567, 2251, 2898, 47, 714, 1961, 368, 2656, 2864, 2029, 1661, 1625, 2758, 1771, 1664, 1720, 1441, 250]. Biologically [1723, 1094]. Biology [1154, 1620, 1464, 117, 646, 1633, 1269, 506, 1635, 2421, 2880, 1811, 2550, 2590, 2036, 1928, 13, 626, 280, 778, 588, 1030, 2865, 765, 766, 1146, 1288, 1900, 2482, 511, 2865]. Bioluminescence [88, 497]. Biomass [505, 2677, 1272, 2391, 1296, 2434, 2413, 820, 1457, 712, 1693, 1894, 1895, 1718, 243, 2246, 1389, 1294]. Biophysical [2550]. Biostatistical [516]. Biosynthesis [1695, 1781]. Biota [1654, 1660, 863]. Biotelemetry [2355]. Biotin [2049]. Biotransformation [1901]. Biphenyl [1710, 1434, 2673, 662]. Biphenyls [624, 1755]. Birge [2243]. bis [247]. Biscayne [2230]. Biting [609]. Bivalve [1817, 1272, 2921, 2525, 660]. Bivalves [2614]. Bivalvia [1731, 2031, 2029]. Black [1538, 66, 1913, 1151, 1129]. Blackbass [1815]. Blackfly [287, 2264]. Blackgill [2662]. Blacknose [202, 1249, 752]. Blackstripe [969]. blainvillei [1633]. Blasting [2692]. Blazka [155]. Bleach [1651]. Bleached [1799, 1099, 919, 1609, 2256, 1362, 1604, 2906, 76, 898, 805, 2304]. Blebs [1473]. Blennioid [2048]. Bligh [848]. Blinded [125]. Bloater [228]. Blood [1406, 1997, 756, 1914, 123, 452, 1839, 2645, 1686, 1855, 2578, 356, 400, 363, 767, 2453, 1595, 210, 922, 1777, 2244, 409, 1501, 2046, 2209, 2443, 135, 354, 1575]. Blood-Pressure [409]. Bloodworm [882]. Bloom [95]. Blooms [1824]. Blowout [2601]. Blubber [1298, 2312]. Blue [2268, 1603, 2937, 712, 2074, 2396, 2447, 2715, 770, 1823, 2892, 1356, 839, 277]. Blue-Green [1823, 2268]. Bluefin [423, 1024]. Bluefish [1060, 617]. Bluegill [1757, 1862, 657, 2023, 2505, 1254, 911, 1395, 888, 328, 2110, 1728, 2161].

Bluegills [1143, 1385]. Bluespotted [476]. Blumer [2595, 2979]. Board [173, 454, 749, 1056, 1676]. Boards [27]. Boat [1801]. BOD [2705]. Bodensee [689]. Bodied [2004]. Bodies [1088, 1171]. Bodonidae [758]. Body [559, 1071, 2198, 40, 1060, 2830, 2065, 1965, 100, 2199, 114, 1789, 2769, 2953, 595, 2538]. Boeckosimus [1809]. Bog [1394]. Bomb [348]. Bomolochidae [504]. Bonavista [585, 1161]. Bone [61, 2673, 1582, 2451, 1750]. Bones [324, 2186]. Book [1440, 2030, 2056, 2082, 2136, 2568, 2546, 1816, 2795, 2794, 2867, 2922, 2216, 2495, 2589, 2468, 2777, 2986, 2687, 2688]. boops [1163]. Boothbay [208]. Borate [1792]. Bordering [1209]. Boreal [2392]. borealis [71, 447, 667]. Boreogadus [2159]. Boreohydra [1455]. Borer [27, 130]. Borers [132]. Boron [1614]. Bothrimonus [1269, 778]. Bothwell [1265]. Bottlenose [1627, 1630]. Bottlenosed [1017]. Bottom [1147, 2087, 1364, 809, 1386, 1605, 2684, 747, 2200]. Bottomfish [1224]. botulinum [431, 1136, 635, 963, 541]. Boulder [797]. Bound [2607]. Bourne [2625]. bouvieri [2850]. Bovine [2248, 2641, 2018]. Bow [870]. Bow-Mount [870]. Bowhead [551]. Boxes [1286]. Brachiella [245]. Brachiopod [518]. Brachydanio [1611]. Brachyura [1162]. Brachyuran [1734, 2123]. Brackish [1219]. Brackish-Water [1219]. Brain [356]. brama [2885]. Branchial [767, 2585]. branchialis [1907]. branchiaux [2585]. Branchiitis [2053]. Break [2936]. Breakdown [1383]. Bream [2885]. Breeding [945, 117, 1867]. Breton [871]. brevirostrum [949]. Brevoortia [1581, 332, 1324]. Briarosaccus [86]. Brief [1012, 1731]. Brienzersee [682]. Brine [1541]. British [520, 1005, 1464, 890, 230, 8, 808, 793, 2341, 2356, 1752, 1442, 2981, 2935, 2557, 144, 369, 133, 2, 36, 278, 850, 147, 942, 1689, 1008, 1784, 2014, 2308, 2854, 2091, 2274, 2015, 461, 519, 561, 2261, 1646, 610, 510, 2342, 1101, 618, 740, 2872, 126, 323, 502, 2817, 2544, 1261, 560, 2440, 94, 504, 982, 2115, 834, 1512, 2013, 1776, 2073, 710, 1697, 500, 404, 476, 1967, 2109, 576, 392, 2779, 2853, 146]. Brood [1280]. Brook [2180, 38, 1795, 1803, 1971, 640, 974, 2131, 821, 1035, 162, 1662, 2011, 428, 584, 2047, 2571, 1700, 2931, 607, 1751, 566, 1143, 1010, 2039, 362, 363, 767, 1122, 331, 2459, 186, 739, 329, 402, 358, 124, 2673, 727, 210, 366, 1317, 2143, 1584, 1833, 2955, 854, 2359, 1018, 1821, 1031, 1467, 434, 340, 51, 263, 1787, 2258, 1579, 2696]. Brown [955, 722, 1115, 2229, 1425, 85, 124, 2955, 1781, 1480, 1412, 2818, 1151, 1959, 2331]. Brown-Water [722]. Bruch [1446]. Brunswick [1751, 1476, 1590, 2042, 1961, 1921, 316, 985, 464]. Bryozoan [234, 201]. Bryozoichthys [265]. Bubble [660, 2216]. Bubbling [812]. Budget [1274, 1892, 1040, 1469, 253, 2273, 705, 2556]. Budgets [1889, 2347, 2120, 2116]. Buffers [734]. Building [490, 263]. Buildup [466]. Bulb [769]. Bulbar [2930]. Bullhead [955, 1115, 2229, 1425, 1538, 1913]. Bullheads [1959]. Bunker [471, 2612, 557, 2599, 768, 863]. Buoyancy [2555, 1996, 202, 1374, 1249, 1538]. Burden [2065, 1965]. Burhot [196]. Burial [1696]. Burlington [2939]. Burrow [427]. Burrowing [25]. Bush [484]. Bute [1752, 1442, 2981]. Butter [532, 823, 822]. Butyl [1547].

Buzzards [2597, 2624]. By-Catches [1647]. By-Pass [1278].

C [2180, 2627, 2193, 2202, 2203, 2850, 2075, 1733, 2580, 471, 1607, 2639, 2612, 557, 1384, 2037, 1806, 907, 2599, 1673, 224, 1054, 1438, 1383, 1691, 768, 300, 812, 2197, 377, 863]. C-Aldrin [2075]. C-benzene [2639]. C-Labeled [1733]. C-Labelled [377]. C-p [2180, 2193]. C. [273, 2123, 2542, 947, 2202, 2203]. Cabot [2249]. Cache [1413]. Caddisfly [2551]. Cadmium [1097, 2574, 2554, 2842, 457, 774, 248, 757, 2582, 2507, 995, 726, 2946, 2679, 2716, 2066, 2579, 897]. Cadmium-Induced [2946]. Caecum [130]. caenaematicus [252, 667]. Cage [2526]. Caged [2264, 964]. Calanoid [257]. Calanoida [1926, 2231, 2013, 1107]. Calanus [1725, 273, 988, 2193]. Calculated [2186]. Calculating [2534]. Calculation [2650, 15]. Calculations [1780]. Calibration [1919, 1605]. California [1816, 1703, 376, 422, 2636, 1782, 154, 2662, 2501, 412, 889, 1645]. californianus [1782, 1782]. californiensis [1002, 2195]. Caligid [313]. Caligidae [313, 438, 807, 1141, 982, 1157, 1250, 1443]. caligorum [1141]. Caligus [807, 1141]. Callorhinus [1514, 1788]. callosus [86]. Caloric [379, 2292]. Calorie [434]. Calorie-to-Protein [434]. Calorific [348]. Calorimeter [348]. Cambios [1212]. Campbell [1107]. Campeloma [149]. Campostoma [969]. camtschatica [2317]. Can [1724]. Canada [1265, 312, 173, 454, 749, 1056, 1522, 2277, 2964, 2965, 1899, 1983, 1331, 175, 602, 1185, 2085, 2879, 1299, 346, 342, 599, 2327, 2263, 1352, 238, 969, 1074, 1751, 1585, 1330, 2656, 2054, 883, 2007, 2005, 621, 735, 2469, 2080, 525, 1856, 1420, 1771, 2873, 1736, 1676, 1417, 2116, 21, 216, 994, 2668, 2564]. canadense [2377, 1006, 1960, 2406, 2067, 2395]. canadensis [514]. Canadian [1029, 1628, 624, 1349, 132, 1272, 1529, 1774, 1568, 2048, 1307, 1808, 806, 1360, 214, 2112, 245, 1444, 2222, 2093, 635, 2031, 1163, 551, 2888, 2869, 1947, 1886, 1927, 1276, 1976, 2518, 2416, 279, 625, 814, 383, 1572]. Canal [2939, 2380, 425]. Canary [730]. Cancer [444, 2636, 10, 2721, 2836, 2162, 2237, 2711, 791, 588, 1441, 2949]. Canned [1525, 247]. Cannibalism [2067]. Canning [1048, 842, 2509, 1044]. Can’t [2775, 2464]. canthaxanthindiylidene [247]. Canyon [1285, 2622]. Capacity [123, 1426, 1375, 1684, 2040]. Cape [871, 1070]. Capelin [2435, 2437, 2661, 2580, 1568, 2436, 1268, 2518, 1747, 1746, 2197, 250, 45, 418]. Caperea [1626]. Captan [1143]. Captive [1905, 112, 721, 2053]. Captivity [835, 873, 1055]. Capture [1644, 956, 1646, 2735, 1499, 2717, 2241, 1264, 1050, 2108, 1645, 325]. Captured [1285]. Capturing [195]. Carapace [642, 2123, 134]. Carassius [761, 1849, 1850, 1690, 764, 251, 1943, 568, 981, 2507, 1392, 1842, 125, 2928]. Carbaryl [1322, 659]. Carbohydrate [1604, 1145]. Carbon [788, 2546, 1292, 2347, 2842, 401, 1064, 1493, 474, 1319, 2833, 466, 1092, 2514, 2787, 1086, 2858, 1297, 2576, 2367, 2674]. Carbonate [300]. Carbonates [1086]. carcharias [741]. Carcharodon [741]. Carcinogens [2973]. Carcinonemertes [2949]. Carcinus [857, 2156]. Cardiac [1396, 338].

Cardiovascular [1122]. Caribbean [1638, 2682]. Caridean [510]. Caristiidae [21]. Caristius [21]. Carnation [2014]. Carnitine [93]. Carnivores [1610]. Carolina [6]. Carotenoid [2268]. Carotenoids [110]. Carp [2248, 475, 1770, 1122, 1003, 2002, 1994, 2928]. Carpentaria [2947]. carpio [2248, 475, 1770, 1122, 2002]. Carps [1940]. Carrier [1785, 1669, 2729, 1579]. Carriers [2839]. Carrying [2907]. Cascade [2762]. Casco [253, 232]. Case [1204, 236, 1223, 1762, 453]. Casein [1020, 1558]. castaneus [207]. Catalogue [2495]. Catalyzed [1749]. cataractae [242, 616, 1249]. Catastrophe [2157]. Catastrophes [2855]. Catch [350, 1638, 2974, 1648, 625, 2169, 87]. Catches [1844, 1922, 1647, 1222, 1420, 1206, 1970, 2461, 2462]. Catching [1796]. Catfish [1130, 1764, 1302, 2706, 1910, 1115, 1992, 2978, 357, 1300, 1675, 1578, 2581, 951, 869, 633, 1942, 2919, 1050, 586, 580, 1109, 1730]. Catheptic [573]. Catheterization [183, 381]. Cation [374]. Cations [823]. Catostomidae [2234]. Catostomus [527, 960, 1251, 2225, 722, 2543, 1710, 2394, 1825, 1835, 224, 470, 133, 2320, 2659, 947, 2002, 2204, 968, 1253]. Caught [881, 1029, 1973, 1698, 1072]. caurinus [473, 239, 355]. Causative [2696]. Cause [1724, 763, 2494]. Caused [2050, 2010, 1903, 595, 2579, 2125, 2232]. Causes [1070, 2336, 2420]. Causing [150, 1769]. Caustic [1651]. Cautionary [1411]. Cayuga [379, 1287, 685]. Cd [2159, 2801]. CDC [15]. CDP [1695]. CDP-Choline [1695]. Cedar [2140]. Cell [1914, 612, 2666, 2583, 1835, 2183, 1777, 2516, 1004, 2046, 880]. Cell-Size [880]. Cells [2851, 16, 339, 1716, 1115, 1466, 541]. Census [1666]. Center [1703]. Centimetres [2684]. Central [1816, 2750, 1899, 1212, 1888, 1858, 2632, 2389, 806, 1208, 1209, 2852, 109, 514, 2116]. Central-Eastern [1212]. Centrarchidae [1160, 2519, 1520]. Centrifugation [1247]. Centro [1212]. Centro-Oriental [1212]. cepedianum [2206]. Cephalopoda [534]. cephalus [1448, 2476, 2230]. Ceratomyxa [1697, 2311]. Ceratomyxosis [1687]. cerebralis [2691, 1123, 1445, 1247, 2635, 1421, 2772]. Ceriantharia [519]. cernua [2399]. Certain [150, 163, 594, 1843, 1664, 2837, 650, 217]. Cestoda [1464, 2276, 1269, 274, 778]. Cestode [1008]. Cetacea [1644, 1636, 1634, 1034, 216, 1391]. Cetacean [1640]. Cetaceans [1623, 1649, 2784, 1632, 1648, 1765, 1645, 1620]. CH [2570, 2570]. Chaetognath [915]. Chaetognatha [599]. Chain [1298, 868, 862, 860, 172, 2449, 2480, 903]. Chains [1755, 1866]. Chairman [1243]. challengeri [477]. chalumnae [2867]. Chamber [155]. Change [2914, 2279, 763, 1538, 1463, 595, 2538]. Changes [1406, 756, 2555, 271, 2163, 1914, 772, 2608, 1051, 2400, 1428, 1348, 1119, 1153, 452, 2342, 1068, 1531, 2415, 933, 1826, 100, 2754, 1585, 695, 1824, 2657, 612, 858, 2119, 1507, 2578, 2405, 1860, 2292, 1714, 2453, 1020, 2719, 2675, 1319, 1578, 2927, 224, 470, 210, 1362, 2418, 1501, 1767, 1691, 2618, 2486, 1664, 985, 2403, 2746, 2197, 1491, 2111, 2579, 1118, 1246, 2200, 2414, 250, 586, 589, 146, 1212]. Channel [1130, 1764, 1302, 2706, 1910, 1115, 1992, 2978, 357, 1300, 1675, 2557, 1578, 2581, 951, 869, 2919, 42, 2446, 586, 580, 1109, 853]. Channels [1387]. Chanos [1938]. Chaoborus [2179, 2434]. Char [1499, 1594, 2509, 2736, 980, 2112, 17, 335, 291, 1044, 1413, 1856, 2745, 1338, 1339, 1336, 1337]. Characteristic [283]. Characteristics [733, 603, 2042, 714, 1248, 1686, 1860, 680, 1740, 2390, 2893, 2732, 802, 1709, 2602, 2790, 108, 1661, 1664, 2215, 569, 1276, 1984, 176, 2747, 345, 804]. Characterization [299, 501, 1656, 2293, 2762, 306, 2230]. Characterized [67]. Characters [610, 2508, 380, 2518]. Charlotte [714]. Check [2435]. Chedabucto [471, 2612, 557, 2599, 768, 2605, 863, 2623]. Chef [230]. Chela [631]. Chemical [31, 1401, 299, 603, 271, 2661, 1342, 37, 2923, 2941, 1585, 553, 2234, 1740, 2007, 984, 1014, 2295, 2215, 569, 301, 1336, 2116, 1984, 2197, 1692, 166, 1011, 1601, 42, 2986]. Chemicals [1726, 2905]. Chemistry [307, 1914, 1369, 363, 1501, 2443, 2497, 135, 354, 1575]. Chemoreception [2665, 1706]. Chemosensory [1748]. Chesapeake [401, 1408]. Chestnut [207]. chez [1079, 2585]. Chicago [2927]. Chick [157, 63]. Chignik [577]. Chile [1640, 432]. Chilko [2015]. Chill [2584]. Chilled [1286]. Chinese [1162]. Chinook [2512, 2940, 862, 861, 860, 1156, 801, 574, 2952, 526, 2907, 2503, 1688, 560, 144, 371, 975, 837, 2563, 1273, 589, 2311]. Chionoecetes [629, 2681, 631, 1596, 2695, 2123, 1605, 2027, 2892, 1128, 181, 636]. Chiridius [1402]. Chirolophinae [265]. Chironomid [2525]. Chironomidae [2525, 2360, 2177, 30, 2068, 2202, 2203, 967, 1530, 1550, 2481, 1853]. Chironomus [2434]. Chloramine [547]. Chlordane [2176]. Chloride [1849, 431, 1890, 647, 2143, 1061, 2245, 1451, 1721, 25, 2919, 2535, 1742, 1743]. Chlorinated [2198, 1782, 1112, 1031, 268, 1765, 2236, 2642]. Chlorination [2923]. Chlorine [2163, 1914, 2914, 2362, 1793, 1919, 1798, 2324]. Chlorophyll [2472, 916, 2840, 753, 1290, 1493, 2557, 1895, 1903, 754, 1399, 2783, 2145, 2779]. Chlorophyll-Light [754]. Chlorouracil [1770]. Choice [2456, 892, 1267]. choix [1267]. Cholesterol [1668, 157, 63]. Choline [1695]. Chondracanthus [219]. Chondrococcus [460, 344, 526, 590, 869]. Chondrophora [507]. Chondrus [2561]. Choriocapillaris [1446]. Chromatograms [930]. Chromatographic [549, 1086]. Chromatography [2222, 420, 1561, 1534]. Chromium [1849, 2785, 2168]. Chromosome [16, 2205]. Chromosomes [2822]. Chronic [1328, 2677, 435, 40, 251, 2861, 1990, 2628, 2820, 2582, 2906, 1833, 2616, 726, 913, 2600, 2034]. Chronically [2621, 2612]. CHSE [2666]. CHSE-214 [2666]. Chub [2555, 117, 917]. Chukchi [1470]. Chum [167, 2643, 26, 1667, 2831, 1016, 1697, 1256, 2832, 1587, 42]. Churchill [1369]. Chydoridae [2243]. Cichlid [2200]. Ciliate [321, 2232]. ciliatus [1047]. cincinnatus [1444]. cinerea [2611]. Circulating [1514, 2784, 2691]. Circulation [2092, 2087, 203, 919, 663, 1834, 1473, 2103, 253, 2678, 2756, 2672, 2332, 1831, 2091]. Circulations [2006]. circulatoire [2867]. Circulatory [509]. Circulus [395, 528]. Cirrutulidae [491]. Cisco [1154, 2141]. Ciscoes [2542]. Ciscos [403]. Citrate [734]. Cl [2585]. Cladocera [2243, 1548]. Cladophora [1829, 2640]. Clair [437, 2419]. Clam [179, 2148, 319, 532, 823, 822, 836, 2613, 1458]. Clams [1322, 1716, 2612, 1657, 1145]. clarki [2850, 654, 1394, 461, 2588, 1593, 230, 607, 75, 2346, 654, 1394, 289, 2032]. Class [2411, 1915, 60, 2422]. Classes [1713, 2820, 169]. Classical [1396]. Claw [1072]. Clay [2602, 1462]. Clean [2622]. Clear [2411, 226]. Clearance [2337, 2193, 2799]. Cleidodiscus [1271]. Cleithrum [2186]. clemensi [807]. Climate [1023, 2564]. Clinal [2508]. Clinocottus [875]. Clipping [2132]. Cloche [731]. Closely [2921, 930]. Closer [2684]. Closing [1242, 2345]. Clostridium [431, 1136, 635, 963, 541]. Clupea [2024, 503, 2584, 2791, 656, 2639, 1612, 2045, 560, 1993, 415, 1515, 347, 2524, 1924, 950, 648, 515, 2494, 2041, 2223, 2488, 2487, 3]. clupeaformis [2313, 1120, 1449, 1525, 2934, 1532, 127, 886, 2929, 1119, 1840, 2263, 1554, 2657, 1284, 1911, 2112, 1252, 2969, 364, 1498, 1750, 1424]. CO [2956, 1092]. CoA [2155]. Coagulation [2046]. Coal [2911, 234]. Coast [1024, 2288, 14, 1585, 400, 1250, 69, 255, 199, 2768, 648, 840, 189, 2669, 208, 2100, 1307, 1248, 1378, 473, 214, 545, 1671, 2802, 477, 2573, 576, 2766, 637, 1816]. Coastal [461, 561, 2826, 2598, 753, 2610, 2877, 809, 710, 2469, 1222, 235, 1689, 654, 1394, 2209, 2846, 1063, 2014, 2109, 2308, 2506, 1831, 2779, 250]. Coastline [2742]. Coccidia [1589]. Cod [399, 1997, 431, 2976, 881, 2517, 129, 713, 759, 1536, 98, 2453, 1749, 1693, 1775, 1325, 1383, 1691, 2186, 1366, 1694, 140, 2491, 1286, 2340, 260, 2159, 1907]. Coded [2455, 2132]. Codfishes [605]. Codworm [2306]. Codworms [2154]. Coefficient [2733]. Coefficients [565]. Coelenterate [2702]. Coelenterates [1033]. Coenzyme [93]. coeruleoalba [1632]. Coexisting [1999]. cognatus [2025, 1841]. Cohabiting [144, 405]. Coherence [2098]. Coho [2723, 596, 1799, 2163, 1914, 2198, 40, 1487, 1592, 1593, 1906, 1904, 141, 2141, 2645, 460, 1826, 1962, 2187, 606, 2503, 1434, 2018, 1688, 1609, 1732, 1860, 2726, 1049, 937, 2704, 144, 549, 1804, 1380, 2033, 922, 1362, 1604, 1777, 2906, 371, 81, 315, 1465, 1466, 479, 2245, 1275, 2504, 2802, 2321, 2679, 975, 2043, 1963, 837, 2167, 2648, 1118, 1687, 267, 354, 554, 846, 904, 990, 2139, 589, 1376]. Cohorts [2310]. Cold [1479, 2627, 1302, 963]. Cold-Smoked [963]. Cold-Water [2627]. Coliform [1439]. Collapsing [2336]. Collected [1262, 477]. Collecting [874, 2770]. Collection [2323]. Collector [930, 620]. Collectors [449]. Colloidal [2888]. Colloids [2889]. Colobomatus [219]. Colonization [2059, 30, 1104, 2451]. Colony [373]. Color [1679, 861, 2724, 1319]. Colorimetric [2254]. Colors [892]. Columbia [576, 461, 519, 520, 561, 152, 2261, 1646, 610, 1005, 1464, 510, 890, 230, 8, 808, 2342, 1101, 618, 793, 740, 2872, 344, 2341, 126, 323, 502, 2817, 2544, 1261, 560, 2440, 2356, 94, 504, 982, 2115, 834, 1512, 2013, 1752, 1776, 1442, 2981, 2935, 2073, 2557, 144, 710, 1697, 369, 133, 2, 36, 278, 850, 500, 147, 942, 404, 476, 1689, 1967, 2273, 1008, 1784, 2014, 2109, 2308, 2854, 2091, 2274, 392, 2779, 2015, 2853, 146]. columbiae [2013]. Columnaris [968, 460, 344, 526, 590, 1688, 1998, 869]. Columns [2474]. Colville [946].

Combination [1434]. Combinations [1552]. Combined [2075]. Comment [2687, 2587, 2685, 2686, 2925, 2964, 2983, 453, 1416, 2926, 2965, 2984, 2688]. Comments [2463, 2461, 2900, 2901, 2902, 1979, 1677, 1978, 1678, 1626, 813, 2462, 2464]. Commercial [1723, 1489, 1114, 1607, 2973, 2408, 316, 1957, 1605, 2863, 2883, 836, 1144, 443, 2766, 662]. Commercially [189, 624, 239, 1072]. commersoni [2225, 527, 960, 1251, 722, 1710, 2394, 1825, 1835, 224, 470, 2320, 2659, 947, 2002, 2204, 968, 1253]. Commission [1230]. Common [2221, 1820, 1312]. Communication [984]. Communications [1875]. Communities [2392, 159, 1349, 2606, 702, 703, 385, 698, 697, 1064, 52, 1485, 69, 255, 671, 2426, 2144, 2208, 2630, 303, 1800, 701, 682, 699, 679, 700, 2489, 1551, 2315, 281, 2623, 408]. Community [693, 2387, 696, 2627, 2748, 2174, 687, 684, 2405, 1773, 2417, 523, 680, 681, 1666, 2427, 692, 683, 688, 1897, 690, 686, 525, 691, 689, 2637, 2956, 2068, 2953, 694, 2224, 685]. Compact [715]. Comparaison [2585, 1267]. Comparative [894, 428, 2571, 1710, 1289, 607, 1674, 1445, 1974, 2802, 1344, 573]. Comparison [1476, 1279, 2196, 204, 2921, 1051, 2611, 1529, 1343, 77, 2974, 907, 2770, 735, 459, 2751, 278, 324, 1698, 317, 2560, 1998, 2210, 817, 2514, 2716, 1963, 1490, 55, 2623, 1286, 977, 1267, 2585]. Comparisons [1005, 339, 983, 1554, 2505, 2666, 2093]. Compartments [2675]. Compendium [488]. Compensation [1996]. Competition [2179, 2413, 1457, 1517, 2067]. Competitive [62]. Competitors [2147]. Compilation [2281]. Complete [1466]. Complex [2587, 2934, 2150, 1756, 1404, 252]. Complexation [1321]. Complexes [1328]. Complexing [1426, 2887]. Complexities [843]. Component [720, 1729, 2315]. Components [1992, 1715, 2321, 408, 48, 662, 1109]. Composition [2743, 32, 353, 629, 444, 1964, 271, 1348, 2554, 1568, 1612, 885, 2859, 100, 1248, 239, 2018, 1268, 1581, 2981, 255, 1319, 1395, 1582, 1894, 2769, 1475, 1276, 985, 1474, 1692, 166, 1359, 2111, 2443, 42, 1346, 2576, 1253, 2969]. Compositional [2608]. Compounds [2180, 2298, 2923, 2294, 730, 1383, 1781, 1756, 569, 975]. Comprehensive [1355]. Comptes [1440, 2030, 2056, 2082, 2136]. Computations [1891, 1890, 1879]. Computer [350, 1494, 15, 1075, 433, 325]. Concentrate [2924, 845, 1082]. Concentrates [61, 1136]. Concentration [862, 861, 743, 1469, 1429, 2065, 753, 557, 2960, 1806, 568, 891, 1123, 1414, 1247, 2295, 1382, 445, 830, 1462, 770, 1389, 2563, 65, 903]. Concentrations [2629, 435, 1060, 1514, 2698, 452, 1571, 1378, 356, 767, 1999, 2557, 72, 611, 1317, 1777, 2450, 2558, 2894, 1917, 650, 1903, 996, 975, 1158, 1392, 1842, 2146, 1763, 993, 805, 940, 2801, 89, 1829, 2324]. Concept [2793, 1818]. Conception [439]. Concepts [1864, 2484]. Conceptual [1461]. Concerning [1979, 41, 1678]. conclusions [2867]. concolor [373]. Concurrently [1521]. Condition [208, 26, 1161, 2941, 911, 15, 2714, 648, 1462, 1827, 250]. Conditioned [613, 1591]. Conditioning [1396]. Conditions [733, 1791, 2201, 1497, 2389, 537, 1598, 1741, 1335, 2270, 2675, 2759, 2361, 2315, 1389, 2395]. Conducted [626, 2699, 1703]. Conductivity [841]. Conference [1244, 1243, 2500]. confervicolus [2073]. Confinement [2573]. Confinement-Reared [2573]. Confirmation [787, 1906]. Conflicts [1184]. Confronting [1239]. Congress [2072]. Connecticut [1148, 1447, 795, 1042, 2273]. Connecting [437]. Consequences [2113, 2617, 2533, 846]. Conservation [2620]. Consideration [492, 1201, 1041]. Considerations [2806, 2260]. Considered [2314]. Constance [2397]. Constant [2266, 1459]. Constituents [2941, 937, 1651, 2295, 441, 1246]. Constraints [1504]. Construction [2692]. Consumers [1965]. Consumption [1071, 1791, 962, 1603, 2069, 926, 795, 571, 1206, 328, 1065, 2395, 804, 1159, 2900, 2901, 2902]. Contained [1019, 23]. Containers [1286]. Containing [2043]. Contaminant [2301, 2238]. Contaminants [1727, 1965, 1049]. Contaminated [2598, 802, 2609]. Contamination [1378, 1714, 2723, 1542, 2068, 2203, 2619, 383, 2307]. Contemporary [1170]. Content [2024, 2856, 2458, 793, 928, 2933, 1152, 757, 15, 2537, 1767, 2043, 1118, 993]. Contents [1620, 2562, 2065, 379, 1581, 1145]. Continental [1973, 877, 1134]. Contingency [661]. Continuous [614, 742, 1654, 1879]. Continuous-Flow [742]. Continuously [1726]. contrôle [2950]. Contribution [2065, 2808, 2273, 1355, 2600]. Contributions [1154, 2401, 169]. Control [953, 34, 1953, 2791, 1255, 1524, 734, 1939, 2974, 1944, 2881, 2873, 301, 338, 2951, 718, 963, 1869, 2777, 1942, 1836, 1523]. Controlled [1791, 2643, 780, 142]. Controlling [2871, 79, 18, 2950]. Controls [2856, 1916, 2877, 458]. Controversy [2804]. Convection [2708]. Conventional [2770]. Conversion [2180, 736, 1065, 805, 940]. Conwy [1817]. Cooked [1542]. Cookeolus [1163]. Cooking [713, 650, 975]. Cooling [2770]. Copepod [2939, 2275, 2193, 165, 1323, 2013, 2485, 1117, 1907]. Copepoda [273, 1926, 2231, 313, 245, 94, 504, 438, 807, 709, 1141, 1002, 1157, 1444, 1250, 1443, 2195, 1402, 2, 1107]. Copepodar [982]. Copepodite [1402]. Copepods [1477, 2322, 1314, 2709, 1725, 2115, 1793, 2710, 257, 219, 2685, 2686]. Copper [149, 1097, 955, 1060, 1571, 974, 1518, 918, 2016, 568, 981, 1991, 210, 366, 1317, 1500, 1748, 1321, 2679, 2716, 2887, 796, 637, 897, 2034, 2824]. Copulatory [446]. Cordgrass [1519]. Corégones [2969]. Coregonids [213, 2969]. coregonoides [447]. Coregonus [1154, 2313, 1120, 1449, 1525, 2934, 1532, 127, 886, 987, 2929, 1119, 228, 1840, 2141, 2263, 1554, 2657, 1284, 2542, 1911, 2112, 1252, 403, 2969, 364, 1498, 2730, 1750, 1424]. Corer [205, 317, 2798]. Corers [1393, 191]. cornutus [2221]. Cornwall [2622]. Cornwallis [1340, 1336, 1338]. Corpuscles [772]. Correction [742]. Corrections [173, 454, 749, 1056, 1419, 1506, 2057]. Correlated [1789]. Correlations [2734, 2820, 998, 2169]. Cortical [339, 612]. Corticosteroid [1571]. Cortisol [256, 62, 337, 2679, 1392, 2563, 336]. Cortisone [62, 258]. Corynebacterial [1496, 904]. Corynebacterium [1769]. Cottidae [405].

Cottus [2025, 1509, 479, 1841]. Couesius [117]. coulteri [2824, 843]. Council [1179]. Count [458]. Counter [716]. Countercurrent [2086]. Counters [419]. Counting [1368, 2124]. Countries [1209, 1216]. Country [1204]. Counts [419, 372, 1777, 443, 1286]. County [1080]. Coupling [1738, 1064]. Course [1824]. court [2950]. Covariability [2840]. Cover [2895, 2449]. Crab [629, 444, 2681, 2636, 631, 10, 1734, 1596, 2721, 2836, 2162, 2237, 2711, 195, 2615, 791, 2317, 86, 1605, 2027, 1162, 588, 2156, 1441, 1359, 2892, 1128, 794, 181, 636, 2949]. Crabs [2456, 2695, 2123, 995, 1796, 134]. Crangonyx [369]. Crassostrea [1274, 2823, 1790, 2948, 1258, 1012, 154, 1602, 2227, 78, 1312, 6, 7]. crassus [35]. Crawling [1140]. Crayfish [136, 1786, 2188, 2931, 2438, 2343, 2439, 1975, 2111]. Creating [1948]. Creek [1265, 2555, 230, 1413, 917, 2014]. Cricotopus [2202, 2203, 2177, 2202, 2203]. crispus [2561]. cristivomeri [2736]. Criteria [1951, 1215]. Critical [920, 2187, 1829, 2950]. Critique [458, 2950]. Croissance [649, 1730]. Crop [2470, 2471, 2533]. Crops [22]. Cross [2575, 2557, 1805]. Cross-Channel [2557]. Cross-Section [1805]. Cross-Shelf [2575]. Crowding [2139]. Crude [2606, 2148, 1111, 2610, 2617, 1729, 2609, 2068, 2203, 1768, 2019]. cruentifer [1900]. Crustacea [572, 2799, 502, 1132, 438, 709, 2115, 1402, 1162, 472, 765, 766]. Crustacean [312, 1349, 2240, 303, 786, 942, 2616, 1800, 340, 1562, 1551, 2315, 1294, 1896]. Crustaceans [121, 2493, 131, 2117]. Cryo [371, 558]. Cryo-Preserved [371, 558]. Cryopreservation [1945]. Cryptobia [758]. crysoleucas [2904, 2366]. Ctenopharyngodon [1994, 2928]. Cu [2159, 2801]. cubensis [2029]. Cucullanellus [1277]. Cucullanidae [1277]. Culaea [821, 402, 2359, 263]. Cultivation [2711, 1946, 2795]. Cultural [1884]. Culturally [680]. Culture [2154, 1401, 1792, 1308, 1817, 1734, 1938, 1725, 2241, 2954, 1219, 597, 2526, 1937, 2583, 647, 2759, 1602, 2739, 2822, 1955, 2514, 2139, 1941]. Cultured [16, 512, 342, 2978, 2716, 1458]. Cultures [1733, 2683, 2334, 880, 1917]. Cumberland [467, 1303]. cuneifer [1250]. Cunner [579, 2747]. Cupric [2521]. Curcumin [1614]. Current [2805, 1949]. Currents [2090, 2249, 2099, 53, 871, 2091, 2274]. Curvatures [559]. Curve [20]. Curves [1864, 2235]. Cutthroat [461, 2588, 1593, 230, 607, 2850, 75, 2346, 654, 1394, 289, 2032]. cyanellus [356, 1003]. Cyanide [2477, 2785, 2530]. Cycle [945, 2261, 890, 1161, 1585, 239, 2383, 1002, 915, 778, 2443, 2376]. Cycles [2546, 1314, 1139]. Cycling [2548, 2951]. Cyclonic [2087]. Cyclopropenoid [1410, 1367]. Cyclops [1793]. Cyclopterus [758, 2183, 1698, 1907]. cylindraceum [1532, 638, 905]. Cylindrical [2943]. Cymatogaster [1277]. cynoglossus [2017, 2693, 1672, 229]. Cypionate [336]. Cyprinid [2957]. Cyprinidae [1249]. Cyprinodontid [913]. Cyprinodontidae [457]. Cyprinus [2248, 475, 1770, 1122, 2002]. Cyril [482, 483]. Cystidicola [2736]. Cytochrome [2619]. Cytological [527]. Cytoplasmic [2643]. Cytosomes [1716]. Cytotaxonomic [2326]. Cytotaxonomy [1535]. Czechoslovakia [2386].

D [2917, 1327]. D-Mannitol [2917]. Dabob [642]. Dace [242, 202, 583, 616, 1249, 752]. Dactylogyridae [1486]. dahli [758]. Daily [1791, 1254, 2751, 1422, 755, 2220]. Dakota [1613, 1969, 1913]. Dam [789, 1078]. Damage [2129, 1410]. Damaged [2496]. Dana [1002]. dans [1267, 857]. Daphnia [831, 1328, 1524, 2303, 2916, 2820, 2582, 1975, 781, 2034, 839]. Dark [270]. Dart [112]. Dart-Tagged [112]. Darters [735]. Data [350, 2344, 2281, 908, 2211, 2974, 326, 2896, 285, 2274, 2473, 2108, 2766]. Davis [447]. Day [983, 922]. Daylength [1310]. Days [911, 1362, 1691]. DDD [287]. DDE [2180, 287, 1990]. DDT [1126, 2075, 2180, 38, 40, 1725, 764, 1495, 552, 287, 935, 1391, 1434, 2193, 2799, 2035, 2442, 368, 1990, 172, 1608, 1591, 1503, 650, 1451, 1275, 1540, 975, 1158, 279]. DDT/DDE [1990]. Dead [2715]. Death [787, 356, 641]. débarquements [2226]. Debris [997]. Decade [1332]. Decapod [502, 131]. Decapoda [510, 1692]. Decay [1585]. December [1894]. Dechlorination [1489, 2514]. decipiens [2154]. decisum [149]. Decline [1050]. Declining [2401]. Decomposed [569]. Decomposition [2541, 1615, 2451]. Decompression [2844]. Decrease [2856]. Deep [1752, 563, 2560, 1720, 1262, 1779]. Deep-Sea [563]. Deepwater [653, 621, 86, 1689, 2765]. Defecation [221]. Deficiency [2954]. Deficient [1491]. Definition [1616]. Deflectors [779]. Degradation [2606, 1147, 2608, 436, 9, 1384, 368, 1756, 2497]. Dehydrase [948]. Dehydratase [1855, 2247]. Dehydroabietic [2299, 1556, 1860]. Dehydrogenase [1912, 127, 414, 886, 1006, 1299, 101, 223, 2478, 177]. Delay [1007]. Delay-Differential [1007]. Delayed [2044]. Deleterious [2530]. Delineating [1563]. Delphinapterus [1126, 467, 1030, 1629]. Delta [1126]. DeLury [815, 1641]. Demand [1182]. Demersal [1060, 77, 2035, 2684, 706]. demissus [343]. Demonstration [397, 171]. Densities [2158]. Density [2452, 2636, 2976, 141, 2862, 2045, 2305, 549, 1605, 2728, 1387, 413, 462, 737]. Density-Dependent [2636, 2728]. denticulatus [2243]. Dentigryps [313]. Deoxyribonucleic [12]. Dependence [1172, 2975]. Dependent [2636, 2848, 2728]. Depletion [2895, 1888, 466]. Deposit [2607]. Deposit-Feeding [2607]. Deposited [2752, 2602]. Deposition [2830]. Depreciation [2417]. Depressed [2878]. Depression [2852]. Depth [762, 141, 1700, 1368, 186, 229, 1377, 2950]. Depths [956, 768]. Derivatives [916, 234]. Dermal [2106]. dermatology [2329]. Derzhavin [516]. Described [1038, 492]. Description [1486, 534, 2177, 1098, 1146, 252]. Descriptions [1864, 495, 502, 1402, 1365, 486, 2495, 630, 484]. Design [956, 2813, 2971, 2806, 2882]. Designation [2225]. Designed [2798]. Destruction [2349, 1570]. Destructive [1983]. Detailed [1269]. Details [1477]. Detect [1136]. Detecting [427]. Detection [1769, 35, 2691, 2164, 1502, 1027, 598, 2238, 1445, 161, 2684, 2972, 1579].

Detector [1264]. Deterioration [1525]. Deteriorative [2119, 1691]. Determinants [1182]. Determination [2435, 1077, 1528, 960, 2358, 565, 1426, 1735, 2076, 1614, 1429, 2254, 2171, 424, 2816, 1152, 2460, 420, 1561, 1086, 84, 893, 1829, 1778, 1989]. Determinations [2751]. Determine [1075]. Determined [631, 1089, 590, 784, 1581, 1443, 2222, 1425, 1666, 2074, 1254, 2682, 2300, 348, 1994, 1377, 2158, 231]. Determining [374, 916, 2698, 62, 1087, 2649, 1368, 1788, 2283, 2890, 1685, 300, 812]. Detrital [2833]. Detritus [1719, 2831, 2833, 2832, 1458, 2255]. Detritus-Associated [1719]. deux [1253]. Developement [1228]. Developing [1193, 1204, 1795, 1216]. Development [1066, 370, 399, 503, 1492, 2189, 1702, 2584, 2829, 987, 2709, 1495, 1684, 1240, 1230, 1333, 1068, 1531, 2540, 2866, 1352, 1234, 1208, 1219, 295, 729, 2356, 1223, 1552, 1566, 1170, 1868, 1183, 1461, 2126, 1214, 2673, 2320, 1415, 2244, 1584, 1190, 319, 1201, 1189, 81, 1215, 2170, 1200, 1927, 1198, 1224, 1978, 1178, 151, 2773, 2870, 1158, 2489, 2384, 1933, 1191, 1217, 249, 1218, 1210]. Developmental [1249, 1926, 807, 2485, 2184]. Developments [2777]. développement [1210, 1211]. Device [138, 2683, 2646, 2490, 2761, 2649, 847, 284, 2257, 1381, 23, 2213]. Devices [817, 231]. DHAA [1860]. Di- [1547]. Di-2-Ethylhexyl [1547, 2128]. Diagnosis [1496, 2777]. Diagnostic [2784, 1650]. Dialysis [1308, 1035, 880]. diaphana [83]. Diaptomus [2, 2968]. Diatom [2252, 306]. Diatoms [2842, 2590, 2253]. dibranchiata [882]. dichloride [247]. Dichloro [1569]. Dicrateria [1312]. Dictionary [2352, 2794]. Dictionnaire [2794]. Diego [1703]. Diel [945, 71, 2660, 2211, 753, 2904, 1139, 2266, 2074, 36, 1918]. Dieldrin [1764, 1710, 356, 206, 2769, 781, 1158, 1391]. Diesel [1542]. Diet [2811, 1764, 1428, 736, 2270, 13, 917, 2520, 1480, 2117, 1118, 1109]. Dietary [1405, 1992, 1116, 2493, 2704, 2893, 2581, 2980, 2049, 2046, 1721, 1372]. Diets [513, 1258, 1951, 1950, 2271, 2577, 1581, 1020, 1558, 2110, 2043, 443, 632, 1109, 2658]. Difference [2962]. Differences [2782, 1590, 2344, 789, 1403, 17, 2382, 291, 223, 591, 347, 2181, 308, 275, 276, 2214, 2318, 833]. Différenciation [2969]. Different [513, 318, 2434, 714, 1860, 142, 2675, 1184, 1923, 2320, 1718, 2781, 1763, 798, 156, 1253]. Differential [282, 1247, 1437, 2956, 1007]. Differentiating [283]. Differentiation [2969, 1542, 1397, 1631]. différents [1253]. Differing [2173, 993]. Diffusion [1091]. Diffusive [1890]. Digenea [376, 152, 400]. digestif [2867]. Digestion [197, 1123, 1313, 1065, 84]. Digestions [1247]. Digestive [160, 1116]. Digging [900]. Digital [1368, 419, 445, 2473]. digitale [945]. Diisopropylfluorophosphate [1759]. Dimension [2656]. Dimensional [1494, 1685]. Dimethyl [720]. Dimethylamine [190, 270, 1051, 1309]. Dinoflagellate [2125, 2494, 2572]. Dinoseb [2032]. Diodes [1735]. Dioxide [788, 474, 1319, 466, 1297]. Diplocotyle [1269, 778]. Diploid [1127]. Diptera [2525, 2360, 2177, 2068, 2202, 2203, 967, 1530, 1550, 2481, 281, 1853]. Direct [2071, 989, 247, 2535]. Disappearance [2645]. Discharge [1825, 425, 1023, 2613, 998]. Discharges [2856]. Discontinuity [1033]. Discovery [1649, 245]. Discrete [2888]. Discreteness [415]. Discriminant [2508, 1563]. Discriminating [1608, 1245]. Discrimination [1006, 59, 1021]. Discs [22]. Discussion [1871, 2284]. Disease [2262, 34, 1406, 1769, 1785, 1496, 1101, 933, 344, 2978, 1578, 660, 1697, 538, 2955, 1075, 2972, 2216, 2504, 1488, 2167, 904, 2121, 968, 952, 586, 1521, 2106, 104, 2696, 2777]. Diseases [1953, 1679, 2328, 1980, 1650, 1545, 2797]. Disk [2724]. dispar [642]. Dispenser [2076, 540]. Dispersal [653, 417, 2603, 2364]. Dispersant [1140]. Dispersants [2622]. Dispersed [895, 2979]. Dispersion [2910, 1654]. Display [1645]. Displays [16]. Dissociation [1686]. Dissolution [2252]. Dissolved [399, 2477, 435, 2076, 1774, 2757, 540, 1429, 2294, 2460, 2979, 161, 319, 1092, 1086, 2576, 1267]. dissoute [1267]. Distance [2665, 41, 1427, 2161]. Distances [536]. Distant [1235, 1234]. Distant-Water [1235, 1234]. Distillation [978, 424]. Distinct [2318]. Distinction [2588, 2346]. Distinguishing [579, 2715]. Distribution [2743, 1492, 2587, 2437, 2017, 1755, 220, 2472, 2643, 2299, 2298, 1887, 572, 2179, 1299, 2780, 531, 653, 1090, 2645, 121, 35, 590, 2333, 1290, 806, 2444, 323, 11, 73, 502, 53, 545, 1261, 2436, 1921, 1741, 1141, 1666, 2305, 2764, 1485, 335, 2682, 69, 20, 802, 1671, 200, 85, 329, 367, 849, 2821, 2896, 2143, 1034, 2644, 2516, 2359, 1100, 2570, 563, 1631, 254, 1967, 880, 822, 1626, 648, 1903, 1144, 304, 1959, 2605, 2316, 2308, 515, 55, 1098, 1146, 487, 2108, 2340, 1346, 1896, 511, 1047, 637]. Distributional [278]. Distributions [762, 2141, 1493, 1364, 186, 229, 1893, 2300, 2788, 2532]. Disturbance [2810, 1967]. Diurnal [1549, 1144, 1092]. Diver [620, 2158]. Diver-Operated [620, 2158]. Divergence [2934]. divergens [2680]. Divergent [1120]. Divers [2560]. Diversion [2113]. Diversity [521, 883, 2191, 2246, 1981]. Division [1310, 706]. DNA [2976, 262, 887, 2175]. dodacaedron [577]. Dodecaceria [491]. Dodecyl [976]. Does [1108]. dofleini [2737]. Dogfish [2633, 1589, 793, 740, 58, 2114, 2440, 2356, 834, 1512, 2043, 2127, 2853]. Dolly [461, 116, 1315, 943, 654]. dolomieui [1046, 2393, 2390, 2185, 46]. Dolphin [2713, 1633, 1638, 2712, 1632, 1630, 1017]. Dolphins [1631]. Domestic [2011]. Dominant [770]. Dorosoma [2206]. Dorsal [2358, 1728]. Dose [1465]. Doses [2075, 442]. Double [2580, 2957]. Double-Stranded [2957]. douce [2585]. Douglas [2644, 2764]. Dover [819]. Dow [2720]. Down [394]. Downstream [722, 748, 763, 1756, 275, 276]. Dr [1979, 2595]. Drainage [2911, 291, 19, 625]. Dredge [797, 836]. Dredges [884]. Dredging [884]. Dressing [650]. Dried [16, 340]. Drift [2050, 41, 2818]. Drift-Feeding [2818]. Drifting [264, 2250]. Driftnets [1647]. Drip [2371, 861, 970]. driven [1637]. droebachiensis [2233, 2892]. Drop [409]. Dry [1543, 1480, 2110]. Drying [426]. Dual [2908]. dubius [827, 927, 902]. Due [1479, 136, 789, 1696, 433, 2053]. Dugesia [271, 446]. dulcicoles [2969]. dumerili [1163]. Dummy [1996, 1577]. dunbari [2031]. Dungeness [444, 2636, 10, 2162, 2237, 2711, 1441, 2949]. Durable [449]. Duration [1317]. During [756, 2677, 1142, 881, 2201, 1514, 2698, 2543, 1607, 785, 452, 713, 1068, 1531, 1966, 1518, 1662, 1826, 100, 1248, 2930, 1585, 1824, 612, 1384, 2119, 2045, 767, 1122, 2852, 2419, 1078, 1604, 479, 1659, 2197, 2357, 1694, 1246, 655, 442, 2839, 2122, 2175, 270, 27, 258, 9, 311, 1818]. Dusky [1047]. Duwamish [2121]. Dwarf [527, 2225]. Dwelling [1455, 616]. Dyer [848]. Dyes [211]. Dynamic [953, 1468, 2778, 1015, 2430]. Dynamical [1879]. Dynamics [2313, 118, 1114, 2194, 258, 2319, 1751, 1559, 2817, 2544, 2419, 2412, 1548, 2582, 2128, 2438, 2406, 2295, 2951, 2672, 2968, 2117, 2402, 2853].

E. [2720]. Early [399, 503, 1492, 1406, 1423, 2060, 1514, 1434, 1737, 2115, 1960, 1552, 876, 2530, 200, 2320, 2338, 1611, 1028, 2707, 249]. Earthy [1837]. East [2886, 1532, 187]. Eastern [1888, 238, 378, 226, 1212, 1983, 1272, 2517, 653, 1029, 1208, 1209, 652, 2222, 1628, 538, 265, 525, 1365, 563, 1631, 486, 216, 994, 2316, 625, 1134, 1113]. eau [857, 2585]. écailles [2867, 2969, 1253]. Ecdysis [1918]. Ecdysterone [1724, 746]. échanges [2585]. Echinorhynchidae [2219]. Echinorhynchus [2219, 204]. Echo [1368, 419, 465, 463, 462]. Echo-Counting [1368]. Ecologic [679]. Ecological [1401, 2423, 2934, 2546, 1954, 2611, 1866, 616, 752, 2720, 1485, 2078, 1870, 1632, 1762, 411, 2210, 2951, 281, 1738]. Ecology [501, 2061, 2655, 1592, 39, 2604, 1333, 678, 1773, 1682, 2112, 2795, 1256, 2670, 1615, 967, 778, 51, 939, 2953, 2854, 2760, 120, 2381, 2702, 1845, 1853]. Economic [941, 1723, 2499, 2871, 1333, 1209, 1916, 1184, 2876, 2890, 2869, 2870, 1933, 1810]. Economics [1181, 1932, 2741, 2547]. Ecosystem [953, 2643, 531, 552, 2748, 2601, 2833, 2846]. Ecosystems [2606, 2568, 2496, 1352, 2428, 2238, 1505, 1701, 2548]. Ectocotyla [2695]. Eddy [1091]. Edema [1802]. Edge [2498]. Edible [713, 1053]. Edith [482, 483]. edition [1816]. Editor [294, 670, 1165, 1329, 1358, 1876, 2083, 2373, 2594, 2569]. Editorial [2192, 2966, 2137]. Editorial/Editorial [2137]. Education [1332]. edulis [2937, 2892, 1312]. Edward [921, 747]. Eel [772, 2241, 928, 652, 563, 25, 971, 1408, 1900]. Eelblenny [1155]. Eelgrass [2233, 1585, 199, 2255]. Eels [2212, 1396, 1080, 966, 1377, 1134]. Effect [2180, 1764, 2458, 2580, 2735, 197, 2724, 1799, 881, 1405, 185, 2705, 2842, 1009, 2706, 256, 1388, 2069, 1780, 1770, 1962, 2265, 2577, 773, 546, 729, 1991, 970, 1396, 1577, 1317, 2722, 1608, 1591, 2302, 888, 1325, 1321, 1465, 2245, 1459, 1480, 2046, 823, 1451, 771, 2296, 1721, 1041, 1037, 770, 2043, 633, 665, 2255, 2012, 548, 1159, 260, 2855, 1424, 2324, 1986]. Effective [1435, 1605]. Effectiveness [2714, 1488]. Effects [2828, 2248, 370, 399, 503, 159, 38, 2072, 1792, 1690, 1322, 149, 2407, 2452, 603, 693, 831, 93, 1097, 395, 528, 853, 2398, 1076, 2667, 2276, 431, 2785, 2643, 435, 10, 2168, 1613, 884, 640, 659, 1409, 74, 1043, 544, 866, 702, 696, 2923, 703, 987, 1745, 1851, 2554, 1302, 919, 1910, 764, 2663, 1012, 1731, 808, 141, 26, 974, 664, 2415, 2952, 2838, 746, 2188, 744, 2148, 1352, 2071, 382, 1259, 1471, 2365, 606, 687, 684, 2625, 1992, 2016, 165, 1140, 695, 613, 566, 926, 1310, 2000, 1598, 2266, 2039, 53, 362, 2803, 2234, 43, 2442, 52]. Effects [2861, 1003, 681, 2526, 1990, 2945, 2422, 780, 2615, 1566, 2727, 206, 142, 692, 683, 688, 1049, 2270, 2530, 2059, 2238, 402, 938, 690, 2630, 1606, 13, 686, 2673, 2064, 337, 1508, 727, 366, 2143, 922, 1604, 2906, 835, 1582, 1584, 1833, 409, 2133, 2771, 677, 691, 2626, 689, 533, 2739, 364, 315, 2616, 2140, 951, 1967, 1917, 701, 650, 1615, 1151, 1182, 682, 2476, 2379, 1549, 2110, 466, 1759, 2447, 78, 2250, 996, 1387, 755, 263, 615, 975, 1158, 2489, 2613, 634, 2109, 2692, 863, 48, 171, 2161, 433, 345, 805, 940, 2725, 694, 2019, 1516, 685, 1372]. Effects [336]. Efficacy [1255, 801]. Efficiencies [191, 1458]. Efficiency [2929, 1866, 389, 390, 2257, 2980, 817, 1065, 805, 940]. Efficient [2896, 2882]. Effluent [1799, 2567, 2303, 1099, 919, 2294, 2365, 1609, 2292, 1425, 2305, 2290, 937, 2293, 2256, 922, 1362, 1604, 1777, 2244, 2450, 2722, 2906, 76, 898, 2291, 2302, 1385, 2295, 2300, 2296, 2109, 2146, 805, 2304]. Effluents [1651, 2652, 1987]. Effort [2871, 2974, 2892, 2868]. Egg [1280, 2709, 141, 642, 2123, 458, 549, 1158, 2258, 2204, 1794, 2475, 181, 2866]. Eggs [2512, 1795, 501, 1971, 2914, 2639, 2162, 2231, 2123, 284, 2545, 2708, 579, 371, 558, 2170, 1474, 827, 1652, 655, 2949]. Egtved [397]. Egypt [1758]. Eight [814]. Ekman [901, 317]. Ekman-Type [901]. Ekofisk [2601]. ELA [1852]. Elasmobranch [2171]. elassodon [1492, 187]. Electivity [1416, 2814]. Electric [966]. Electrical [769, 2930, 349]. Electrode [374, 2521, 2535]. Electrodes [972]. Electrofishing [2050]. Electrolyte [2324]. Electrolytes [1910]. Electrometric [1077]. Electron [1557, 1264, 1443, 2909, 1100, 1253]. Electronic [326, 2213]. électronique [1253]. Electropherograms [1006]. Electrophoresis [976, 734, 238]. Electrophoretic [2588, 246, 1035, 101, 4, 2346]. Electrophysiological [1904, 59]. Electroshock [1828]. Electrotaxis [809]. elegans [411, 705, 915]. Elemental [129, 713]. Elements [37, 2065, 1123, 72, 1247]. Elevated [2659]. Elevating [34]. Elicited [2679]. Eliminate [594]. Elimination [179, 178, 802, 2143, 1959]. Ellesmere [2094, 762]. elongatus [520, 2358, 1141]. Elution [930]. Embayment [2308]. Embiotoca [804, 803]. Embiotocid [376, 1277]. Embryo [1010, 2320, 442]. Embryogenesis [2357]. Embryonic [370, 987, 729, 2356, 834, 1552, 925, 319]. Embryos [1613, 1409, 1902, 284, 2545, 996, 804, 1516]. Emerald [1923]. Emergence [141, 26]. Emersion [47]. Emitting [1735]. emphaeus [1670]. Emphasis [1774, 1632, 169, 2355]. Empirical [565, 1436, 411]. Emptying [2023, 140, 2491]. Emulsification [1375]. Enclosed [808, 2474]. Enclosures [1411, 2332]. Encounter [2174]. Encrustation [1128]. Encyclopaedic [2352]. End [1519]. Endangered [949]. Endocrine [1944]. endocrines [2867]. Endosymbiont [2232]. Endotoxin [1465]. Endrin [1130, 251, 856]. Energetic [1692]. Energetics [1865, 281]. Energy [1274, 1142, 2633, 1040, 2830, 1772, 1135, 2933, 2656, 1749, 925, 838, 1124, 369, 1383, 23, 705, 2120, 2741, 2117, 2932]. Engineering [1932, 2827, 1847]. Engineers [2352]. England [1546]. English [1712, 1858, 2617, 1472, 2121]. Engraulis [1203, 609]. Enhancement [39, 1418, 2750, 2506]. Enhancing [2064]. Enolase [398]. Enriched [1096]. Enrichment [2319, 385, 1085]. Entanglements [2834]. Enteric [1785, 2164, 2002, 2236]. Entering [41, 2558]. entomon [2616]. Entrained [2770, 1042]. Entrainment [2764, 2864, 1787]. Entry [2874, 2872, 2876, 2875]. Enumerating [350]. Environment [475, 2701, 1786, 2290, 2981, 1252, 380, 147, 168, 1460, 1549, 2741, 2376, 2466, 903, 2461, 2462]. Environment/Selected [2466]. Environmental [1522, 604, 1713, 1524, 2923, 1241, 26, 2342, 2389, 2595, 2937, 2282, 1962, 2135, 695, 2134, 178, 1741, 2810, 1993, 2553, 1775, 2771, 1463, 1436, 1901, 2361, 1967, 2278, 1129, 2403, 2441, 1722, 2619, 2079, 2169, 1717, 2463, 2464, 2546, 2589]. Environmentally [2126]. Environments [1308, 1844, 2289, 1837, 917, 42, 2596]. Enzymatic [2168, 357, 1247]. Enzyme [2735, 651, 1116, 2837]. Enzymes [160, 2784, 43, 130, 7]. Eopsetta [370, 517, 1761]. Eosinophilic [1835]. Ephemeroptera [1262]. Epibenthic [1667, 2013, 986]. Epibiotic [2162, 2237]. epifauna [1001]. Epifaunal [449]. épifaune [1001]. Epilithiphyton [1093]. Epithelial [1716]. Epithelium [1446, 349]. Epizootic [2015]. Epizootiology [1712]. Equation [1857, 87, 1007]. eques [54]. Equilibrium [1780]. Erie [1343, 1894, 171, 1294, 1121, 1885, 1889, 1888, 1892, 1898, 1887, 1713, 2325, 1293, 437, 1291, 1289, 1290, 695, 1884, 1891, 1890, 1897, 200, 1895, 816, 2418, 1886, 169, 1893, 2430, 1879, 1878, 2692, 1881, 1880, 1896, 1882, 1883]. erinacea [261]. Eriocheir [1162]. Erosion [2121]. errans [2949]. Error [1025]. Erythorbate [2791]. Erythrocyte [1127, 2247]. Erythrocytes [1061]. Erythrocytic [1997, 2647, 2912, 2524, 2209, 2340]. Erythropoiesis [277]. Escanaba [2412]. Escape [2717]. Escapement [2781, 1282]. Esocidae [658]. Esocids [2479]. Esox [1446, 1711, 909, 2479, 2830, 1943, 2390, 60, 645, 802, 224, 470, 410, 2537, 1039, 2980, 2403, 996]. Especially [270, 2419, 198]. Essay [1418]. Essays [2701]. Established [1004]. Establishing [1317]. Establishment [1204, 2940, 1785, 2325]. Ester [420]. Esterase [272]. Esterases [357]. Esters [2981, 1561]. Estimate [1135, 815, 836, 1282, 2532, 1567]. Estimated [2699, 705, 2815, 2781]. Estimates [2330, 2862, 989, 465, 22, 2638, 1758, 2260, 754, 2920, 977, 1805, 2223]. Estimating [962, 1309, 1543, 2520, 1025, 1092, 973, 2783, 536, 2905, 1760, 1575, 1778, 2900, 2901, 2902]. Estimation [2939, 61, 203, 28, 1395, 2710, 2631, 463, 2967, 662]. Estimator [1341]. Estradiol [336]. Estrogens [337]. estuaire [1001, 1267]. estuarien [2950]. Estuaries [2073, 2787, 2932]. Estuarine [1968, 248, 918, 1581, 2833, 234, 1798, 2091, 2950]. Estuarine-Type [2091]. Estuary [1272, 882, 2513, 560, 1256, 1517, 1967, 1540, 939, 268, 2121, 1267, 2743, 2437, 1001, 2562, 2840, 2444, 1526, 2831, 2436, 2833, 2938, 1903, 2832, 1982, 2690, 1779, 2145]. étage [1001]. Etheostoma [735]. Ethological [2382]. Ethylene [896]. Ethylhexyl [1547, 2128]. Etiology [2681]. étude [1253, 857]. Eualus [510]. Eubacterium [2230]. Eubalaena [790].

Eubothrium [1464, 2276, 1008]. Eucumbene [2331]. Eukrohnia [599]. Eupera [2029]. Euphausia [2963, 1763]. Euphausiid [2240, 2199]. Euphausiids [2061, 1540, 1903, 2120]. Eurasian [2407, 2405, 2399, 2386, 2385, 2403, 2384]. Europe [2389, 591, 673]. European [2865, 697, 1312]. Euryhaline [143, 571]. Eutrophic [1899, 539, 2404, 2269]. Eutrophication [675, 693, 2398, 696, 703, 687, 2397, 692, 688, 2429, 689, 786, 942, 682, 2804, 530, 1085, 1340, 1354, 1705, 694]. Evacuated [288]. Evaluate [1099]. Evaluation [167, 733, 1407, 962, 183, 2680, 2522, 887, 222, 2457, 2247, 1450, 645, 1776, 2144, 2338, 2876, 119, 1762, 1341, 2908, 896]. Even [610, 2718]. Even-Year [610, 2718]. Events [2890]. Evidence [2435, 1278, 708, 1590, 961, 2021, 105, 507, 2413, 875, 714, 2718, 2732, 2359, 2086, 2491, 177, 829]. Evidenced [1759]. Evolución [1212]. Evolution [1212, 507, 1977, 2869, 778, 412]. Evolutionary [2850]. Exact [20]. Examination [1293, 2310, 2889, 1698, 2951]. Examining [2172, 2707]. Examples [2810]. excavata [2125, 2572]. Excerpts [1871]. Excess [1374]. Exchange [123, 1089, 1988, 2765, 1297]. Exchanges [2228, 1752, 2585]. Excitation [2935]. Excluding [153]. Excreted [2887]. Excretion [1791, 2849, 1732, 1988, 328, 1062, 1763]. Exercise [743, 1839, 1860, 1604, 155, 595]. Exercises [2552]. Exhibiting [960]. Existence [156]. Exoskeleton [2681]. Exosphaeroma [2720]. Expansion [1236, 1635]. Expatriate [829]. Expected [2890]. Expenditure [1142, 2633, 1135]. Experience [2874, 1235, 1185, 2872, 1195, 606, 1859, 1129, 548, 1817]. Experiment [2314, 106, 458, 977, 2737]. Experimental [1690, 2050, 2660, 1411, 1906, 531, 2647, 1811, 2037, 295, 1537, 1020, 2439, 1974, 2729, 530, 1084, 654, 586, 299, 307, 297, 296, 305, 303, 301, 308, 298, 300, 304, 1093, 1354, 2533, 306, 302]. Experimentally [1268, 1578, 2617, 951]. Experiments [389, 2915, 1354, 451, 619]. Expired [183]. Explain [308]. Explanation [576]. Exploitation [1723, 693, 2398, 2387, 696, 2408, 2899, 687, 684, 695, 1820, 1634, 681, 692, 683, 688, 124, 686, 2343, 689, 701, 676, 1175, 682, 2431, 694, 685]. Exploited [2313, 1658, 1559, 2657, 1190, 2336, 973]. Exploration [1179]. Exponential [20]. Exposed [2643, 2198, 1487, 1495, 1571, 1716, 2930, 1806, 2578, 1609, 1552, 1049, 911, 1362, 76, 995, 1151, 2357, 965, 442]. Exposure [2677, 1914, 2914, 659, 764, 2671, 594, 1855, 2247, 2039, 1732, 1860, 2442, 729, 2861, 1990, 210, 922, 1604, 1501, 2618, 2616, 2246, 1768]. Exposures [1793, 366, 2324]. Expressing [2793]. Expression [1962, 465, 2376]. Extended [2895, 2499, 2706, 67, 1160]. Extending [2234]. Extension [359, 66, 404, 1324, 2012, 906, 667, 999]. Extensive [1201]. Extent [2300]. External [2942, 2907]. Extracellular [1852]. Extract [1568, 1437]. Extractability [546]. Extractable [74, 1051, 1043, 881]. Extraction [864, 848, 1264, 1651, 1421, 2783]. Extracts [1968]. Extrusion [2223]. Eye [1444]. Eyed [442]. Eyestalk [746, 2064, 1108, 2774].

F. [1873, 1868]. Fabrication [1326]. fabricii [1155]. Facilitation [1448]. Facilities [1023, 718]. Facing [2564]. Factor [2334, 2712, 15, 1245, 2280, 648, 2905]. Factors [2401, 2591, 2939, 1320, 1713, 2400, 351, 26, 2211, 2641, 2413, 526, 1962, 1822, 1350, 2271, 2422, 835, 2470, 2471, 1112, 2386, 1422, 301, 1909, 679, 641, 2169, 556, 1952, 904, 2051]. Failure [772, 657, 1722]. Fall [720, 2571, 2042, 1891, 837]. Fall-Fingerling [2571]. Fall-Run [2042]. fallai [1964]. Fallout [2273, 2307]. Falls [1369]. Falmouth [2611, 2624]. Family [2375, 658, 1657, 491]. Farming [2985, 2498, 1936, 1934, 1482]. Farran [1402]. fasciatus [240]. Fat [1568, 1152, 1367, 1767]. Fate [2608, 2598, 2639]. Fathead [1849, 1850, 547, 2477, 2785, 435, 1613, 659, 2663, 1143, 2442, 911, 2820, 2128, 1582, 726, 1037, 2366, 2579, 433]. Fatigue [1839]. Fatty [189, 32, 56, 353, 1476, 1298, 1477, 868, 444, 1795, 93, 862, 860, 1263, 1568, 885, 2155, 848, 1581, 1536, 1410, 1367, 1986, 2449, 792, 1118, 1312, 580, 1109]. Fauna [2228, 1478, 789, 763, 30, 1720, 2624, 120, 673, 496]. faune [496]. Fauvel [493]. Fe [2159]. Features [2249, 199, 432, 2806]. Feces [428]. Fecundity [2207, 1280, 422, 375, 943, 2693, 1270, 1554, 2657, 1045, 834, 209, 330, 1963, 2892, 442, 1408, 418]. Fed [1279, 2065, 2577, 1020, 1410, 1367, 1044, 2769, 393, 965, 1145]. Federal [2651]. Feed [1964, 1326, 2049, 1359]. Feeder [1541]. Feeding [395, 528, 853, 490, 1258, 1524, 2717, 2736, 1161, 2607, 2233, 2831, 2265, 1284, 2977, 560, 2070, 2542, 264, 2440, 1443, 2505, 2206, 1254, 626, 36, 609, 533, 1448, 2385, 1480, 1597, 2559, 2818, 2609, 340, 654, 927, 855, 1041, 1036, 2961, 2953, 1313, 1065, 1458, 2063, 2814, 1312, 1073]. Feeding-Behavior [2542]. Feeds [896]. felis [359]. Female [1994, 1942]. Females [283]. Fenitrothion [1371, 1467, 965, 1487]. Fennoscandian [698]. Feral [2955, 2573]. Fermentation [2755]. ferruginea [254, 330, 1475]. Fertilizability [1902]. Fertilization [808, 564, 2750, 371, 558]. Fertilized [2638]. FHM [2666]. Fiber [22]. Fibrillar [2889]. Fibrinolysis [2844]. Fibroma [327]. Fibrosis [2118]. Fiddler [995, 794]. Field [1985, 568, 2520, 1102, 33, 448, 1829]. Fields [966]. Fifteen [305]. Fifth [185]. Fight [1812]. Filamental [1473]. Filamentous [2229]. Fillet [2024]. Filleting [1268]. Fillets [1051, 1309, 1607, 2119, 2483, 1673, 180, 311, 970, 1286]. Filter [2151]. Filtering [1524, 2303, 609]. Filtration [2151, 22, 812]. fimbria [512, 195, 112, 721]. Fin [2358, 740, 1588, 2121]. Financing [1183]. Finback [790]. find [107]. Finding [2330]. Fine [2789, 1446, 1711, 1269, 2602]. Finfish [897]. Finger [197]. Fingerling [1130, 2571, 2769, 951, 2049, 2447, 2844]. Fingerlings [535, 1558, 1742, 1743, 1150, 1372]. Finland [2628, 2399, 2885]. finmarchicus [273, 2193]. Finned [2454, 1641, 1637]. Fiords [762]. First [520, 1803, 1260, 342, 1101, 618, 969, 1074, 1431, 2718, 2523, 1162, 476, 719, 1281, 1417, 723, 19, 54, 1146, 1117, 2500]. Fish [978, 2154, 2180, 1850, 1370, 2461, 2462, 1204, 1489, 1479, 1702, 1077, 1528, 143, 731, 2028, 2228, 1535, 1754, 184, 1995, 262, 1194, 962, 74, 190, 270, 864, 1309, 16, 1719, 1565, 2201, 976, 1375, 1263, 1348, 2705, 1951, 1060, 61, 183, 578, 2445, 608, 77, 2717, 2924, 2069, 2211, 2748, 136, 983, 512, 1966, 2048, 2522, 2975, 287, 1307, 344, 1844, 594, 382, 537, 2365, 236, 44, 887, 1351, 1950, 1435, 2035, 678, 1136, 1543, 1704, 2119, 2483, 1557, 1855, 2247, 2405, 1945, 1737, 847, 2417, 948, 1219, 1741, 1450, 2666, 2797, 1368, 1666, 2764, 390, 2432, 1323, 1152]. Fish [474, 1123, 426, 2583, 1897, 2732, 2770, 598, 2962, 1789, 2794, 802, 2264, 2553, 757, 290, 284, 1653, 2912, 419, 172, 122, 970, 425, 1042, 1606, 1247, 1708, 538, 2257, 2338, 1987, 2244, 2450, 2644, 4, 465, 2516, 409, 571, 2052, 1561, 1669, 1854, 1416, 2471, 1448, 1131, 2684, 1075, 1357, 2822, 1421, 1707, 2618, 1685, 2957, 2510, 2637, 845, 1082, 1276, 650, 2213, 1978, 2132, 1182, 972, 771, 1013, 1699, 1341, 1758, 1482, 1106, 445, 715, 718, 1564, 2186, 2514, 2480, 1909, 155, 211, 679, 913, 2451, 1652, 1021, 276, 973, 870, 2124, 2619, 595, 625, 2169, 1601, 2891]. Fish [536, 54, 1098, 462, 673, 2908, 443, 783, 896, 84, 383, 448, 1580, 2307, 1572, 2905, 2340, 2755, 1511, 1794, 2538, 1952, 2139, 2236, 1805, 2482, 1430, 2224, 619, 89, 586, 1113, 1109, 1941, 1838, 145, 2900, 2901, 2902, 453, 1545]. Fish-Holding [718]. Fish-Population [2553]. Fish-to-Fish [1448]. Fished [1605]. Fisheries [1235, 1236, 2669, 1723, 173, 454, 749, 1056, 1620, 2277, 2964, 2407, 1212, 1494, 1638, 2499, 2879, 1229, 2880, 1240, 1600, 1772, 2408, 2327, 1195, 1234, 1208, 1209, 1304, 1232, 2974, 1867, 1560, 1911, 2157, 1221, 316, 1868, 2054, 2651, 1171, 1184, 1214, 2080, 1642, 2876, 415, 2863, 1228, 1190, 2883, 1222, 2869, 1189, 1215, 1762, 1206, 1198, 701, 1175, 1460, 1736, 1976, 1676, 1680, 2870, 2536, 2547, 2668, 2409, 2777, 1191, 1504, 2766, 1007, 1760, 1845, 1818, 2535, 2882, 2465, 1218, 2965, 1210, 1230]. Fisherman [2882]. Fishermen [2593, 1223]. Fishery [941, 2401, 2261, 1646, 1203, 1231, 2636, 1658, 1635, 1114, 2947, 1241, 744, 2872, 2484, 1181, 2314, 695, 1176, 1820, 1216, 1916, 331, 1015, 1677, 106, 1170, 1179, 326, 1183, 200, 1461, 2878, 2044, 2886, 13, 1624, 2404, 2439, 1420, 169, 2834, 1200, 1846, 924, 1178, 451, 1226, 1217, 1645]. Fishes [150, 163, 2219, 376, 2856, 2113, 113, 1912, 1616, 1121, 956, 674, 1679, 2328, 2178, 628, 2349, 1127, 1006, 2375, 2279, 653, 1038, 436, 1607, 2866, 2649, 344, 590, 2333, 912, 1374, 2217, 1285, 1373, 899, 381, 1364, 1935, 94, 504, 982, 1250, 1443, 2115, 2301, 2305, 389, 2205, 809, 2429, 1929, 72, 1804, 410, 1544, 539, 514, 877, 2191, 1321, 500, 1104, 2501, 1059, 873, 672, 830, 1462, 1650, 1908, 1036, 279, 2468, 2573, 2692, 632, 408, 706, 1277, 2801, 2502, 464, 2687, 2688, 1523, 2329]. Fishexpo [2654]. Fishing [1618, 1617, 894, 2871, 1239, 2415, 2838, 1433, 1857, 2886, 2868, 2010, 2527, 1810, 1205, 1836]. Fishway [452]. Fistulicola [274]. Five [2349, 2954, 43, 1335, 1112, 303, 1676]. Fixation [2233, 1557, 2353, 1001]. Fixed [1733]. Fjord [2092, 2544, 2093, 1709, 2308, 2562, 2938, 2765, 1779]. Fjords [2093]. flagfish [2066]. Flameless [84]. Flat [1322, 1312]. Flatfish [1124, 1955, 2158]. Flathead [1492, 2242, 187]. Flats [421]. flavescens [1079, 2391, 2060, 2141, 2415, 2310, 649, 2786, 2394, 2419, 1825, 2424, 1054, 1718, 2751, 2406, 1767, 1854, 1039, 1463, 2409, 120, 2381, 2414, 839]. flavidus [708, 2816]. Flavor [860, 1048, 1837, 142]. Flavored [1484]. flavus [19]. Flea [3]. Fleet [2817, 2873]. Fleets [2881, 2877]. Flesh [862, 861, 2543, 1607, 35, 2119, 2483, 2234, 1673, 1775, 1986, 2304]. fleuve [1730]. Flexibacter [1688, 1998]. Flin [2307]. Floating [102]. Flon [2307]. Flora [501, 1908, 1439]. Florida [1630, 2624, 1017, 2230]. floridae [913, 2066]. flos [1824]. flos-aquae [1824]. Flounder [2017, 2693, 2448, 105, 1686, 222, 1268, 1364, 439, 1672, 103, 254, 330, 1475, 229, 268, 1158, 524, 1827, 556, 2158, 2121]. Flounders [459, 1073]. Flow [1573, 742, 1064, 2656, 780, 1124, 369, 2699, 655]. Flow-Through [1573, 2699]. Flowing [1734, 2559, 2086]. Floy [640]. Fluctuating [789, 1844]. Fluctuations [2341, 2266, 1222, 2420, 2779, 2640]. Fluid [2767, 1474]. Flukes [400]. Fluorescence [506, 1027, 2222]. Fluorescent [1769, 1502, 598, 2635, 2972, 211]. Fluoride [1378, 424, 2739, 2963]. Flushing [2714]. fluviatilis [2407, 2405, 2399, 2386, 2385, 2403, 2384, 2381, 2376]. Foam [1984]. Following [1914, 772, 471, 2160, 1153, 2264, 1501, 1463, 20]. fontinalis [2180, 1795, 954, 1280, 1423, 118, 1283, 1971, 640, 974, 2131, 1035, 162, 1662, 2011, 428, 2047, 2571, 1700, 1848, 2931, 607, 566, 1143, 1010, 2039, 362, 363, 767, 1122, 923, 331, 2459, 739, 329, 124, 2673, 727, 210, 366, 1317, 2143, 1584, 1833, 2955, 1018, 1821, 1031, 1467, 434, 340, 51, 2105, 1787, 5, 2258, 1579, 1521, 2696]. Food [56, 2685, 2686, 1204, 116, 2660, 271, 1755, 322, 2322, 810, 962, 1527, 909, 1866, 1161, 892, 2229, 2442, 2440, 795, 172, 2, 187, 854, 1303, 571, 405, 850, 2825, 1058, 2385, 1660, 1913, 927, 1564, 2480, 2832, 263, 1610, 2384, 634, 641, 1065, 2067, 2395, 706, 805, 940, 903, 2900, 2901, 2902, 1619]. Forage [2748]. Foraging [2904]. Force [565, 965]. Force-Fed [965]. Foreign [2877]. Foreshortened [2775]. Foresight [1522]. Forest [2568]. Foreword [294, 670, 1165, 1329, 1358, 1621, 1876, 2083, 2373, 2594]. Form [714]. Formaldehyde [1043, 864]. Formalin [615, 554, 1273]. Formalin-Treated [615]. Format [15]. Formation [2371, 395, 528, 190, 1915, 1536]. Former [1642]. Forming [2851]. formosa [2554]. Forms [596, 527, 2477, 1885, 1094, 1480, 2318, 1882, 1883]. Formulated [896]. Formulating [1950]. Formulations [1327, 783]. Fort [2202, 2203]. Fortran [350, 15, 325]. Fortran/Format [15]. Forty [303]. Forty-Five [303]. fossilis [633, 1942]. Fouling [2162, 2237, 2949]. Found [221, 2120, 314]. Four [1121, 2179, 653, 2048, 1270, 1350, 613, 2393, 504, 1163, 538, 1969, 2406, 1365, 1821, 1462, 156, 2034]. Fourth [185, 1085]. FPC [1136]. Fraction [549, 1768]. Fractions [2757, 2558, 2894, 822]. Frames [1607]. Framework [2469]. France [1024]. Franciscana [1633]. franciscanus [2261]. Fraser [1350, 1019, 1063]. Frazer [2940]. FRB [29, 50, 68, 90, 115, 139, 158, 174, 188, 215, 241, 293, 310, 334, 361, 388, 407, 430, 456, 481, 543, 292, 309, 333, 360, 387, 406, 429, 455, 480, 542, 582, 601, 623, 644, 669, 725, 751, 776, 800, 825, 852, 581, 600, 622, 643, 668, 724, 750, 775, 799, 824, 851, 879, 914, 936, 957, 979, 1000, 1026, 1057, 1083, 1110]. Free [189, 1454, 862, 860, 2477, 2362, 1263, 1839, 2941, 2103, 1536, 1986, 2124, 2175].

Free-Ranging [2941]. Free-Swimming [1454, 1839, 2124]. Freedom [1703]. Freeze [2458]. Freeze-Up [2458]. Freezing [2580, 713, 180, 1383]. French [1210, 1079, 910, 2226, 1001, 1267, 2950, 649, 496, 1730, 857, 2969, 2585, 1253]. Frequencies [428, 2120]. Frequency [2097, 2099, 2896, 1827, 2108, 2532, 1579]. Frequency-of-Capture [2108]. Fresh [2154, 785, 100, 82, 757, 2068, 1086]. Freshwater [2568, 271, 772, 2842, 2709, 346, 2543, 983, 959, 191, 1270, 1374, 594, 382, 1373, 612, 1682, 368, 1547, 1368, 2290, 426, 1509, 1988, 2354, 2807, 970, 539, 291, 2910, 2637, 2805, 348, 830, 2715, 1036, 1657, 2489, 783, 383, 2224, 2586, 464, 1523, 2969, 2585]. Freshwater-Resident [291]. Freund [1466]. Front [2936, 2970]. Frozen [2371, 2584, 2791, 862, 861, 860, 190, 270, 1051, 881, 1309, 2543, 713, 1612, 222, 2119, 2483, 2234, 1673, 180, 1386, 1325, 1694, 1384]. Frozen-In [1386]. Frozen-Stored [1309]. Frustules [2252]. Fry [350, 2075, 1795, 559, 167, 853, 1613, 2411, 587, 722, 26, 1822, 2187, 1870, 1863, 1256, 1804, 2545, 119, 479, 1558, 2447, 1041, 1768, 2258, 42, 2015, 1742, 839, 1424, 1873, 1868]. Fuca [520, 2557]. Fucus [2626]. Fuel [2598, 557, 1111, 2625, 2597]. Fulton [853]. Function [2377, 2629, 2555, 2172, 2508, 1563, 1759, 2619, 2959, 2747]. Functional [1808, 2631]. Functionality [1375]. Functions [2470, 1436]. Fundulus [16, 2671, 457, 969, 43, 1902, 2848, 2256, 2719, 2959, 413, 48, 1516]. Fundy [1627, 790, 656, 192, 818, 935, 2763, 1073, 2855, 464]. Fungal [771, 1013, 1699, 1128]. Fungus [844]. Furanace [1516]. furcatus [422]. Further [105, 157, 492, 85, 329, 849, 1970]. Furunculosis [1255, 933, 266]. fusca [486]. Future [698, 697, 1170, 2805, 699, 700, 2812, 1949]. fyviei [1486].

G. [325]. Gadidae [833]. Gadoid [270, 1051, 1043, 1309]. Gadus [399, 1997, 605, 2976, 2517, 759, 98, 2453, 1749, 1693, 1775, 1325, 2186, 1366, 1907, 2491, 2340]. Gaetanus [1107]. Gaffkemia [1490, 1488]. Gaffkya [641]. Gain [2478]. gairdneri [1265, 1406, 164, 1279, 2851, 2588, 559, 604, 1899, 416, 788, 1097, 2660, 1656, 2735, 1925, 2785, 2168, 787, 1785, 1592, 123, 1428, 1904, 1920, 2160, 2706, 203, 509, 865, 351, 67, 1839, 2767, 1135, 2641, 2859, 892, 1822, 856, 2691, 2503, 2016, 2930, 2129, 2849, 2457, 2577, 2118, 926, 2000, 2578, 2266, 1688, 1695, 1513, 2155, 707, 2803, 1261, 773, 1048, 1837, 1943, 264, 417, 1502, 546, 397, 2909, 2107, 1990, 378, 1552, 1566, 2727, 1537, 1807, 1651, 2270, 1020, 1991, 286, 1595, 367, 849, 1859, 938, 1410, 1367, 1986, 1432, 1576, 535, 878, 1415, 2722, 1608]. gairdneri [1591, 2955, 1969, 2545, 2771, 1501, 1014, 944, 1100, 1028, 2570, 1974, 558, 1421, 2585, 1901, 2769, 2837, 1451, 2346, 2478, 2181, 1151, 1129, 2946, 1558, 1452, 1998, 2504, 1474, 1062, 2773, 1828, 1387, 615, 1491, 2815, 1569, 2357, 1587, 2214, 2928, 2318, 2063, 1687, 964, 993, 548, 737, 986, 2725, 97, 135, 554, 846, 990, 1273, 1575, 2139, 2634, 2847, 442, 2304, 952, 1742, 1743, 1521, 2658, 1150, 1372, 2122, 2175, 2324]. Galatheidae [1692]. galbana [1312]. galeata [2582]. Galvanonarcosis [2706, 1537]. Galway [2865]. Gametes [1902]. Gamitagama [2368]. Gamma [364]. Gammarellus [766]. Gammarid [2073]. gammaridean [1546].

Gammarus [149, 547, 1809, 3, 1068]. Gap [1174]. Garrick [1146, 1125]. Gas [2592, 123, 1089, 1403, 2222, 660, 420, 1561, 1534, 1086, 5, 2216]. Gas-Bubble [660]. Gas-Chromatographic [1086]. Gas-Exchange [1089]. Gas-Stripping [1086]. Gases [1374]. Gaspereau [999]. Gasterosteidae [324]. Gasterosteus [1379, 14, 96, 1049, 79, 413]. Gastric [197, 2023, 888, 1065, 140, 2491]. Gastrointestinal [2928]. Gastropod [1272]. Gastropoda [2611]. Gastrulae [1495]. Gatherers [894]. Gatnmarellus [765]. Gauge [134]. Gear [1393, 1618, 1617]. Geiger [977]. Gel [976, 734]. Gelatin [1020]. gen [2360, 2481]. Gene [1261]. Genera [605, 2551, 833]. General [2627]. générales [2867]. Generalization [87]. Generated [2098]. Generating [1666, 2764, 2644, 1652]. Generation [2710]. Generations [341, 1270, 2039, 2143, 365]. Genes [1912, 2142]. Genetic [604, 1423, 1790, 1954, 2611, 2838, 2263, 1956, 1403, 1957, 878, 2359, 1955, 2181, 1753, 1580]. Genetics [2178, 127, 414, 886, 1958, 1704, 2227, 1753]. Genic [423]. Genie [1024]. Genotype [223]. Genotypes [2167]. Genotypically [2318]. Genus [113, 921, 1127, 1038, 2695, 583, 752, 238, 313, 245, 2013, 489, 735, 2177, 1631, 967, 1530, 1550, 1488, 54, 1277]. Geochemist [2595]. Geochemistry [1884]. Geographic [318, 2517, 2731, 545, 372, 1574]. Geographical [1299, 2359]. Geography [297]. Geological [1347]. Geomagnetic [1102]. George [2720]. Georges [1658]. Georgia [890, 793, 2341, 2356, 147, 2854, 2090, 988, 473, 2099, 201, 950, 637]. Georgian [1265, 572, 1283, 136]. Geosmin [1823, 1484]. Geostrophic [663, 859]. Geotechnology [1847]. German [2922]. Geryon [1596]. Get [2775, 2464]. Ghana [1235]. Giant [473, 358, 1948, 2443]. Gibbons [1277]. gibbosus [997, 2505, 2519, 1459]. giganteus [532, 823, 822, 1145]. gigas [1274, 1790, 2948, 154]. Gila [2326]. gilae [2326]. Gilbert [1444]. Gill [2929, 509, 865, 2952, 58, 911, 2746, 589, 2909]. gilli [1098]. Gillnet [815, 989, 1739, 2314, 1648, 392]. Gills [1097, 184, 1278, 578, 1271, 1473, 1826, 409, 1100, 120, 177, 770]. Girard [446]. Girth [847]. Given [941]. Gizzard [2206]. Glacial [672, 472, 673]. glaciale [11]. glacialis [790, 273]. Glaciated [674]. Glacier [11]. Glaciers [1709]. gladius [1053]. Gland [2021, 611, 336]. glandes [2867]. Glands [2448]. Glass [374, 22, 1021]. Globicephala [2454, 1641, 1637]. globiceps [875]. Glomerular [2118]. Glucose [82, 130]. Glucuronidase [214]. Glue [1698]. Glugea [816]. Glutamic [206]. Glycera [882]. glycero [2155]. Glycerol [39, 886, 1513]. Glycerol-3-Phosphate [1513, 886]. Glycerophosphate [101]. Glycogen [2000, 470, 1437]. Glycolysis [1384, 1383]. Glycolytic [224, 1325]. Glycoproteins [441]. Glyptocephalus [2017, 2693, 1672, 229]. Goals [2173]. Gobius [592]. Goby [592]. Golden [2904, 2366]. Goldeye [1905]. Goldfish [761, 1849, 1850, 1690, 764, 251, 1943, 568, 981, 2507, 1392, 1842, 125, 2928]. Golfe [2226]. Gonad [739, 838, 2537, 1408]. Gonadectomized [1388, 337, 1508, 336]. Gonadectomy [256, 258]. Gonadotropic [2160]. Gonadotropin [564, 1388, 633]. Gonadotropin-Injected [564]. Gonads [2830, 2018, 1003, 2537, 833]. Gonatid [534]. Gonyaulax [2125, 2572]. gorbuscha [733, 1407, 2142, 2942, 1553, 564, 1016, 2718, 1776, 2126, 1063, 1768, 392, 365]. Gould’s [453]. Gourami [277]. Governments [1239]. Grab [901, 317]. Grabs [1393, 191]. gracilis [1402]. Graded [2577]. Gradient [2333, 568, 2727, 2191, 2894]. Gradients [2761, 2095, 535]. Gram [2349, 1650]. Gram-Negative [2349, 1650]. Grand [2693, 83, 330, 1058, 1475, 1664, 418]. Grande [649]. grandis [2256]. Grands [2969]. Granules [2859]. Graphical [28, 1760]. Graphics [1494]. Grass [1994, 2928]. Grassy [2528]. Gravel [1407, 2789, 141, 1859]. Gravity [732, 2486]. Grayling [1499, 1594, 372, 2053]. Grazing [1983, 165, 2, 2485, 2814]. Great [1589, 358, 1175, 679, 700, 2153, 1344, 1288, 1346, 1342, 1295, 1348, 1345, 1038, 67, 1741, 1740, 681, 1825, 2750, 72, 2644, 786, 1275, 2804, 2420, 1347, 2348, 1963, 2576]. Green [356, 1003, 2156, 1823, 2268]. Greenland [2903, 1634, 209, 1647, 1533, 551, 1420, 176]. Greenling [404]. gregarium [1033]. Grey [1298, 2312]. groenlandicus [958, 1029, 2339, 2819, 21, 855, 872, 2697]. Gros [2745]. Gross [1274, 1103]. Ground [884, 730, 1519]. Groundfishery [2878]. Group [664, 2177]. Groups [2505, 1184, 2896]. Growth [2248, 777, 929, 1792, 1764, 2437, 1308, 2330, 960, 1251, 1120, 831, 853, 1076, 2677, 2017, 513, 1754, 1865, 467, 2976, 262, 640, 659, 2511, 39, 2733, 2554, 2842, 230, 1594, 656, 1333, 244, 585, 2363, 2838, 1270, 1662, 1786, 2188, 282, 2937, 2612, 1555, 2077, 1350, 28, 1067, 100, 1248, 887, 11, 2625, 73, 1585, 473, 2018, 926, 2266, 1046, 1045, 2436, 905, 389, 390, 391, 834, 1512, 2526, 1552, 832, 1414, 20, 2673, 2064, 2317, 403, 727, 1923, 2320, 611, 366, 1362, 835, 2537, 1582, 1584, 2027, 1520, 1969, 2739, 315, 1761, 2616, 1459, 2769, 1475, 2049]. Growth [880, 1917, 2181, 2010, 2441, 648, 151, 1462, 1017, 855, 1721, 1722, 1041, 1037, 148, 2043, 2146, 842, 65, 2147, 2774, 46, 1827, 42, 1145, 1511, 1794, 345, 1952, 805, 940, 442, 893, 2125, 2304, 649, 1730]. Growth-Enhancing [2064]. grypus [1298, 2312]. Guide [2655, 2829]. Guiding [779]. Guilds [1616, 2423]. Guinea [484]. Gulf [2437, 2226, 2436, 1693, 415, 728, 1563, 1515, 1540, 2788, 2041, 484, 958, 2061, 2240, 2947, 2087, 2780, 656, 663, 1834, 935, 2123, 791, 1993, 895, 2256, 1070, 877, 2885, 1830, 229, 1720, 1903, 2120, 1564, 1399, 998, 1831, 2169, 2001, 2858]. Gulls [2723]. Gunnel [240]. Guppies [711]. Guppy [781]. Gustatory [1050]. Gut [2065, 926, 1581]. Guthion(R) [1849]. Gymnoblastic [1117]. Gymnocephalus [2399]. Gynogenesis [1994]. gyrinus [1538].

H [2075, 2645]. H-p [2075]. H-Testosterone [2645]. Habitat [1593, 574, 2188, 607, 400, 1825, 2425, 144, 108, 1601, 2297, 1312, 2224]. Habitats [2378, 1581, 2429, 2388, 2531]. Habits [2736, 983, 133]. Habituate [657]. Habrobaenus [2481]. Haddock [1414, 311]. Hagerman [1279]. Hagfish [108]. Hake [2782, 2119, 2483, 1673, 49, 850, 463]. Halargyreus [44]. Half [1441]. Half-life [1441]. Halibut [862, 861, 860, 209, 1533, 1226, 998, 176, 1055]. Halichoerus [1298, 2312]. Halifax [2089]. Halosphaera [32]. Hamell [2659]. Hampshire [2802, 2573]. Hand [2323]. Hand-Operated [2323]. Handbook [1980]. Handling [657, 544, 866, 362, 363, 180, 595, 2563, 846, 2139, 1430]. Haplosporidan [154]. Harbor [1298, 2611, 1049, 2619, 2760]. Harbors [494]. Harbour [221, 741, 373, 220, 818, 935, 1647, 2449, 314]. Hard [2507, 1458]. Hard-Clam [1458]. Hardshell [319, 836]. harengus [2024, 2584, 2791, 2639, 1612, 2045, 560, 1993, 415, 1515, 347, 2524, 1924, 950, 2494, 2041, 2223, 2488, 2487, 2024, 656, 1612, 1993, 415, 1515, 347, 2524, 1924, 648, 2494, 2041, 2223, 2488, 2487, 3]. Harmful [1855, 2247]. Harnisch [2360]. Harp [958, 1029, 2339, 2819, 855, 872, 2697, 2983, 2984]. Harpacticoid [2322, 2685, 2686]. Harvest [1995, 1717]. Harvested [2561, 239]. Harvesting [2652, 1698, 1949]. Harvests [2403]. Hatchability [2258]. Hatched [1495]. Hatcheries [1360, 655]. Hatchery [167, 733, 2705, 1156, 339, 2047, 2513, 2118, 707, 109, 1415, 2181, 2773, 2961, 1011, 2063, 1311, 2311]. Hatchery-Reared [2047, 2513, 2118, 707, 2773, 2961]. Hatchetfish [83]. Hatching [1766, 878]. Hawaii [1818]. head [2585]. Headwater [1265]. Healing [2488]. Health [2306]. Heart [1298, 972]. Hearts [533]. Heat [253, 2652, 2094, 2321, 1286, 990, 1424]. Heat-Sealed [1286]. Heat-Treated [2321]. Heated [1385]. Heavy [2267, 2734, 1782, 2598, 2659, 1715, 1382, 383, 2307]. Heavy-Metal [2307]. Hecate [98]. Hedriodiscus [281]. Heinz [2198]. Held [2119, 1319, 1438, 2361, 2216, 1818]. Helisoma [1270]. Helminth [376, 485, 934, 902, 1277]. Helminths [2713]. Hemagglutinogens [450]. Hematocrit [933, 1372]. Hematologic [2163]. Hematological [248, 1860, 2719, 2450, 2167]. Hematology [1808]. Hematopoietic [34, 150, 163, 1406, 1843, 2361, 289, 2642, 2015, 1113]. Hematozoa [1472]. Hemitripterus [338]. Hemiuridae [152]. Hemlock [1968]. Hemocytes [1363]. Hemoglobin [933]. Hemoglobins [1962, 156]. Hemolymph [2740]. Hemolytic [2198]. Hemorrhage [721]. Hemorrhagic [1974]. Hemosiderin [277]. Hepatic [1513, 1675, 1901]. hepatopancreas [857, 857]. herbarum [1699]. Herbicides [2497, 2032]. Herd [2713, 2712]. Here [2775, 2464]. Hermaphroditic [355, 1498]. Hermaphroditism [49]. Herring [189, 2024, 503, 2584, 2791, 208, 1428, 987, 656, 2639, 1612, 1232, 2045, 560, 729, 1020, 1993, 415, 728, 1563, 1515, 2723, 533, 347, 2524, 950, 648, 2730, 515, 2473, 2494, 2041, 2223, 2488, 2487, 3]. hertwigi [816]. heteroclitus [2671, 457, 43, 1902, 2848, 2719, 2959, 413, 48, 1516]. Heterogeneity [168]. Heteropneustes [633, 1942]. Heterotanytarsus [1530]. Heterotrophic [1517, 1096]. Heterotrophy [1852]. Hexagrammos [404]. hexapterus [1924]. Hexavalent [1849, 2785, 2168]. Hexose [858]. Hg [2186]. HgCl [2570]. High [1236, 320, 2861, 739, 2770, 549, 119, 1383, 2758, 393, 2146, 980, 2149, 1340, 2674]. High-Energy [1383]. High-Seas [1236]. High-Speed [119]. High-Zinc [393]. Highlights [1243]. Highly [2801]. Hindsight [1522]. Hiodon [1905, 1555, 2077, 1281]. Hippoglossoides [1492, 2448, 2242, 222, 838, 187, 1058, 1664, 1574, 1564, 209, 1533, 176]. Hippoglossus [862, 861, 860, 998, 1055]. Hirudinoidea [346]. Hirundichthys [1163]. hispida [2337, 1808, 2071, 2941, 1105]. Histochemistry [651]. Histologic [289]. Histological [339, 248, 2018, 336]. Histopathologic [586]. Histopathological [2671, 2719, 1578]. Histopathology [2696]. Historical [2085, 1714, 676]. Histories [2811, 2810, 2202, 2805, 1562, 2812]. History [2423, 2809, 1638, 1295, 2025, 988, 1737, 1960, 791, 378, 2775, 832, 2732, 2807, 1044, 2689, 1670, 1630, 2808, 2806, 1663, 2745, 411, 281, 2269, 2967]. Hjälmaren [2403]. hoferi [844]. Holding [102, 1837, 2270, 2239, 718]. Hole [2611]. Holman [1610]. Holt [1114]. homard [2226]. homari [1596, 766, 641]. Homarus [1066, 1724, 272, 900, 160, 1405, 1668, 398, 185, 1497, 2272, 2021, 128, 664, 2836, 585, 1068, 1161, 1531, 1766, 2456, 746, 2665, 1598, 1116, 2893, 2044, 2064, 161, 835, 984, 992, 898, 1371, 1500, 1438, 1748, 2740, 1363, 1072, 33, 427, 2020, 1108, 634, 641, 1490, 1488, 998, 1918, 1753, 2774, 2019, 3]. Home [708]. Homeostasis [1827]. Homing [2142, 708, 1590, 1906, 1920, 1858, 1148, 1447, 320, 875, 59, 75, 1018, 1011, 2331]. Homogamety [1994]. Homogenates [431]. Homoiotes [1157]. Homotetramers [1022]. Honolulu [1818]. Horizontal [2141, 535]. Hormone [2248, 2021, 2160, 2018, 1807, 611]. Hormones [1940, 1939, 1471, 1941]. hornorum [246]. Host [2219, 512, 2363, 844, 2195]. Hosts [204, 485, 934]. Hour [1972]. Housing [664, 1598]. Howe [2308]. hoyi [228, 2542]. hr [922]. Hudson [1030, 2101, 1285, 859, 19]. hudsoni [2481]. hudsonius [1039]. Hull [883]. Human [2669, 1603, 1206]. Humboldt [154]. Humic [569]. humicola [1013]. Humidity [142]. Humoral [2851]. Humpback [2682, 1567]. Humus [2215]. Hunger [1499, 737]. Hunters [894]. Hunting [1643]. Huron [1343, 437, 1294, 118, 693, 550, 1478, 1553, 2929, 1271, 136, 1911, 2764, 1498, 2105, 2416]. Hurricanes [2855]. Hyalella [73, 369]. Hyas [2695]. Hybrid [1006, 1160]. Hybridization [1867, 1160, 1520]. Hybrids [954, 1280, 1423, 246, 2479, 2011, 1848, 923, 735, 947, 1974, 2918, 2346, 5]. Hydroacoustic [2473, 2920]. Hydrocarbon [2629, 1782, 2973, 268, 2605, 2614, 2747]. Hydrocarbon-Utilizing [2605]. Hydrocarbons [978, 2608, 2337, 2610, 2222, 2602, 1112, 1542, 2449, 1031, 2609, 2613, 1765, 2597]. Hydroelectric [1522, 789, 1352, 1078]. Hydroelectrical [2113]. Hydrogen [761]. Hydrogenated [533]. Hydroid [1117]. Hydroids [175, 602]. Hydrolase [1149]. hydrophila [2503, 1909]. Hydrophobic [1726]. Hydrophone [870]. Hydrosoil [1547]. Hydrostatic [1068, 1531]. Hydroxide [1037]. Hydroxycorticosteroids [612]. Hydroxylase [2614]. Hydrozoan [1455]. hyperborea [243]. Hyperecdysonism [1724]. Hyperiid [1671]. Hyperoodon [1627].

Hyperosmotic [2196, 2641]. Hyperplasia [1432, 234]. Hypocalcemia [2946]. Hypochlorite [1536]. Hypochlorous [2362]. Hypocholesterolemic [63]. Hyponeustonic [1966]. Hypophysectomized [2448]. Hypophysectomy [261]. Hypotheses [1827]. Hypothesis [341, 787, 1906, 458, 308]. Hypoxia [787, 338].

I.B.P. [2656]. IBM [350]. IBP [1351]. Ice [2606, 2096, 2895, 2661, 2490, 991, 2103, 9, 2119, 122, 1386, 730, 1325, 2449, 1746, 2001]. Ice-Free [2103]. Iced [1051, 1659]. Icelanonchohaptor [1486]. Icelus [561]. Ichthyocodium [1117]. Ichthyofauna [1163]. Ichthyology [2566]. Ichthyomyzon [207]. Ichthyophonus [844]. Ichthyoplankton [2743]. Icing [713]. ICNAF [1231]. Icosahedral [2957]. Ictaluridae [1050]. Ictalurus [1130, 1764, 955, 1302, 2706, 1910, 1115, 2229, 1992, 2978, 357, 1425, 1300, 1675, 1538, 1730, 1578, 2581, 951, 1913, 869, 1959, 2919, 580, 1109]. Idaho [574, 607]. idella [1994, 2928]. Identification [501, 610, 1769, 37, 1006, 2212, 2048, 937, 1651, 69, 1510, 2518, 2366, 1823, 1128, 2755, 1783, 662, 1113, 1484]. Identify [2132]. Identifying [728]. Identity [1629]. IHN [1406, 2666]. IHNV [2642]. II [163, 2621, 1188, 376, 2540, 2670, 2832]. III [2795, 2867, 1202, 286, 2630, 2833, 1277]. IJssel [2402]. illecebrosus [1641]. Illex [1641]. Illinois [2927]. Illumination [2727, 1151]. Illustrated [2354]. Immature [225, 2933, 1262, 508]. Immatures [2481]. Immediate [2625]. Immersion [2851, 2792, 222]. Immobilization [1537]. Immune [1406, 460, 1466, 2357]. Immunization [2196]. Immunodiffusion [579]. Immunology [1707]. Imol [1561]. Impact [1522, 1240, 1333, 2282, 2931, 1884, 2886, 2749, 2873, 2278, 2624, 2474, 1205, 2568]. Impacted [2611]. Impacto [1205]. Impacts [2499]. Impaired [1495]. Impairment [2304]. impi [126]. Implanted [972]. Implication [2731]. Implications [189, 2592, 2435, 1229, 2871, 2880, 313, 2881, 2206, 2710, 324, 2619, 1430]. Importance [1423, 1342, 2380, 2607, 1660, 2367, 2833]. Important [1094]. Imported [1908]. Imposed [2877]. Imposition [2856]. Impossible [1059]. Impoundment [1719, 275, 276, 2700]. Impoundments [60]. Imprinted [1906, 1904, 1920]. Improve [2564]. Improved [1734, 901, 2076, 2260, 754, 1021]. Improvement [381, 1673]. Improvements [1956]. improvisus [1503]. Inability [1136]. Inactivation [1653]. Inadequacy [783]. Inbred [164, 559]. Inbreeding [2107]. Incidence [2681, 86, 1075, 1574, 1564, 814, 1572]. Incidental [1648]. Incidents [1076]. incisa [885, 2255]. Including [1407, 713, 502, 2553, 1804, 1777, 2450, 514, 219, 321, 2232]. Incomplete [321]. Inconnu [929, 2182]. inconstans [821, 402, 2359, 263]. Incorporated [1733]. Incorporating [2539]. Incorporation [1513, 2155, 2255]. Increase [1724, 1497]. Increased [2050, 479, 1768]. Increases [1012, 1731, 1032]. Increasing [1938, 393]. Incubating [501, 2541, 2708]. Incubation [2512, 655]. Incubators [1407]. Independence [2491]. Index [269, 562, 2028, 2265, 2933, 1160, 2703, 1416, 2441, 1341, 1728]. Indexing [2937]. India [1195, 1219, 1935]. Indian [484, 1214, 1928, 1215, 906]. Indicated [2736]. Indication [1378]. Indicator [1855, 2247, 1133, 2441, 2679]. Indicators [1076, 262, 974, 2954, 2723, 2315]. Indices [1140, 2825, 2814]. indicus [493]. Indies [1638, 1567]. Indirect [2071, 1502, 598, 2972]. Individual [810, 820, 1457, 2004, 1805]. Individuals [820, 1457, 2132]. Indonesia [1218]. Induced [2851, 2448, 1940, 918, 1507, 2578, 1252, 2126, 2946, 2321, 2919, 1805]. Inducible [2747]. Inducing [809, 1490]. Induction [2726, 1990, 2729, 2959]. Industria [1205]. Industrial [1333, 1378, 1206]. Industrialized [883]. Industries [1239]. Industry [1204, 1192, 597, 1947, 1205]. inermis [56]. Inert [2164]. Inexpensive [1456, 2022]. Infauna [69, 255]. Infected [2691, 397, 2453, 2719, 1578, 951, 641]. Infecting [1360, 1751]. Infection [1785, 1596, 590, 1688, 2306, 1974, 1490, 1907]. Infections [2503, 2311]. Infectious [34, 163, 1406, 1971, 1570, 2131, 428, 2960, 1843, 1653, 2361, 2729, 478, 2843, 2642, 2015, 1113, 1311, 1579, 1521, 2839]. Inferences [20, 1907]. Inferred [927]. Infestation [1323]. Infiltration [2196, 2641]. Inflammatory [24]. Influence [143, 1757, 40, 743, 2060, 909, 2363, 1270, 2936, 344, 2941, 385, 2018, 948, 114, 1425, 925, 2151, 2553, 2714, 2449, 1467, 1206, 1129, 2176, 2946, 2609, 2300, 151, 1722, 2619, 2959, 2067, 737, 2967, 1150, 1376]. Influenced [761, 1031, 2038]. Influences [2248, 2811, 1775, 2769, 249]. Influencing [351, 2641, 2422, 1422]. Information [2958, 2809, 1624, 2808, 1047]. infralittoral [1001]. Infrared [1735, 578]. Infusion [746]. Ingested [121]. Ingredient [1359]. Inhabiting [1413]. Inherent [1805]. Inheritance [923]. Inhibiting [2021]. Inhibition [1536, 1749, 1517, 2919, 783, 2125]. Inhomogeneities [1088]. Initial [296, 593, 896, 2063, 2223]. Initiation [1869]. Injected [1724, 1985, 564, 2455, 211]. Injection [1406, 2196, 2160, 1573]. Inland [2879, 1813, 1935, 626, 420, 2922, 2501]. Inlet [572, 2342, 1752, 1442, 2981, 500, 1709]. Inlets [841, 1378, 432]. Inoculated [596, 1279]. Inorganic [2842, 1092, 1441, 1086, 1763, 2576]. inornata [2720, 1312]. Inosine [436]. Input [2533]. Inputs [2600]. Inquiry [2564]. Insect [2206, 1807, 2806]. Insectan [2807]. Insecticide [1, 1322, 10, 1487, 2176, 268, 965]. Insecticides [613, 566]. Insects [2811, 2796, 2059, 2808, 1262, 2250]. Inserted [2223, 2488]. Inshore [2580, 1618, 744, 2886, 2080, 2834, 216, 1747, 1746, 2197, 1389, 1907]. insignis [1417]. Instability [1401, 1607]. Instar [2179, 2275]. Institute [626]. Institution [1703]. Institutional [1504]. Instrument [841, 1919, 2883]. Instrumentation [614]. Instruments [2490]. Intake [322, 2893]. integra [149]. Integral [1087]. Integrated [462]. Integrating [23]. Integrator [465, 463]. Integument [2189]. Intensified [2415]. Intensity [2683, 1552, 1861, 1138]. Intensive [1201, 2139]. Inter [2413]. Inter- [2413]. Interaction [1263, 1241, 574, 981, 1993, 1641, 532, 823, 654, 837]. Interactions [2606, 2914, 961, 1310, 2396, 2426, 2418, 1039, 2531, 314]. Interchangeable [956]. Intercomparison [2979]. Interdependent [1723]. Interest [1184]. Interference [1669]. Interferes [2917]. Interior [1976]. Interlake [2113]. Intermediate [204, 485, 934]. intermedius [1107]. Intermingling [1366]. Intermolt [1497]. Internal [136, 1374, 2907, 2098, 1919, 2145]. Internally [1996]. International [2433, 1229, 2374, 1482, 1810, 2500, 1847, 1230, 1179]. Interpretation [2028, 1663]. Interpreting [2211]. Interrelation [609, 1036]. Interrelations [2393, 2505]. Interrenal [1388, 137, 336]. Interspecific [1320, 2179, 961, 1006]. Interstitial [421, 496]. interstitielle [496]. Interstrain [2011]. Intertidal [1322, 1076, 2655, 2627, 421, 47, 2342, 1067, 1140, 729, 1816, 2630, 1708, 2004, 2597, 863, 2623, 524, 1073, 2855]. Intervals [1702, 1978]. Intestinal [902]. Intestine [617]. Intoxicated [2000]. Intragravel [141]. Intraperitoneal [2196]. Intraperitoneally [1466]. Intraspecific [335, 1034, 2413]. Intrinsic [1524]. Introduced [2942, 1478, 1090, 417, 201]. Introducing [1599]. Introduction [2374, 550, 2796, 1813, 1330, 1252, 671, 2749, 2868, 676, 2278, 2805, 2596]. Introductions [693, 2398, 702, 696, 687, 684, 692, 683, 686, 677, 2447, 694, 685]. Introductory [2829, 2778]. Intrusion [2789, 576]. Intrusions [2970]. Inuit [1610]. Invasion [2492, 1092]. Invertebrate [2050, 2911, 2805]. Invertebrates [1719, 386, 2625, 1972, 2035, 214, 1682, 264, 368, 1741, 1816, 2807, 626, 2560, 2004, 65]. Investigation [353, 1035, 1643, 338, 2474]. Investigations [29, 50, 68, 90, 115, 139, 158, 174, 188, 215, 241, 292, 309, 333, 360, 387, 406, 429, 455, 480, 542, 581, 600, 622, 643, 668, 724, 750, 775, 799, 824, 851, 879, 914, 936, 957, 979, 1026, 1110, 1635, 462, 1000, 1057, 1083]. Involved [301, 472]. iode [1079]. Iodine [189, 2024, 1971, 2803, 1079]. Iodine-Treated [1971]. Iodophors [570, 771]. Iodothyronine [1035]. Ion [2458, 2521, 1100, 2535]. Ion-Selective [2535]. Ionic [743, 1694]. Ionization [2667]. Ionized [798]. Ions [74, 70, 981]. Iowa [2411]. IPN [2960, 2666, 1669]. IPNV [2642]. Iranian [2610]. Ireland [1637]. Iron [1885, 1060, 1716, 1037, 637]. Irradiance [1310]. Irradiation [364, 906]. Irrigation [865, 767]. irroratus [2721, 2836, 791, 588]. Island [2312, 2940, 373, 1638, 230, 762, 1148, 649, 2454, 1338, 187, 850, 2026, 1422, 2014, 2600, 921, 2159, 467, 2100, 1447, 553, 871, 1303, 66, 1761, 2094, 1340, 1336, 840, 2506, 747, 2274]. islandica [1476]. Islands [2226]. Isochrysis [1312]. Isolated [512, 2334, 1044, 1413, 1602, 569, 2745, 2585, 2370]. Isolates [1998, 1909]. Isolation [242, 428, 2519, 1123, 1651, 1247, 1669, 1823, 2230, 1113]. isolée [2585]. Isopod [2616]. Isopoda [2720]. Isotope [2001, 2858]. Isotopic [2576]. Isozymes [272, 127, 414, 886, 1006]. issu [1253]. Items [892, 1610]. IV [350, 1207, 1213, 1220, 1227, 1233, 1238]. Ivlev [1416].

J [1873, 1868]. Japan [1936, 1946, 1222, 1949]. Japanese [1187, 2948, 1529, 1318, 1639]. Jar [2365]. Javex [1536]. Jaw [1156, 345]. Jeddore [2449]. Jensen [21]. John [1590, 2042]. Johnny [735]. Johnson [492, 486]. johnsonii [44]. Johnston [1141]. Johnstone [2091, 2274]. Joint [1216, 2915, 1810]. Jolicoeur [1678]. Jolly [325]. Jordan [1444]. Jordanella [913, 2066]. jordani [370, 517, 1761]. journal [1847, 2797]. Journey [612]. JPU [2259]. Juan [2935, 520, 2557]. Jurisdiction [2499]. Juvenile [1265, 1712, 2437, 1464, 2276, 1799, 1487, 1592, 1593, 1668, 1497, 1009, 1302, 1594, 574, 2515, 1667, 437, 526, 1700, 606, 2831, 926, 2266, 1046, 1609, 560, 1581, 1016, 937, 1807, 1651, 1859, 1804, 2064, 1032, 1054, 922, 1362, 1604, 1777, 2722, 2906, 2714, 1729, 2833, 1501, 500, 235, 1465, 1466, 2631, 2038, 2181, 950, 2832, 1008, 837, 2147, 2648, 1918, 2774, 1145, 354, 846, 990, 1273, 2139, 2634, 2184, 2866]. Juveniles [564, 2662, 2616].

Kachemak [71]. Kalamalka [1005]. kanagurta [906]. Kanamycin [1576]. Kananaskis [947]. Kansas [2029]. Karyotypes [1127, 658]. Karyotypic [2850]. Katsuwonus [184, 231]. Keeping [382, 180, 1796]. Keller [680]. Kelp [1983, 2342, 666, 1948]. Kemano [2540]. Kern [412]. keta [2643, 26, 1667, 2831, 1016, 1697, 1256, 2832, 1587, 42]. Ketones [2154]. Ketotestosterone [336]. Key [346, 2218, 2354, 2532]. Keynote [1168]. Keys [967, 1530, 1550, 2925, 2926]. Kidney [1769, 2921, 1496, 1101, 1300, 337, 2955, 2822, 2972, 2504, 2167, 904, 1521, 2696]. Kidneys [2859]. Kieffer [967, 1550]. Kill [1479, 1788, 2494]. Killer [1646]. Killifish [43, 2256, 450, 48]. Killifishes [16]. Kiln [142]. Kinetics [2151, 2208, 2558, 2894, 2986]. King [195, 2317, 86]. kisutch [596, 1799, 2163, 1914, 2198, 1487, 1592, 1593, 1906, 1904, 141, 2141, 2645, 1826, 1962, 2187, 2503, 1434, 769, 2018, 1688, 1609, 1732, 1860, 2726, 1049, 937, 2704, 144, 549, 1804, 1380, 2033, 922, 1362, 1604, 1777, 2906, 2723, 1974, 81, 315, 1465, 1466, 479, 2245, 1275, 2504, 2802, 2321, 2679, 975, 2043, 1963, 837, 2167, 2648, 1118, 1687, 267, 354, 554, 846, 904, 990, 2139, 1376]. Klicava [2386, 2405]. Knowledge [2536]. Known [1788, 404, 1324]. Kodiak [2940]. Kokanee [350, 550, 1478, 2749, 1615, 993]. Kootenay [691]. Korea [1223]. Korean [1236]. Krabbe [1510]. Kraft [1799, 2299, 2298, 1099, 919, 2294, 1609, 937, 1651, 2256, 922, 1362, 1604, 2906, 76, 898, 2304]. Kraftmill [805]. Krill [56, 1199, 2963]. Kuhlia [571]. Kuwait [2148]. kyphosus [219].

l [2155, 2018, 101]. l-acyl- [2155]. L-Alpha-Glycerophosphate [101]. L-Thyroxine [2018]. L. [2114, 2505, 2519, 2381]. Labeled [1733]. Labeling [236, 453]. Labelled [377]. Labidesthes [1803]. Laboratory [2683, 2237, 1111, 1284, 389, 2943, 2283, 1602, 2552, 630, 33, 2773, 2250, 451, 1387, 913, 2190, 2019]. Laboratory-Maintained [2773]. Labrador [1369, 96, 1431, 1999, 209, 447, 66, 1539]. Labrifer [376]. lac [649, 117, 2408, 692]. Lack [1278]. Lacs [2408]. Lactate [1912, 788, 1097, 127, 1839, 2000, 223, 2478, 177]. Lactic [452]. Lacustrine [1786]. Ladders [344]. Lag [742, 2699]. lageniformis [245]. Lagenorhynchus [2713, 2712]. Lago [684]. Lagodon [2256, 2476]. Lagoon [738]. Lagoonal [863]. Lagoons [1984]. Lake [1690, 2856, 1121, 2458, 1251, 2267, 2313, 1120, 1449, 1525, 2940, 2934, 1464, 220, 327, 117, 297, 1532, 1888, 1713, 1040, 2411, 1409, 127, 886, 987, 2929, 1119, 8, 808, 1469, 1684, 92, 1840, 1090, 1089, 136, 2141, 437, 1915, 2413, 649, 2263, 616, 1259, 379, 1351, 2397, 1554, 2657, 1091, 1284, 980, 53, 1045, 652, 2417, 70, 1403, 1911, 295, 1741, 1740, 2112, 2390, 2394, 2301, 1338, 712, 2074, 2396, 2764, 2412, 391, 2738, 17, 335, 378, 186, 739, 832, 2750, 1094, 2151, 1586, 1252, 122, 13, 369, 133, 193, 2659, 2, 2749, 36, 278, 1018, 1080, 2471]. Lake [2368, 364, 1498, 2049, 2046, 1025, 2637, 1275, 2273, 2956, 2638, 2300, 280, 2296, 51, 226, 1340, 1354, 1336, 2116, 2447, 1462, 1262, 2730, 1394, 2430, 1008, 2105, 2416, 1784, 2395, 2968, 2692, 2269, 1096, 1287, 46, 1356, 1853, 2015, 2367, 1750, 839, 2032, 1579, 2839, 1424, 146, 2674, 2219, 1454, 2401, 1479, 2521, 1803, 2942, 118, 693, 2398, 2472, 646, 1532, 1292, 1885, 1889, 1892, 1898, 1887, 2325, 2347, 1426, 2400, 696, 550, 1478, 1553, 2929, 2251, 1119, 608, 1271, 2434, 1293, 2748, 567, 1087, 2188, 2333, 1290, 687, 2862, 2103, 1139, 695, 2397, 748, 1911, 521]. Lake [523, 522, 680, 2419, 1338, 681, 1884, 17, 2718, 1891, 1890, 876, 1509, 683, 2750, 1897, 2492, 200, 2927, 2886, 13, 686, 75, 2689, 1894, 1895, 1718, 2191, 2751, 816, 2418, 2723, 691, 1886, 2152, 2252, 2253, 169, 2558, 2894, 1385, 1498, 1266, 1893, 650, 1913, 1339, 280, 2210, 2951, 530, 1092, 1085, 2410, 472, 1562, 2364, 1879, 1878, 975, 2105, 1389, 2067, 2395, 2968, 1881, 1880, 1282, 2920, 2331, 1287, 2200, 1896, 1297, 1337, 694, 2414, 2402, 1882, 1883, 685]. Lake-Dwelling [616]. Lakes [159, 31, 312, 1349, 461, 299, 2113, 1899, 2895, 603, 731, 674, 2824, 2387, 2734, 1331, 307, 297, 1995, 1613, 2391, 1296, 702, 296, 703, 1369, 983, 584, 2047, 2571, 1350, 2931, 806, 698, 697, 305, 678, 553, 1176, 1814, 2852, 52, 2112, 2393, 2390, 1335, 2040, 2676, 2675, 2889, 2888, 2438, 1969, 1852, 1767, 677, 2744, 1357, 303, 942, 2510, 701, 676, 1095, 699, 672, 301, 2556, 530, 308, 298, 300, 304, 1093, 1354, 2533, 2409, 166, 1551, 306, 302, 1355, 1823, 673, 2307, 2051, 2402, 1005, 1342, 1295, 1348, 1345, 1038, 1343, 67, 1291]. Lakes [1289, 1825, 671, 72, 358, 2644, 786, 1175, 1275, 2403, 2804, 2420, 1347, 679, 700, 2348, 1963, 2153, 171, 1344, 1288, 1294, 1346, 2576]. Lakewater [2757]. Lakewide [1894]. Lamellibranch [64]. Laminaria [1946, 380]. lamottei [358, 854]. Lampetra [358, 854, 950, 889]. Lamprey [2189, 207, 959, 2363, 876, 194, 854, 950, 1869, 889]. Lampreys [748, 1586, 358, 1266]. Lampricide [1147, 2698, 748, 1674, 2208, 1028]. Lampsilis [806]. LAN [1656]. Lance [1268, 827, 927, 902]. Lances [828]. Land [1738, 625]. Landings [998, 2226]. Landlocked [1, 339, 2363, 186, 194, 345]. Lanternfish [11, 148, 829]. Lanternfishes [88]. Lapse [784]. Large [2024, 527, 956, 1192, 1296, 1922, 102, 2078, 131, 1215, 155, 2332, 673, 146, 2969]. Large-Scale [1192, 1296, 2078, 1215]. Largemouth [1468, 639, 1409, 1302, 2265, 651, 2526, 1759, 2579, 2108]. Largescale [133]. Larus [2723]. Larva [1510]. Larvae [1454, 2178, 1734, 2976, 1409, 1012, 1731, 2639, 1770, 1966, 2048, 2237, 2551, 53, 1414, 2770, 2257, 2662, 630, 648, 1144, 827, 996, 2366, 1907, 2019, 1783].

Larval [639, 208, 185, 656, 1068, 1531, 1766, 1326, 1284, 2436, 2764, 1674, 403, 854, 319, 2227, 2204, 1794, 2866]. Larvicide [287, 2264]. Last [1380]. Lat [1340, 1336]. Lateral [416, 2380, 372, 1787]. Lateral-Line [2380]. lateralis [1731, 804, 803]. Latimeria [2867]. latipes [1944, 458]. latipinna [206]. Latitudinal [2732]. Launce [353, 237]. Laurent [2226, 1001, 1267, 1730]. laurentiae [2153]. Laurentian [1288, 1767, 1175, 679, 700, 2153, 1344]. laurettae [1154]. lavalaei [1323]. Lawrence [2743, 2437, 2226, 2562, 2840, 2444, 2436, 1730, 2938, 1903, 1982, 2690, 1779, 2145, 1346, 958, 2437, 2061, 2240, 1001, 1295, 1267, 1345, 2780, 663, 1834, 2123, 2436, 1693, 1993, 895, 415, 728, 1563, 1515, 786, 1830, 229, 1540, 2120, 1564, 2788, 1399, 998, 2001, 2858, 2041]. Layer [930, 1779]. Layers [1033, 782, 1903]. Laying [181]. LDH [1261, 1022]. LDH-B2033 [1261]. Leachate [2140]. Lead [2903, 2677, 1855, 2247, 2578, 2039, 948]. Lead-Exposed [2578]. Leaf [2541, 638, 905, 2130]. Leaping [2648]. Learn [613]. Learning [1591]. Leaving [747]. Lecithin [1695]. Leech [1260]. Leeches [2036]. Left [2962, 418]. Leiostomus [2476]. Léman [692]. Length [604, 2975, 740, 583, 2199, 642, 2123, 847, 2218, 1789, 380, 717, 2534, 1564, 1305, 2532, 2925, 2926, 649]. Lengths [2710, 1266, 2186]. leniusculus [1786]. Lens [1036]. Lepeophtheirus [313, 982, 1250]. Lepidion [54]. Lepocreadiidae [376]. Lepomis [1757, 1862, 657, 997, 2023, 1143, 356, 1003, 2505, 2519, 1254, 911, 1395, 1520, 888, 1459, 1385, 328, 2110, 770, 2220, 1728, 2161]. Leptocottus [1398, 577]. Lernaeocera [1907]. Lernaeoceridae [1444]. Lernaeopodidae [245, 94, 709, 1002, 2195]. Lesion [24]. Lesions [2681, 918, 286, 1750]. Leslie [1641]. Lesser [1391, 1638]. Lessons [2813]. Lethal [2785, 2201, 1806, 1046, 770, 1841, 343]. Lethality [2184]. Letters [2055, 2081, 2166, 2287, 2350]. Letters/Letters [2166]. Letters/Lettres [2055, 2081, 2287, 2350]. Lettres [2055, 2081, 2287, 2350]. leucas [1126, 467, 1030, 1629]. leucichthys [929, 2182, 1535]. Leucine [1656]. Leucocrit [2450]. leucopsarus [148]. Leukemia [289]. Leukocytolytic [2334]. leurs [2969]. Level [395, 1754, 2627, 1433, 1456, 2581, 1625, 1542, 2049]. Levels [189, 868, 958, 1795, 1914, 1405, 2434, 1839, 789, 1518, 2577, 82, 1860, 2861, 2270, 2704, 1020, 2244, 1718, 1854, 147, 441, 2186, 393, 798, 2487, 6, 1372]. Levulinic [1855, 2247]. lewisi [2850]. License [1187, 1185, 1186]. Life [2811, 2809, 1295, 1314, 2025, 1774, 988, 1434, 2483, 1737, 2977, 2234, 1002, 1960, 791, 378, 832, 2530, 2732, 2810, 2807, 1380, 1044, 2338, 2689, 1670, 1611, 1630, 2808, 1028, 2806, 2745, 2202, 2805, 411, 915, 778, 1562, 2111, 1286, 906, 2812, 2967, 2376, 1441]. Light [528, 1735, 468, 1068, 1531, 1138, 1443, 1552, 107, 23, 2379, 1861, 2527, 298, 78, 451, 754, 1389, 1728, 2161, 1816]. Light-Emitting [1735]. Lightweight [2798]. lignorum [2026]. Likelihood [661]. Limanda [2262, 254, 330, 1475]. limatula [885]. Limb [2027]. Lime [1037]. Limitation [1187, 1185, 1186, 2253, 2707]. Limitations [451]. Limited [2874, 2872, 2876, 1654, 2875]. Limits [2330]. Limnetic [1314, 1548, 36, 2953]. Limnocalanus [1793]. Limnocorral [1411]. Limnological [1801, 2390, 1335, 2315, 1355]. Limnology [31, 1813, 678, 553, 1330, 672, 1084, 1336, 1347, 673, 1332, 2564, 1356, 146]. Limnoria [2026]. lindbergi [152]. Line [416, 2380, 2583, 2183, 372, 1004, 1787]. linéaires [2226]. Linear [820, 1677, 924, 2226]. linearis [25]. Lines [2666]. ling [197]. Lingcod [2358]. Link [721]. Lions [1782, 840]. Lipase [416]. Lipid [2024, 629, 1964, 1568, 379, 1264, 2537, 760, 1698, 2176]. Lipid-Iodine [2024]. Lipids [56, 353, 1476, 1298, 1477, 1513, 1442, 1312]. Lipogenys [1098]. liquefaciens [598]. Liquid [2222, 300, 377, 977]. Liquids [1306]. Liquor [745]. List [1137, 1164, 1874]. Literature [2915]. Lithodes [86]. Littoral [789, 2333, 2191, 2603, 496]. littorale [496]. littorea [1140]. Littorina [1140]. Live [1644, 138, 1646, 1543, 2715, 1645]. Live-Bait [138]. Live-Capture [1646, 1645]. Liver [1298, 2830, 129, 2265, 651, 1695, 2155, 43, 2933, 1536, 1020, 1410, 1367, 337, 2537, 12, 1022, 177]. Livers [2129, 533]. Living [1169, 1925, 2257, 2787, 448]. Loaded [2652]. Loading [2267, 2793, 2675, 2470, 2744, 2210]. Loadings [1886]. Lobster [1066, 272, 160, 1405, 1668, 398, 2574, 185, 1596, 2021, 128, 664, 2836, 1068, 1531, 1766, 744, 2665, 1598, 1116, 666, 2044, 161, 835, 1500, 2876, 525, 2740, 1542, 1363, 1144, 427, 2020, 1108, 1796, 634, 641, 1490, 998, 1753, 2019, 3, 2226]. Lobsters [1724, 900, 1497, 2272, 585, 1161, 2456, 2893, 2064, 76, 161, 984, 992, 898, 1371, 1438, 1748, 1072, 33, 1796, 1488, 1918, 1753, 2774]. Local [868, 998]. Locality [1107]. Localization [1656, 1675, 1394]. Locally [2142]. Located [1023]. Location [2731, 2488]. Locations [945]. Loch [690]. Loci [2915]. Locomotor [1735, 764, 981, 125]. Locomotory [438]. Locus [2948]. Logic [1025]. Logistic [1433]. Loliginid [212]. Lomond [690]. Long [1298, 862, 860, 2580, 955, 2608, 659, 2598, 1447, 2454, 887, 2039, 2442, 2615, 2630, 13, 210, 366, 2143, 2338, 1641, 1222, 2626, 1637, 1427, 494, 1660, 2088, 2692, 2564, 619, 146, 1148, 439, 2269]. Long-Chain [1298, 862, 860]. Long-Distance [1427]. Long-Finned [2454, 1641, 1637]. Long-Term [2580, 955, 2608, 659, 2598, 887, 2039, 2442, 2615, 2630, 13, 210, 366, 2143, 2338, 1222, 2626, 1660, 2088, 619]. Long-Terra [146]. Longear [2519]. Longer [1689]. Longevity [2958]. longicornis [492]. Longnose [527, 616, 1249, 752, 947, 153]. longueur [649]. Loss [2830, 2359, 243, 1072, 2753, 2860]. Lota [196]. Lotic [552]. Louis [649]. Louver [779]. Low [1862, 1925, 435, 2627, 1571, 2099, 2459, 2151, 2744, 2758, 1542, 1917, 300, 1694, 2258, 2204]. Low-Carbonate [300]. Low-Frequency [2099]. Low-Level [2627]. Low-Pressure [2151]. Low-Virulence [1925]. Lower [1046, 848, 790, 587, 2637, 1982, 2690, 1288]. Lowering [2704]. Lucia [1391]. lucioperca [2387, 2399, 2382, 2388, 2385, 2403]. lucius [1446, 1711, 909, 2830, 1943, 2390, 60, 645, 802, 224, 470, 410, 2537, 1039, 2980, 2403, 996]. Lumpenidae [2048]. Lumpenus [1155]. Lumpfish [758, 2183, 1698, 1907]. lumpi [1117]. lumpus [758, 2183, 1698, 1907]. Lunar [1804]. Lycodapus [2768]. Lycodes [1323]. Lymnaea [2677]. Lymphocystis [2262, 538, 1385, 2106]. Lyrodiscus [1038]. Lysolecithin [546]. Lysosomal [416].

M [2627, 325]. M.S. [1259, 2003, 290, 1591, 97]. Macdonaldia [477]. Macellicephaloides [489]. Mackenzie [1126, 821, 1450]. mackenziensis [2202, 2203]. Mackerel [720, 2239, 1661, 66, 235, 1427, 906, 2539, 2475]. Macoma [1067]. Macrobenthic [939, 2932, 1981]. Macrobenthos [2813]. Macrobrachium [2910]. macrocephalus [399, 98]. macrocheilus [133]. macrochirus [1757, 1862, 657, 2023, 1143, 2505, 1254, 911, 1395, 888, 1385, 328, 2110, 770, 1728, 2161]. Macrocystis [1948]. Macrofauna [2628, 2626]. Macroinvertebrate [2809]. Macroinvertebrates [1345, 521, 522, 1674, 2297]. macrolepidotum [2176]. Macrophthalmia [971]. Macrophytes [2897, 8]. Macrozooplankton [146]. macrurus [1793]. Mactaquac [1078]. maculosus [320]. Madeleine [2226]. Madtom [1538, 1417]. maenas [857, 2156]. Magdalen [2226]. Magellan [2604]. magellanicus [2836, 64, 9, 396, 858, 1384, 2763, 1659, 2443]. Magelona [492]. Maggiore [684]. magister [444, 2636, 10, 2162, 2237, 2711, 1441, 2949]. magna [831, 1328, 2820, 781]. Magnesium [637]. Magnetic [966, 2223, 2488]. Magnification [368]. Magnitude [2830]. Main [2871, 2406]. Maine [939, 375, 208, 882, 656, 949, 2513, 935, 253, 232, 791, 2602, 1070, 877, 1720, 2802, 2209, 2573, 1831, 2169]. Mainland [2100]. Mainstream [276]. Maintain [1059]. Maintained [162, 2773]. Maintaining [873]. Maintenance [2823, 736, 2404, 715]. Major [2458, 610, 549, 662]. Make [1059]. Makers [1599]. Malaŵi [2200]. Malacoceros [493, 493]. Mälaren [2403]. Malate [414, 1006, 1299]. Maldanid [2977]. Male [2580, 631, 2767, 170, 2126, 1508, 2773, 1747, 1746, 2197, 833]. Males [631, 273, 1530]. Malformation [1052]. Mallotus [2435, 2437, 2661, 2580, 2436, 2518, 1747, 1746, 2197, 45, 418]. malma [461, 1315, 943, 654]. Malnutrition [2924]. Mammal [2194]. Mammalian [137]. Mammals [2650, 2899]. Man [691, 2449, 690]. Management [2964, 2965, 1203, 2826, 1229, 1241, 2880, 2194, 1240, 1239, 2540, 1173, 1234, 1181, 1208, 1559, 1232, 2173, 1867, 1560, 1237, 1221, 106, 1170, 2775, 1179, 2339, 1868, 2651, 1184, 1461, 2878, 1214, 2469, 2863, 1228, 2869, 1200, 676, 1178, 1225, 2536, 2547, 2668, 1226, 1504, 1217, 1818, 1949, 1218]. Managing [1169, 1167, 2890, 2868]. mandibularis [2768]. Manefish [21]. Manganese [1885, 1060]. Manifestations [289]. Manipulation [509]. Manitoba [242, 567, 752, 712, 2074, 2396, 402, 917, 2447, 830, 1823, 1356, 839]. Manitoulin [553]. Mankind [1167]. Mannitol [2917]. Mantle [1716]. Manual [1816]. Margaree [2972]. Margaret [2035, 2977, 2835, 411, 1665]. marginata [1626]. Margined [1417]. margoi [1260]. Mariculture [1954, 1947]. marina [2607, 1585, 199, 2255]. Marine [624, 2828, 777, 1401, 1792, 2606, 921, 475, 1316, 152, 2328, 132, 1755, 39, 531, 1725, 1260, 2701, 2650, 1614, 2194, 457, 774, 2008, 2899, 27, 512, 386, 1811, 2917, 121, 2550, 2776, 2625, 2193, 2799, 2752, 2754, 214, 1310, 1196, 1560, 2115, 1583, 2222, 1999, 1323, 2036, 2013, 2981, 2183, 130, 1124, 1505, 2630, 1703, 2033, 2602, 2710, 1706, 1386, 1701, 2739, 2822, 2618, 243, 2616, 2728, 2485, 880, 1917, 569, 1660, 2609, 1549, 2865, 2846, 2332, 2777, 2109, 2146, 2308, 1601, 2753, 932, 706, 2614, 2596, 1875, 1845, 2466, 2125, 2232, 903, 2465, 573, 464, 1989, 2370, 2795]. Marine [1847]. marinus [180, 2189, 756, 1071, 959, 2363, 876, 1586, 1319, 194, 854, 1266, 1869]. Marion [8, 808, 1351, 369, 2, 278]. maritime [1267, 1360, 1572]. marjorius [265]. Marketing [2241]. Markets [1193]. Marking [1649, 37, 639, 1301, 2004, 211, 619]. Marks [1649]. Maros [474]. Maros-Schulek [474]. Marsh [2625, 1519, 2615, 2619]. Masked [404]. Masking [1048]. Mass [2546, 2454]. Massachusetts [2611, 2625, 496, 268, 2597]. Masses [1709]. Mastigophora [758]. Mastitis [2712]. Matamek [1379, 2212, 1373, 1018, 1821, 1800, 1104, 51]. Material [1733, 61, 2752, 180, 2451, 625, 2753]. Materials [1617, 1386, 394, 2887]. Maternal [2312, 1423]. Mathematical [2540, 2174, 411]. Mathematics [2352]. Matière [1267]. Mating [2721, 181, 636]. Matter [2562, 2754, 1334, 2790, 2762, 1830, 2938, 792, 217, 2858, 171, 1665, 1267]. Maturation [2448, 256, 105, 258, 137, 1010, 1943, 2726, 739, 838, 79, 1515, 633, 2700, 1783]. Mature [1274, 1063]. Maturing [1508]. Maturity [422, 960, 2017, 631, 2650, 256, 244, 239, 905, 834, 1017, 181]. Max [2595]. Maximization [1952]. Maximum [1723, 322, 1433, 1789, 1383, 2781]. Maybe [2775]. meadi [1926, 2231]. Meal [1428, 888, 1698, 1558, 2043, 1827]. Meals [1268, 1276]. Mean [1844]. Means [1938, 819]. Measure [887, 2787]. Measured [907]. Measurement [1292, 864, 2097, 474, 23, 1095, 2908, 2887, 2535]. Measurements [2096, 1309, 2087, 2898, 2933, 1583, 95, 2190, 134]. Measures [1827, 1205]. Measuring [1308, 991, 77, 841, 847, 1987, 2244, 2450, 1381, 2905]. Meat [1542]. Mechanical [2603]. Mechanically [2561]. Mechanism [79, 2708, 2336]. Mechanisms [438, 1059, 338, 2041, 2589]. Medaka [1944, 458]. media [1253, 2720]. Mediated [1448]. Mediating [1320]. Medidas [1205]. Medium [1401, 1437, 842, 2755]. Medusa [2544]. Meet [1209]. Meeting [1818, 1623]. megalotis [2519]. Meganyctiphanes [56, 2963]. Megaptera [2682, 1567]. Meigs [2627]. Meiofauna [2255, 2190, 2832]. melaena [2454, 1641, 1637]. Melanin [532, 823, 822]. Melanogrammus [1414, 311]. melanopleura [246]. melanostomus [2662, 252]. melas [1538, 1913]. Melbourne [2654]. Membrane [1446, 2151]. MEMORIAM [2595]. Memphremagog [2434, 1718, 2191, 2751, 2333, 2558, 2894]. Menaquinone [2046]. Mending [1618]. mendotae [2582]. Menhaden [1581, 332, 1698, 1324]. Mensural [2344]. mentella [180]. Mercenaria [1716, 178, 319, 2012]. Merchant [1529]. Merchant-Ship [1529]. Mercuric [537, 647, 1061, 1742]. Mercury [2856, 2643, 1060, 793, 1716, 1053, 928, 1307, 818, 935, 1391, 711, 2016, 1714, 938, 1100, 1019, 845, 1082, 830, 1462, 994, 2246, 1441, 1610, 1657, 2474, 2716, 2043, 1881, 84, 964, 794, 903, 897, 89, 1742, 1743]. Meretta [1338, 1340]. Meristic [1590, 1379, 2344, 1552, 291, 223, 347, 2768, 828, 2518, 176]. Meristics [237]. Merluccius [2782, 2119, 2483, 245, 1673, 49, 850, 463]. Meromictic [2968]. Meromixis [92]. Mesangial [2118]. Mesidotea [2616]. Mesodinium [321, 2232]. Mesopelagic [877]. Metabolic [920, 2633, 1527, 858, 938, 1436, 2110, 2190]. Metabolism [2075, 831, 1865, 1725, 351, 2000, 523, 130, 838, 2144, 1367, 1604, 2570, 2956, 1659, 1062, 1569, 1511, 2122, 857, 1079]. Métabolisme [1079, 857]. Metabolites [772, 1823]. Metal [1328, 2734, 1782, 74, 2562, 43, 1715, 383, 2307]. Metal-NTA [1328]. Metals [2267, 831, 2159, 2546, 2921, 948, 2659, 1019, 1382, 1287, 2586, 1619]. Metamorphic [2435]. Metamorphosis [756, 194]. Metavermilia [484]. Metazoan [1856, 153, 213]. Metering [1306]. Methane [362, 767]. Methane-sulphonate [362]. Methemoglobinemia [2919]. Method [167, 733, 374, 1308, 2330, 2939, 57, 962, 1296, 991, 593, 916, 183, 2698, 1684, 2097, 1614, 2541, 1089, 2941, 2254, 815, 2561, 848, 1264, 261, 424, 1160, 2816, 1254, 1788, 1123, 426, 2682, 647, 2896, 1987, 2052, 1685, 516, 972, 300, 812, 377, 1021, 2124, 5, 2908, 2772]. méthodes [1267]. Methodology [1727, 1820, 1762]. Methods [1035, 1773, 907, 1170, 419, 1445, 2283, 2751, 1201, 1198, 2560, 2514, 973, 1760, 1575, 2900, 2901, 2902, 1267]. Methoxychlor [2264]. Methyl [64, 1061, 1742, 1743]. Methyladenosine [1149]. Methylation [2861]. Methylisoborneol [1823]. Methylmercuric [2143, 1451]. Methylmercury [928, 802, 1854, 2980, 464, 1989]. Methyltestosterone [336, 1032]. Metres [2684]. Metridia [1763]. Metula [2604]. Michigan [2219, 2472, 646, 2251, 1087, 2931, 2862, 748, 1045, 1911, 1714, 2492, 2927, 2438, 2689, 2252, 2253, 650, 975, 694, 2414]. michiganensis [2153]. Micro [37]. Micro-Tags [37]. Microalgae [1470]. Microalgal [907]. Microanalysis [2921]. Microbial [2606, 2608, 2705, 2604, 842]. Microbiological [1293]. Microcultures [1115]. Microflora [73, 617, 377, 2928]. Micromesistius [1603]. microns [394]. Micronutrient [2554]. Microprobe [1100]. Micropterus [1468, 1409, 2362, 1302, 2265, 651, 1046, 2393, 2390, 2526, 2185, 1759, 2579, 46, 2108]. Microscale [930]. Microscopic [2909]. microscopie [1253]. Microscopy [2681, 1557, 1443, 1253]. Microsomes [1695]. Microstomus [819]. Mid [2866]. Mid-Atlantic [2866]. Midcontinental [1023]. Midsummer [1551]. Midwater [1509, 257, 2662, 873]. Migrant [1822]. Migrating [350, 874, 1973, 1439, 1408]. Migration [2182, 945, 116, 1315, 71, 1283, 587, 785, 1518, 2513, 2907, 784, 748, 2369, 1427, 1422, 1626, 872, 1246]. Migrations [437, 1661, 1366, 1907, 2127, 2487]. Migratory [2736, 710, 611, 2185]. Mile [1810]. milieux [1253]. Milk [2312]. Milkfish [1938]. Mill [1799, 2299, 2298, 2303, 1099, 919, 2294, 2292, 2305, 2290, 2293, 922, 76, 898, 2291, 2302, 2296, 2304, 2295]. Mille [2408]. Miller [119]. Mimic [1421]. Minas [2855]. Minced [1607, 2119, 2483, 2234, 1673]. Mincing [1325]. Mine [2903, 52]. Mineral [861, 1413, 2548, 2122, 2969, 1253]. minérale [2969, 1253]. Minimal [1774]. Mink [752, 917]. Minke [1624, 1625]. Minnesota [2395, 2393, 2394, 2409, 2067]. Minnow [547, 2477, 2785, 435, 1613, 659, 2663, 2820, 726, 1037, 2366]. Minnows [1849, 1850, 1143, 2442, 911, 2128, 1582, 2579, 433]. minor [2903]. Minute [1793]. minutus [592, 1597]. Miramichi [1333, 1961, 316, 985]. Mirex [2704]. Missouri [2217, 60, 2406]. Mitchill [2381]. Mitochondria [93]. Mitochondrial [1006]. Mitten [1162]. Mixed [797, 2881, 1924, 2619, 2959, 2747]. Mixed-Function [2619, 2747]. Mixing [1411, 2970, 2575]. Mixture [1020, 662, 2586]. Mixtures [2785, 2896]. Mn [2801]. Mobilization [537, 711]. Möbius [758]. MOCNESS [2345]. Mode [2977, 1443, 2497]. Model [2330, 1658, 1114, 2947, 2733, 2100, 2445, 1469, 1965, 2540, 1087, 1433, 2174, 2575, 820, 391, 2424, 2422, 2339, 2553, 2144, 2208, 2317, 1854, 2756, 2295, 738, 411, 2210, 2951, 2260, 2430, 2474, 2539, 2707, 325, 1778]. Modèle [2226]. Modeling [2101, 2006]. Modelling [1088, 1705, 2986]. Models [1370, 2636, 1015, 1457, 80, 2044, 1395, 1641, 2819, 2744, 2210, 2556, 2589, 451, 1007, 1760, 2226]. Modem [1618]. Modes [2070]. Modification [2165, 2767, 2744]. Modified [2649, 1326, 874, 739, 2582, 1641, 33, 348, 155, 1796, 2755]. Modiolus [343]. Moist [2769, 1480]. Moisture [1568]. Mola [313]. Molecular [848, 549, 2758, 156, 1989]. mollis [723]. Mollusca [1731, 883]. Molluscan [2921, 627]. Molluscs [1817, 1076, 2525, 660]. Mollusks [1957, 1599]. Molt [1066, 2021, 2916, 2064, 2027]. Molt- [2064]. Molt-Inhibiting [2021]. Molt-Related [2916]. Molting [746, 2910, 2616, 2120, 1108]. Molts [185]. Molybdenum [993]. Monitoring [2567, 1919, 445, 2747]. Monocelididae [2695]. monoceros [1628]. Monodiamesa [967]. Monodon [1628]. Monoethylenic [1298]. Monofilament [2929]. Monogenea [1141, 153, 2053]. Monogenoidea [1038, 1271, 1486]. Monona [1385]. Monophosphate [436, 858]. Month [2618]. Months [2197]. Montreal [1623]. Montsweag [949]. Mooneye [1555, 2077, 1281]. Moonlight [479]. mordax [868, 2344, 162, 200, 816, 609, 2730, 2348]. morhua [1997, 2976, 2517, 759, 2453, 1749, 1693, 1775, 1325, 2186, 1366, 1907, 2491, 2340]. Morid [44, 54, 2482]. Moringua [25]. Morne [2745]. Morone [2914, 2325, 544, 618, 2731, 1729, 2766]. Morphoedaphic [2471, 2441, 1341]. Morpholine [1904, 1435, 2930]. Morpholine-Imprinted [1904]. Morphological [2189, 2934, 806, 17, 2382, 2889, 735, 196, 828, 833, 286]. Morphology [150, 163, 756, 2942, 2178, 1914, 1269, 732, 612, 651, 1300, 1675, 2516, 2381, 99, 1900]. Morphometric [2518]. Morphometries [605, 237]. Mortalities [731, 136, 2162]. Mortality [1899, 2677, 1636, 2445, 2363, 1270, 1433, 2937, 1696, 28, 1067, 2266, 1857, 112, 1788, 2534, 1380, 2033, 816, 951, 2010, 2946, 973, 721, 2230, 433, 1794, 2949, 2223, 2697, 2983, 2984]. Moscow [2384]. mossambica [246, 474, 2220]. Mosshead [875]. Motion [1102]. Moult [585]. Mount [870]. Mountain [159, 31, 731, 2104, 1135, 264, 739, 288, 2762]. Mountains [1349]. Mouth [1443]. Movement [1265, 2028, 141, 1662, 2817, 70, 1425, 2074, 1825, 2305, 1804, 2152, 2910]. Movements [722, 1594, 2509, 2104, 2513, 320, 1526, 1094, 402, 710, 36, 1761, 1588, 1106, 1036, 1063]. Moxostoma [2176]. Mozambique [488]. Mp [2948]. Mp-1 [2948]. Mud [1322, 900, 1455, 797, 421, 1572]. Mud-Dwelling [1455]. Muddy [773, 1048, 1837, 2885, 1823, 1484]. Muddy-Flavored [1484]. Mudflats [1019]. Muds [70]. Mueller [1486]. Mugil [1448, 2476, 2230]. Muir [1709]. Mulinia [1731]. Müller [977]. Mullet [1448, 2476, 2230]. Multichannel [1573]. Multifilament [2929]. Multilingual [2794]. multilingue [2794]. Multiple [127, 414, 1173, 1962, 1456, 205, 1820, 2345, 2370]. Multipurpose [2881]. Multispecies [1993, 1175]. Multivariate [318, 2344, 806]. Mummichog [413]. Mummichogs [2719]. Municipal [1914]. murmanensis [2453]. Murphy [87]. Muscle [2154, 1316, 2948, 270, 1043, 881, 864, 1375, 1263, 127, 414, 886, 1060, 436, 129, 713, 1053, 238, 2018, 64, 9, 396, 858, 1384, 82, 290, 224, 470, 2537, 730, 1325, 1383, 1691, 1659, 1759, 1462, 99, 448, 573]. Muscles [74, 912, 1103]. Mussel [2937, 2892, 343]. Mutagenesis [1704]. Mya [179, 2148, 2612, 2613]. Myctophidae [88]. Myctophids [732]. Myctophum [829]. Myofibrillar [976, 1694]. Myogen [735, 99]. Myoxocephalus [2587, 244, 227, 1589, 621, 2150, 1303]. Mysidacea [1132]. Mysis [204, 1295, 1132, 2862, 832, 2749, 2689]. mysticetus [551]. Mytilus [2937, 2892]. Myxosoma [2691, 1123, 1445, 1247, 2635, 1421, 2772]. Myxosporida [1697].

N [64, 510, 1486, 2360, 492, 491, 496, 2031, 486, 2481, 1340, 1336]. N-Methyl [64]. N. [2153]. N.I.O. [2345]. N.S. [768]. N.W.T. [1126, 762, 1610]. namaycush [954, 1280, 1423, 118, 1283, 327, 1409, 1259, 1848, 2657, 923, 1403, 2112, 2390, 2738, 335, 1586, 193, 2368, 2049, 2046, 1275, 280, 2105, 5, 1287, 2032]. Names [2495]. Nanaimo [2831, 2833, 2832]. Nanticoke [2644]. Naphthalene [2671, 2848]. Nares [2094]. Narragansett [1364, 2600]. Narrows [2528]. Narwhal [1628]. Nash [864]. Nasser [1758]. National [31, 1230, 1335, 2745, 1349]. Nations [2877]. Native [1921]. Natura [1863]. Natural [733, 1733, 567, 2647, 2011, 1433, 385, 165, 1837, 1338, 1788, 1919, 1380, 2388, 2708, 1112, 1542, 95, 1460, 2116, 2451, 2716, 281, 1601, 42, 1794, 2015]. Naturally [2861, 109, 1807, 1578, 2485]. Naturans [1863]. Nature [1997, 1733, 2574, 390]. Navigation [2593]. Near [2903, 2087, 1519, 735, 1654, 168, 1349, 2267, 1666, 883, 1730, 1470, 577, 2384]. Near-Bottom [2087]. Near-Shore [168]. Nearctic [2360, 2177, 967, 1530, 1550]. Nearshore [1666, 1825, 2644, 1549]. Nebraska [539]. nebulosus [955, 1115, 2229, 1425, 1959]. Necessary [594]. Nécrologie [2845, 2944, 2982]. Necrosis [34, 150, 163, 1406, 1997, 1971, 1570, 2131, 2647, 428, 2960, 1843, 1653, 2912, 2361, 2729, 2524, 2209, 478, 2843, 2340, 2642, 2015, 1113, 1311, 1579, 1521, 2839]. Need [2669, 1189]. Needs [2327, 2054, 2805, 2812]. Negative [2349, 1650]. Nematoda [2178, 514, 1510, 1277]. Nematode [2736, 926]. Nematodes [107]. Neoamphitrite [490]. Neoplasm [289]. Nephroblastoma [944]. Nephrocalcinosis [2859]. Nephthys [2255]. Nephtys [885]. Nereis [1458]. nerka [350, 395, 528, 341, 529, 853, 2940, 1464, 2276, 197, 513, 322, 1142, 920, 1791, 1754, 1865, 318, 587, 550, 1478, 110, 919, 468, 256, 1388, 1571, 258, 137, 1518, 225, 1822, 1350, 2792, 1843, 784, 1526, 100, 769, 2195, 2664, 499, 2750, 710, 337, 1508, 611, 1245, 2749, 2133, 1615, 1663, 2638, 2504, 1041, 2364, 1008, 1439, 1282, 2920, 392, 42, 993, 805, 940, 2015, 336]. Nest [997, 263]. Net [2929, 874, 1139, 440, 1642, 2834, 1294]. Netherlands [2402]. Nets [2770, 1133, 2257, 2345]. Netting [1617]. Neuromuscular [1759]. Neutralized [1099, 1037]. Neutralizing [1048, 397]. Newfoundland [2024, 2017, 2693, 2511, 244, 227, 585, 1161, 27, 240, 96, 439, 1999, 1335, 2036, 209, 1776, 1973, 1641, 1624, 1661, 235, 347, 2372, 2834, 254, 330, 1058, 2745, 1366, 1907, 2443, 2892, 213, 237, 999, 2487]. Newly [900]. Newly-Settled [900]. Next [1332]. Niche [2427]. Nicotinic [64]. Night [983]. Nights [479]. nilotica [143]. Nine [833]. Ninespine [1040]. Nipigon [2296, 2301, 2305, 2297]. Nipissing [2401]. Niskin [2979]. Nitrate [2705, 1517, 530, 1085, 2772]. Nitrification [1745, 1744, 1851]. Nitrilotriacetate [1328, 1270, 911, 738, 1023]. Nitrilotriacetic [420]. Nitrite [2705, 2245, 1452, 2919, 2634]. Nitrite-Induced [2919]. Nitrogen [1077, 1791, 2546, 1995, 2857, 384, 401, 2233, 2954, 2676, 2675, 1078, 1501, 328, 1145, 2674, 857]. Nitrophenol [1147, 1732, 1674, 2208]. Nitrosalicylanilide [1569]. No [2529, 2625, 2797, 1317, 2597]. Nobska [2611]. Nomenclature [813, 2495]. Non [1662, 2807]. Non-Insectan [2807]. Non-Trout [1662]. Nonadaptive [533]. Noncatch [2010]. Nonconventional [1199, 1191]. Nonfusion [424]. Noninflammable [1152]. Nonlinear [80]. Nonnative [1921]. Nonparasitic [889]. Nonsalmonids [266]. Nonselective [1597]. Nonspawning [2661]. Nontoxic [1152]. Nontreated [1971]. Nordmann [1141]. Normal [1575]. North [1455, 2517, 423, 1024, 14, 400, 1250, 265, 1642, 1034, 2768, 412, 2777, 55, 54, 1098, 1134, 2482, 1113, 56, 1477, 2587, 2591, 2592, 2178, 1613, 772, 2391, 1229, 2479, 1260, 436, 534, 2551, 2796, 2242, 1374, 385, 1221, 2149, 2150, 1228, 514, 257, 1856, 1969, 1162, 591, 1361, 2957, 1539, 1680, 699, 672, 792, 1629, 1036, 2086, 55, 54, 932, 1305, 897, 6]. Northeast [2884, 1636, 1493, 1179, 1775, 1708, 1017, 723, 252, 198, 893, 1783]. Northeastern [945, 782, 2393, 2114]. Northern [2159, 1446, 1711, 909, 1514, 2830, 2048, 1519, 178, 1943, 2390, 60, 645, 1788, 802, 224, 470, 410, 2537, 1627, 1039, 609, 66, 1324, 2980, 2176, 2403, 827, 927, 902, 996, 1313, 706, 2596, 511, 999, 2636, 175, 602, 2408, 401, 1643, 607, 1248, 52, 1357, 2372]. Northumberland [1966, 332, 588, 1144]. Northwest [1230, 821, 1741, 1740, 2112, 1338, 1413, 2068, 2202, 2203, 777, 2983, 2984, 1635, 1333, 1559, 844, 1232, 274, 212, 254, 828, 1574, 855, 217, 176, 1146, 1117, 2697, 829]. Northwestern [299, 307, 297, 296, 305, 1852, 303, 301, 530, 308, 298, 300, 304, 1093, 306, 302, 11, 2436, 295, 2339, 1095, 830, 2788, 765, 766]. norvegica [56, 2963]. Norway [2985, 1643, 2568]. Norwegian [1635]. Notacanthid [477]. notatus [969]. Note [1979, 1411, 1678, 1970, 2569]. Notemigonus [2904, 2366]. Notes [946, 2025, 616, 2525, 126, 2720, 1634, 2029, 934, 2359, 1146]. Notholca [2153]. Notropis [2362, 2221, 2062, 1074, 1923, 1039, 2366]. Noturus [1538, 1417, 19]. nouvel [496]. nov [1250, 514, 219, 2230]. Nova [2312, 373, 2454, 2612, 2035, 2977, 2835, 2599, 1335, 1627, 2449, 411, 915, 2605, 863, 2623, 1665, 353, 720, 1476, 868, 2936, 1832, 2970, 1698, 229, 1155, 648, 2892, 464]. novaeangliae [2682, 1567]. NTA [1328, 911, 1023]. Nuclear [425, 1042]. Nuclei [12]. Nucleotidase [225]. Nucleotide [1384]. Nucleotides [9]. Number [2587, 1912, 642, 2123, 820, 1457, 1789, 2150]. Numerical [2100, 1088, 1087, 2101, 1890, 2756, 2006, 1382]. Nursery [678, 1481]. Nutrient [2458, 1889, 1887, 1951, 593, 1469, 2319, 2333, 385, 2444, 232, 1268, 2652, 1886, 226, 2788, 1982, 2506, 1829, 1838]. Nutrient-Loaded [2652]. Nutrient/Production [2333]. Nutrients [2780, 1343, 1291, 2557, 91, 1895, 2079, 625, 2145]. Nutrition [513, 1405, 2500]. Nutritional [1258, 2206, 2893, 904]. Nutritive [1276]. Nyctiphanes [1964]. Nylon [2929].

O [2851, 2956]. O-Antigen [2851]. O. [2940, 2503, 1688, 1016, 144, 2504, 975, 837, 392, 42]. Obituary [1819, 2845, 2944, 2982, 233, 259, 352, 555, 575, 811, 867, 1069, 1257, 1390, 1453, 1683, 2565]. Obituary/Nécrologie [2845, 2944, 2982]. objective [910]. Objectives [1169, 1209, 1198]. Obligate [2928]. obliquus [781]. Observation [1972, 122]. Observations [1690, 2178, 373, 884, 2627, 587, 1529, 2632, 1766, 784, 1493, 2909, 85, 329, 849, 1034, 1670, 2152, 2372, 2520, 648, 1008, 321, 2473, 1408]. Observe [772]. Observed [785, 1020, 440]. Obstacle [1787]. Obtain [2822]. Obtained [2211, 532]. Obtaining [448]. Occa [113]. Occella [113, 126, 577]. occidentale [1210]. occidentalis [369]. Occupied [917]. Occurrence [1126, 475, 1803, 1638, 1553, 2647, 821, 599, 949, 64, 1634, 170, 660, 551, 1472, 947, 816, 235, 1398, 577, 2026, 2209, 1909, 2913, 1134, 1311, 104]. Occurrences [306]. Occurring [1807, 1661, 2485]. Ocean [1476, 313, 1934, 2009, 2369, 1319, 1199, 2006, 1963, 2891, 2707, 2884, 2958, 1636, 614, 1529, 2048, 323, 714, 1248, 1493, 101, 2114, 1214, 1928, 1215, 563, 2088, 723, 217, 2086, 252, 198, 893, 1305, 1783, 484]. Oceanic [2095, 1703, 760, 2768]. oceanicus [3]. Oceanographic [2249, 432, 2088]. Oceanography [2085, 2778, 2007, 2005, 1703, 1771, 2552, 2549, 1831]. Oceans [1167, 1197]. ocellata [261]. ochraceous [1149]. October [2216, 2865]. octogrammus [404]. Octopine [396]. Octopus [2737]. Ocular [2452]. Odd [868, 610, 2449]. Odd-Chain [868, 2449]. Odd-Year [610].

Odobenus [2680]. Odontocete [1034]. Odontoceti [1391]. Odor [720, 773, 992, 898, 2885, 1823]. Oegopsida [534]. Off [2936, 2890, 2741, 732, 2517, 740, 323, 1248, 488, 545, 2628, 69, 255, 1671, 1672, 2662, 1761, 2372, 229, 21, 754, 755, 477, 576, 484]. Offered [2456]. Offshore [2661, 1293, 850, 2253, 1389, 515, 1907]. Ohio [2206]. Oil [2592, 2606, 2627, 2611, 471, 2598, 1612, 2148, 2071, 557, 2607, 2610, 2625, 1140, 1027, 2599, 2615, 1020, 2149, 2601, 2630, 2617, 1729, 2626, 533, 2603, 1542, 1917, 2609, 2068, 2203, 2624, 768, 2597, 863, 1768, 2019]. Oil-Impacted [2611]. Oiled [2621, 2620, 2612, 2617, 2623, 2596]. Oils [189, 353, 624, 1568, 1111]. Okanagan [1005, 942, 1355]. Old [2186]. Olfaction [1447]. Olfactory [1906, 59, 769, 899, 1435, 2016, 2930, 161, 1371, 349, 1050]. Oligochaeta [220, 2495]. Oligochaetes [221, 1320, 961, 314]. Oligochaetologica [2495]. Oligocottus [320]. Oligotrophic [702, 703, 1335, 671, 701, 2638, 699, 672, 679, 673]. omiscomaycus [1271, 1045]. On-Site [1797]. Oncorhynchus [350, 2512, 596, 733, 1407, 2142, 501, 2942, 395, 528, 341, 529, 853, 2940, 862, 861, 860, 1464, 2276, 197, 513, 322, 1142, 920, 1791, 1754, 1865, 1799, 2643, 2163, 1914, 2198, 1487, 1592, 1593, 318, 587, 550, 1478, 1553, 1906, 1904, 1156, 110, 919, 801, 141, 26, 468, 256, 564, 1388, 1571, 2141, 2952, 258, 137, 2645, 1667, 1518, 225, 1826, 1962, 1822, 2187, 1350, 2792, 2907, 1843, 2503, 784, 1526, 100, 1434, 769, 2831, 2018, 1688, 1609, 1732, 102, 560, 1860, 2726, 2195, 1016, 2664, 2718, 499, 2750, 1049, 937, 1776, 2704, 144, 2126, 710, 1697, 549, 1256]. Oncorhynchus [1804, 1380, 2033, 337, 1508, 611, 922, 1362, 1604, 1777, 2906, 1245, 2749, 2133, 2723, 1974, 81, 315, 1465, 1466, 479, 2245, 1275, 1615, 1663, 2638, 950, 2504, 2802, 2321, 2679, 1041, 2832, 2364, 975, 1008, 2043, 1063, 1963, 837, 2563, 1439, 2167, 2648, 1587, 1768, 1282, 2920, 1118, 392, 1687, 267, 42, 365, 993, 805, 940, 354, 554, 846, 904, 990, 1273, 2139, 2015, 1376, 2311, 336]. One [1298, 1021, 1312]. One-way [1021]. Oneida [1915, 2413, 53]. Onisimus [1809]. onitis [925]. Onset [1986]. Ontario [1343, 1291, 1289, 2392, 1454, 2401, 221, 299, 2856, 731, 2267, 220, 307, 297, 1292, 572, 2347, 696, 296, 653, 608, 2632, 92, 2408, 983, 584, 2571, 305, 1139, 652, 2417, 2852, 52, 521, 523, 522, 295, 2390, 2519, 876, 2651, 124, 13, 1852, 2723, 2368, 303, 1266, 1095, 301, 2210, 2951, 2556, 226, 530, 308, 298, 300, 304, 1093, 830, 1462, 478, 166, 1389, 306, 302, 171, 1294, 314, 1297, 2051]. Ontogenesis [2122, 2175]. Ontogenetic [1962]. Oocyte [2448, 105, 1943, 2726]. Oocytes [633]. Open [1723, 1156, 568, 2678, 2756]. Open-Access [1723]. Open-Jaw [1156]. Opening [2345]. Opeongo [13, 686]. Operated [620, 27, 2323, 2158]. Operations [1698]. Opercular [183, 1588, 1768]. Operculum [972]. Ophichthidaeo [1900]. Ophiodon [2358]. opilio [629, 1596, 2695, 2123, 1605, 2027, 2892, 181, 636]. Opinicon [1454]. Opossum [2749]. Opportunities [1240, 2564]. Optima [2927]. Optimal [953, 733, 1114, 2813, 1820, 1916, 1717, 1511]. Optimization [106, 1481]. Optimum [941, 2534, 2049, 1818]. Oral [40, 2645, 1743]. Orbiniid [496]. Orbiniidé [496]. orca [1646]. Orcas [187]. Orcinus [1646]. Orconectes [2931, 2438, 2343, 2439, 2111]. Order [510]. Oregon [732, 753, 545, 1671, 2769, 754, 755, 477, 837, 889]. oregonensis [2, 1313]. Organ [709, 1827]. Organic [2724, 2298, 1292, 2923, 2595, 2294, 1334, 2301, 1988, 2762, 1781, 1830, 2938, 1086, 2858, 1726, 1665, 2801, 2674, 1267]. organique [1267]. organisations [1211]. Organisms [2828, 2621, 2646, 1968, 1735, 2761, 457, 774, 2008, 2917, 2610, 1999, 2945, 1706, 1701, 2140, 2323, 2370, 2865, 2072]. Organization [145]. Organochlorine [624, 958, 2312, 1029, 2723]. Organophosphate [1487, 965]. Organs [446, 2380, 428]. Oriental [1212]. Orientation [568, 992, 2213, 2280]. Orienting [981]. Origin [2042, 2752, 2981, 1245, 2359]. Örsted [495]. Orthophosphate [2309]. Oryzias [1944, 458]. Oscillations [2099]. Osmerus [868, 2344, 162, 200, 816, 2730, 2348]. Osmolality [743]. Osmoregulation [128, 1132]. Osmotic [343]. Osoyoos [1005]. Osteichthyes [447]. Ostrea [1312]. Ostreidae [2948]. Other [1298, 29, 50, 68, 90, 115, 139, 158, 174, 188, 215, 241, 292, 309, 333, 360, 387, 406, 429, 455, 480, 542, 581, 600, 622, 643, 668, 724, 750, 775, 799, 824, 851, 879, 914, 936, 957, 979, 1026, 1110, 1322, 733, 495, 510, 1428, 2525, 381, 70, 1957, 1199, 224, 2915, 1987, 199, 1019, 1059, 1072, 55, 176, 2340, 2772, 1000, 1057, 1083, 457, 948, 324, 300]. Otolith [1533]. Otoliths [2435, 2782, 638, 2816, 1415, 728, 2220]. Ottawa [883, 735, 1854]. Otter [1801, 1973]. Ouabain [1826]. Ouabain-Sensitive [1826]. Our [1167]. Outbreak [2131, 968, 952]. Outer [473, 840]. Outgrowth [431]. Outlet [1562]. Outplant [2561]. Output [941]. Ova [564, 1945]. ovaire [857]. Ovarian [16, 2767, 1474]. Ovaries [418]. Ovary [1749, 804, 803, 857]. Over-Wintering [250]. Overdominance [2948]. Overhead [2727, 1861]. Overhead-Light [1861]. Overlying [1134]. Overseas [1815]. Overview [2621, 2277, 2608, 2630, 1886, 2278, 1084]. Overwintering [2487]. Ovulation [2448, 1944, 1943, 2726]. Oxaloacetic [206]. Oxidants [2898]. Oxidase [423, 1024, 2747]. Oxidation [93, 788, 1097, 2755]. Oxidation/Fermentation [2755]. Oxidative [1525]. Oxide [896]. Oxygen [761, 399, 2895, 1071, 1791, 435, 1888, 1887, 123, 1409, 544, 866, 2076, 1774, 540, 1429, 2148, 2754, 1686, 907, 891, 2460, 319, 571, 2708, 466, 298, 445, 996, 2001, 804, 1159]. Oxygenase [2959]. Oxygenases [2619]. Oxygenating [138]. Oxygenation [817]. Oxytetracycline [1301]. Oyster [2823, 1790, 2948, 2574, 1012, 154, 1602, 78, 1458]. Oysters [1274, 1258, 1957, 1318, 2227, 1382, 393, 1312, 637]. Ozobranchus [1260]. Ozonated [2236, 2642]. Ozonation [2705]. Ozone [2467, 2847]. Ozonized [2898].

P [614, 1529, 2088, 2075, 2180, 2193, 1451, 2619, 2075, 2180, 2193, 2799, 2180, 2799]. P-450 [2619]. p’-DDE [2180]. p’-DDT [2075, 2180, 2193, 2799, 2180]. P. [2381]. pêche [1211, 2794]. pêches [1210]. Pachycerianthus [519].

Pacifastacus [1786]. Pacific [2072, 2884, 1493, 2114, 563, 723, 2086, 252, 198, 893, 1305, 1783, 399, 503, 1985, 2196, 945, 2782, 2958, 501, 1274, 2371, 2584, 2791, 862, 861, 860, 2655, 1790, 1229, 1099, 782, 2639, 1052, 2680, 512, 342, 2647, 534, 282, 2242, 385, 323, 714, 1248, 14, 58, 769, 1378, 473, 400, 2045, 102, 560, 101, 729, 245, 1444, 1250, 1221, 154, 98, 2775, 1776, 1708, 1032, 538, 108, 265, 49, 257, 1365, 850, 1398, 2873, 1631, 1762, 2728, 486, 577, 441, 1059, 2010, 950, 1680, 1226, 515, 463, 1055, 1246, 2707, 198, 893, 1305, 249]. pacifica [1763, 889]. pacificus [819]. Packaged [1286]. Packed [2046]. Pagophilus [958, 1029, 2339, 2819, 855, 872, 2697]. Pagurus [502]. Palaearctic [2360, 967, 1530, 1550]. Paleogeographical [2423]. pallasi [503, 2584, 2791, 2639, 2045, 560, 950, 515]. Pallette [2141]. palliata [1157]. palustris [2677]. Panacea [2924]. Pancreas [2018, 337]. Pancreatic [1971, 1570, 2131, 428, 2960, 1653, 2729, 478, 2843, 2642, 1311, 1579, 1521, 2839]. Pandalidae [283]. Pandalopsis [642]. Pandalus [71, 2455, 630]. Panel [2284]. Panels [494]. Paper [2299, 2298, 2294, 2292, 2305, 2290, 2293, 2291, 2302, 2295]. Paracladopelma [2360]. Paradigm [1870, 1863]. parahaemolyticus [814, 1572]. Paralepididae [447]. Paralepis [447]. Paralithodes [2317]. Paralytic [2763, 532, 823, 822, 2572]. Parameter [20]. Parameters [2330, 593, 178, 1625, 634, 1760, 1717, 2779]. Parasite [1464, 2276, 1478, 1260, 1444, 1250, 2195, 120]. Parasites [376, 2113, 608, 35, 96, 1373, 1360, 1751, 1961, 1443, 485, 934, 1523, 1856, 902, 1574, 1564, 153, 213, 1277]. Parasitic [2712, 2525, 844, 926, 2115, 1323, 107, 219]. Parasiticides [1745]. Parasitism [86, 950]. Parent [1172]. Parental [954, 249]. Parents [804]. Park [31, 1335, 2745, 2556]. Parks [1349]. Parophrys [1712, 1858, 2617, 2121]. Parr [2515, 1826, 613, 566, 114, 1722, 2961, 218, 589]. Part [2922, 2795]. Partial [1428, 261]. Participants [1874]. Particle [2164]. Particles [2293, 2888, 2152, 2485]. Particular [2405, 30]. Particularly [1460]. Particulate [1292, 401, 2754, 895, 2149, 2790, 2762, 792, 217, 625, 2858, 171, 2367]. Partitioning [2224]. Parts [956]. Pass [1278]. Passage [452, 1078, 1652]. Passamaquoddy [2207, 2232]. Paternal [2142]. Pathogen [1596, 512, 1013, 1699]. Pathogenesis [2843]. Pathogenic [2349, 2792, 2321, 2573]. Pathogens [2705, 771, 1650, 783, 2755, 2913, 2236]. Pathological [2617, 2618, 615, 1491]. Pathology [1360, 1929, 904]. Pathophysiology [1406]. Pathway [1695, 409]. Pathways [1061, 2367]. Patinopecten [473, 239, 355]. Pattern [2165, 1791, 1834, 2295]. Patterns [159, 1887, 1735, 323, 1034, 4, 278, 1422, 2185, 2108, 1981]. Pb [2801]. PCB [1755, 935, 1391, 1434, 1992, 1854, 975]. PCBs [2704]. Peak [1059]. Peak-Period [1059]. Pectinaria [243]. Pectoral [740, 324, 1588, 1759]. Peel [2228]. Pelagic [497, 1064, 1368, 809, 2662, 1222, 147, 2480, 1511, 2538, 903]. pelamis [184, 231]. Pelecypod [1957]. Pelecypoda [806, 1067]. Pelecypods [1103]. Pellet [2110]. Pelleted [1480, 896]. Pellets [2769].

Pelvic [324, 2359]. Pen [1985, 2043, 1997, 2524, 2209]. Pen-Reared [1985, 2043]. Penaeid [1937]. Penaeus [1412]. Penetration [1779]. Peninsula [213]. Pennsylvania [109]. Penobscot [2513, 939]. Pens [2891]. Pentachlorophenate [1556, 940]. Pentachlorophenol [1849]. Pepsin [888]. Peptone [1437]. Perca [2378, 2407, 1079, 2391, 2060, 2141, 2415, 2310, 649, 2786, 2405, 2394, 2419, 1825, 2424, 2399, 1054, 1718, 2751, 2406, 1767, 1854, 1039, 1463, 2386, 2385, 2403, 2409, 2384, 120, 2381, 2376, 2414, 839]. Percent [2024, 798]. Perception [468, 125]. Perch [2407, 2958, 2325, 2391, 2060, 1271, 2141, 2415, 2310, 323, 714, 1248, 2731, 2786, 2405, 1045, 101, 2394, 2419, 1825, 2424, 2399, 1319, 1054, 1718, 2751, 2406, 1767, 119, 1854, 1039, 1463, 2386, 2385, 2403, 2409, 2384, 120, 1277, 2414, 198, 893, 1305, 839, 649]. perchaudes [649]. Perches [2378]. Percid [2392, 2374, 2389, 2396, 2427, 2425, 2429, 2426, 2418, 2410, 2433]. Percidae [2375]. Percids [2423, 2398, 2421, 2380, 2397, 2383, 2428, 2422, 2416, 2431, 2402]. Perciform [1454]. Perciformes [400]. PERCIS [2433, 2374]. Percomorphi [21]. Percopsis [1271, 1045]. Perennial [2509]. Performance [1142, 2303, 1609, 1450, 891, 2105, 2214, 940]. Performances [2476]. perfused [2585]. perfusée [2585]. Perfusion [58, 1557]. Period [528, 1497, 1662, 1059, 2839]. Periodic [408]. Periodicals [29, 50, 68, 90, 115, 139, 158, 174, 188, 215, 241, 292, 309, 333, 360, 387, 406, 429, 455, 480, 542, 581, 600, 622, 643, 668, 724, 750, 775, 799, 824, 851, 879, 914, 936, 957, 979, 1026, 1110, 1000, 1057, 1083]. Periodicities [2766]. Periodicity [753, 1804, 2253, 1065, 2640]. Periods [1871, 529]. Periphytic [1153]. Periphyton [2302, 302]. Periscopic [1685]. Permanent [2674]. Perrot [649]. Persistence [596, 552, 2294]. Persistent [2627]. Perspective [2748, 1771, 1927]. Perspectives [1400, 2880, 2550]. Perturbations [1993]. Peru [1204]. Peruana [1205]. Peruvian [1203, 1205]. Pesquera [1205]. Pesquerías [1212]. Pesticides [624]. petites [1211]. Petrale [370, 517, 1761]. Petroleum [978, 2629, 2608, 2337, 895, 2149, 2979, 2602, 2618, 2616, 1756, 2068, 2613, 2600, 2747]. Petroleum-Inducible [2747]. Petromyzon [2189, 756, 1071, 959, 2363, 876, 1586, 194, 854, 1266, 1869]. Petromyzontidae [889]. pF [424]. pH [861, 2667, 734, 1780, 2578, 2852, 2459, 1606, 1833, 2258, 2204, 798, 1022]. pH-Induced [2578]. Phaffia [2493]. Phagocytosis [1363]. Phanerodon [422]. pharao [55]. Phenolic [1756]. Phenological [2561]. Phenols [1534]. Phenotypes [2478]. Phenotypic [1252]. Pheopigment [1399]. Pheopigments [1290]. Phialidium [1033]. Phillipson [348]. Phoca [1298, 373, 2337, 935, 1808, 2071, 2941, 1105]. Phocoena [741, 818, 1647]. Pholis [240]. Phoma [1699]. Phosphamidon [1438, 1748]. Phosphatase [1675, 541]. Phosphate [1513, 70, 1988, 1561, 530, 1085, 171, 886]. Phosphatidylcholine [2155]. Phosphoglucomutase [1412, 2156]. Phospholipase [1986]. Phosphorus [2546, 1995, 1892, 2897, 2793, 2842, 1469, 2757, 129, 713, 162, 386, 759, 2954, 1891, 2040, 1094, 1988, 2151, 2191, 1383, 2470, 2744, 2152, 2758, 2558, 2894, 2273, 2210, 2951, 2556, 2804, 2533, 2489, 1763, 1882, 1883, 3, 2674]. Phosphorylase [470]. phosphorylcholine [2155]. Photic [466]. Photochemical [1489, 2514]. photogenis [1074]. Photographic [578]. Photography [88, 784, 1543, 1605]. Photoperiod [2248, 2221, 2952, 739, 79, 2722, 151, 1942, 1159]. Photoreceptor [1711]. Photoreceptors [2378]. Photosynthesis [385, 1139, 300, 1337]. Photosynthetic [2554, 1138]. Phototrophic [39]. Phrixocephalus [1444]. Phthalate [1547, 2128]. Phylogenetic [324]. Physa [149]. physalus [790]. Physical [31, 2085, 553, 1740, 1891, 2005, 2470, 432, 2215, 1336, 2788, 1831, 2395, 2779]. Physics [1342, 1369]. Physiological [1864, 2621, 1401, 628, 2279, 2671, 1824, 362, 2382, 2699, 1828, 1491, 634, 1022, 846, 990, 2139, 2847]. Physiology [1681, 2701, 2008, 2612, 1977, 2848, 2381]. Phytomacrofauna [469]. phytoplancton [2950]. Phytoplankters [1792, 2739]. Phytoplankton [2434, 2936, 385, 165, 1493, 232, 52, 1338, 2302, 1894, 1895, 1718, 2470, 2471, 168, 1917, 301, 2951, 300, 812, 2533, 2690, 2787, 166, 1784, 2109, 2146, 2079, 2308, 2854, 2760, 1763, 1344, 2950, 1138]. Picloram [2032]. Pierre [2693]. Pigment [2452, 1446, 2554, 2493, 1397, 2232]. Pigment-Producing [1397]. Pigmentation [1034, 247, 340]. Pigments [2378, 2268]. Pike [1446, 1711, 909, 2479, 2830, 1943, 2390, 60, 645, 802, 224, 470, 410, 2537, 1039, 2980, 2403, 996]. Pikeperch [2387, 2399, 2382, 2388, 2385, 2403]. Pikeperches [2378]. pileus [1033]. Pilot [2454, 1641, 1637]. Pimephales [1849, 1850, 547, 2477, 2785, 435, 1613, 659, 2663, 1143, 2442, 911, 2820, 2128, 1582, 726, 1037, 2366, 2579, 433]. Pinfish [2256, 2476]. Pink [167, 733, 1407, 2142, 2942, 610, 1553, 564, 1016, 2718, 1776, 2126, 1063, 1768, 392, 365]. pinniger [730]. Pisaster [1149]. Pisces [457, 126, 245, 1444, 324, 1365, 2768, 723, 833, 1900]. Piscicolid [1260]. Piscicolidae [2036]. Piscine [1997, 2524, 2209, 2340]. Pisodonophis [1900]. Pituitaries [1003]. Pituitary [2448, 2018, 611, 336]. Place [2529, 917]. Placopecten [2836, 64, 9, 396, 858, 1384, 2763, 1659, 2443]. Plaice [2448, 838, 1058, 1664, 1574, 1564]. Planes [513]. Planetary [2099]. planipes [1692, 1359]. Planktivorous [2717, 2069, 1416]. Plankton [312, 1349, 808, 1369, 620, 2319, 716, 1139, 2544, 2770, 1133, 2927, 303, 786, 942, 95, 1800, 2345, 2788, 2480, 2475, 1294]. Planktonic [1033, 572, 39, 2799, 1414, 880]. Planned [1460]. Planning [2964, 2829, 1209, 2651, 1461, 1189, 1215, 2668, 1355, 2965]. Plans [1195]. Plant [2864, 2770, 2628, 425, 1042, 1699]. Planted [1899, 1283, 584, 2571, 2368]. Planting [2047]. Plants [1651]. Plaque [2851]. Plaque-Forming [2851]. Plasma [2735, 743, 2160, 1910, 62, 1035, 1518, 2941, 612, 357, 2841, 2270, 549, 103, 1246, 2324]. Plastein [1131]. Plastic [1286]. Plasticity [1249]. platessoides [2448, 222, 838, 1058, 1664, 1574, 1564]. Platichthys [2121]. platyceros [2455, 630]. Platyhelminthes [271, 446]. Play [1418]. Plecoglossus [1944].

Plecoptera [1262]. Pleistocene [2228]. Pleomerism [1789, 2800]. Pleurobrachia [1033]. Pleuroncodes [1692, 1359]. Pleuronectiformes [1444]. Pleuroxus [2243]. plicatus [274]. Plotting [285]. plumbeus [117]. plumchrus [988]. Plume [2291, 2295, 2815]. Plumes [2294, 2292, 2293]. Poacher [126, 476, 577]. Poecilia [711, 206, 781]. Poeciliopsis [1127]. poids [649]. Poikilotherms [1864]. Poinsett [1913]. Point [2611, 2764, 2692, 2644]. Poison [532, 823, 822]. Poisoning [457, 251, 43]. Poisons [1619]. Poisson [20]. poissons [2794]. Poland [2388]. Polar [1338, 2335, 1853]. Polarized [468]. Polarographic [2574]. Policies [2194, 1239, 2881]. Policy [2871, 2172, 1762]. Polish [2387]. Polistotrema [108]. Political [2592]. Pollachius [2933]. Pollock [2933]. Pollutant [2828, 2761, 2776, 1854]. Pollutants [2945, 2079, 433]. Polluted [129, 2610, 1338, 2652]. Pollution [2072, 1792, 1076, 39, 2008, 162, 1682, 2628, 2149, 647, 2739, 2072]. polyacanthocephalus [1589]. Polyacrylamide [976]. Polychaeta [495, 2884, 490, 492, 486, 484]. Polychaete [493, 2607, 491, 489, 243, 1458, 2255]. Polychaetes [762, 497, 2977, 485, 934, 498, 487]. Polychaetous [488, 494]. polychete [496, 496]. Polychlorinated [624, 1755, 1710, 1434, 2673, 662]. Polyculture [1935]. Polycyclic [2973, 2609]. Polymerase [12]. Polymorphism [423, 1024, 182, 1412, 2156, 99, 267, 177]. Polymorphisms [272, 2837]. Polypeptides [156]. polyspinis [511]. Pomatomus [1060, 617]. Pomfret [717]. Pond [1314, 2978, 439, 1937, 1548, 2801]. Ponds [312, 1349, 1655, 2011, 1353]. Pontella [1926, 2231]. Pontoporeia [204, 379, 472]. Pontoporia [1633]. Pool [452, 1015, 405, 2176]. Population [1712, 2313, 118, 2178, 467, 646, 1636, 1658, 2411, 2400, 2344, 2025, 1119, 1858, 2445, 67, 2194, 2838, 1662, 1786, 1915, 2341, 616, 1067, 1751, 592, 1773, 820, 2419, 712, 2412, 1548, 1509, 2732, 2339, 2553, 1124, 2582, 1413, 2659, 2890, 1641, 2819, 2438, 2406, 1625, 243, 103, 2510, 516, 2638, 2300, 2745, 836, 280, 411, 51, 2430, 2968, 463, 632, 655, 1827, 2402, 1567, 2853, 1579]. Population-Genetic [2838]. Populations [1370, 2142, 1733, 960, 1120, 1449, 1272, 57, 1590, 2911, 1299, 2611, 808, 887, 165, 1559, 2657, 1403, 331, 2112, 2394, 2396, 2615, 124, 1044, 199, 415, 1563, 2955, 278, 2418, 2227, 223, 1075, 1357, 1361, 1631, 2336, 1025, 1629, 2409, 2716, 276, 973, 2067, 2892, 1753, 2200, 2307, 433, 2414, 1521, 2839, 829, 2969]. Pore [2151]. Porpoise [741, 1647]. Porpoises [818]. Port [2935, 463]. Portable [1133]. Portion [612, 883]. Portugal [484]. Portuguese [1318]. Porvoo [2885]. Position [1527]. Positioning [1861]. Positions [1685]. Possibilities [1655, 1956]. Possible [204, 106, 1669, 2837, 338, 2420, 1787, 721]. Post [1326]. Post-Larval [1326]. Posterior [1787]. Postexposure [1972]. Postglacial [653, 335, 1104]. Postlarvae [1924]. Postmortem [2129, 9, 396, 224, 470, 1659]. Postrigor [1691]. Postscript [704]. Postsmolt [1159]. Postsmolts [2700]. Potamodromous [2185]. Potamogeton [2640]. Potential [1850, 2113, 1449, 1304, 1360, 1961, 597, 1937, 525, 1698, 2273, 1758, 2679, 1797, 783, 2596, 2304]. Potentials [1176]. Potentiated [1255]. Potentiometric [2535]. Pots [195]. poutassou [1603]. Power [779, 2864, 2770, 425, 1042]. Practical [1570, 1950, 2271, 2577, 426]. Practice [1174]. Practices [2653]. Prairie [1655, 2458, 2521, 2895, 1353, 1969]. Prawn [2947, 2876, 2910, 2455]. Prawns [1937]. Pre [81, 315]. Pre-Smolt [81, 315]. Preamble [1288]. Precambrian [2047, 2368, 1095, 2556, 2116, 46]. Precipitating [164]. Precipitation [2267, 2568, 2632, 1035]. Precocious [2726, 2126, 2773, 2700]. Precopulatory [631]. Predation [1690, 741, 2275, 1009, 1302, 1866, 2836, 959, 1667, 1822, 2931, 566, 2544, 2738, 410, 500, 479, 1800, 2637, 2818, 2730, 2335, 2067, 548, 737, 986, 1424]. Predator [909, 2522, 1457, 2825, 2520, 2579, 665]. Predators [2069, 243, 2631]. Predatory [1586]. Predicting [1469, 1684, 1965, 2338]. Prediction [1066, 1655, 2709, 390, 2710, 796, 2226]. predictive [2226]. Predominance [170, 2386]. Preface [1622, 1877, 1930, 2084, 1166]. Preference [2281, 2649, 2282, 2731, 2283, 81, 2825, 2918, 2278, 1728]. Preferences [1593]. Preferenda [143]. Preferred [2378, 2201, 1848, 2727, 193, 2052, 1841]. Pregnancy [1514]. Preheat [1673]. Preliminary [299, 1040, 1258, 1635, 2611, 983, 534, 1634, 597, 2601, 1642, 1019, 306, 383, 1047]. Premigrant [1025]. Prenatal [803]. Preparation [2851, 16, 1326, 931]. Preparations [2448, 2205]. Prerigor [1384, 1325, 1383]. près [1730]. Presence [2435, 2142, 1755, 270, 2021, 2736, 1473, 2601, 549, 2888, 898, 1748, 427, 472]. Present [1529, 1224, 1629]. Preservation [2353, 760, 1747, 1746]. Preserved [371, 558]. Prespawning [110, 2221, 402]. Pressure [614, 865, 1068, 1531, 1374, 2151, 1081, 409]. Pressure-Sensing [1081]. Pretreatment [2180]. Prevalence [1843, 758, 109]. Prevention [1953, 2634]. Previously [1038]. Prey [2976, 909, 2069, 2456, 1667, 2522, 2904, 2665, 2070, 1457, 1414, 1387, 2579, 1728, 737, 986, 1430]. Priacanthus [1163]. Price [941, 2536]. Prickleback [265]. Primary [991, 1088, 1087, 1289, 2037, 1470, 1895, 1852, 1517, 95, 1967, 226, 300, 1092, 754, 755, 1389, 1784, 2357, 1344, 977]. Primer [2593]. Prince [747, 921]. Principal [9, 1610, 2315]. Principally [2884]. Principles [1773, 2810, 2549, 1846]. Printer [285]. Probabilistic [2793]. Probabilities [2174]. Probability [2717]. Probable [2712, 1070, 2494, 999]. probatocephalus [2583]. Probe [1077]. Problem [675, 14, 1993, 472, 1797]. problèmes [1211]. Problems [1169, 2212, 1240, 1239, 2134, 1196, 1189, 2463, 2464]. Procedural [2278]. Procedure [381, 1086]. Procedures [1099, 1650]. Proceedings [2654, 2703, 2216, 1818, 704, 2374, 671, 2500]. Process [178, 1705]. Processed [2119, 2483]. Processes [1342, 1898, 1048, 1436, 2765]. Processing [350, 713, 2241, 773, 180, 326, 1359]. Processor [2473]. Procladius [2434]. Produced [189, 624, 733, 1268, 730, 1118, 42].

Producing [311, 1397, 1823]. Production [1265, 505, 2983, 2984, 2240, 2677, 2939, 431, 1272, 1954, 2391, 2411, 270, 1719, 2434, 1772, 2748, 1088, 1087, 2011, 1433, 460, 2333, 1289, 1519, 2831, 2037, 907, 820, 1015, 522, 1338, 712, 2432, 474, 1353, 1124, 369, 1641, 525, 2759, 2438, 1895, 2833, 2191, 1852, 1080, 1821, 2471, 147, 95, 1967, 1339, 301, 411, 915, 869, 226, 300, 812, 1092, 1093, 2533, 2260, 2447, 2480, 2832, 754, 755, 1389, 1784, 2014, 2109, 2308, 2506, 217, 2968, 2781, 171, 1344, 1288, 1717, 2707, 2967, 2446, 2697, 2831]. Productivity [1272, 991, 8, 2712, 1197, 1470, 2302, 2470, 1517, 243, 855, 2781, 977]. Products [2068, 1572, 2535, 2794]. productus [2782, 49, 850, 463]. produits [2794]. Proecdysis [1066]. Professor [1678]. Profile [1873, 1292, 2160]. Profiles [841]. profondeur [2950]. Profundal [205]. Progeny [954]. Progesterone [1514]. Program [1075, 2873, 2875, 325]. Programming [2540]. Programs [350, 2971, 15, 2882]. Progress [1640, 1633, 1946]. Project [2540]. Projects [1522]. Prolactin [611, 1842]. Prolarvae [2914]. Prolarval [925]. Proliferative [1750, 2053]. Prolonged [529, 2197, 1743]. promelas [1849, 1850, 547, 2477, 2785, 435, 1613, 659, 2663, 1143, 2442, 911, 2820, 2128, 1582, 726, 1037, 2366, 2579, 433]. Promoting [1223, 1032]. Propensity [2142]. Properties [2392, 1370, 357, 2583, 931, 247, 1689, 2788]. propinquus [2111]. Propionate [2803, 1003]. Proportional [18]. Proposal [1616]. Proposals [2134, 2463, 2464]. Proposed [2113, 1296]. Prosobranchia [2611]. Prosopium [2824, 1532, 2104, 638, 905, 843, 288]. Prospects [1196, 1219, 1199, 1949]. Protanypus [1550]. Protection [1279, 1925, 1241, 2321]. Protective [2245]. Protein [2948, 74, 1051, 881, 1405, 1938, 61, 1607, 2924, 62, 1518, 238, 1326, 1136, 1367, 2537, 4, 1131, 103, 845, 1082, 434, 328, 48, 1150, 1372, 1376, 2175]. Proteins [1043, 246, 976, 1035, 549, 1694, 2658]. Proteolytic [286]. Provided [419]. Provincial [2651]. Proximate [1395, 2280, 2111]. Prudhoe [2606, 2609, 1768]. Pseudobranchectomy [2452]. Pseudocalanus [1597]. pseudoharengus [646, 452, 1284, 2070, 2542, 1422, 2209, 2348, 999]. pseudolimnaeus [149, 547]. Pseudomonad [286, 2719]. Pseudophyllidea [1464, 2276, 1269, 778]. Pseudopleuronectes [2448, 105, 1686, 1364, 439, 103, 268, 1158, 524, 1827, 556, 1073]. Pteropods [2494]. Ptychocheilus [1313]. Public [2306, 1645]. Publications [269, 293, 310, 334, 361, 388, 407, 430, 456, 481, 543, 562, 582, 601, 623, 644, 669, 725, 751, 776, 800, 825, 852, 1137, 1164, 483, 173, 454, 749, 1056, 1419, 1506, 2057, 1419, 1506, 2057]. Published [29, 50, 68, 90, 115, 139, 158, 174, 188, 215, 241, 292, 309, 333, 360, 387, 406, 429, 455, 480, 542, 581, 600, 622, 643, 668, 724, 750, 775, 799, 824, 851, 879, 914, 936, 957, 979, 1026, 1110, 1000, 1057, 1083]. Puget [1858, 1712, 1667, 492, 1016, 69, 2912, 1380, 2033, 49, 1670]. pugettensis [1527]. pugilator [794]. pugnax [2615]. pulex [2916]. pulicaria [839]. Pulp [2299, 2298, 2303, 1099, 919, 2294, 2292, 2305, 2290, 2293, 922, 2291, 2302, 2296]. Pulping [937]. Pulpmill [2365, 1609, 1651, 2256, 1362, 1604, 1777, 1987, 2244, 2450, 2722, 2906, 2109, 2146]. Pulse [809]. Pump [2770]. Pumped [1652]. Pumpkinseed [997, 2505, 2519, 1459]. punctatum [829]. punctatus [1130, 1764, 1302, 2706, 1910, 1115, 1992, 2978, 357, 1300, 1675, 1730, 1578, 2581, 951, 869, 2919, 580, 1109]. punctulatus [2362]. Pungitius [1040]. Pup [2312]. Pupae [1530]. Pups [1788]. Pure [1733]. Purging [1837]. Purification [1247]. Purines [1401]. Purposes [2113]. Purse [2817]. Purse-Seine [2817]. Pygmy [2824, 843, 1626]. pyridiniummethylcarbonylhydrazino [247]. Pyridoxine [1491]. Pyridoxine-Deficient [1491]. Pyruvate [1659, 1694].

quadricornis [621, 2587, 2150]. quadricornis-Complex [2150]. quadruplicatum [2268]. Quahaug [1476, 178]. Quahog [1716]. Qualicum [144, 365]. Qualitative [1602]. Quality [2580, 1309, 2567, 2923, 1343, 382, 2135, 2134, 2463, 2464]. Quantification [2897, 2841, 2839]. Quantifying [1838]. Quantitation [1115]. Quantitative [733, 1735, 916, 469, 2883, 1602, 1720]. Quebec [2434, 1718, 2191, 2751, 327, 1379, 653, 2212, 1373, 638, 905, 1730, 1018, 1080, 1821, 1104, 2637, 51, 1730]. Queen [629, 714]. Questions [1178]. Quick [1384, 419]. Quick-frozen [1384]. Quinaldine [290]. quinquedens [1596]. quinqueradiata [1936]. Quinte [2417, 521, 523, 522]. Quotas [1230]. Quotient [474, 2664].

R. [242, 1249, 261]. Rabbit [1998]. Races [1415]. Racial [1245]. radiata [806, 261]. Radiation [755]. Radio [2907]. Radioactive [531]. Radioactivity [2645]. Radioimmunoassay [2841]. Radioisotope [506, 1254]. Radiometer [23]. Radionuclides [2589]. Radiostrontium [639]. Radium [1090]. Radium-226 [1090]. Radon [1089]. Railliet [514]. Rainbow [868, 1265, 1406, 164, 1279, 2851, 2588, 559, 604, 1899, 93, 416, 788, 1097, 2660, 1656, 2735, 2785, 2168, 787, 1785, 123, 1428, 1904, 1920, 2344, 2160, 2706, 203, 509, 865, 351, 67, 1839, 2767, 1135, 2641, 584, 111, 460, 526, 2859, 892, 1822, 856, 2691, 2016, 2930, 2129, 2849, 2457, 2577, 2118, 926, 2000, 2578, 2266, 1695, 1513, 2155, 707, 2803, 1261, 773, 1048, 1837, 1943, 264, 417, 1502, 546, 397, 2909, 2107, 1990, 378, 1552, 1566, 2727, 1537, 1807, 1651, 2270, 1020, 1991, 286, 1595, 367, 849, 1859, 938, 1410, 1367, 1986, 1432, 535, 2722, 1608]. Rainbow [1591, 2955, 1969, 816, 2771, 1014, 223, 944, 1100, 1028, 2570, 1974, 1901, 2769, 1451, 2346, 2478, 1151, 1129, 2946, 1558, 1452, 1998, 1474, 1062, 1828, 2730, 1387, 615, 1491, 2348, 2815, 1569, 2357, 1587, 2214, 2928, 2318, 2063, 1687, 964, 993, 548, 737, 986, 2725, 97, 135, 554, 1575, 2139, 2847, 442, 2304, 177, 952, 1742, 1743, 2658, 1150, 1372, 2122, 2175, 2324]. Rainforest [2014, 2506]. Rainhow [1421]. Rainy [2400]. Raja [282, 261, 1125]. Ram [184]. Rancidity [2584, 2791, 1986]. Random [2813, 836]. Randomly [2890]. Range [960, 1519, 359, 2888, 1691, 66, 404, 1324, 667, 999]. Ranges [1934, 1575]. Ranging [2941]. Rank [1757]. Rapeseed [2658]. Rapid [374, 2939, 2698, 261, 2205, 2682, 1595, 284, 1987, 913, 1050, 2844]. Rapidly [2244]. Rare [563]. Rarely [1285]. raschii [2240, 2199]. Rastrelliger [906]. Rate [2180, 1864, 221, 197, 1754, 2633, 2976, 2303, 256, 2069, 2363, 2838, 282, 2607, 2265, 938, 1380, 288, 2010, 705, 1041, 65, 2491, 940, 2860, 1159, 1778]. Rates [1076, 2677, 2939, 920, 262, 2411, 962, 1524, 1089, 1555, 2077, 28, 2831, 1310, 2266, 316, 809, 1254, 2369, 2534, 2, 2759, 888, 1383, 1459, 1025, 2120, 2110, 973, 1313, 2753, 1768, 140, 556, 1145, 1883, 1138]. Rathke [2026]. Ratings [707]. Ratio [2976, 2212, 887, 434, 1399]. Ration [1754, 2751, 1827, 2538]. Rational [2878]. Rationale [295]. Rationalization [1195, 2869]. Ratios [262, 1272, 385, 1266, 2858]. Rats [533, 393]. rauschi [2053]. Raven [338]. Raw [180, 533]. Ray [37, 506, 318, 2921, 2171]. Rayed [2132]. Rays [2358, 442]. Reaction [1850, 1749, 1131]. Reactions [2279, 2898]. Reactive [2757, 2758, 2161]. Reactivity [2923]. Reading [98]. Reagent [864, 2076]. Reagents [540]. really [2464]. Reared [1985, 2237, 2047, 2513, 2118, 707, 2773, 2961, 2043, 2573, 2700]. Rearing [1859, 630, 715, 2446]. Reason [2800]. Recapture [325]. Receiving [2298, 796]. récentes [1419, 1506, 2057]. Recently [674, 748, 492]. Receptacle [421]. Receptor [1050]. Reciprocal [1159]. Recirculating [1744, 1851, 2272, 2370]. Recirculation [715, 718]. Recognition [2635, 196]. Recolonization [2622]. Recommendations [1244, 2285, 2278]. Reconciling [1184]. Reconsideration [467]. Reconstruction [2310]. Record [520, 1260, 1101, 1431, 560, 876, 332, 476, 83, 719, 1281, 1155, 1417, 667, 45]. Recorded [313]. Recorder [445]. Recording [614, 2649, 972, 1106, 2370]. Records [790, 510, 653, 342, 618, 969, 1074, 502, 538, 1162, 498, 1626, 216, 723, 515, 54, 1146, 1117]. Recoveries [2571, 1921]. Recovering [1131]. Recovery [978, 159, 2911, 2496, 584, 2960, 2625, 767, 1122, 1990, 2628, 2619, 2596]. Recreational [2465]. Recruitment [2629, 1172, 2235, 1693, 1993, 2044, 2539, 2041]. Recycling [2252]. Red [1596, 2492, 2140, 1359, 321, 2261, 1070, 2409, 2232]. Red-Water [321]. Redescription [1486, 493, 492, 1107, 563]. Redfish [1268, 180, 1117]. Redhorse [2176]. Redmouth [1279, 1785, 2164, 2236, 952]. Reduced [1409, 891, 1833, 2227, 996]. Reductase [508]. Reduction [2233, 780, 763, 1420, 2514]. Reductions [2246]. Reevaluation [2114]. Reexamination [2268]. Reference [1850, 467, 1638, 1556, 273, 1209, 1559, 2405, 1741, 1993, 647, 409, 571, 30, 1215, 2068, 412, 2564, 2814]. References [2286]. Refined [1254]. Refinery [2628, 1756]. Reflex [38]. Refrigerated [2272, 2119, 311, 2239, 1319, 4, 1747]. Refrozen [2580, 2483]. Regarding [1827]. Regeneration [2251, 1891, 2027]. Regime [763, 1041, 1159]. Regimes [2170, 298]. Region [1369, 818, 2089, 2878, 1070, 2153, 1572]. Regional [99, 1210]. régionales [1210]. Regions [318, 2336]. Regression [1395]. Regressions [1677, 924]. Regulating [2879, 2413, 2883]. Regulation [2221, 2157, 2870, 25, 1942, 2882]. Regulations [1187, 1231, 1185, 1186]. Regulatorias [1205]. Regulatory [1205]. Rehabilitation [2054]. reicherti [2479]. Reinforcement [1396]. Reinhardtius [209, 1533, 176]. Related [2180, 2378, 1401, 1964, 1865, 787, 2921, 1343, 1700, 126, 858, 1237, 2916, 1789, 1957, 99, 2779, 249, 1109]. Relates [1347]. Relating [1436]. Relation [2512, 505, 1071, 920, 1754, 490, 1636, 270, 1527, 1060, 1052, 2937, 111, 2187, 2503, 1234, 323, 1304, 2199, 73, 642, 2123, 1688, 707, 253, 232, 157, 264, 1015, 1857, 2390, 1368, 2305, 2764, 2775, 180, 1414, 380, 666, 2537, 2581, 243, 2441, 2504, 830, 1462, 1564, 1389, 2395, 2117, 462, 1287, 1827, 1687, 2307, 2297, 1717, 1794, 2475, 343, 2376, 1430, 277]. Relations [2387, 471, 401, 2385, 625, 1022, 1305]. Relationship [2643, 2235, 1433, 58, 717, 2946, 1903, 2480, 65]. Relationships [32, 2024, 2976, 865, 2162, 2522, 2195, 2850, 1895, 2471, 1144]. Relative [1423, 2477, 1593, 2929, 1239, 545, 63, 316, 1385, 2046, 1660, 2167]. Release [1130, 1393, 1766, 1852]. Released [1921, 2223, 2488]. Relevant [2327, 2747]. Reliability [1077, 2190]. Relict [379, 472]. relicta [204, 1295, 2862, 832, 2749, 2689]. Remains [2206]. Remarks [2212]. Remote [125]. Removal [1995, 772, 1662, 746, 2394, 284, 845, 1082, 171, 1484, 2737]. Renal [1732, 1300, 2696]. Rendus [1440, 2056, 2082, 2136, 2030]. Renewable [2540, 1460]. Renewal [232]. Repeat [170, 2331]. Repeated [1595]. Repertoire [2542]. Replacement [1428]. Replications [1109]. Reply [2462, 2688, 2686, 2926, 2965, 2984, 2464]. Report [1640, 1242, 1180, 1188, 1202, 1207, 1213, 1220, 1227, 1233, 1238, 1623, 2568, 1633, 1635, 2654, 2750, 1642]. Representation [848]. Representing [2975]. Reproduction [831, 467, 659, 882, 227, 1939, 2356, 366, 2368, 1037, 148, 1942, 46, 1783]. Reproductive [1616, 2423, 242, 2261, 1449, 631, 323, 1304, 239, 2519, 2732, 2443, 2892]. Reptant [502]. Reptiles [1679]. Republic [1758]. Required [1075]. Requirements [1258, 1668, 544, 866, 1951, 1774, 1010, 2383, 1046, 1960, 403, 319, 1150, 1376, 1838]. rerio [1611]. Rescue [2620]. Research [1640, 2669, 173, 454, 749, 1056, 2568, 1600, 1916, 1922, 1677, 2775, 326, 1703, 924, 1676, 2805, 2536, 1347, 2564, 1356, 1970]. Reserves [271]. Reservoir [2206, 763, 2404, 947, 30, 2386, 1758, 2269, 2405]. Reservoirs [1176, 2406, 2384]. Resident [152, 943, 2509, 1261, 2526, 1380, 291, 408]. Residual [2163, 1914, 2362, 1111, 1793]. Residuals [1919]. Residue [958, 2264, 2128, 1854]. Residues [624, 1126, 2312, 1, 1782, 1029, 287, 2035, 2301, 2599, 895, 2149, 2979, 290, 2208, 1275, 1540, 268, 279]. Resistance [2276, 1596, 923, 770, 1490]. Resistances [2167]. Resonance [192]. Resource [1331, 2540, 2327, 2333, 1355, 706, 1949]. Resources [1169, 2826, 1173, 695, 2351, 1183, 1199, 1224, 1460, 1182, 1177]. Respiration [961, 919, 2199, 1140, 980, 832, 705, 2120, 1692, 2613, 556, 1337, 2051]. respiratoire [2867]. Respiratory [1122, 474, 338, 803]. Respirometer [155]. Respirometry [742]. Response [1864, 1406, 2621, 2521, 2555, 529, 1809, 2362, 2342, 137, 2333, 2016, 613, 1860, 331, 568, 80, 1538, 898, 1371, 1748, 1591, 1503, 2343, 1465, 1466, 2559, 1129, 2679, 2357, 1728, 2139]. Responses [2378, 513, 628, 509, 2445, 2671, 1068, 1531, 460, 248, 769, 2930, 2394, 2877, 1003, 2429, 2722, 1718, 2699, 2631, 2213, 2068, 2203, 338, 1828, 2431, 218, 349, 2167, 125]. Responsible [342]. Responsiveness [1862]. Rest [1604]. Resting [2231]. Restoration [2496]. Restoring [1948]. Restraint [1392, 1842]. Result [933]. Resultant [2163]. Resulting [731, 1839, 2246, 2597, 2200]. Results [2423, 296, 2915, 2439, 365]. Resuspended [1049]. Retain [613, 5]. Retained [1733]. Retarded [1495]. Retention [2148, 1403, 2040, 2744, 2455, 393]. reticulata [711, 781]. Retinal [2377, 1446]. Retinas [2688, 2687, 2468]. Retrieving [2490]. retrocurva [2303]. Return [341, 2890]. Returns [1407, 1778]. Reuse [2705]. Revealed [1526, 710, 1063]. Reversible [612]. Review [2568, 2546, 506, 1774, 1739, 59, 627, 1816, 2795, 2399, 2794, 1957, 1775, 485, 2306, 2915, 2867, 626, 1639, 147, 1771, 2922, 2216, 2495, 196, 676, 2010, 1341, 778, 2589, 1347, 963, 2355, 2468, 2777, 321, 54, 1344, 1645, 1356, 2986, 2030, 2687, 2688, 1523, 1620]. review/Comptes [2030]. Reviews [1440, 2056, 2082, 2136]. Reviews/ [1440, 2056, 2082, 2136]. Revised [167, 733, 1663]. Revision [2695, 484]. Revisited [1931]. Rhabdochona [514]. Rhabdochonidae [514]. Rhacochilus [804, 803]. Rheotropic [2221, 2767]. Rhinichthys [242, 202, 583, 616, 1249, 752]. Rhode [1422, 2600]. rhodozyma [2493]. rhomboides [2256, 2476]. rhotheus [479]. Rhynchobdellae [2036]. Rhythms [1918]. Ribbed [343]. Ribonucleic [65]. Ribonucleotide [508]. Richards [1979]. richardsoni [1146, 1125]. Richland [2216]. richmondensis [369]. Ricker [1857]. Rigging [1618]. Right [790, 2962, 551, 1626, 418]. Rights [2883]. Ringed [2337, 1808, 2071, 2941, 1105]. ringens [1203]. Rings [2220]. Ripening [1055]. Rising [2201]. Risk [1655, 2895, 2890, 986]. River [946, 2824, 2228, 1532, 1590, 1379, 2509, 2104, 2212, 821, 2513, 590, 2042, 752, 1526, 1373, 1961, 638, 905, 1921, 2519, 316, 2425, 2130, 144, 2264, 91, 402, 291, 1044, 917, 2972, 1104, 1715, 950, 2068, 985, 412, 939, 268, 837, 998, 747, 2121, 152, 341, 2511, 587, 1148, 1447, 287, 344, 1555, 2077, 1350, 2173, 1450, 60, 2172, 499, 795, 883, 1730, 1042, 1472, 2406, 1854, 1019, 1800, 1422, 1663, 1063, 392, 964, 365, 89]. Riverine [2513]. Rivers [375, 2560, 2781, 2640]. RNA [2976, 262, 887, 2861, 2957, 2175]. Roach [2386]. robusta [490]. Rock [2721, 2836, 791, 2876, 2441, 588]. Rockfish [708, 2816, 730, 2662, 1670, 1047]. Rockfishes [400, 198]. Rockflsh [511]. Rocky [1349, 2622]. Rodlet [2516]. Roe [1698]. Role [56, 2857, 997, 1447, 2717, 417, 1179, 1693, 1868, 1859, 1947, 1200, 2804, 1933, 1787]. rondeleti [1163]. Ronge [117]. Room [1987]. rosaceus [376]. Roseau [402]. rosenbergii [2910]. rosmarus [2680]. rostrata [772, 1060, 2212, 928, 652, 1396, 719, 966, 1377, 971, 1134, 1408, 2482]. Rosyface [2362]. Rotenone [159, 594]. Rotifers [2153]. Rough [596]. Round [1532, 2119, 2483, 638, 905, 2197]. Round-Frozen [2483]. RTG [2666]. RTG-2 [2666]. rubellus [2362]. rubrum [321, 2232]. ruckeri [2851].

Rudolphi [1510]. Ruff [2399]. Rugophilic [1503]. Run [2042]. Runs [610, 2940]. rupestris [2441]. Rutilus [2386].

S [1561]. S-140 [1561]. S. [2377, 2588, 954, 1280, 1423, 118, 1283, 1006, 1848, 923, 2112, 2850, 2382, 2406, 1974, 486, 2346, 2105, 2067, 2395, 5, 252]. Sabellidae [2884]. Sable [2312, 373, 2454]. Sablefish [512, 195, 112, 597, 721]. Sabussowia [921]. Sac [1613, 1052]. Sæther [2481]. Saetheria [2360]. Safe [2529]. Sagavanirktok [291]. Saginaw [2415]. Sagitta [411, 705, 915]. Saguenay [1080, 2562, 2938, 2765, 1779]. saida [2159]. Saint [2226, 1001, 649, 1730]. Saint-Laurent [2226, 1001, 1730]. Saint-Lawrence [1001]. Saint-Louis [649]. saisonnière [910]. saisonniers [1001]. Salamander [278]. salar [2075, 777, 1, 2452, 375, 743, 2511, 339, 1495, 779, 1333, 2515, 759, 933, 1996, 2513, 1700, 1471, 874, 1751, 1961, 613, 566, 612, 1507, 1806, 2841, 1921, 114, 316, 891, 1973, 2508, 2530, 2130, 170, 1472, 1577, 182, 2523, 1004, 591, 1301, 1420, 1361, 1031, 1588, 2038, 2486, 2170, 1539, 2559, 966, 985, 1558, 2746, 2802, 1721, 1722, 2961, 218, 349, 1011, 2700, 2843, 965, 2147, 2648, 2531, 2781, 345, 2913, 745, 1159, 3, 2184]. salina [2759]. Saline [1690, 1613, 2968, 1823]. Salinities [628, 1500, 2227]. Salinity [370, 399, 503, 1033, 143, 2667, 1809, 743, 544, 614, 1910, 2649, 2187, 2009, 2848, 571, 995, 81, 315, 1722, 2012, 794, 1150, 1372, 1376]. Salinity-Dependent [2848]. Salmincola [1002, 2195]. Salmo [2075, 1265, 777, 1406, 1, 164, 1279, 2851, 461, 2588, 559, 604, 1899, 2452, 375, 2326, 416, 788, 1097, 2660, 1656, 2735, 1925, 2785, 2168, 787, 1785, 1592, 1593, 743, 123, 2511, 1428, 230, 1904, 1920, 2160, 2706, 203, 509, 339, 865, 351, 1495, 67, 1839, 779, 1333, 2767, 2515, 1135, 2641, 759, 933, 1996, 2513, 1700, 2859, 892, 1822, 1471, 856, 238, 2691, 607, 2503, 2016, 2930, 874, 1751, 1961, 613, 566, 2129, 612, 2849, 2457, 2577, 2118, 926, 1507, 1806, 2000, 2578, 2266, 1688, 1695, 1513, 2155, 707, 2803, 1261, 2841, 773, 1048, 1837, 1943, 264, 417, 1921]. Salmo [1502, 114, 546, 397, 2909, 316, 2107, 1990, 891, 378, 1552, 1566, 2727, 1537, 1807, 1651, 1973, 2508, 2270, 2530, 1020, 2130, 1991, 286, 1595, 2850, 85, 367, 849, 1859, 2821, 170, 938, 1410, 1367, 124, 1986, 1432, 1576, 1472, 535, 75, 1577, 878, 1415, 2722, 1608, 1591, 2955, 182, 1969, 2523, 2545, 2771, 1501, 1014, 1004, 591, 944, 1301, 1100, 1028, 2570, 1974, 558, 1420, 1421, 2585, 1361, 1901, 1781, 1031, 1588, 2038, 2486, 2170, 2918, 2769, 1480, 1539, 2837, 1451, 2346, 2478, 2181, 2559, 2818, 1151, 1129, 2946, 966, 985, 1558, 1452, 1998, 2504, 1474]. Salmo [151, 2746, 2802, 1062, 2773, 412, 1828, 654, 1721, 1722, 1394, 1387, 289, 615, 1491, 2961, 2815, 1569, 2357, 218, 349, 1011, 2700, 2843, 965, 2147, 2648, 2531, 2781, 1587, 2214, 2331, 2928, 2318, 2063, 1687, 964, 993, 548, 737, 986, 345, 2725, 2913, 97, 135, 554, 846, 990, 1273, 1575, 2139, 2634, 2847, 442, 2304, 745, 1159, 952, 1742, 1743, 2032, 1521, 2658, 1150, 1372, 2122, 2175, 2324, 3, 2184]. salmoides [1468, 1409, 1302, 2265, 651, 2526, 1759, 2579, 2108]. Salmon [350, 2075, 2512, 777, 1, 596, 1985, 2196, 2985, 2452, 167, 733, 1407, 2142, 375, 501, 2942, 2371, 395, 528, 341, 529, 610, 853, 2940, 862, 861, 860, 1464, 2276, 197, 513, 322, 1142, 920, 1791, 1754, 1865, 1799, 2643, 2163, 1914, 2198, 40, 1487, 1592, 1593, 743, 318, 2511, 587, 550, 1478, 1553, 1906, 1904, 1156, 110, 339, 1099, 919, 801, 141, 26, 468, 1495, 256, 564, 1388, 1571, 779, 1333, 1052, 2141, 512, 342, 1980, 2647, 574, 2952, 258, 137, 2645, 2515, 1667, 759, 1518, 933, 2872, 1996, 2513, 460, 526, 225, 1700, 1826, 1962, 1822, 2187, 1471]. Salmon [1350, 2792, 2907, 606, 1843, 2503, 784, 1526, 1434, 59, 769, 874, 1751, 1961, 1802, 678, 613, 566, 2831, 612, 2018, 2173, 2817, 1507, 1806, 1688, 1609, 1732, 102, 2841, 560, 1860, 2726, 1921, 114, 1934, 2195, 1016, 2172, 316, 891, 2664, 2718, 186, 499, 1418, 2775, 2750, 1049, 937, 1647, 1776, 1973, 2508, 2704, 2530, 2130, 144, 2821, 170, 2126, 710, 1697, 549, 1256, 1804, 1380, 2033, 337, 1032, 1508, 1472, 1577, 611, 288, 922, 1362, 1604, 1777, 2906, 1245, 182, 2523, 2833, 36, 2133, 1004, 2723, 591, 2708, 1301, 1648, 371, 81, 315, 1420, 1465, 1466, 479]. Salmon [1361, 2873, 2245, 1762, 2631, 247, 1031, 1588, 2038, 2486, 2170, 1539, 441, 1275, 1615, 1059, 1663, 2010, 2559, 2638, 966, 950, 1225, 985, 1558, 2504, 151, 2746, 2802, 2321, 2679, 1721, 1722, 1041, 2832, 2364, 975, 1008, 2961, 963, 2043, 1063, 1963, 837, 2563, 842, 1439, 633, 218, 349, 1011, 2700, 2167, 2843, 665, 965, 2147, 2648, 2531, 2781, 508, 1587, 1768, 1282, 2920, 1118, 392, 1687, 267, 42, 365, 1717, 2707, 993, 345, 805, 940, 354, 554, 846, 904, 990, 1273, 2139, 2015, 745, 1159, 249, 1812, 2844, 589, 1376, 2311, 3, 2184, 336, 2913]. Salmon-Canning [842]. Salmonella [364]. salmonicida [164, 596, 2196, 512, 2334, 2503, 109, 1465, 1466, 2913, 2236]. Salmonid [150, 163, 570, 1912, 674, 693, 1535, 702, 696, 703, 1556, 1101, 62, 698, 687, 697, 684, 381, 680, 681, 692, 683, 688, 671, 690, 686, 1987, 2244, 2450, 691, 689, 500, 2336, 701, 682, 699, 700, 2891, 655, 156, 2642, 694, 2446, 685, 2311, 1481]. Salmonidae [2915, 813]. Salmonids [1953, 1255, 1769, 1527, 1496, 1009, 1956, 1360, 2493, 1327, 2714, 677, 2972, 2728, 478, 213]. salmonis [2219, 204]. Salmons [1246]. Salpidae [545]. Salt [1043, 881, 785, 2625, 2615, 2010, 2094]. Salt-Extractable [881]. Salt-Marsh [2615]. saltatrix [1060, 617]. Salting [713]. Saltwater [1377, 2370]. salvelini [1464, 2276, 1008]. Salvelinus [2180, 461, 1315, 1795, 954, 1280, 1423, 118, 1283, 943, 327, 1971, 640, 1409, 1594, 2509, 974, 2736, 2131, 1035, 162, 1662, 2011, 428, 2047, 2571, 1700, 1259, 1848, 2931, 607, 566, 2657, 1143, 1010, 2039, 980, 362, 363, 767, 1122, 923, 1403, 331, 2112, 2459, 2390, 2738, 335, 739, 1586, 329, 124, 2673, 2915, 291, 1044, 1413, 193, 727, 210, 366, 1317, 2143, 1584, 1833, 2955, 1856, 1018, 1821, 2368, 1031, 1467, 2918, 2049, 2046, 1404, 1275, 434, 2745, 340, 280, 51, 654, 2105, 1787, 5, 1287, 2258, 2032, 1579, 1521, 2696]. Sam [1418]. Same [1318]. Sample [2813, 1075, 1549, 2694, 1778]. Sampler [2165, 797, 1456, 469, 119, 2345, 2323, 2158]. Samplers [2979]. Samples [2534, 1561, 348, 84, 448, 1989]. Sampling [2646, 421, 183, 2971, 2211, 191, 1053, 2941, 205, 1922, 1595, 290, 2807, 2257, 394, 119, 317, 2560, 2806, 836, 1392, 1842, 814, 1970, 2798]. San [2935, 1703]. Sand [353, 592, 1268, 2059, 827, 828, 927, 902, 237]. sandvicensis [571].

Sanguinicolid [400]. sapidissima [1590, 866, 785, 1148, 1447, 2042, 1431, 440, 2369, 2732, 795, 24]. Saprophyte [1699]. Sarcoma [2106]. sarcotretis [1117]. Sarsia [1033]. Saskatchewan [964, 117, 287, 2659, 2968, 89, 968, 952]. Satiation [322, 909]. Saturating [1726]. Sauger [2377, 1006, 1960, 2406, 2067, 2395]. Sawdust [939]. Sawdust-Bearing [939]. saxatilis [2914, 544, 618, 1729, 2766]. Saxidomus [532, 823, 822, 1145]. Saxitoxin [2254]. Scale [610, 1192, 1296, 2095, 1195, 2078, 2508, 372, 1215, 1198, 345]. Scales [2358, 395, 528, 819, 98, 2969, 1253]. Scallop [57, 884, 1658, 2574, 473, 355, 64, 9, 396, 858, 1384, 2763, 157, 63, 1659, 2443]. Scallops [2836, 239]. Scanning [1443, 2909]. Scans [1100]. Scattered [1976]. Scattering [782, 1903, 1805]. Scenedesmus [781]. Schaefer [2260]. Schedule [1396]. Schizoporella [201]. Schooling [2069, 1924, 536, 413]. Schools [1685, 145]. Schroederella [496]. Schulek [474]. Sciadonus [1365]. Science [2072, 1494, 2327, 2054, 2080, 626, 1846, 1736, 1976, 2777, 1875, 1847]. Sciences [2669, 2277]. Scientific [2327, 1237, 1171, 2703]. Scientist [2595]. Scientists [2352]. Scintillation [300, 377, 977]. SCOL [704, 676]. Scolecobasidium [1013]. Scomber [2239, 1661, 66, 1427, 2539, 2475]. scombrus [2239, 1661, 66, 1427, 2539, 2475]. Scope [1865, 351]. Scorpaenid [400, 252]. Scorpaenidae [1783]. scorpius [244, 227, 1303]. Scotia [464, 353, 720, 1476, 2312, 373, 2936, 1832, 2454, 2612, 2035, 2970, 2977, 2835, 2599, 1627, 1698, 2449, 229, 1155, 411, 648, 915, 2605, 863, 2623, 2892, 1665]. Scotia-Caught [1698]. Scotian [868, 2319, 2575, 2694, 827, 1564, 2672, 1831]. Scrap [1268]. Scripps [1703]. Sculpin [561, 2025, 244, 227, 1589, 320, 875, 1509, 621, 1303, 479, 1398, 577, 1841]. Sculpins [2587, 2150]. Sea [2189, 1401, 2096, 2498, 862, 861, 860, 1983, 1782, 190, 991, 1529, 894, 2076, 129, 1600, 384, 2836, 2456, 959, 2363, 1111, 126, 2233, 2193, 473, 222, 1232, 64, 396, 858, 1384, 2119, 748, 2763, 1027, 876, 2239, 2628, 1586, 1319, 194, 666, 854, 1602, 563, 1654, 1266, 1539, 1917, 1659, 2004, 1747, 840, 1869, 2001, 2892, 1970, 903, 2262, 2592, 2261, 2071, 2288, 2123, 2102, 447, 1179, 1470, 1398, 2885, 1224, 338, 2316]. Sea-Surface [1529]. Seabird [2620]. Seaboard [635]. seafloor [1847]. Seafoods [2973]. Seal [1298, 1514, 2337, 1788, 2339, 2819, 2637, 872, 1105, 515]. Sealed [2365, 1286]. Sealed-Jar [2365]. Seals [1298, 958, 2312, 373, 1229, 1029, 935, 1808, 2071, 2941, 2637, 855, 994, 1610, 2335, 2316, 2697, 2983, 2984]. Seaperch [422, 804, 803]. Searsport [2602]. Seas [1236, 2080, 1736, 2149]. Seashores [2655]. Season [881, 67, 1045]. Seasonal [189, 868, 2219, 1454, 1899, 2437, 271, 2472, 890, 208, 2342, 1161, 590, 1291, 2077, 323, 73, 2754, 1585, 2383, 2436, 1364, 712, 2764, 186, 2675, 1775, 91, 1708, 1709, 2722, 1767, 2253, 494, 1422, 348, 304, 1399, 998, 2108, 1294, 1896, 2640, 2184, 910, 1001]. Seasonally [1121, 2733]. Seattle [1509, 2121]. Seaward [2513, 1804, 1982]. Seawater [2667, 2272, 2898, 2515, 162, 2799, 2460, 2979, 2700]. Seawater-Maintained [162]. Seaweed [386, 525, 1947]. Seaweeds [2233].

Sebastes [2958, 708, 714, 1248, 2816, 180, 1319, 730, 2662, 1670, 252, 511, 667, 893, 1047, 1305, 1783]. Sebastodes [323, 101, 219]. Secchi [2724]. Second [1407, 1085]. Secondary [369, 2967]. Secretion [256, 258, 1116]. Secretory [888]. Section [1207, 1213, 1220, 1227, 1233, 1238, 1805]. Sections [2782]. Sédentaire [496]. sedentary [496]. Sediment [2629, 221, 2606, 603, 2897, 2857, 537, 2607, 73, 2835, 1049, 1886]. Sediment-Bound [2607]. Sedimentary [916, 2269]. Sedimentation [2943, 2059, 1665, 2674]. Sedimented [2752, 2753]. Sediments [2789, 1147, 2299, 307, 1293, 711, 2754, 1583, 1884, 635, 2602, 2617, 1112, 1654, 1830, 2938, 2609, 2296, 768, 1086, 2605, 306, 2597, 1881, 1880, 2600, 314, 1882, 1883, 2798]. Seen [1285]. Segregation [461, 752, 400, 2915]. Seiche [136, 2935]. Seiches [2103]. Seine [2817]. Selected [1871, 1121, 2159, 297, 772, 1565, 584, 91, 1674, 1798, 298, 2297, 2466, 2351]. Selecting [5]. Selection [2660, 909, 997, 2899, 2456, 574, 2062, 1700, 1259, 2904, 2314, 2665, 2070, 2459, 1859, 535, 1054, 2, 1031, 2038, 2815, 839]. Selective [2717, 2838, 1667, 1992, 1867, 1437, 500, 2818, 2535]. Selectivity [815, 989, 1739, 1913, 392]. Selenium [1528, 2546, 2861, 1611]. Self [124, 23]. Self-Contained [23]. Self-Sustaining [124]. Semiautomatic [1393]. Seminar [2654]. Semiroutine [916]. Semotilus [2555, 917]. Sensing [1081]. Sensitive [1826]. Sensitivities [1050]. Sensitivity [2916, 1445, 1777, 2819, 966]. Sensor [47, 427]. Sensory [2661, 2380, 2197]. Separating [1415]. Separation [2757, 549]. sepioidea [212]. Sepioteuthis [212]. Sept [2654]. September [1818]. Septicemia [1974]. Sequence [1175]. Sequences [419]. Sequential [1247]. Sera [441]. Serial [1392, 1842]. Seriata [2695]. Series [2088]. Seriola [1936, 1163]. Serological [1496, 1397, 1998]. Serpulidae [484]. Serum [2641, 397, 2704, 611, 1392, 1842, 48, 1372]. Serums [164]. Session [1242, 1180, 1188, 1202, 1207, 1213, 1220, 1227, 1233, 1238, 2630, 2621, 2608]. Seston [22]. Set [2490]. setacea [130]. Setal [1066]. Settled [900, 2754]. Settlement [494, 78]. Settling [1292, 1503]. Seven [1364]. Seven-Year [1364]. Seventy [1676]. Seventy-Five [1676]. Several [1316, 270, 436, 1607, 386, 1450, 2820, 835, 1397, 2918, 830, 634, 2307, 1430]. Severity [1797]. Sevin [10, 659]. Sewage [738, 1023, 171]. Sex [2212, 1788, 1266, 2111, 249]. Sex-Related [249]. Sexual [2017, 2650, 2212, 256, 244, 258, 137, 239, 2126, 79, 2773, 2700, 1055]. Sexually [1508]. Shad [1590, 866, 785, 1148, 1447, 2042, 1431, 2206, 440, 2369, 2732, 795, 24]. Shade [1700]. Shagawa [2395]. Shallow [568, 394, 317, 768, 304, 1546, 2640]. Shallow-Water [317, 1546]. Shark [741, 1589]. Shasta [1575, 1697, 2311]. Shawanaga [572]. Sheep [2104]. Sheepscot [882]. Sheepshead [2583]. Shelf [175, 602, 2936, 2288, 2970, 2575, 2234, 2553, 1736, 906, 2319, 2575, 1973, 2694, 827, 1564, 2672, 1831, 1134]. Shelf-Break [2936]. Shelf-Life [906]. Shelf-Seas [1736]. Shelf/Slope [2970]. Shell [179, 2148, 2612, 2613, 1599, 1128]. Shellfish [1658, 1614, 1307, 2763, 532, 823, 822, 814, 1572, 2572]. Shellfishes [279]. Shelters [2020, 770]. Shield [1095]. Shift [2386]. Shifts [1708]. Shiner [2362, 2062, 1074, 2904, 2366, 1277]. Shiners [2221, 1923, 1039]. Ship [1529]. Shipping [1908]. Shock [1479, 1009, 1302, 1424, 2324]. Shocks [2250]. Shore [1832, 168]. Shores [2622]. Short [2840, 974, 2247, 2022, 210, 1641, 1222, 1392, 1842, 2950]. Short-Finned [1641]. Short-Term [2840, 2247, 2022, 210, 1222, 1392, 1842, 974, 2950]. Shorthorn [244, 227, 1303]. Shortnose [949]. Shrimp [642, 1219, 2749, 630, 1412, 1541]. Shrimps [510]. Shunting [1278]. sicculus [1803]. Side [2274]. Sided [2713, 2712]. Sides [2962]. Sidestripe [642]. Sigalionidae [486]. Sight [790]. Signals [1014]. Significance [1401, 39, 1719, 2362, 35, 245, 172, 2837, 1692, 1022, 2063]. Significant [178]. Silicon [2251, 2252, 2253]. Silver [1074, 2119, 2483, 1673, 637, 2772]. Silvering [114]. Silverside [1803]. Similar [2533]. Similarities [308]. Similarity [1024]. Simple [978, 38, 1306, 449, 2490, 930, 1469, 1684, 613, 2575, 381, 1015, 1595, 2553, 2450, 1591, 317, 2336, 972, 451, 973, 448, 1760, 619]. simplex [2178, 1455, 1510]. Simplified [1614]. Simpson [2202, 2203]. Simulacra [2259]. Simulate [2683]. Simulated [1593, 2705, 888]. Simulates [733]. Simulating [1102]. Simulation [2484, 391, 499, 1586, 1854, 2250, 2364, 2474, 433, 2846]. Simultaneous [1309, 2370, 2213]. Simultaneously [62]. Since [2856]. sinensis [1162]. Single [2075, 2160, 1916, 2264, 442]. Sinus [2021]. Siphon [822]. Siphons [2012]. Site [708, 997, 1708, 1797]. Sites [2062, 2459]. Situ [1292, 2751, 2004, 991, 107, 977]. Situated [930]. Six [495, 1028, 1901]. Sixes [837]. Sixteen [307]. Size [1280, 960, 1757, 920, 1754, 631, 141, 1667, 111, 1248, 239, 2070, 114, 736, 1485, 834, 2151, 458, 2896, 2888, 394, 278, 888, 500, 880, 2176, 830, 1041, 2953, 2111, 2200, 737, 1794, 2475, 2538, 1430, 45, 839, 1778]. Size-Frequency [2896]. Sizes [2660, 1127, 1075]. Skaha [1005]. Skate [282, 261, 1125]. Skeena [341, 1526, 2173, 2172, 1663, 392]. Skeletal [414, 436, 1123, 1247]. Skeleton [2359]. Skeletons [324]. Sketch [2085]. Skin [286, 459, 337]. Skipjack [184, 231]. Slave [681]. Slender [1964, 1155]. Slimy [2025, 1841]. Slope [2089, 2970, 1672, 1856, 877, 2346, 2672, 2850]. Sluggish [2520]. Small [1640, 461, 1690, 1899, 2895, 956, 2267, 1649, 1638, 1192, 2511, 1635, 2087, 1594, 1801, 1153, 2095, 1090, 1089, 1135, 2522, 1643, 2047, 1195, 607, 1091, 1335, 2738, 186, 1256, 2626, 2368, 1198, 2556, 348, 2116, 1394, 211, 448, 1639]. Small-Boat [1801]. Small-Scale [1192, 2095, 1195, 1198]. small-type [1639]. Smaller [1620, 1623, 1644, 1636, 1634, 1632, 1645, 2402]. Smallmouth [1046, 2393, 2390, 2185, 46]. Smelt [868, 520, 2344, 162, 186, 200, 816, 2730]. Smelts [2348]. Smoke [773]. Smoke-Processing [773]. Smoked [142, 364, 963]. Smoking [1048]. Smolt [2515, 1826, 2187, 1977, 114, 81, 315, 151, 2746, 2781, 589]. Smoltification [2930]. Smolts [116, 779, 759, 1996, 2513, 874, 1472, 1577, 1376, 1961].

Smooth [596]. Snail [1270, 1140]. Snake [1900]. Snow [1596, 2892, 1128]. Snowmelt [2852]. Social [1757, 1497, 2871, 1209, 417, 837]. Society [1818]. Sockeye [395, 528, 341, 529, 853, 2940, 1464, 2276, 197, 513, 322, 1142, 920, 1791, 1754, 1865, 318, 587, 110, 919, 468, 256, 1388, 1571, 258, 137, 1518, 225, 1822, 1350, 2792, 1843, 784, 1526, 100, 678, 2195, 2664, 499, 2750, 710, 337, 1508, 611, 288, 1245, 36, 2133, 1663, 2638, 2504, 1041, 2364, 1008, 1439, 665, 1282, 2920, 392, 42, 805, 940, 2015, 336]. Sodium [1849, 2791, 431, 976, 1556, 1536, 911, 970, 1721, 1694, 940]. Sodium-Dodecyl [976]. Soft [149, 179, 2148, 2612, 2507, 2956, 2004, 2613, 2798]. Soft-Bodied [2004]. Soft-Shell [2148, 2612, 2613]. Soft-Water [2956]. Softwater [2801]. Soil [1583, 2354]. Soils [424]. Solar [755, 2821, 151]. Sole [370, 1492, 2262, 1712, 1858, 819, 517, 2242, 2617, 187, 1761, 2121]. Solid [2205]. Solitude [185]. Solubility [2986]. Solubilization [1694]. Solubilizing [546]. Soluble [9, 1895, 2046, 1768]. Solution [1536, 898, 1797]. Solutions [2451, 798]. Solvent [1152]. Somatic [2265, 1827]. Some [1477, 2024, 624, 503, 1864, 1322, 2178, 1005, 2939, 1342, 1865, 2322, 2298, 2163, 884, 602, 471, 2898, 808, 26, 1132, 974, 664, 1568, 1727, 121, 1307, 526, 988, 1374, 1111, 806, 1064, 357, 2383, 362, 1860, 1677, 94, 2115, 1999, 2036, 791, 2093, 2293, 1991, 802, 1319, 970, 2888, 410, 539, 2710, 760, 1103, 2762, 1014, 1039, 2140, 2918, 432, 569, 1382, 1978, 1036, 973, 766, 65, 625, 120, 2473, 813, 2747, 314, 135, 354, 846, 990, 1430, 2446, 897, 573, 145, 150, 2685, 2686, 2900, 2901, 2902, 1968, 1127, 214, 283, 324, 1601, 1783]. some [2402]. Sonar [784]. Sonic [77, 131]. Sorption [2554, 2842, 2986]. Sound [1712, 2159, 782, 187, 49, 1903, 2308, 467, 1858, 1148, 1447, 1667, 714, 492, 1016, 69, 2912, 1380, 2033, 1303, 1670, 2600]. Sound-Scattering [782]. Source [2229, 2493, 1654, 2885]. Sources [604, 1423, 2784]. South [2089, 2970, 1913, 2331, 2632, 1186, 1626, 964]. South-Central [2632]. Southeast [488, 1937, 1327, 1908, 1933, 2443]. Southeastern [2548, 2024, 116, 1315, 943, 1968, 2123, 1045, 235]. Southern [92, 1155, 648, 437, 2102, 1693, 415, 728, 1563, 1515, 2662, 1398, 1689, 166, 1645, 2041]. Southwest [850, 2628, 1761, 2487]. Southwestern [2519, 347]. Soybean [1428, 1558]. sp [510, 1271, 1486, 492, 1250, 491, 496, 2031, 514, 1365, 486, 2769, 2481, 219, 2230]. Spacing [413]. Spanish [1212, 1205]. Spärck [1530]. Spartina [1519]. Spat [78]. Spatial [1658, 574, 1290, 2644, 168, 1981]. Spawn [729, 365]. Spawning [1532, 1590, 207, 1940, 110, 67, 1518, 2062, 2907, 1802, 612, 1010, 2045, 1045, 2459, 2718, 170, 1708, 133, 1515, 917, 2523, 1014, 235, 347, 254, 1587, 1246, 42, 2475, 250, 999]. Special [1638, 273, 1209, 1559, 1993, 647, 409, 1215, 2068, 2814]. Specialization [810]. Speciation [2986]. Species [56, 1193, 376, 921, 495, 1316, 2934, 220, 628, 2921, 2179, 270, 702, 1348, 1774, 1038, 2415, 512, 1662, 949, 1391, 2365, 126, 14, 502, 695, 844, 313, 1310, 1916, 261, 521, 1741, 2112, 492, 1450, 1237, 982, 2738, 2013, 489, 255, 1789, 1163, 2426, 660, 2915, 459, 538, 265, 1318, 1190, 514, 1894, 2191, 2418, 1365, 2177, 2372, 2918, 1924, 1798, 2609, 2068, 967, 1530, 1550, 2120, 830, 1462, 1036, 1191, 2220, 252, 156, 889, 1277, 1145, 1346, 804, 803, 1783, 897, 2034, 573, 484, 2311]. Specific [1468, 2636, 2734, 732, 590, 2266, 2732, 1383, 2486, 1041, 2913]. Specificity [758, 2321]. Specimen [355, 2003]. Specimens [1365, 2495, 19, 55, 1098]. Spectra [565, 2098]. Spectral [2790, 2088]. Spectrometric [318]. Spectrometry [506, 2171]. Spectrophotometric [1614, 1429]. Spectrophotometry [84]. Spectroradiometer [908]. Spectroscopy [37, 2921, 1027, 2222]. Spectrum [2095, 79]. Speed [111, 912, 606, 1133, 119, 1511, 2538]. Speeds [920, 2187, 1463]. Spent [745]. Sperm [2045, 371, 558]. Spermatozoa [564, 1945]. Sphaeriidae [2029]. Sphaeromatidae [2720]. Sphyrion [33, 1117]. Spider [1605, 2027, 181, 636]. Spill [2627, 2604, 2598, 2625, 2599, 2615, 2626, 2624, 2622]. Spill-Contaminated [2598]. Spills [2608, 2630, 2597]. spilopterus [2062, 2366]. Spine [2132]. Spines [283]. spiniger [561]. Spiny [2633, 1589, 793, 2114, 2440, 2356, 834, 2132, 2127, 2853]. Spiny-Rayed [2132]. Spionid [493]. Splake [954, 1280, 1423, 118, 1283, 584, 2105]. Spontaneous [944]. Spore [1445]. Spores [1136, 1421, 541, 2772]. Sporozoa [1421]. Sport [2408, 316]. Sporulation [154]. Spot [2455, 630, 2476]. Spotfin [2062, 2366]. Spots [930]. Spottail [1039]. Spotted [2903, 2362]. spp [2378, 1725, 2543, 2065, 102, 274, 1948, 364]. Spring [2509, 801, 2065, 2415, 1044, 1413, 415, 1563, 347, 281, 1273, 589]. Spring- [347]. Spring-Fed [2065, 1044]. Spruce [1968]. Squalus [2633, 1589, 793, 740, 58, 2114, 2171, 2440, 2356, 834, 1512, 2043, 2127, 2853]. Squawfish [1313]. Squid [534, 359, 212, 1641]. St [958, 2437, 2061, 2240, 2226, 1638, 2780, 663, 1834, 437, 2123, 2436, 2419, 1693, 1993, 895, 415, 728, 1563, 1515, 1830, 229, 1540, 1903, 2120, 1564, 2788, 1982, 2690, 1399, 998, 2001, 2858, 2041, 1267]. St-Laurent [1267]. St. [2743, 2437, 2693, 1295, 1590, 1267, 1345, 2562, 2840, 2042, 1391, 2444, 2035, 2977, 2835, 2436, 786, 2938, 411, 1779, 2145, 1346, 1665]. Staaland [2649]. stabilité [2950]. Stability [881, 2543, 713, 2271, 1673, 2581, 2336, 2950]. Stable [2789, 531, 6]. Stage [733, 2111, 442, 1001]. Stages [10, 2060, 137, 1926, 2187, 1434, 807, 2115, 1960, 2530, 1402, 2338, 1611, 1028, 2485, 229, 630, 2772, 2866]. Staghorn [1398, 577]. Stain [2004]. Staining [2772]. Stake [874]. Stamina [912, 707, 1577, 2906, 2049, 2318]. Standard [1850, 351, 1159]. Standing [2936, 199, 2470, 2471, 2533, 217]. Stannius [772]. Starch [734]. Starfish [1149]. Starks [1670]. Starry [2121]. Starvation [528, 529, 853, 634, 641]. Starved [1799]. State [1785, 1457, 2729, 757]. States [1911, 1225, 104]. Static [1520]. Station [614, 1529, 1666, 2764, 2088, 1652, 1389]. Stations [2644]. Statistical [2330, 2028, 806, 2218, 20, 419, 1797, 1118, 1575]. Statistics [1638, 67, 712, 2351, 2185]. Status [1684, 342, 273, 2457, 2720, 1157, 1937, 1509, 1630, 1224, 1404, 2745, 2746, 1629, 1949]. Stave [133]. Steam [978]. Steelhead [1925, 1592, 1920, 574, 2503, 1688, 1576, 878, 1415, 2545, 1501, 558, 2837, 2181, 2504, 2773, 2318, 846, 990, 1273, 2634, 589]. stellatus [2121]. Stem [2406]. Stenella [1632, 1631]. Stenobrachius [148]. Stenodus [929, 2182, 1535]. stenolepis [862, 861, 860, 1132, 1055]. Sterilization [896]. Sterilizers [1489]. Sternal [283]. Sternoptyx [83]. Sterols [157, 63, 518, 627]. Stevens [2113]. Sthenelais [486]. Stichaeidae [2048]. Stickleback [1040, 1379, 821, 96, 79, 2359, 263]. Sticklebacks [1049, 402, 413]. Still [2559]. Stimpson [2073]. Stimulant [1435]. Stimulation [88, 1362]. Stimulatory [899]. Stimuli [981, 1448]. Stimulus [1396]. Stipe [380]. Stizostedion [2378, 2377, 2401, 2387, 1713, 2391, 2411, 2400, 414, 1006, 1299, 2408, 437, 1915, 2413, 2310, 989, 2393, 2390, 2394, 2419, 1714, 736, 712, 2074, 2412, 2424, 1960, 2399, 2382, 2404, 2388, 2406, 2385, 2403, 2379, 466, 1861, 2420, 2447, 2430, 2409, 1065, 2067, 2395, 99, 2106]. Stochastic [2339, 325]. Stock [1280, 1203, 1172, 2235, 2936, 1433, 1693, 2185, 985, 217, 2539]. Stocking [2411]. Stocks [1554, 1820, 2881, 1237, 2553, 1415, 199, 728, 1175, 1144, 2518, 1366, 1059]. Stomach [15, 337, 288, 2714]. Stomachs [2211, 2310, 515]. stomias [1444]. Stonecat [19]. Stoneflies [1975]. Stoneroller [969]. Storage [2371, 2580, 270, 881, 2543, 1607, 713, 9, 1384, 2119, 2483, 1673, 180, 311, 838, 2581, 1659, 2197, 1652, 1694, 1286]. Stored [2661, 190, 1309, 2272, 2483, 1902, 730, 1325, 1691, 1747, 1746]. Storm [2103, 2102]. Storms [2855]. stouti [108]. Stowage [2584]. Strain [1925, 512, 1474, 1575]. Strains [2792, 1901, 2802, 2214, 2311]. Strait [520, 2557, 2094, 2091, 2274, 890, 2090, 2249, 1966, 793, 988, 2341, 473, 2099, 2356, 98, 447, 332, 147, 201, 950, 588, 1144, 2854]. Straits [2604]. Stranded [2713, 2603, 2957]. Stranding [1636, 2454, 1627]. Strategic [2964, 2591, 2651, 1461, 2668, 2965]. Strategies [894, 2732, 2807, 1717]. Strategists [2899]. Strategy [2651, 1178]. Strathcona [2159]. Stratification [1858]. Stratified [2813, 1091, 1517, 836]. Stream [501, 1283, 2050, 2660, 2857, 722, 1594, 2025, 2509, 1153, 2065, 1135, 1662, 264, 780, 144, 2059, 2144, 2208, 124, 41, 223, 1025, 1615, 1387, 2014, 2506, 2147, 2531, 1981, 1481, 2087]. Stream-Incubating [501]. Stream-Resident [2509]. Streams [350, 2541, 574, 607, 1261, 1334, 2852, 394, 2762, 1821, 2558, 2209, 2246, 2961, 1829]. Strength [1713, 2422, 1694, 2908]. Strengths [2975]. Strenuous [1839]. Stress [2096, 2735, 2582, 2659, 2450, 2906, 859, 995, 2441, 2679, 2308, 2563, 665, 794, 343, 97, 554, 846, 990, 1273, 2139, 2844]. Stressed [2911, 1551]. Stresses [2621, 2079]. Stressful [2244, 2450]. Stressors [1777]. Striped [2914, 544, 618, 1729, 1632, 1448, 2476, 2230, 2766]. Stripping [1086]. Stromberg [2720]. Strongylocentrotus [2261, 2233, 2892]. Structural [1189]. Structure [2392, 2377, 1370, 1005, 2387, 1446, 1711, 1269, 2174, 2018, 102, 709, 2427, 2432, 255, 1505, 1701, 1103, 278, 2510, 2480, 1389, 2332, 1753, 2200, 2953]. Structured [1927]. Studies [2182, 150, 163, 1792, 1527, 246, 2574, 2604, 1904, 2100, 593, 1801, 1738, 47, 507, 2282, 1111, 2561, 1972, 1091, 2037, 2835, 157, 295, 499, 2022, 286, 2264, 1957, 626, 1602, 2739, 1357, 2524, 2806, 2278, 1746, 1355, 2001, 1580, 2232]. Study [159, 1204, 527, 2326, 1258, 894, 2541, 2838, 2263, 2089, 1289, 2625, 597, 1223, 1537, 2943, 2979, 1894, 1061, 1762, 1382, 654, 2953, 573, 1703, 1253, 857]. Studying [2761, 578, 899, 2213]. Sturgeon [152, 949, 572, 2090]. Subadult [2379]. Subalpine [312, 1786, 697, 2749, 1615, 2367]. Subarctic [1755, 2086]. Subcellular [1656, 2570]. Subcutaneous [327, 721]. Subgenera [813]. Subject [1825]. Subjected [1392, 1842]. Sublethal [38, 2072, 2785, 1914, 919, 764, 974, 1770, 1140, 568, 1777, 2771, 1501, 805, 940, 990, 442]. Sublethally [2357]. Sublittoral [2932]. Submarine [1285, 23]. Submersible [57, 908]. Submersion [47]. Submissive [770]. Subsequent [853, 2580, 1384, 363]. Subspecies [2225, 196]. Substances [2820, 161, 2038]. Substrate [1052, 1859, 394, 1129, 939]. Substrates [797, 567]. Subsurface [2090, 1493]. Subtidal [2323]. Subunits [177]. Succession [1454, 2423, 2179, 2764, 679, 1784]. Successive [341]. Sucker [527, 960, 2225, 722, 2543, 1710, 2234, 2394, 1825, 1835, 224, 470, 133, 2320, 947]. Suckers [1251, 2659, 2176, 2002, 153, 2204, 968]. suckleyi [58]. Sucrose [1096]. Suction [2323]. Sudbury [2267]. Sugar [2244]. Suitable [1296, 1264, 619]. Sulfate [976]. Sulfate-Polyacrylamide [976]. Sulfide [720, 761, 2477, 537]. Sulfite [2514]. Sulfobromophthalein [1062]. Sulfonamide [1255]. Sulfur [2546]. Sulfuric [1733]. Sulphite [745]. Sulphonate [767, 362]. Sumithion [1487]. Summary [1343, 2656, 2422, 2750, 1243, 2812, 2621, 2608, 2630]. Summer [1966, 1666, 838, 1415, 2470, 2788, 2690]. Summerkill [1655]. Sunfish [657, 997, 313, 356, 1003, 2519, 1459, 328, 92]. Sunfishes [1160, 1520, 1430]. Sunlight [2683]. superba [2963]. Superchill [1691]. Superchilled [236, 453]. Superimposed [433]. Superior [1911, 1803, 2942, 1343, 2301, 17, 378, 2718, 683, 280, 2296, 1262, 1562, 2730, 2395]. Supernatant [1006, 1299]. Supernumerary [446]. Supersaturated [2133, 1975, 2545]. Supersaturation [1078, 1501]. Supersede [113]. Supplemental [2271, 2577, 2274]. Supplementary [1624]. Supplementation [1558]. Supply [1489, 1938, 226]. Suppression [2357]. Surface [501, 307, 2092, 1529, 509, 1834, 1291, 1290, 2199, 871, 2206, 80, 1830, 792, 217, 1779]. Surfaces [1864]. Surfacing [1386]. Surficial [2296, 1881, 1880, 1882]. Surgery [363, 1122]. Surges [2103, 2102, 524]. Surplus [820, 1015, 2539]. Surplus-Production [820]. Survey [132, 296, 2562, 2813, 2680, 757, 2888, 1019, 383, 198]. Surveying [57]. Surveys [1296, 2073, 1797]. Survival [954, 1423, 733, 2142, 1251, 831, 853, 1799, 1613, 2914, 640, 659, 2060, 2275, 2272, 1570, 26, 1662, 2289, 1010, 707, 729, 2526, 1902, 1566, 1414, 1653, 911, 425, 1042, 2673, 539, 403, 727, 1923, 2320, 366, 2714, 1969, 1975, 2545, 315, 2728, 2361, 1025, 2181, 1144, 1037, 2563, 2774, 2204, 2236, 2642, 1812, 1579, 1778]. Susan [463]. Susceptibility [526, 2912, 1974, 2802, 2311]. Suspected [1006].

Suspended [2562, 2293, 2753]. Suspensions [1037]. Sustainable [1433, 1818, 2697, 2983, 2984]. Sustained [1142, 1304]. Sustaining [124]. Swimbladder [732, 2736, 583, 1374, 1403, 5]. Swimming [1454, 920, 1839, 111, 2187, 606, 1609, 1450, 891, 571, 1463, 1467, 2049, 2476, 1377, 2124, 2891, 2214, 2318, 2491, 1511, 2538, 940, 260]. Swordfish [1053, 274]. Syllidae [495]. Syllides [495]. Symbionts [321]. Sympatric [242, 605]. Symposium [2072, 1871, 2433, 2374, 1772, 1863, 671, 2865, 1818, 2868, 2596]. Syndrome [1156]. Synonymy [667]. Synthesis [247, 778]. Synthetic [1951, 532, 2451]. Syringe [1381, 1086]. System [1187, 2911, 1379, 1744, 1851, 2272, 1573, 2380, 2212, 192, 908, 2540, 1186, 1350, 1373, 1961, 1922, 1368, 1586, 91, 1800, 873, 718, 1692, 392, 2370, 2824, 2972]. Systematic [1509]. Systematics [674, 2375, 2517, 412, 1047]. Systems [1738, 77, 1727, 2484, 2974, 1935, 2078, 2432, 2770, 2699, 1436, 1762, 1836, 2986, 177].

T [2914]. T. [246]. tête [2585]. Tables [661]. Tadpole [1538]. Tag [640, 1921, 2022, 131, 2455, 619, 2223, 1778]. Tagged [585, 1996, 112, 1761]. Tagging [2522, 345, 2223]. Tags [37, 1961, 645, 2455, 33, 2223, 2488]. Tahoe [2188]. Tail [282]. Tailbeats [2124]. Taint [1837]. Tainted [978, 2962]. Tainting [2304]. Taiwan [2653]. Taken [2662, 1539, 1645]. Tank [445]. Tanker [471]. Tanks [1489, 1077]. Tanner [2681, 631]. tanneri [2681]. tantilla [1696]. Tanypleurus [1323]. Tar [234]. tarantellus [2230]. Target [2975, 2908]. Tasmanian [1382]. Taste [773, 2962, 970, 218]. Tautog [925, 579]. Tautoga [925]. Tautogolabrus [2747]. Taxonomic [1006, 342, 273, 2114, 1157, 380, 1404, 2745]. Taxonomical [32]. Taxonomy [2243, 758, 2036, 843, 2808, 967, 2953, 511]. TCA [864]. Technical [1180, 1188, 1202, 1207, 1213, 1220, 1227, 1233, 1238, 1240, 2703]. Technique [350, 978, 578, 2522, 2037, 1502, 474, 598, 1595, 2450, 2822, 2972, 2004, 754, 211, 448]. Techniques [761, 1644, 2028, 785, 2649, 440, 2635, 1948, 1762, 1955, 2124, 2355, 1567]. Technologically [1723]. teguments [2867]. Teil [2922]. Telemetry [1996, 1577, 870, 2124]. Teleost [1278, 658, 248, 381, 458, 2329]. Teleostei [245]. Teleosts [959, 918]. Tellinidae [1067]. Temperate [2544, 832, 2710]. Temperature [761, 2248, 1724, 370, 399, 1864, 2512, 34, 143, 1071, 2667, 197, 920, 1887, 2914, 1809, 743, 2201, 2060, 987, 614, 1529, 2281, 2279, 2840, 1012, 1731, 2221, 228, 1840, 2095, 1052, 1780, 2952, 2363, 744, 1291, 2282, 2187, 1259, 2365, 606, 2503, 2731, 2265, 1310, 1507, 2000, 1598, 2266, 2383, 1688, 1046, 1003, 736, 981, 2422, 1552, 1566, 2727, 142, 925, 2009, 2270, 1693, 402, 938, 2927, 193, 535, 1054, 2283, 403, 727, 1987, 2052, 1691, 995, 1463, 1465, 1459, 1031, 2038, 2170, 2918, 951, 1451, 2280, 2278, 2946, 2476, 2504, 298, 718, 1041, 2815, 1389]. Temperature [2332, 2959, 634, 1942, 1687, 48, 2012, 794, 2725, 2640, 1424]. Temperature-Induced [1507]. Temperatures [1121, 1565, 2201, 2415, 1848, 2266, 1046, 2727, 180, 1923, 2320, 1500, 1383, 1841, 18, 798, 343]. Temperautre [503]. Temporal [1290, 1981]. Tendency [1789]. Teratogenicity [801]. Terebellidae [490]. Terebratalia [518]. tergisus [1555, 2077, 1281]. Term [2580, 955, 2608, 659, 2598, 2840, 887, 2247, 2039, 2442, 2615, 2022, 2630, 13, 210, 366, 2143, 2338, 1222, 2626, 1689, 1660, 2088, 1392, 1842, 2564, 619, 974, 2950]. terme [2950]. Terminology [1702, 1978]. Terms [2793]. Terra [146, 1335]. Terranova [2154]. Terrestrial [1610]. Terrigenous [1830, 2938]. Territorial [1387]. Territories [821, 1559, 1741, 1740, 2112, 1338, 1413, 2068, 2202, 2203]. Territory [2113, 2228, 2934, 1532]. Tessellated [735]. Test [1469, 1965, 27, 2820, 494, 2609]. Tested [167]. Testes [225, 508]. Testing [1985, 2916, 2728]. Testosterone [2645, 2803, 1003]. Tests [2809, 2164, 1317, 2338, 2478, 33, 1778]. Tetraonchus [2053]. Tetrazolium [423, 1024]. TFM [1674, 2208, 1028]. Thailand [1217]. Thames [2519]. Thaw [2371, 861, 970]. Thawed [222, 858, 2483]. Thawing [713]. Thecate [175]. Their [56, 1850, 2113, 1964, 1309, 2554, 2380, 1240, 1962, 1115, 1234, 9, 2429, 2044, 735, 1415, 2918, 1206, 2346, 1182, 648, 1903, 1908, 2167, 1022, 2887, 2297, 1812, 159, 246, 1518, 323, 1559, 313, 2396, 1523, 549, 1641, 2969, 324, 1112, 1357, 1800, 1144, 1355, 1601]. Theoretical [2069, 2037, 2260, 2538]. Theory [953, 1174, 2157, 1485, 2259, 2427, 41, 2870]. Therapeutic [2237]. There [2775, 2804, 2464]. Therefrom [1907]. Thermal [1862, 1005, 1009, 1010, 923, 1425, 1666, 1825, 1960, 1902, 763, 2670, 2644, 2250, 2815, 281, 665]. Thermocline [1091, 910]. Thermoselection [1757]. Thin [930]. Thin-Layer [930]. Thiobarbituric [860]. Thiourea [1508]. Third [1816, 1085]. thomasi [1793]. thompsoni [621]. Those [55]. Three [221, 2824, 605, 628, 1494, 1314, 105, 2931, 502, 1310, 2039, 261, 1581, 661, 660, 15, 459, 2143, 1500, 2438, 2372, 1685, 967, 2120, 2802, 2167, 2147, 2220, 156, 2614, 365, 1145, 2913, 484]. Three-Dimensional [1494, 1685]. Three-Way [661]. Three-Year [105]. Threespine [1379, 96, 1049, 79, 413]. Threshold [1805]. Threshold-Induced [1805]. Throughout [1161]. Thunersee [682]. Thunnus [423, 1024]. Thyasira [2031]. Thyasiridae [2031]. Thymallus [1594, 372, 2053]. Thymus [1432]. thynnus [423, 1024]. Thyroid [1388, 1471, 2018, 2348, 336]. Thyroxine [2018, 2270, 2704, 260, 1079]. Thysanoëssa [56, 2240, 2199]. Tidal [192, 1549, 2091, 2932]. Tide [320, 405, 1070, 2232]. Tide-Pool [405]. Tidepool [320]. Tides [2100, 2101, 2145]. Tidewater [1709]. tigrina [271, 446]. Tilapia [143, 246, 82, 474, 2132, 2220]. Time [2512, 322, 2914, 185, 203, 105, 2941, 784, 180, 878, 1520, 917, 2088, 641, 2475, 1430]. Time-Growth [1520]. Time-Lapse [784]. Times [2709, 594, 1990]. Timing [756, 2830, 26, 1921, 2063]. Tipula [2065]. Tissue [1528, 416, 2168, 787, 127, 110, 129, 713, 2541, 137, 2784, 363, 647, 730, 1561, 1100, 2570, 1421, 1458, 336]. Tissues [444, 788, 1612, 2645, 121, 546, 2205, 922, 993, 89]. Titration [2460]. Tocopherol [1612, 2581]. Tolerance [1864, 228, 1840, 1902, 2227, 2258].

Tomales [422]. Tome [2867]. tonsa [2716]. Tool [2536]. Topics [2466]. Topminnow [969]. Topographic [2099]. Törnquist [1277]. Toronto [221, 220, 314]. Torrent [479]. Torrey [2622]. Total [1077, 860, 1914, 1060, 1518, 1053, 928, 757, 1372]. Toxaphene [1582, 1584]. Toxic [2643, 2289, 2442, 937, 1651, 2722, 2038, 822, 2622, 1619, 2125]. Toxicant [1573, 2144]. Toxicants [1850, 1556, 1987, 2079]. Toxicities [1968, 1434]. Toxicity [761, 1849, 547, 1328, 2477, 2163, 40, 2809, 787, 1099, 801, 162, 759, 2365, 1972, 1143, 2849, 1507, 1793, 1807, 2916, 2848, 938, 1674, 2820, 1606, 1576, 2507, 1317, 2143, 2338, 1987, 1371, 1500, 1438, 1327, 1729, 1611, 1028, 1321, 2245, 726, 1798, 2946, 1452, 2589, 1984, 2066, 2887, 2634, 2847, 2304, 745, 796, 2034, 1742, 2586, 2032, 3]. Toxicological [2716]. Toxicology [2135]. Toxicosis [856]. Toxin [431]. Toxins [2763, 2494, 2572]. Trace [2159, 2562, 1965, 1027, 72, 2238, 1534, 1287]. Tracking [785, 1526, 1425, 2074, 2022, 440, 1081, 710, 75, 1063, 870, 1377, 231]. Tracks [884]. Tract [2713, 631]. Trade [2890]. Trade-Off [2890]. Trading [2741]. Traditionally [1190]. Trail [2068]. Training [1223]. Transaminase [206]. Transatlantic [872, 2127]. Transcarbamylase [1749]. Transdisciplinary [1460]. Transect [2557, 285]. Transects [1864]. Transennella [1696]. Transfer [2113, 2546, 2776, 2861, 804]. Transferrin [182, 1361, 2478, 2167, 267]. Transferring [1837]. Transfers [365]. Transformation [2971, 2515, 1826, 1977, 114, 2432, 589]. Transformations [2857, 2676]. Transformed [748]. Transition [1192, 74]. Transmission [1971, 2647, 2909, 2912]. Transmitter [1081]. Transmitters [1996, 2907, 1577]. Transparency [2724]. Transplant [1776]. Transplanted [2142]. Transport [2249, 1891, 1982]. Transported [2762]. Transporting [2708]. Transports [1890, 2094, 1879]. transversa [518]. Trap [1605]. Trapnet [956]. Trapped [747]. Trappers [894]. Traps [2490, 2943, 2073, 1796]. Travelled [41]. Trawl [1801, 1922, 1973, 1970]. Trawler [881]. Trawler-Caught [881]. Trawling [2200]. Trawls [884, 257, 2662]. Tray [655]. Treated [1525, 1971, 2906, 2321, 615]. Treatment [2349, 2160, 2237, 748, 1003, 1673, 1508, 1023, 171, 1273, 2634, 1523]. Treatments [2371, 2698, 554]. Trematoda [274, 2053]. Trematode [152, 1574, 1564]. Trench [489]. Trends [189, 1170, 2403, 2410]. Tri [1561]. Tri-aryl [1561]. triacanthus [476]. Triaenophorus [35]. Trials [1258, 716]. Tributaries [1590, 1044]. Tributary [1803, 748, 1886, 1266]. Tricaine [362, 767]. Trichogaster [277]. Trichoptera [1262, 2551]. trichopterus [277]. Triclad [921, 271, 932]. Trifluoromethyl [1147, 1732, 931, 1674, 2208]. Triglyceride [416]. Triglycerides [1568]. Trigonelline [64]. Triiodothyronine [2704]. Trimethyl [420]. Trimethylamine [1051, 1309, 311]. Trimethylamine-Producing [311]. Triphosphatase [1826]. Triphosphate [1583, 1095]. Triploid [1127]. Tripolyphosphate [970]. Trisodium [1270, 738]. Tristoma [274]. Tritiated [2357]. triventricosus [219]. trivolvis [1270]. Trophic [1527, 1684, 1016, 2894].

Trophogenic [2676]. Tropical [1813, 212, 1631]. Trout [2180, 1265, 1406, 38, 164, 1279, 2851, 461, 2588, 1795, 559, 604, 1899, 2985, 2326, 93, 416, 788, 1097, 2660, 1656, 2735, 1925, 2785, 327, 810, 2168, 1971, 787, 1785, 1592, 1593, 123, 640, 1409, 1428, 230, 1904, 1920, 2160, 2706, 203, 509, 865, 1271, 351, 67, 1839, 974, 2131, 2767, 1980, 574, 1035, 2838, 1135, 2641, 162, 1662, 2011, 428, 584, 2047, 2571, 1473, 111, 460, 526, 1700, 2859, 892, 1822, 1259, 2931, 856, 2691, 607, 2503, 2016, 2930, 566, 2657, 2129, 1143, 2849, 2457, 2271, 2577, 2118, 926, 2000, 2578, 1010, 2266, 2039, 1688, 1695, 1513]. Trout [707, 1045, 362, 363, 767, 1122, 2803, 1261, 1403, 773, 1048, 1837, 1943, 264, 417, 331, 1502, 2112, 2459, 2390, 546, 397, 2861, 2909, 391, 2738, 2107, 1990, 780, 109, 378, 1552, 1566, 2727, 1537, 186, 739, 1353, 1807, 1651, 2270, 1020, 1586, 1991, 286, 1595, 2850, 85, 329, 367, 849, 1859, 938, 1410, 1367, 124, 13, 2673, 1986, 1432, 1576, 193, 535, 75, 727, 878, 1415, 210, 366, 1317, 2143, 2722, 1608, 1591, 1584, 1833, 2955, 1969, 2545, 2771, 1501, 1014, 223, 1018, 1080, 1821, 944, 1100, 1028, 1061, 2570, 2368, 1974, 558, 1421, 1901, 1781, 1031]. Trout [1467, 2769, 1480, 2049, 2046, 1025, 2837, 1451, 1275, 2346, 2478, 2181, 2729, 2818, 434, 1151, 1129, 2946, 1558, 1452, 340, 280, 1998, 2504, 1474, 51, 1062, 2773, 1828, 654, 1394, 1387, 289, 615, 1491, 2105, 2815, 1569, 2357, 1787, 5, 12, 1587, 2214, 2331, 1287, 2258, 2928, 1022, 2318, 2063, 1687, 964, 993, 548, 737, 986, 2725, 97, 135, 554, 846, 990, 1273, 1575, 2139, 2634, 2847, 442, 2304, 177, 952, 1742, 1743, 2032, 1579, 104, 2696, 1484, 2658, 589, 1150, 1372, 2122, 2175, 2324, 2155, 1751]. Trout-Perch [1271, 1045]. Trouts [412]. truncatus [1642, 1630, 1017]. truquii [281]. trutta [85, 124, 2955, 1781, 1480, 2818, 1151, 2331]. Trypanosoma [2453]. Trypsin [1123]. tshawytscha [2512, 2940, 862, 861, 860, 1156, 801, 2952, 2907, 2503, 1688, 560, 144, 975, 837, 2563, 1273, 2311]. Tube [490]. Tube-Building [490]. Tubifex [2769]. Tubificid [221, 220, 1320, 2857, 961, 2769, 314]. Tubular [1300]. tubulosa [1033]. Tubulovesicula [152]. Tumors [1712, 459]. Tuna [1964, 184, 423, 1024, 231]. Tunas [1237, 409]. Tunnel [900, 707]. Tunnel-Digging [900]. Turbellaria [271, 446]. Turbellarian [921, 2695]. Turbidity [2724, 2161]. Turbines [779, 1078]. Turbulence [232, 2125]. Tursiops [1642, 1630, 1017]. Turtles [1260]. Twelve [809]. Twenty [287]. Twilight [2786]. Two [56, 1298, 570, 159, 31, 945, 2142, 603, 242, 375, 1120, 1449, 2660, 449, 1809, 2921, 16, 2842, 762, 2380, 77, 2757, 2235, 512, 574, 1391, 2792, 1378, 165, 566, 844, 1403, 52, 60, 645, 2073, 2553, 1044, 1500, 2644, 2819, 405, 1852, 278, 2227, 1767, 498, 1112, 2046, 2560, 2213, 1059, 771, 1023, 1095, 1530, 1550, 2321, 304, 1093, 1984, 281, 625, 2117, 2147, 2597, 1022, 2814, 804, 2236, 803, 1253]. Type [431, 1649, 901, 2825, 1107, 2495, 2609, 541, 2091, 1639]. tyrannus [1581, 332, 1324].

Uca [2615, 794]. Udonella [1141]. Udonellidae [1141]. Ultrasonic [785, 1526, 1425, 2074, 440, 1081, 710, 75, 1063, 870, 1377, 2124, 2355, 231].

Ultrasound [35]. Ultrastructural [2524, 321, 2696]. Ultrastructurally [2888]. Ultrastructure [2340, 2232]. Ultraviolet [2349, 107]. Un-Ionized [798]. Unanesthetized [1543]. Uncertain [1717]. Unconditioned [1396]. Undergoing [2397]. Understanding [1347]. Underwater [884, 2355, 2692]. Underyearling [2727, 36, 2531, 805, 940]. Underyearlings [144, 2364]. Undesirable [2890]. Undifferentiated [289]. Unexpected [1076]. Unexploited [2112, 712, 2396, 1124, 1177]. Unfamiliar [1193]. Unfrozen [858]. Ungava [638, 905, 2130]. Ungutted [2661]. Unialgal [1917]. Unicellular [2268, 1310]. unicornis [201]. Unionid [2525]. Unionidae [806, 1657]. Unit [715]. United [104, 1911, 1225]. Units [23]. Univalent [374]. University [1703, 1474]. Unoiled [2623]. Unperturbed [2611, 2428]. Unstressed [680]. Untreated [2906, 2236, 2642]. Unusual [2521, 821, 534]. Upon [664, 2616, 1182, 1598, 362, 1641, 2967, 260]. Upper [1841, 343, 358, 947, 1540, 412]. Upstream [275, 276]. Uptake [1130, 1764, 774, 1090, 2337, 2641, 2148, 1716, 1710, 2193, 2799, 2754, 1806, 995, 1100, 1061, 2570, 2609, 1959, 1657, 2613, 964, 794]. Upwelling [1832, 2768, 755, 1692]. Urchin [2261, 2892]. Urchins [1983, 2456, 2233, 666]. Uric [1401]. urogénital [2867]. Urohyal [1544]. Urosalpinx [2611]. ursinus [1514, 1788]. Uruguayan [1633]. USA [496, 472]. Use [2791, 2809, 1309, 1940, 614, 1951, 1556, 1115, 1326, 815, 2037, 1264, 1133, 838, 1032, 1415, 728, 1416, 1131, 2744, 107, 972, 836, 1092, 2679, 2715, 2622, 2108, 2755]. Used [1368, 1602, 2560]. Useful [578]. Uses [1194]. Using [1308, 2358, 2782, 1649, 2921, 864, 1258, 1735, 1035, 1160, 2816, 1605, 465, 2822, 2972, 1998, 445, 754, 2474, 1021, 2315, 306, 463, 2491, 2532, 287]. USSR [626, 2385, 2384]. utawana [2225]. Uteriporus [932]. Utility [2172, 2518]. Utilization [2251, 2554, 2842, 1603, 2333, 607, 316, 925, 144, 1199, 1698, 2049, 2020, 1359, 1096]. Utilize [2974]. Utilizing [2605].

vacca [804, 803]. Vaccination [2792]. Vaccines [1490]. Vacuum [930]. Validation [1533]. Validity [2225, 2816]. Valley [942]. Value [2024, 860, 1844, 2784, 1916]. Values [189, 2820, 348, 135, 354]. Vancomycin [1488]. Vancouver [230, 1761, 2100, 850, 840, 2014, 2506, 2760, 2274]. var [1596]. Varden [461, 116, 1315, 943, 654]. variabilité [2950]. Variability [2683, 2243, 1844, 2816, 1775, 2806, 1549, 2690, 2672, 1717, 1109, 2950]. Variable [941]. Variables [1457, 1436, 2213]. Variance [465, 1575, 1109]. Variants [101]. Variation [958, 1, 604, 1423, 1790, 1379, 2517, 2188, 583, 1962, 806, 1116, 1552, 2508, 2732, 1252, 91, 372, 1034, 2523, 1361, 2768, 103, 196, 1800, 2694, 1092, 828, 1574, 1399, 1753, 99, 2184]. Variations [2024, 868, 1809, 208, 1291, 363, 1364, 335, 1689, 2379]. Varied [2411, 2170]. Various [831, 862, 628, 10, 1565, 137, 191, 1554, 546, 1775, 1327, 2759, 2470, 1031, 2038, 2361, 2385, 229, 817, 1490, 1989]. Varying [956, 123, 2733, 1020, 2890]. Vascular [381]. Vascularization [578]. Vättern [687]. Vector [2494]. Vegetative [541]. Velella [507]. Velocities [1467, 2147]. Velocity [2555, 1292, 1052, 142, 1538, 2476]. VEN [2912].

Ventilation [184, 58]. Ventilatory [509]. Ventures [1216, 1810]. Vermont [2434, 1718, 2191, 2751]. Vers [910]. Versatile [2896, 325]. Version [155]. Versus [859, 1201, 2758]. Vertebral [2587, 458, 1789, 2150, 176]. Vertebrate [2911]. Vertebrates [485, 934]. Vertical [2437, 71, 2472, 2179, 2141, 1088, 11, 2970, 1091, 2727, 535, 36, 1144, 1861, 304, 655, 2950]. verticale [2950]. Very [2798]. Vessels [1473, 1922]. vetulus [1712, 1858, 2617, 2121]. vi [1271]. Vibrio [1985, 2196, 1925, 342, 2792, 2802, 814, 1572]. Vibrios [2321, 2573]. Vibriosis [1925, 342]. Vicinity [1529, 2644]. Victoria [2886]. View [2875]. villosus [2435, 2437, 2661, 2580, 2436, 2518, 1747, 1746, 2197, 45, 418]. Vincent [1638]. Viral [1953, 1997, 2647, 2912, 1974]. virens [2933, 1458]. Virginal [1625]. Virginia [1672, 1408]. virginica [2823, 1258, 1012, 1602, 2227, 78, 1312, 6, 7]. viridans [1596]. viridis [32]. virilis [2931, 2438, 2343, 2439]. Virucidal [570]. Virulence [596, 1925]. Virus [34, 163, 1406, 1570, 428, 1115, 2960, 1843, 397, 1653, 1578, 1669, 2361, 951, 2957, 2340, 586, 1311, 1579, 1521, 2839]. Viruses [150, 163, 570, 2666, 1653, 2957, 2642, 586, 1113, 2106]. Vision [1447, 2665]. Visual [2378, 2452, 2022, 122, 1448, 2152, 1021, 171, 1567]. Vital [1331, 67, 712, 2185, 2004]. Vitamin [890]. Vitellogenin [2841]. vitreum [2377, 2401, 1713, 2391, 2411, 2400, 414, 1006, 1299, 2408, 437, 1915, 2413, 2310, 989, 2393, 2390, 2394, 2419, 1714, 736, 712, 2074, 2412, 2424, 1960, 2382, 2404, 2406, 2379, 466, 1861, 2420, 2447, 2430, 2409, 1065, 2067, 2395, 99, 2106, 2377, 2401, 1713, 2391, 2411, 2400, 414, 1006, 1299, 2408, 437, 1915, 2413, 2310, 989, 2393, 2390, 2394, 2419, 1714, 736, 712, 2074, 2412, 2424, 1960, 2382, 2404, 2406, 2379, 466, 1861, 2420, 2447, 2430, 2409, 1065, 2067, 2395, 99, 2106]. Vitro [2823, 1255, 1263, 1943, 1363, 633, 150, 1061, 2585]. vitulina [373, 1298, 935]. Viviparous [804, 803]. Vivo [1255, 129, 163, 2585]. Vocalization [1105]. Vol [2797, 2795]. Volatile [2154, 1306, 730]. Volcanic [603]. Voltage [809, 462]. Volume [2546, 583, 2550, 2770, 2922, 2046]. Volumetric [1381]. vulgaris [932]. Vulnerability [1009, 1302, 566, 2738, 1042, 410, 1424].

Wabamun [1479]. Walbaum [2747]. Waldsea [2968]. Wales [2331]. Walleye [2377, 2401, 1713, 2391, 2411, 2400, 414, 1006, 1299, 2408, 1915, 2413, 2310, 989, 53, 2393, 2390, 2394, 2419, 736, 712, 2074, 2412, 2424, 1960, 2382, 2404, 2406, 119, 169, 466, 1861, 2420, 2447, 2430, 2409, 1065, 2067, 2395, 99, 2106]. Walleyes [437, 1714, 2379]. Walruses [2680]. Warm [1413]. Warming [1384]. Warmwater [2417, 576]. Warning [2079]. Washing [382]. Washington [1248, 1509, 2216, 472, 1712, 519, 1646, 2748, 323, 642, 473, 1509, 69, 255, 187, 49, 498, 1761, 1474, 2269, 463, 1282, 2920, 2121]. Waste [2652, 1131, 340, 842]. Wastes [52, 1698]. Wastewater [2163, 1914, 2198]. Watasenia [359]. Water [1235, 34, 1401, 1489, 149, 2113, 2606, 374, 1306, 2555, 1331, 1734, 2349, 1613, 2627, 2567, 722, 2705, 421, 593, 2076, 183, 2698, 1469, 1738, 1343, 785, 1429, 129, 1052, 2415, 2363, 1088, 789, 2089, 1700, 1456, 1111, 2960, 2503, 100, 1234, 2193, 222, 2970, 2119, 1598, 1688, 53, 2442, 1219, 1027, 1094, 2151, 2239, 647, 1319, 2351, 1078, 1538, 1709, 2507, 1602, 1561, 1534, 2133, 1975, 2545, 2684, 2744, 317, 1654, 2558, 2046, 1917, 1715, 1660, 2956, 2010, 2559, 2300, 2476, 2094, 2504, 2116, 817, 2514, 1262, 1747, 1879, 2474, 1086, 1355, 2357, 842, 321, 18, 1768, 1687]. Water [1572, 1726, 2860, 903, 2640, 1372, 384, 1752, 857, 1619, 1546]. Water-Sampling [421]. Water-Soluble [2046, 1768]. Waterfall [1413, 223]. Waters [353, 299, 561, 132, 2017, 1633, 2298, 1885, 175, 602, 1635, 244, 227, 2389, 437, 599, 618, 1813, 1291, 753, 1290, 557, 1431, 502, 1935, 2440, 1444, 2861, 2222, 834, 1512, 1566, 2009, 757, 551, 1919, 2652, 539, 728, 626, 1661, 420, 1112, 2449, 235, 2922, 2972, 2372, 1689, 2385, 1155, 2068, 300, 21, 2518, 1908, 1909, 2416, 2109, 217, 1779, 171, 1645, 993, 2236, 2642, 1134, 2779, 796, 2853, 2521, 1426, 1029, 1628, 2152]. Watershed [2228, 2104, 1821]. Watersheds [1886]. Wave [565, 2097]. Waves [2099, 2098]. Wax [2981]. Way [661, 1021]. Weak [966]. Weanling [533]. Weather [1529, 2088]. Weathering [2607, 2602]. Weathervane [239, 355]. Web [56, 1527]. Weight [604, 1071, 40, 1060, 2199, 1543, 2003, 848, 820, 1457, 2933, 717, 1032, 2758, 2478, 595, 1305, 649]. Weights [631, 549]. Weir [452]. Wells [138]. West [2090, 2611, 1210, 1638, 712, 2074, 2396, 1647, 2850, 2768, 2346, 2624, 2447, 1356, 1567, 839]. West-Slope [2850, 2346]. Western [2140, 312, 2437, 1121, 1713, 1455, 1299, 653, 423, 1024, 2263, 1352, 2093, 2553, 1642, 2876, 1034, 2296, 412, 1262, 2730, 1030, 2430, 2395, 55, 54, 1098, 2482, 104, 484]. Westward [2086]. Westward-Flowing [2086]. Wet [84]. Weweantic [268]. Whale [1646, 467, 1649, 1638, 1643, 2454, 551, 1641, 1627, 1624, 1625, 1626, 1567]. Whales [1126, 790, 1635, 2682, 1637, 2372, 2834, 1030, 1629]. Whaling [1229, 1635, 1639]. Wheeler [1926]. Whelping [373]. Which [733, 1421, 288]. Whirling [104]. White [741, 422, 527, 960, 1251, 2225, 152, 2713, 467, 2325, 886, 722, 1060, 1710, 2712, 2731, 2394, 1825, 1835, 224, 470, 2320, 2659, 947, 1276, 1151, 1129, 1462, 1030, 1629, 2002, 153, 2204, 968, 2736]. White-Sided [2713, 2712]. Whitebait [520]. Whitefish [946, 2313, 1449, 1525, 2824, 2934, 1532, 127, 886, 2929, 1119, 2104, 1840, 2263, 35, 1559, 1554, 2657, 1284, 1911, 638, 905, 2112, 142, 843, 1252, 288, 364, 1498, 1750, 1424]. Whiteflsh [1120]. Whiteftsh [1532]. Whiteside [2165]. Whiting [1603]. Whole [2782, 1557, 1268, 1061, 1354, 805]. Whole-Lake [1354]. Wide [960, 488]. Widely [993]. Widespread [1789]. Width [2123]. Wild [339, 1101, 1662, 2011, 2893, 2264, 1415, 2972, 2181, 1151, 2961, 2619, 1050, 2611]. Williams [2165]. williamsoni [2104, 288]. Wilson [1157, 1444]. Wind [2096, 2092, 2060, 2098, 859]. Wind-Generated [2098]. Wind-Stress [859]. Windermere [2407, 688]. Winkler [540]. Winnipeg [567, 1119, 2103]. Winter [2024, 1592, 1593, 2448, 105, 663, 2936, 2047, 1686, 871, 1364, 439, 838, 1380, 1415, 103, 268, 1158, 524, 1827, 556, 2051, 1073]. Wintering [840, 250]. Winterkill [2458, 2895, 122]. Wire [1133, 2455]. Wisconsin [2141, 2412, 1385]. Witch [2017, 2693, 1672, 229]. Within [2934, 1411, 1898, 2676, 1933, 804, 1109, 14]. Without [1319, 972]. Wolffish [2903]. Wood [1005]. Woodborer [2026]. Wooden [1286]. Woods [2611, 608, 2067, 2395, 142, 2410]. Woody [997]. Working [2607]. Workshop [2216]. World [1240, 488, 678, 1237, 2009, 2351, 1199, 85, 329, 367, 849, 2821, 2502]. World-Wide [488]. Worm [25]. Worms [2857, 2769]. Wounds [1072, 2488]. Wreck [471]. Wulker [2737]. Wulp [2177].

X [37, 506, 318, 2921, 2171, 442]. X-Ray [37, 506, 318, 2921, 2171]. X-rays [442]. xanthogrammica [2004]. xanthurus [2476]. Xeneretmus [476]. Xenobiotics [1901]. Xenomystax [563]. Xiphias [1053]. XXVI [2922].

Yarmouth [1832]. Year [610, 1713, 2411, 2445, 105, 1840, 1915, 2625, 1364, 60, 2422, 2718, 1923, 169, 2476, 2839]. Year-Class [2411, 1915, 60, 2422]. Year-Classes [1713, 169]. Yearling [1283, 468, 2131, 2571, 1471, 784, 2704, 193, 1129, 2379, 2563, 2531]. Yearly [705]. Years [2627, 287, 2599, 1676, 1085, 1817]. Yeast [2493]. Yellow [2391, 2060, 2141, 2415, 162, 386, 759, 2310, 2786, 2394, 2419, 1825, 2424, 1054, 1718, 2751, 2406, 1767, 119, 1854, 1039, 1463, 2409, 120, 2414, 839, 3]. Yellowfin [2262]. Yellowstone [2850, 75]. Yellowtail [708, 1936, 2816, 254, 330, 1475]. Yersinia [2851]. Yield [2392, 941, 1723, 2387, 1866, 1433, 1304, 2266, 1015, 1857, 2471, 1341, 1758, 2539, 1818, 2697, 2983, 2984]. Yields [2391, 2534, 1886, 1059, 2410]. Yoldia [885]. Yolk [1052]. Yolk-Sac [1052]. York [44, 757, 1915, 2413, 379, 53]. Young [1423, 395, 528, 529, 513, 1754, 1527, 2445, 1840, 100, 891, 795, 425, 1042, 1672, 727, 1923, 2476, 393, 1387, 1652, 965, 140, 2108, 803]. Young-Of-The-Year [1923, 2445, 1840, 2476]. Yukon [2113, 2228, 1532, 2934].

zaddachi [1809]. Zalembius [376]. Zalophus [1782]. Zealand [491]. Zebrafish [1611]. zillii [246]. Zinc [2903, 2276, 2785, 787, 531, 1060, 1518, 121, 1507, 1806, 2000, 1777, 1660, 393, 770, 2066, 637, 897, 7, 2184]. Zinc-65 [121]. Zinc-Intoxicated [2000]. Zn [2159, 2186, 2801, 6]. Zoarcidae [2768]. Zone [246, 2333, 2676, 2469, 2191, 2603, 524]. Zones [1810]. Zoobenthos [305]. Zoogeographic [472, 2376]. Zoogeography [2375, 2263, 245, 843]. Zooplankter [2717]. Zooplankton [159, 1690, 2159, 2643, 2275, 471, 2341, 2174, 1741, 2738, 1442, 2353, 36, 2922, 1339, 1549, 2694, 348, 304, 2715, 1562, 2474, 1551, 2315, 2953, 2968, 2814, 1294, 1346, 1896]. Zostera [1585, 199, 2255].

References

Anderson:1970:AVI

[1] Richard B. Anderson and Owen C. Fenderson. An analysis of variation of insecticide residues in landlocked Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):1–11, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-001.

McQueen:1970:GRF

[2] Donald J. McQueen. Grazing rates and food selection in Diaptomus oregonensis (Copepoda) from Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):13–20, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-002.

Zitko:1970:TYP

[3] V. Zitko, D. E. Aiken, S. N. Tibbo, K. W. T. Besch, and J. M. Anderson. Toxicity of yellow phosphorus to herring (Clupea harengus), Atlantic salmon (Salmo salar), lobster (Homarus americanus), and beach flea (Gammarus oceanicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):21–29, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-003.

Moore:1970:AEP

[4] G. S. Moore, H. A. Peters, and R. E. Levin. Alterations in the electrophoretic protein patterns of refrigerated fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):31–38, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-004.

Tait:1970:MST

[5] J. S. Tait. A method of selecting trout hybrids (Salvelinus fontinalis × S. namaycush) for ability to retain swimbladder gas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):39–45, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-005.

Wolfe:1970:LSZ

[6] Douglas A. Wolfe. Levels of stable Zn and ^{65}Zn in Crassostrea virginica from North Carolina. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1): 47–57, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-006.

Wolfe:1970:ZEC

[7] Douglas A. Wolfe. Zinc enzymes in Crassostrea virginica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):59–69, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-007.

Davies:1970:PMM

[8] Gordon S. Davies. Productivity of macrophytes in Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):71–81, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-008.

Hiltz:1970:PAS

[9] Doris Fraser Hiltz and W. J. Dyer. Principal acid-soluble nucleotides in adductor muscle of the scallop Placopecten magellanicus and their degradation during postmortem storage in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):83–92, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-009.

Buchanan:1970:EIS

[10] David V. Buchanan, Raymond E. Millemann, and Nelson E. Stewart. Effects of the insecticide Sevin on various stages of the Dungeness crab, Cancer magister. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (1):93–104, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-010.

Halliday:1970:GVD

[11] R. G. Halliday. Growth and vertical distribution of the glacier lanternfish, Benthosema glaciale, in the Northwestern Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):105–116, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-011.

Tarr:1970:DAP

[12] H. L. A. Tarr and Linda Gardner. Deoxyribonucleic acid polymerase of trout liver nuclei. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (1):117–123, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-012.

Martin:1970:LTE

[13] N. V. Martin. Long-term effects of diet on the biology of the lake trout and the fishery in Lake Opeongo, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):125–146, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-013.

Hagen:1970:SPW

[14] D. W. Hagen and J. D. McPhail. The species problem within Gasterosteus aculeatus on the Pacific Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):147–155, January 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-014.

Mawson:1970:TCP

[15] J. C. Mawson and R. J. Reed. Three computer programs: Backcalculation, condition factor, and stomach content, CDC 3600 Fortran/ Format. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):156–157, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-015.

Chen:1970:FCP

[16] T. R. Chen. Fish chromosome preparation: Air-dried displays of cultured ovarian cells in two killifishes (Fundulus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):158–161, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-016.

Khan:1970:MDL

[17] N. Y. Khan and S. U. Qadri. Morphological differences in Lake Superior lake char. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):161–167, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-017.

Thiel:1970:PCW

[18] V. Thiel and H. H. Harvey. Proportional controlling of water temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):167–170, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-018.

Stewart:1970:FSS

[19] K. W. Stewart and C. C. Lindsey. First specimens of the stonecat, Noturus flavus, from the Hudson Bay drainage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):170–172, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-019.

Lindsey:1970:ESI

[20] J. K. Lindsey. Exact statistical inferences about the parameter for an exponential growth curve following a Poisson distribution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):172–174, January 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-020.

Scott:1970:MCG

[21] W. B. Scott, A. C. Kohler, and R. E. Zurbrigg. The manefish, Caristius groenlandicus Jensen (Percomorphi: Caristiidae), in Atlantic waters off Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):174–179, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-021.

Reed:1970:ESC

[22] Dale F. Reed and Edward B. Reed. Estimates of seston crops by filtration with glass fiber discs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):

180–185, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-022.

Platt:1970:IRS

[23] T. Platt, E. Larsen, and R. Vine. Integrating radiometer: A selfcontained device for measurement of submarine light energy in absolute units. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):185–191, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-023.

Warner:1970:ILA

[24] Ronald W. Warner and Stanley C. Katkansky. An inflammatory lesion in an American shad, Alosa sapidissima. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):191–193, January 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-024.

Subramanian:1970:CRB

[25] A. Subramanian. Chloride regulation in the burrowing worm eel Moringua linearis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (1):193–196, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-025.

Dill:1970:ESE

[26] L. M. Dill and T. G. Northcote. Effects of some environmental factors on survival, condition, and timing of emergence of chum salmon fry (Oncorhynchus keta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (1):196–201, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-026.

Evans:1970:MBA

[27] John W. Evans. Marine borer activity in test boards operated in the Newfoundland area during 1967–68. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):201–203, January 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-027.

Green:1970:GER

[28] Roger H. Green. Graphical estimation of rates of mortality and growth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):204–208, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-028.

Anonymous:1970:RAFa

[29] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(1):209–211, January 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-029.

Paterson:1970:BFC

[30] C. G. Paterson and C. H. Fernando. Benthic fauna colonization of a new reservoir with particular reference to the Chironomidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):213–232, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-030.

Anderson:1970:PCL

[31] R. Stewart Anderson. Physical and chemical limnology of two mountain lakes in Banff National Park, Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):233–249, February 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-031.

Ackman:1970:HVF

[32] R. G. Ackman, R. F. Addison, S. N. Hooper, and A. Prakash. Halosphaera viridis: Fatty acid composition and taxonomical relationships. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):251–255, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-032.

Scarratt:1970:LFT

[33] D. J. Scarratt. Laboratory and field tests of modified sphyrion tags on lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):257–264, February 1970. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f70-033.

Amend:1970:CIH

[34] Donald F. Amend. Control of infectious hematopoietic necrosis virus disease by elevating the water temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):265–270, February 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-034.

Freese:1970:DTC

[35] M. Freese. Distribution of Triaenophorus crassus parasites in whitefish flesh and its significance to automatic detection of the parasites with ultrasound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):271–279, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-035.

Narver:1970:DVM

[36] David W. Narver. Diel vertical movements and feeding of underyearling sockeye salmon and the limnetic zooplankton in Babine Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):281–316, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-036.

Calaprice:1970:MAM

[37] J. R. Calaprice and F. P. Calaprice. Marking animals with micro-tags of chemical elements for identification by X-ray spectroscopy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):317–330, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-037.

Anderson:1970:ESD

[38] J. M. Anderson and H. B. Prins. Effects of sublethal DDT on a simple reflex in brook trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2): 331–334, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-038.

Cheng:1970:EGP

[39] Joseph Y. Cheng and Naval J. Antia. Enhancement by glycerol of phototrophic growth of marine planktonic algae and its significance to the ecology of glycerol pollution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):335–346, February 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-039.

Buhler:1970:IBW

[40] Donald R. Buhler and W. E. Shanks. Influence of body weight on chronic oral DDT toxicity in Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):347–358, February 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-040.

McLay:1970:TCD

[41] Colin McLay. A theory concerning the distance travelled by animals entering the drift of a stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):359–370, February 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-041.

Vanstone:1970:GCC

[42] W. E. Vanstone, J. R. Markert, D. B. Lister, and M. A. Giles. Growth and chemical composition of chum (Oncorhynchus keta) and sockeye (O. nerka) salmon fry produced in spawning channel and natural environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):371–382, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-042.

Jackim:1970:EMP

[43] Eugene Jackim, Janice M. Hamlin, and Stephen Sonis. Effects of metal poisoning on five liver enzymes in the killifish (Fundulus heteroclitus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):383–390, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-043.

Haedrich:1970:MFH

[44] Richard L. Haedrich and Michael H. Horn. The morid fish Halargyreus johnsonii from the New York Bight. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):391–393, February 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-044.

Winters:1970:RSA

[45] G. H. Winters. Record size and age of Atlantic capelin, Mallotus villosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):393–394, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-045.

Turner:1970:RGS

[46] Gary E. Turner and Hugh R. MacCrimmon. Reproduction and growth of smallmouth bass, Micropterus dolomieui, in a Precambrian lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):395–400, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-046.

Druehl:1970:SES

[47] Louis D. Druehl and John M. Green. A submersion–emersion sensor, for intertidal biological studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):401–403, February 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-047.

Umminger:1970:ETS

[48] Bruce L. Umminger. Effects of temperature on serum protein components in the killifish, Fundulus heteroclitus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):404–409, February 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-048.

Millikan:1970:HPH

[49] Alan E. Millikan and Bradley H. Pattie. Hermaphroditism in a Pacific hake, Merluccius productus, from Puget Sound, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):409–410, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-049.

Anonymous:1970:RAFb

[50] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(2):411–412, February 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-050.

Saunders:1970:PEB

[51] Lloyd H. Saunders and G. Power. Population ecology of the brook trout, Salvelinus fontinalis, in Matamek Lake, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):413–424, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-051.

Johnson:1970:EAM

[52] M. G. Johnson, M. F. P. Michalski, and A. E. Christie. Effects of acid mine wastes on phytoplankton communities of two Northern Ontario lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):425–444, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-052.

Houde:1970:EWC

[53] Edward D. Houde and John L. Forney. Effects of water currents on distribution of walleye larvae in Oneida Lake, New York. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):445–456, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-053.

Templeman:1970:RMF

[54] Wilfred Templeman. A review of the morid fish genus Lepidion of the North Atlantic with first records of Lepidion eques from the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3): 457–498, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-054.

Templeman:1970:DAP

[55] Wilfred Templeman. Distribution of Anotopterus pharao in the North Atlantic and comparison of specimens of A. pharao from the Western North Atlantic with those from other areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):499–512, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-055.

Ackman:1970:LFA

[56] R. G. Ackman, C. A. Eaton, J. C. Sipos, S. N. Hooper, and John D. Castell. Lipids and fatty acids of two species of North Atlantic krill (Meganyctiphanes norvegica and Thysanoëssa inermis) and their role in the aquatic food web. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3): 513–533, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-056.

Caddy:1970:MSS

[57] J. F. Caddy. A method of surveying scallop populations from a submersible. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):535–549, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-057.

Hanson:1970:RGV

[58] David Hanson and Kjell Johansen. Relationship of gill ventilation and perfusion in Pacific dogfish, Squalus suckleyi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):551–564, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-058.

Hara:1970:EBO

[59] Toshiaki J. Hara. An electrophysiological basis for olfactory discrimination in homing salmon: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):565–586, March 1970. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f70-059.

June:1970:AYC

[60] Fred C. June. Atresia and year-class abundance of northern pike, Esox lucius, in two Missouri River impoundments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):587–591, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-060.

Dambergs:1970:EBM

[61] N. Dambergs and L. W. Regier. Estimation of bone material in fish protein concentrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3): 591–595, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-061.

Fagerlund:1970:DCC

[62] U. H. M. Fagerlund. Determining cortisol and cortisone simultaneously in salmonid plasma by competitive protein binding. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):596–601, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-062.

Idler:1970:SSH

[63] D. R. Idler, L. M. Safe, G. A. MacKinnon, and F. G. Proudfoot. Scallop sterols are hypocholesterolemic relative to cholesterol for the chick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):601–603, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-063.

Hiltz:1970:OTM

[64] Doris Fraser Hiltz. Occurrence of trigonelline (n-methyl nicotinic acid) in the adductor muscle of a lamellibranch, the sea scallop (Placopecten magellanicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3): 604–606, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-064.

Sutcliffe:1970:RBG

[65] W. H. Sutcliffe, Jr. Relationship between growth rate and ribonucleic acid concentration in some invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):606–609, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-065.

Parsons:1970:NRE

[66] L. S. Parsons. Northern range extension of the Atlantic mackerel, Scomber scombrus, to black Island, Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):610–613, March 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-066.

Dodge:1970:VSP

[67] Douglas P. Dodge and Hugh R. MacCrimmon. Vital statistics of a population of Great Lakes rainbow trout (Salmo gairdneri) characterized by an extended spawning season. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):613–618, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-067.

Anonymous:1970:RAFc

[68] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(3):619, March 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-068.

Lie:1970:BIC

[69] Ulf Lie and James C. Kelley. Benthic infauna communities off the coast of Washington and in Puget Sound: Identification and distribution of the communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4): 621–651, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-069.

Hynes:1970:MPO

[70] H. B. N. Hynes and B. J. Greib. Movement of phosphate and other ions from and through lake muds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):653–668, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-070.

Barr:1970:DVM

[71] Louis Barr. Diel vertical migration of Pandalus borealis in Kachemak Bay, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):

669–676, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-071.

Lucas:1970:CTE

[72] Henry F. Lucas, Jr., David N. Edgington, and Peter J. Colby. Concentrations of trace elements in Great Lakes fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):677–684, April 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-072.

Hargrave:1970:DGS

[73] Barry T. Hargrave. Distribution, growth, and seasonal abundance of Hyalella azteca (Amphipoda) in relation to sediment microflora. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):685–699, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-073.

Castell:1970:ETM

[74] C. H. Castell, Barbara Smith, and Wanda Neal. Effects of transition metal ions on the extractable protein of fish muscles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):701–714, April 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-074.

McCleave:1970:UTH

[75] James D. McCleave and Ross M. Horrall. Ultrasonic tracking of homing cutthroat trout (Salmo clarki) in Yellowstone Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):715–730, April 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-075.

McLeese:1970:BLE

[76] D. W. McLeese. Behaviour of lobsters exposed to bleached kraft mill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):731– 736, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-076.

Dowd:1970:CBT

[77] R. G. Dowd, E. Bakken, and O. Nakken. A comparison between two sonic measuring systems for demersal fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):737–742, April 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-077.

Shaw:1970:ELS

[78] Rosslyn Shaw, David C. Arnold, and Wilson B. Stallworthy. Effects of light on spat settlement of the American oyster (Crassostrea virginica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):743–748, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-078.

McInernev:1970:ASP

[79] John E. McInernev and David O. Evans. Action spectrum of the photoperiod mechanism controlling sexual maturation in the threespine stickleback, Gasterosteus aculeatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):749–763, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-079.

Lindsey:1970:ANM

[80] J. K. Lindsey, D. F. Alderdice, and L. V. Pienaar. Analysis of nonlinear models — the nonlinear response surface. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):765–791, April 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-080.

Otto:1970:DSP

[81] Robert G. Otto and John E. McInerney. Development of salinity preference in pre-smolt Coho salmon, Oncorhynchus kisutch. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):793–800, April 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-081.

Hoffman:1970:GLF

[82] A. Hoffman, J. G. Disney, B. E. Grimwood, and N. R. Jones. Glucose levels in fresh Tilapia muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):801–803, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-082.

Pitt:1970:RHS

[83] T. K. Pitt and G. H. Winters. A record of the hatchetfish Sternoptyx diaphana from the Grand Bank. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):804–805, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-083.

Uthe:1970:MDF

[84] J. F. Uthe, F. A. J. Armstrong, and M. P. Stainton. Mercury determination in fish samples by wet digestion and flameless atomic absorption spectrophotometry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4): 805–811, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-084.

MacCrimmon:1970:WDB

[85] Hugh R. MacCrimmon, T. Larry Marshall, and Barra L. Gots. World distribution of brown trout, Salmo trutta: Further observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):811–818, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-085.

McMullen:1970:IPD

[86] John C. McMullen and Harvey T. Yoshihara. An incidence of parasitism of deepwater king crab, Lithodes aequispina, by the barnacle Briarosaccus callosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4): 818–821, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-086.

Tomlinson:1970:GMC

[87] Patrick K. Tomlinson. A generalization of the Murphy catch equation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):821–825, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-087.

Anctil:1970:SPB

[88] M. Anctil and C. G. Gruchy. Stimulation and photography of bioluminescence in lanternfishes (Myctophidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):826–829, April 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-088.

Wobeser:1970:MCT

[89] G. Wobeser, N. O. Nielsen, R. H. Dunlop, and F. M. Atton. Mercury concentrations in tissues of fish from the Saskatchewan River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):830–834, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-089.

Anonymous:1970:RAFd

[90] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(4):835, April 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-090.

MacCrimmon:1970:SVS

[91] H. R. MacCrimmon and J. R. M. Kelso. Seasonal variation in selected nutrients of a river system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):837–846, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-091.

Duthie:1970:MSL

[92] Hamish C. Duthie and John C. H. Carter. The meromixis of Sunfish Lake, Southern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5): 847–856, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-092.

Bilinski:1970:ECC

[93] E. Bilinski and R. E. E. Jonas. Effects of coenzyme a and carnitine on fatty acid oxidation by rainbow trout mitochondria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):857–864, May 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-093.

Kabata:1970:SLC

[94] Z. Kabata. Some Lernaeopodidae (Copepoda) from fishes of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):865– 885, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-094.

Platt:1970:PPM

[95] Trevor Platt and D. V. Subba Rao. Primary production measurements on a natural plankton Bloom. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):887–899, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-095.

Hanek:1970:PTS

[96] George Hanek and William Threlfall. Parasites of the threespine stickleback (Gasterosteus aculeatus) in Newfoundland and Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):901–907, May 1970. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f70-096.

Wedemeyer:1970:SAM

[97] Gary Wedemeyer. Stress of anesthesia with M.S. 222 and benzocaine in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):909–914, May 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-097.

Kennedy:1970:RSA

[98] W. A. Kennedy. Reading scales to age Pacific cod (Gadus macrocephalus) from Hecate Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5): 915–922, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-098.

Uthe:1970:RVM

[99] J. F. Uthe and R. A. Ryder. Regional variation in muscle myogen polymorphism in walleye (Stizostedion vitreum vitreum) as related to morphology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):923– 927, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-099.

Groves:1970:BCC

[100] T. D. D. Groves. Body composition changes during growth in young sockeye (Oncorhynchus nerka) in fresh water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):929–942, May 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-100.

Johnson:1970:EVA

[101] A. G. Johnson, F. M. Utter, and H. O. Hodgins. Electrophoretic variants of L-alpha-glycerophosphate dehydrogenase in Pacific Ocean perch (Sebastodes alutus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5): 943–945, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-101.

Hunter:1970:LFS

[102] Charles J. Hunter and Winston E. Farr. Large floating structure for holding adult Pacific salmon (Oncorhynchus spp.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):947–950, May 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-102.

Pesch:1970:PPV

[103] Gerald Pesch. Plasma protein variation in a winter flounder (Pseudopleuronectes americanus) population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):951–954, May 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-103.

Yasutake:1970:OWD

[104] W. T. Yasutake and Harold Wolf. Occurrence of whirling disease of trout in Western United States. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):955–956, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-104.

Dunn:1970:FET

[105] R. S. Dunn. Further evidence for a three-year oocyte maturation time in the winter flounder (Pseudopleuronectes americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):957–960, May 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-105.

Knight:1970:POE

[106] William Knight. A possible optimization experiment in fishery management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):961– 963, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-106.

Pippy:1970:UUL

[107] John H. C. Pippy. Use of ultraviolet light to find parasitic nematodes in situ. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):963– 965, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-107.

McInerney:1970:HCP

[108] J. E. McInerney and D. O. Evans. Habitat characteristics of the Pacific hagfish, Polistotrema stouti. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):966–968, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-108.

Krantz:1970:PNA

[109] George E. Krantz and Carroll E. Heist. Prevalence of naturally acquired agglutinating antibodies against Aeromonas salmonicida in hatchery trout in Central Pennsylvania. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):969–973, May 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-109.

Crozier:1970:TCP

[110] George F. Crozier. Tissue carotenoids in prespawning and spawning sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):973–975, May 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-110.

Fry:1970:RSS

[111] F. E. J. Fry and E. T. Cox. A relation of size to swimming speed in rainbow trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5): 976–978, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-111.

Kennedy:1970:MDT

[112] W. A. Kennedy. Mortality of dart-tagged captive sablefish (Anoplopoma fimbria). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):979– 981, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-112.

Bailey:1970:OSO

[113] Reeve M. Bailey and C. G. Gruchy. Occella to supersede Occa for a genus of agonid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5): 981–983, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-113.

Johnston:1970:IBS

[114] C. E. Johnston and J. G. Eales. Influence of body size on silvering of Atlantic salmon (Salmo salar) at Parr–smolt transformation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):983–987, May 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-114.

Anonymous:1970:RAFe

[115] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(5):989, May 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-115.

Armstrong:1970:AFM

[116] Robert H. Armstrong. Age, food, and migration of Dolly Varden smolts in Southeastern Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):991–1004, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-116.

Brown:1970:BBL

[117] Joseph H. Brown, U. Theodore Hammer, and Gordon D. Koshinsky. Breeding biology of the lake chub, Couesius plumbeus, at Lac la Ronge, Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6): 1005–1015, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-117.

Berst:1970:PDS

[118] A. H. Berst and G. R. Spangler. Population dynamics of F_{1} splake (Salvelinus fontinalis × S. namaycush) in Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1017–1032, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-118.

Noble:1970:EMH

[119] Richard L. Noble. Evaluation of the Miller high-speed sampler for sampling yellow perch and walleye fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1033–1044, June 1970. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f70-119.

Tedla:1970:SAE

[120] Shibru Tedla and C. H. Fernando. Some aspects of the ecology of the parasite fauna of the gills of yellow perch, Perca flavescens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1045–1050, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-120.

Fowler:1970:DIZ

[121] Scott W. Fowler, Lawrence F. Small, and John M. Dean. Distribution of ingested zinc-65 in the tissues of some marine crustaceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1051–1058, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-121.

Magnuson:1970:VOF

[122] John J. Magnuson and Duane J. Karlen. Visual observation of fish beneath the ice in a winterkill lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1059–1068, June 1970. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f70-122.

Cameron:1970:GER

[123] James N. Cameron and John C. Davis. Gas exchange in rainbow trout (Salmo gairdneri) with varying blood oxygen capacity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1069–1085, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-123.

Marshall:1970:ESS

[124] T. Lawrence Marshall and Hugh R. MacCrimmon. Exploitation of selfsustaining Ontario stream populations of brown trout (Salmo trutta) and brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1087–1102, June 1970. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f70-124.

Timms:1970:LRB

[125] A. M. Timms and H. Kleerekoper. Locomotor responses of blinded goldfish (Carassius auratus) to remote perception of barriers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1103–1107, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-125.

Gruchy:1970:OIN

[126] C. G. Gruchy. Occella impi, a new species of sea poacher from British Columbia with notes on related species (Agonidae, Pisces). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1109–1114, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-126.

Clayton:1970:GML

[127] J. W. Clayton and W. G. Franzin. Genetics of multiple lactate dehydrogenase isozymes in muscle tissue of lake whitefish (Coregonus clupeaformis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1115–1121, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-127.

Dall:1970:OLH

[128] W. Dall. Osmoregulation in the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1123–1130, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-128.

Dyer:1970:VAC

[129] W. J. Dyer, D. F. Hiltz, R. G. Ackman, J. Hingley, and G. L. Fletcher. In vivo assimilation by cod muscle and liver tissue of elemental phosphorus from polluted sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1131–1139, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-129.

Liu:1970:EGM

[130] D. L. Liu and C. C. Walden. Enzymes of glucose metabolism in the caecum of the marine borer Bankia setacea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1141–1146, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-130.

Lund:1970:STL

[131] William A. Lund, Jr. and Russell C. Lockwood, Jr. Sonic tag for large decapod crustaceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (6):1147–1151, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-131.

Bohn:1970:SMB

[132] A. Bohn and C. C. Walden. Survey of marine borers in Canadian Atlantic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6): 1151–1154, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-132.

McCart:1970:SHL

[133] P. McCart and Nevin Aspinwall. Spawning habits of the largescale sucker, Catostomus macrocheilus, at Stave Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1154–1158, June 1970.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-133.

Watson:1970:GCM

[134] J. Watson and P. G. Wells. A gauge for carapace measurements of crabs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1158–1161, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-134.

Wedemeyer:1970:SBC

[135] Gary Wedemeyer and K. Chatterton. Some blood chemistry values for the rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1162–1164, June 1970. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f70-135.

Emery:1970:FCM

[136] Alan R. Emery. Fish and crayfish mortalities due to an internal seiche in Georgian Bay, Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1165–1168, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-136.

Fagerlund:1970:RMA

[137] U. H. M. Fagerlund. Response to mammalian ACTH of the interrenal tissue of sockeye salmon (Oncorhynchus nerka) at various stages of sexual maturation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (6):1169–1172, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-137.

Baldwin:1970:ODL

[138] Wayne J. Baldwin. Oxygenating device for live-bait Wells. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1172–1174, June 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-138.

Anonymous:1970:RAFf

[139] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(6):1175, June 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-139.

Tyler:1970:RGE

[140] A. V. Tyler. Rates of gastric emptying in young cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1177–1189, July 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-140.

Dill:1970:EGS

[141] L. M. Dill and T. G. Northcote. Effects of gravel size, egg depth, and egg density on intragravel movement and emergence of Coho salmon (Oncorhynchus kisutch) alevins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1191–1199, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-141.

Lantz:1970:FED

[142] A. W. Lantz and M. Vaisey. Flavor effects of different Woods on whitefish smoked in a kiln with controlled temperature, humidity, and air velocity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1201–1207, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-142.

Beamish:1970:ITS

[143] F. W. H. Beamish. Influence of temperature and salinity acclimation on temperature preferenda of the euryhaline fish Tilapia nilotica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1209–1214, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-143.

Lister:1970:SHU

[144] D. B. Lister and H. S. Genoe. Stream habitat utilization by cohabiting underyearlings of Chinook (Oncorhynchus tshawytscha) and Coho (O. kisutch) salmon in the Big Qualicum River, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1215–1224, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-144.

vanOlst:1970:SAO

[145] Jon C. van Olst and John R. Hunter. Some aspects of the organization of fish schools. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7): 1225–1238, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f70-145.

Zyblut:1970:LTC

[146] E. R. Zyblut. Long-terra changes in the limnology and macrozooplankton of a large British Columbia lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1239–1250, July 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-146.

Parsons:1970:LPP

[147] T. R. Parsons, R. J. LeBrasseur, and W. E. Barraclough. Levels of production in the pelagic environment of the Strait of Georgia, British Columbia: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27 (7):1251–1264, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-147.

Smoker:1970:GRL

[148] William Smoker and William G. Pearcy. Growth and reproduction of the lanternfish Stenobrachius leucopsarus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1265–1275, July 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-148.

Arthur:1970:ECG

[149] John W. Arthur and Edward N. Leonard. Effects of copper on Gammarus pseudolimnaeus, Physa integra, and Campeloma decisum in soft water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1277–1283, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-149.

Amend:1970:MCVa

[150] Donald F. Amend and Velma C. Chambers. Morphology of certain viruses of salmonid fishes. I. In vitro studies of some viruses causing hematopoietic necrosis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1285–1293, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-150.

Saunders:1970:IPS

[151] Richard L. Saunders and Eugene B. Henderson. Influence of photoperiod on smolt development and growth of Atlantic salmon (Salmo solar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1295–1311, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-151.

Becker:1970:MTT

[152] C. Dale Becker. Marine trematode Tubulovesicula lindbergi (Digenea: Hemiuridae) from resident white sturgeon in the Columbia River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1313–1316, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-152.

Threlfall:1970:MPE

[153] William Threlfall and George Hanek. Metazoan parasites, excluding Monogenea, from longnose and white suckers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1317–1319, July 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-153.

Katkansky:1970:SHP

[154] Stanley C. Katkansky and Ronald W. Warner. Sporulation of a haplosporidan in a Pacific oyster (Crassostrea gigas) in Humboldt Bay, California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1320– 1321, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-154.

Smith:1970:MVB

[155] Lynwood S. Smith and Timothy W. Newcomb. A modified version of the blazka respirometer and exercise chamber for large fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1321–1324, July 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-155.

Tsuyuki:1970:ESH

[156] H. Tsuyuki and A. P. Ronald. Existence in salmonid hemoglobins of molecular species with three and four different polypeptides. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1325–1328, July 1970. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f70-156.

Idler:1970:FSS

[157] D. R. Idler, G. A. MacKinnon, and Allan C. Cox. Further studies on scallop sterols in relation to cholesterol in the chick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7):1329–1331, July 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-157.

Anonymous:1970:RAFg

[158] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(7): 1333, July 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-158.

Anderson:1970:ERZ

[159] R. Stewart Anderson. Effects of rotenone on zooplankton communities and a study of their recovery patterns in two mountain lakes in Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1335–1356, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-159.

Brockerhoff:1970:DEA

[160] H. Brockerhoff, R. J. Hoyle, and P. C. Hwang. Digestive enzymes of the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1357–1370, August 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-160.

McLeese:1970:DDS

[161] D. W. McLeese. Detection of dissolved substances by the American lobster (Homarus americanus) and olfactory attraction between lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1371–1378, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-161.

Fletcher:1970:YPP

[162] G. L. Fletcher, R. J. Hoyle, and D. A. Horne. Yellow phosphorus pollution: its toxicity to seawater-maintained brook trout (Salvelinus fontinalis) and smelt (Osmerus mordax). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1379–1384, August 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-162.

Amend:1970:MCVb

[163] Donald F. Amend and Velma C. Chambers. Morphology of certain viruses of salmonid fishes. II. In vivo studies of infectious hematopoietic necrosis virus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8): 1385–1388, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-163.

Anderson:1970:PAA

[164] Douglas P. Anderson and George W. Klontz. Precipitating antibody against Aeromonas salmonicida in serums of inbred albino rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8): 1389–1393, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-164.

Hargrave:1970:ECG

[165] Barry T. Hargrave and Glen H. Geen. Effects of copepod grazing on two natural phytoplankton populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1395–1403, August 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-165.

Sparling:1970:CCP

[166] J. H. Sparling and C. Nalewajko. Chemical composition and phytoplankton of lakes in Southern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1405–1428, August 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-166.

Bams:1970:ERH

[167] R. A. Bams. Evaluation of a revised hatchery method tested on pink and chum salmon fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8): 1429–1452, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-167.

Platt:1970:SHP

[168] Trevor Platt, Lloyd M. Dickie, and Ronald W. Trites. Spatial heterogeneity of phytoplankton in a near-shore environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1453–1473, August 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-168.

Parsons:1970:WFL

[169] John W. Parsons. Walleye fishery of Lake Erie in 1943–62 with emphasis on contributions of the 1942–61 year-classes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1473–1489, August 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-169.

MacDonald:1970:OMP

[170] J. Rod MacDonald. An occurrence of male predominance among repeat spawning Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1491–1492, August 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-170.

Vallentyne:1970:VDB

[171] J. R. Vallentyne, W. E. Johnson, and A. J. Harris. A visual demonstration of the beneficial effects of sewage treatment for phosphate removal on particulate matter production in waters of Lakes Erie and Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1493–1496, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-171.

Macek:1970:SFC

[172] Kenneth J. Macek and Sidney Korn. Significance of the food chain in DDT accumulation by fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1496–1498, August 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-172.

Anonymous:1970:CRP

[173] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1499, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-173.

Anonymous:1970:RAFh

[174] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(8):1500, August 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-174.

Calder:1970:THS

[175] Dale R. Calder. Thecate hydroids from the shelf waters of Northern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1501–1547, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-175.

Templeman:1970:VOM

[176] Wilfred Templeman. Vertebral and other meristic characteristics of Greenland halibut, Reinhardtius hippoglossoides, from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1549–1562, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-176.

Williscroft:1970:LDS

[177] S. N. Williscroft and H. Tsuyuki. Lactate dehydrogenase systems of rainbow trout — evidence for polymorphism in liver and additional subunits in gills. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1563–1567, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-177.

Heffernan:1970:EBN

[178] W. Paul Heffernan and Victor J. Cabelli. Elimination of bacteria by the northern quahaug (Mercenaria mercenaria): Environmental parameters significant to the process. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1569–1577, September 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-178.

Cabelli:1970:EBS

[179] Victor J. Cabelli and W. Paul Heffernan. Elimination of bacteria by the soft shell clam, Mya arenaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1579–1587, September 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-179.

Lauder:1970:KTF

[180] J. T. Lauder, W. A. MacCallum, and D. R. Idler. Keeping time of frozen redfish (Sebastes marinus mentella) fillets in relation to handling of the raw material and storage temperatures after processing and freezing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1589–1605, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-180.

Watson:1970:MME

[181] J. Watson. Maturity, mating, and egg laying in the spider crab, Chionoecetes opilio. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9): 1607–1616, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-181.

Moller:1970:TPA

[182] Dag Møller. Transferrin polymorphism in Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1617–1625, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-182.

Davis:1970:EOC

[183] John C. Davis and Kenneth Watters. Evaluation of opercular catheterization as a method for sampling water expired by fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1627–1635, September 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-183.

Brown:1970:ARV

[184] Clinton E. Brown and Barry S. Muir. Analysis of ram ventilation of fish gills with application to skipjack tuna (Katsuwonus pelamis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1637–1652, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-184.

Cobb:1970:EST

[185] J. Stanley Cobb. Effect of solitude on time between fourth and fifth larval molts in the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1653–1655, September 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-185.

Lackey:1970:SDD

[186] Robert T. Lackey. Seasonal depth distributions of landlocked Atlantic salmon, brook trout, landlocked alewives, and American smelt in a small lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1656–1661, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-186.

Miller:1970:FFS

[187] Bruce S. Miller. Food of flathead sole (Hippoglossoides elassodon) in East Sound, Orcas Island, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1661–1665, September 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-187.

Anonymous:1970:RAFi

[188] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(9):1667, September 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-188.

Ackman:1970:BIS

[189] R. G. Ackman and C. A. Eaton. Biochemical implications of seasonal trends in the iodine values and free fatty acid levels of commercially produced Atlantic Coast herring oils. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1669–1683, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-189.

Castell:1970:FDS

[190] C. H. Castell, Wanda Neal, and Barbara Smith. Formation of dimethylamine in stored frozen sea fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1685–1690, October 1970. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f70-190.

Flannagan:1970:EVG

[191] John F. Flannagan. Efficiencies of various grabs and corers in sampling freshwater benthos. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10): 1691–1700, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-191.

Duff:1970:TRT

[192] G. F. D. Duff. Tidal resonance and tidal barriers in the Bay of Fundy system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1701–1728, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-192.

McCauley:1970:PTY

[193] R. W. McCauley and J. S. Tait. Preferred temperature of yearling lake trout, Salvelinus namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1729–1733, October 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-193.

Manion:1970:MLS

[194] Patrick J. Manion and Thomas M. Stauffer. Metamorphosis of the landlocked sea lamprey, Petromyzon marinus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1735–1746, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-194.

Hughes:1970:AKC

[195] Steven E. Hughes, Fred W. Hipkins, and Donald D. Worlund. Adaptation of king crab pots for capturing sablefish (Anoplopoma fimbria). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1747–1755, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-195.

Pivnicka:1970:MVB

[196] Karel Pivnička. Morphological variation in the burhot (Lota lota) and recognition of the subspecies: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1757–1765, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-196.

Brett:1970:ETR

[197] J. R. Brett and D. A. Higgs. Effect of temperature on the rate of gastric digestion in finger ling sockeye salmon, Oncorhynchus nerka. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1767–1779, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f70-197.

Westrheim:1970:SRE

[198] S. J. Westrheim. Survey of rockfishes, especially Pacific Ocean perch, in the Northeast Pacific Ocean, 1963–66. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1781–1809, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-198.

McRoy:1970:SSO

[199] C. Peter McRoy. Standing stocks and other features of eelgrass (Zostera marina) populations on the coast of Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1811–1821, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-199.

MacCallum:1970:DSO

[200] Wayne R. MacCallum and Henry A. Regier. Distribution of smelt, Osmerus mordax, and the smelt fishery in Lake Erie in the early 1960’s. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1823–1846, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-200.

Powell:1970:SUA

[201] N. A. Powell. Schizoporella unicornis — an alien bryozoan introduced into the Strait of Georgia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1847–1853, October 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-201.

Gee:1970:ABB

[202] J. H. Gee. Adjustment of buoyancy in blacknose dace, Rhinichthys atratulus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1855–1859, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-202.

Davis:1970:ECT

[203] John C. Davis. Estimation of circulation time in rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1860–1863, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-203.

Brownell:1970:CMR

[204] Willard N. Brownell. Comparison of Mysis relicta and Pontoporeia affinis as possible intermediate hosts for the acanthocephalan Echinorhynchus salmonis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1864–1866, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-204.

Hamilton:1970:MCS

[205] Andrew L. Hamilton, William Burton, and John F. Flannagan. A multiple corer for sampling profundal benthos. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1867–1869, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-205.

Lane:1970:EDG

[206] Charles E. Lane and Edward D. Scura. Effects of dieldrin on glutamic oxaloacetic transaminase in Poecilia latipinna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1869–1871, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-206.

Case:1970:SBC

[207] Brian Case. Spawning behaviour of the chestnut lamprey (Ichthyomyzon castaneus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1872–1874, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-207.

Chenoweth:1970:SVC

[208] Stanley B. Chenoweth. Seasonal variations in condition of larval herring in Boothbay area of the Maine Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1875–1879, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-208.

Lear:1970:FGH

[209] W. H. Lear. Fecundity of Greenland halibut (Reinhardtius hippoglossoides) in the Newfoundland–Labrador area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1880–1882, October 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-209.

McKim:1970:CBB

[210] J. M. McKim, G. M. Christensen, and Evelyn P. Hunt. Changes in the blood of brook trout (Salvelinus fontinalis) after short-term and longterm exposure to copper. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1883–1889, October 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-210.

Smith:1970:TMS

[211] R. J. F. Smith. A technique for marking small fish with injected fluorescent dyes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10): 1889–1891, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-211.

Mercer:1970:TLS

[212] M. C. Mercer. The tropical loliginid squid Sepioteuthis sepioidea from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10): 1892–1893, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-212.

Threlfall:1970:MPS

[213] William Threlfall and George Hanek. Metazoan parasites of salmonids and coregonids from the Avalon Peninsula, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1894–1897, October 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-213.

Hiltz:1970:GAS

[214] Doris Fraser Hiltz and Thérèse E. Lightle. β-glucuronidase and arylsulphatase in some marine invertebrates of the Canadian Atlantic Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1898–1900, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-214.

Anonymous:1970:RAFj

[215] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(10):1901, October 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-215.

Sergeant:1970:IRC

[216] D. E. Sergeant, A. W. Mansfield, and B. Beck. Inshore records of cetacea for Eastern Canada, 1949–68. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1903–1915, November 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-216.

Sutcliffe:1970:CAP

[217] W. H. Sutcliffe, Jr., R. W. Sheldon, and A. Prakash. Certain aspects of production and standing stock of particulate matter in the surface waters of the Northwest Atlantic Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1917–1926, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-217.

Sutterlin:1970:TRA

[218] A. M. Sutterlin and N. Sutterlin. Taste responses in Atlantic salmon (Salmo salar) Parr. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11): 1927–1942, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-218.

Sekerak:1970:PCS

[219] A. D. Sekerak. Parasitic copepods of Sebastodes alutus including Chondracanthus triventricosus sp. nov. and Colobomatus kyphosus sp. nov. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1943–1960, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-219.

Brinkhurst:1970:DAT

[220] R. O. Brinkhurst. Distribution and abundance of tubificid (Oligochaeta) species in Toronto Harbour, Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1961–1969, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-220.

Appleby:1970:DRT

[221] A. G. Appleby and R. O. Brinkhurst. Defecation rate of three tubificid oligochaetes found in the sediment of Toronto Harbour, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1971–1982, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-221.

Hayward:1970:BEF

[222] M. J. Hayward, W. A. MacCallum, and D. H. Shaw. Bacteriological evaluation of flounder (Hippoglossoides platessoides) frozen at sea and thawed by water immersion. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1983–1986, November 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-222.

Northcote:1970:MLD

[223] T. G. Northcote, S. N. Williscroft, and H. Tsuyuki. Meristic and lactate dehydrogenase genotype differences in stream populations of rainbow trout below and above a waterfall. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1987–1995, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-223.

Manohar:1970:PGO

[224] S. V. Manohar. Postmortem glycolytic and other biochemical changes in white muscle of white sucker (Catostomus commersoni) and northern pike (Esox lucius) at 0 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):1997–2002, November 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-224.

Gardner:1970:NTI

[225] Linda J. Gardner. 5’-nucleotidase of testes of immature sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2003–2007, November 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-225.

Schindler:1970:NSP

[226] D. W. Schindler and J. E. Nighswander. Nutrient supply and primary production in Clear Lake, Eastern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2009–2036, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-226.

Ennis:1970:RAB

[227] G. P. Ennis. Reproduction and associated behaviour in the shorthorn sculpin, Myoxocephalus scorpius in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2037–2045, November 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-227.

Edsall:1970:TTB

[228] Thomas A. Edsall, Donald V. Rottiers, and Edward H. Brown. Temperature tolerance of bloater (Coregonus hoyi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2047–2052, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-228.

Powles:1970:DDV

[229] P. M. Powles and A. C. Kohler. Depth distributions of various stages of witch flounder (Glyptocephalus cynoglossus) off Nova Scotia and in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11): 2053–2062, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-229.

Cooper:1970:GCT

[230] Edwin L. Cooper. Growth of cutthroat trout (Salmo clarki) in Chef Creek, Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2063–2070, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-230.

Yuen:1970:BST

[231] Heeny S. H. Yuen. Behavior of skipjack tuna, Katsuwonus pelamis, as determined by tracking with ultrasonic devices. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2071–2079, November 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-231.

Hulburt:1970:RPT

[232] Edward M. Hulburt and Nathaniel Corwin. Relation of the phytoplankton to turbulence and nutrient renewal in Casco Bay, Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2081–2090, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-232.

Anonymous:1970:Oa

[233] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2091–2093, November 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-233.

Powell:1970:HEB

[234] N. A. Powell, C. S. Sayce, and D. F. Tufts. Hyperplasia in an estuarine bryozoan attributable to coal tar derivatives. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2095–2096, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-234.

Parsons:1970:OJS

[235] L. S. Parsons and V. M. Hodder. Occurrence of juvenile and spawning Atlantic mackerel in Southeastern Newfoundland coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2097–2100, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-235.

Gould:1970:CAL

[236] Edith Gould. The case for accurate labeling of superchilled fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2101–2103, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-236.

Winters:1970:MMS

[237] G. H. Winters. Meristics and morphometries of sand launce in the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11): 2104–2108, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-237.

Gray:1970:MPE

[238] R. W. Gray and J. A. McKenzie. Muscle protein electrophoresis in the genus Salmo of Eastern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2109–2112, November 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-238.

Hennick:1970:RCS

[239] Daniel P. Hennick. Reproductive cycle, size at maturity, and sexual composition of commercially harvested weathervane scallops (Patinopecten caurinus) in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11): 2112–2119, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-239.

Green:1970:BGP

[240] John M. Green. The banded gunnel, Pholis fasciatus, in Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2120–2121, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-240.

Anonymous:1970:RAFk

[241] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(11):2123, November 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-241.

Bartnik:1970:RIB

[242] Victor G. Bartnik. Reproductive isolation between two sympatric dace, Rhinichthys atratulus and R. cataractae, in Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2125–2141, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-242.

Peer:1970:RBB

[243] D. L. Peer. Relation between biomass, productivity, and loss to predators in a population of a marine benthic polychaete, Pectinaria hyperborea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2143–2153, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-243.

Ennis:1970:AGS

[244] G. P. Ennis. Age, growth, and sexual maturity of the shorthorn sculpin, Myoxocephalus scorpius, in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2155–2158, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-244.

Kabata:1970:DBL

[245] Z. Kabata. Discovery of Brachiella lageniformis (Copepoda: Lernaeopodidae) in the Canadian Pacific and its significance to zoogeography of the genus Merluccius (Pisces: Teleostei). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2159–2165, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-245.

Chen:1970:ZES

[246] F. Y. Chen and H. Tsuyuki. Zone electrophoretic studies on the proteins of Tilapia mossambica and T. hornorum and their F_{1} hybrids, T. zillii, and T. melanopleura. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12): 2167–2177, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-246.

Peters:1970:ADP

[247] M. D. Peters and R. J. Bose. Approaches to the direct pigmentation of canned salmon: Synthesis and properties of 4,4’-bis (pyridiniummethylcarbonylhydrazino) canthaxanthindiylidene dichloride. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2179–2184, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-247.

Gardner:1970:HHR

[248] George R. Gardner and Paul P. Yevich. Histological and hematological responses of an estuarine teleost to cadmium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2185–2196, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-248.

Withler:1970:SRP

[249] F. C. Withler and R. B. Morley. Sex-related parental influences on early development of Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2197–2214, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-249.

Winters:1970:BCC

[250] G. H. Winters. Biological changes in coastal capelin from the overwintering to the spawning condition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2215–2224, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-250.

Grant:1970:CEP

[251] Blake F. Grant and Paul M. Mehrle. Chronic endrin poisoning in goldfish, Carassius auratus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12): 2225–2232, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-251.

Tsuyuki:1970:ASA

[252] H. Tsuyuki and S. J. Westrheim. Analyses of the Sebastes aleutianus– S. melanostomus complex, and description of a new scorpaenid species, Sebastes caenaematicus, in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2233–2254, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-252.

Hulburt:1970:RHB

[253] Edward M. Hulburt. Relation of heat budget to circulation in Casco Bay, Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2255–2260, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-253.

Pitt:1970:DAS

[254] T. K. Pitt. Distribution, abundance, and spawning of yellowtail flounder, Limanda ferruginea, in the Newfoundland area of the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2261–2271, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-254.

Lie:1970:SCS

[255] Ulf Lie and Dale S. Kisker. Species composition and structure of benthic infauna communities off the coast of Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2273–2285, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-255.

Donaldson:1970:ESM

[256] Edward M. Donaldson and Ulf H. M. Fagerlund. Effect of sexual maturation and gonadectomy at sexual maturity on cortisol secretion rate in sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2287–2296, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-256.

Morris:1970:CCM

[257] Byron Morris. Calanoid copepods from midwater trawls in the North Pacific along 160^{o}E. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12): 2297–2321, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-257.

Fagerlund:1970:DCS

[258] Ulf H. M. Fagerlund and Edward M. Donaldson. Dynamics of cortisone secretion in sockeye salmon (Oncorhynchus nerka) during sexual maturation and after gonadectomy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2323–2331, December 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-258.

Anonymous:1970:Ob

[259] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2333–2335, December 1970. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f70-259.

Woodhead:1970:ETU

[260] P. M. J. Woodhead. An effect of thyroxine upon the swimming of cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2337–2338, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-260.

Idler:1970:RMP

[261] D. R. Idler, B. Truscott, and M. J. O’Halloran. A rapid method for partial hypophysectomy in three species of skate (Raja radiata, R. ocellata, and R. erinacea). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12): 2339–2342, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-261.

Bulow:1970:RDR

[262] Frank J. Bulow. RNA–DNA ratios as indicators of recent growth rates of a fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2343–2349, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f70-262.

Smith:1970:EFA

[263] R. J. F. Smith. Effects of food availability on aggression and nest building in brook stickleback (Culaea inconstans). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2350–2355, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-263.

Jenkins:1970:FRT

[264] T. M. Jenkins, Jr., G. V. Elliott, and C. R. Feldmeth. Feeding of rainbow trout (Salmo gairdneri) in relation to abundance of drifting invertebrates in a mountain stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12): 2356–2361, December 1970. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f70-264.

McPhail:1970:NSP

[265] J. D. McPhail. A new species of prickleback, Bryozoichthys marjorius (Chirolophinae), from the Eastern North Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2362–2365, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-265.

McFadden:1970:FN

[266] T. W. McFadden. Furunculosis in nonsalmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2365–2370, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-266.

Utter:1970:TPC

[267] Fred M. Utter, Warren E. Ames, and Harold O. Hodgins. Transferrin polymorphism in Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2371–2373, December 1970. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f70-267.

Smith:1970:CHI

[268] Roderick M. Smith and Charles F. Cole. Chlorinated hydrocarbon insecticide residues in winter flounder, Pseudopleuronectes americanus, from the Weweantic River Estuary, Massachusetts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2374–2380, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-268.

Anonymous:1970:IP

[269] Anonymous. Index to publications for 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 27(12):2381–2487, December 1970. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f70-269.

Castell:1971:PDM

[270] C. H. Castell, Barbara Smith, and Wanda Neal. Production of dimethylamine in muscle of several species of gadoid fish during frozen storage, especially in relation to presence of dark muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):1–5, January 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-001.

Boddington:1971:SCC

[271] M. J. Boddington and D. F. Mettrick. Seasonal changes in the chemical composition and food reserves of the freshwater triclad Dugesia tigrina (Platyhelminthes: Turbellaria). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):7–14, January 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-002.

Barlow:1971:PEI

[272] John Barlow and George J. Ridgway. Polymorphisms of esterase isozymes in the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):15–21, January 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-003.

Frost:1971:TSC

[273] Bruce W. Frost. Taxonomic status of Calanus finmarchicus and C. glacialis (Copepoda), with special reference to adult males. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):23–30, January 1971. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f71-004.

Iles:1971:FPC

[274] Christine Iles. Fistulicola plicatus (Cestoda) and Tristoma spp. (Trematoda) on swordfish from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):31–38, January 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-005.

Spence:1971:DBU

[275] J. A. Spence and H. B. N. Hynes. Differences in benthos upstream and downstream of an impoundment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):35–43, January 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-006.

Spence:1971:DFP

[276] J. A. Spence and H. B. N. Hynes. Differences in fish populations upstream and downstream of a mainstream impoundment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):45–46, January 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-007.

Yu:1971:RHE

[277] Man-Lim Yu, Charles W. Kiley, Daniel A. Sarot, and Alfred Perlmutter. Relation of hemosiderin to erythropoiesis in the blue gourami, Trichogaster trichopterus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):47–48, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-008.

Neish:1971:CSS

[278] I. C. Neish. Comparison of size, structure, and distributional patterns of two Salamander populations in Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):49–58, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-009.

Sprague:1971:DRC

[279] J. B. Sprague and J. R. Duffy. DDT residues in Canadian Atlantic fishes and shellfishes in 1967. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):59–64, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-010.

Sakagawa:1971:PBL

[280] Gary T. Sakagawa and Richard L. Pycha. Population biology of lake trout (Salvelinus namaycush) of Lake Superior before 1950. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):65–71, January 1971. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f71-011.

Stockner:1971:EEN

[281] John G. Stockner. Ecological energetics and natural history of Hedriodiscus truquii (Diptera) in two thermal spring communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):73–94, January 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-012.

Ford:1971:DGR

[282] Peter Ford. Differential growth rate in the tail of the Pacific big skate, Raja binoculata. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (1):95–98, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-013.

McCrary:1971:SSC

[283] Jerry A. McCrary. Sternal spines as a characteristic for differentiating between females of some pandalidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):98–100, January 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-014.

MacKelvie:1971:DRR

[284] R. M. MacKelvie. A device for rapid removal of embryos from fish eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):100–102, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-015.

Reynolds:1971:PBT

[285] James B. Reynolds and Douglas R. McLain. Plotting of bathythermograph transect data on a printer. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):102–103, January 1971. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f71-016.

Li:1971:PPS

[286] M. F. Li and G. S. Traxler. A proteolytic pseudomonad from skin lesions of rainbow trout (Salmo gairdneri). III. Morphological studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):104–105, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-017.

Fredeen:1971:RDD

[287] F. J. H. Fredeen, J. G. Saha, and L. M. Royer. Residues of DDT, DDE, and DDD in fish in the Saskatchewan River after using DDT as a blackfly larvicide for twenty years. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):105–109, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-018.

McKone:1971:RWS

[288] D. McKone. Rate at which sockeye salmon alevins are evacuated from the stomach of mountain whitefish (Prosopium williamsoni). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):110–111, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-019.

Smith:1971:UHN

[289] Charlie E. Smith. An undifferentiated hematopoietic neoplasm with histologic manifestations of leukemia in a cutthroat trout (Salmo clarki). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):112–113, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-020.

Luhning:1971:SFM

[290] Charles W. Luhning and Paul D. Harman. Sampling of fish muscle for M.S. 222 and quinaldine residues. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):113–115, January 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-021.

McCart:1971:MDB

[291] P. McCart and P. Craig. Meristic differences between anadromous and freshwater-resident Arctic char (Salvelinus alpinus) in the Sagavanirktok River Drainage, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):115–118, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-022.

Anonymous:1971:RAFa

[292] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):119, January 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-023.

Anonymous:1971:AFPa

[293] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(1):120, January 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-024.

Anonymous:1971:EF

[294] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):121, February 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-025.

Johnson:1971:RBD

[295] W. E. Johnson and J. R. Vallentyne. Rationale, background, and development of experimental lake studies in Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):123–128, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-026.

Cleugh:1971:RIS

[296] T. R. Cleugh and B. W. Hauser. Results of the initial survey of the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):129–137, February 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-027.

Brunskill:1971:GBS

[297] G. J. Brunskill and D. W. Schindler. Geography and bathymetry of selected lake basins, Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):139–155, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-028.

Schindler:1971:LTO

[298] D. W. Schindler. Light, temperature, and oxygen regimes of selected lakes in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):157–169, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-029.

Armstrong:1971:PCC

[299] F. A. J. Armstrong and D. W. Schindler. Preliminary chemical characterization of waters in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):171–187, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-030.

Schindler:1971:PPP

[300] D. W. Schindler and S. K. Holmgren. Primary production and phytoplankton in the Experimental Lakes Area, Northwestern Ontario, and other low-carbonate waters, and a liquid scintillation method for determining ^{14}C activity in photosynthesis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):189–201, February 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-031.

Sakamoto:1971:CFI

[301] Mitsuru Sakamoto. Chemical factors involved in the control of phytoplankton production in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):203–213, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-032.

Stockner:1971:PEL

[302] John G. Stockner and F. A. J. Armstrong. Periphyton of the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):215–229, February 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-033.

Patalas:1971:CPC

[303] Kazimierz Patalas. Crustacean plankton communities in forty-five lakes in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):231–244, February 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-034.

Schindler:1971:VDS

[304] D. W. Schindler and Bengt Novén. Vertical distribution and seasonal abundance of zooplankton in two shallow lakes of the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):245–256, February 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-035.

Hamilton:1971:ZFL

[305] Andrew L. Hamilton. Zoobenthos of fifteen lakes in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):257–263, February 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-036.

Stockner:1971:PCL

[306] John G. Stockner. Preliminary characterization of lakes in the Experimental Lakes Area, Northwestern Ontario, using diatom occurrences in sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):265–275, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-037.

Brunskill:1971:CSS

[307] G. J. Brunskill, D. Povoledo, B. W. Graham, and M. P. Stainton. Chemistry of surface sediments of sixteen lakes in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):277–294, February 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-038.

Schindler:1971:HED

[308] D. W. Schindler. A hypothesis to explain differences and similarities among lakes in the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):295–301, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-039.

Anonymous:1971:RAFb nal de l’Office des recherches sur les pêcheries du Canada, 28(2):303, February 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-040.

Anonymous:1971:AFPb

[310] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(2):304, February 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-041.

Laycock:1971:TPB

[311] R. A. Laycock and L. W. Regier. Trimethylamine-producing bacteria on haddock (Melanogrammus aeglefinus) fillets during refrigerated storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):305–309, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-042.

Anderson:1971:CPA

[312] R. Stewart Anderson. Crustacean plankton of 146 alpine and subalpine lakes and ponds in Western Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):311–321, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-043.

Hewitt:1971:SLC

[313] G. C. Hewitt. Species of Lepeophtheirus (Copepoda, Caligidae) recorded from the ocean sunfish (Mola mola) and their implications for the Caligid genus Dentigryps. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (3):323–334, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-044.

Wavre:1971:IBS

[314] M. Wavre and R. O. Brinkhurst. Interactions between some tubificid oligochaetes and bacteria found in the sediments of Toronto Harbour, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):335– 341, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-045.

Otto:1971:ESS

[315] Robert G. Otto. Effects of salinity on the survival and growth of presmolt Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):343–349, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-046.

Kerswill:1971:RRU

[316] C. J. Kerswill. Relative rates of utilization by commercial and sport fisheries of Atlantic salmon (Salmo salar) from the Miramichi River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (3):351–363, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-047.

Paterson:1971:CSC

[317] C. G. Paterson and C. H. Fernando. A comparison of a simple corer and an Ekman grab for sampling shallow-water benthos. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):365–368, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-048.

Calaprice:1971:XRS

[318] J. R. Calaprice. X-ray spectrometric and multivariate analysis of sockeye salmon (Oncorhynchus nerka) from different geographic regions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):369–377, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-049.

Morrison:1971:DOR

[319] George Morrison. Dissolved oxygen requirements for embryonic and larval development of the hardshell clam, Mercenaria mercenaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):379–381, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-050.

Green:1971:HTM

[320] John M. Green. High tide movements and homing behaviour of the tidepool sculpin Oligocottus maculosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):383–389, March 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-051.

Taylor:1971:RWC

[321] F. J. R. Taylor, D. J. Blackbourn, and Janice Blackbourn. The redwater ciliate Mesodinium rubrum and its “incomplete symbionts”: A review including new ultrastructural observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):391–407, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-052.

Brett:1971:STA

[322] J. R. Brett. Satiation time, appetite, and maximum food intake of sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):409–415, March 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-053.

Gunderson:1971:RPP

[323] Donald R. Gunderson. Reproductive patterns of Pacific Ocean perch (Sebastodes alutus) off Washington and British Columbia and their relation to bathymetric distribution and seasonal abundance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):417–425, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-054.

Nelson:1971:CPP

[324] Joseph S. Nelson. Comparison of the pectoral and pelvic skeletons and of some other bones and their phylogenetic implications in the aulorhynchidae and gasterosteidae (Pisces). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):427–442, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-055.

White:1971:VFC

[325] E. G. White. A versatile Fortran computer program for the capture– recapture stochastic model of G. M. Jolly. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):443–445, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-056.

Lenarz:1971:EPA

[326] William H. Lenarz and James H. Green. Electronic processing of acoustical data for fishery research. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):446–447, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-057.

Broughton:1971:SFL

[327] E. Broughton and L. P. E. Choquette. Subcutaneous fibroma in lake trout (Salvelinus namaycush) in Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):448–449, March 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-058.

Savitz:1971:NEP

[328] J. Savitz. Nitrogen excretion and protein consumption of the bluegill sunfish (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):449–451, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-059.

MacCrimmon:1971:WDB

[329] Hugh R. MacCrimmon, Barra L. Gots, and J. Scott Campbell. World distribution of brook trout, Salvelinus fontinalis: Further observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):452–456, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-060.

Pitt:1971:FYF

[330] T. K. Pitt. Fecundity of the yellowtail flounder (Limanda ferruginea) from the Grand Bank, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):456–457, March 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-061.

Jensen:1971:RBT

[331] A. L. Jensen. Response of brook trout (Salvelinus fontinalis) populations to a fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (3):458–460, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-062.

Lear:1971:RAM

[332] W. H. Lear and J. H. C. Pippy. A record of the Atlantic menhaden (Brevoortia tyrannus) from Northumberland Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):461–462, March 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-063.

Anonymous:1971:RAFc

[333] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):463, March 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-064.

Anonymous:1971:AFPc

[334] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(3):464, March 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-065.

Khan:1971:IVP

[335] N. Y. Khan and S. U. Qadri. Intraspecific variations and postglacial distribution of lake char (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):465–476, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-066.

vanOverbeeke:1971:HEK

[336] A. P. van Overbeeke and J. R. McBride. Histological effects of 11ketotestosterone, 17α-methyltestosterone, estradiol, estradiol cypionate, and cortisol on the interrenal tissue, thyroid gland, and pituitary gland of gonadectomized sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):477–484, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-067.

McBride:1971:EAE

[337] J. R. McBride and A. P. van Overbeeke. Effects of androgens, estrogens, and cortisol on the skin, stomach, liver, pancreas, and kidney in gonadectomized adult sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):485–490, April 1971. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f71-068.

Saunders:1971:CRR

[338] R. L. Saunders and A. M. Sutterlin. Cardiac and respiratory responses to hypoxia in the Sea Raven, Hemitripterus americanus, and an investigation of possible control mechanisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):491–503, April 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-069.

Davis:1971:HCA

[339] Robert M. Davis and Owen C. Fenderson. Histological comparisons of the adrenal-cortical cells of hatchery and wild landlocked Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4): 505–508, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-070.

Saito:1971:PBT

[340] Akira Saito and L. W. Regier. Pigmentation of brook trout (Salvelinus fontinalis) by feeding dried crustacean waste. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):509–512, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-071.

Bilton:1971:HAA

[341] H. T. Bilton. A hypothesis of alternation of age of return in successive generations of Skeena River sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):513–516, April 1971.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-072.

Evelyn:1971:FRV

[342] T. P. T. Evelyn. First records of vibriosis in Pacific salmon cultured in Canada, and taxonomic status of the responsible bacterium, Vibrio anguillarum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4): 517–525, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-073.

Waugh:1971:ULT

[343] D. L. Waugh and E. T. Garside. Upper lethal temperatures in relation to osmotic stress in the ribbed mussel Modiolus demissus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):527–532, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-074.

Fujihara:1971:CCD

[344] M. P. Fujihara and F. P. Hungate. Chondrococcus columnaris disease of fishes: Influence of Columbia River fish ladders. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):533–536, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-075.

Warner:1971:EJT

[345] Kendall Warner. Effects of jaw tagging on growth and scale characteristics of landlocked Atlantic salmon, Salmo salar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):537–542, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-076.

Davies:1971:KFH

[346] Ronald W. Davies. A key to the freshwater hirudinoidea of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):543–552, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-077.

Parsons:1971:MDB

[347] L. S. Parsons and V. M. Hodder. Meristic differences between springand autumn-spawning Atlantic herring (Clupea harengus harengus) from Southwestern Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):553–558, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-078.

Schindler:1971:SCV

[348] D. W. Schindler, A. S. Clark, and J. R. Gray. Seasonal calorific values of freshwater zooplankton, as determined with a phillipson Bomb calorimeter modified for small samples. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):559–564, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-079.

Sutterlin:1971:ERO

[349] A. M. Sutterlin and N. Sutterlin. Electrical responses of the olfactory epithelium of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):565–572, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-080.

Acara:1971:TEK

[350] A. H. Acara and H. D. Smith. A technique for enumerating Kokanee salmon (Oncorhynchus nerka) fry migrating through streams, with an appendix for processing catch data by IBM 360 Fortran IV computer programs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):573– 585, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-081.

Dickson:1971:FIS

[351] I. W. Dickson and R. H. Kramer. Factors influencing scope for activity and active and standard metabolism of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):587–596, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-082.

Anonymous:1971:Oa

[352] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):597–599, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-083.

Ackman:1971:IFA

[353] R. G. Ackman and C. A. Eaton. Investigation of the fatty acid composition of oils and lipids from the sand launce (Ammodytes americanus) from Nova Scotia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):601–606, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-084.

Wedemeyer:1971:SBC

[354] Gary Wedemeyer and K. Chatterton. Some blood chemistry values for the juvenile Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):606–608, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-085.

Hennick:1971:HSW

[355] Daniel Hennick. A hermaphroditic specimen of weathervane scallop, Patinopecten caurinus, in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):608–609, April 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-086.

Hogan:1971:CDB

[356] Roger L. Hogan and Eugene W. Roelofs. Concentrations of dieldrin in the blood and brain of the green sunfish, Lepomis cyanellus, at death. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):610–612, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-087.

Hogan:1971:SEP

[357] James W. Hogan. Some enzymatic properties of plasma esterases from channel catfish (Ictalurus punctatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):613–616, April 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-088.

Manion:1971:GAB

[358] Patrick J. Manion and Harold A. Purvis. Giant American brook lampreys, Lampetra lamottei, in the Upper Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):616–620, April 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-089.

MacAskie:1971:RES

[359] I. B. MacAskie. A range extension for the squid Abraliopsis (Watasenia) felis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):620– 621, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-090.

Anonymous:1971:RAFd

[360] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):623, April 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-091.

Anonymous:1971:AFPd

[361] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(4):624, April 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-092.

Houston:1971:SPE

[362] Arthur H. Houston, Jane A. Madden, R. John Woods, and Harry M. Miles. Some physiological effects of handling and tricaine methanesulphonate anesthetization upon the brook trout, Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):625–633, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-093.

Houston:1971:VBT

[363] Arthur H. Houston, Jane A. Madden, R. John Woods, and Harry M. Miles. Variations in the blood and tissue chemistry of brook trout, Salvelinus fontinalis, subsequent to handling, anesthesia, and surgery.

Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):635–642, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-094.

Ostovar:1971:EGI

[364] K. Ostovar, R. R. Pereira, and R. A. Gallop. Effects of gamma irradiation on Salmonella spp. in smoked lake whitefish (Coregonus clupeaformis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):643–646, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-095.

Walker:1971:RTG

[365] C. E. Walker and D. B. Lister. Results for three generations from transfers of pink salmon (Oncorhynchus gorbuscha) spawn to the Qualicum River in 1963 and 1964. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):647–654, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-096.

McKim:1971:ELT

[366] J. M. McKim and D. A. Benoit. Effects of long-term exposures to copper on survival, growth, and reproduction of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):655– 662, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-097.

MacCrimmon:1971:WDR

[367] Hugh R. MacCrimmon. World distribution of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):663– 704, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-098.

Johnson:1971:BMD

[368] B. Thomas Johnson, C. Richard Saunders, Herman O. Sanders, and Robert S. Campbell. Biological magnification and degradation of DDT and aldrin by freshwater invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):705–709, May 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-099.

Mathias:1971:EFS

[369] Jack A. Mathias. Energy flow and secondary production of the amphipods Hyalella azteca and Crangonyx richmondensis occidentalis in Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):711–726, May 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-100.

Alderdice:1971:ESTa

[370] D. F. Alderdice and C. R. Forrester. Effects of salinity and temperature on embryonic development of the petrale sole (Eopsetta jordani). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):727–744, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-101.

Ott:1971:FCC

[371] A. G. Ott and H. F. Horton. Fertilization of Chinook and Coho salmon eggs with cryo-preserved sperm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):745–748, May 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-102.

McCart:1971:GVL

[372] P. McCart and V. A. Pepper. Geographic variation in the lateral line scale counts of the Arctic grayling, Thymallus arcticus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):749–754, May 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-103.

Boulva:1971:OCW

[373] Jean Boulva. Observations on a colony of whelping harbour seals, Phoca vitulina concolor, on Sable Island, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):755–759, May 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-104.

Barica:1971:RMD

[374] Jan Barica. A rapid method of determining ammonium in water with a univalent cation glass electrode. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):759–764, May 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-105.

Baum:1971:FAS

[375] E. T. Baum and A. L. Meister. Fecundity of Atlantic salmon (Salmo salar) from two Maine rivers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):764–767, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-106.

Arai:1971:HPE

[376] Hisao P. Arai. Helminth parasites of embiotocid fishes. II. A new species of Labrifer (Digenea:Lepocreadiidae) from Zalembius rosaceus from California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):767– 769, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-107.

Skauen:1971:LSM

[377] D. M. Skauen, N. Marshall, and R. J. Fragala. A liquid scintillation method for assaying^{14}C-labelled benthic microflora. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):769–770, May 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-108.

Kwain:1971:LHR

[378] Wen hwa Kwain. Life history of rainbow trout (Salmo gairdneri) in Batchawana Bay, Eastern Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):771–775, May 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-109.

Green:1971:LCC

[379] Roger H. Green. Lipid and caloric contents of the relict amphipod Pontoporeia affinis in Cayuga Lake, New York. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):776–777, May 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-110.

Mann:1971:RBS

[380] K. H. Mann. Relation between stipe length, environment, and the taxonomic characters of Laminaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):778–780, May 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-111.

Houston:1971:SIV

[381] Arthur H. Houston. A simple improvement in the vascular catheterization procedure for salmonid and other teleost fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):781–783, May 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-112.

Gillespie:1971:EWF

[382] D. C. Gillespie and Kurosh Ostovar. Effects of washing freshwater fish on keeping quality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5): 783–785, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-113.

Uthe:1971:PSH

[383] J. F. Uthe and E. G. Bligh. Preliminary survey of heavy metal contamination of Canadian freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):786–788, May 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-114.

Ellender:1971:AAN

[384] Rudolph Dennis Ellender, Carl Lee Armour, and Bennie Joe Camp. Analysis of ammonium–nitrogen in artificial sea water aquaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):788–789, May 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-115.

Glooschenko:1971:INE

[385] Walter A. Glooschenko and Herbert Curl, Jr. Influence of nutrient enrichment on photosynthesis and assimilation ratios in natural North Pacific phytoplankton communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):790–793, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-116.

Fletcher:1971:AYP

[386] G. L. Fletcher. Accumulation of yellow phosphorus by several marine invertebrates and seaweed. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):793–796, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-117.

Anonymous:1971:RAFe

[387] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):797– 798, May 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-118.

Anonymous:1971:AFPe

[388] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(5):799, May 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-119.

Kerr:1971:ALE

[389] S. R. Kerr. Analysis of laboratory experiments on growth efficiency of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):801– 808, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-120.

Kerr:1971:PFG

[390] S. R. Kerr. Prediction of fish growth efficiency in nature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):809–814, June 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-121.

Kerr:1971:SML

[391] S. R. Kerr. A simulation model of lake trout growth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):815–819, June 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-122.

Todd:1971:GSS

[392] Ian St. P. Todd and P. A. Larkin. Gillnet selectivity on sockeye (Oncorhynchus nerka) and pink salmon (O. gorbuscha) of the Skeena River system, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):821–842, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-123.

Shah:1971:ZRY

[393] B. G. Shah, J. C. Meranger, B. Belonje, and D. S. Deshmukh. Zinc retention in young rats fed increasing levels of high-zinc oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):843–848, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-124.

Mundie:1971:SBS

[394] J. H. Mundie. Sampling benthos and substrate materials, down to 50 microns in size, in shallow streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):849–860, June 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-125.

Bilton:1971:EFL

[395] H. T. Bilton and G. L. Robins. Effects of feeding level on circulus formation on scales of young sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):861–868, June 1971. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f71-126.

Hiltz:1971:OPA

[396] Doris Fraser Hiltz and W. J. Dyer. Octopine in postmortem adductor muscle of the sea scallop (Placopecten magellanicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):869–874, June 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-127.

Jorgensen:1971:EVD

[397] P. E. Vestergård Jørgensen. Egtved virus: Demonstration of neutralizing antibodies in serum from artificially infected rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):875– 877, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-128.

Chapman:1971:ELH

[398] M. John Chapman, Christopher Chin, and Finn Wold. Enolase from lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):879–882, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-129.

Alderdice:1971:ESTb

[399] D. F. Alderdice and C. R. Forrester. Effects of salinity, temperature, and dissolved oxygen on early development of the Pacific cod (Gadus macrocephalus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6): 883–902, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-130.

Holmes:1971:HSS

[400] John C. Holmes. Habitat segregation in sanguinicolid blood flukes (Digenea) of scorpaenid rockfishes (Perciformes) on the Pacific Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6): 903–909, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-131.

Flemer:1971:PCN

[401] David A. Flemer and Robert B. Biggs. Particulate carbon : nitrogen relations in Northern Chesapeake Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):911–918, June 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-132.

MacLean:1971:ETM

[402] J. A. MacLean and John H. Gee. Effects of temperature on movements of prespawning brook sticklebacks, Culaea inconstans, in the Roseau River, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6): 919–923, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-133.

McCormick:1971:TRG

[403] J. Howard McCormick, Bernard R. Jones, and Roll F. Syrett. Temperature requirements for growth and survival of larval ciscos (Coregonus artedii). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6): 924–927, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-134.

Peden:1971:EKR

[404] Alex E. Peden. Extension of the known range of the masked greenling, Hexagrammos octogrammus, to British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):927–928, June 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-135.

Nakamura:1971:FTC

[405] Royden Nakamura. Food of two cohabiting tide-pool Cottidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):928–932, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-136.

Anonymous:1971:RAFf

[406] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):933, June 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-137.

Anonymous:1971:AFPf

[407] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(6):934, June 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-138.

Tyler:1971:PRC

[408] A. V. Tyler. Periodic and resident components in communities of Atlantic fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7): 935–946, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-139.

Muir:1971:EBP

[409] Barry S. Muir and Clinton E. Brown. Effects of blood pathway on the blood-pressure drop in fish gills, with special reference to tunas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):947–955, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-140.

Mauck:1971:VSF

[410] Wilbur L. Mauck and Daniel W. Coble. Vulnerability of some fishes to northern pike (Esox lucius) predation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):957–969, July 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-141.

Sameoto:1971:LHE

[411] D. D. Sameoto. Life history, ecological production, and an empirical mathematical model of the population of Sagitta elegans in St. Margaret’s Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):971–985, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-142.

Schreck:1971:TUK

[412] Carl B. Schreck and Robert J. Behnke. Trouts of the Upper Kern River Basin, California, with reference to systematics and evolution of Western North American Salmo. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):987–998, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-143.

Symons:1971:SDS

[413] Philip E. K. Symons. Spacing and density in schooling threespine sticklebacks (Gasterosteus aculeatus) and mummichog (Fundulus heteroclitus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):999–1004, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-144.

Clayton:1971:GMM

[414] J. W. Clayton, D. N. Tretiak, and A. H. Kooyman. Genetics of multiple malate dehydrogenase isozymes in skeletal muscle of walleye (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (7):1005–1008, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-145.

Messieh:1971:DAH

[415] S. N. Messieh and S. N. Tibbo. Discreteness of Atlantic herring (Clupea harengus harengus) populations in spring and autumn fisheries in the Southern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1009–1014, July 1971. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f71-146.

Bilinski:1971:LTL

[416] E. Bilinski, R. E. E. Jonas, and Y. C. Lau. Lysosomal triglyceride lipase from the lateral line tissue of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1015–1018, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-147.

Jenkins:1971:RSB

[417] Thomas M. Jenkins, Jr. Role of social behavior in dispersal of introduced rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1019–1027, July 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-148.

Winters:1971:FLR

[418] G. H. Winters. Fecundity of the left and right ovaries of Grand Bank capelin (Mallotus villosus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1029–1033, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-149.

MacNeill:1971:QSM

[419] I. B. MacNeill. Quick statistical methods for analysing the sequences of fish counts provided by digital echo counters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1035–1042, July 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-150.

Murray:1971:DNA

[420] D. Murray and D. Povoledo. Determination of nitrilotriacetic acid in inland waters by gas chromatography of the trimethyl ester. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1043–1047, July 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-151.

Creaser:1971:IWS

[421] Edwin P. Creaser, Jr. An interstitial water-sampling receptacle for intertidal mud flats. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (7):1049–1051, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-152.

Banerjee:1971:MFW

[422] Tapan Banerjee. Maturity and fecundity of the white seaperch (Phanerodon furcatus) from Tomales Bay, California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1051–1053, July 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-153.

Edmunds:1971:GPT

[423] Philip H. Edmunds and James I. Sammons. Genic polymorphism of tetrazolium oxidase in bluefin tuna, Thunnus thynnus, from the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7): 1053–1055, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-154.

Ke:1971:DFS

[424] Paul J. Ke and Lloyd W. Regier. Determination of fluoride in soils by a nonfusion distillation pF method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1055–1056, July 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-155.

Marcy:1971:SYF

[425] Barton C. Marcy, Jr. Survival of young fish in the discharge canal of a nuclear power plant. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28 (7):1057–1060, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-156.

Lantz:1971:PMD

[426] A. W. Lantz and D. G. Iredale. A practical method for drying freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1061– 1062, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-157.

Scrivener:1971:SDP

[427] J. C. Scrivener and L. D. B. Terhune. A sensor for detecting the presence of a lobster (Homarus americanus) in a burrow. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1062–1063, July 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-158.

Frantsi:1971:IPN

[428] C. Frantsi and M. Savan. Infectious pancreatic necrosis virus: Comparative frequencies of isolation from feces and organs of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7): 1064–1065, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f71-159.

Anonymous:1971:RAFg

[429] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1067– 1068, July 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-160.

Anonymous:1971:AFPg

[430] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(7):1069, July 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-161.

Boyd:1971:ESC

[431] J. W. Boyd and B. A. Southcott. Effects of sodium chloride on outgrowth and toxin production of Clostridium botulinum type E in cod homogenates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8): 1071–1075, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-162.

Pickard:1971:SPO

[432] G. L. Pickard. Some physical oceanographic features of inlets of Chile. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1077–1106, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-163.

Waller:1971:CSE

[433] William T. Waller, Richard E. Sparks, John Cairns, Jr., and M. L. Dahlberg. A computer simulation of the effects of superimposed mortality due to pollutants on populations of fathead minnows (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1107–1112, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-164.

Ringrose:1971:CPR

[434] R. C. Ringrose. Calorie-to-protein ratio for brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1113–1117, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-165.

Brungs:1971:CEL

[435] William A. Brungs. Chronic effects of low dissolved oxygen concentrations on the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1119–1123, August 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-166.

Dingle:1971:DIM

[436] J. R. Dingle and J. A. Hines. Degradation of inosine 5’-monophosphate in the skeletal muscle of several North Atlantic fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1125–1131, August 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-167.

Ferguson:1971:MAJ

[437] R. G. Ferguson and A. J. Derksen. Migrations of adult and juvenile walleyes (Stizostedion vitreum vitreum) in Southern Lake Huron, Lake St. Clair, Lake Erie, and connecting waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1133–1142, August 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-168.

Kabata:1971:LMC

[438] Z. Kabata and G. C. Hewitt. Locomotory mechanisms in Caligidae (Crustacea: Copepoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8): 1143–1151, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-169.

Kennedy:1971:WFP

[439] V. S. Kennedy and D. H. Steele. The winter flounder (Pseudopleuronectes americanus) in Long Pond, Conception Bay, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1153–1165, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-170.

Leggett:1971:NAB

[440] William C. Leggett and Robert A. Jones. Net avoidance behavior in American shad (Alosa sapidissima) as observed by ultrasonic tracking techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1167–1171, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-171.

Qureshi:1971:LCG

[441] M. D. Qureshi, R. V. Hledin, P. A. Anastassiadis, and W. E. Vanstone. Levels of constituents of glycoproteins in the sera of Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1173–1179, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-172.

Welander:1971:GFR

[442] Arthur D. Welander, Gerald W. Wadley, and Douglas K. Dysart. Growth and fecundity of rainbow trout (Salmo gairdneri) exposed to single sublethal doses of X-rays during the eyed embryo stage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1181–1184, August 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-173.

Trust:1971:BCC

[443] T. J. Trust. Bacterial counts of commercial fish diets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1185–1189, August 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-174.

Allen:1971:AAF

[444] W. V. Allen. Amino acid and fatty acid composition of tissues of the Dungeness crab (Cancer magister). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1191–1195, August 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-175.

Scott:1971:MFT

[445] K. R. Scott. Monitoring of fish tank oxygen concentration using a digital recorder. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1196–1197, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-176.

Boddington:1971:SCO

[446] M. J. Boddington and D. F. Mettrick. Supernumerary copulatory organs in Dugesia tigrina (Girard 1850) (Platyhelminthes: Turbellaria). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1198–1199, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-177.

Lear:1971:PCB

[447] W. H. Lear and A. W. May. Paralepis coregonoides borealis (Osteichthyes: Paralepididae) from Davis Strait and Labrador Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1199–1203, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-178.

Uthe:1971:SFT

[448] J. F. Uthe. A simple field technique for obtaining small samples of muscle from living fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8): 1203–1204, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-179.

Bourget:1971:TSD

[449] Edwin Bourget and Guy Lacroix. Two simple durable epifaunal collectors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1205–1207, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-180.

Metroka:1971:KH

[450] C. E. Metroka, J. V. Chuba, W. J. Kuhns, R. F. Nigrelli, and G. D. Ruggieri. Killifish “hemagglutinogens”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1208–1210, August 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-181.

Silliman:1971:ALS

[451] R. P. Silliman. Advantages and limitations of “simple” fishery models in light of laboratory experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1211–1214, August 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-182.

Dominy:1971:CBL

[452] C. L. Dominy. Changes in blood lactic acid concentrations in alewives (Alosa pseudoharengus) during passage through a pool and weir fishway. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1215–1217, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-183.

Freese:1971:CGC

[453] M. Freese. Comment: Gould’s the case for accurate labeling of superchilled fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1217–1218, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-184.

Anonymous:1971:CRP

[454] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1219, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-185.

Anonymous:1971:RAFh

[455] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1221–1222, August 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-186.

Anonymous:1971:AFPh

[456] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(8):1223, August 1971. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f71-187.

Eisler:1971:CPF

[457] Ronald Eisler. Cadmium poisoning in Fundulus heteroclitus (Pisces: Cyprinodontidae) and other marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1225–1234, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-188.

Lindsey:1971:EMO

[458] C. C. Lindsey and M. Y. Ali. An experiment with medaka, Oryzias latipes, and a critique of the hypothesis that teleost egg size controls vertebral count. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9): 1235–1240, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-189.

McArn:1971:CST

[459] G. E. McArn and S. R. Wellings. A comparison of skin tumors in three species of flounders. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9): 1241–1251, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-190.

Fujihara:1971:API

[460] M. P. Fujihara and R. E. Nakatani. Antibody production and immune responses of rainbow trout and Coho salmon to Chondrococcus columnaris. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1253–1258, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-191.

Andrusak:1971:SBA

[461] H. Andrusak and T. G. Northcote. Segregation between adult cutthroat trout (Salmo clarki) and Dolly Varden (Salvelinus malma) in small coastal British Columbia lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1259–1268, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-192.

Thorne:1971:IRB

[462] Richard E. Thorne. Investigations into the relation between integrated echo voltage and fish density. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1269–1273, September 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-193.

Thorne:1971:EPH

[463] Richard E. Thorne, Jerry E. Reeves, and Alan E. Millikan. Estimation of the Pacific hake (Merluccius productus) population in Port Susan, Washington, using an echo integrator. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1275–1284, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-194.

Zitko:1971:MFM

[464] V. Zitko, B. J. Finlayson, D. J. Wildish, J. M. Anderson, and A. C. Kohler. Methylmercury in freshwater and marine fishes in New Brunswick, in the Bay of Fundy, and on the Nova Scotia Banks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1285–1291, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-195.

Moose:1971:EVA

[465] Paul H. Moose and John E. Ehrenberg. An expression for the variance of abundance estimates using a fish echo integrator. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1293–1301, September 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-196.

Scherer:1971:EOD

[466] E. Scherer. Effects of oxygen depletion and of carbon dioxide buildup on the photic behavior of the walleye (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1303–1307, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-197.

Brodie:1971:RAG

[467] P. F. Brodie. A reconsideration of aspects of growth, reproduction, and behavior of the white whale (Delphinapterus leucas), with reference to the Cumberland Sound, Baffin Island, population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1309–1318, September 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-198.

Dill:1971:PPL

[468] P. A. Dill. Perception of polarized light by yearling sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9): 1319–1322, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-199.

Mackie:1971:QSA

[469] G. L. Mackie and S. U. Qadri. A quantitative sampler for aquatic phytomacrofauna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):

1322–1324, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-200.

Manohar:1971:PCG

[470] S. V. Manohar and Helen Boese. Postmortem changes in the glycogen phosphorylase activity of the muscle of white sucker (Catostomus commersoni) and northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1325–1326, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-201.

Conover:1971:SRB

[471] R. J. Conover. Some relations between zooplankton and bunker C oil in Chedabucto Bay following the wreck of the tanker Arrow. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1327–1330, September 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-202.

Segerstraale:1971:ZPI

[472] Sven G. Segerstråle. The zoogeographic problem involved in the presence of the glacial relict Pontoporeia affinis (Crustacea Amphipoda) in Lake Washington, USA. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9): 1331–1334, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-203.

Haynes:1971:AGG

[473] Evan B. Haynes and Charles R. Hitz. Age and growth of the giant Pacific sea scallop, Patinopecten caurinus, from the Strait of Georgia and outer Washington Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9): 1335–1341, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-204.

Kutty:1971:MST

[474] M. N. Kutty, N. V. Karuppannan, M. Narayanan, and M. Peer Mohamed. Maros-schulek technique for measurement of carbon dioxide production in fish and respiratory quotient in Tilapia mossambica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1342–1344, September 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-205.

Barraclough:1971:AOC

[475] W. E. Barraclough and D. G. Robinson. Anomalous occurrence of carp (Cyprinus carpio) in the marine environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1345–1347, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-206.

Peden:1971:FRB

[476] Alex E. Peden and C. G. Gruchy. First record of the bluespotted poacher, Xeneretmus triacanthus, in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1347–1348, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-207.

Stein:1971:NMC

[477] David Stein and John Butler. A notacanthid Macdonaldia challengeri collected off the Oregon Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1349–1350, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-208.

Sonstegard:1971:IPN

[478] R. A. Sonstegard and L. A. McDermott. Infectious pancreatic necrosis of salmonids in Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9): 1350–1351, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-209.

Patten:1971:IPT

[479] Benjamin G. Patten. Increased predation by the torrent sculpin, Cottus rhotheus, on Coho salmon fry, Oncorhynchus kisutch, during moonlight nights. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1352–1354, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-210.

Anonymous:1971:RAFi

[480] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1355–1356, September 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-211.

Anonymous:1971:AFPi

[481] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(9):1357, September 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-212.

Anonymous:1971:ECB

[482] Anonymous. Edith and cyril Berkeley — an appreciation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1359–1364, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-213.

Arai:1971:PEC

[483] Mary Needler Arai. Publications by Edith and/or Cyril Berkeley. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1365–1372, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-214.

Zibrowius:1971:RMB

[484] Helmut W. Zibrowius. Revision of Metavermilia Bush (Polychaeta, Serpulidae), with descriptions of three new species from off Portugal, Gulf of Guinea, and Western Indian Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1373–1383, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-215.

Margolis:1971:PIH

[485] L. Margolis. Polychaetes as intermediate hosts of helminth parasites of vertebrates: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1385–1392, October 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-216.

Pettibone:1971:DSF

[486] Marian H. Pettibone. Descriptions of Sthenelais fusca Johnson 1897 and S. berkeleyi n. sp. (Polychaeta: Sigalionidae) from the Eastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1393–1401, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-217.

Uschakov:1971:ADP

[487] P. V. Uschakov. Amphipacific distribution of polychaetes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1403–1406, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-218.

Hartman:1971:APA

[488] Olga Hartman. Abyssal polychaetous annelids from the Mozambique Basin off Southeast Africa, with a compendium of abyssal polychaetous annelids from world-wide areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1407–1428, October 1971. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f71-219.

Levenstein:1971:NPS

[489] R. Ya. Levenstein. A new polychaete species of the genus Macellicephaloides from the Aleutian Trench. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1429–1431, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-220.

Brown:1971:RBT

[490] P. L. Brown and D. V. Ellis. Relation between tube-building and feeding in Neoamphitrite robusta (Polychaeta: Terebellidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1433–1435, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-221.

Knox:1971:DBS

[491] G. A. Knox. Dodecaceria berkeleyi n. sp., a polychaete (Family Cirrutulidae) from New Zealand. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1437–1443, October 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-222.

Jones:1971:MBS

[492] Meredith L. Jones. Magelona berkeleyi n. sp. from Puget Sound (Annelida: Polychaeta), with a further redescription of Magelona longicornis Johnson and a consideration of recently described species of Magelona. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1445–1454, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-223.

Foster:1971:RSP

[493] Nancy M. Foster. Redescription of the spionid polychaete Malacoceros (Malacoceros) indicus (Fauvel 1928). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1455–1457, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-224.

Reish:1971:SSP

[494] Donald J. Reish. Seasonal settlement of polychaetous annelids on test panels in Los Angeles–Long Beach Harbors 1950–1951. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1459–1467, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-225.

Banse:1971:NSA

[495] Karl Banse. A new species, and additions to the descriptions of six other species of Syllides Örsted (Syllidae: Polychaeta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1469–1481, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-226.

Laubier:1971:SBS

[496] Lucien Laubier. Schroederella berkeleyi sp. n., un nouvel Orbiniidé (Polychète Sédentaire) de la faune interstitielle littorale du Massachusetts (USA). (French) [Schroederella berkeleyi sp. n., a new Orbiniid (polychete sedentary) from the interstitial littoral fauna of Massachusetts (USA)]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1483–1486, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-227.

Dales:1971:BPP

[497] R. Phillips Dales. Bioluminescence in pelagic polychaetes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1487–1489, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-228.

Nichols:1971:TNR

[498] Frederic H. Nichols. Two new records of benthic polychaetes from Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1491–1492, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-229.

Larkin:1971:SSA

[499] P. A. Larkin. Simulation studies of the Adams River sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10): 1493–1502, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-230.

Parker:1971:SSP

[500] Robert R. Parker. Size selective predation among juvenile salmonid fishes in a British Columbia inlet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1503–1510, October 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-231.

Bell:1971:ACI

[501] Gordon R. Bell, Gary E. Hoskins, and W. Hodgkiss. Aspects of the characterization, identification, and ecology of the bacterial flora associated with the surface of stream-incubating Pacific salmon (Oncorhynchus) eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1511–1525, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-232.

Hart:1971:NDR

[502] Josephine F. L. Hart. New distribution records of reptant decapod crustacea, including descriptions of three new species of Pagurus, from the waters adjacent to British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1527–1544, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-233.

Alderdice:1971:SES

[503] D. F. Alderdice and F. P. J. Velsen. Some effects of salinity and temperautre on early development of Pacific herring (Clupea pallasi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1545–1562, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-234.

Kabata:1971:FBC

[504] Z. Kabata. Four Bomolochidae (Copepoda) from fishes of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1563–1572, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-235.

Allen:1971:RBP

[505] K. Radway Allen. Relation between production and biomass. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1573–1581, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-236.

Calaprice:1971:RXR

[506] J. R. Calaprice, H. M. McSheffrey, and L. A. Lapi. Radioisotope X-ray fluorescence spectrometry in aquatic biology: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1583–1594, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-237.

Fields:1971:ECE

[507] W. Gordon Fields and G. O. Mackie. Evolution of the chondrophora: Evidence from behavioural studies on Velella. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1595–1602, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-238.

Tarr:1971:RRI

[508] H. L. A. Tarr and Linda Gardner. Ribonucleotide reductase in immature testes of salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10): 1603–1608, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-239.

Davis:1971:CVR

[509] John C. Davis. Circulatory and ventilatory responses of rainbow trout (Salmo gairdneri) to artificial manipulation of gill surface area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1609–1614, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-240.

Butler:1971:EBS

[510] T. H. Butler. Eualus berkeleyorum n. sp., and records of other caridean shrimps (order Decapoda) from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1615–1620, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-241.

Westrheim:1971:TDB

[511] S. J. Westrheim and H. Tsuyuki. Taxonomy, distribution, and biology of the northern rockflsh, Sebastes polyspinis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1621–1627, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-242.

Evelyn:1971:ASB

[512] T. P. T. Evelyn. An aberrant strain of the bacterial fish pathogen Aeromonas salmonicida isolated from a marine host, the sablefish (Anoplopoma fimbria), and from two species of cultured Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1629–1634, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-243.

Brett:1971:GRY

[513] J. R. Brett. Growth responses of young sockeye salmon (Oncorhynchus nerka) to different diets and planes of nutrition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1635–1643, October 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-244.

Moravec:1971:NCA

[514] F. Moravec and Hisao P. Arai. The North and Central American species of Rhabdochona Railliet 1916 (Nematoda: Rhabdochonidae) of fishes, including Rhabdochona canadensis sp. nov. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1645–1662, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-245.

Taylor:1971:ODP

[515] F. H. C. Taylor. Offshore distribution of Pacific herring (Clupea pallasi) on basis of records from fur seal stomachs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1663–1665, October 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-246.

Ricker:1971:DBM

[516] W. E. Ricker. Derzhavin’s biostatistical method of population analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1666–1672, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-247.

Forrester:1971:APS

[517] C. R. Forrester and M. S. Smith. Ambicoloration in a petrale sole (Eopsetta jordani). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10): 1672–1674, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-248.

Idler:1971:SBT

[518] D. R. Idler and P. Wiseman. Sterols of a brachiopod, Terebratalia transversa. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1675–1676, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-249.

Arai:1971:PCB

[519] Mary Needler Arai. Pachycerianthus (Ceriantharia) from British Columbia and Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1677–1680, October 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-250.

Barraclough:1971:FRW

[520] W. E. Barraclough and R. M. Wilson. First record of whitebait smelt (Allosmerus elongatus) from Juan de Fuca Strait, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(10):1681, October 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f71-251.

Johnson:1971:ASD

[521] M. G. Johnson and R. O. Brinkhurst. Associations and species diversity in benthic macroinvertebrates of Bay of Quinte and Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1683–1697, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-252.

Johnson:1971:PBM

[522] M. G. Johnson and R. O. Brinkhurst. Production of benthic macroinvertebrates of Bay of Quinte and Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1699–1714, November 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-253.

Johnson:1971:BCM

[523] M. G. Johnson and R. O. Brinkhurst. Benthic community metabolism in Bay of Quinte and Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1715–1725, November 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-254.

Tyler:1971:SWF

[524] A. V. Tyler. Surges of winter flounder, Pseudopleuronectes americanus, into the intertidal zone. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1727–1732, November 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-255.

Miller:1971:PPS

[525] R. J. Miller, K. H. Mann, and D. J. Scarratt. Production potential of a seaweed–lobster community in Eastern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1733–1738, November 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-256.

Fujihara:1971:SFS

[526] M. P. Fujihara, P. A. Olson, and R. E. Nakatani. Some factors in susceptibility of juvenile rainbow trout and Chinook salmon to Chondrococcus columnaris. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11): 1739–1743, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-257.

Beamish:1971:BCS

[527] R. J. Beamish and H. Tsuyuki. A biochemical and cytological study of the longnose sucker (Catostomus catostomus) and large and dwarf forms of the white sucker (Catostomus commersoni). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1745–1748, November 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-258.

Bilton:1971:ESF

[528] H. T. Bilton and G. L. Robins. Effects of starvation, feeding, and light period on circulus formation on scales of young sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11): 1749–1755, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-259.

Bilton:1971:RYS

[529] H. T. Bilton and G. L. Robins. Response of young sockeye salmon (Oncorhynchus nerka) to prolonged periods of starvation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1757–1761, November 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-260.

Schindler:1971:ELE

[530] D. W. Schindler, F. A. J. Armstrong, S. K. Holmgren, and G. J. Brunskill. Eutrophication of Lake 227, Experimental Lakes Area, Northwestern Ontario, by addition of phosphate and nitrate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1763–1782, November 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-261.

Cross:1971:DRS

[531] Ford A. Cross, James N. Willis, and John P. Baptist. Distribution of radioactive and stable zinc in an experimental marine ecosystem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1783–1788, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-262.

Price:1971:IBP

[532] R. J. Price and J. S. Lee. Interaction between paralytic shellfish poison and melanin obtained from butter clam (Saxidomus giganteus) and synthetic melanin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11): 1789–1792, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-263.

Odense:1971:EAN

[533] P. H. Odense and H. Brockerhoff. Effects of adaptive and nonadaptive feeding of hydrogenated and raw herring oil on hearts and livers of weanling rats. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1793–1795, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-264.

Fields:1971:PDU

[534] W. Gordon Fields and Veronica A. Gauley. Preliminary description of an unusual gonatid squid (Cephalopoda: Oegopsida) from the North Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1796–1801, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-265.

McCauley:1971:TSR

[535] R. W. McCauley and W. L. Pond. Temperature selection of rainbow trout (Salmo gairdneri) fingerlings in vertical and horizontal gradients. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1801–1804, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-266.

Symons:1971:EDB

[536] Philip E. K. Symons. Estimating distances between fish schooling in an aquarium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1805–1806, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-267.

Gillespie:1971:MMS

[537] D. C. Gillespie and D. P. Scott. Mobilization of mercuric sulfide from sediment into fish under aerobic conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1807–1808, November 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-268.

McCosker:1971:NRL

[538] John E. McCosker and Ross F. Nigrelli. New records of lymphocystis disease in four Eastern Pacific fish species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1809–1810, November 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-269.

McCarraher:1971:SSF

[539] D. B. McCarraher. Survival of some freshwater fishes in the alkaline eutrophic waters of Nebraska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1811–1814, November 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-270.

Duedall:1971:DWD

[540] I. W. Duedall, A. R. Coote, D. F. Knox, and G. F. Connolly. A dispenser for Winkler’s dissolved oxygen reagents. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1815–1816, November 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-271.

Strasdine:1971:APA

[541] G. A. Strasdine and Joanne M. Melville. Acid phosphatase activity in vegetative cells and spores of Clostridium botulinum type E. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1817–1820, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-272.

Anonymous:1971:RAFj

[542] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1821, November 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-273.

Anonymous:1971:AFPj

[543] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(11):1822, November 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-274.

Chittenden:1971:EHS

[544] Mark E. Chittenden, Jr. Effects of handling and salinity on oxygen requirements of the striped bass, Morone saxatilis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1823–1830, December 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-275.

Hubbard:1971:GDR

[545] Lyle T. Hubbard, Jr. and William G. Pearcy. Geographic distribution and relative abundance of salpidae off the Oregon Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1831–1836, December 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-276.

Jonas:1971:SEL

[546] R. E. E. Jonas. The solubilizing effect of lysolecithin on the extractability of various tissues of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1837–1840, December 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-277.

Arthur:1971:CTA

[547] John W. Arthur and John G. Eaton. Chloramine toxicity to the amphipod Gammarus pseudolimnaeus and the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1841–1845, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-278.

Ware:1971:PRT

[548] D. M. Ware. Predation by rainbow trout (Salmo gairdneri): the effect of experience. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12): 1847–1852, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-279.

Markert:1971:EPC

[549] J. R. Markert and W. E. Vanstone. Egg proteins of Coho salmon (Oncorhynchus kisutch): Chromatographic separation and molecular weights of the major proteins in the high density fraction and their presence in salmon plasma. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12): 1853–1856, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-280.

Collins:1971:IKS

[550] John J. Collins. Introduction of Kokanee salmon (Oncorhynchus nerka) into Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12): 1857–1871, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-281.

Mansfield:1971:OBG

[551] A. W. Mansfield. Occurrence of the bowhead or Greenland right whale (Balaena mysticetus) in Canadian Arctic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1873–1875, December 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-282.

Dimond:1971:APD

[552] J. B. Dimond, A. S. Getchell, and J. A. Blease. Accumulation and persistence of DDT in a lotic ecosystem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1877–1882, December 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-283.

Harvey:1971:PCL

[553] H. H. Harvey and J. F. Coombs. Physical and chemical limnology of the lakes of Manitoulin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1883–1897, December 1971. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f71-284.

Wedemeyer:1971:SFT

[554] Gary Wedemeyer. The stress of formalin treatments in rainbow trout (Salmo gairdneri) and Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1899–1904, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f71-285.

Anonymous:1971:Ob

[555] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1905, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-286.

Voyer:1971:FAR

[556] R. A. Voyer and G. E. Morrison. Factors affecting respiration rates of winter flounder (Pseudopleuronectes americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1907–1911, December 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-287.

Gordon:1971:CBC

[557] Donald C. Gordon, Jr. and Paula A. Michalik. Concentration of bunker C fuel oil in the waters of Chedabucto Bay, April 1971. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1912–1914, December 1971. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f71-288.

Ott:1971:FST

[558] A. G. Ott and H. F. Horton. Fertilization of steelhead trout (Salmo gairdneri) eggs with cryo-preserved sperm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1915–1918, December 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-289.

Aulstad:1971:ABC

[559] Dag Aulstad and Arne Kittelsen. Abnormal body curvatures of rainbow trout (Salmo gairdneri) inbred fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1918–1920, December 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-290.

Ito:1971:RPH

[560] Jun Ito and R. R. Parker. A record of Pacific herring (Clupea harengus pallasi) feeding on juvenile Chinook salmon (Oncorhynchus tshawytscha) in a British Columbia estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1921, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-291.

Barraclough:1971:SIS

[561] W. E. Barraclough. A sculpin (Icelus spiniger) new to the coastal waters of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12): 1922–1924, December 1971. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f71-292.

Anonymous:1971:IP

[562] Anonymous. Index to publications for 1971. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 28(12):1925–2056, December 1971. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f71-293.

Peden:1972:RDR

[563] Alex E. Peden. Redescription and distribution of the rare deep-sea eel Xenomystax atrarius in the Eastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):1–12, January 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-001.

Donaldson:1972:FPS

[564] Edward M. Donaldson, James D. Funk, F. C. Withler, and R. B. Morley. Fertilization of pink salmon (Oncorhynchus gorbuscha) ova by spermatozoa from gonadotropin-injected juveniles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):13–18, January 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-002.

Bell:1972:DSE

[565] W. H. Bell. Determination of the spectra of empirical wave force coefficients. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):19–25, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-003.

Hatfield:1972:ETI

[566] C. T. Hatfield and J. M. Anderson. Effects of two insecticides on the vulnerability of Atlantic salmon (Salmo salar) Parr to brook trout (Salvelinus fontinalis) predation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):27–29, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-004.

Evans:1972:AAA

[567] D. Evans and John G. Stockner. Attached algae on artificial and natural substrates in Lake Winnipeg, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):31–44, January 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-005.

Kleerekoper:1972:OGC

[568] H. Kleerekoper, G. F. Westlake, J. H. Matis, and P. J. Gensler. Orientation of goldfish (Carassius auratus) in response to a shallow gradient of a sublethal concentration of copper in an open field. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):45–54, January 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-006.

Rashid:1972:CCH

[569] M. A. Rashid and A. Prakash. Chemical characteristics of humic compounds isolated from some decomposed marine algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):55–60, January 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-007.

Amend:1972:VAT

[570] Donald F. Amend and John P. Pietsch. Virucidal activity of two iodophors to salmonid viruses. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):61–65, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-008.

Muir:1972:OCE

[571] B. S. Muir and A. J. Niimi. Oxygen consumption of the euryhaline fish aholehole (Kuhlia sandvicensis) with reference to salinity, swimming, and food consumption. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29 (1):67–77, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-009.

Carter:1972:DAP

[572] John C. H. Carter. Distribution and abundance of planktonic crustacea in Sturgeon Bay and Shawanaga Inlet, Georgian Bay, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):79–83, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-010.

Wojtowicz:1972:CSM

[573] M. B. Wojtowicz and P. H. Odense. Comparative study of the muscle catheptic activity of some marine species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):85–90, January 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-011.

Everest:1972:HSS

[574] F. H. Everest and D. W. Chapman. Habitat selection and spatial interaction by juvenile Chinook salmon and steelhead trout in two Idaho streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):91–100, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-012.

Anonymous:1972:Oa

[575] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):101, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-013.

Thomson:1972:EWI

[576] Richard E. Thomson. An explanation for the warmwater intrusion off the British Columbia Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):103–107, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-014.

Phinney:1972:OBP

[577] Duane E. Phinney. Occurrence of the Bering poacher (Occella dodacaedron) and the Pacific staghorn sculpin (Leptocottus armatus) near Chignik, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1): 107–108, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-015.

Davis:1972:IPT

[578] John C. Davis. An infrared photographic technique useful for studying vascularization of fish gills. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):109–111, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-016.

Orlowski:1972:DCT

[579] Stanley J. Orlowski, Sidney S. Herman, Richard G. Malsberger, and Hayden N. Pritchard. Distinguishing cunner and tautog eggs by immunodiffusion. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1): 111–112, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-017.

Worthington:1972:FAC

[580] R. E. Worthington, T. S. Boggess, Jr., and E. K. Heaton. Fatty acids of channel catfish (Ictalurus punctatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):113–115, January 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-018.

Anonymous:1972:RAFa

[581] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):117, January 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-019.

Anonymous:1972:AFPa

[582] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(1):118, January 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-020.

Gee:1972:AVS

[583] John H. Gee. Adaptive variation in swimbladder length and volume in dace, genus Rhinichthys. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):119–127, February 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-021.

Fraser:1972:RPB

[584] J. M. Fraser. Recovery of planted brook trout, splake, and rainbow trout from selected Ontario lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):129–142, February 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-022.

Ennis:1972:GPM

[585] G. P. Ennis. Growth per moult of tagged lobsters (Homarus americanus) in Bonavista Bay, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):143–148, February 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-023.

Wolf:1972:FVH

[586] Ken Wolf, Roger L. Herman, and C. P. Carlson. Fish viruses: Histopathologic changes associated with experimental channel catfish virus disease. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):149–150, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-024.

Clarke:1972:OMS

[587] W. Craig Clarke and H. D. Smith. Observations on the migration of sockeye salmon fry (Oncorhynchus nerka) in the Lower Babine River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):151–159, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-025.

Scarratt:1972:BRC

[588] D. J. Scarratt and Robert Lowe. Biology of rock crab (Cancer irroratus) in Northumberland Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):161–166, February 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-026.

Zaugg:1972:CGA

[589] W. S. Zaugg and L. R. McLain. Changes in gill adenosinetriphosphatase activity associated with Parr–smolt transformation in steelhead trout, Coho, and spring Chinook salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):167–171, February 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-027.

Fujihara:1972:SDC

[590] M. P. Fujihara and F. P. Hungate. Seasonal distribution of Chondrococcus columnaris infection in river fishes as determined by specific agglutinins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):173–178, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-028.

Nyman:1972:DAS

[591] O. L. Nyman and J. H. C. Pippy. Differences in Atlantic salmon, Salmo salar, from North America and Europe. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):179–185, February 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-029.

Healey:1972:BSG

[592] M. C. Healey. Bioenergetics of a sand goby (Gobius minutus) population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):187–194, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-030.

Cullimore:1972:ISM

[593] D. R. Cullimore and A. McCann. Initial studies on a method of algal assay for nutrient parameters in water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):195–198, February 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-031.

Gilderhus:1972:ETN

[594] Philip A. Gilderhus. Exposure times necessary for antimycin and rotenone to eliminate certain freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):199–202, February 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-032.

Stevens:1972:CBW

[595] E. Don Stevens. Change in body weight caused by handling and exercise in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):202–203, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-033.

Anderson:1972:VPR

[596] Douglas P. Anderson. Virulence and persistence of rough and smooth forms of Aeromonas salmonicida inoculated into Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2): 204–206, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-034.

Kennedy:1972:PSS

[597] W. A. Kennedy. Preliminary study of sablefish culture, a potential new industry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):207–210, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-035.

Lewis:1972:DAA

[598] D. H. Lewis and N. L. Savage. Detection of antibodies to Aeromonas liquefaciens in fish by an indirect fluorescent antibody technique. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):211–212, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-036.

Figueira:1972:OEB

[599] Armando J. G. Figueira. Occurrence of Eukrohnia bathypelagica Alvariño 1962 (Chaetognatha) in the Atlantic waters of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):213–214, February 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-037.

Anonymous:1972:RAFb

[600] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):215, February 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-038.

Anonymous:1972:AFPb

[601] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(2):216, February 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-039.

Calder:1972:SAH

[602] Dale R. Calder. Some athecate hydroids from the shelf waters of Northern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):217– 228, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-040.

Barsdate:1972:EVA

[603] Robert J. Barsdate and Richard C. Dugdale. Effects of volcanic ashfalls on chemical and sediment characteristics of two Alaskan lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):229–236, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-041.

Aulstad:1972:GES

[604] Dag Aulstad, Trygve Gjedrem, and Harald Skjervold. Genetic and environmental sources of variation in length and weight of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29 (3):237–241, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-042.

Boulva:1972:MTS

[605] Jean Boulva. Morphometries of three sympatric Arctic codfishes of the genera Arctogadus and Gadus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):243–249, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-043.

Griffiths:1972:EAA

[606] J. S. Griffiths and D. F. Alderdice. Effects of acclimation and acute temperature experience on the swimming speed of juvenile Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):251–264, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-044.

Griffith:1972:CBH

[607] J. S. Griffith, Jr. Comparative behavior and habitat utilization of brook trout (Salvelinus fontinalis) and cutthroat trout (Salmo clarki) in small streams in Northern Idaho. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):265–273, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-045.

Dechtiar:1972:PFL

[608] Alex O. Dechtiar. Parasites of fish from Lake of the Woods, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):275–283, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-046.

OConnell:1972:IBF

[609] Charles P. O’Connell. The interrelation of biting and filtering in the feeding activity of the northern anchovy (Engraulis mordax). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):285–293, March 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-047.

Bilton:1972:IMB

[610] H. T. Bilton. Identification of major British Columbia and Alaska runs of even-year and odd-year pink salmon from scale characters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):295–301, March 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-048.

McKeown:1972:PGH

[611] B. A. McKeown and A. P. van Overbeeke. Prolactin and growth hormone concentrations in the serum and pituitary gland of adult migratory sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):303–309, March 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-049.

Heyl:1972:RCA

[612] Henry L. Heyl and Stanley J. Carpenter. Reversible changes in adrenal cortical cell morphology and plasma hydroxycorticosteroids during freshwater portion of the spawning journey of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):311–314, March 1972.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-050.

Hatfield:1972:EFI

[613] C. T. Hatfield and P. H. Johansen. Effects of four insecticides on the ability of Atlantic salmon Parr (Salmo salar) to learn and retain a simple conditioned response. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3): 315–321, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-051.

Collins:1972:UCR

[614] C. A. Collins and D. A. Healey. Use of continuous recording salinity– temperature–pressure instrumentation at Ocean Station P (50N, 145W). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):323–327, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-052.

Smith:1972:PEF

[615] Charlie E. Smith and Robert G. Piper. Pathological effects in formalintreated rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):328–329, March 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-053.

Gee:1972:ENL

[616] John H. Gee and Kazimierz Machniak. Ecological notes on a lakedwelling population of longnose dace (Rhinichthys cataractae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):330–332, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-054.

Newman:1972:AMB

[617] Joseph T. Newman, Jr., Benjamin J. Cosenza, and John D. Buck. Aerobic microflora of the bluefish (Pomatomus saltatrix) intestine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):333–336, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-055.

Forrester:1972:FRS

[618] C. R. Forrester, Alex E. Peden, and R. M. Wilson. First records of the striped bass, Morone saxatilis, in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):337–339, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-056.

White:1972:SFT

[619] W. J. White and R. J. Beamish. A simple fish tag suitable for longterm marking experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):339–341, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-057.

Ennis:1972:DOP

[620] G. P. Ennis. A diver-operated plankton collector. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):341–343, March 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-058.

McAllister:1972:DSM

[621] D. E. McAllister and J. C. Ward. The deepwater sculpin, Myoxocephalus quadricornis thompsoni, new to Alberta, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):344–345, March 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-059.

Anonymous:1972:RAFc

[622] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):347, March 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-060.

Anonymous:1972:AFPc

[623] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(3):348, March 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-061.

Addison:1972:ROP

[624] R. F. Addison, M. E. Zinck, and R. G. Ackman. Residues of organochlorine pesticides and polychlorinated biphenyls in some commercially produced Canadian marine oils. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):349–355, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-062.

Sutcliffe:1972:SRL

[625] W. H. Sutcliffe, Jr. Some relations of land drainage, nutrients, particulate material, and fish catch in two Eastern Canadian bays. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):357–362, April 1972. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f72-063.

Monakov:1972:RSF

[626] A. V. Monakov. Review of studies on feeding of aquatic invertebrates conducted at the Institute of Biology of Inland Waters, Academy of Science, USSR. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4): 363–383, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-064.

Idler:1972:MSR

[627] D. R. Idler and P. Wiseman. Molluscan sterols: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):385–398, April 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-065.

Brocksen:1972:PRT

[628] Robert W. Brocksen and Ronald E. Cole. Physiological responses of three species of fishes to various salinities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):399–405, April 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-066.

Addison:1972:LCQ

[629] R. F. Addison, R. G. Ackman, and J. Hingley. Lipid composition of the Queen crab (Chionoecetes opilio). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):407–411, April 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-067.

Price:1972:LRS

[630] Vincent A. Price and Kenneth K. Chew. Laboratory rearing of spot shrimp larvae (Pandalus platyceros) and descriptions of stages. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):413–422, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-068.

Brown:1972:SMM

[631] Ralph B. Brown and Guy C. Powell. Size at maturity in the male Alaskan Tanner crab, Chionoecetes bairdi, as determined by chela allometry, reproductive tract weights, and size of precopulatory males. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):423–427, April 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-069.

Trust:1972:BPD

[632] T. J. Trust and Vida G. Money. Bacterial population of diets for aquarium fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):429– 433, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-070.

Sundararaj:1972:ESG

[633] Bangalore I. Sundararaj, Shashi V. Goswami, and Edward M. Donaldson. Effect of salmon gonadotropin on in vitro maturation of oocytes of a catfish, Heteropneustes fossilis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):435–437, April 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-071.

Stewart:1972:ETF

[634] James E. Stewart, G. W. Horner, and B. Arie. Effects of temperature, food, and starvation on several physiological parameters of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):439–442, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-072.

Laycock:1972:CBS

[635] R. A. Laycock and A. A. Longard. Clostridium botulinum in sediments from the Canadian Atlantic Seaboard. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):443–446, April 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-073.

Watson:1972:MBS

[636] J. Watson. Mating behavior in the spider crab, Chionoecetes opilio. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):447–449, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-074.

Windom:1972:DIM

[637] Herbert L. Windom and Ralph G. Smith. Distribution of iron, magnesium, copper, zinc, and silver in oysters along the Georgia Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):450–452, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-075.

Jessop:1972:ARW

[638] B. M. Jessop. Aging round whitefish (Prosopium cylindraceum) of the Leaf River, Ungava, Quebec, by otoliths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):452–454, April 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-076.

Carlson:1972:MLL

[639] Clarence A. Carlson and M. H. Shealy, Jr. Marking larval largemouth bass with radiostrontium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):455–458, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-077.

Carline:1972:EFA

[640] R. F. Carline and O. M. Brynildson. Effects of the floy anchor tag on the growth and survival of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):458–460, April 1972.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-078.

Stewart:1972:FSF

[641] James E. Stewart, B. M. Zwicker, B. Arie, and G. W. Horner. Food and starvation as factors affecting the time to death of the lobster Homarus americanus infected with Gaffkya homari. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):461–464, April 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-079.

Harris:1972:REN

[642] Colin Harris, Kenneth K. Chew, and Vincent Price. Relation of egg number to carapace length of sidestripe shrimp (Pandalopsis dispar) from dabob Bay, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):464–465, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-080.

Anonymous:1972:RAFd

[643] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):467, April 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-081.

Anonymous:1972:AFPd

[644] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(4):468, April 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-082.

Koshinsky:1972:ETT

[645] Gordon D. Koshinsky. An evaluation of two tags with northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5): 469–476, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-083.

Brown:1972:PBA

[646] Edward H. Brown, Jr. Population biology of alewives, Alosa pseudoharengus, in Lake Michigan, 1949–70. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):477–500, May 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-084.

Li:1972:TCB

[647] M. F. Li and G. S. Traxler. Tissue culture bioassay method for water pollution with special reference to mercuric chloride. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):501–505, May 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-085.

Sameoto:1972:DHC

[648] D. D. Sameoto. Distribution of herring (Clupea harengus) larvae along the southern coast of Nova Scotia with observations on their growth and condition factor. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5): 507–515, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-086.

Fortin:1972:CLP

[649] Réjean Fortin and Etienne Magnin. Croissance en longueur et en poids des perchaudes Perca flavescens de la Grande Anse de l’île Perrot au lac Saint-Louis. (French) [Growth in length and weight of perch Perca flavescens from Grande Anse on Perrot Island in Lake Saint-Louis]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):517–523, May 1972. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f72-087.

Reinert:1972:EDC

[650] Robert E. Reinert, Donald Stewart, and Harry L. Seagran. Effects of dressing and cooking on DDT concentrations in certain fish from Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):525– 529, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-088.

Hinton:1972:MEH

[651] David E. Hinton, Robert L. Snipes, and Michael W. Kendall. Morphology and enzyme histochemistry in the liver of largemouth bass (Micropterus salmoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):

531–534, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-089.

Hurley:1972:AEA

[652] Donal A. Hurley. The American eel (Anguilla rostrata) in Eastern Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):535– 543, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-090.

Dadswell:1972:PDF

[653] M. J. Dadswell. Postglacial dispersal of four deepwater fishes on the basis of new distribution records in Eastern Ontario and Western Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):545–553, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-091.

Schutz:1972:ESF

[654] D. C. Schutz and T. G. Northcote. An experimental study of feeding behavior and interaction of coastal cutthroat trout (Salmo clarki clarki) and Dolly Varden (Salvelinus malma). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):555–565, May 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-092.

Trust:1972:BPV

[655] T. J. Trust. The bacterial population in vertical flow tray hatcheries during incubation of salmonid eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):567–571, May 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-093.

Das:1972:GLH

[656] Naresh Das. Growth of larval herring (Clupea harengus) in the Bay of Fundy and Gulf of Maine area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):573–575, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-094.

Chiszar:1972:FBS

[657] David Chiszar, Mark Moody, and J. T. Windell. Failure of bluegill sunfish, Lepomis macrochirus, to habituate to handling. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):576–578, May 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-095.

Davisson:1972:KTF

[658] Muriel Trask Davisson. Karyotypes of the teleost family Esocidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):579–582, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-096.

Carlson:1972:ELT

[659] A. R. Carlson. Effects of long-term exposure to Carbaryl (Sevin) on survival, growth, and reproduction of the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):583– 587, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-097.

Malouf:1972:OGB

[660] Robert Malouf, Richard Keck, Don Maurer, and Charles Epifanio. Occurrence of gas-bubble disease in three species of bivalve molluscs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):588–589, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-098.

Lindsey:1972:LAT

[661] J. K. Lindsey and F. W. Nash. Likelihood analysis of three-way contingency tables. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5): 590–591, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-099.

Willis:1972:IEM

[662] D. E. Willis and R. F. Addison. Identification and estimation of the major components of a commercial polychlorinated biphenyl mixture, Aroclor 1221. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5): 592–595, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-100.

El-Sabh:1972:WGC

[663] M. I. El-Sabh and O. M. Johannessen. Winter geostrophic circulation in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):595–598, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-101.

Dunham:1972:SEG

[664] Philip J. Dunham. Some effects of group housing upon the aggressive behavior of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):598–601, May 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-102.

Sylvester:1972:ETS

[665] J. R. Sylvester. Effect of thermal stress on predator avoidance in sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):601– 603, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-103.

Mann:1972:RBL

[666] K. H. Mann and P. A. Breen. The relation between lobster abundance, sea urchins, and kelp beds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):603–605, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-104.

Westrheim:1972:SSC

[667] S. J. Westrheim and H. Tsuyuki. Synonymy of Sebastes caenaematicus with Sebastes borealis, and range extension record. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):606–607, May 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-105.

Anonymous:1972:RAFe

[668] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):609, May 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-106.

Anonymous:1972:AFPe

[669] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(5):610, May 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-107.

Anonymous:1972:EF

[670] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):611, June 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-108.

Loftus:1972:IPS

[671] Kenneth H. Loftus and Henry A. Regier. Introduction to the Proceedings of the 1971 Symposium on Salmonid Communities in Oligotrophic Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):613–616, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-109.

Ryder:1972:LFO

[672] R. A. Ryder. The limnology and fishes of oligotrophic glacial lakes in North America (about 1800 A.D.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):617–628, June 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-110.

Toivonen:1972:FFL

[673] Jorma Toivonen. The fish fauna and limnology of large oligotrophic glacial lakes in Europe (about 1800 A.D.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):629–637, June 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-111.

Behnke:1972:SSF

[674] Robert J. Behnke. The systematics of salmonid fishes of recently glaciated lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6): 639–671, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-112.

Beeton:1972:EP

[675] A. M. Beeton and W. T. Edmondson. The eutrophication problem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):673–682, June 1972. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f72-113.

Regier:1972:HRM

[676] Henry A. Regier and Vernon C. Applegate. Historical review of the management approach to exploitation and introduction in SCOL lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):683–692, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-114.

Nilsson:1972:EIS

[677] Nils-Arvid Nilsson. Effects of introductions of salmonids into barren lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):693– 697, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-115.

Hartman:1972:LFE

[678] W. L. Hartman and R. L. Burgner. Limnology and fish ecology of sockeye salmon nursery lakes of the world. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):699–715, June 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-116.

Smith:1972:FES

[679] Stanford H. Smith. Factors of ecologic succession in oligotrophic fish communities of the Laurentian Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):717–730, June 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-117.

Johnson:1972:KLC

[680] Lionel Johnson. Keller Lake: Characteristics of a culturally unstressed salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6): 731–740, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-118.

Keleher:1972:GSL

[681] J. J. Keleher. Great Slave Lake: Effects of exploitation on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):741– 753, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-119.

Roth:1972:BTB

[682] H. Roth and W. Geiger. Brienzersee, thunersee, and bielersee: Effects of exploitation and eutrophication on the salmonid communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):755–764, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-120.

Lawrie:1972:LSE

[683] A. H. Lawrie and J. F. Rahrer. Lake Superior: Effects of exploitation and introductions on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):765–776, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-121.

Grimaldi:1972:LME

[684] Ettore Grimaldi. Lago maggiore: Effects of exploitation and introductions on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):777–785, June 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-122.

Youngs:1972:CLE

[685] William D. Youngs and Ray T. Oglesby. Cayuga Lake: Effects of exploitation and introductions on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):787–794, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-123.

Martin:1972:LOE

[686] N. V. Martin and F. E. J. Fry. Lake Opeongo: Effects of exploitation and introductions on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):795–805, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-124.

Grimaas:1972:LVE

[687] Ulf Grimås, Nils-Arvid Nilsson, and Curt Wendt. Lake Vättern: Effects of exploitation, eutrophication, and introductions on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):807– 817, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-125.

LeCren:1972:WEE

[688] E. D. Le Cren, Charlotte Kipling, and Jean C. McCormack. Windermere: Effects of exploitation and eutrophication on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):819–832, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-126.

Numann:1972:BEE

[689] W. Numann. The bodensee: Effects of exploitation and eutrophication on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):833–847, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-127.

Maitland:1972:LLM

[690] Peter S. Maitland. Loch Lomond: Man’s effects on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):849– 860, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-128.

Northcote:1972:KLM

[691] T. G. Northcote. Kootenay Lake: Man’s effects on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):861– 865, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-129.

Laurent:1972:LLE

[692] P. J. Laurent. Lac Léman: Effects of exploitation, eutrophication, and introductions on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):867–875, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-130.

Berst:1972:LHE

[693] A. H. Berst and G. R. Spangler. Lake Huron: Effects of exploitation, introductions, and eutrophication on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):877–887, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-131.

Wells:1972:LME

[694] LaRue Wells and Alberton L. McLain. Lake Michigan: Effects of exploitation, introductions, and eutrophication on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):889–898, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-132.

Hartman:1972:LEE

[695] W. L. Hartman. Lake Erie: Effects of exploitation, environmental changes and new species on the fishery resources. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):899–912, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-133.

Christie:1972:LOE

[696] W. J. Christie. Lake Ontario: Effects of exploitation, introductions, and eutrophication on the salmonid community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):913–929, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-134.

Grimaldi:1972:FSC

[697] E. Grimaldi and W. Nümann. The future of salmonid communities in the European subalpine lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):931–936, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-135.

Grimaas:1972:FSC

[698] Ulf Grimås, Nils-Arvid Nilsson, Jorma Toivonen, and Curt Wendt. The future of salmonid communities in fennoscandian lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):937–940, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-136.

Ryder:1972:FSC

[699] R. A. Ryder and L. Johnson. The future of salmonid communities in North American oligotrophic lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):941–949, June 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-137.

Smith:1972:FSC

[700] Stanford H. Smith. The future of salmonid communities in the Laurentian Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6): 951–957, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-138.

Regier:1972:EFE

[701] Henry A. Regier and Kenneth H. Loftus. Effects of fisheries exploitation on salmonid communities in oligotrophic lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):959–968, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-139.

Christie:1972:ESI

[702] W. J. Christie, J. M. Fraser, and S. J. Nepszy. Effects of species introductions on salmonid communities in oligotrophic lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):969–973, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-140.

Colby:1972:EES

[703] P. J. Colby, G. R. Spangler, D. A. Hurley, and A. M. McCombie. Effects of eutrophication on salmonid communities in oligotrophic lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):975–983, June 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-141.

Anonymous:1972:PSP

[704] Anonymous. Postscript to SCOL Proceedings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(6):985–986, June 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-142.

Sameoto:1972:YRR

[705] D. D. Sameoto. Yearly respiration rate and estimated energy budget for Sagitta elegans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7): 987–996, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-143.

Tyler:1972:FRD

[706] A. V. Tyler. Food resource division among northern, marine, demersal fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):997– 1003, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-144.

Horak:1972:SHR

[707] Donald L. Horak. Survival of hatchery-reared rainbow trout (Salmo gairdneri) in relation to stamina tunnel ratings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1005–1009, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-145.

Carlson:1972:EHS

[708] H. Richard Carlson and R. E. Haight. Evidence for a home site and homing of adult yellowtail rockfish, Sebastes flavidus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1011–1014, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-146.

Kabata:1972:SAO

[709] Z. Kabata and B. Cousens. The structure of the attachment organ of Lernaeopodidae (Crustacea: Copepoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1015–1023, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-147.

Madison:1972:MMA

[710] D. M. Madison, R. M. Horrall, A. B. Stasko, and A. D. Hasler. Migratory movements of adult sockeye salmon (Oncorhynchus nerka) in coastal British Columbia as revealed by ultrasonic tracking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1025–1033, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-148.

Gillespie:1972:MMS

[711] D. C. Gillespie. Mobilization of mercury from sediments into guppies (Poecilia reticulata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29 (7):1035–1041, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-149.

Kelso:1972:VSB

[712] J. R. M. Kelso and F. J. Ward. Vital statistics, biomass, and seasonal production of an unexploited walleye (Stizostedion vitreum vitreum) population in West Blue Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1043–1052, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-150.

Dyer:1972:SEP

[713] W. J. Dyer, Doris F. Hiltz, R. G. Ackman, Joyce Hingley, G. L. Fletcher, and R. F. Addison. Stability of elemental phosphorus in edible muscle tissue of cod during processing including icing, freezing and thawing, frozen storage, salting, and cooking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1053–1060, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-151.

Gunderson:1972:EPO

[714] Donald R. Gunderson. Evidence that Pacific Ocean perch (Sebastes alutus) in Queen Charlotte Sound form aggregations that have different biological characteristics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1061–1070, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-152.

Scott:1972:CRU

[715] K. R. Scott and D. C. Gillespie. A compact recirculation unit for the rearing and maintenance of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1071–1074, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-153.

Fulton:1972:TAP

[716] John Fulton. Trials with an automated plankton counter. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1075–1078, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-154.

Manzer:1972:LWR

[717] J. I. Manzer. Length–weight relationship for pomfret. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1079–1081, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-155.

Scott:1972:TCS

[718] K. R. Scott. Temperature control system for recirculation fish-holding facilities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1082– 1083, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-156.

Radford:1972:FRA

[719] D. S. Radford. The first record of Anguilla rostrata from Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1084, July 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-157.

Ackman:1972:DSO

[720] R. G. Ackman, J. Hingley, and K. T. MacKay. Dimethyl sulfide as an odor component in Nova Scotia fall mackerel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1085–1088, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-158.

Thorson:1972:SHC

[721] Kenneth N. Thorson. Subcutaneous hemorrhage in captive sablefish (Anoplopoma fimbria): A possible link to mortality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1089–1090, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-159.

Clifford:1972:DMW

[722] Hugh F. Clifford. Downstream movements of white sucker, Catostomus commersoni, fry in a brown-water stream of Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1091–1093, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-160.

Stein:1972:FRB

[723] David Stein and John Butler. First records of Bathysaurus mollis (Pisces: Bathysauridae) from the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1093–1094, July 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-161.

Anonymous:1972:RAFf

[724] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1095– 1096, July 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-162.

Anonymous:1972:AFPf

[725] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(7):1097, July 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-163.

Pickering:1972:ACT

[726] Quentin H. Pickering and Marion H. Gast. Acute and chronic toxicity of cadmium to the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1099–1106, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-164.

McCormick:1972:ETG

[727] J. Howard McCormick, Kenneth E. F. Hokanson, and Bernard R. Jones. Effects of temperature on growth and survival of young brook trout, Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8): 1107–1112, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-165.

Messieh:1972:UOI

[728] Shoukry N. Messieh. Use of otoliths in identifying herring stocks in the Southern Gulf of St. Lawrence and adjacent waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1113–1118, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-166.

Jones:1972:EIE

[729] Barry C. Jones. Effect of intertidal exposure on survival and embryonic development of Pacific herring spawn. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1119–1124, August 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-167.

Miller:1972:VCP

[730] A. Miller III, R. A. Scanlan, J. S. Lee, and L. M. Libbey. Volatile compounds produced in ground muscle tissue of canary rockfish (Sebastes pinniger) stored on ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1125–1129, August 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-168.

Beamish:1972:ACM

[731] Richard J. Beamish and Harold H. Harvey. Acidification of the La Cloche Mountain Lakes, Ontario, and resulting fish mortalities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1131–1143, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-169.

Butler:1972:SMS

[732] John L. Butler and William G. Pearcy. Swimbladder morphology and specific gravity of myctophids off Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1145–1150, August 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-170.

Bams:1972:QES

[733] R. A. Bams. A quantitative evaluation of survival to the adult stage and other characteristics of pink salmon (Oncorhynchus gorbuscha) produced by a revised hatchery method which simulates optimal natural conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1151–1167, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-171.

Clayton:1972:ACB

[734] J. W. Clayton and D. N. Tretiak. Amine-citrate buffers for pH control in starch gel electrophoresis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1169–1172, August 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-172.

McAllister:1972:MMC

[735] Don E. McAllister, Pierre Jolicoeur, and Hiroshi Tsuyuki. Morphological and myogen comparison of Johnny and tessellated darters and their hybrids, genus Etheostoma, near Ottawa, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1173–1180, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-173.

Kelso:1972:CMA

[736] John R. M. Kelso. Conversion, maintenance, and assimilation for walleye, Stizostedion vitreum vitreum, as affected by size, diet, and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1181–1192, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-174.

Ware:1972:PRT

[737] D. M. Ware. Predation by rainbow trout (Salmo gairdneri): the influence of hunger, prey density, and prey size. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1193–1201, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-175.

Rudd:1972:BTN

[738] J. W. M. Rudd and R. D. Hamilton. Biodegradation of trisodium nitrilotriacetate in a model aerated sewage lagoon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1203–1208, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-176.

Larson:1972:GMB

[739] Gary Lee Larson. Gonad maturation of brook trout (Salvelinus fontinalis) in a high mountain lake under a modified photoperiod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1209–1211, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-177.

Forrester:1972:PFL

[740] C. R. Forrester. Pectoral fin length in the dogfish Squalus acanthias off British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8): 1211–1212, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-178.

Arnold:1972:PHP

[741] P. W. Arnold. Predation on harbour porpoise, Phocoena phocoena, by a white shark, Carcharodon carcharias. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1213–1214, August 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-179.

Evans:1972:CLC

[742] David O. Evans. Correction for lag in continuous-flow respirometry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1214–1216, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-180.

Byrne:1972:IST

[743] J. M. Byrne, F. W. H. Beamish, and R. L. Saunders. Influence of salinity, temperature, and exercise on plasma osmolality and ionic concentration in Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1217–1220, August 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-181.

Flowers:1972:ATE

[744] J. M. Flowers and S. B. Saila. An analysis of temperature effects on the inshore lobster fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1221–1225, August 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-182.

Wilson:1972:ATS

[745] R. C. H. Wilson. Acute toxicity of spent sulphite liquor to Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8): 1225–1228, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-183.

Flint:1972:EER

[746] R. Warren Flint. Effects of eyestalk removal and ecdysterone infusion on molting in Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1229–1233, August 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-184.

Thomas:1972:BTL

[747] Martin L. H. Thomas and Elizabeth Jelley. Benthos trapped leaving the bottom in Bideford River, Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1234–1237, August 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-185.

Hodges:1972:DMR

[748] John W. Hodges. Downstream migration of recently transformed sea lampreys before and after treatment of a Lake Michigan tributary with a lampricide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8): 1237–1240, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-186.

Anonymous:1972:CRP

[749] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1241, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-187.

Anonymous:1972:RAFg

[750] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1243, August 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-188.

Anonymous:1972:AFPg

[751] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(8):1244, August 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-189.

Gibbons:1972:ESB

[752] John R. H. Gibbons and John H. Gee. Ecological segregation between longnose and blacknose dace (genus Rhinichthys) in the Mink River, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1245–1252, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-190.

Glooschenko:1972:DPC

[753] Walter A. Glooschenko, Herbert Curl, Jr., and Lawrence F. Small. Diel periodicity of chlorophyll a concentration in Oregon coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1253–1259, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-191.

Small:1972:EPP

[754] Lawrence F. Small, Herbert Curl, Jr., and Walter A. Glooschenko. Estimates of primary production off Oregon using an improved chlorophylllight technique. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9): 1261–1267, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-192.

Small:1972:ESR

[755] Lawrence F. Small, Herbert Curl, Jr., and Walter A. Glooschenko. Effects of solar radiation and upwelling on daily primary production off Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1269–1275, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-193.

Beamish:1972:TCB

[756] F. W. H. Beamish and I. C. Potter. Timing of changes in the blood, morphology, and behavior of Petromyzon marinus during metamorphosis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1277–1282, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-194.

Lovett:1972:STC

[757] Raymond J. Lovett, Walter H. Gutenmann, Irene S. Pakkala, William D. Youngs, Donald J. Lisk, George E. Burdick, and Earl J. Harris. A survey of the total cadmium content of 406 fish from 49 New York State fresh waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1283–1290, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-195.

Khan:1972:TPS

[758] R. A. Khan and E. R. Noble. Taxonomy, prevalence, and specificity of Cryptobia dahli (Möbius) (Mastigophora: Bodonidae) in lumpfish, Cyclopterus lumpus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9): 1291–1294, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-196.

Fletcher:1972:ATY

[759] G. L. Fletcher and R. J. Hoyle. Acute toxicity of yellow phosphorus to Atlantic cod (Gadus morhua) and Atlantic salmon (Salmo salar) smolts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1295–1301, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-197.

Morris:1972:PSO

[760] Robert J. Morris. The preservation of some oceanic animals for lipid analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1303–1307, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-198.

Adelman:1972:THS

[761] Ira R. Adelman and Lloyd L. Smith, Jr. Toxicity of hydrogen sulfide to goldfish (Carassius auratus) as influenced by temperature, oxygen, and bioassay techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9): 1309–1317, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-199.

Curtis:1972:DDB

[762] Mark A. Curtis. Depth distributions of benthic polychaetes in two fiords on Ellesmere Island, N.W.T. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1319–1327, September 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-200.

Lehmkuhl:1972:CTR

[763] D. M. Lehmkuhl. Change in thermal regime as a cause of reduction of benthic fauna downstream of a reservoir. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1329–1332, September 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-201.

Davy:1972:EES

[764] F. B. Davy, H. Kleerekoper, and P. Gensler. Effects of exposure to sublethal DDT on the locomotor behavior of the goldfish (Carassius auratus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1333–1336, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-202.

Steele:1972:BGA

[765] D. H. Steele and V. J. Steele. Biology of Gatnmarellus angulosus (Crustacea, Amphipoda) in the Northwestern Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1337–1340, September 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-203.

Steele:1972:SAB

[766] D. H. Steele. Some aspects of the biology of Gammarellus homari (Crustacea, Amphipoda) in the Northwestern Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1340–1343, September 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-204.

Houston:1972:BCT

[767] A. H. Houston and R. J. Woods. Blood concentrations of tricaine methane sulphonate in brook trout, Salvelinus fontinalis, during anesthetization, branchial irrigation, and recovery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1344–1346, September 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-205.

Scarratt:1972:BCO

[768] D. J. Scarratt and V. Zitko. Bunker C oil in sediments and benthic animals from shallow depths in Chedabucto Bay, N.S. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1347–1350, September 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-206.

Hara:1972:ERO

[769] Toshiaki J. Hara. Electrical responses of the olfactory bulb of Pacific salmon Oncorhynchus nerka and Oncorhynchus kisutch. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1351–1355, September 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-207.

Sparks:1972:ESR

[770] Richard E. Sparks, William T. Waller, and John Cairns, Jr. Effect of shelters on the resistance of dominant and submissive blue gills (Lepomis macrochirus) to a lethal concentration of zinc. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1356–1358, September 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-208.

Ross:1972:ETI

[771] A. J. Ross and Cathy A. Smith. Effect of two iodophors on bacterial and fungal fish pathogens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1359–1361, September 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-209.

Butler:1972:FOC

[772] David Gordon Butler. Failure to observe changes in selected metabolites following removal of the Stannius corpuscles from the freshwater North American eel, Anguilla rostrata. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1362–1364, September 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-210.

Iredale:1972:ESP

[773] D. G. Iredale and D. Rigby. Effect of smoke-processing on muddy odor and taste in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1365–1366, September 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-211.

Eisler:1972:CUM

[774] R. Eisler, G. E. Zaroogian, and R. J. Hennekey. Cadmium uptake by marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9): 1367–1369, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-212.

Anonymous:1972:RAFh

[775] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1371, September 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-213.

Anonymous:1972:AFPh

[776] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(9):1372, September 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-214.

Allen:1972:MGA

[777] K. Radway Allen, Richard L. Saunders, and Paul F. Elson. Marine growth of Atlantic salmon (Salmo salar) in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1373–1380, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-215.

Sandeman:1972:BBD

[778] I. M. Sandeman and M. D. B. Burt. Biology of Bothrimonus (= Diplocotyle) (Pseudophyllidea: Cestoda): Ecology, life cycle, and evolution; a review and synthesis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):

1381–1395, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-216.

Ducharme:1972:ALD

[779] L. J. A. Ducharme. An application of louver deflectors for guiding Atlantic salmon (Salmo salar) smolts from power turbines. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1397–1404, October 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-217.

Kraft:1972:ECF

[780] Melvin E. Kraft. Effects of controlled flow reduction on a trout stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1405–1411, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-218.

Reinert:1972:ADA

[781] Robert E. Reinert. Accumulation of dieldrin in an alga (Scenedesmus obliquus), Daphnia magna, and the guppy (Poecilia reticulata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1413–1418, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-219.

Donaldson:1972:SSL

[782] Henry A. Donaldson and William G. Pearcy. Sound-scattering layers in the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1419–1423, October 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-220.

Trust:1972:IAA

[783] T. J. Trust. Inadequacy of aquarium antibacterial formulations for the inhibition of potential pathogens of freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1425–1430, October 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-221.

Groot:1972:MYS

[784] C. Groot. Migration of yearling sockeye salmon (Oncorhynchus nerka) as determined by time-lapse photography of sonar observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1431–1444, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-222.

Dodson:1972:BAA

[785] Julian J. Dodson, William C. Leggett, and Robert A. Jones. The behavior of adult American shad (Alosa sapidissima) during migration from salt to fresh water as observed by ultrasonic tracking techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1445–1449, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-223.

Patalas:1972:CPE

[786] K. Patalas. Crustacean plankton and the eutrophication of St. Lawrence Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1451– 1462, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-224.

Burton:1972:AZT

[787] Dennis T. Burton, Alma H. Jones, and John Cairns, Jr. Acute zinc toxicity to rainbow trout (Salmo gairdneri): Confirmation of the hypothesis that death is related to tissue hypoxia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1463–1466, October 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-225.

Bilinski:1972:OLC

[788] E. Bilinski and R. E. E. Jonas. Oxidation of lactate to carbon dioxide by rainbow trout (Salmo gairdneri) tissues. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1467–1471, October 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-226.

Fisher:1972:DLF

[789] Stuart G. Fisher and Alison Lavoy. Differences in littoral fauna due to fluctuating water levels below a hydroelectric dam. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1472–1476, October 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-227.

Arnold:1972:SRR

[790] P. W. Arnold and D. E. Gaskin. Sight records of right whales (Eubalaena glacialis) and finback whales (Balaenoptera physalus) from the Lower Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10): 1477–1478, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-228.

Krouse:1972:SLH

[791] Jay S. Krouse. Some life history aspects of the rock crab, Cancer irroratus, in the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1479–1482, October 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-229.

Schultz:1972:FAS

[792] David M. Schultz and James G. Quinn. Fatty acids in surface particulate matter from the North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1482–1486, October 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-230.

Forrester:1972:MCS

[793] C. R. Forrester, K. S. Ketchen, and C. C. Wong. Mercury content of spiny dogfish (Squalus acanthias) in the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1487– 1490, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f72-231.

Vernberg:1972:TSS

[794] Winona B. Vernberg and James O’Hara. Temperature–salinity stress and mercury uptake in the fiddler crab, Uca pugilator. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1491–1494, October 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-232.

Levesque:1972:FAC

[795] Raymond C. Levesque and Roger J. Reed. Food availability and consumption by young Connecticut River shad Alosa sapidissima. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1495–1499, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-233.

Wilson:1972:PCT

[796] R. C. H. Wilson. Prediction of copper toxicity in receiving waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1500–1502, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f72-234.

Clarke:1972:ADB

[797] A. H. Clarke. The Arctic dredge, a benthic biological sampler for mixed boulder and mud substrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1503–1505, October 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-235.

Trussell:1972:PIA

[798] R. P. Trussell. The percent un-ionized ammonia in aqueous ammonia solutions at different pH levels and temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1505–1507, October 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-236.

Anonymous:1972:RAFi

[799] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1509–1510, October 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-237.

Anonymous:1972:AFPi

[800] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(10):1511, October 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-238.

DeCew:1972:ATE

[801] Mark G. DeCew. Antibiotic toxicity, efficacy, and teratogenicity in adult spring Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1513–1517, November 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-239.

Lockhart:1972:MNP

[802] W. L. Lockhart, J. F. Uthe, A. R. Kenney, and P. M. Mehrle. Methylmercury in northern pike (Esox lucius): Distribution, elimination, and some biochemical characteristics of contaminated fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1519–1523, November 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-240.

Webb:1972:RAP

[803] P. W. Webb and J. R. Brett. Respiratory adaptations of prenatal young in the ovary of two species of viviparous seaperch, Rhacochilus vacca and Embiotoca lateralis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11): 1525–1542, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-241.

Webb:1972:OCE

[804] P. W. Webb and J. R. Brett. Oxygen consumption of embryos and parents, and oxygen transfer characteristics within the ovary of two species of viviparous seaperch, Rhacochilus vacca and Embiotoca lateralis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1543–1553, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-242.

Webb:1972:ESC

[805] P. W. Webb and J. R. Brett. The effects of sublethal concentrations of whole bleached kraftmill effluent on the growth and food conversion efficiency of underyearling sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1555–1563, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-243.

Green:1972:DMV

[806] Roger H. Green. Distribution and morphological variation of Lampsilis radiata (Pelecypoda, Unionidae) in some Central Canadian lakes: A multivariate statistical approach. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1565–1570, November 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-244.

Kabata:1972:DSC

[807] Z. Kabata. Developmental stages of Caligus clemensi (Copepoda: Caligidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1571–1593, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-245.

Dickman:1972:SEA

[808] Mike Dickman and Ian E. Efford. Some effects of artificial fertilization on enclosed plankton populations in Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1595–1604, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-246.

Klima:1972:VPR

[809] Edward F. Klima. Voltage and pulse rates for inducing electrotaxis in twelve coastal pelagic and bottom fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1605–1614, November 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-247.

Bryan:1972:FSI

[810] James E. Bryan and P. A. Larkin. Food specialization by individual trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1615–1624, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-248.

Anonymous:1972:Ob

[811] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1625, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-249.

Schindler:1972:ABA

[812] D. W. Schindler, R. V. Schmidt, and R. A. Reid. Acidification and bubbling as an alternative to filtration in determining phytoplankton production by the ^{14}C method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1627–1631, November 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-250.

Vladykov:1972:CNS

[813] V. D. Vladykov and C. G. Gruchy. Comments on the nomenclature of some subgenera of Salmonidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1631–1632, November 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-251.

Thomson:1972:IVP

[814] W. K. Thomson and C. L. Thacker. Incidence of Vibrio parahaemolyticus in shellfish from eight Canadian Atlantic sampling areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1633–1635, November 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-252.

Hamley:1972:UDM

[815] John M. Hamley. Use of the DeLury method to estimate gillnet selectivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1636–1638, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-253.

Nepszy:1972:OGH

[816] Stephen J. Nepszy and Alex O. Dechtiar. Occurrence of Glugea hertwigi in Lake Erie rainbow smelt (Osmerus mordax) and associated mortality of adult smelt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11): 1639–1641, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-254.

Scott:1972:CEV

[817] K. R. Scott. Comparison of the efficiency of various aeration devices for oxygenation of water in aquaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1641–1643, November 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-255.

Gaskin:1972:MHP

[818] D. E. Gaskin, K. Ishida, and R. Frank. Mercury in harbour porpoises (Phocoena phocoena) from the Bay of Fundy region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1644–1646, November 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-256.

Demory:1972:SMA

[819] Robert L. Demory. Scales as a means of aging Dover sole (Microstomus pacificus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1647–1650, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-257.

Jensen:1972:PBN

[820] A. L. Jensen. Population biomass, number of individuals, average individual weight, and the linear surplus-production model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1651–1655, November 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-258.

Falk:1972:UOB

[821] M. R. Falk. Unusual occurrence of the brook stickleback (Culaea inconstans) in the Mackenzie River, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1655–1656, November 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-259.

Price:1972:PSP

[822] R. J. Price and J. S. Lee. Paralytic shellfish poison and melanin distribution in fractions of toxic butter clam (Saxidomus giganteus) siphon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1657–1658, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-260.

Price:1972:ECI

[823] R. J. Price and J. S. Lee. Effect of cations on the interaction between paralytic shellfish poison and butter clam (Saxidomus giganteus) melanin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1659–1661, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-261.

Anonymous:1972:RAFj

[824] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1663, November 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-262.

Anonymous:1972:AFPj

[825] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(11):1664, November 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-263.

Anonymous:1972:A

[826] Anonymous. Announcement. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1665, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-264.

Scott:1972:ELN

[827] J. S. Scott. Eggs and larvae of northern sand lance (Ammodytes dubius) from the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1667–1671, December 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-265.

Scott:1972:MMV

[828] J. S. Scott. Morphological and meristic variation in Northwest Atlantic sand lances (Ammodytes). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1673–1678, December 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-266.

Zurbrigg:1972:EEP

[829] R. E. Zurbrigg and W. B. Scott. Evidence for expatriate populations of the lanternfish Myctophum punctatum in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1679–1683, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-267.

Scott:1972:MCR

[830] D. P. Scott and F. A. J. Armstrong. Mercury concentration in relation to size in several species of freshwater fishes from Manitoba and Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12): 1685–1690, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-268.

Biesinger:1972:EVM

[831] Kenneth E. Biesinger and Glenn M. Christensen. Effects of various metals on survival, growth, reproduction, and metabolism of Daphnia magna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1691–1700, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-269.

Lasenby:1972:GLH

[832] D. C. Lasenby and R. R. Langford. Growth, life history, and respiration of Mysis relicta in an Arctic and temperate lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1701–1708, December 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-270.

Vladykov:1972:MDM

[833] V. D. Vladykov. Morphological differences in male gonads among nine genera of Gadidae (Pisces). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1709–1716, December 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-271.

Ketchen:1972:SMF

[834] K. S. Ketchen. Size at maturity, fecundity, and embryonic growth of the spiny dogfish (Squalus acanthias) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1717–1723, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-272.

McLeese:1972:ESF

[835] D. W. McLeese. Effects of several factors on the growth of the American lobster (Homarus americanus) in captivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1725–1730, December 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-273.

Russell:1972:UCD

[836] Howard J. Russell, Jr. Use of a commercial dredge to estimate a hardshell clam population by stratified random sampling. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1731–1735, December 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-274.

Stein:1972:SIB

[837] Roy A. Stein, Paul E. Reimers, and James D. Hall. Social interaction between juvenile Coho (Oncorhynchus kisutch) and fall Chinook salmon (O. tshawytscha) in Sixes River, Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1737–1748, December 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-275.

MacKinnon:1972:SSE

[838] J. C. MacKinnon. Summer storage of energy and its use for winter metabolism and gonad maturation in American plaice (Hippoglossoides platessoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12): 1749–1759, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-276.

Wong:1972:SSD

[839] Brian Wong and F. J. Ward. Size selection of Daphnia pulicaria by yellow perch (Perca flavescens) fry in West Blue Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1761–1764, December 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-277.

Smith:1972:SLW

[840] Ian D. Smith. Sea lions wintering along the outer coast of Vancouver Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1764–1766, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-278.

Farmer:1972:IMC

[841] D. M. Farmer and T. R. Osborn. An instrument for measuring conductivity profiles in inlets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12): 1767–1769, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-279.

Strasdine:1972:SCW

[842] G. A. Strasdine and Joanne M. Melville. Salmon-canning waste water as a microbial growth medium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1769–1771, December 1972. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f72-280.

Lindsey:1972:NCZ

[843] C. C. Lindsey and W. G. Franzin. New complexities in zoogeography and taxonomy of the pygmy whitefish (Prosopium coulteri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1772–1775, December 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-281.

Hendricks:1972:TNH

[844] Jerry D. Hendricks. Two new host species for the parasitic fungus Ichthyophonus hoferi in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1776–1777, December 1972. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f72-282.

Regier:1972:MRF

[845] L. W. Regier. Mercury removal from fish protein concentrate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1777–1779, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-283.

Wedemeyer:1972:SPC

[846] Gary Wedemeyer. Some physiological consequences of handling stress in the juvenile Coho salmon (Oncorhynchus kisutch) and steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12): 1780–1783, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-284.

Hunter:1972:DML

[847] Charles J. Hunter and Chester O. Wheeler. Device for measuring length and girth of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12): 1784–1785, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f72-285.

Hurst:1972:RLM

[848] R. E. Hurst. The representation of lower molecular weight fatty acids by the bligh and Dyer extraction method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1786–1787, December 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-286.

MacCrimmon:1972:WDR

[849] Hugh R. MacCrimmon. World distribution of rainbow trout (Salmo gairdneri): Further observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1788–1791, December 1972. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f72-287.

Outram:1972:FPH

[850] Donald N. Outram and C. Haegele. Food of Pacific hake (Merluccius productus) on an offshore bank southwest of Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1792–1795, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-288.

Anonymous:1972:RAFk

[851] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1797–1798, December 1972. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f72-289.

Anonymous:1972:AFPk

[852] Anonymous. Additional FRB publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 29(12):1799, December 1972. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f72-290.

Bilton:1973:ESS

[853] H. T. Bilton and G. L. Robins. The effects of starvation and subsequent feeding on survival and growth of Fulton Channel sockeye salmon fry (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):1–5, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-001.

Moore:1973:FLS

[854] J. W. Moore and F. W. H. Beamish. Food of larval sea lamprey (Petromyzon marinus) and American brook lamprey (Lampetra lamottei). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):7–15, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-002.

Sergeant:1973:FGP

[855] D. E. Sergeant. Feeding, growth, and productivity of Northwest Atlantic harp seals (Pagophilus groenlandicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):17–29, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-003.

Grant:1973:ETR

[856] Blake F. Grant and Paul M. Mehrle. Endrin toxicosis in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30 (1):31–40, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-004.

Martin:1973:EDM

[857] J.-L. M. Martin. Étude du métabolisme de l’eau et de l’azote dans l’hépatopancréas et l’ovaire de Carcinus maenas. (French) [Study of water and nitrogen metabolism in the hepatopancreas and ovary of Carcinus maenas]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):41–44, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-005.

Hiltz:1973:HMA

[858] Doris Fraser Hiltz and W. J. Dyer. Hexose monophosphate accumulation and related metabolic changes in unfrozen and thawed adductor muscle of the sea scallop (Placopecten magellanicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):45–52, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-006.

Murty:1973:BVG

[859] T. S. Murty and K. B. Yuen. Balanced versus geostrophic wind-stress for Hudson Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30 (1):53–62, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-007.

Botta:1973:TAV

[860] J. R. Botta, J. F. Richards, and N. Tomlinson. Thiobarbituric acid value, total long-chain free fatty acids, and flavor of Pacific halibut (Hippoglossus stenolepis) and Chinook salmon (Oncorhynchus tshawytscha) frozen at sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):63–69, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-008.

Botta:1973:FPC

[861] J. R. Botta, J. F. Richards, and N. Tomlinson. Flesh pH, color, thaw drip, and mineral concentration of Pacific halibut (Hippoglossus stenolepis) and Chinook salmon (Oncorhynchus tshawytscha) frozen at sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):71–77, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-009.

Botta:1973:FCV

[862] J. R. Botta, J. F. Richards, and N. Tomlinson. Flesh concentration of various long-chain free fatty acids of Pacific halibut (Hippoglossus stenolepis) and Chinook salmon (Oncorhynchus tshawytscha) frozen at sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):79–82, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-010.

Thomas:1973:EBC

[863] Martin L. H. Thomas. Effects of bunker C oil on intertidal and lagoonal biota in Chedabucto Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):83–90, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-011.

Castell:1973:MFF

[864] C. H. Castell and Barbara Smith. Measurement of formaldehyde in fish muscle using TCA extraction and the Nash reagent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):91–98, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-012.

Davis:1973:GIP

[865] John C. Davis and David J. Randall. Gill irrigation and pressure relationships in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):99–104, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-013.

Chittenden:1973:EHO

[866] Mark E. Chittenden, Jr. Effects of handling on oxygen requirements of American shad (Alosa sapidissima). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):105–110, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-014.

Anonymous:1973:Oa

[867] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):111, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-015.

Addison:1973:SLV

[868] R. F. Addison, R. G. Ackman, and J. Hingley. Seasonal and local variations in odd-chain fatty acid levels in Nova Scotian rainbow smelt (Osmerus mordax). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1): 113–115, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-016.

Schachte:1973:PAA

[869] John H. Schachte, Jr. and Emilio C. Mora. Production of agglutinating antibodies in the channel catfish (Ictalurus punctatus) against Chondrococcus columnaris. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30 (1):116–118, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-017.

Stasko:1973:HBM

[870] A. B. Stasko and S. M. Polar. Hydrophone and bow-mount for tracking fish by ultrasonic telemetry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):119–121, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-018.

Ingram:1973:WSC

[871] R. Grant Ingram. Winter surface currents around Cape Breton Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):121–123, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-019.

Sergeant:1973:TMH

[872] D. E. Sergeant. Transatlantic migration of a harp seal, Pagophilus groenlandicus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):124–125, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-020.

Robison:1973:SMM

[873] Bruce H. Robison. A system for maintaining midwater fishes in captivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):126–128, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-021.

Hare:1973:MSN

[874] G. M. Hare. A modified stake net for collecting migrating smolts of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):128–129, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-022.

Green:1973:EHM

[875] John M. Green. Evidence for homing in the mosshead sculpin (Clinocottus globiceps). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1): 129–130, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-023.

Lark:1973:ERS

[876] John G. I. Lark. An early record of the sea lamprey (Petromyzon marinus) from Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):131–133, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-024.

Musick:1973:MFG

[877] John A. Musick. Mesopelagic fishes from the Gulf of Maine and the adjacent continental slope. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):134–137, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-025.

McIntyre:1973:GAH

[878] J. D. McIntyre and J.-M. Blanc. A genetic analysis of hatching time in steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):137–139, January 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-026.

Anonymous:1973:RAFa

[879] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(1):141, January 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-027.

Prakash:1973:GCS

[880] A. Prakash, Liv Skoglund, Britt Rystad, and Arne Jensen. Growth and cell-size distribution of marine planktonic algae in batch and dialysis cultures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):143–155, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-028.

Castell:1973:ESS

[881] C. H. Castell and Doris M. Bishop. Effect of season on salt-extractable protein in muscle from trawler-caught cod and on its stability during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2): 157–160, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-029.

Creaser:1973:RBG

[882] Edwin P. Creaser, Jr. Reproduction of the bloodworm (Glycera dibranchiata) in the Sheepscot Estuary, Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):161–166, February 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-030.

Mackie:1973:ADM

[883] G. L. Mackie and S. U. Qadri. Abundance and diversity of Mollusca in an industrialized portion of the Ottawa River near Ottawa–Hull, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):167–172, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-031.

Caddy:1973:UOT

[884] J. F. Caddy. Underwater observations on tracks of dredges and trawls and some effects of dredging on a scallop ground. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):173–180, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-032.

Farrington:1973:FAC

[885] John W. Farrington, James G. Quinn, and Wayne R. Davis. Fatty acid composition of Nephtys incisa and Yoldia limatula. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):181–185, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-033.

Clayton:1973:GGP

[886] J. W. Clayton, W. G. Franzin, and D. N. Tretiak. Genetics of glycerol-3-phosphate dehydrogenase isozymes in white muscle of lake whitefish (Coregonus clupeaformis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):187–193, February 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-034.

Haines:1973:ERD

[887] Terry A. Haines. An evaluation of RNA–DNA ratio as a measure of long-term growth in fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):195–199, February 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-035.

Norms:1973:ESM

[888] James S. Norms, David O. Norris, and John T. Windell. Effect of simulated meal size on gastric acid and pepsin secretory rates in bluegill (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2): 201–204, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-036.

Vladykov:1973:LPN

[889] V. D. Vladykov. Lampetra pacifica, a new nonparasitic species of lamprey (Petromyzontidae) from Oregon and California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):205–213, February 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-037.

Cattell:1973:SCV

[890] S. Allen Cattell. The seasonal cycle of vitamin B_{12} in the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):215–222, February 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-038.

Kutty:1973:SPY

[891] M. N. Kutty and R. L. Saunders. Swimming performance of young Atlantic salmon (Salmo salar) as affected by reduced ambient oxygen concentration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):223–227, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-039.

Ginetz:1973:CCF

[892] R. M. Ginetz and P. A. Larkin. Choice of colors of food items by rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):229–234, February 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-040.

Westrheim:1973:ADG

[893] S. J. Westrheim. Age determination and growth of Pacific Ocean perch (Sebastes alutus) in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):235–247, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-041.

Cove:1973:HTG

[894] J. J. Cove. Hunters, trappers, and gatherers of the sea: a comparative study of fishing strategies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):249–259, February 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-042.

Levy:1973:DPP

[895] E. M. Levy and A. Walton. Dispersed and particulate petroleum residues in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):261–267, February 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-043.

Trust:1973:IEE

[896] T. J. Trust and A. J. Wood. An initial evaluation of ethylene oxide for the sterilization of formulated and pelleted fish feeds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):269–274, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-044.

Windom:1973:ACC

[897] H. Windom, R. Stickney, R. Smith, D. White, and F. Taylor. Arsenic, cadmium, copper, mercury, and zinc in some species of North Atlantic finfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):275–279, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-045.

McLeese:1973:RLH

[898] D. W. McLeese. Response of lobsters Homarus americanus to odor solution in the presence of bleached kraft mill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):279–282, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-046.

Hara:1973:SAS

[899] T. J. Hara, Y. M. C. Law, and E. Van der Veen. A stimulatory apparatus for studying the olfactory activity in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):283–285, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-047.

Berrill:1973:TDM

[900] Michael Berrill and Robert Stewart. Tunnel-digging in mud by newlysettled American lobsters, Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):285–287, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-048.

Burton:1973:IET

[901] W. Burton and John F. Flannagan. An improved Ekman-type grab. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):287–290, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-049.

Scott:1973:IHP

[902] J. S. Scott. Intestinal helminth parasites of northern sand lance (Ammodytes dubius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2): 291–292, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-050.

Williams:1973:MME

[903] P. M. Williams and H. V. Weiss. Mercury in the marine environment: Concentration in sea water and in a pelagic food chain. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):293–295, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-051.

Wedemeyer:1973:NFB

[904] Gary A. Wedemeyer and A. J. Ross. Nutritional factors in the biochemical pathology of corynebacterial kidney disease in the Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2): 296–298, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-052.

Jessop:1973:AGM

[905] B. M. Jessop and G. Power. Age, growth, and maturity of round whitefish (Prosopium cylindraceum) from the Leaf River, Ungava, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):299–304, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-053.

Venugopal:1973:ESL

[906] V. Venugopal, K. A. Savagaon, U. S. Kumta, and A. Sreenivasan. Extension of shelf-life of Indian mackerel (Rastrelliger kanagurta) by irradiation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):305–309, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-054.

Hunding:1973:CBM

[907] C. Hunding and B. T. Hargrave. A comparison of benthic microalgal production measured by C^{14} and oxygen methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):309–312, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-055.

Duval:1973:SSD

[908] W. S. Duval, T. J. Brown, and G. H. Geen. A submersible spectroradiometer and data acquisition system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):313–316, February 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-056.

Coble:1973:IAP

[909] Daniel W. Coble. Influence of appearance of prey and satiation of predator on food selection by northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):317–320, February 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-057.

Boudreault:1973:VAO

[910] F.-Robert Boudreault and J. Pierre Laprise. Vers une analyse objective de la thermocline saisonnière. (French) [Towards an objective analysis of the seasonal thermocline]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):320–322, February 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-058.

Macek:1973:SGC

[911] Kenneth J. Macek and Robert N. Sturm. Survival and gill condition of bluegill (Lepomis macrochirus) and fathead minnows (Pimephales promelas) exposed to sodium nitrilotriacetate (NTA) for 28 days. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):323–325, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-059.

Gatz:1973:SSM

[912] A. John Gatz, Jr. Speed, stamina, and muscles in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):325–328, February 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-060.

Smith:1973:CFJ

[913] Wesley E. Smith. A cyprinodontid fish, Jordanella floridae, as a laboratory animal for rapid chronic bioassays. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):329–330, February 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-061.

Anonymous:1973:RAFb

[914] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(2):331, February 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-062.

Sameoto:1973:ALC

[915] D. D. Sameoto. Annual life cycle and production of the chaetognath Sagitta elegans in Bedford Basin, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):333–344, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-063.

Daley:1973:QSM

[916] Ralph J. Daley, C. B. J. Gray, and S. R. Brown. A quantitative, semiroutine method for determining algal and sedimentary chlorophyll derivatives. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):345–356, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-064.

Moshenko:1973:DTP

[917] Robert W. Moshenko and John H. Gee. Diet, time and place of spawning, and environments occupied by creek chub (Semotilus atromaculatus) in the Mink River, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):357–362, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-065.

Gardner:1973:CIL

[918] George R. Gardner and Gilles LaRoche. Copper induced lesions in estuarine teleosts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30 (3):363–368, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-066.

Davis:1973:SEB

[919] John C. Davis. Sublethal effects of bleached kraft pulp mill effluent on respiration and circulation in sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):369–377, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-067.

Brett:1973:MRC

[920] J. R. Brett and N. R. Glass. Metabolic rates and critical swimming speeds of sockeye salmon (Oncorhynchus nerka) in relation to size and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):379–387, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-068.

Ball:1973:NSM

[921] Ian R. Ball. A new species of marine triclad turbellarian of the genus Sabussowia from Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):389–394, March 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-069.

McLeay:1973:EHD

[922] D. J. McLeay. Effects of a 12-hr and 25-day exposure to kraft pulp mill effluent on the blood and tissues of juvenile Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):395– 400, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-070.

Ihssen:1973:ITR

[923] Peter Ihssen. Inheritance of thermal resistance in hybrids of Salvelinus fontinalis and S. namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):401–408, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-071.

Ricker:1973:LRF

[924] W. E. Ricker. Linear regressions in fishery research. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):409–434, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-072.

Laurence:1973:ITE

[925] Geoffrey C. Laurence. Influence of temperature on energy utilization of embryonic and prolarval tautog, Tautoga onitis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):435–442, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-073.

Hiscox:1973:EPG

[926] J. I. Hiscox and R. W. Brocksen. Effects of a parasitic gut nematode on consumption and growth in juvenile rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):443–450, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-074.

Scott:1973:FIF

[927] J. S. Scott. Food and inferred feeding behavior of northern sand lance (Ammodytes dubius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3): 451–454, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-075.

Freeman:1973:TMM

[928] H. C. Freeman and D. A. Horne. Total mercury and methylmercury content of the American eel (Anguilla rostrata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):454–456, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-076.

Alt:1973:AGI

[929] Kenneth T. Alt. Age and growth of the inconnu (Stenodus leucichthys) in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):457– 459, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-077.

Cheng:1973:SMV

[930] Joseph Y. Cheng. A simple microscale vacuum collector for the elution of closely situated spots from thin-layer chromatograms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):460–461, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-078.

Lech:1973:PPT

[931] John J. Lech. Preparation and properties of 3-trifluoromethyl-4-aminophenol. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3): 461–463, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-079.

Tomkiewicz:1973:UVB

[932] Warren Tomkiewicz and Ian R. Ball. Uteriporus vulgaris Bergendal, 1890, a marine triclad new to North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):464–466, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-080.

Foda:1973:CHH

[933] A. Foda. Changes in hematocrit and hemoglobin in Atlantic salmon (Salmo salar) as a result of furunculosis disease. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):467–468, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-081.

Margolis:1973:ANP

[934] L. Margolis. Additional notes on polychaetes as intermediate hosts of helminth parasites of vertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):469–470, March 1973. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f73-082.

Gaskin:1973:MDP

[935] D. E. Gaskin, R. Frank, M. Holdrinet, K. Ishida, C. J. Walton, and M. Smith. Mercury, DDT, and PCB in harbour seals (Phoca vitulina) from the Bay of Fundy and Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):471–475, March 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-083.

Anonymous:1973:RAFc

[936] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(3):477, March 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-084.

Leach:1973:ICK

[937] J. M. Leach and A. N. Thakore. Identification of the constituents of kraft pulping effluent that are toxic to juvenile Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):479– 484, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-085.

MacLeod:1973:TEM

[938] J. C. MacLeod and E. Pessah. Temperature effects on mercury accumulation, toxicity, and metabolic rate in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):485–492, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-086.

Shorey:1973:MES

[939] Wayne K. Shorey. Macrobenthic ecology of a sawdust-bearing substrate in the Penobscot River Estuary (Maine). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):493–497, April 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-087.

Webb:1973:ESC

[940] P. W. Webb and J. R. Brett. Effects of sublethal concentrations of sodium pentachlorophenate on growth rate, food conversion efficiency, and swimming performance in underyearling sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4): 499–507, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-088.

Anderson:1973:OEY

[941] Lee G. Anderson. Optimum economic yield of a fishery given a variable price of output. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4): 509–518, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-089.

Patalas:1973:CPE

[942] Kazimierz Patalas and Alex Salki. Crustacean plankton and the eutrophication of lakes in the Okanagan Valley, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):519–542, April 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-090.

Blackett:1973:FRA

[943] R. F. Blackett. Fecundity of resident and anadromous Dolly Varden (Salvelinus malma) in Southeastern Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):543–548, April 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-091.

Odense:1973:SNR

[944] P. H. Odense, V. H. Logan, and S. R. Baker. Spontaneous nephroblastoma in a rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):549–551, April 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-092.

Arai:1973:DMB

[945] Mary Needler Arai and John Fulton. Diel migration and breeding cycle of Aglantha digitale from two locations in the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):551–553, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-093.

Alt:1973:NWC

[946] Kenneth T. Alt and Dennis R. Kogl. Notes on the whitefish of the Colville River, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4): 554–556, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-094.

Nelson:1973:OHB

[947] Joseph S. Nelson. Occurrence of hybrids between longnose sucker (Catostomus catostomus) and white sucker (C. commersoni) in Upper Kananaskis Reservoir, Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):557–560, April 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-095.

Jackim:1973:ILO

[948] Eugene Jackim. Influence of lead and other metals on fish δaminolevulinate dehydrase activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):560–562, April 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-096.

Fried:1973:OSS

[949] Stephen M. Fried and James D. McCleave. Occurrence of the shortnose sturgeon (Acipenser brevirostrum), an endangered species, in Montsweag Bay, Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):563– 564, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-097.

Roos:1973:PJP

[950] J. F. Roos, P. Gilhousen, S. R. Killick, and E. R. Zyblut. Parasitism on juvenile Pacific salmon (Oncorhynchus) and Pacific herring (Clupea harengus pallasi) in the Strait of Georgia by the river lamprey (Lampetra ayresi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4): 565–570, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-098.

Plumb:1973:ETM

[951] John A. Plumb. Effects of temperature on mortality of fingerling channel catfish (Ictalurus punctatus) experimentally infected with channel catfish virus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4): 568–570, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-099.

Wobeser:1973:ORD

[952] G. Wobeser. An outbreak of redmouth disease in rainbow trout (Salmo gairdneri) in Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):571–575, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-100.

Ahmed:1973:OCT

[953] N. U. Ahmed and N. D. Georganas. Optimal control theory applied to a dynamic aquatic ecosystem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):576–579, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-101.

Ayles:1973:PAS

[954] G. B. Ayles and A. H. Berst. Parental age and survival of progeny in splake hybrids (Salvelinus fontinalis × S. namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):579–582, April 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-102.

Brungs:1973:ALT

[955] William A. Brungs, Edward N. Leonard, and James M. McKim. Acute and long-term accumulation of copper by the brown bullhead, Ictalurus nebulosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4): 583–586, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-103.

Beamish:1973:DTI

[956] R. J. Beamish. Design of a trapnet with interchangeable parts for the capture of large and small fishes from varying depths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):587–590, April 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-104.

Anonymous:1973:RAFd

[957] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(4):591– 592, April 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-105.

Addison:1973:VOR

[958] R. F. Addison, S. R. Kerr, J. Dale, and D. E. Sergeant. Variation of organochlorine residue levels with age in Gulf of St. Lawrence harp seals (Pagophilus groenlandicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):595–600, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-106.

Farmer:1973:SLP

[959] G. J. Farmer and F. W. H. Beamish. Sea lamprey (Petromyzon marinus) predation on freshwater teleosts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):601–605, May 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-107.

Beamish:1973:DAG

[960] Richard J. Beamish. Determination of age and growth of populations of the white sucker (Catostomus commersoni) exhibiting a wide range in size at maturity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5): 607–616, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-108.

Chua:1973:EII

[961] K. E. Chua and R. O. Brinkhurst. Evidence of interspecific interactions in the respiration of tubificid oligochaetes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):617–622, May 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-109.

Carline:1973:EME

[962] Robert F. Carline and James D. Hall. Evaluation of a method for estimating food consumption rates of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):623–629, May 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-110.

Southcott:1973:CBC

[963] B. A. Southcott and W. E. Razzell. Clostridium botulinum control in cold-smoked salmon: a review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):631–641, May 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-111.

Uthe:1973:UMC

[964] J. F. Uthe, F. M. Atton, and L. M. Royer. Uptake of mercury by caged rainbow trout (Salmo gairdneri) in the South Saskatchewan River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):643–650, May 1973. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f73-112.

Symons:1973:BYA

[965] P. E. K. Symons. Behavior of young Atlantic salmon (Salmo salar) exposed to or force-fed fenitrothion, an organophosphate insecticide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):651–655, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-113.

Rommel:1973:SAE

[966] S. A. Rommel, Jr. and J. D. McCleave. Sensitivity of American eels (Anguilla rostrata) and Atlantic salmon (Salmo salar) to weak electric and magnetic fields. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5): 657–663, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-114.

Saether:1973:TET

[967] Ole A. Sæther. Taxonomy and ecology of three new species of Monodiamesa Kieffer, with keys to Nearctic and Palaearctic species of the genus (Diptera: Chironomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):665–679, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-115.

Wobeser:1973:OCD

[968] G. Wobeser and F. M. Atton. An outbreak of columnaris disease in white suckers (Catostomus commersoni) in Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):681–683, May 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-116.

Gruchy:1973:FRSa

[969] C. G. Gruchy, R. H. Bowen, and I. M. Gruchy. First records of the stoneroller (Campostoma anomalum) and the blackstripe topminnow (Fundulus notatus) from Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):683–684, May 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-117.

Manohar:1973:EST

[970] S. V. Manohar, D. L. Rigby, and L. C. Dugal. Effect of sodium tripolyphosphate on thaw drip and taste of fillets of some freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):685–688, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-118.

Vladykov:1973:MAE

[971] Vadim D. Vladykov. Macrophthalmia in the American eel (Anguilla rostrata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):689– 693, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-119.

Rommel:1973:SMR

[972] S. A. Rommel, Jr. A simple method for recording fish heart and operculum beats without the use of implanted electrodes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):693–694, May 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-120.

Ssentongo:1973:SSM

[973] G. W. Ssentongo and P. A. Larkin. Some simple methods of estimating mortality rates of exploited fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):695–698, May 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-121.

Drummond:1973:SST

[974] R. A. Drummond, W. A. Spoor, and G. F. Olson. Some short-term indicators of sublethal effects of copper on brook trout, Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):698–701, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-122.

Smith:1973:ECC

[975] Waldina E. Smith, Kaye Funk, and Mary E. Zabik. Effects of cooking on concentrations of PCB and DDT compounds in Chinook (Oncorhynchus tshawytscha) and Coho (O. kisutch) salmon from Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):702–706, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-123.

Childs:1973:EFM

[976] E. A. Childs. Electrophoresis of fish myofibrillar proteins with sodiumdodecyl sulfate-polyacrylamide gel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):706–707, May 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-124.

Ward:1973:CLS

[977] F. J. Ward and Masami Nakanishi. A comparison of liquid scintillation and Geiger–Müller estimates of primary productivity in an in situ experiment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):708– 711, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-125.

Ackman:1973:SDS

[978] R. G. Ackman and D. Noble. Steam distillation: A simple technique for recovery of petroleum hydrocarbons from tainted fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):711–714, May 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-126.

Anonymous:1973:RAFe

[979] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(5):715, May 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-127.

Holeton:1973:RAC

[980] G. F. Holeton. Respiration of Arctic char (Salvelinus alpinus) from a High Arctic lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 717–723, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-128.

Kleerekoper:1973:ITC

[981] H. Kleerekoper, J. B. Waxman, and J. Matis. Interaction of temperature and copper ions as orienting stimuli in the locomotor behavior of the goldfish (Carassius auratus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):725–728, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-129.

Kabata:1973:SLC

[982] Z. Kabata. The species of Lepeophtheirus (Copepodar Caligidae) from fishes of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):729–759, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-130.

Emery:1973:PCD

[983] Alan R. Emery. Preliminary comparisons of day and night habits of freshwater fish in Ontario lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):761–774, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-131.

McLeese:1973:CCA

[984] D. W. McLeese. Chemical communication among lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 775–778, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-132.

Ruggles:1973:RCS

[985] C. P. Ruggles and G. E. Turner. Recent changes in stock composition of Atlantic salmon (Salmo salar) in the Miramichi River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):779–786, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-133.

Ware:1973:REP

[986] D. M. Ware. Risk of epibenthic prey to predation by rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 787–797, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-134.

Colby:1973:ETE

[987] Peter J. Colby and L. T. Brooke. Effects of temperature on embryonic development of lake herring (Coregonus artedii). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):799–810, June 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-135.

Fulton:1973:SAL

[988] John Fulton. Some aspects of the life history of Calanus plumchrus in the Strait of Georgia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 811–815, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-136.

Hamley:1973:DEG

[989] John M. Hamley and Henry A. Regier. Direct estimates of gillnet selectivity to walleye (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):817–830, June 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-137.

Wedemeyer:1973:SPA

[990] Gary Wedemeyer. Some physiological aspects of sublethal heat stress in the juvenile steelhead trout (Salmo gairdneri) and Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 831–834, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-138.

Clasby:1973:SMM

[991] R. C. Clasby, Rita Horner, and Vera Alexander. An in situ method for measuring primary productivity of Arctic sea ice algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):835–838, June 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-139.

McLeese:1973:OLH

[992] D. W. McLeese. Orientation of lobsters (Homarus americanus) to odor. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):838–840, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-140.

Ward:1973:MCT

[993] James V. Ward. Molybdenum concentrations in tissues of rainbow trout (Salmo gairdneri) and Kokanee salmon (Oncorhynchus nerka) from waters differing widely in molybdenum content. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):841–842, June 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-141.

Sergeant:1973:MSE

[994] D. E. Sergeant and F. A. J. Armstrong. Mercury in seals from Eastern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):843– 846, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-142.

OHara:1973:CUF

[995] James O’Hara. Cadmium uptake by fiddler crabs exposed to temperature and salinity stress. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 846–848, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-143.

Siefert:1973:ERO

[996] Richard E. Siefert, William A. Spoor, and Roll F. Syrett. Effects of reduced oxygen concentrations on northern pike (Esox lucius) embryos and larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):849– 852, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-144.

Colgan:1973:RWD

[997] Patrick Colgan and David Ealey. Role of woody debris in nest site selection by pumpkinseed sunfish, Lepomis gibbosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):853–856, June 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-145.

Sutcliffe:1973:CBS

[998] W. H. Sutcliffe, Jr. Correlations between seasonal river discharge and local landings of American lobster (Homarus americanus) and Atlantic halibut (Hippoglossus hippoglossus) in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6):856–859, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-146.

Winters:1973:NRE

[999] G. H. Winters, J. A. Moores, and R. Chaulk. Northern range extension and probable spawning of gaspereau (Alosa pseudoharengus) in the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(6): 860–861, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-147.

Anonymous:1973:RAFf nal de l’Office des recherches sur les pêcheries du Canada, 30(6):863, June 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-148.

Bourget:1973:ASF

[1001] E. Bourget and G. Lacroix. Aspects saisonniers de la fixation de l’épifaune benthique de l’étage infralittoral de l’estuaire du Saint-Laurent. (French) [Seasonal aspects of the fixation of the benthic epifauna of the infralittoral stage of the Saint-Lawrence Estuary]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):867–880, July 1973. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f73-149.

Kabata:1973:LCS

[1002] Z. Kabata and B. Cousens. Life cycle of Salmincola californiensis (Dana 1852) (Copepoda: Lernaeopodidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):881–903, July 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-150.

Kaya:1973:ETR

[1003] Calvin M. Kaya. Effects of temperature on responses of the gonads of green sunfish (Lepomis cyanellus) to treatment with carp pituitaries and testosterone propionate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):905–912, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-151.

Nicholson:1973:ECL

[1004] Bruce L. Nicholson and Charles Byrne. An established cell line from the Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):913–916, July 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-152.

Blanton:1973:SCT

[1005] J. O. Blanton. Some comparisons in the thermal structure of Lakes Wood, Kalamalka, Okanagan, Skaha, and Osoyoos, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):917–925, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-153.

Clayton:1973:ISM

[1006] J. W. Clayton, R. E. K. Harris, and D. N. Tretiak. Identification of supernatant and mitochondrial isozymes of malate dehydrogenase on electropherograms applied to the taxonomic discrimination of walleye (Stizostedion vitreum vitreum), sauger (S. canadense), and suspected interspecific hybrid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7): 927–938, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-154.

Walter:1973:DDE

[1007] Gilbert G. Walter. Delay-differential equation models for fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):939–945, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-155.

Smith:1973:OCE

[1008] H. D. Smith. Observations on the cestode Eubothrium salvelini in juvenile sockeye salmon (Oncorhynchus nerka) at Babine Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):947–964, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-156.

Coutant:1973:ETS

[1009] Charles C. Coutant. Effect of thermal shock on vulnerability of juvenile salmonids to predation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):965–973, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-157.

Hokanson:1973:TRM

[1010] Kenneth E. F. Hokanson, J. Howard McCormick, Bernard R. Jones, and James H. Tucker. Thermal requirements for maturation, spawning, and embryo survival of the brook trout, Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):975–984, July 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-158.

Sutterlin:1973:CBH

[1011] A. M. Sutterlin and R. Gray. Chemical basis for homing of Atlantic salmon (Salmo salar) to a hatchery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):985–989, July 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-159.

Diaz:1973:EBT

[1012] Robert J. Diaz. Effects of brief temperature increases on larvae of the American oyster (Crassostrea virginica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):991–993, July 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-160.

Ross:1973:SHF

[1013] A. J. Ross and W. T. Yasutake. Scolecobasidium humicola, a fungal pathogen of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7): 994–995, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-161.

Newcombe:1973:SCS

[1014] C. Newcombe and G. Hartman. Some chemical signals in the spawning behavior of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):995–997, July 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-162.

Jensen:1973:RBS

[1015] A. L. Jensen. Relation between simple dynamic pool and surplus production models for yield from a fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):998–1002, July 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-163.

Kaczynski:1973:TAJ

[1016] V. W. Kaczynski, R. J. Feller, J. Clayton, and R. J. Gerke. Trophic analysis of juvenile pink and chum salmon (Oncorhynchus gorbuscha and O. keta) in Puget Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1003–1008, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-164.

Sergeant:1973:AGM

[1017] D. E. Sergeant, David K. Caldwell, and Melba C. Caldwell. Age, growth, and maturity of bottlenosed dolphin (Tursiops truncatus) from Northeast Florida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1009– 1011, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-165.

OConnor:1973:HBT

[1018] J. F. O’Connor and G. Power. Homing of brook trout (Salvelinus fontinalis) in Matamek Lake, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1012–1014, July 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-166.

Parsons:1973:PSM

[1019] T. R. Parsons, C. A. Bawden, and W. A. Heath. Preliminary survey of mercury and other metals contained in animals from the Fraser River mudflats. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1014– 1016, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-167.

Lee:1973:OLC

[1020] D. J. Lee and J. H. Wales. Observed liver changes in rainbow trout (Salmo gairdneri) fed varying levels of a casein–gelatin mixture and herring oil in experimental diets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1017–1020, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-168.

Sosiak:1973:IMU

[1021] A. J. Sosiak. An improved method of using one-way glass to assess visual discrimination in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1021–1022, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-169.

Tsuyuki:1973:PAR

[1022] H. Tsuyuki and S. N. Williscroft. The pH activity relations of two LDH homotetramers from trout liver and their physiological significance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1023–1026, July 1973. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f73-170.

Rudd:1973:DNN

[1023] John W. M. Rudd, B. E. Townsend, and R. D. Hamilton. Discharge of nitrilotriacetate (NTA) from two sewage treatment facilities located in a midcontinental climate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1026–1030, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-171.

Edmunds:1973:SGP

[1024] Philip H. Edmunds and James I. Sammons III. Similarity of Genie polymorphism of tetrazolium oxidase in bluefin tuna (Thunnus thynnus) from the Atlantic Coast of France and the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1031–1032, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-172.

Power:1973:ELE

[1025] G. Power and Lloyd H. Saunders. An error in logic in estimating stream populations of premigrant trout from stream survival rates and lake populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1033– 1034, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-173.

Anonymous:1973:RAFg

[1026] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(7):1035– 1036, July 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-174.

Keizer:1973:DTA

[1027] P. D. Keizer and D. C. Gordon, Jr. Detection of trace amounts of oil in sea water by fluorescence spectroscopy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1039–1046, August 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-175.

Olson:1973:TTL

[1028] L. E. Olson and L. L. Marking. Toxicity of TFM (Lampricide) to six early life stages of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1047–1052, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-176.

Frank:1973:ORH

[1029] R. Frank, K. Ronald, and H. E. Braun. Organochlorine residues in harp seals (Pagophilus groenlandicus) caught in Eastern Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1053–1063, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-177.

Sergeant:1973:BWW

[1030] D. E. Sergeant. Biology of white whales (Delphinapterus leucas) in Western Hudson Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8): 1065–1090, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-178.

Peterson:1973:TSA

[1031] R. H. Peterson. Temperature selection of Atlantic salmon (Salmo salar) and brook trout (Salvelinus fontinalis) as influenced by various chlorinated hydrocarbons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8): 1091–1097, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-179.

McBride:1973:UMP

[1032] J. R. McBride and U. H. M. Fagerlund. The use of 17α-methyltestosterone for promoting weight increases in juvenile Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1099–1104, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-180.

Arai:1973:BPC

[1033] Mary Needler Arai. Behavior of the planktonic coelenterates, Sarsia tubulosa, Phialidium gregarium, and Pleurobrachia pileus in salinity discontinuity layers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8): 1105–1110, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-181.

Mercer:1973:ODI

[1034] M. C. Mercer. Observations on distribution and intraspecific variation in pigmentation patterns of odontocete cetacea in the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1111–1130, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-182.

Falkner:1973:IIB

[1035] N. W. Falkner and J. G. Eales. Investigation of iodothyronine binding to plasma proteins in brook trout, Salvelinus fontinalis, using precipitation, dialysis, and electrophoretic methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1131–1140, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-183.

Sivak:1973:IFB

[1036] Jacob G. Sivak. Interrelation of feeding behavior and accommodative lens movements in some species of North American freshwater fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1141–1146, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-184.

Smith:1973:ELN

[1037] E. J. Smith, J. L. Sykora, and M. A. Shapiro. Effect of lime neutralized iron hydroxide suspensions on survival, growth, and reproduction of the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1147–1153, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-185.

Dechtiar:1973:NPD

[1038] Alex O. Dechtiar. New and previously described species of the genus Lyrodiscus (Monogenoidea: Ancyrocephalinae) from fishes of the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1155–1160, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-186.

Nursall:1973:SBI

[1039] J. R. Nursall. Some behavioral interactions of spottail shiners (Notropis hudsonius), yellow perch (Perca flavescens), and northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1161–1178, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-187.

Cameron:1973:PEB

[1040] J. N. Cameron, J. Kostoris, and P. A. Penhale. Preliminary energy budget of the ninespine stickleback (Pungitius pungitius) in an Arctic lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1179–1189, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-188.

Shelbourn:1973:ETF

[1041] J. E. Shelbourn, J. R. Brett, and S. Shirahata. Effect of temperature and feeding regime on the specific growth rate of sockeye salmon fry (Oncorhynchus nerka), with a consideration of size effect. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1191–1194, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-189.

Marcy:1973:VSY

[1042] Barton C. Marcy, Jr. Vulnerability and survival of young Connecticut River fish entrained at a nuclear power plant. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1195–1203, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-190.

Castell:1973:EFS

[1043] C. H. Castell, B. Smith, and W. J. Dyer. Effects of formaldehyde on salt extractable proteins of gadoid muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1205–1213, August 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-191.

McCart:1973:LHT

[1044] P. McCart and P. Craig. Life history of two isolated populations of Arctic char (Salvelinus alpinus) in spring-fed tributaries of the Canning River, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8): 1215–1220, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-192.

House:1973:AGS

[1045] Robert House and Larue Wells. Age, growth, spawning season, and fecundity of the trout-perch (Percopsis omiscomaycus) in Southeastern Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8): 1221–1225, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-193.

Horning:1973:GTR

[1046] W. B. Horning II and R. E. Pearson. Growth temperature requirements and lower lethal temperatures for juvenile smallmouth bass (Micropterus dolomieui). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1226– 1230, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f73-194.

Westrheim:1973:PIS

[1047] S. J. Westrheim. Preliminary information on the systematics, distribution, and abundance of the dusky rockfish, Sebastes ciliatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1230–1234, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-195.

Iredale:1973:MNM

[1048] D. G. Iredale and K. J. Shaykewich. Masking or neutralizing muddy flavor in rainbow trout (Salmo gairdneri) by smoking and canning processes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1235–1239, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-196.

LeGore:1973:AAE

[1049] Richard S. LeGore and David M. DesVoigne. Absence of acute effects on threespine sticklebacks (Gasterosteus aculeatus) and Coho salmon (Oncorhynchus kisutch) exposed to resuspended harbor sediment contaminants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1240–1242, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-197.

Tucker:1973:RDO

[1050] D. Tucker. Rapid decline of olfactory and gustatory receptor sensitivities of wild catfish (Ictaluridae) after capture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1243–1245, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-198.

Castell:1973:CCT

[1051] C. H. Castell, W. E. Neal, and J. Dale. Comparison of changes in trimethylamine, dimethylamine, and extractable protein in iced and frozen gadoid fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8): 1246–1248, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-199.

Emadi:1973:YSM

[1052] Hossein Emadi. Yolk-sac malformation in Pacific salmon in relation to substrate, temperature, and water velocity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1249–1250, August 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-200.

Freeman:1973:SEM

[1053] H. C. Freeman and D. A. Horne. Sampling the edible muscle of the swordfish (Xiphias gladius) for total mercury analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1251–1252, August 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-201.

McCauley:1973:TSJ

[1054] R. W. McCauley and L. A. A. Read. Temperature selection by juvenile and adult yellow perch (Perca flavescens) acclimated to 24 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1253–1255, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-202.

Tomlinson:1973:SRP

[1055] N. Tomlinson and E. G. Baker. Sexual ripening of Pacific halibut (Hippoglossus stenolepis) in captivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1255–1256, August 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-203.

Anonymous:1973:CRP

[1056] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1257, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-204.

Anonymous:1973:RAFh

[1057] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(8):1259, August 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-205.

Pitt:1973:FAP

[1058] T. K. Pitt. Food of American plaice (Hippoglossoides platessoides) from the Grand Bank, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1261–1273, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-206.

Ricker:1973:TMM

[1059] W. E. Ricker. Two mechanisms that make it impossible to maintain peak-period yields from Stocks of Pacific salmon and other fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1275–1286, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-207.

Cross:1973:RBT

[1060] F. A. Cross, L. H. Hardy, N. Y. Jones, and R. T. Barber. Relation between total body weight and concentrations of manganese, iron, copper, zinc, and mercury in white muscle of bluefish (Pomatomus saltatrix) and a bathyl-demersal fish Antimora rostrata. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1287–1291, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-208.

Olson:1973:UMM

[1061] Kenneth R. Olson, Harold L. Bergman, and Paul O. Fromm. Uptake of methyl mercuric chloride and mercuric chloride by trout: A study of uptake pathways into the whole animal and uptake by erythrocytes in vitro. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1293–1299, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-209.

Schmidt:1973:MBE

[1062] Dana C. Schmidt and L. J. Weber. Metabolism and biliary excretion of sulfobromophthalein by rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1301–1308, September 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-210.

Stasko:1973:CMM

[1063] A. B. Stasko, R. M. Horrall, A. D. Hasler, and D. Stasko. Coastal movements of mature Fraser River pink salmon (Oncorhynchus gorbuscha) as revealed by ultrasonic tracking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1309–1316, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-211.

Hargrave:1973:CCF

[1064] Barry T. Hargrave. Coupling carbon flow through some pelagic and benthic communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9): 1317–1326, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-212.

Swenson:1973:GDF

[1065] William A. Swenson and Lloyd L. Smith, Jr. Gastric digestion, food consumption, feeding periodicity, and food conversion efficiency in walleye (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1327–1336, September 1973. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f73-213.

Aiken:1973:PSD

[1066] D. E. Aiken. Proecdysis, setal development, and molt prediction in the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1337–1344, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-214.

Green:1973:GMA

[1067] Roger H. Green. Growth and mortality in an Arctic intertidal population of Macoma balthica (Pelecypoda, Tellinidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1345–1348, September 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-215.

Ennis:1973:BRC

[1068] G. P. Ennis. Behavioral responses to changes in hydrostatic pressure and light during larval development of the lobster Homarus gammarus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1349–1360, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-216.

Anonymous:1973:Ob

[1069] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1361, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-217.

Mulligan:1973:PCR

[1070] Hugh Francis Mulligan. Probable causes for the 1972 Red Tide in the Cape Ann region of the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1363–1366, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-218.

Beamish:1973:OCA

[1071] F. W. H. Beamish. Oxygen consumption of adult Petromyzon marinus in relation to body weight and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1367–1370, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-219.

Scakratt:1973:CLO

[1072] D. J. Scakratt. Claw loss and other wounds in commercially caught lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1370–1373, September 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-220.

Wells:1973:IFW

[1073] Barbara Wells, D. H. Steele, and A. V. Tyler. Intertidal feeding of winter flounders (Pseudopleuronectes americanus) in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1374–1378, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-221.

Gruchy:1973:FRSb

[1074] C. G. Gruchy, R. H. Bowen, and I. M. Gruchy. First records of the silver shiner, Notropis photogenis, from Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1379–1382, September 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-222.

Ossiander:1973:CPS

[1075] Frank J. Ossiander and Gary Wedemeyer. Computer program for sample sizes required to determine disease incidence in fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1383–1384, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-223.

Black:1973:GRI

[1076] Robert Black. Growth rates of intertidal molluscs as indicators of effects of unexpected incidents of pollution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1385–1388, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-224.

Barica:1973:RAP

[1077] Jan Barica. Reliability of an ammonia probe for electrometric determination of total ammonia nitrogen in fish tanks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1389–1392, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-225.

Macdonald:1973:SNW

[1078] J. Rod Macdonald and R. A. Hyatt. Supersaturation of nitrogen in water during passage through hydroelectric turbines at Mactaquac Dam. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1392–1394, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-226.

Bibor:1973:MIT

[1079] V. Bibor and C. Leray. Métabolisme de l’iode et de la thyroxine chez Perca flavescens. (French) [Metabolism of iodine and thyroxine in Perca flavescens]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1395–1398, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-227.

OConnor:1973:TPE

[1080] J. F. O’Connor and G. Power. Trout production and eels in Bill Lake, Saguenay County, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1398–1401, September 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-228.

Luke:1973:PSU

[1081] D. McG. Luke, D. G. Pincock, and A. B. Stasko. Pressure-sensing ultrasonic transmitter for tracking aquatic animals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1402–1404, September 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-229.

Regier:1973:AMR

[1082] L. W. Regier. Addendum: Mercury removal from fish protein concentrate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1404, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-230.

Anonymous:1973:RAFi

[1083] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(9):1405–1406, September 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-231.

Schindler:1973:EAL

[1084] D. W. Schindler. Experimental approaches to limnology — an overview. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1409–1413, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-232.

Schindler:1973:ELA

[1085] D. W. Schindler, H. Kling, R. V. Schmidt, J. Prokopowich, V. E. Frost, R. A. Reid, and M. Capel. Eutrophication of Lake 227 by addition of phosphate and nitrate: the second, third, and fourth years of enrichment, 1970, 1971, and 1972. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10): 1415–1440, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-233.

Stainton:1973:SGS

[1086] Michael P. Stainton. A syringe gas-stripping procedure for gaschromatographic determination of dissolved inorganic and organic carbon in fresh water and carbonates in sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1441–1445, October 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-234.

Fee:1973:NMD

[1087] Everett J. Fee. A numerical model for determining integral primary production and its application to Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1447–1468, October 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-235.

Fee:1973:MPP

[1088] Everett J. Fee. Modelling primary production in water bodies: A numerical approach that allows vertical inhomogeneities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1469–1473, October 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-236.

Emerson:1973:GER

[1089] Steve Emerson, Wallace Broecker, and D. W. Schindler. Gas-exchange rates in a small lake as determined by the radon method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1475–1484, October 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-237.

Emerson:1973:DUA

[1090] S. Emerson and R. Hesslein. Distribution and uptake of artificially introduced radium-226 in a small lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1485–1490, October 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-238.

Hesslein:1973:VED

[1091] R. Hesslein and P. Quay. Vertical eddy diffusion studies in the thermocline of a small stratified lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1491–1500, October 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-239.

Schindler:1973:DVD

[1092] D. W. Schindler and E. J. Fee. Diurnal variation of dissolved inorganic carbon and its use in estimating primary production and CO_{2} invasion in Lake 227. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1501–1510, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-240.

Schindler:1973:PET

[1093] D. W. Schindler, V. E. Frost, and R. V. Schmidt. Production of epilithiphyton in two lakes of the Experimental Lakes Area, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1511–1524, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-241.

Lean:1973:MPB

[1094] D. R. S. Lean. Movements of phosphorus between its biologically important forms in lake water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1525–1536, October 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-242.

Rudd:1973:MAT

[1095] John W. M. Rudd and R. D. Hamilton. Measurement of adenosine triphosphate (ATP) in two Precambrian Shield lakes of Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1537–1546, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-243.

Thompson:1973:HUS

[1096] B. M. Thompson and R. D. Hamilton. Heterotrophic utilization of sucrose in an artificially enriched lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1547–1552, October 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-244.

Bilinski:1973:ECC

[1097] E. Bilinski and R. E. E. Jonas. Effects of cadmium and copper on the oxidation of lactate by rainbow trout (Salmo gairdneri) gills. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1553–1558, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-245.

Templeman:1973:DDN

[1098] Wilfred Templeman. Description and distribution of new specimens of the fish Lipogenys gilli from the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1559–1564, October 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-246.

Davis:1973:BPE

[1099] John C. Davis and Barbara J. Mason. Bioassay procedures to evaluate acute toxicity of neutralized bleached kraft pulp mill effluent to Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1565–1573, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-247.

Olson:1973:MUI

[1100] Kenneth R. Olson and Paul O. Fromm. Mercury uptake and ion distribution in gills of rainbow trout (Salmo gairdneri): Tissue scans with an electron microprobe. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10): 1575–1578, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-248.

Evelyn:1973:FRB

[1101] T. P. T. Evelyn, G. E. Hoskins, and G. R. Bell. First record of bacterial kidney disease in an apparently wild salmonid in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1578–1580, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-249.

Rommel:1973:ASM

[1102] S. A. Rommel, Jr. Apparatus for simulating motion of animals through the geomagnetic field. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10): 1581–1582, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-250.

Morrison:1973:GSA

[1103] Carol M. Morrison and Paul H. Odense. Gross structure of the adductor muscles of some pelecypods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1583–1585, October 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-251.

Power:1973:PCM

[1104] G. Power, G. F. Pope, and B. W. Coad. Postglacial colonization of the Matamek River, Quebec, by fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1586–1591, October 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-252.

Stirling:1973:VRS

[1105] Ian Stirling. Vocalization in the ringed seal (Phoca hispida). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1592–1594, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-253.

Scherer:1973:ARA

[1106] E. Scherer and S. Nowak. Apparatus for recording avoidance movements of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1594–1596, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-254.

Park:1973:RGI

[1107] Taisoo Park. Redescription of Gaetanus intermedius Campbell (Calanoida: Copepoda) from the type locality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1597–1600, October 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-255.

Sochasky:1973:DEA

[1108] Jeffrey B. Sochasky, D. E. Aiken, and D. W. McLeese. Does eyestalk ablation accelerate molting in the lobster Homarus americanus? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1600–1603, October 1973.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f73-256.

Worthington:1973:FAC

[1109] R. E. Worthington and R. T. Lovell. Fatty acids of channel catfish (Ictalurus punctatus): Variance components related to diet, replications within diets, and variability among fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1604–1608, October 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-257.

Anonymous:1973:RAFj

[1110] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(10):1609, October 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-258.

Gordon:1973:LSA

[1111] Donald C. Gordon, Jr., Paul D. Keizer, and Nicholas J. Prouse. Laboratory studies of the accommodation of some crude and residual fuel oils in sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11): 1611–1618, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-259.

Oloffs:1973:FAB

[1112] P. C. Oloffs, L. J. Albright, S. Y. Szeto, and J. Lau. Factors affecting the behavior of five chlorinated hydrocarbons in two natural waters and their sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11): 1619–1623, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-260.

Wolf:1973:FVI

[1113] Ken Wolf, M. C. Quimby, L. L. Pettijohn, and M. L. Landolt. Fish viruses: Isolation and identification of infectious hematopoietic necrosis in Eastern North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1625–1627, November 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-261.

Clark:1973:BHM

[1114] Colin Clark, Gordon Edwards, and Michael Friedlaender. Beverton– Holt model of a commercial fishery: Optimal dynamics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1629–1640, November 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-262.

Gratzek:1973:MBB

[1115] John B. Gratzek, Merrill H. McGlamery, Donald L. Dawe, and Trudy Scott. Microcultures of brown bullhead (Ictalurus nebulosus) cells: Their use in quantitation of channel catfish (Ictalurus punctatus) virus and antibody. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1641–1645, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-263.

Hoyle:1973:DES

[1116] R. J. Hoyle. Digestive enzyme secretion after dietary variation in the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1647–1653, November 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-264.

Templeman:1973:FRG

[1117] Wilfred Templeman. First records of the gymnoblastic hydroid, Ichthyocodium sarcotretis, on the copepod, Sphyrion lumpi, from redfish of the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11): 1655–1660, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-265.

Tinsley:1973:FAC

[1118] Ian J. Tinsley, Hugo M. Krueger, and James B. Saddler. Fatty acid content of Coho salmon, Oncorhynchus kisutch — a statistical approach to changes produced by diet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1661–1666, November 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-266.

Davidoff:1973:CPL

[1119] E. B. Davidoff, R. W. Rybicki, and K. H. Doan. Changes in the population of lake whitefish (Coregonus clupeaformis) in Lake Winnipeg from 1944 to 1969. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11): 1667–1682, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-267.

Bidgood:1973:DGT

[1120] B. F. Bidgood. Divergent growth in two lake whiteflsh (Coregonus clupeaformis) populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1683–1696, November 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-268.

Barans:1973:TSS

[1121] Charles A. Barans and Richard A. Tubb. Temperatures selected seasonally by four fishes from Western Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1697–1703, November 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-269.

Houston:1973:CRA

[1122] A. H. Houston, C. L. Czerwinski, and R. J. Woods. Cardiovascular– respiratory activity during recovery from anesthesia and surgery in brook trout (Salvelinus fontinalis) and carp (Cyprinus carpio). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1705–1712, November 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-270.

Landolt:1973:MCI

[1123] Marsha L. Landolt. Myxosoma cerebralis: Isolation and concentration from fish skeletal elements — trypsin digestion method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1713–1716, November 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-271.

MacKinnon:1973:AEF

[1124] J. C. MacKinnon. Analysis of energy flow and production in an unexploited marine flatfish population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1717–1728, November 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-272.

Templeman:1973:SRR

[1125] Wilfred Templeman. The skate, Raja richardsoni Garrick 1961, assigned to Bathyraja. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11): 1729–1732, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-273.

Addison:1973:ODR

[1126] R. F. Addison and P. F. Brodie. Occurrence of DDT residues in beluga whales (Delphinapterus leucas) from the Mackenzie Delta, N.W.T. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1733–1736, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-274.

Cimino:1973:KES

[1127] Michael C. Cimino. Karyotypes and erythrocyte sizes of some diploid and triploid fishes of the genus Poeciliopsis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1736–1737, November 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-275.

VanHyning:1973:IFE

[1128] J. M. Van Hyning and A. M. Scarborough. Identification of fungal encrustation on the shell of the snow crab (Chionoecetes bairdi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1738–1739, November 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-276.

Ritter:1973:IEE

[1129] John A. Ritter and Hugh R. MacCrimmon. Influence of environmental experience on response of yearling rainbow trout (Salmo gairdneri) to a black and white substrate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1740–1742, November 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-277.

Argyle:1973:EUR

[1130] Ray L. Argyle, George C. Williams, and Harry K. Dupree. Endrin uptake and release by fingerling channel catfish (Ictalurus punctatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1743–1744, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-278.

Onoue:1973:UPR

[1131] Y. Onoue and V. M. Riddle. Use of plastein reaction in recovering protein from fish waste. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11): 1745–1747, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-279.

Dormaar:1973:SAO

[1132] K. A. Dormaar and S. Corey. Some aspects of osmoregulation in Mysis stenolepis (Crustacea, Mysidacea). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1747–1749, November 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-280.

Lusz:1973:PWS

[1133] Larry D. Lusz and K. D. Waldron. A portable wire–speed indicator for use with plankton nets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1749–1751, November 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-281.

Wenner:1973:OAE

[1134] Charles A. Wenner. Occurrence of American eels, Anguilla rostrata, in waters overlying the Eastern North American Continental Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1752–1755, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-282.

Feldmeth:1973:EEE

[1135] C. Robert Feldmeth and Thomas M. Jenkins, Jr. An estimate of energy expenditure by rainbow trout (Salmo gairdneri) in a small mountain stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1755–1759, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-283.

Hauschild:1973:IDS

[1136] A. H. W. Hauschild, H. Pivnick, and L. W. Regier. Inability to detect spores of Clostridium botulinum in fish protein concentrates (FPC). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1760–1762, November 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-284.

Anonymous:1973:LRPa

[1137] Anonymous. List of recent publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(11):1765–1769, November 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-285.

Harris:1973:LIP

[1138] Graham P. Harris and John N. A. Lott. Light intensity and photosynthetic rates in phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1771–1778, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-286.

Harris:1973:DAC

[1139] G. P. Harris. Diel and annual cycles of net plankton photosynthesis in Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 1779–1787, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-287.

Hargrave:1973:CRI

[1140] B. T. Hargrave and C. P. Newcombe. Crawling and respiration as indices of sublethal effects of oil and a dispersant on an intertidal snail Littorina littorea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1789–1792, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-288.

Kabata:1973:DUC

[1141] Z. Kabata. Distribution of Udonella caligorum Johnston, 1835 (Monogenea: Udonellidae) on Caligus elongatus Nordmann, 1832 (Copepoda: Caligidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1793–1798, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-289.

Brett:1973:EES

[1142] J. R. Brett. Energy expenditure of sockeye salmon, Oncorhynchus nerka, during sustained performance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1799–1809, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-290.

Hermanutz:1973:CTF

[1143] Roger O. Hermanutz, Leonard H. Mueller, and Kenneth D. Kempfert. Captan toxicity to fathead minnows (Pimephales promelas), bluegills (Lepomis macrochirus), and brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1811–1817, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-291.

Scarratt:1973:ASV

[1144] D. J. Scarratt. Abundance, survival, and vertical and diurnal distribution of lobster larvae in Northumberland Strait, 1962–63, and their relationships with commercial stocks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1819–1824, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-292.

Walne:1973:GRN

[1145] P. R. Walne. Growth rates and nitrogen and carbohydrate contents of juvenile clams, Saxidomus giganteus, fed three species of algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1825–1830, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-293.

Templeman:1973:FRD

[1146] Wilfred Templeman. First records, description, distribution, and notes on the biology of Bathyraja richardsoni (Garrick) from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1831–1840, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-294.

Bothwell:1973:DLT

[1147] Max L. Bothwell, A. M. Beeton, and John J. Lech. Degradation of the lampricide 3-trifluoromethyl-4-nitrophenol by bottom sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1841–1846, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-295.

Dodson:1973:BAA

[1148] Julian J. Dodson and William C. Leggett. Behavior of adult American shad (Alosa sapidissima) homing to the Connecticut River from Long Island Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 1847–1860, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-296.

Tarr:1973:MHS

[1149] H. L. A. Tarr. 1-methyladenosine hydrolase of starfish (Pisaster ochraceous). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1861–1866, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-297.

Zeitoun:1973:ISP

[1150] I. H. Zeitoun, J. E. Halver, D. E. Ullrey, and P. I. Tack. Influence of salinity on protein requirements of rainbow trout (Salmo gairdneri) fingerlings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1867–1873, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-298.

Ritter:1973:EIB

[1151] John A. Ritter and Hugh R. MacCrimmon. Effects of illumination on behavior of wild brown (Salmo trutta) and rainbow trout (Salmo gairdneri) exposed to black and white backgrounds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1875–1880, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-299.

Korn:1973:DFC

[1152] Sid Korn and Dan Macedo. Determination of fat content in fish with a nontoxic, noninflammable solvent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1880–1881, December 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-300.

Dickman:1973:CPA

[1153] M. Dickman. Changes in periphytic algae following bicarbonate additions to a small stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 1882–1884, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-301.

Alt:1973:CBB

[1154] Kenneth T. Alt. Contributions to the biology of the Bering cisco (Coregonus laurettae) in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1885–1888, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-302.

Rojo:1973:SEL

[1155] Alfonso Rojo. The slender eelblenny, Lumpenus fabricii, a southern record in Nova Scotia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1888–1889, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-303.

Crouch:1973:OJS

[1156] Dennis E. Crouch, William T. Yasutake, and Robert R. Rucker. Openjaw syndrome in Chinook salmon (Oncorhynchus tshawytscha) at a hatchery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1890–1892, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-304.

Kabata:1973:TSH

[1157] Z. Kabata. Taxonomic status of Homoiotes palliata Wilson, 1905 (Copepoda: Caligidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 1892–1893, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-305.

Smith:1973:EEC

[1158] Roderick M. Smith and Charles F. Cole. Effects of egg concentrations of DDT and dieldrin on development in winter flounder (Pseudopleuronectes americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 1894–1898, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-306.

Withey:1973:ERP

[1159] Kathryn G. Withey and Richard L. Saunders. Effect of a reciprocal photoperiod regime on standard rate of oxygen consumption of postsmolt Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1898–1900, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-307.

Keenleyside:1973:EAH

[1160] Miles H. A. Keenleyside, R. K. Misra, and David W. Bateson. Extended analysis of hybridization in sunfishes (Centrarchidae) using an adjusted hybrid index method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 1901–1904, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-308.

Ennis:1973:FFC

[1161] G. P. Ennis. Food, feeding, and condition of lobsters, Homarus americanus, throughout the seasonal cycle in Bonavista Bay, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1905–1909, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-309.

Nepszy:1973:FRC

[1162] S. J. Nepszy and J. H. Leach. First records of the Chinese mitten crab, Eriocheir sinensis, (Crustacea: Brachyura) from North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1909–1910, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-310.

MacKay:1973:HRC

[1163] K. T. MacKay and J. Gilhen. Hirundichthys rondeleti, Cookeolus boops, Priacanthus arenatus, Seriola dumerili, four species new to the Canadian Atlantic ichthyofauna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1911–1913, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-311.

Anonymous:1973:LRPb

[1164] Anonymous. List of recent publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1916–1919, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-312.

Anonymous:1973:EF

[1165] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1925, December 1973. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f73-313.

Anonymous:1973:P

[1166] Anonymous. PREFACE. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1926, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-314.

C:1973:MOO

[1167] M. P. Hon. Jack Davis P. C. Managing our oceans for mankind. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1928–1930, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-315.

Popper:1973:KA

[1168] F. E. Popper. Keynote address. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1931–1934, December 1973. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f73-316.

Alverson:1973:OPM

[1169] D. L. Alverson and G. J. Paulik. Objectives and problems of managing aquatic living resources. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1936–1947, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-317.

Larkin:1973:CMF

[1170] Peter A. Larkin and Norman J. Wilimovsky. Contemporary methods and future trends in fishery management and development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1948–1957, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-318.

Lucas:1973:SAF

[1171] C. E. Lucas. Scientific advice to fisheries bodies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1958–1964, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-319.

Cushing:1973:DRP

[1172] D. H. Cushing. Dependence of recruitment on parent stock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1965–1976, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-320.

Garrod:1973:MMR

[1173] D. J. Garrod. Management of multiple resources. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1977–1985, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-321.

Fukuda:1973:GBT

[1174] Y. Fukuda. A gap between theory and practice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1986–1991, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-322.

Regier:1973:SES

[1175] H. A. Regier. Sequence of exploitation of stocks in multispecies fisheries in the Laurentian Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):1992–1999, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-323.

Henderson:1973:AFP

[1176] H. F. Henderson, R. A. Ryder, and A. W. Kudhongania. Assessing fishery potentials of lakes and reservoirs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2000–2009, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-324.

Saetersdal:1973:AUR

[1177] G. Saetersdal. Assessment of unexploited resources. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2010–2016, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-325.

Rothschild:1973:QSF

[1178] B. J. Rothschild. Questions of strategy in fishery management and development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2017–2030, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-326.

Lee:1973:FMN

[1179] A. J. Lee. Fishery management in the Northeast Atlantic: the role of the International Council for the Exploration of the Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2031–2037, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-327.

Anonymous:1973:RTSa

[1180] Anonymous. Report on technical session I. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2038–2040, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-328.

Gulland:1973:EFM

[1181] J. A. Gulland and M. A. Robinson. Economics of fishery management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2042–2050, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-329.

Robinson:1973:DDF

[1182] M. A. Robinson. Determinants of demand for fish and their effects upon resources. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2051–2058, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-330.

Luna-Munoz:1973:FRF

[1183] J. Luna-Muñoz. Financing resources for fishery development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2059–2064, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-331.

MacKenzie:1973:RCA

[1184] W. C. MacKenzie. Reconciling conflicts among different economic interest groups in the management of fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2065–2069, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-332.

Campbell:1973:LLR

[1185] B. A. Campbell. License limitation regulations: Canada’s experience. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2070–2076, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-333.

Gertenbach:1973:LLR

[1186] L. P. D. Gertenbach. License limitation regulations: The South African system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2077–2084, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-334.

Asada:1973:LLR

[1187] Y. Asada. License limitation regulations: The Japanese system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2085–2095, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-335.

Anonymous:1973:RTSb

[1188] Anonymous. Report on technical session II. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2096–2097, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-336.

Onarheim:1973:SPN

[1189] T. Onarheim. Structural problems and the need for structural planning in fisheries development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2100–2108, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-337.

Moiseev:1973:DFT

[1190] P. A. Moiseev. Development of fisheries for traditionally exploited species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2109–2120, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-338.

Suda:1973:DFN

[1191] A. Suda. Development of fisheries for nonconventional species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2121–2158, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-339.

Campleman:1973:TSS

[1192] G. Campleman. The transition from small-scale to large-scale industry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2159–2165, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-340.

Anderson:1973:DMU

[1193] A. M. Anderson. Developing markets for unfamiliar species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2166–2169, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-341.

Burgess:1973:AUF

[1194] G. H. O. Burgess. The alternative uses of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2170–2171, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-342.

George:1973:EPR

[1195] P. C. George. Experience and plans for rationalization of small-scale fisheries in India. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2172–2177, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-343.

Idyll:1973:MAP

[1196] C. P. Idyll. Marine aquaculture: Problems and prospects. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2178–2183, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-344.

Hempel:1973:PO

[1197] G. Hempel. Productivity of the oceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2184–2189, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-345.

Proude:1973:OMS

[1198] P. D. Proude. Objectives and methods of small-scale fisheries development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2190–2195, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-346.

Lyubimova:1973:PUK

[1199] T. G. Lyubimova, A. G. Naumov, and L. L. Lagunov. Prospects of the utilization of krill and other nonconventional resources of the world ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2196–2201, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-347.

Pillay:1973:RAF

[1200] T. V. R. Pillay. The role of aquaculture in fishery development and management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2202–2217, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-348.

Neal:1973:AAD

[1201] R. A. Neal. Alternatives in aquacultural development: Consideration of extensive versus intensive methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2218–2222, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-349.

Anonymous:1973:RTSc

[1202] Anonymous. Report on technical session III. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2223–2224, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-350.

Boerema:1973:SAP

[1203] L. K. Boerema and J. A. Gulland. Stock assessment of the Peruvian anchovy (Engraulis ringens) and management of the fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2226–2235, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-351.

Appleyard:1973:PCS

[1204] W. P. Appleyard. Peru: A case study in the establishment of a food fish industry for a developing country. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2236–2241, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-352.

Valdez-Zamudio:1973:IMR

[1205] F. Valdez-Zamudio. Impacto de medidas regulatorias en la industria pesquera peruana. (Spanish) [Impact of regulatory measures on the Peruvian fishing industry]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2242–2253, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-353.

Popiel:1973:IFT

[1206] J. Popiel and J. Sosinski. Industrial fisheries and their influence on catches for human consumption. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2254–2259, December 1973. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f73-354.

Anonymous:1973:RTSd

[1207] Anonymous. Report on technical session IV — section 1. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2260–2261, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-355.

Gulland:1973:MDF

[1208] J. A. Gulland, J. P. Troadec, and E. O. Bayagbona. Management and development of fisheries in the Eastern Central Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2264–2275, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-356.

Hamlisch:1973:FPM

[1209] R. Hamlisch. Fisheries planning to meet economic and social objectives with special reference to countries bordering the Eastern Central Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2276–2281, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-357.

Aubray:1973:DPR

[1210] R. Aubray, M. I. Dia, and B. Diop. La développement des pêches régionales en Afrique occidentale. (French) [The development of regional fisheries in West Africa]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2282–2292, December 1973. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f73-358.

Lacour:1973:PDP

[1211] P. Lacour. Les problèmes de développement de petites organisations de pêche. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2293–2300, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-359.

Bas:1973:CEP

[1212] C. Bas. Cambios y evolución de las pesquerías en al Atlántico Centro-Oriental. (Spanish) [Changes and evolution of fisheries in the Central-Eastern Atlantic]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2301–2307, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-360.

Anonymous:1973:RTSe

[1213] Anonymous. Report on technical session IV — section 2. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2308–2309, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-361.

Marr:1973:MDF

[1214] J. C. Marr. Management and development of fisheries in the Indian Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2312–2320, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-362.

Payne:1973:PCL

[1215] R. L. Payne. Planning criteria for large-scale fisheries development with special reference to the Indian Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2321–2327, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-363.

Honda:1973:FJV

[1216] K. Honda. Fishery joint ventures in developing countries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2328–2332, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-364.

Tiews:1973:FDM

[1217] K. Tiews. Fishery development and management in Thailand. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2333–2334, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-365.

Zachman:1973:FDM

[1218] N. Zachman. Fisheries development and management in Indonesia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2335–2340, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-366.

Jhingran:1973:PDB

[1219] V. G. Jhingran and V. Gopalakrishnan. Prospects for the development of brackish-water fish and shrimp culture in India. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2341–2343, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-367.

Anonymous:1973:RTSf

[1220] Anonymous. Report on technical session IV — section 3. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2344–2345, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-368.

Kasahara:1973:MFN

[1221] H. Kasahara. Management of fisheries in the North Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2348–2360, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-369.

Nagasaki:1973:LTS

[1222] F. Nagasaki. Long-term and short-term fluctuations in the catches of coastal pelagic fisheries around Japan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2361–2367, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-370.

Koh:1973:KCS

[1223] Yoon Ho Koh. Korea: A case study in promoting development through Fishermen’s training. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2368–2372, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-371.

Pruter:1973:DPS

[1224] A. T. Pruter. Development and present status of bottomfish resources in the Bering Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2373–2385, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-372.

Royce:1973:SMU

[1225] W. F. Royce. Salmon management in the United States. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2386–2392, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-373.

Skud:1973:MPH

[1226] B. E. Skud. Management of the Pacific halibut fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2393–2398, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-374.

Anonymous:1973:RTSg

[1227] Anonymous. Report on technical session IV — section 4. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2399–2400, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-375.

Mocklinghoff:1973:MDF

[1228] G. Möcklinghoff. Management and developement of fisheries in the North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2402–2418, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-376.

Chapman:1973:MIW

[1229] D. G. Chapman. Management of international whaling and North Pacific fur seals: Implications for fisheries management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2419–2426, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-377.

Edelman:1973:DNQ

[1230] M. S. Edelman and I. I. Dokuchaev. Development of national quotas by the International Commission for the Northwest Atlantic Fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2427–2435, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-378.

Bogdanov:1973:FRI

[1231] A. S. Bogdanov and K. G. Konstantinov. Fishery regulations in the ICNAF area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2436–2443, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-379.

Hennemuth:1973:MSH

[1232] R. C. Hennemuth. Management of sea herring fisheries in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2444–2451, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-380.

Anonymous:1973:RTSh

[1233] Anonymous. Report on technical session IV — section 5. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2452–2453, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-381.

Gulland:1973:DWF

[1234] J. A. Gulland. Distant-water fisheries and their relation to development and management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2456–2462, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-382.

Adjetey:1973:GED

[1235] J. N. N. Adjetey. Ghana’s experience in distant-water fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2463–2467, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-383.

Ahn:1973:EKH

[1236] Myong Nam Ahn. Expansion of Korean high-seas fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2468–2470, December 1973. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f73-384.

Joseph:1973:SMW

[1237] J. Joseph. Scientific management of the world stocks of tunas, billfishes, and related species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12): 2471–2482, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f73-385.

Anonymous:1973:RTSi

[1238] Anonymous. Report on technical session IV — section 6. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2483, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-386.

Elliot:1973:PCF

[1239] G. H. Elliot. Problems confronting fishing industries relative to management policies adopted by governments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2486–2489, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-387.

Eddie:1973:ITD

[1240] G. C. Eddie and A. D. Insull. Impact of technical development on the problems and opportunities in world fisheries and their management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2490–2495, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-388.

Dickie:1973:IBF

[1241] L. M. Dickie. Interaction between fishery management and environmental protection. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2496–2506, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-389.

Anonymous:1973:RCS

[1242] Anonymous. Report on the closing session. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2507, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-390.

Needler:1973:CSH

[1243] A. W. H. Needler. Chairman’s summary of the highlights of the conference. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2508–2511, December 1973. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f73-391.

Anonymous:1973:CR

[1244] Anonymous. Conference recommendations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 30(12):2512–2513, December 1973. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f73-392.

Messinger:1974:FAD

[1245] H. B. Messinger and H. T. Bilton. Factor analysis in discriminating the racial origin of sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):1–10, January 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-001.

Triplett:1974:CPC

[1246] E. Triplett and J. R. Calaprice. Changes in plasma constituents during spawning migration of Pacific salmons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):11–14, January 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-002.

Markiw:1974:MCI

[1247] Maria E. Markiw and Ken Wolf. Myxosoma cerebralis: Isolation and concentration from fish skeletal elements — sequential enzymatic digestions and purification by differential centrifugation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):15–20, January 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-003.

Gunderson:1974:ASC

[1248] Donald R. Gunderson. Availability, size composition, age composition, and growth characteristics of Pacific Ocean perch (Sebastes alutus) off the Northern Washington Coast during 1967–72. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):21–34, January 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-004.

Gee:1974:BDP

[1249] John H. Gee. Behavioral and developmental plasticity of buoyancy in the longnose, Rhinichthys cataractae, and blacknose, R. atratulus (Cyprinidae) dace. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (1):35–41, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-005.

Kabata:1974:LCS

[1250] Z. Kabata. Lepeophtheirus cuneifer sp. nov. (Copepoda: Caligidae), a parasite of fishes from the Pacific Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):43–47, January 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-006.

Beamish:1974:GSW

[1251] Richard J. Beamish. Growth and survival of white suckers (Catostomus commersoni) in an acidified lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):49–54, January 1974. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f74-007.

Loch:1974:PVL

[1252] J. S. Loch. Phenotypic variation in the lake whitefish, Coregonus clupeaformis, induced by introduction into a new environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):55–62, January 1974. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f74-008.

vanCoillie:1974:CME

[1253] R. van Coillie and A. Rousseau. Composition minérale des écailles du Catostomus commersoni issu de deux milieux différents: étude par microscopie électronique analytique. (French) [Mineral composition of Catostomus commersoni scales from two different media: study by analytical electron microscopy]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1): 63–66, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-009.

Kolehmainen:1974:DFR

[1254] Seppo E. Kolehmainen. Daily feeding rates of bluegill (Lepomis macrochirus) determined by a refined radioisotope method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):67–74, January 1974. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f74-010.

Bullock:1974:VVE

[1255] G. L. Bullock, H. M. Stuckey, Diane Collis, R. L. Herman, and G. Maestrone. In vitro and in vivo efficacy of a potentiated sulfonamide in control of furunculosis in salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):75–82, January 1974. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f74-011.

Mason:1974:BEC

[1256] J. C. Mason. Behavioral ecology of chum salmon fry (Oncorhynchus keta) in a small estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (1):83–92, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-012.

Anonymous:1974:Oa

[1257] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):93, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-013.

Castell:1974:PFT

[1258] J. D. Castell and D. J. Trider. Preliminary feeding trials using artificial diets to study the nutritional requirements of oysters (Crassostrea virginica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):95–99, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-014.

Goddard:1974:EMA

[1259] C. I. Goddard, J. W. Lilley, and J. S. Tait. Effects of M.S. 222 anesthetization on temperature selection in lake trout, Salvelinus namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):100–103, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-015.

Davies:1974:FRN

[1260] Ronald W. Davies and C. G. Chapman. First record from North America of the piscicolid leech, Ozobranchus margoi, a parasite of marine turtles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):104–106, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-016.

Huzyk:1974:DLB

[1261] L. Huzyk and H. Tsuyuki. Distribution of LDH-b˘2033 gene in resident and anadromous rainbow trout (Salmo gairdneri) from streams in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):106–108, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-017.

Selgeby:1974:IIP

[1262] James H. Selgeby. Immature insects (Plecoptera, Trichoptera, and Ephemeroptera) collected from deep water in Western Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):109–111, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-018.

Childs:1974:IFF

[1263] E. A. Childs. Interaction of free fatty acids with fish muscle in vitro. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):111–113, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-019.

Hurst:1974:MLE

[1264] R. E. Hurst. A method of lipid extraction suitable for use with an electron capture detector. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1): 113–116, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-020.

Alexander:1974:PMJ

[1265] D. Ross Alexander and Hugh R. MacCrimmon. Production and movement of juvenile rainbow trout (Salmo gairdneri) in a headwater of Bothwell’s Creek, Georgian Bay, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):117–121, January 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-021.

Potter:1974:SRL

[1266] I. C. Potter, F. W. H. Beamish, and B. G. H. Johnson. Sex ratios and lengths of adult sea lampreys (Petromyzon marinus) from a Lake Ontario tributary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(1):122–124, January 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-022.

Cauchois:1974:MOD

[1267] D. Cauchois and M. Khalil. Matière organique dissoute dans l’estuaire maritime du St-Laurent — comparaison et choix des méthodes. (French) [Organic matter dissolved in the maritime estuary of the St. Lawrence — comparison and choice of methods]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):133–139, February 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-023.

Jangaard:1974:NCE

[1268] P. M. Jangaard, L. W. Regier, F. G. Claggett, B. E. March, and J. Biely. Nutrient composition of experimentally produced meals from whole Argentine, capelin, sand lance, and from flounder and redfish filleting scrap. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):141–146, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-024.

Burt:1974:BBD

[1269] M. D. B. Burt and I. M. Sandeman. The biology of Bothrimonus (= Diplocotyle) (Pseudophyllidea: Cestoda): Detailed morphology and fine structure. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):147–153, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-025.

Flannagan:1974:ITN

[1270] John F. Flannagan. Influence of trisodium nitrilotriacetate on the mortality, growth, and fecundity of the freshwater snail (Helisoma trivolvis) through four generations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):155–161, February 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-026.

Dechtiar:1974:CBV

[1271] Alex O. Dechtiar. Cleidodiscus baldwini vi. sp. (Monogenoidea: Ancyrocephalinae) from the gills of the trout-perch (Percopsis omiscomaycus) of Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2): 163–165, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-027.

Burke:1974:PPB

[1272] Michael V. Burke and K. H. Mann. Productivity and production: Biomass ratios of bivalve and gastropod populations in an Eastern Canadian estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):167–177, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-028.

Wedemeyer:1974:SFT

[1273] Gary Wedemeyer and W. T. Yasutake. Stress of formalin treatment in juvenile spring Chinook salmon (Oncorhynchus tshawytscha) and steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):179–184, February 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-029.

Bernard:1974:ABG

[1274] F. R. Bernard. Annual biodeposition and Gross energy budget of mature Pacific oysters, Crassostrea gigas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):185–190, February 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-030.

Reinert:1974:RDL

[1275] Robert E. Reinert and Harold L. Bergman. Residues of DDT in lake trout (Salvelinus namaycush) and Coho salmon (Oncorhynchus kisutch) from the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2): 191–199, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-031.

Regier:1974:CNC

[1276] L. W. Regier, P. M. Jangaard, H. E. Power, B. E. March, and J. Biely. Composition and nutritive characteristics of Atlantic Canadian white fish meals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):201–204, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-032.

Walder:1974:HPE

[1277] Gordon L. Walder and Hisao P. Arai. The helminth parasites of embiotocid fishes. III. A new species of the genus Cucullanellus Törnquist, 1931 (Nematoda: Cucullanidae) from the shiner perch, Cymatogaster aggregata Gibbons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2): 205–209, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-033.

Cameron:1974:ELP

[1278] James N. Cameron. Evidence for the lack of by-pass shunting in teleost gills. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):211–213, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-034.

Anderson:1974:CPR

[1279] Douglas P. Anderson and John R. Nelson. Comparison of protection in rainbow trout (Salmo gairdneri) inoculated with and fed Hagerman Redmouth Bacterins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2): 214–216, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-035.

Ayles:1974:FES

[1280] G. B. Ayles. Fecundity and egg size of a brood stock of Salvelinus fontinalis × S. namaycush hybrids (splake). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):217–220, February 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-036.

Roberts:1974:FRM

[1281] Wayne E. Roberts. First record of the mooneye (Hiodon tergisus) in Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):220–221, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-037.

Thorne:1974:AEE

[1282] Richard E. Thorne and James J. Dawson. An acoustic estimate of the escapement of sockeye salmon (Oncorhynchus nerka) into Lake Washington in 1971. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2): 222–225, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-038.

Berst:1974:SMY

[1283] A. H. Berst and N. R. Payne. Stream migration of yearling splake (Salvelinus fontinalis × S. namaycush) planted in Georgian Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):226–229, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-039.

Hoagman:1974:FAA

[1284] Walter J. Hoagman. Feeding by alewives (Alosa pseudoharengus) on larval lake whitefish (Coregonus clupeaformis) in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):229–230, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-040.

Haedrich:1974:RSF

[1285] R. L. Haedrich and P. T. Polloni. Rarely seen fishes captured in Hudson submarine canyon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2): 231–234, February 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-041.

Varga:1974:CSL

[1286] S. Varga, W. A. Hirtle, and W. E. Anderson. A comparison of the storage life and bacteria counts of chilled cod fillets packaged in wooden boxes and heat-sealed plastic containers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):234–237, February 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-042.

Tong:1974:TML

[1287] Steven Shap Chow Tong, William D. Youngs, Walter H. Gutenmann, and Donald J. Lisk. Trace metals in Lake Cayuga lake trout (Salvelinus namaycush) in relation to age. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(2):238–239, February 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-043.

Vollenweider:1974:PBL

[1288] Richard A. Vollenweider. The production biology of the Lower Laurentian Great Lakes — a preamble. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):251–252, March 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-044.

Glooschenko:1974:PPL

[1289] Walter A. Glooschenko, James E. Moore, Mohiuddin Munawar, and R. A. Vollenweider. Primary production in Lakes Ontario and Erie: A comparative study. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (3):253–263, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-045.

Glooschenko:1974:STD

[1290] Walter A. Glooschenko, James E. Moore, and Richard A. Vollenweider. Spatial and temporal distribution of chlorophyll a and pheopigments in surface waters of Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):265–274, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-046.

Gachter:1974:SVT

[1291] R. Gächter, R. A. Vollenweider, and W. A. Glooschenko. Seasonal variations of temperature and nutrients in the surface waters of Lakes Ontario and Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):275– 290, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-047.

Burns:1974:SMS

[1292] N. M. Burns and A. E. Pashley. In situ measurement of the settling velocity profile of particulate organic carbon in Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):291–297, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-048.

Dutka:1974:MEO

[1293] B. J. Dutka, J. B. Bell, and D. L. S. Liu. Microbiological examination of offshore Lake Erie sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):299–308, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-049.

Watson:1974:SAC

[1294] N. H. F. Watson and G. F. Carpenter. Seasonal abundance of crustacean zooplankton and net plankton biomass of Lakes Huron, Erie, and Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):309– 317, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-050.

Carpenter:1974:ALH

[1295] G. F. Carpenter, E. L. Mansey, and N. H. F. Watson. Abundance and life history of Mysis relicta in the St. Lawrence Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):319–325, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-051.

Carpenter:1974:PMS

[1296] George F. Carpenter. A proposed method suitable for large-scale surveys of biomass in lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (3):327–328, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-052.

Weiler:1974:ECD

[1297] R. R. Weiler. Exchange of carbon dioxide between the atmosphere and Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3): 329–332, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-053.

Ackman:1974:LCM

[1298] R. G. Ackman and S. N. Hooper. Long-chain monoethylenic and other fatty acids in heart, liver, and blubber lipids of two harbor seals (Phoca vitulina) and one grey seal (Halichoerus grypus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):333–341, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-054.

Clayton:1974:GDA

[1299] J. W. Clayton, R. E. K. Harris, and D. N. Tretiak. Geographical distribution of alleles for supernatant malate dehydrogenase in walleye (Stizostedion vitreum vitreum) populations from Western Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):342–345, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-055.

Kendall:1974:RTM

[1300] M. W. Kendall and D. E. Hinton. Renal tubular morphology in the channel catfish (Ictalurus punctatus) kidney. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):346–347, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-056.

Odense:1974:MAS

[1301] P. H. Odense and V. H. Logan. Marking Atlantic salmon (Salmo salar) with oxytetracycline. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3): 348–350, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-057.

Coutant:1974:ECS

[1302] C. C. Coutant, H. M. Ducharme, Jr., and J. R. Fisher. Effects of cold shock on vulnerability of juvenile channel catfish (Ictalurus punctatus) and largemouth bass (Micropterus salmoides) to predation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):351–354, March 1974. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f74-058.

Moore:1974:FSS

[1303] I. A. Moore and J. W. Moore. Food of shorthorn sculpin, Myoxocephalus scorpius, in the Cumberland Sound area of Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):355–359, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-059.

Hannesson:1974:RBR

[1304] Rögnvaldur Hannesson. Relation between reproductive potential and sustained yield of fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):359–362, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-060.

Westrheim:1974:LWR

[1305] S. J. Westrheim and V. A. Snytko. Length–weight relations of Pacific Ocean perch (Sebastes alutus) in the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):363–366, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-061.

Benville:1974:SAM

[1306] Pete E. Benville, Jr. and Sid Korn. A simple apparatus for metering volatile liquids into water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):367–368, March 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-062.

Freeman:1974:MSC

[1307] H. C. Freeman, D. A. Horne, B. McTague, and M. McMenemy. Mercury in some Canadian Atlantic Coast fish and shellfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):369–372, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-063.

Baskett:1974:MMB

[1308] Russell C. Baskett and William J. Lulves. A method of measuring bacterial growth in aquatic environments using dialysis culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(3):372–374, March 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-064.

Castell:1974:SMT

[1309] C. H. Castell, Barbara Smith, and W. J. Dyer. Simultaneous measurements of trimethylamine and dimethylamine in fish, and their use for estimating quality of frozen-stored gadoid fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):383–389, April 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-065.

Hobson:1974:EII

[1310] Louis A. Hobson. Effects of interactions of irradiance, daylength, and temperature on division rates of three species of marine unicellular algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):391–395, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-066.

Yamamoto:1974:IPN

[1311] T. Yamamoto. Infectious pancreatic necrosis virus occurrence at a hatchery in Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4): 397–402, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-067.

Watanabe:1974:LFA

[1312] T. Watanabe and R. G. Ackman. Lipids and fatty acids of the American (Crassostrea virginica) and European flat (Ostrea edulis) oysters from a common habitat, and after one feeding with Dicrateria inornata or Isochrysis galbana. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4): 403–409, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-068.

Steigenberger:1974:FAR

[1313] Lance W. Steigenberger and P. A. Larkin. Feeding activity and rates of digestion of northern squawfish (Ptychocheilus oregonensis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):411–420, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-069.

Carter:1974:LCT

[1314] John C. H. Carter. Life cycles of three limnetic copepods in a beaver pond. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):421–434, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-070.

Armstrong:1974:MAD

[1315] Robert H. Armstrong. Migration of anadromous Dolly Varden (Salvelinus malma) in Southeastern Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):435–444, April 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-071.

Bauer:1974:MAA

[1316] Beverly A. Bauer and R. R. Eitenmiller. Muscle arylamidase activity of several marine species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):445–449, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-072.

McKim:1974:DTT

[1317] James M. McKim and Duane A. Benoit. Duration of toxicity tests for establishing “no effect” concentrations for copper with brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4): 449–452, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-073.

Menzel:1974:PJO

[1318] R. W. Menzel. Portuguese and Japanese oysters are the same species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):453–456, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-074.

Longard:1974:CSC

[1319] A. A. Longard and L. W. Regier. Color and some composition changes in ocean perch (Sebastes marinus) held in refrigerated sea water with and without carbon dioxide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):456–460, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-075.

Brinkhurst:1974:FMI

[1320] R. O. Brinkhurst. Factors mediating interspecific aggregation of tubificid oligochaetes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4): 460–462, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-076.

Pagenkopf:1974:ECT

[1321] Gordon K. Pagenkopf, Rosemarie C. Russo, and Robert V. Thurston. Effect of complexation on toxicity of copper to fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):462–465, April 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-077.

Armstrong:1974:EIC

[1322] David A. Armstrong and Raymond E. Millemann. Effects of the insecticide carbaryl on clams and some other intertidal mud flat animals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):466–470, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-078.

Khan:1974:IMF

[1323] R. A. Khan and T. K. Pitt. An infestation of the marine fish Lycodes lavalaei by the parasitic copepod Tanypleurus alcicornis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):470–471, April 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-079.

Peppar:1974:NEK

[1324] John L. Peppar. Northern extension of known range of Atlantic menhaden (Brevoortia tyrannus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):471–472, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-080.

Nowlan:1974:EMG

[1325] Sandra S. Nowlan and W. J. Dyer. Effect of mincing on glycolytic activity in prerigor Atlantic cod (Gadus morhua) muscle stored in ice or frozen. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):473–476, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-081.

Groninger:1974:PUM

[1326] Herman S. Groninger, Jr. and Ruth Miller. Preparation and use of a modified protein Binder in the fabrication of feed for post-larval aquacultural animals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4): 477–479, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-082.

Meehan:1974:TVF

[1327] William R. Meehan, Logan A. Norris, and Howard S. Sears. Toxicity of various formulations of 2,4-D to salmonids in Southeast Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):480–485, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-083.

Biesinger:1974:CTN

[1328] K. E. Biesinger, R. W. Andrew, and J. W. Arthur. Chronic toxicity of NTA (nitrilotriacetate) and metal-NTA complexes to Daphnia magna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(4):486–490, April 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-084.

Anonymous:1974:EFa

[1329] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):499–500, May 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-085.

Harvey:1974:ILC

[1330] Harold H. Harvey and Glen H. Geen. Introduction to limnology in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):501– 504, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-086.

Bruce:1974:LVW

[1331] J. P. Bruce. Lakes — the vital water resource in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):505–512, May 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-087.

Vallentyne:1974:LEN

[1332] J. R. Vallentyne. Limnology and education in the next decade. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):513–519, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-088.

Elson:1974:IRE

[1333] Paul F. Elson. Impact of recent economic growth and industrial development on the ecology of Northwest Miramichi Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):521– 544, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-089.

Hynes:1974:BAO

[1334] H. B. N. Hynes, N. K. Kaushik, M. A. Lock, D. L. Lush, Z. S. J. Stocker, R. R. Wallace, and D. D. Williams. Benthos and allochthonous organic matter in streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5): 545–553, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-090.

Kerekes:1974:LCF

[1335] Joseph J. Kerekes. Limnological conditions in five small oligotrophic lakes in Terra Nova National Park, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):555–583, May 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-091.

Schindler:1974:PCL

[1336] D. W. Schindler, H. E. Welch, J. Kalff, G. J. Brunskill, and N. Kritsch. Physical and chemical limnology of Char Lake, Cornwallis Island (75^{o}N lat.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):585–607, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-092.

Welch:1974:BPR

[1337] H. E. Welch and J. Kalff. Benthic photosynthesis and respiration in Char Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):609– 620, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-093.

Kalff:1974:PPC

[1338] J. Kalff and H. E. Welch. Phytoplankton production in Char Lake, a natural polar lake, and in Meretta Lake, a polluted polar lake, Cornwallis Island, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):621–636, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-094.

Rigler:1974:PZC

[1339] F. H. Rigler, M. E. MacCallum, and J. C. Roff. Production of zooplankton in Char Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5): 637–646, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-095.

Schindler:1974:EHA

[1340] D. W. Schindler, J. Kalff, H. E. Welch, G. J. Brunskill, H. Kling, and N. Kritsch. Eutrophication in the High Arctic — Meretta Lake, Cornwallis Island (75^{o}N lat.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5): 647–662, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-096.

Ryder:1974:MIF

[1341] R. A. Ryder, S. R. Kerr, K. H. Loftus, and H. A. Regier. The morphoedaphic index, a fish yield estimator — review and evaluation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):663–688, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-097.

Boyce:1974:SAG

[1342] F. M. Boyce. Some aspects of Great Lakes physics of importance to biological and chemical processes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):689–730, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-098.

Dobson:1974:SCN

[1343] Hugh F. H. Dobson, Michael Gilbertson, and Peter G. Sly. A summary and comparison of nutrients and related water quality in Lakes Erie, Ontario, Huron, and Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):731–738, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-099.

Vollenweider:1974:CRP

[1344] R. A. Vollenweider, M. Munawar, and P. Stadelmann. A comparative review of phytoplankton and primary production in the Laurentian Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):739– 762, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-100.

Cook:1974:BMS

[1345] David G. Cook and Murray G. Johnson. Benthic macroinvertebrates of the St. Lawrence Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):763–782, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-101.

Watson:1974:ZSL

[1346] N. H. F. Watson. Zooplankton of the St. Lawrence Great Lakes — species composition, distribution, and abundance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):783–794, May 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-102.

Sly:1974:RGR

[1347] P. G. Sly and R. L. Thomas. Review of geological research as it relates to an understanding of Great Lakes limnology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):795–825, May 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-103.

Christie:1974:CFS

[1348] W. J. Christie. Changes in the fish species composition of the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):827– 854, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-104.

Anderson:1974:CPC

[1349] R. Stewart Anderson. Crustacean plankton communities of 340 lakes and ponds in and near the National Parks of the Canadian Rocky Mountains. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):855–869, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-105.

Goodlad:1974:FAS

[1350] J. C. Goodlad, T. W. Gjernes, and E. L. Brannon. Factors affecting sockeye salmon (Oncorhynchus nerka) growth in four lakes of the Fraser River system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5): 871–892, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-106.

Hall:1974:MLI

[1351] Ken J. Hall and Kim D. Hyatt. Marion Lake (IBP) — from bacteria to fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):893– 911, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-107.

Geen:1974:EHD

[1352] Glen H. Geen. Effects of hydroelectric development in Western Canada on aquatic ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):913–927, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-108.

Lawler:1974:TPP

[1353] G. H. Lawler, L. A. Sunde, and J. Whitaker. Trout production in prairie ponds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5): 929–936, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-109.

Schindler:1974:ELA

[1354] D. W. Schindler and E. J. Fee. Experimental Lakes Area: Whole-lake experiments in eutrophication. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):937–953, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-110.

Stockner:1974:RLS

[1355] J. G. Stockner and T. G. Northcote. Recent limnological studies of Okanagan Basin lakes and their contribution to comprehensive water resource planning. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5): 955–976, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-111.

Ward:1974:RRL

[1356] F. J. Ward and G. G. C. Robinson. A review of research on the limnology of West Blue Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(5):977–1005, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-112.

Paetz:1974:SNA

[1357] M. J. Paetz and K. A. Zelt. Studies of Northern Alberta lakes and their fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (5):1007–1020, May 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-113.

Anonymous:1974:EFb

[1358] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1023–1024, June 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-114.

Spinelli:1974:CPU

[1359] John Spinelli, Lawrence Lehman, and Dave Wieg. Composition, processing, and utilization of red crab (Pleuroncodes planipes) as an aquacultural feed ingredient. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (6):1025–1029, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-115.

Hare:1974:APP

[1360] G. M. Hare and C. Frantsi. Abundance and potential pathology of parasites infecting salmonids in Canadian Maritime hatcheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1031–1036, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-116.

Payne:1974:TVN

[1361] R. H. Payne. Transferrin variation in North American populations of the Atlantic salmon, Salmo salar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1037–1041, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-117.

McLeay:1974:GSB

[1362] D. J. McLeay and D. A. Brown. Growth stimulation and biochemical changes in juvenile Coho salmon (Oncorhynchus kisutch) exposed to bleached kraft pulpmill effluent for 200 days. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1043–1049, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-118.

Paterson:1974:VPH

[1363] W. D. Paterson and James E. Stewart. In vitro phagocytosis by hemocytes of the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1051–1056, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-119.

Jeffries:1974:SDB

[1364] H. Perry Jeffries and William C. Johnson. Seasonal distributions of bottom fishes in the Narragansett Bay area: Seven-year variations in the abundance of winter flounder (Pseudopleuronectes americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1057–1066, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-120.

Nielsen:1974:DNS

[1365] Jørgen G. Nielsen and Rodney J. Eagle. Descriptions of a new species of Barathronus (Pisces, Aphyonidae) and four specimens of Sciadonus sp. from the Eastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1067–1072, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-121.

Templeman:1974:MIA

[1366] Wilfred Templeman. Migrations and intermingling of Atlantic cod (Gadus morhua) stocks of the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1073–1092, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-122.

Malevski:1974:LFP

[1367] Y. Malevski, M. W. Montgomery, and R. O. Sinnhuber. Liver fat and protein metabolism in rainbow trout (Salmo gairdneri) fed cyclopropenoid fatty acids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1093– 1100, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-123.

Kelso:1974:DEC

[1368] J. R. M. Kelso, E. E. Pickett, and R. G. Dowd. A digital echo-counting system used in determining abundance of freshwater pelagic fish in relation to depth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (6):1101–1104, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-124.

Duthie:1974:PCP

[1369] H. C. Duthie and M. L. Ostrofsky. Plankton, chemistry, and physics of lakes in the Churchill Falls region of Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1105–1117, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-125.

Allen:1974:PAS

[1370] Robin L. Allen and Pereti Basasibwaki. Properties of age structure models for fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1119–1125, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-126.

McLeese:1974:ORF

[1371] D. W. McLeese. Olfactory response and fenitrothion toxicity in American lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1127–1131, June 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-127.

Zeitoun:1974:EWS

[1372] Ibrahim H. Zeitoun, Duane E. Ullrey, and Peter I. Tack. Effects of water salinity and dietary protein levels on total serum protein and hematocrit of rainbow trout (Salmo gairdneri) fingerlings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1133–1134, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-128.

Hanek:1974:PFA

[1373] George Hanek and Kalman Molnar. Parasites of freshwater and anadromous fishes from Matamek River System, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1135–1139, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-129.

Gee:1974:ABE

[1374] John H. Gee, Kazimierz Machniak, and Sandra M. Chalanchuk. Adjustment of buoyancy and excess internal pressure of swimbladder gases in some North American freshwater fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1139–1141, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-130.

Childs:1974:FFM

[1375] E. A. Childs. Functionality of fish muscle: Emulsification capacity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1142–1144, June 1974. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f74-131.

Zeitoun:1974:ISP

[1376] Ibrahim H. Zeitoun, Duane E. Ullrey, John E. Halver, Peter I. Tack, and W. T. Magee. Influence of salinity on protein requirements of Coho salmon (Oncorhynchus kisutch) smolts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1145–1148, June 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-132.

Stasko:1974:SDA

[1377] A. B. Stasko and S. A. Rommel, Jr. Swimming depth of adult American eels (Anguilla rostrata) in a saltwater bay as determined by ultrasonic tracking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (6):1148–1150, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-133.

Harbo:1974:FCT

[1378] R. M. Harbo, F. T. McComas, and J. A. J. Thompson. Fluoride concentrations in two Pacific Coast inlets — an indication of industrial contamination. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1151– 1154, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-134.

Coad:1974:MVT

[1379] Brian W. Coad and G. Power. Meristic variation in the threespine stickleback, Gasterosteus aculeatus, in the Matamek River System, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1155–1157, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-135.

Mathews:1974:NMR

[1380] Stephen B. Mathews and Raymond Buckley. Natural mortality rate in last winter of life of Coho salmon (Oncorhynchus kisutch) resident in Puget Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (6):1158–1160, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-136.

Myers:1974:SVM

[1381] G. L. Myers and J. J. Peterka. A syringe volumetric measuring device. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(6):1160–1161, June 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-137.

Ratkowsky:1974:NSC

[1382] D. A. Ratkowsky, S. J. Thrower, I. J. Eustace, and June Olley. A numerical study of the concentration of some heavy metals in Tasmanian oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1165–1171, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-138.

Nowlan:1974:MRG

[1383] Sandra S. Nowlan and W. J. Dyer. Maximum rates of glycolysis and breakdown of high-energy phosphorus compounds in prerigor cod muscle at specific freezing temperatures between −1 and −4 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1173–1179, July 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-139.

Hiltz:1974:AND

[1384] Doris Fraser Hiltz, Lucy J. Bishop, and W. J. Dyer. Accelerated nucleotide degradation and glycolysis during warming to and subsequent storage at −5 C of prerigor, quick-frozen adductor muscle of the sea scallop (Placopecten magellanicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1181–1187, July 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-140.

Petty:1974:LAB

[1385] Lorna L. Petty and John J. Magnuson. Lymphocystis in age 0 bluegills (Lepomis macrochirus) relative to heated effluent in Lake Monona, Wisconsin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1189– 1193, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-141.

Medcof:1974:SIF

[1386] J. C. Medcof and M. L. H. Thomas. Surfacing on ice of frozen-in marine bottom materials. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (7):1195–1200, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-142.

Slaney:1974:EPA

[1387] P. A. Slaney and T. G. Northcote. Effects of prey abundance on density and territorial behavior of young rainbow trout (Salmo gairdneri) in laboratory stream channels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1201–1209, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-143.

Donaldson:1974:EAS

[1388] Edward M. Donaldson and J. R. McBride. Effect of ACTH and salmon gonadotropin on interrenal and thyroid activity of gonadectomized adult sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1211–1214, July 1974. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f74-144.

Stadelmann:1974:PPR

[1389] P. Stadelmann, J. E. Moore, and E. Pickett. Primary production in relation to temperature structure, biomass concentration, and light conditions at an inshore and offshore station in Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1215–1232, July 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-145.

Anonymous:1974:Ob

[1390] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1233–1234, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-146.

Gaskin:1974:MDD

[1391] D. E. Gaskin, G. J. D. Smith, P. W. Arnold, M. V. Louisy, R. Frank, M. Holdrinet, and J. W. McWade. Mercury, DDT, Dieldrin, and PCB in two species of Odontoceti (Cetacea) from St. Lucia, Lesser Antilles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1235–1239, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-147.

Spieler:1974:STS

[1392] Richard E. Spieler. Short-term serum cortisol concentrations in goldfish (Carassius auratus) subjected to serial sampling and restraint. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1240–1242, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-148.

Burton:1974:SRG

[1393] William Burton. A semiautomatic release gear for grabs and corers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1244–1246, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-149.

Shepherd:1974:ALC

[1394] B. G. Shepherd. Activity localization in coastal cutthroat trout (Salmo clarki clarki) in a small bog lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1246–1249, July 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-150.

McComish:1974:EBL

[1395] Thomas S. McComish, Richard O. Anderson, and F. Glenn Goff. Estimation of bluegill (Lepomis macrochirus) proximate composition with regression models. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31 (7):1250–1254, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-151.

McCleave:1974:ERS

[1396] James D. McCleave, Kristin A. Stred, and Enoch H. Albert. Effect of reinforcement schedule and unconditioned stimulus on classical cardiac conditioning in American eels (Anguilla rostrata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1254–1258, July 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-152.

Paterson:1974:BSD

[1397] W. D. Paterson. Biochemical and serological differentiation of several pigment-producing aeromonads. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1259–1261, July 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-153.

Paulson:1974:OPS

[1398] Alan C. Paulson and Ronald L. Smith. Occurrence of Pacific staghorn sculpin (Leptocottus armatus) in the Southern Bering Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1262, July 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-154.

Spence:1974:SVC

[1399] C. Spence and D. M. Steven. Seasonal variation of the chlorophyll a : pheopigment ratio in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1263–1268, July 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-155.

Anonymous:1974:P

[1400] Anonymous. Perspectives. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(7):1269–1322, July 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-156.

Antia:1974:PES

[1401] Naval J. Antia and Arthur F. Landymore. Physiological and ecological significance of the chemical instability of uric acid and related purines in sea water and marine algal culture medium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1327–1335, August 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-157.

Maclellan:1974:DCS

[1402] Delphine C. Maclellan and Chang tai Shih. Descriptions of copepodite stages of Chiridius gracilis Farran 1908 (Crustacea: Copepoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1337–1349, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-158.

Ihssen:1974:GDR

[1403] P. Ihssen and J. S. Tait. Genetic differences in retention of swimbladder gas between two populations of lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1351–1354, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-159.

Qadri:1974:TSS

[1404] S. U. Qadri. Taxonomic status of the Salvelinus alpinus complex. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1355–1361, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-160.

Castell:1974:LNE

[1405] J. D. Castell and S. D. Budson. Lobster nutrition: The effect on Homarus americanus of dietary protein levels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1363–1370, August 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-161.

Amend:1974:PIH

[1406] Donald F. Amend and Lynnwood Smith. Pathophysiology of infectious hematopoietic necrosis virus disease in rainbow trout (Salmo gairdneri): Early changes in blood and aspects of the immune response after injection of IHN virus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8): 1371–1378, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-162.

Bams:1974:GIS

[1407] Robert A. Bams. Gravel incubators: a second evaluation on pink salmon, Oncorhynchus gorbuscha, including adult returns. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1379–1385, August 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-163.

Wenner:1974:FGO

[1408] Charles A. Wenner and J. A. Musick. Fecundity and gonad observations of the American eel, Anguilla rostrata, migrating from Chesapeake Bay, Virginia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1387–1391, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-164.

Carlson:1974:ERO

[1409] Anthony R. Carlson and Richard E. Siefert. Effects of reduced oxygen on the embryos and larvae of lake trout (Salvelinus namaycush) and largemouth bass (Micropterus salmoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1393–1396, August 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-165.

Malevski:1974:LDR

[1410] Y. Malevski, J. H. Wales, and M. W. Montgomery. Liver damage in rainbow trout (Salmo gairdneri) fed cyclopropenoid fatty acids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1397–1400, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-166.

Boyce:1974:MWE

[1411] F. M. Boyce. Mixing within experimental enclosures: a cautionary note on the limnocorral. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8): 1400–1405, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-167.

Proctor:1974:PPB

[1412] R. R. Proctor, K. T. Marvin, L. M. Lansford, and R. C. Benton. Phosphoglucomutase polymorphism in brown shrimp, Penaeus aztecus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1405–1407, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-168.

McCart:1974:IPA

[1413] P. McCart and H. Bain. An isolated population of Arctic char (Salvelinus alpinus) inhabiting a warm mineral spring above a waterfall at Cache Creek, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1408–1414, August 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-169.

Laurence:1974:GSH

[1414] Geoffrey C. Laurence. Growth and survival of haddock (Melanogrammus aeglefinus) larvae in relation to planktonic prey concentration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1415–1419, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-170.

McKern:1974:DST

[1415] J. L. McKern, H. F. Horton, and K. V. Koski. Development of steelhead trout (Salmo gairdneri) otoliths and their use for age analysis and for separating summer from winter races and wild from hatchery stocks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1420–1426, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-171.

OBrien:1974:CUI

[1416] W. John O’Brien and Gary L. Vinyard. Comment on the use of Ivlev’s electivity index with planktivorous fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1427–1429, August 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-172.

Rubec:1974:FRM

[1417] P. J. Rubec and B. W. Coad. First record of the margined madtom (Noturus insignis) from Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1430–1431, August 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-173.

Larkin:1974:PIA

[1418] P. A. Larkin. Play it again Sam — an essay on salmon enhancement. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8):1433–1456, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-174.

Anonymous:1974:CRPa

[1419] Anonymous. Corrections to recent publications… corrections sur récentes publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(8): 1457, August 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-175.

Paloheimo:1974:RAS

[1420] J. E. Paloheimo and P. F. Elson. Reduction of Atlantic salmon (Salmo salar) catches in Canada attributed to the Greenland fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1467–1480, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-176.

Pauley:1974:FSE

[1421] Gilbert B. Pauley. Fish sporozoa: Extraction of antigens from Myxosoma cerebralis spores which mimic tissue antigens of rainhow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1481–1484, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-177.

Richkus:1974:FIS

[1422] W. A. Richkus. Factors influencing the seasonal and daily patterns of alewife (Alosa pseudoharengus) migration in a Rhode Island River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1485–1497, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-178.

Ayles:1974:RIA

[1423] G. Burton Ayles. Relative importance of additive genetic and maternal sources of variation in early survival of young splake hybrids (Salvelinus fontinalis × S. namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1499–1502, September 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-179.

Yocom:1974:EAT

[1424] Thomas G. Yocom and Thomas A. Edsall. Effect of acclimation temperature and heat shock on vulnerability of fry of lake whitefish (Coregonus clupeaformis) to predation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1503–1506, September 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-180.

Kelso:1974:ITE

[1425] John R. M. Kelso. Influence of a thermal effluent on movement of brown bullhead (Ictalurus nebulosus) as determined by ultrasonic tracking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1507–1513, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-181.

Chau:1974:DAC

[1426] Y. K. Chau, R. Gächter, and K. Lum-Shue-Chan. Determination of the apparent complexing capacity of Lake Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1515–1519, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-182.

Parsons:1974:LDM

[1427] L. S. Parsons and J. A. Moores. Long-distance migration of an Atlantic mackerel (Scomber scombrus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1521–1522, September 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-183.

Cho:1974:PRH

[1428] C. Y. Cho, H. S. Bayley, and S. J. Slinger. Partial replacement of herring meal with soybean meal and other changes in a diet for rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9): 1523–1528, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-184.

Duval:1974:SDD

[1429] W. S. Duval, P. J. Brockington, M. S. von Melville, and G. H. Geen. Spectrophotometric determination of dissolved oxygen concentration in water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1529–1530, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-185.

Werner:1974:FSP

[1430] Earl E. Werner. The fish size, prey size, handling time relation in several sunfishes and some implications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1531–1536, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-186.

Hare:1974:FRA

[1431] G. M. Hare and H. P. Murphy. First record of the American shad (Alosa sapidissima) from Labrador waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1536–1537, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-187.

McArdle:1974:BHT

[1432] J. F. McArdle and R. J. Roberts. Bilateral hyperplasia of the thymus in a rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1537–1539, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-188.

Francis:1974:RFM

[1433] Robert C. Francis. Relationship of fishing mortality to natural mortality at the level of maximum sustainable yield under the logistic stock production model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9): 1539–1542, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-189.

Halter:1974:ATP

[1434] Mark T. Halter and Howard E. Johnson. Acute toxicities of a polychlorinated biphenyl (PCB) and DDT alone and in combination to early life stages of Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1543–1547, September 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-190.

Hara:1974:MEO

[1435] Toshiaki J. Hara. Is morpholine an effective olfactory stimulant in fish? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1547–1550, September 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-191.

Parker:1974:EFR

[1436] Richard A. Parker. Empirical functions relating metabolic processes in aquatic systems to environmental variables. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1550–1552, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-192.

McCoy:1974:PBE

[1437] R. H. McCoy and K. S. Pilcher. Peptone beef extract glycogen agar, a selective and differential Aeromonas medium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1553–1555, September 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-193.

McLeese:1974:TPA

[1438] D. W. McLeese. Toxicity of phosphamidon to American lobsters (Homarus americanus) held at 4 and 12 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1556–1558, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-194.

Strasdine:1974:CFM

[1439] G. A. Strasdine and Lyn Dubetz. Coliform flora of migrating sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1559–1560, September 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-195.

Anonymous:1974:BRC

[1440] Anonymous. Book Reviews/ comptes rendus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(9):1561, September 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-196.

Sloan:1974:BHL

[1441] J. P. Sloan, J. A. J. Thompson, and P. A. Larkin. The biological half-life of inorganic mercury in the Dungeness crab (Cancer magister). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1571–1576, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-197.

Lee:1974:LZB

[1442] Richard F. Lee. Lipids of zooplankton from Bute Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1577–1582, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-198.

Kabata:1974:MMF

[1443] Z. Kabata. Mouth and mode of feeding of Caligidae (Copepoda), parasites of fishes, as determined by light and scanning electron microscopy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1583–1588, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-199.

Kabata:1974:ASJ

[1444] Z. Kabata and C. R. Forrester. Atheresthes stomias (Jordan and Gilbert 1880) (Pisces: Pleuronectiformes) and its eye parasite Phrixocephalus cincinnatus Wilson 1908 (Copepoda: Lernaeoceridae) in Canadian Pacific waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1589– 1595, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-200.

Markiw:1974:MCC

[1445] Maria E. Markiw and Ken Wolf. Myxosoma cerebralis: Comparative sensitivity of spore detection methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1597–1600, October 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-201.

Braekevelt:1974:FSR

[1446] C. R. Braekevelt. Fine structure of the retinal pigment epithelium, Bruch’s membrane, and choriocapillaris in the northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1601–1605, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-202.

Dodson:1974:ROV

[1447] Julian J. Dodson and William C. Leggett. Role of olfaction and vision in the behavior of American shad (Alosa sapidissima) homing to the Connecticut River from long Island Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1607–1619, October 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-203.

Olla:1974:FFA

[1448] Bori L. Olla and Carol Samet. Fish-to-fish attraction and the facilitation of feeding behavior as mediated by visual stimuli in striped mullet, Mugil cephalus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1621–1630, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-204.

Bidgood:1974:RPT

[1449] B. F. Bidgood. Reproductive potential of two lake whitefish (Coregonus clupeaformis) populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1631–1639, October 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-205.

Jones:1974:ESP

[1450] D. R. Jones, J. W. Kiceniuk, and O. S. Bamford. Evaluation of the swimming performance of several fish species from the Mackenzie River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1641–1647, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-206.

Reinert:1974:ETA

[1451] Robert E. Reinert, Linda J. Stone, and Wayne A. Willford. Effect of temperature on accumulation of methylmercuric chloride and p,p’ DDT by rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1649–1652, October 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-207.

Russo:1974:ATN

[1452] Rosemarie C. Russo, Charlie E. Smith, and Robert V. Thurston. Acute toxicity of nitrite to rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1653–1655, October 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-208.

Anonymous:1974:Oc

[1453] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1657–1659, October 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-209.

Amundrud:1974:SSF

[1454] James R. Amundrud, Daniel J. Faber, and Allen Keast. Seasonal succession of free-swimming Perciform larvae in Lake Opinicon, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1661–1665, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-210.

Calder:1974:MDH

[1455] Dale R. Calder. The mud-dwelling hydrozoan Boreohydra simplex in the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1666–1667, October 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-211.

Goodwin:1974:IML

[1456] M. H. Goodwin and C. I. Goddard. An inexpensive multiple level water sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1667–1668, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-212.

Jensen:1974:PPC

[1457] A. L. Jensen. Predator–prey and competition models with state variables: Biomass, number of individuals, and average individual weight. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1669–1674, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f74-213.

Tenore:1974:FEP

[1458] Kenneth R. Tenore and U. K. Gopalan. Feeding efficiencies of the polychaete Nereis virens cultured on hard-clam tissue and oyster detritus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1675–1678, October 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-214.

Pessah:1974:ECT

[1459] E. Pessah and P. M. Powles. Effect of constant temperature on growth rates of pumpkinseed sunfish (Lepomis gibbosus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1678–1682, October 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-215.

Regier:1974:PTA

[1460] Henry A. Regier, Patrick L. Bishop, and David J. Rapport. Planned transdisciplinary approaches: Renewable resources and the natural environment, particularly fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1683–1703, October 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-216.

MacKenzie:1974:CAS

[1461] W. C. MacKenzie. Conceptual aspects of strategic planning for fishery management and development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(10):1705–1712, October 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-217.

Scott:1974:MCW

[1462] D. P. Scott. Mercury concentration of white muscle in relation to age, growth, and condition in four species of fishes from Clay Lake, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1723–1729, November 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-218.

Otto:1974:SSY

[1463] Robert G. Otto and John O’Hara Rice. Swimming speeds of yellow perch (Perca flavescens) following an abrupt change in environmental temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1731–1734, November 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-219.

Boyce:1974:BES

[1464] Norbert P. J. Boyce. Biology of Eubothrium salvelini (Cestoda: Pseudophyllidea), a parasite of juvenile sockeye salmon (Oncorhynchus nerka) of Babine Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1735–1742, November 1974. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f74-220.

Paterson:1974:ETA

[1465] W. D. Paterson and J. L. Fryer. Effect of temperature and antigen dose on the antibody response of juvenile Coho salmon (Oncorhynchus kisutch) to Aeromonas salmonicida endotoxin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1743–1749, November 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-221.

Paterson:1974:IRJ

[1466] W. D. Paterson and J. L. Fryer. Immune response of juvenile Coho salmon (Oncorhynchus kisutch) to Aeromonas salmonicida cells administered intraperitoneally in Freund’s complete adjuvant. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1751–1755, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-222.

Peterson:1974:IFS

[1467] R. H. Peterson. Influence of fenitrothion on swimming velocities of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1757–1762, November 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-223.

Beamish:1974:ASD

[1468] F. W. H. Beamish. Apparent specific dynamic action of largemouth bass, Micropterus salmoides. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1763–1769, November 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-224.

Dillon:1974:TSN

[1469] P. J. Dillon and F. H. Rigler. A test of a simple nutrient budget model predicting the phosphorus concentration in lake water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1771–1778, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-225.

Matheke:1974:PPB

[1470] G. E. M. Matheke and Rita Horner. Primary productivity of the benthic microalgae in the Chukchi Sea near Barrow, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1779–1786, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-226.

Godin:1974:ETH

[1471] J.-G. Godin, P. A. Dill, and D. E. Drury. Effects of thyroid hormones on behavior of yearling Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1787–1790, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-227.

McCarthy:1974:OHA

[1472] D. H. McCarthy. Occurrence of hematozoa in Atlantic salmon (Salmo salar) smolts and adults in an English River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1790–1792, November 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-228.

Fromm:1974:CTG

[1473] Paul O. Fromm. Circulation in trout gills: Presence of “blebs” in afferent filamental vessels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11): 1793–1796, November 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-229.

Satia:1974:COF

[1474] Benedict P. Satia, Lauren R. Donaldson, Lynwood S. Smith, and John N. Nightingale. Composition of ovarian fluid and eggs of the University of Washington strain of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1796–1799, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-230.

Pitt:1974:ACG

[1475] T. K. Pitt. Age composition and growth of yellowtail flounder (Limanda ferruginea) from the Grand Bank. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1800–1802, November 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-231.

Ackman:1974:CLF

[1476] R. G. Ackman, S. Epstein, and M. Kelleher. A comparison of lipids and fatty acids of the ocean quahaug, Arctica islandica, from Nova Scotia and New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11): 1803–1811, November 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-232.

Ackman:1974:SDF

[1477] K. G. Ackman, B. A. Linke, and J. Hingley. Some details of fatty acids and alcohols in the lipids of North Atlantic copepods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1812–1818, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-233.

Collins:1974:PFK

[1478] J. J. Collins and A. O. Dechtiar. Parasite fauna of Kokanee salmon (Oncorhynchus nerka) introduced into Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1818–1821, November 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-234.

Ash:1974:FKD

[1479] G. R. Ash, N. R. Chymko, and D. N. Gallup. Fish kill due to “cold shock” in Lake Wabamun, Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1822–1824, November 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-235.

Poston:1974:EFB

[1480] Hugh A. Poston. Effect of feeding brown trout (Salmo trutta) a diet pelleted in dry and moist forms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1824–1826, November 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-236.

Mundie:1974:OSN

[1481] J. H. Mundie. Optimization of the salmonid nursery stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1827–1837, November 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-237.

Saunders:1974:FFI

[1482] Richard L. Saunders. Fish farming international. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(11):1839, November 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-238.

Anonymous:1974:AA

[1483] Anonymous. Announcement/Avis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1849, December 1974. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f74-239.

Yurkowski:1974:IAR

[1484] M. Yurkowski and Jo-Anne L. Tabachek. Identification, analysis, and removal of geosmin from muddy-flavored trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1851–1858, December 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-240.

Kerr:1974:TSD

[1485] S. R. Kerr. Theory of size distribution in ecological communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1859–1862, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-241.

Dechtiar:1974:RAA

[1486] Alex O. Dechtiar and William A. Dillon. Redescription of Anonchohaptor anomalum Mueller 1938 and a description of Icelanonchohaptor fyviei n. sp. (Monogenoidea: Dactylogyridae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1863–1866, December 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-242.

Bull:1974:BJC

[1487] Christopher J. Bull and John E. McInerney. Behavior of juvenile Coho salmon (Oncorhynchus kisutch) exposed to Sumithion (Fenitrothion), an organophosphate insecticide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1867–1872, December 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-243.

Stewart:1974:EVA

[1488] James E. Stewart and B. Arie. Effectiveness of vancomycin against gaffkemia, the bacterial disease of lobsters (genus Homarus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1873–1879, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-244.

Armstrong:1974:PDW

[1489] F. A. J. Armstrong and D. P. Scott. Photochemical dechlorination of water supply for fish tanks, with commercial water sterilizers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1881–1885, December 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-245.

Stewart:1974:CVV

[1490] James E. Stewart and B. M. Zwicker. Comparison of various vaccines for inducing resistance in the lobster Homarus americanus to the bacterial infection, gaffkemia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12): 1887–1892, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-246.

Smith:1974:BPP

[1491] Charlie E. Smith, Myron Brin, and John E. Halver. Biochemical, physiological, and pathological changes in pyridoxine-deficient rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12): 1893–1898, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-247.

Alderdice:1974:EDD

[1492] D. F. Alderdice and C. R. Forrester. Early development and distribution of the flathead sole (Hippoglossoides elassodon). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1899–1918, December 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-248.

Hobson:1974:OSD

[1493] Louis A. Hobson and Dawn E. Ketcham. Observations on subsurface distributions of chlorophyll a and phytoplankton carbon in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12): 1919–1925, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-249.

Brown:1974:TDC

[1494] Daniel J. Brown and Loh Lee Low. Three-dimensional computer graphics in fisheries science. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12): 1927–1935, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-250.

Dill:1974:RBD

[1495] P. A. Dill and R. C. Saunders. Retarded behavioral development and impaired balance in Atlantic salmon (Salmo salar) alevins hatched from gastrulae exposed to DDT. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1936–1938, December 1974. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f74-251.

Chen:1974:SDC

[1496] Peter K. Chen, G. L. Bullock, H. M. Stuckey, and A. C. Bullock. Serological diagnosis of corynebacterial kidney disease of salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1939–1940, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-252.

Cobb:1974:SCI

[1497] J. Stanley Cobb and George R. Tamm. Social conditions increase intermolt period in juvenile lobsters Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1941–1943, December 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-253.

Porter:1974:HLW

[1498] T. R. Porter and S. Corey. A hermaphroditic lake whitefish, Coregonus clupeaformis, from Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1944–1945, December 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-254.

Bryan:1974:HCG

[1499] J. E. Bryan. Hunger and the capture of grayling and char. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1945–1948, December 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-255.

McLeese:1974:TCT

[1500] D. W. McLeese. Toxicity of copper at two temperatures and three salinities to the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1949–1952, December 1974. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f74-256.

Newcomb:1974:CBC

[1501] T. W. Newcomb. Changes in blood chemistry of juvenile steelhead trout, Salmo gairdneri, following sublethal exposure to nitrogen supersaturation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1953–1957, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-257.

Johnson:1974:IFA

[1502] G. R. Johnson, G. Wobeser, and B. T. Rouse. Indirect fluorescent antibody technique for detection of RM bacterium of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1957–1959, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-258.

Meith-Avcin:1974:DRR

[1503] N. Meith-Avčin. DDT and the rugophilic response of settling barnacles Balanus improvisus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12): 1960–1963, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f74-259.

Thompson:1974:ICF

[1504] Peter C. Thompson. Institutional constraints in fisheries management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1965–1981, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f74-260.

Mann:1974:SME

[1505] Kennethe H. Mann. The structure of marine ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12):1983, December 1974. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f74-261.

Anonymous:1974:CRPb

[1506] Anonymous. Corrections to recent publications… corrections sur récentes publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 31(12): 1984, December 1974. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f74-262.

Hodson:1975:TIC

[1507] Peter V. Hodson and John B. Sprague. Temperature-induced changes in acute toxicity of zinc to Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):1–10, January 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-001.

McBride:1975:ETT

[1508] J. R. McBride and A. P. van Overbeeke. Effects of thiourea treatment on sexually maturing and gonadectomized male sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32 (1):11–19, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-002.

Larson:1975:SSM

[1509] K. W. Larson and G. W. Brown, Jr. Systematic status of a midwater population of freshwater sculpin (Cottus) from Lake Washington, Seattle, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32 (1):21–28, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-003.

Pippy:1975:IAL

[1510] John H. C. Pippy and Paul van Banning. Identification of Anisakis larva (I) as Anisakis simplex (Rudolphi, 1809, det. Krabbe 1878) (Nematoda: Ascaridata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):29–32, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-004.

Ware:1975:GMO

[1511] D. M. Ware. Growth, metabolism, and optimal swimming speed of a pelagic fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):33–41, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-005.

Ketchen:1975:AGD

[1512] K. S. Ketchen. Age and growth of dogfish Squalus acanthias in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32 (1):43–59, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-006.

Holub:1975:IGP

[1513] B. J. Holub, J. T. H. Connor, and S. J. Slinger. Incorporation of glycerol-3-phosphate into the hepatic lipids of rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):61–64, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-007.

Daniel:1975:CCP

[1514] Joseph C. Daniel, Jr. Concentrations of circulating progesterone during early pregnancy in the northern fur seal, Callorhinus ursinus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):65–66, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-008.

Messieh:1975:MSA

[1515] S. N. Messieh. Maturation and spawning of Atlantic herring (Clupea harengus harengus) in the Southern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):66–68, January 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-009.

Williams:1975:EAF

[1516] George Williams, Jr., Richard C. Simmonds, and John F. Boyd. Effects of the antimicrobial Furanace on Fundulus heteroclitus embryos. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):69–71, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-010.

Parker:1975:IPP

[1517] R. R. Parker, J. Sibert, and T. J. Brown. Inhibition of primary productivity through heterotrophic competition for nitrate in a stratified estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):72–77, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-011.

Fletcher:1975:CZT

[1518] G. L. Fletcher, E. G. Watts, and M. J. King. Copper, zinc, and total protein levels in the plasma of sockeye salmon (Oncorhynchus nerka) during their spawning migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):78–82, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-012.

Hatcher:1975:AGP

[1519] B. G. Hatcher and K. H. Mann. Above-ground production of marsh cordgrass (Spartina alterniflora) near the northern end of its range. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):83–87, January 1975. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f75-013.

Misra:1975:HSL

[1520] R. K. Misra. Hybridization in sunfishes (Lepomis, Centrarchidae) based on time-growth and static allometry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):88–91, January 1975. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f75-014.

Yamamoto:1975:IPN

[1521] T. Yamamoto. Infectious pancreatic necrosis virus and bacterial kidney disease appearing concurrently in populations of Salmo gairdneri and Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32 (1):92–95, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-015.

Anonymous:1975:EIA

[1522] Anonymous. Environmental impact assessment and hydroelectric projects: Hindsight and foresight in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1):97–209, January 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-016.

Margolis:1975:PFF

[1523] L. Margolis. Parasites of freshwater fishes — a review of their control and treatment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(1): 211, January 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f75-017.

Chisholm:1975:EIC

[1524] S. W. Chisholm, Raymond G. Stross, and Paul A. Nobbs. Environmental and intrinsic control of filtering and feeding rates in Arctic Daphnia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):219–226, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-018.

Biggar:1975:ODC

[1525] C. A. Biggar, N. A. M. Eskin, M. Vaisey, and B. Johnston. Oxidative deterioration of canned lake whitefish (Coregonus clupeaformis) treated with antioxidants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2): 227–232, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-019.

Groot:1975:MSS

[1526] C. Groot, K. Simpson, I. Todd, P. D. Murray, and G. A. Buxton. Movements of sockeye salmon (Oncorhynchus nerka) in the Skeena River Estuary as revealed by ultrasonic tracking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):233–242, February 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-020.

Chang:1975:MSA

[1527] B. D. Chang and T. R. Parsons. Metabolic studies on the amphipod Anisogammarus pugettensis in relation to its trophic position in the food web of young salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):243–247, February 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-021.

Beal:1975:SDF

[1528] Albert R. Beal. Selenium determination on fish tissue. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):249–252, February 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-022.

Collins:1975:CCJ

[1529] C. A. Collins, L. F. Giovando, and K. B. Abbott-Smith. Comparison of Canadian and Japanese merchant-ship observations of sea-surface temperature in the vicinity of present Ocean Weather Station “P,” 1927– 33. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):253–258, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-023.

Saether:1975:TNSa

[1530] Ole A. Sæther. Two new species of Heterotanytarsus Spärck, with keys to Nearctic and Palaearctic males and pupae of the genus (Diptera: Chironomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):259–270, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-024.

Ennis:1975:BRC

[1531] G. P. Ennis. Behavioral responses to changes in hydrostatic pressure and light during larval development of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):271–281, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-025.

Bryan:1975:SLW

[1532] J. E. Bryan and D. A. Kato. Spawning of lake whitefish, Coregonus clupeaformis, and round whiteftsh, Prosopium cylindraceum, in Aishihik Lake and East Aishihik River, Yukon Territory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):283–288, February 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-026.

Lear:1975:OAV

[1533] W. H. Lear and T. K. Pitt. Otolith age validation of Greenland halibut (Reinhardtius hippoglossoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):289–292, February 1975. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f75-027.

Murray:1975:TAP

[1534] Derek A. J. Murray. Trace analysis of phenols in water by gas chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):292–294, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-028.

Booke:1975:CSF

[1535] Henry E. Booke. Cytotaxonomy of the salmonid fish Stenodus leucichthys. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):295–296, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-029.

Ke:1975:IFF

[1536] P. J. Ke and R. G. Ackman. Inhibition of formation of free fatty acids in cod liver with sodium hypochlorite solution (Javex). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):297–300, February 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-030.

Kynard:1975:ESG

[1537] Boyd Kynard and Edward Lonsdale. Experimental study of galvanonarcosis for rainbow trout (Salmo gairdneri) immobilization. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):300–302, February 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-031.

Machniak:1975:ABT

[1538] Kazimierz Machniak and John H. Gee. Adjustment of buoyancy by tadpole madtom, Noturus gyrinus, and black bullhead, Ictalurus melas, in response to a change in water velocity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):303–307, February 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-032.

Power:1975:ASS

[1539] G. Power and Douglas Cressman. Atlantic salmon (Salmo salar) taken in the sea in North Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):307–309, February 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-033.

Sameoto:1975:DRE

[1540] D. D. Sameoto, D. C. Darrow, and S. Guildford. DDT residues in euphausiids in the Upper Estuary of the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):310–314, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-034.

Schimmel:1975:ABS

[1541] Steven C. Schimmel and David J. Hansen. An automatic brine shrimp feeder for aquatic bioassays. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):314–316, February 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-035.

Paradis:1975:DBN

[1542] M. Paradis and R. G. Ackman. Differentiation between natural hydrocarbons and low level Diesel oil contamination in cooked lobster meat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):316–320, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-036.

Hawkes:1975:EDW

[1543] Clifford L. Hawkes. Estimating dry weight of live, unanesthetized fish by photography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2): 321–324, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-037.

McAllister:1975:UF

[1544] D. E. McAllister. The urohyal in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):325, February 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-038.

Wolf:1975:DF

[1545] Ken Wolf. Diseases of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):325, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-039.

Stock:1975:SWG

[1546] J. H. Stock. Shallow-water gammaridean Amphipoda of New England. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(2):326, February 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-040.

Johnson:1975:BBP

[1547] B. Thomas Johnson and William Lulves. Biodegradation of di-n-butyl phthalate and di-2-ethylhexyl phthalate in freshwater hydrosoil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):333–339, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-041.

Kwik:1975:PDL

[1548] Judith K. Kwik and John C. H. Carter. Population dynamics of limnetic Cladocera in a beaver pond. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):341–346, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-042.

Sameoto:1975:TDE

[1549] D. D. Sameoto. Tidal and diurnal effects on zooplankton sample variability in a nearshore marine environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):347–366, March 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-043.

Saether:1975:TNSb

[1550] Ole A. Sæther. Two new species of Protanypus Kieffer, with keys to Nearctic and Palaearctic species of the genus (Diptera: Chironomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):367–388, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-044.

Sprules:1975:MCZ

[1551] W. Gary Sprules. Midsummer crustacean zooplankton communities in acid-stressed lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32 (3):389–395, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-045.

Kwain:1975:EDE

[1552] Wen hwa Kwain. Embryonic development, early growth, and meristic variation in rainbow trout (Salmo gairdneri) exposed to combinations of light intensity and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):397–402, March 1975. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f75-046.

Collins:1975:OPS

[1553] John J. Collins. Occurrence of pink salmon (Oncorhynchus gorbuscha) in Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3): 402–404, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-047.

Healey:1975:FCV

[1554] M. C. Healey and C. W. Nicol. Fecundity comparisons for various stocks of lake whitefish, Coregonus clupeaformis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):404–407, March 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-048.

Glenn:1975:AGR

[1555] C. L. Glenn. Annual growth rates of mooneye, Hiodon tergisus, in the Assiniboine River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32 (3):407–410, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-049.

Davis:1975:USP

[1556] John C. Davis and Richard A. W. Hoos. Use of sodium pentachlorophenate and dehydroabietic acid as reference toxicants for salmonid bioassays. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):411–416, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-050.

Hinton:1975:PFW

[1557] D. E. Hinton. Perfusion fixation of whole fish for electron microscopy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):416–422, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-051.

Rumsey:1975:AAS

[1558] G. L. Rumsey and H. G. Ketola. Amino acid supplementation of casein in diets of Atlantic salmon (Salmo salar) fry and of soybean meal for rainbow trout (Salmo gairdneri) fingerlings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):422–426, March 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-052.

Healey:1975:DEW

[1559] M. C. Healey. Dynamics of exploited whitefish populations and their management with special reference to the Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):427–448, March 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-053.

Iles:1975:MMF

[1560] T. Derrick Iles. The management of marine fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(3):449–450, March 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-054.

Murray:1975:ATA

[1561] Derek A. J. Murray. Analysis of tri-aryl phosphate esters and the determination of Imol S-140 in fish tissue and water samples by gas chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4): 457–460, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-055.

Selgeby:1975:LHA

[1562] James H. Selgeby. Life histories and abundance of crustacean zooplankton in the outlet of Lake Superior, 1971–72. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):461–470, April 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-056.

Messieh:1975:DSA

[1563] S. N. Messieh. Delineating spring and autumn herring populations in the Southern Gulf of St. Lawrence by discriminant function analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):471–477, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-057.

Scott:1975:ITP

[1564] J. S. Scott. Incidence of trematode parasites of American plaice (Hippoglossoides platessoides) of the Scotian Shelf and Gulf of St. Lawrence in relation to fish length and food. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):479–483, April 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-058.

Cherry:1975:TSA

[1565] Donald S. Cherry, Kenneth L. Dickson, and John Cairns, Jr. Temperatures selected and avoided by fish at various acclimation temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):485–491, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-059.

Kwain:1975:ETD

[1566] Wen hwa Kwain. Effects of temperature on development and survival of rainbow trout, Salmo gairdneri, in acid waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):493–497, April 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-060.

Winn:1975:PEH

[1567] H. E. Winn, R. K. Edel, and A. G. Taruski. Population estimate of the humpback whale (Megaptera novaeangliae) in the West Indies by visual and acoustic techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):499–506, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-061.

Eaton:1975:CCF

[1568] C. A. Eaton, R. G. Ackman, C. S. Tocher, and K. D. Spencer. Canadian capelin 1972–1973. fat and moisture composition, and fatty acids of some oils and lipid extract triglycerides. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):507–513, April 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-062.

Statham:1975:MDN

[1569] Charles N. Statham and John J. Lech. Metabolism of 2’ ,5-Dichloro-4’- nitrosalicylanilide (Bayer 73) in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):515–522, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-063.

Desautels:1975:PAS

[1570] Denise Desautels and R. M. MacKelvie. Practical aspects of survival and destruction of infectious pancreatic necrosis virus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):523–531, April 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-064.

Donaldson:1975:CCS

[1571] Edward M. Donaldson and Helen M. Dye. Corticosteroid concentrations in sockeye salmon (Oncorhynchus nerka) exposed to low concentrations of copper. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4): 533–539, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-065.

Varga:1975:IVP

[1572] S. Varga and W. A. Hirtle. Incidence of Vibrio parahaemolyticus in fish, shellfish, mud, water, and fish products in the Canadian Maritime region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):541–544, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-066.

DeFoe:1975:MTI

[1573] David L. DeFoe. Multichannel toxicant injection system for flow-through bioassays. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):544– 546, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-067.

Scott:1975:GVI

[1574] J. S. Scott. Geographic variation in incidence of trematode parasites of American plaice (Hippoglossoides platessoides) in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):547–550, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-068.

Wedemeyer:1975:SME

[1575] Gary A. Wedemeyer and Nancy C. Nelson. Statistical methods for estimating normal blood chemistry ranges and variance in rainbow trout (Salmo gairdneri), Shasta strain. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):551–554, April 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-069.

McBride:1975:ATK

[1576] J. McBride, G. Strasdine, and U. H. M. Fagerlund. Acute toxicity of kanamycin to steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):554–558, April 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-070.

McCleave:1975:EDT

[1577] James D. McCleave and Kristin A. Stred. Effect of dummy telemetry transmitters on stamina of Atlantic salmon (Salmo salar) smolts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):559–563, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-071.

Major:1975:HCC

[1578] Ronald D. Major, Joseph P. McCraren, and Charlie E. Smith. Histopathological changes in channel catfish (Ictalurus punctatus) experimentally and naturally infected with channel catfish virus disease. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):563–567, April 1975.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-072.

Yamamoto:1975:FDS

[1579] T. Yamamoto. Frequency of detection and survival of infectious pancreatic necrosis virus in a carrier population of brook trout (Salvelinus fontinalis) in a lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4): 568–570, April 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-073.

Utter:1975:GSF

[1580] Fred M. Utter. Genetic studies of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(4):571, April 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-074.

Jeffries:1975:DJA

[1581] H. Perry Jeffries. Diets of juvenile Atlantic menhaden (Brevoortia tyrannus) in three estuarine habitats as determined from fatty acid composition of gut contents. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5): 587–592, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-075.

Mehrle:1975:TEGa

[1582] Paul M. Mehrle and Foster L. Mayer, Jr. Toxaphene effects on growth and bone composition of fathead minnows, Pimephales promelas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):593–598, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-076.

Karl:1975:ATM

[1583] David M. Karl and Paul A. Larock. Adenosine triphosphate measurements in soil and marine sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):599–607, May 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-077.

Mehrle:1975:TEGb

[1584] Paul M. Mehrle and Foster L. Mayer, Jr. Toxaphene effects on growth and development of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):609–613, May 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-078.

Harrison:1975:CCD

[1585] P. G. Harrison and K. H. Mann. Chemical changes during the seasonal cycle of growth and decay in eelgrass (Zostera marina) on the Atlantic Coast of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5): 615–621, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-079.

Lett:1975:SSP

[1586] Patrick F. Lett, F. W. H. Beamish, and G. J. Farmer. System simulation of the predatory activities of sea lampreys (Petromyzon marinus) on lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):623–631, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-080.

Tautz:1975:SBC

[1587] A. F. Tautz and C. Groot. Spawning behavior of chum salmon (Oncorhynchus keta) and rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):633–642, May 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-081.

Peterson:1975:PFO

[1588] Richard H. Peterson. Pectoral fin and opercular movements of Atlantic salmon (Salmo salar) alevins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):643–647, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-082.

Fitzgerald:1975:NCS

[1589] Paul R. Fitzgerald. New coccidia from the spiny dogfish shark (Squalus acanthias) and great sculpin (Myoxocephalus polyacanthocephalus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):649–651, May 1975. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f75-083.

Carscadden:1975:MDS

[1590] James E. Carscadden and William C. Leggett. Meristic differences in spawning populations of American shad, Alosa sapidissima: Evidence for homing to tributaries in the St. John River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):653–660, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-084.

McNicholl:1975:EDM

[1591] Patricia G. McNicholl and W. C. Mackay. Effect of DDT and M.S. 222 on learning a simple conditioned response in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):661– 665, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-085.

Bustard:1975:AWE

[1592] David R. Bustard and David W. Narver. Aspects of the winter ecology of juvenile Coho salmon (Oncorhynchus kisutch) and steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5): 667–680, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-086.

Bustard:1975:PJC

[1593] David R. Bustard and David W. Narver. Preferences of juvenile Coho salmon (Oncorhynchus kisutch) and cutthroat trout (Salmo clarki) relative to simulated alteration of winter habitat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):681–687, May 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-087.

Craig:1975:MGA

[1594] P. C. Craig and V. A. Poulin. Movements and growth of Arctic grayling (Thymallus arcticus) and juvenile Arctic char (Salvelinus alpinus) in a small Arctic stream, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):689–697, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-088.

Lied:1975:SRT

[1595] E. Lied, J. Gjerde, and O. R. Braekkan. Simple and rapid technique for repeated blood sampling in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):699–701, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-089.

Cornick:1975:RCG

[1596] John W. Cornick and James E. Stewart. Red crab (Geryon quinquedens) and snow crab (Chionoecetes opilio) resistance to infection by the lobster pathogen Aerococcus viridans (var.) homari. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):702–706, May 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-090.

Poulet:1975:NFP

[1597] S. A. Poulet and J. P. Chanut. Nonselective feeding of Pseudocalanus minutus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):706– 713, May 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-091.

Hoffman:1975:EWT

[1598] Ronald S. Hoffman, Philip J. Dunham, and Patrick V. Kelly. Effects of water temperature and housing conditions upon the aggressive behavior of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):713–717, May 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-092.

Turner:1975:SMI

[1599] Ruth D. Turner. The shell makers, introducing mollusks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):719–720, May 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-093.

Edwards:1975:SFR

[1600] R. L. Edwards. Sea fisheries research. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(5):720, May 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-094.

Sutterlin:1975:CAS

[1601] A. M. Sutterlin. Chemical attraction of some marine fish in their natural habitat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):729– 738, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-095.

Murchelano:1975:QQS

[1602] R. A. Murchelano, C. Brown, and J. Bishop. Quantitative and qualitative studies of bacteria isolated from sea water used in the laboratory culture of the American oyster, Crassostrea virginica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):739–745, June 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-096.

Dagbjartsson:1975:UBW

[1603] Björn Dagbjartsson. Utilization of blue whiting, Micromesistius poutassou, for human consumption. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):747–751, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-097.

McLeay:1975:EAE

[1604] D. J. McLeay and D. A. Brown. Effects of acute exposure to bleached kraft pulpmill effluent on carbohydrate metabolism of juvenile Coho salmon (Oncorhynchus kisutch) during rest and exercise. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):753–760, June 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-098.

Miller:1975:DCS

[1605] R. J. Miller. Density of the commercial spider crab, Chionoecetes opilio, and calibration of effective area fished per trap using bottom photography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):761– 768, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-099.

Marking:1975:EPT

[1606] Leif L. Marking. Effects of pH on toxicity of antimycin to fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):769–773, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-100.

Dingle:1975:PIM

[1607] J. R. Dingle and J. A. Hines. Protein instability in minced flesh from fillets and frames of several commercial Atlantic fishes during storage at −5 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):775– 783, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-101.

McNicholl:1975:EDD

[1608] Patricia G. McNicholl and W. C. Mackay. Effect of DDT on discriminating ability of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):785–788, June 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-102.

Howard:1975:SPJ

[1609] T. E. Howard. Swimming performance of juvenile Coho salmon (Oncorhynchus kisutch) exposed to bleached kraft pulpmill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):789–793, June 1975. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f75-103.

Smith:1975:MST

[1610] Thomas G. Smith and F. A. J. Armstrong. Mercury in seals, terrestrial carnivores, and principal food items of the Inuit, from Holman, N.W.T. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):795–801, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-104.

Niimi:1975:STE

[1611] A. J. Niimi and Q. N. LaHam. Selenium toxicity on the early life stages of zebrafish (Brachydanio rerio). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):803–806, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-105.

Engelhardt:1975:TCF

[1612] F. R. Engelhardt, J. R. Geraci, and B. L. Walker. Tocopherol composition of frozen Atlantic herring (Clupea harengus harengus) tissues and oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):807–809, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-106.

Burnham:1975:ESW

[1613] B. L. Burnham and J. J. Peterka. Effects of saline water from North Dakota lakes on survival of fathead minnow (Pimephales promelas) embryos and sac fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6): 809–812, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-107.

Drew:1975:SSC

[1614] R. E. Drew. Simplified spectrophotometric curcumin method for the determination of boron in marine shellfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):813–816, June 1975. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f75-108.

Richey:1975:EKS

[1615] Jeffrey E. Richey, Michael A. Perkins, and Charles R. Goldman. Effects of Kokanee salmon (Oncorhynchus nerka) decomposition on the ecology of a subalpine stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6): 817–820, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-109.

Balon:1975:RGF

[1616] Eugene K. Balon. Reproductive guilds of fishes: A proposal and definition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):821– 864, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-110.

Carrothers:1975:NMF

[1617] P. J. G. Carrothers. Netting materials for fishing gear. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):865–866, June 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-111.

Carrothers:1975:MIF

[1618] P. J. G. Carrothers. Modem inshore fishing gear — rigging and mending. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):866, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-112.

Sprague:1975:PAU

[1619] John B. Sprague. The poisons around us. toxic metals in food, air, and water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(6):867, June 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-113.

Anonymous:1975:RBF

[1620] Anonymous. Review of Biology and Fisheries for Smaller Cetaceans contents. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):875– 877, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-114.

Anonymous:1975:F

[1621] Anonymous. Foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):879–882, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-115.

Anonymous:1975:P

[1622] Anonymous. Preface. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7): 883–886, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-116.

Anonymous:1975:RMS

[1623] Anonymous. Report of the Meeting on Smaller Cetaceans Montreal, April 1–11, 1974. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7): 889–983, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-117.

Mitchell:1975:SIM

[1624] Edward Mitchell and V. Michael Kozicki. Supplementary information on minke whale (Balaenoptera acutorostrata) from Newfoundland fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):985–994, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-118.

Ohsumi:1975:BPA

[1625] Seiji Ohsumi and Yasuaki Masaki. Biological parameters of the Antarctic minke whale at the virginal population level. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):995–1004, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-119.

Ross:1975:NRP

[1626] G. J. B. Ross, P. B. Best, and B. G. Donnelly. New records of the pygmy right whale (Caperea marginata) from South Africa, with comments on distribution, migration, appearance, and behavior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1005–1017, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-120.

Mitchell:1975:ASN

[1627] Edward Mitchell and V. Michael Kozicki. Autumn stranding of a northern bottlenose whale (Hyperoodon ampullatus) in the Bay of Fundy, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1019– 1040, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-121.

Mansfield:1975:NMM

[1628] A. W. Mansfield, T. G. Smith, and B. Beck. The narwhal, Monodon monoceros, in Eastern Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1041–1046, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-122.

Sergeant:1975:IAP

[1629] D. E. Sergeant and P. F. Brodie. Identity, abundance, and present status of populations of white whales, Delphinapterus leucas, in North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1047–1054, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-123.

Odell:1975:SAL

[1630] Daniel K. Odell. Status and aspects of the life history of the bottlenose dolphin, Tursiops truncatus, in Florida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1055–1058, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-124.

Perrin:1975:DDP

[1631] William F. Perrin. Distribution and differentiation of populations of dolphins of the genus Stenella in the Eastern tropical Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1059–1067, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-125.

Nishiwaki:1975:EAS

[1632] Masaharu Nishiwaki. Ecological aspects of smaller cetaceans, with emphasis on the striped dolphin (Stenella coeruleoalba). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1069–1072, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-126.

Brownell:1975:PRB

[1633] Robert L. Brownell, Jr. Progress report on the biology of the franciscana dolphin, Pontoporia blainvillei, in Uruguayan waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1073–1078, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-127.

Kapel:1975:PNO

[1634] Finn O. Kapel. Preliminary notes on the occurrence and exploitation of smaller cetacea in Greenland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1079–1082, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-128.

Christensen:1975:PRN

[1635] Ivar Christensen. Preliminary report on the Norwegian fishery for small whales: Expansion of Norwegian whaling to Arctic and Northwest Atlantic waters, and Norwegian investigations of the biology of small whales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1083– 1094, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-129.

Brown:1975:RBS

[1636] S. G. Brown. Relation between stranding mortality and population abundance of smaller cetacea in the Northeast Atlantic Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1095–1099, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-130.

ORiordan:1975:LFP

[1637] C. E. O’Riordan. Long-finned pilot whales, Globicephala melaena, driven ashore in Ireland, 1800–1973. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1101–1103, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-131.

Caldwell:1975:DSW

[1638] David K. Caldwell and Melba C. Caldwell. Dolphin and small whale fisheries of the Caribbean and West Indies: Occurrence, history, and catch statistics — with special reference to the Lesser Antillean Island of St. Vincent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7): 1105–1110, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-132.

Ohsumi:1975:RJS

[1639] Seiji Ohsumi. Review of Japanese small-type whaling. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1111–1121, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-133.

AguayoL:1975:PRS

[1640] Anelio Aguayo L. Progress report on small cetacean research in Chile. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1123–1143, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-134.

Mercer:1975:MLD

[1641] M. C. Mercer. Modified Leslie–DeLury population models of the longfinned pilot whale (Globicephala melaena) and annual production of the short-finned squid (Illex illecebrosus) based upon their interaction at Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7): 1145–1154, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-135.

Mead:1975:PRF

[1642] James G. Mead. Preliminary report on the former net fisheries for Tursiops truncatus in the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1155–1162, July 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-136.

Foote:1975:ISW

[1643] Don Charles Foote. Investigation of small whale hunting in Northern Norway, 1964. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7): 1163–1189, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-137.

Asper:1975:TLC

[1644] Edward D. Asper. Techniques of live capture of smaller cetacea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1191–1196, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-138.

Walker:1975:RLC

[1645] William A. Walker. Review of the live-capture fishery for smaller cetaceans taken in Southern California waters for public display, 1966–73. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1197–1211, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-139.

Bigg:1975:LCK

[1646] Michael A. Bigg and Allen A. Wolman. Live-capture killer whale (Orcinus orca) fishery, British Columbia and Washington, 1962–73. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1213–1221, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-140.

Lear:1975:CHP

[1647] W. H. Lear and O. Christensen. By-catches of harbour porpoise (Phocoena phocoena) in salmon driftnets at West Greenland in 1972. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1223–1228, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-141.

Ohsumi:1975:ICC

[1648] Seiji Ohsumi. Incidental catch of cetaceans with salmon gillnet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1229–1235, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-142.

Brown:1975:MSC

[1649] S. G. Brown. Marking of small cetaceans using “discovery” type whale marks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(7):1237– 1240, July 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-143.

Shotts:1975:BDF

[1650] Emmett B. Shotts, Jr. and G. L. Bullock. Bacterial diseases of fishes: Diagnostic procedures for Gram-negative pathogens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1243–1247, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-144.

Leach:1975:IIC

[1651] J. M. Leach and A. N. Thakore. Isolation and identification of constituents toxic to juvenile rainbow trout (Salmo gairdneri) in caustic extraction effluents from kraft pulpmill bleach plants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1249–1257, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-145.

Snyder:1975:PFE

[1652] Darrel E. Snyder. Passage of fish eggs and young through a pumped storage generating station. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1259–1266, August 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-146.

MacKelvie:1975:FVS

[1653] R. M. MacKelvie and Denise Desautels. Fish viruses — survival and inactivation of infectious pancreatic necrosis virus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1267–1273, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-147.

Penrose:1975:LAD

[1654] William R. Penrose, Robert Black, and Michael J. Hayward. Limited arsenic dispersion in sea water, sediments, and biota near a continuous source. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1275–1281, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-148.

Barica:1975:SRP

[1655] Jan Barica. Summerkill risk in prairie ponds and possibilities of its prediction. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1283–1288, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-149.

Bouck:1975:CSL

[1656] G. R. Bouck, P. W. Schneider, Jr., Janet Jacobson, and R. C. Ball. Characterization and subcellular localization of leucine aminonaphthylamidase (LAN) in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1289–1295, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-150.

Smith:1975:UMF

[1657] Annette L. Smith, Roger H. Green, and Andrew Lutz. Uptake of mercury by freshwater clams (family Unionidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1297–1303, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-151.

Caddy:1975:SME

[1658] J. F. Caddy. Spatial model for an exploited shellfish population, and its application to the Georges Bank scallop fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1305–1328, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-152.

Sakaguchi:1975:MPP

[1659] M. Sakaguchi, Doris Fraser Hiltz, and W. J. Dyer. Metabolism of pyruvate in postmortem adductor muscle of the sea scallop (Placopecten magellanicus) during iced storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1329–1337, August 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-153.

Renfro:1975:RIF

[1660] W. C. Renfro, S. W. Fowler, M. Heyraud, and J. La Rosa. Relative importance of food and water in long-term zinc^{65} accumulation by marine biota. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1339–1345, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-154.

Moores:1975:MBC

[1661] J. A. Moores, G. H. Winters, and L. S. Parsons. Migrations and biological characteristics of Atlantic mackerel (Scomber scombrus) occurring in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1347–1357, August 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-155.

Flick:1975:MGS

[1662] William A. Flick and Dwight A. Webster. Movement, growth, and survival in a stream population of wild brook trout (Salvelinus fontinalis) during a period of removal of non-trout species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1359–1367, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-156.

Ricker:1975:RIH

[1663] W. E. Ricker and H. D. Smith. A revised interpretation of the history of the Skeena River sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1369–1381, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-157.

Pitt:1975:CAC

[1664] T. K. Pitt. Changes in abundance and certain biological characteristics of Grand Bank American plaice, Hippoglossoides platessoides. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1383–1398, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-158.

Webster:1975:SOM

[1665] T. J. M. Webster, Madhu A. Paranjape, and K. H. Mann. Sedimentation of organic matter in St. Margaret’s Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1399–1407, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-159.

Kelso:1975:SDN

[1666] John R. M. Kelso and C. K. Minns. Summer distribution of the nearshore fish community near a thermal generating station as determined by acoustic census. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8): 1409–1418, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-160.

Feller:1975:SSP

[1667] Robert J. Feller and Victor W. Kaczynski. Size selective predation by juvenile chum salmon (Oncorhynchus keta) on epibenthic prey in Puget Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1419–1429, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-161.

Castell:1975:CRJ

[1668] John D. Castell, Edgar G. Mason, and Jane F. Covey. Cholesterol requirements of juvenile American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1431–1435, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-162.

Nicholson:1975:PII

[1669] Bruce L. Nicholson and Roger Dexter. Possible interference in the isolation of IPN virus from carrier fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1437–1439, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-163.

Moulton:1975:LHO

[1670] Lawrence L. Moulton. Life history observations on the Puget Sound rockfish, Sebastes emphaeus (Starks, 1911). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1439–1442, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-164.

Lorz:1975:DHA

[1671] H. (Van Arsdale) Lorz and W. G. Pearcy. Distribution of hyperiid amphipods off the Oregon Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1442–1447, August 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-165.

Markle:1975:YWF

[1672] Douglas F. Markle. Young witch flounder, Glyptocephalus cynoglossus, on the slope off Virginia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8): 1447–1450, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-166.

Lall:1975:PTI

[1673] Barbara Smith Lall, Alison R. Manzer, and Doris Fraser Hiltz. Preheat treatment for improvement of frozen storage stability at −10 C in fillets and minced flesh of silver hake (Merluccius bilinearis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1450–1454, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-167.

Maki:1975:CTL

[1674] Alan W. Maki, Lester Geissel, and Howard E. Johnson. Comparative toxicity of larval lampricide (TFM: 3-trifluoromethyl-4-nitrophenol) to selected benthic macroinvertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1455–1459, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-168.

Kendall:1975:HMA

[1675] M. W. Kendall and W. E. Hawkins. Hepatic morphology and acid phosphatase localization in the channel catfish (Ictalurus punctatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1459–1464, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-169.

Ricker:1975:FRB

[1676] W. E. Ricker. The Fisheries Research Board of Canada — seventy-five years of achievements. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8): 1465–1490, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-170.

Jolicoeur:1975:LRF

[1677] Pierre Jolicoeur. Linear regressions in fishery research: Some comments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1491–1494, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-171.

Ricker:1975:NCP

[1678] W. E. Ricker. A note concerning Professor Jolicoeur’s comments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1494–1498, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-172.

Bell:1975:CAD

[1679] Gordon R. Bell. Color atlas of the diseases of fishes, amphibians and reptiles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1499, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-173.

Royce:1975:FNP

[1680] William F. Royce. Fisheries of the North Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1500, August 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-174.

Antia:1975:APB

[1681] Naval J. Antia. Algal physiology and biochemistry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1500–1501, August 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-175.

Hynes:1975:PEF

[1682] H. B. N. Hynes. Pollution ecology of freshwater invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1501–1502, August 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-176.

Anonymous:1975:Oa

[1683] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(8):1503–1504, August 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-177.

Dillon:1975:SMP

[1684] P. J. Dillon and F. H. Rigler. A simple method for predicting the capacity of a lake for development based on lake trophic status. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1519–1531, September 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-178.

Pitcher:1975:PMD

[1685] T. J. Pitcher. A periscopic method for determining the three-dimensional positions of fish in schools. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1533–1538, September 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-179.

Hayden:1975:BOD

[1686] Jonathan B. Hayden, Joseph J. Cech, Jr., and David W. Bridges. Blood oxygen dissociation characteristics of the winter flounder, Pseudopleuronectes americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9): 1539–1544, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-180.

Udey:1975:RWT

[1687] L. R. Udey, J. L. Fryer, and K. S. Pilcher. Relation of water temperature to ceratomyxosis in rainbow trout (Salmo gairdneri) and Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1545–1551, September 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-181.

Holt:1975:RWT

[1688] R. A. Holt, J. E. Sanders, J. L. Zinn, J. L. Fryer, and K. S. Pilcher. Relation of water temperature to Flexibacter columnaris infection in steelhead trout (Salmo gairdneri), Coho (Oncorhynchus kisutch) and Chinook (O. tshawytscha) salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9): 1553–1559, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-182.

Pickard:1975:ALT

[1689] G. L. Pickard. Annual and longer term variations of deepwater properties in the coastal waters of Southern British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1561–1587, September 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-183.

Archibald:1975:EOE

[1690] Chris P. Archibald. Experimental observations on the effects of predation by goldfish (Carassius auratus) on the zooplankton of a small saline lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1589–1594, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-184.

Nowlan:1975:TDC

[1691] Sandra S. Nowlan, W. J. Dyer, and R. A. Keith. Temperature and deteriorative changes in postrigor cod muscle stored up to 14 days in the superchill range, −1 to −4 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1595–1605, September 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-185.

Smith:1975:RCC

[1692] K. L. Smith, Jr., G. R. Harbison, G. T. Rowe, and C. H. Clifford. Respiration and chemical composition of Pleuroncodes planipes (Decapoda: Galatheidae): Energetic significance in an upwelling system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1607–1612, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-186.

Lett:1975:RSB

[1693] P. F. Lett, A. C. Kohler, and D. N. Fitzgerald. Role of stock biomass and temperature in recruitment of Southern Gulf of St. Lawrence Atlantic cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9): 1613–1627, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-187.

Tran:1975:SCM

[1694] V. D. Tran. Solubilization of cod myofibrillar proteins at low ionic strength by sodium pyruvate during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1629–1632, September 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-188.

Holub:1975:BLC

[1695] B. J. Holub, K. Nilsson, J. Piekarski, and S. J. Slinger. Biosynthesis of lecithin by the CDP-choline pathway in liver microsomes of rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1633–1637, September 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-189.

Gallucci:1975:MTT

[1696] Vincent F. Gallucci and Robert K. Kawaratani. Mortality of Transennella tantilla due to burial. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1637–1640, September 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-190.

Margolis:1975:CSM

[1697] L. Margolis and T. P. T. Evelyn. Ceratomyxa shasta (Myxosporida) disease in chum salmon (Oncorhynchus keta) in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1640–1643, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-191.

Paradis:1975:UWL

[1698] M. Paradis, R. G. Ackman, J. Hingley, and C. A. Eaton. Utilization of wastes from lumpfish, Cyclopterus lumpus, roe harvesting operations: An examination of the lipid and glue potential, and comparison of meal with that from Nova Scotia-caught menhaden. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1643–1648, September 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-192.

Ross:1975:PHF

[1699] A. J. Ross, W. T. Yasutake, and Steve Leek. Phoma herbarum, a fungal plant saprophyte, as a fish pathogen. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1648–1652, September 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-193.

Gibson:1975:SBT

[1700] R. John Gibson and G. Power. Selection by brook trout (Salvelinus fontinalis) and juvenile Atlantic salmon (Salmo salar) of shade related to water depth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9): 1652–1656, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-194.

Mills:1975:BOS

[1701] Eric L. Mills. Benthic organisms and the structure of marine ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1657–1663, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-195.

Balon:1975:TIF

[1702] Eugene K. Balon. Terminology of intervals in fish development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1663–1670, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-196.

Marr:1975:FOR

[1703] John C. Marr. Freedom of oceanic research: A study conducted by the Center for Marine Affairs of the Scripps Institution of Oceanography, University of California, San Diego. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1671–1672, September 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-197.

Hershberger:1975:GMF

[1704] William K. Hershberger. Genetics and mutagenesis of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1672–1673, September 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-198.

Schindler:1975:MEP

[1705] D. W. Schindler. Modelling the eutrophication process. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1673–1674, September 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-199.

McLeese:1975:CMO

[1706] D. W. McLeese and A. M. Sutterlin. Chemoreception in marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1674, September 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-200.

Pauley:1975:FI

[1707] Gilbert B. Pauley. Fish immunology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(9):1674–1675, September 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-201.

Marliave:1975:SSS

[1708] Jeffrey B. Marliave. Seasonal shifts in the spawning site of a Northeast Pacific intertidal fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10): 1687–1691, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-202.

Matthews:1975:SCW

[1709] J. B. Matthews and A. V. Quinlan. Seasonal characteristics of water masses in Muir Inlet, a fjord with tidewater glaciers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1693–1703, October 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-203.

Frederick:1975:CUP

[1710] Lynn L. Frederick. Comparative uptake of a polychlorinated biphenyl and dieldrin by the white sucker (Catostomus commersoni). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1705–1709, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-204.

Braekevelt:1975:PFS

[1711] C. R. Braekevelt. Photoreceptor fine structure in the northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1711–1721, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-205.

Angell:1975:ETP

[1712] Charles L. Angell, Bruce S. Miller, and S. R. Wellings. Epizootiology of tumors in a population of juvenile English sole (Parophrys vetulus) from Puget Sound, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1723–1732, October 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-206.

Busch:1975:EFA

[1713] Wolf-Dieter N. Busch, Russell L. Scholl, and Wilbur L. Hartman. Environmental factors affecting the strength of walleye (Stizostedion vitreum vitreum) year-classes in Western Lake Erie, 1960–70. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1733–1743, October 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-207.

Kelly:1975:HCM

[1714] T. M. Kelly, J. D. Jones, and G. R. Smith. Historical changes in mercury contamination in Michigan walleyes (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1745–1754, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-208.

Ramamoorthy:1975:HMB

[1715] S. Ramamoorthy and D. J. Kushner. Heavy metal binding components of river water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10): 1755–1766, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-209.

Fowler:1975:MIU

[1716] Bruce A. Fowler, Douglas A. Wolfe, and William F. Hettler. Mercury and iron uptake by cytosomes in mantle epithelial cells of quahog clams (Mercenaria mercenaria) exposed to mercury. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1767–1775, October 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-210.

Walters:1975:OHS

[1717] Carl J. Walters. Optimal harvest strategies for salmon in relation to environmental variability and uncertain production parameters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1777–1784, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-211.

Nakashima:1975:YPP

[1718] Brian S. Nakashima and William C. Leggett. Yellow perch (Perca flavescens) biomass responses to different levels of phytoplankton and benthic biomass in Lake Memphremagog, Quebec–Vermont. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1785–1797, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-212.

Cherry:1975:SDD

[1719] Donald S. Cherry and Rufus K. Guthrie. Significance of detritus or detritus-associated invertebrates to fish production in a new impoundment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1799–1804, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-213.

Rowe:1975:QBA

[1720] Gilbert T. Rowe, Pamela T. Polloni, and Richard L. Haedrich. Quantitative biological assessment of the benthic fauna in deep basins of the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10): 1805–1812, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-214.

Shaw:1975:EDS

[1721] H. M. Shaw, R. L. Saunders, H. C. Hall, and E. B. Henderson. Effect of dietary sodium chloride on growth of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1813–1819, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-215.

Shaw:1975:ESF

[1722] H. M. Shaw, R. L. Saunders, and H. C. Hall. Environmental salinity: Its failure to influence growth of Atlantic salmon (Salmo salar) Parr. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1821–1824, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-216.

Anderson:1975:AOA

[1723] Lee G. Anderson. Analysis of open-access commercial exploitation and maximum economic yield in biologically and technologically interdependent fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10): 1825–1842, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-217.

Aiken:1975:TIC

[1724] D. E. Aiken and S. L. Waddy. Temperature increase can cause hyperecdysonism in American lobsters (Homarus americanus) injected with ecdysterone. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1843– 1845, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-218.

Darrow:1975:AAA

[1725] D. C. Darrow and G. C. H. Hakding. Accumulation and apparent absence of DDT metabolism by marine copepods, Calanus spp., in culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1845–1849, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-219.

Veith:1975:ACS

[1726] G. D. Veith and V. M. Comstock. Apparatus for continuously saturating water with hydrophobic organic chemicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1849–1851, October 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-220.

Eberhardt:1975:SMA

[1727] L. L. Eberhardt. Some methodology for appraising contaminants in aquatic systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10): 1852–1859, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-221.

Vinyard:1975:DLR

[1728] Gary L. Vinyard and W. John O’Brien. Dorsal light response as an index of prey preference in bluegill (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1860–1863, October 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-222.

Meyerhoff:1975:ATB

[1729] Roger D. Meyerhoff. Acute toxicity of benzene, a component of crude oil, to juvenile striped bass (Morone saxatilis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1864–1866, October 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-223.

Magnin:1975:CBI

[1730] Etienne Magnin and Catherine Fradette. Croissance de la barbue (Ictalurus punctatus) du fleuve Saint-Laurent près de Québec. (French) [Growth of catfish (Ictalurus punctatus) of the Saint Lawrence River near Quebec]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1867–1870, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-224.

Diaz:1975:EBT

[1731] Robert J. Diaz. Effects of brief temperature increases on larvae of Mulinia lateralis (Mollusca: Bivalvia). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1870–1872, October 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-225.

Hunn:1975:REC

[1732] Joseph B. Hunn and John L. Allen. Renal excretion in Coho salmon (Oncorhynchus kisutch) after acute exposure to 3-trifluoromethyl-4-nitrophenol. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1873– 1876, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-226.

Baross:1975:NIL

[1733] John A. Baross, F. J. Hanus, Robert P. Griffiths, and Richard Y. Morita. Nature of incorporated ^{14}C-labeled material retained by sulfuric acid fixed bacteria in pure cultures and in natural aquatic populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1876–1879, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-227.

Buchanan:1975:IFW

[1734] David V. Buchanan, Michael J. Myers, and Richard S. Caldwell. Improved flowing water apparatus for the culture of brachyuran crab larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1880–1883, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-228.

Cripe:1975:AQD

[1735] Claude R. Cripe, Jesse H. Cripe, and Robert J. Livingston. Apparatus for the quantitative determination of locomotor activity patterns of aquatic organisms using infrared light-emitting diodes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1884–1885, October 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-229.

Regier:1975:SCS

[1736] H. A. Regier and F. D. McCracken. Science for Canada’s shelf-seas fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1887–1932, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-230.

Houde:1975:ELH

[1737] Edward D. Houde. The early life history of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1933–1934, October 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-231.

Dillon:1975:CLW

[1738] P. J. Dillon. Coupling of land and water systems. Ecological studies 10. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(10):1934–1935, October 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-232.

Hamley:1975:RGS

[1739] John M. Hamley. Review of gillnet selectivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):1943–1969, November 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-233.

Johnson:1975:PCC

[1740] Lionel Johnson. Physical and chemical characteristics of Great Bear Lake, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):1971–1987, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-234.

Johnson:1975:DFS

[1741] Lionel Johnson. Distribution of fish species in Great Bear Lake, Northwest Territories, with reference to zooplankton, benthic invertebrates, and environmental conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):1989–2004, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-235.

Wobeser:1975:ATM

[1742] G. Wobeser. Acute toxicity of methyl mercury chloride and mercuric chloride for rainbow trout (Salmo gairdneri) fry and fingerlings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2005–2013, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-236.

Wobeser:1975:POA

[1743] G. Wobeser. Prolonged oral administration of methyl mercury chloride to rainbow trout (Salmo gairdneri) fingerlings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2015–2023, November 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-237.

Collins:1975:NAR

[1744] M. T. Collins, J. B. Gratzek, E. B. Shotts, Jr., D. L. Dawe, Lucy M. Campbell, and D. R. Senn. Nitrification in an aquatic recirculating system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2025–2031, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-238.

Collins:1975:EPN

[1745] M. T. Collins, J. B. Gratzek, D. L. Dawe, and T. G. Nemetz. Effects of parasiticides on nitrification. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2033–2037, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-239.

Shaw:1975:PSI

[1746] D. H. Shaw and J. R. Botta. Preservation studies of inshore male capelin (Mallotus villosus) stored in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2039–2046, November 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-240.

Shaw:1975:PIM

[1747] D. H. Shaw and J. R. Botta. Preservation of inshore male capelin (Mallotus villosus) stored in refrigerated sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2047–2053, November 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-241.

McLeese:1975:CRA

[1748] D. W. McLeese. Chemosensory response of American lobsters (Homarus americanus) in the presence of copper and phosphamidon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2055–2060, November 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-242.

King:1975:AEA

[1749] Deborah King and Dwight L. Baker. Activation energy and allosteric inhibition of reaction catalyzed by aspartate transcarbamylase from Atlantic cod (Gadus morhua) ovary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2061–2064, November 1975. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f75-243.

Wobeser:1975:PBL

[1750] G. Wobeser and N. W. Smith. Proliferative bone lesions in lake whitefish (Coregonus clupeaformis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2065–2068, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-244.

Hare:1975:APD

[1751] G. M. Hare and M. D. B. Burt. Abundance and population dynamics of parasites infecting Atlantic salmon (Salmo salar) in Trout Brook, New Brunswick, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11): 2069–2074, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-245.

Lafond:1975:DWE

[1752] C. A. Lafond and G. L. Pickard. Deep water exchanges in Bute Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11): 2075–2089, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-246.

Tracey:1975:BGL

[1753] Martin L. Tracey, Keith Nelson, Dennis Hedgecock, Robert A. Shleser, and Mary Lou Pressick. Biochemical genetics of lobsters: Genetic variation and the structure of American lobster (Homarus americanus) populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2091–2101, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-247.

Brett:1975:GRY

[1754] J. R. Brett and J. E. Shelbourn. Growth rate of young sockeye salmon, Oncorhynchus nerka, in relation to fish size and ration level. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2103–2110, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-248.

Bowes:1975:PDP

[1755] Gerald W. Bowes and Charles J. Jonkel. Presence and distribution of polychlorinated biphenyls (PCB) in Arctic and subarctic marine food chains. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2111–2123, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-249.

Polisois:1975:DPC

[1756] G. Polisois, A. Tessier, P. G. C. Campbell, and J. P. Villeneuve. Degradation of phenolic compounds downstream from a petroleum refinery complex. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2125–2131, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-250.

Beitinger:1975:ISR

[1757] Thomas L. Beitinger and John J. Magnuson. Influence of social rank and size on thermoselection behavior of bluegill (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2133–2136, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-251.

Ryder:1975:EPF

[1758] R. A. Ryder and H. F. Henderson. Estimates of potential fish yield for the Nasser Reservoir, Arab Republic of Egypt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2137–2151, November 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-252.

Schneider:1975:NFA

[1759] Phillip W. Schneider, Jr. and Lavern J. Weber. Neuromuscular function and acetylcholinesterase in the pectoral abductor muscle of largemouth bass (Micropterus salmoides) as evidenced by the effects of diisopropylfluorophosphate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11): 2153–2161, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-253.

Walter:1975:GME

[1760] Gilbert G. Walter. Graphical methods for estimating parameters in simple models of fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2163–2168, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-254.

Pedersen:1975:MGP

[1761] Mark G. Pedersen. Movements and growth of petrale sole (Eopsetta jordani) tagged off Washington and Southwest Vancouver Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2169–2177, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-255.

Peterman:1975:NTP

[1762] Randall M. Peterman. New techniques for policy evaluation in ecological systems: Methodology for a case study of Pacific salmon fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2179–2188, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-256.

Takahashi:1975:EAI

[1763] M. Takahashi and T. Ikeda. Excretion of ammonia and inorganic phosphorus by Euphausia pacifica and Metridia pacifica at different concentrations of phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2189–2195, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-257.

Argyle:1975:DDC

[1764] Ray L. Argyle, George C. Williams, and Clara B. Daniel. Dieldrin in the diet of channel catfish (Ictalurus punctatus): Uptake and effect on growth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2197–2204, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-258.

Taruski:1975:CHC

[1765] A. G. Taruski, C. E. Olney, and Howard E. Winn. Chlorinated hydrocarbons in cetaceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11): 2205–2209, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-259.

Ennis:1975:OHL

[1766] G. P. Ennis. Observations on hatching and larval release in the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2210–2213, November 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-260.

Newsome:1975:SCF

[1767] George E. Newsome and Gérard Leduc. Seasonal changes of fat content in the yellow perch (Perca flavescens) of two Laurentian lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2214–2221, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-261.

Thomas:1975:IOR

[1768] Robert E. Thomas and Stanley D. Rice. Increased opercular rates of pink salmon (Oncorhynchus gorbuscha) fry after exposure to the watersoluble fraction of Prudhoe Bay crude oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2221–2224, November 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-262.

Bullock:1975:FAI

[1769] G. L. Bullock and H. M. Stuckey. Fluorescent antibody identification and detection of the Corynebacterium causing kidney disease of salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2224–2227, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-263.

Eyman:1975:SEC

[1770] L. D. Eyman, C. W. Gehrs, and J. J. Beauchamp. Sublethal effect of 5-chlorouracil on carp (Cyprinus carpio) larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2227–2229, November 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-264.

Parsons:1975:BOC

[1771] T. R. Parsons. Biological oceanography in Canada: A perspective and review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2231–2283, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-265.

Efford:1975:FEP

[1772] Ian E. Efford. Fisheries and energy production. A symposium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2285, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f75-266.

Horn:1975:PCE

[1773] Henry S. Horn. Population and community ecology: Principles and methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(11):2285–2286, November 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-267.

Davis:1975:MDO

[1774] John C. Davis. Minimal dissolved oxygen requirements of aquatic life with emphasis on Canadian species: a review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2295–2332, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-268.

Love:1975:VAC

[1775] R. Malcolm Love. Variability in Atlantic cod (Gadus morhua) from the Northeast Atlantic: a review of seasonal and environmental influences on various attributes of the flesh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2333–2342, December 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-269.

Lear:1975:ETP

[1776] W. H. Lear. Evaluation of the transplant of Pacific pink salmon (Oncorhynchus gorbuscha) from British Columbia to Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2343–2356, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-270.

McLeay:1975:SBC

[1777] D. J. McLeay. Sensitivity of blood cell counts in juvenile Coho salmon (Oncorhynchus kisutch) to stressors including sublethal concentrations of pulpmill effluent and zinc. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2357–2364, December 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-271.

Youngs:1975:ESR

[1778] W. D. Youngs and D. S. Robson. Estimating survival rate from tag returns: Model tests and sample size determination. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2365–2371, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-272.

Therriault:1975:PDL

[1779] Jean Claude Therriault and Guy Lacroix. Penetration of the deep layer of the Saguenay Fjord by surface waters of the St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2373–2377, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-273.

Emerson:1975:AAE

[1780] Kenneth Emerson, Rosemarie C. Russo, Richard E. Lund, and Robert V. Thurston. Aqueous ammonia equilibrium calculations: Effect of pH and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2379– 2383, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f75-274.

Penrose:1975:BOA

[1781] W. R. Penrose. Biosynthesis of organic arsenic compounds in brown trout (Salmo trutta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2385–2390, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-275.

Buhler:1975:HMC

[1782] Donald R. Buhler, Robert R. Claeys, and Bruce R. Mate. Heavy metal and chlorinated hydrocarbon residues in California sea lions (Zalophus californianus californianus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2391–2397, December 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-276.

Westrheim:1975:RMI

[1783] S. J. Westrheim. Reproduction, maturation, and identification of larvae of some Sebastes (Scorpaenidae) species in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2399–2411, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-277.

Stockner:1975:PSP

[1784] John G. Stockner and K. R. S. Shortreed. Phytoplankton succession and primary production in Babine Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2413–2427, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-278.

Busch:1975:EAC

[1785] R. A. Busch and A. J. Lingg. Establishment of an asymptomatic carrier state infection of enteric redmouth disease in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2429–2432, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-279.

Flint:1975:GPC

[1786] R. Warren Flint. Growth in a population of the crayfish Pacifastacus leniusculus from a subalpine lacustrine environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2433–2440, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-280.

Sutterlin:1975:PRP

[1787] A. M. Sutterlin and Susan Waddy. Possible role of the posterior lateral line in obstacle entrainment by brook trout (Salvelinus fontinalis).

Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2441–2446, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-281.

Lander:1975:MDN

[1788] Robert H. Lander. Method of determining natural mortality in the northern fur seal (Callorhinus ursinus) from known pups and kill by age and sex. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2447–2452, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-282.

Lindsey:1975:PWT

[1789] C. C. Lindsey. Pleomerism, the widespread tendency among related fish species for vertebral number to be correlated with maximum body length. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2453–2469, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-283.

Buroker:1975:GVP

[1790] Norman E. Buroker, William K. Hershberger, and Kenneth K. Chew. Genetic variation in the Pacific oyster, Crassostrea gigas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2471–2477, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-284.

Brett:1975:DPN

[1791] J. R. Brett and C. A. Zala. Daily pattern of nitrogen excretion and oxygen consumption of sockeye salmon (Oncorhynchus nerka) under controlled conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2479–2486, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-285.

Antia:1975:CSE

[1792] Naval J. Antia and Joseph Y. Cheng. Culture studies on the effects from borate pollution on the growth of marine phytoplankters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2487–2494, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-286.

Latimer:1975:TME

[1793] David L. Latimer, Arthur S. Brooks, and Alfred M. Beeton. Toxicity of 30-minute exposures of residual chlorine to the copepods Limnocalanus macrurus and Cyclops bicuspidatus thomasi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2495–2501, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-287.

Ware:1975:RBE

[1794] D. M. Ware. Relation between egg size, growth, and natural mortality of larval fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2503–2512, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-288.

Atchison:1975:FAL

[1795] G. J. Atchison. Fatty acid levels in developing brook trout (Salvelinus fontinalis) eggs and fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2513–2515, December 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-289.

Stasko:1975:MLT

[1796] Aivars B. Stasko. Modified lobster traps for catching crabs and keeping lobsters out. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2515–2520, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-290.

Sinclair:1975:SBP

[1797] William F. Sinclair and R. W. Morley. Statistical bias problem in onsite surveys: the severity of the problem and its potential for solution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2520–2524, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-291.

Roberts:1975:ATC

[1798] Morris H. Roberts, Jr., Robert J. Diaz, Michael E. Bender, and Robert J. Huggett. Acute toxicity of chlorine to selected estuarine species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2525–2528, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-292.

Brown:1975:EBK

[1799] D. A. Brown and D. J. McLeay. Effect of bleached kraft mill effluent on the survival of starved juvenile Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2528–2530, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-293.

Pope:1975:CPC

[1800] G. F. Pope and J. C. H. Carter. Crustacean plankton communities of the Matamek River system and their variation with predation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2530–2535, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-294.

Dadswell:1975:SBO

[1801] M. J. Dadswell. A small-boat otter trawl for limnological studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2535–2538, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-295.

Hargens:1975:ESS

[1802] Alan R. Hargens and M. Perez. Edema in spawning salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2538–2541, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-296.

Berg:1975:FOB

[1803] Richard E. Berg, Philip A. Doepke, and Paul R. Hannuksela. First occurrence of the brook silverside, Labidesthes sicculus, in a tributary of Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2541–2542, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-297.

Mason:1975:SMJ

[1804] John C. Mason. Seaward movement of juvenile fishes, including lunar periodicity in the movement of Coho salmon (Oncorhynchus kisutch) fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2542–2547, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-298.

Weimer:1975:ATI

[1805] R. T. Weimer and J. E. Ehrenberg. Analysis of threshold-induced bias inherent in acoustic scattering cross-section estimates of individual fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2547–2551, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-299.

Hodson:1975:ZUA

[1806] Peter V. Hodson. Zinc uptake by Atlantic salmon (Salmo salar) exposed to a lethal concentration of zinc at 3, 11, and 19 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2552–2556, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-300.

Leach:1975:ATJ

[1807] J. M. Leach, A. N. Thakore, and J. F. Manville. Acute toxicity to juvenile rainbow trout (Salmo gairdneri) of naturally occurring insect juvenile hormone analogues. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2556–2559, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-301.

Geraci:1975:FHR

[1808] J. R. Geraci and T. G. Smith. Functional hematology of ringed seals (Phoca hispida) in the Canadian Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2559–2564, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-302.

Busdosh:1975:RTA

[1809] M. Busdosh and Ronald M. Atlas. Response of two Arctic amphipods, Gammarus zaddachi and Boeckosimus (Onisimus) affinis, to variations in temperature and salinity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2564–2568, December 1975. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f75-303.

Tomlinson:1975:IJV

[1810] J. W. C. Tomlinson and I. Vertinsky. International joint ventures in fishing and 200-mile economic zones. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2569–2579, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-304.

Fletcher:1975:EMB

[1811] G. L. Fletcher. Experimental marine biology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2581, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-305.

Withler:1975:STF

[1812] F. C. Withler. The salmon: Their fight for survival. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2582, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-306.

Fryer:1975:IWT

[1813] Geoffrey Fryer. The inland waters of tropical Africa. An introduction to tropical limnology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12): 2582–2584, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-307.

Hiltunen:1975:BL

[1814] Jarl Hiltunen. The benthos of lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2584, December 1975. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f75-308.

Christie:1975:BAO

[1815] W. J. Christie. The blackbass in America and overseas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2584–2585, December 1975. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f75-309.

Kozloff:1975:BRL

[1816] Eugene N. Kozloff. Book review: Light’s Manual: Intertidal Invertebrates of the Central California Coast. Third edition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2585–2586, December 1975. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f75-310.

Bernard:1975:CBM

[1817] F. R. Bernard. Culture of bivalve molluscs: 50 years experience at Conwy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2586, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-311.

Westrheim:1975:OSY

[1818] S. J. Westrheim. Optimum sustainable yield as a concept in fisheries management. (proceedings of a symposium held during the 104th Annual Meeting of the American Fisheries Society, Honolulu, Hawaii, September 9, 1974). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2586, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f75-312.

Anonymous:1975:Ob

[1819] Anonymous. Obituary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 32(12):2589, December 1975. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f75-313.

Hilborn:1976:OEM

[1820] R. Hilborn. Optimal exploitation of multiple stocks by a common fishery: A new methodology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1): 1–5, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f76-001.

OConnor:1976:PBT

[1821] J. F. O’Connor and G. Power. Production by brook trout (Salvelinus fontinalis) in four streams in the Matamek Watershed, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):6–18, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-002.

Ginetz:1976:FAR

[1822] R. M. Ginetz and P. A. Larkin. Factors affecting rainbow trout (Salmo gairdneri) predation on migrant fry of sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):19–24, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-003.

Tabachek:1976:IIB

[1823] Jo-Anne L. Tabachek and M. Yurkowski. Isolation and identification of blue-green algae producing muddy odor metabolites, geosmin, and 2-methylisoborneol, in saline lakes in Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):25–35, January 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-004.

Healey:1976:PCD

[1824] F. P. Healey and L. L. Hendzel. Physiological changes during the course of blooms of Aphanizomenon flos-aquae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):36–41, January 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-005.

Kelso:1976:MYP

[1825] J. R. M. Kelso. Movement of yellow perch (Perca flavescens) and white sucker (Catostomus commersoni) in a nearshore Great Lakes habitat subject to a thermal discharge. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):42–53, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-006.

Giles:1976:COS

[1826] M. A. Giles and W. E. Vanstone. Changes in ouabain-sensitive adenosine triphosphatase activity in gills of Coho salmon (Oncorhynchus kisutch) during Parr–smolt transformation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):54–62, January 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-007.

Tyler:1976:RGM

[1827] A. V. Tyler and R. S. Dunn. Ration, growth, and measures of somatic and organ condition in relation to meal frequency in winter flounder, Pseudopleuronectes americanus, with hypotheses regarding population homeostasis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):63–75, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-008.

Schreck:1976:PRR

[1828] Carl B. Schreck, Roy A. Whaley, Michael L. Bass, O. Eugene Maughan, and Mario Solazzi. Physiological responses of rainbow trout (Salmo gairdneri) to electroshock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1): 76–84, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-009.

Wong:1976:FDC

[1829] S. L. Wong and B. Clark. Field determination of the critical nutrient concentrations for Cladophora in streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):85–92, January 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-010.

Pocklington:1976:TOM

[1830] Roger Pocklington. Terrigenous organic matter in surface sediments from the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):93–97, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-011.

Sutcliffe:1976:CCP

[1831] W. H. Sutcliffe, Jr., R. H. Loucks, and K. F. Drinkwater. Coastal circulation and physical oceanography of the Scotian Shelf and the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):98–115, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-012.

Garrett:1976:UAY

[1832] C. J. R. Garrett and H. Loucks. Upwelling along the yarmouth shore of Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1): 116–117, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-013.

Menendez:1976:CER

[1833] Raymond Menendez. Chronic effects of reduced pH on brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):118–123, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-014.

El-Sabh:1976:SCP

[1834] Mohammed I. El-Sabh. Surface circulation pattern in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):124–138, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-015.

Lester:1976:ECW

[1835] R. J. G. Lester and B. A. Daniels. The eosinophilic cell of the white sucker, Catostomus commersoni. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):139–144, January 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-016.

Walters:1976:ACF

[1836] Carl J. Walters and Ray Hilborn. Adaptive control of fishing systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):145–159, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-017.

Iredale:1976:PME

[1837] D. G. Iredale and R. K. York. Purging a muddy–earthy flavor Taint from rainbow trout (Salmo gairdneri) by transferring to artificial and natural holding environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):160–166, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-018.

Zeitoun:1976:QNR

[1838] Ibrahim H. Zeitoun, Duane E. Ullrey, William T. Magee, John L. Gill, and Werner G. Bergen. Quantifying nutrient requirements of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):167–172, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-019.

Driedzic:1976:BLL

[1839] William R. Driedzic and Joe W. Kiceniuk. Blood lactate levels in freeswimming rainbow trout (Salmo gairdneri) before and after strenuous exercise resulting in fatigue. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):173–176, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-020.

Edsall:1976:TTY

[1840] Thomas A. Edsall and Donald V. Rottiers. Temperature tolerance of young-of-the-year lake whitefish, Coregonus clupeaformis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):177–180, January 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-021.

Symons:1976:ULP

[1841] P. E. K. Symons, Janice L. Metcalfe, and G. D. Harding. Upper lethal and preferred temperatures of the slimy sculpin, Cottus cognatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):180–183, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-022.

Spieler:1976:STS

[1842] Richard E. Spieler and Albert H. Meier. Short-term serum prolactin concentrations in goldfish (Carassius auratus) subjected to serial sampling and restraint. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (1):183–186, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-023.

Grischkowsky:1976:IHN

[1843] Roger S. Grischkowsky and Donald F. Amend. Infectious hematopoietic necrosis virus: Prevalence in certain Alaskan sockeye salmon, Oncorhynchus nerka. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (1):186–188, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-024.

Gatto:1976:MVV

[1844] M. Gatto and S. Rinaldi. Mean value and variability of fish catches in fluctuating environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):189–193, January 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-025.

Ware:1976:MEF

[1845] D. M. Ware. Marine ecology and fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):194, January 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-026.

Raleigh:1976:PFS

[1846] Robert F. Raleigh. Principles of fishery science. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):194–195, January 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-027.

Milliman:1976:MGI

[1847] John D. Milliman. Marine Geotechnology. An international journal of seafloor science and engineering. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(1):195, January 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-028.

Goddard:1976:PTH

[1848] C. I. Goddard and J. S. Tait. Preferred temperatures of F_{3} to F_{5} hybrids of Salvelinus fontinalis × S. namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):197–202, February 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-029.

Adelman:1976:ATS

[1849] Ira R. Adelman, Lloyd L. Smith, Jr., and Gary D. Siesennop. Acute toxicity of sodium chloride, pentachlorophenol, Guthion(R), and hexavalent chromium to fathead minnows (Pimephales promelas) and goldfish (Carassius auratus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2): 203–208, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-030.

Adelman:1976:FMP

[1850] Ira R. Adelman and Lloyd L. Smith, Jr. Fathead minnows (Pimephales promelas) and goldfish (Carassius auratus) as standard fish in bioassays and their reaction to potential reference toxicants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):209–214, February 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-031.

Collins:1976:EAA

[1851] M. T. Collins, J. B. Gratzek, D. L. Dawe, and T. G. Nemetz. Effects of antibacterial agents on nitrification in an aquatic recirculating system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):215–218, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-032.

Nalewajko:1976:PPE

[1852] C. Nalewajko and D. W. Schindler. Primary production, extracellular release, and heterotrophy in two lakes in the ELA, Northwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):219–226, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-033.

Welch:1976:ECD

[1853] H. E. Welch. Ecology of Chironomidae (Diptera) in a polar lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):227–247, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-034.

Norstrom:1976:BBM

[1854] R. J. Norstrom, A. E. McKinnon, and A. S. W. deFreitas. A bioenergetics-based model for pollutant accumulation by fish. simulation of PCB and methylmercury residue levels in Ottawa River yellow perch (Perca flavescens). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2): 248–267, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-035.

Hodson:1976:ALA

[1855] Peter V. Hodson. δ-amino levulinic acid dehydratase activity of fish blood as an indicator of a harmful exposure to lead. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):268–271, February 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-036.

Mudry:1976:MPA

[1856] Dwight R. Mudry and Peter J. McCart. Metazoan parasites of Arctic char (Salvelinus alpinus) from the North slope of Canada and Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):271–275, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-037.

Jensen:1976:RBY

[1857] A. L. Jensen. Relation between yield and fishing mortality for Ricker’s yield equation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2): 275–277, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-038.

Day:1976:HBP

[1858] Dwane E. Day. Homing behavior and population stratification in Central Puget Sound English sole (Parophrys vetulus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):278–282, February 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-039.

MacCrimmon:1976:RRE

[1859] Hugh R. MacCrimmon and Christopher M. Hawkins. Role of rearing experience in gravel substrate selection by juvenile rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):282–284, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-040.

Iwama:1976:CSH

[1860] George K. Iwama, G. L. Greer, and P. A. Larkin. Changes in some hematological characteristics of Coho salmon (Oncorhynchus kisutch) in response to acute exposure to dehydroabietic acid (DHAA) at different exercise levels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2): 285–289, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-041.

Scherer:1976:OLI

[1861] E. Scherer. Overhead-light intensity and vertical positioning of the walleye, Stizostedion vitreum vitreum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):289–292, February 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-042.

Beitinger:1976:LTR

[1862] Thomas L. Beitinger and John J. Magnuson. Low thermal responsiveness in the bluegill, Lepomis macrochirus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):293–295, February 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-043.

Kerr:1976:NNS

[1863] S. R. Kerr and A. H. Lawrie. Natura Naturans: A Symposium on the Fry Paradigm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2): 296–299, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-044.

Alderdice:1976:SCD

[1864] D. F. Alderdice. Some concepts and descriptions of physiological tolerance: Rate–temperature curves of poikilotherms as transects of response surfaces. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):299–307, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-045.

Brett:1976:SMG

[1865] J. R. Brett. Scope for metabolism and growth of sockeye salmon, Oncorhynchus nerka, and some related energetics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):307–313, February 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-046.

Dickie:1976:PYE

[1866] L. M. Dickie. Predation, yield, and ecological efficiency in aquatic food chains. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):313–316, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-047.

Ihssen:1976:SBH

[1867] Peter Ihssen. Selective breeding and hybridization in fisheries management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):316–321, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-048.

Loftus:1976:NAF

[1868] K. H. Loftus. A new approach to fisheries management and F. E. J. Fry’s role in its development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):321–325, February 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-049.

Sprules:1976:ISL

[1869] W. M. Sprules. Initiation of sea lamprey (Petromyzon marinus) control. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):326–329, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-050.

Kerr:1976:EAF

[1870] S. R. Kerr. Ecological analysis and the fry paradigm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):329–335, February 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-051.

Anonymous:1976:SES

[1871] Anonymous. Selected excerpts from symposium discussion periods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):336–339, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-052.

Anonymous:1976:A

[1872] Anonymous. Acknowledgments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):339, February 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-053.

Anonymous:1976:PFJ

[1873] Anonymous. Profile of F. E. J. Fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):340–344, February 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-054.

Anonymous:1976:LP

[1874] Anonymous. List of participants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):344–345, February 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-055.

Waldichuk:1976:MSC

[1875] M. Waldichuk. Marine science communications — basic and applied. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(2):346, February 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-056.

Anonymous:1976:EFa

[1876] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):349–350, March 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-057.

Anonymous:1976:Pa

[1877] Anonymous. Preface. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3): 351–354, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-058.

Sly:1976:LEB

[1878] P. G. Sly. Lake Erie and its basin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):355–370, March 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-059.

Simons:1976:CDC

[1879] T. J. Simons. Continuous dynamical computations of water transports in Lake Erie for 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (3):371–384, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-060.

Thomas:1976:SSL

[1880] R. L. Thomas, J.-M. Jaquet, A. L. W. Kemp, and C. F. M. Lewis. Surficial sediments of Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):385–403, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-061.

Thomas:1976:MSS

[1881] R. L. Thomas and J.-M. Jaquet. Mercury in the surficial sediments of Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3): 404–412, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-062.

Williams:1976:FPS

[1882] J. D. H. Williams, J-M. Jaquet, and R. L. Thomas. Forms of phosphorus in the surficial sediments of Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):413–429, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-063.

Williams:1976:RAP

[1883] J. D. H. Williams, T. P. Murphy, and T. Mayer. Rates of accumulation of phosphorus forms in Lake Erie sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):430–439, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-064.

Kemp:1976:CIG

[1884] A. L. W. Kemp, R. L. Thomas, C. I. Dell, and J.-M. Jaquet. Cultural impact on the geochemistry of sediments in Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):440–462, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-065.

Burns:1976:FIM

[1885] N. M. Burns and J. O. Nriagu. Forms of iron and manganese in Lake Erie waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):463– 470, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-066.

Ongley:1976:SYN

[1886] Edwin D. Ongley. Sediment yields and nutrient loadings from Canadian watersheds tributary to Lake Erie: an overview. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):471–484, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-067.

Burns:1976:TON

[1887] N. M. Burns. Temperature, oxygen, and nutrient distribution patterns in Lake Erie, 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (3):485–511, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-068.

Burns:1976:ODC

[1888] N. M. Burns. Oxygen depletion in the Central and Eastern Basins of Lake Erie, 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (3):512–519, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-069.

Burns:1976:NBL

[1889] N. M. Burns. Nutrient budgets for Lake Erie, 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):520–536, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-070.

Lam:1976:NCA

[1890] D. C. L. Lam and T. J. Simons. Numerical computations of advective and diffusive transports of chloride in Lake Erie, 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):537–549, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-071.

Lam:1976:CPT

[1891] D. C. L. Lam and J.-M. Jaquet. Computations of physical transport and regeneration of phosphorus in Lake Erie, fall 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):550–563, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-072.

Burns:1976:PBL

[1892] N. M. Burns, J. D. H. Williams, J.-M. Jaquet, A. L. W. Kemp, and D. C. L. Lam. A phosphorus budget for Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):564–573, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-073.

Rao:1976:BDL

[1893] S. S. Rao and B. K. Burnison. Bacterial distributions in Lake Erie (1967, 1970). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):574–580, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-074.

Munawar:1976:LSP

[1894] Mohiuddin Munawar and Iftekhar F. Munawar. A lakewide study of phytoplankton biomass and its species composition in Lake Erie, April– December 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (3):581–600, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-075.

Munawar:1976:RPB

[1895] M. Munawar and N. M. Burns. Relationships of phytoplankton biomass with soluble nutrients, primary production, and chlorophyll a in Lake Erie, 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):601– 611, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-076.

Watson:1976:SDA

[1896] N. H. F. Watson. Seasonal distribution and abundance of crustacean zooplankton in Lake Erie, 1970. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):612–621, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-077.

Leach:1976:FCL

[1897] J. H. Leach and S. J. Nepszy. The fish community in Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):622–638, March 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-078.

Burns:1976:PWL

[1898] N. M. Burns, J.-M. Jaquet, A. L. W. Kemp, D. C. L. Lam, J. H. Leach, M. Munawar, T. J. Simons, P. G. Sly, R. L. Thomas, N. H. F. Watson, and J. D. H. Williams. Processes within Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(3):639–643, March 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-079.

Ayles:1976:SMR

[1899] G. Burton Ayles, J. G. I. Lark, J. Barica, and Hedy Kling. Seasonal mortality of rainbow trout (Salmo gairdneri) planted in small eutrophic lakes of Central Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):647–655, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-080.

Wenner:1976:ABM

[1900] Charles A. Wenner. Aspects of the biology and morphology of the snake eel, Pisodonophis cruentifer (Pisces, Ophichthidaeo). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):656–665, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-081.

Pedersen:1976:HBE

[1901] Mark G. Pedersen, William K. Hershberger, Prince K. Zachariah, and Mont R. Juchau. Hepatic biotransformation of environmental xenobiotics in six strains of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):666–675, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-082.

Kuchnow:1976:TTS

[1902] Karl P. Kuchnow and Robert S. Foster IV. Thermal tolerance of stored Fundulus heteroclitus gametes: Fertilizability and survival of embryos. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):676–680, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-083.

Sameoto:1976:DSS

[1903] D. D. Sameoto. Distribution of sound scattering layers caused by euphausiids and their relationship to chlorophyll a concentrations in the Gulf of St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):681–687, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-084.

Cooper:1976:ESM

[1904] Jon C. Cooper and Arthur D. Hasler. Electrophysiological studies of morpholine-imprinted Coho salmon (Oncorhynchus kisutch) and rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):688–694, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-085.

Fernet:1976:ABC

[1905] D. A. Fernet and R. J. F. Smith. Agonistic behavior of captive goldeye (Hiodon alosoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 695–702, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-086.

Cooper:1976:ECO

[1906] Jon C. Cooper, Allan T. Scholz, Ross M. Horrall, Arthur D. Hasler, and Dale M. Madison. Experimental confirmation of the olfactory hypothesis with homing, artificially imprinted Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):703–710, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-087.

Templeman:1976:ILC

[1907] Wilfred Templeman, V. M. Hodder, and A. M. Fleming. Infection of lumpfish (Cyclopterus lumpus) with larvae and of Atlantic cod (Gadus morhua) with adults of the copepod, Lernaeocera branchialis, in and adjacent to the Newfoundland area, and inferences therefrom on inshore– offshore migrations of cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):711–731, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-088.

Shotts:1976:BFA

[1908] E. B. Shotts, Jr., A. L. Kleckner, J. B. Gratzek, and J. L. Blue. Bacterial flora of aquarium fishes and their shipping waters imported from Southeast Asia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 732–735, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-089.

Shotts:1976:OFA

[1909] E. B. Shotts, Jr., V. L. Vanderwork, and L. M. Campbell. Occurrence of R factors associated with Aeromonas hydrophila isolates from aquarium fish and waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 736–740, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-090.

Davis:1976:SEP

[1910] Kenneth B. Davis and Bill A. Simco. Salinity effects on plasma electrolytes of channel catfish, Ictalurus punctatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):741–746, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-091.

Jensen:1976:AUS

[1911] A. L. Jensen. Assessment of the United States lake whitefish (Coregonus clupeaformis) fisheries of Lake Superior, Lake Michigan, and Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):747–759, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-092.

Bailey:1976:NGL

[1912] G. S. Bailey, H. Tsuyuki, and A. C. Wilson. The number of genes for lactate dehydrogenase in salmonid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):760–767, April 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-093.

Repsys:1976:FFS

[1913] Andrew J. Repsys, Richard L. Applegate, and Donald C. Hales. Food and food selectivity of the black bullhead, Ictalurus melas, in Lake Poinsett, South Dakota. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 768–775, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-094.

Buckley:1976:CBC

[1914] James A. Buckley, Cecil M. Whitmore, and Robert I. Matsuda. Changes in blood chemistry and blood cell morphology in Coho salmon (Oncorhynchus kisutch) following exposure to sublethal levels of total residual chlorine in municipal wastewater. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):776–782, April 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-095.

Forney:1976:YCF

[1915] J. L. Forney. Year-class formation in the walleye (Stizostedion vitreum vitreum) population of Oneida Lake, New York, 1966–73. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):783–792, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-096.

Huang:1976:OCS

[1916] C. C. Huang, Ilan B. Vertinsky, and Norman J. Wilimovsky. Optimal controls for a single species fishery and the economic value of research. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):793–809, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-097.

Prouse:1976:ELC

[1917] Nicholas J. Prouse, Donald C. Gordon, Jr., and Paul D. Keizer. Effects of low concentrations of oil accommodated in sea water on the growth of unialgal marine phytoplankton cultures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):810–818, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-098.

Tamm:1976:DER

[1918] George R. Tamm and J. Stanley Cobb. Diel ecdysis rhythms in juvenile lobsters Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):819–821, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-099.

Marinenko:1976:IIC

[1919] George Marinenko, Robert J. Huggett, and Dale G. Friend. An instrument with internal calibration for monitoring chlorine residuals in natural waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):

822–826, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-100.

Cooper:1976:HAI

[1920] Jon C. Cooper and Allan T. Scholz. Homing of artificially imprinted steelhead (rainbow) trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):826–829, April 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-101.

Jessop:1976:DTT

[1921] B. M. Jessop. Distribution and timing of tag recoveries from native and nonnative Atlantic salmon (Salmo salar) released into Big Salmon River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 829–833, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-102.

Hughes:1976:SSL

[1922] S. E. Hughes. System for sampling large trawl catches of research vessels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):833–839, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-103.

McCormick:1976:GSY

[1923] J. Howard McCormick and Charles F. Kleiner. Growth and survival of young-of-the-year emerald shiners (Notropis atherinoides) at different temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 839–842, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-104.

Richards:1976:MSS

[1924] Sarah W. Richards. Mixed species schooling of postlarvae of Ammodytes hexapterus and Clupea harengus harengus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):843–844, April 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-105.

Braaten:1976:PST

[1925] Bruce A. Braaten and Harold O. Hodgins. Protection of steelhead trout (Salmo gairdneri) against vibriosis with a living low-virulence strain of Vibrio anguillarum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 845–847, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-106.

Gibson:1976:DSP

[1926] Victoria R. Gibson and George D. Grice. Developmental stages of Pontella meadi Wheeler (Copepoda: Calanoida). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):847–854, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-107.

Pritchard:1976:SAD

[1927] G. I. Pritchard. Structured aquaculture development with a Canadian perspective. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):855– 870, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-108.

Marr:1976:BIO

[1928] John C. Marr. The biology of the Indian Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):871, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-109.

Leibovitz:1976:PF

[1929] Louis Leibovitz. The pathology of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):872, April 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-110.

Anonymous:1976:Pb

[1930] Anonymous. Preface. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 875–878, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-111.

Bardach:1976:AR

[1931] John E. Bardach. Aquaculture revisited. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):880–887, April 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-112.

Milne:1976:EEA

[1932] P. H. Milne. Engineering and the economics of aquaculture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):888–898, April 1976. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f76-113.

Sribhibhadh:1976:RAE

[1933] Arporna Sribhibhadh. Role of aquaculture in economic development within Southeast Asia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):899, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-114.

Joyner:1976:FOR

[1934] Timothy Joyner. Farming ocean ranges for salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):902–904, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-115.

Jhingran:1976:SPF

[1935] V. G. Jhingran. Systems of polyculture of fishes in the inland waters of India. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):905– 910, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-116.

Fujiya:1976:YSQ

[1936] Masaru Fujiya. Yellowtail (Seriola quinqueradiata) farming in Japan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):911–915, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-117.

Kuronuma:1976:SPP

[1937] Katsuzo Kuronuma. Status and potential of pond culture of penaeid prawns in Southeast Asia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):916, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-118.

Chen:1976:CMC

[1938] T. P. Chen. Culture of milkfish (Chanos chanos) as a means of increasing animal protein supply. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):917–919, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-119.

Fontaine:1976:HCR

[1939] M. Fontaine. Hormones and the control of reproduction in aquaculture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):922–939, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-120.

Chaudhuri:1976:UHI

[1940] H. Chaudhuri. Use of hormones in induced spawning of carps. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):940–947, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-121.

Yamazaki:1976:AHF

[1941] Fumio Yamazaki. Application of hormones in fish culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):948–958, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-122.

Sundararaj:1976:PTC

[1942] Bangalore I. Sundararaj and Sudhir Vasal. Photoperiod and temperature control in the regulation of reproduction in the female catfish Heteropneustes fossilis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 959–973, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-123.

Jalabert:1976:VOM

[1943] Bernard Jalabert. In vitro oocyte maturation and ovulation in rainbow trout (Salmo gairdneri), northern pike (Esox lucius), and goldfish (Carassius auratus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4): 974–988, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-124.

Hirose:1976:ECO

[1944] Keiji Hirose. Endocrine control of ovulation in medaka (Oryzias latipes) and ayu (Plecoglossus altivelis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):989–994, April 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-125.

Horton:1976:CFS

[1945] Howard F. Horton and Alvin G. Ott. Cryopreservation of fish spermatozoa and ova. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (4):995–1000, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-126.

Hasegawa:1976:PLC

[1946] Yoshio Hasegawa. Progress of Laminaria cultivation in Japan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1002–1006, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-127.

Neish:1976:RMC

[1947] Iain C. Neish. Role of mariculture in the Canadian seaweed industry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1007–1014, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-128.

North:1976:ATC

[1948] Wheeler J. North. Aquacultural techniques for creating and restoring beds of giant kelp, Macrocystis spp. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1015–1023, April 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-129.

Yamada:1976:CSF

[1949] Nobuo Yamada. Current status and future prospects for harvesting and resource management of the agarophyte in Japan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1024–1030, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-130.

Halver:1976:FPD

[1950] J. E. Halver. Formulating practical diets for fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1032–1039, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-131.

Cowey:1976:USD

[1951] C. B. Cowey. Use of synthetic diets and biochemical criteria in the assessment of nutrient requirements of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1040–1045, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-132.

Weatherley:1976:FAM

[1952] A. H. Weatherley. Factors affecting maximization of fish growth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1046–1058, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-133.

Amend:1976:PCV

[1953] Donald F. Amend. Prevention and control of viral diseases of salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1059–1066, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-134.

Calaprice:1976:MEG

[1954] J. R. Calaprice. Mariculture — ecological and genetic aspects of production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1068– 1087, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-135.

Purdom:1976:GTF

[1955] C. E. Purdom. Genetic techniques in flatfish culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1088–1093, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-136.

Gjedrem:1976:PGI

[1956] Trygve Gjedrem. Possibilities for genetic improvements in salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1094–1099, April 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-137.

Longwell:1976:RGR

[1957] A. Crosby Longwell. Review of genetic and related studies on commercial oysters and other pelecypod mollusks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1100–1107, April 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-138.

Hedgecock:1976:ABG

[1958] Dennis Hedgecock, Robert A. Shleser, and Keith Nelson. Applications of biochemical genetics to aquaculture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(4):1108–1119, April 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-139.

Schultz:1976:AUD

[1959] Donald P. Schultz and Paul D. Harman. Antimycin: Uptake, distribution, and elimination in brown bullheads (Ictalurus nebulosus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1121–1129, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-140.

Koenst:1976:TRE

[1960] Walter M. Koenst and Lloyd L. Smith, Jr. Thermal requirements of the early life history stages of walleye, Stizostedion vitreum vitreum, and sauger, Stizostedion canadense. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1130–1138, May 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-141.

Hare:1976:PPB

[1961] G. M. Hare and M. D. B. Burt. Parasites as potential biological tags of Atlantic salmon (Salmo salar) smolts in the Miramichi River System, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (5):1139–1143, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-142.

Giles:1976:OVM

[1962] M. A. Giles and W. E. Vanstone. Ontogenetic variation in the multiple hemoglobins of Coho salmon (Oncorhynchus kisutch) and effect of environmental factors on their expression. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1144–1149, May 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-143.

Stauffer:1976:FCS

[1963] Thomas M. Stauffer. Fecundity of Coho salmon (Oncorhynchus kisutch) from the Great Lakes and a comparison with ocean salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1150–1155, May 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-144.

Bishop:1976:LCS

[1964] David G. Bishop, David G. James, and June Olley. Lipid composition of slender tuna (Allothunnus fallai) as related to lipid composition of their feed (Nyctiphanes australis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1156–1161, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-145.

Elwood:1976:TMP

[1965] Jerry W. Elwood and L. Dean Eyman. Test of a model for predicting the body burden of trace contaminants in aquatic consumers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1162–1166, May 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-146.

Faber:1976:HFL

[1966] Daniel J. Faber. Hyponeustonic fish larvae in the Northumberland Strait during summer 1962. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5): 1167–1174, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-147.

Pomeroy:1976:EED

[1967] W. M. Pomeroy and J. G. Stockner. Effects of environmental disturbance on the distribution and primary production of benthic algae on a British Columbia estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (5):1175–1187, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-148.

Buchanan:1976:ATS

[1968] David V. Buchanan, Paul S. Tate, and John R. Moring. Acute toxicities of spruce and hemlock bark extracts to some estuarine organisms in Southeastern Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5): 1188–1192, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-149.

Myers:1976:SGR

[1969] Gary L. Myers and John J. Peterka. Survival and growth of rainbow trout (Salmo gairdneri) in four prairie lakes, North Dakota. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1192–1195, May 1976. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f76-150.

Westrheim:1976:FNS

[1970] S. J. Westrheim. A further note on sampling research trawl catches at sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1196, May 1976. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f76-151.

Bullock:1976:IPN

[1971] G. L. Bullock, R. R. Rucker, D. Amend, K. Wolf, and H. M. Stuckey. Infectious pancreatic necrosis: Transmission with iodine-treated and nontreated eggs of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1197–1198, May 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-152.

Hansen:1976:AHP

[1972] Carl R. Hansen, Jr. and Joseph A. Kawatski. Application of 24-hour postexposure observation to acute toxicity studies with invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1198–1201, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-153.

Lear:1976:MAS

[1973] W. H. Lear. Migrating Atlantic salmon (Salmo salar) caught by otter trawl on the Newfoundland Continental Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1202–1205, May 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-154.

Ord:1976:VHS

[1974] Wilma M. Ord, Monique Le Berre, and Pierre de Kinkelin. Viral hemorrhagic septicemia: Comparative susceptibility of rainbow trout (Salmo gairdneri) and hybrids (S. gairdneri × Oncorhynchus kisutch) to experimental infection. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (5):1205–1208, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-155.

Nebeker:1976:SDC

[1975] Alan V. Nebeker. Survival of Daphnia, crayfish, and stoneflies in airsupersaturated water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5): 1208–1212, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-156.

Regier:1976:SSF

[1976] Henry A. Regier. Science for the scattered fisheries of the Canadian interior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1213– 1232, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-157.

Hoar:1976:STE

[1977] William S. Hoar. Smolt transformation: Evolution, behavior, and physiology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1233– 1252, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-158.

Richards:1976:SCB

[1978] W. J. Richards. Some comments on Balon’s terminology of fish development intervals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (5):1253–1254, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-159.

Balon:1976:NCD

[1979] Eugene K. Balon. A note concerning Dr Richards’ comments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1254–1256, May 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-160.

Evelyn:1976:HTS

[1980] T. P. T. Evelyn. Handbook of trout and salmon diseases. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(5):1257–1258, May 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-161.

deMarch:1976:STP

[1981] Brigitte G. E. de March. Spatial and temporal patterns in macrobenthic stream diversity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (6):1261–1270, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-162.

Sinclair:1976:SNT

[1982] Michael Sinclair, Mohammed El-Sabh, and Jean-Rock Brindle. Seaward nutrient transport in the Lower St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1271–1277, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-163.

Breen:1976:DGK

[1983] P. A. Breen and K. H. Mann. Destructive grazing of kelp by sea urchins in Eastern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (6):1278–1283, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-164.

Servizi:1976:CCA

[1984] J. A. Servizi, R. W. Gordon, I. H. Rogers, and H. W. Mahood. Chemical characteristics and acute toxicity of foam on two aerated lagoons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1284–1290, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-165.

Antipa:1976:FTI

[1985] Ross Antipa. Field testing of injected Vibrio anguillarum bacterins in pen-reared Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1291–1296, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-166.

Mazeaud:1976:FFA

[1986] F. Mazeaud and E. Bilinski. Free fatty acids and the onset of rancidity in rainbow trout (Salmo gairdneri) flesh. Effect of phospholipase A. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1297–1302, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-167.

McLeay:1976:RMM

[1987] Donald J. McLeay. A rapid method for measuring the acute toxicity of pulpmill effluents and other toxicants to salmonid fish at ambient room temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (6):1303–1311, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-168.

Lean:1976:PEO

[1988] D. R. S. Lean and C. Nalewajko. Phosphate exchange and organic phosphorus excretion by freshwater algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1312–1323, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-169.

deJong:1976:MAA

[1989] I. G. de Jong and D. R. Wiles. Molecular activation analysis and its application to methylmercury determination in various marine samples. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1324–1330, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-170.

Klaverkamp:1976:ECD

[1990] J. F. Klaverkamp, W. L. Lockhart, D. Metner, and N. Grift. Effects of chronic DDT/DDE exposure on anesthetic induction and recovery times in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1331–1334, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-171.

Lett:1976:ECS

[1991] Patrick F. Lett, G. J. Farmer, and F. W. H. Beamish. Effect of copper on some aspects of the bioenergetics of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1335–1342, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-172.

Hansen:1976:EDA

[1992] L. G. Hansen, W. B. Wiekhorst, and J. Simon. Effects of dietary Aroclor(R) 1242 on channel catfish (Ictalurus punctatus) and the selective accumulation of PCB components. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1343–1352, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-173.

Lett:1976:RPM

[1993] P. F. Lett and A. C. Kohler. Recruitment: a problem of multispecies interaction and environmental perturbations, with special reference to Gulf of St. Lawrence Atlantic herring (Clupea harengus harengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1353–1371, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-174.

Stanley:1976:FHG

[1994] Jon G. Stanley. Female homogamety in grass carp (Ctenopharyngodon idella) determined by gynogenesis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1372–1374, June 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-175.

Bull:1976:NPR

[1995] C. J. Bull and W. C. Mackay. Nitrogen and phosphorus removal from lakes by fish harvest. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (6):1374–1376, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-176.

Fried:1976:BCA

[1996] Stephen M. Fried, James D. McCleave, and Kristin A. Stred. Buoyancy compensation by Atlantic salmon (Salmo salar) smolts tagged internally with dummy telemetry transmitters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1377–1380, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-177.

Appy:1976:VNP

[1997] R. G. Appy, M. D. B. Burt, and T. J. Morris. Viral nature of piscine erythrocytic necrosis (PEN) in the blood of Atlantic cod (Gadus morhua). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1380–1385, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-178.

Sanders:1976:SCF

[1998] J. E. Sanders, R. A. Holt, and J. L. Fryer. Serological comparison of Flexibacter columnaris isolates using rabbit and rainbow trout (Salmo gairdneri) antisera. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (6):1386–1388, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-179.

Kennedy:1976:ACS

[1999] Victor S. Kennedy. Arsenic concentrations in some coexisting marine organisms from Newfoundland and Labrador. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1388–1393, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-180.

Hodson:1976:TEL

[2000] Peter V. Hodson. Temperature effects on lactate–glycogen metabolism in zinc-intoxicated rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1393–1397, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-181.

Tan:1976:OIS

[2001] F. C. Tan and W. D. Fraser. Oxygen isotope studies on sea ice in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33 (6):1397–1401, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-182.

Souter:1976:EBC

[2002] B. W. Souter, R. A. Sonstegard, and L. A. McDermott. Enteric bacteria in carp (Cyprinus carpio) and white suckers (Catostomus commersoni). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1401–1403, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-183.

Houston:1976:SWM

[2003] A. H. Houston, J. T. Corlett, and R. J. Woods. Specimen weight and M.S. 222. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1403– 1407, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-184.

Sebens:1976:IMS

[2004] Kenneth P. Sebens. Individual marking of soft-bodied intertidal invertebrates in situ: a vital stain technique applied to the sea anemone, Anthopleura xanthogrammica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1407–1410, June 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-185.

Mann:1976:POC

[2005] C. R. Mann. Physical oceanography in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1411–1433, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-186.

Pond:1976:NMO

[2006] S. Pond. Numerical modeling of ocean circulations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1434–1450, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-187.

Mann:1976:COC

[2007] C. R. Mann. Chemical oceanography in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1451–1469, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-188.

Eisler:1976:PPM

[2008] R. Eisler. Pollution and physiology of marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1470–1471, June 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-189.

LeBlond:1976:TSA

[2009] P. H. LeBlond. Temperature–salinity analysis of world ocean waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(6):1471, June 1976. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f76-190.

Ricker:1976:RRG

[2010] W. E. Ricker. Review of the rate of growth and mortality of Pacific salmon in salt water, and noncatch mortality caused by fishing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1483–1524, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-191.

Flick:1976:PWD

[2011] William A. Flick and Dwight A. Webster. Production of wild, domestic, and interstrain hybrids of brook trout (Salvelinus fontinalis) in natural ponds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1525– 1539, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-192.

VanWinkle:1976:ETS

[2012] W. Van Winkle, S. Y. Feng, and H. H. Haskin. Effect of temperature and salinity on extension of siphons by Mercenaria mercenaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1540–1546, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-193.

Koeller:1976:ACN

[2013] Peter A. Koeller and Jack L. Littlepage. Azygokeras columbiae, a new genus and species of marine epibenthic copepod (Calanoida: Aetideidae) from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1547–1552, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-194.

Stockner:1976:APC

[2014] John G. Stockner and K. R. S. Shortreed. Autotrophic production in Carnation Creek, a coastal rainforest stream on Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1553– 1563, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-195.

Williams:1976:NEI

[2015] Ian V. Williams and Donald F. Amend. A natural epizootic of infectious hematopoietic necrosis in fry of sockeye salmon (Oncorhynchus nerka) at Chilko Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1564–1567, July 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-196.

Hara:1976:EMC

[2016] Toshiaki J. Hara, Y. M. C. Law, and S. Macdonald. Effects of mercury and copper on the olfactory response in rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1568–1573, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-197.

Bowering:1976:DAG

[2017] W. R. Bowering. Distribution, age and growth, and sexual maturity of witch flounder (Glyptocephalus cynoglossus) in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1574–1584, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-198.

Higgs:1976:IBG

[2018] David A. Higgs, Edward M. Donaldson, Helen M. Dye, and J. R. McBride. Influence of bovine growth hormone and L-thyroxine on growth, muscle composition, and histological structure of the gonads, thyroid, pancreas, and pituitary of Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1585–1603, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-199.

Wells:1976:ECO

[2019] Peter G. Wells and John B. Sprague. Effects of crude oil on American lobster (Homarus americanus) larvae in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1604–1614, July 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-200.

Sheehy:1976:UAS

[2020] Daniel J. Sheehy. Utilization of artificial shelters by the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1615–1622, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-201.

Couch:1976:EPM

[2021] Ernest F. Couch, Carol S. Daily, and William B. Smith. Evidence for the presence of molt-inhibiting hormone in the sinus gland of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1623–1627, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-202.

Leaman:1976:ITS

[2022] Bruce M. Leaman. An inexpensive tag for short-term visual tracking studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1628– 1629, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-203.

El-Shamy:1976:AGE

[2023] F. M. El-Shamy. Analyses of gastric emptying in bluegill (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7): 1630–1633, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-204.

Ackman:1976:VFL

[2024] R. G. Ackman and C. A. Eaton. Variations in fillet lipid content and some percent lipid-iodine value relationships for large winter Atlantic herring (Clupea harengus harengus) from Southeastern Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1634–1638, July 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-205.

Craig:1976:LHN

[2025] P. C. Craig and J. Wells. Life history notes for a population of slimy sculpin (Cottus cognatus) in an Alaskan Arctic stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1639–1642, July 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-206.

Richards:1976:OWL

[2026] B. R. Richards and C. I. Belmore. Occurrence of the woodborer, Limnoria lignorum (Rathke) at Amchitka Island, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1642–1644, July 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-207.

Miller:1976:GPM

[2027] R. J. Miller and J. Watson. Growth per molt and limb regeneration in the spider crab, Chionoecetes opilio. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1644–1649, July 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-208.

Bergersen:1976:IFM

[2028] Eric P. Bergersen and Thomas J. Keefe. An index of fish movement with statistical techniques for its interpretation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1649–1652, July 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-209.

Mackie:1976:BNE

[2029] G. L. Mackie and D. G. Huggins. Biological notes on Eupera cubensis (Bivalvia: Sphaeriidae) from Kansas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1652–1656, July 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-210.

Anonymous:1976:BRCa

[2030] Anonymous. Book review/comptes rendus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(7):1657–1658, July 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-211.

Lubinsky:1976:TDS

[2031] Irene Lubinsky. Thyasira dunbari n. sp. (Bivalvia: Thyasiridae) from the Canadian Arctic Archipelago. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1667–1670, August 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-212.

Woodward:1976:THD

[2032] D. F. Woodward. Toxicity of the herbicides dinoseb and picloram to cutthroat (Salmo clarki) and lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1671–1676, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-213.

Mathews:1976:MMP

[2033] Stephen B. Mathews and Raymond Buckley. Marine mortality of Puget Sound Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1677–1684, August 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-214.

Winner:1976:ACT

[2034] Robert W. Winner and Michael P. Farrell. Acute and chronic toxicity of copper to four species of Daphnia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1685–1691, August 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-215.

Hargrave:1976:DRB

[2035] Barry T. Hargrave and Georgina A. Phillips. DDT residues in benthic invertebrates and demersal fish in St. Margaret’s Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1692–1698, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-216.

Khan:1976:TBS

[2036] R. A. Khan and M. C. Meyer. Taxonomy and biology of some Newfoundland marine leeches (Rhynchobdellae: Piscicolidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1699–1714, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-217.

Hobson:1976:TEA

[2037] Louis A. Hobson, W. James Morris, and Kathleen T. Pirquet. Theoretical and experimental analysis of the ^{14}C technique and its use in studies of primary production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8): 1715–1721, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-218.

Peterson:1976:TSJ

[2038] R. H. Peterson. Temperature selection of juvenile Atlantic salmon (Salmo salar) as influenced by various toxic substances. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1722–1730, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-219.

Holcombe:1976:LTE

[2039] G. W. Holcombe, D. A. Benoit, E. N. Leonard, and J. M. McKim. Longterm effects of lead exposure on three generations of brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8): 1731–1741, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-220.

Larsen:1976:PRC

[2040] D. P. Larsen and H. T. Mercier. Phosphorus retention capacity of lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1742–1750, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-221.

Winters:1976:RMS

[2041] G. H. Winters. Recruitment mechanisms of Southern Gulf of St. Lawrence Atlantic herring (Clupea harengus harengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1751–1763, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-222.

Gabriel:1976:OCF

[2042] W. L. Gabriel, W. C. Leggett, J. E. Carscadden, and B. D. Glebe. Origin and characteristics of “fall-run” American shad (Alosa sapidissima) from the St. John River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1764–1770, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-223.

Spinelli:1976:EDC

[2043] John Spinelli and Conrad Mahnken. Effect of diets containing dogfish (Squalus acanthias) meal on the mercury content and growth of pen-reared Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1771–1778, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-224.

Marchesseault:1976:DRM

[2044] Guy D. Marchesseault, Saul B. Saila, and William J. Palm. Delayed recruitment models and their application to the American lobster (Homarus americanus) fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1779–1787, August 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-225.

Hourston:1976:SDD

[2045] A. S. Hourston and H. Rosenthal. Sperm density during active spawning of Pacific herring (Clupea harengus pallasi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1788–1790, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-226.

Poston:1976:RET

[2046] H. A. Poston. Relative effect of two dietary water-soluble analogues of menaquinone on coagulation and packed cell volume of blood of lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1791–1793, August 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-227.

Fraser:1976:AWP

[2047] J. M. Fraser. Assessment of winter planting of hatchery-reared brook trout (Salvelinus fontinalis) in small Precambrian lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1794–1797, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-228.

Faber:1976:IFN

[2048] Daniel J. Faber. Identification of four northern blennioid fish larvae in the Canadian Atlantic Ocean (Stichaeidae, Lumpenidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1798–1802, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-229.

Poston:1976:OLD

[2049] H. A. Poston. Optimum level of dietary biotin for growth, feed utilization, and swimming stamina of fingerling lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1803–1806, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-230.

Bisson:1976:IID

[2050] Peter A. Bisson. Increased invertebrate drift in an experimental stream caused by electrofishing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1806–1808, August 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-231.

Welch:1976:FAW

[2051] H. E. Welch, P. J. Dillon, and A. Sreedharan. Factors affecting winter respiration in Ontario lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1809–1815, August 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-232.

Muller:1976:PTF

[2052] Rudolf Müller and F. E. J. Fry. Preferred temperature of fish: a new method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1815–1817, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-233.

Wobeser:1976:PBD

[2053] G. Wobeser, L. F. Kratt, R. J. F. Smith, and G. Acompanado. Proliferative branchiitis due to Tetraonchus rauschi (Trematoda: Monogenea) in captive Arctic grayling (Thymallus arcticus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1817–1821, August 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-234.

Loftus:1976:SCF

[2054] K. H. Loftus. Science for Canada’s fisheries rehabilitation needs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1822–1857, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f76-235.

Anonymous:1976:LLa

[2055] Anonymous. Letters/lettres. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1858–1864, August 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-236.

Anonymous:1976:BRCb

[2056] Anonymous. Book Reviews/ comptes rendus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8):1865–1866, August 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-237.

Anonymous:1976:CRP

[2057] Anonymous. Corrections to recent publications… corrections sur récentes publications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(8): 1867–1868, August 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-238.

Anonymous:1976:AAa

[2058] Anonymous. Announcement/Avis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1879, September 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-239.

Luedtke:1976:ESS

[2059] R. J. Luedtke and M. A. Brusven. Effects of sand sedimentation on colonization of stream insects. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1881–1886, September 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-240.

Clady:1976:ITW

[2060] Michael D. Clady. Influence of temperature and wind on the survival of early stages of yellow perch, Perca flavescens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1887–1893, September 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-241.

Berkes:1976:EEG

[2061] F. Berkes. Ecology of euphausiids in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1894–1905, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-242.

Gale:1976:SAS

[2062] William F. Gale and Cynthia A. Gale. Selection of artificial spawning sites by the spotfin shiner (Notropis spilopterus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1906–1913, September 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-243.

Twongo:1976:STI

[2063] Timothy K. Twongo and Hugh R. MacCrimmon. Significance of the timing of initial feeding in hatchery rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1914–1921, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-244.

Mauviot:1976:MGE

[2064] Jean C. Mauviot and John D. Castell. Molt- and growth-enhancing effects of bilateral eyestalk ablation on juvenile and adult American lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1922–1929, September 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-245.

Elwood:1976:CGC

[2065] Jerry W. Elwood, Stephen G. Hildebrand, and John J. Beauchamp. Contribution of gut contents to the concentration and body burden of elements in Tipula spp. from a spring-fed stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1930–1938, September 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-246.

Spehar:1976:CZT

[2066] R. L. Spehar. Cadmium and zinc toxicity to flagfish, Jordanella floridae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1939–1945, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-247.

Swenson:1976:IFC

[2067] William A. Swenson and Lloyd L. Smith, Jr. Influence of food competition, predation, and cannibalism on walleye (Stizostedion vitreum vitreum) and sauger (S. canadense) populations in Lake of the Woods, Minnesota. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1946–1954, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-248.

Rosenberg:1976:CSR

[2068] D. M. Rosenberg and A. P. Wiens. Community and species responses of Chironomidae (Diptera) to contamination of fresh waters by crude oil and petroleum products, with special reference to the Trail River, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1955–1963, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-249.

Eggers:1976:TES

[2069] Douglas M. Eggers. Theoretical effect of schooling by planktivorous fish predators on rate of prey consumption. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1964–1971, September 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-250.

Janssen:1976:FMP

[2070] John Janssen. Feeding modes and prey size selection in the alewife (Alosa pseudoharengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):

1972–1975, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-251.

Geraci:1976:DIE

[2071] Joseph R. Geraci and Thomas G. Smith. Direct and indirect effects of oil on ringed seals (Phoca hispida) of the Beaufort Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1976–1984, September 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-252.

Anonymous:1976:PSP

[2072] Anonymous. Pollution Symposium — 13th Pacific Science Congress Sublethal Effects of Pollution on Aquatic Organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):1985–2065, September 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-253.

Levings:1976:BTS

[2073] C. D. Levings. Basket traps for surveys of a gammarid amphipod, Anisogammarus confervicolus (Stimpson), at two British Columbia estuaries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2066–2069, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-254.

Kelso:1976:DMW

[2074] J. R. M. Kelso. Diel movement of walleye, Stizostedion vitreum vitreum, in West Blue Lake, Manitoba, as determined by ultrasonic tracking. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2070–2072, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-255.

Addison:1976:MSC

[2075] R. F. Addison, M. E. Zinck, and J. R. Leahy. Metabolism of single and combined doses of ^{14}C-Aldrin and ^{3}H-p,p’-DDT by Atlantic salmon (Salmo salar) fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9): 2073–2076, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-256.

Cunningham:1976:IRD

[2076] C. C. Cunningham, D. D. W. Conrad, J. D. Moffatt, and E. M. Levy. Improved reagent dispenser for the determination of dissolved oxygen in sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2076–2078, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-257.

Glenn:1976:SGR

[2077] C. L. Glenn. Seasonal growth rates of mooneye (Hiodon tergisus) from the Assiniboine River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9): 2078–2082, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-258.

Kerr:1976:ALS

[2078] S. R. Kerr and M. W. Neal. Analysis of large-scale ecological systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2083–2089, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-259.

Stockner:1976:PAE

[2079] John G. Stockner and Naval J. Antia. Phytoplankton adaptation to environmental stresses from toxicants, nutrients, and pollutants — a warning. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2089–2096, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-260.

McCracken:1976:SCA

[2080] F. D. McCracken and R. D. S. Macdonald. Science for Canada’s Atlantic inshore seas fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9): 2097–2139, September 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-261.

Anonymous:1976:LLb

[2081] Anonymous. Letters/lettres. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2140–2141, September 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-262.

Anonymous:1976:BRCc

[2082] Anonymous. Book Reviews/ comptes rendus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(9):2142–2144, September 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-263.

Anonymous:1976:EFb

[2083] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2149–2150, October 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-264.

Anonymous:1976:Pc

[2084] Anonymous. Preface. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10): 2151–2154, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-265.

Campbell:1976:HSP

[2085] Neil J. Campbell. An historical sketch of physical oceanography in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2155–2167, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-266.

Tabata:1976:EWF

[2086] S. Tabata. Evidence of a westward-flowing “Subarctic Countercurrent” in the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2168–2196, October 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-267.

Clarke:1976:MSN

[2087] R. A. Clarke. Measurements of a small near-bottom cyclonic circulation near the Gulf Stream axis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2197–2202, October 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-268.

Ross:1976:SAL

[2088] C. K. Ross and G. T. Needler. Spectral analysis of the long-term oceanographic time series at Ocean Weather Station “P”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2203–2212, October 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-269.

Gatien:1976:SSW

[2089] M. Germaine Gatien. A study in the slope water region south of Halifax. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2213–2217, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-270.

Chang:1976:SCS

[2090] P. Chang, S. Pond, and S. Tabata. Subsurface currents in the Strait of Georgia, west of Sturgeon Bank. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2218–2241, October 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-271.

Thomson:1976:TCE

[2091] Richard E. Thomson. Tidal currents and estuarine-type circulation in Johnstone Strait, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2242–2264, October 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-272.

Buckley:1976:WSC

[2092] J. R. Buckley and S. Pond. Wind and the surface circulation of a fjord. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2265–2271, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-273.

Lake:1976:CAF

[2093] R. A. Lake and E. R. Walker. A Canadian Arctic fjord with some comparisons to fjords of the Western Americas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2272–2285, October 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-274.

Sadler:1976:WHS

[2094] H. Eric Sadler. Water, heat, and salt transports through Nares Strait, Ellesmere Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10): 2286–2295, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-275.

Elliott:1976:SSS

[2095] J. A. Elliott and N. S. Oakey. Spectrum of small-scale oceanic temperature gradients. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10): 2296–2306, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-276.

Banke:1976:RMW

[2096] E. G. Banke, S. D. Smith, and R. J. Anderson. Recent measurements of wind stress on Arctic sea ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2307–2317, October 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-277.

Donelan:1976:MAM

[2097] M. Donelan. A method for the automatic measurement of wave frequency. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2318–2322, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-278.

Kase:1976:SCW

[2098] R. H. Käse and C. L. Tang. Spectra and coherence of wind-generated internal waves. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10): 2323–2328, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-279.

Helbig:1976:SGO

[2099] J. A. Helbig and L. A. Mysak. Strait of Georgia oscillations: Lowfrequency currents and topographic planetary waves. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2329–2339, October 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-280.

Crean:1976:NMS

[2100] P. B. Crean. Numerical model studies of the tides between Vancouver Island and the Mainland Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2340–2344, October 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-281.

Freeman:1976:NMT

[2101] N. G. Freeman and T. S. Murty. Numerical modeling of tides in Hudson Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2345–2361, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-282.

Henry:1976:SSS

[2102] R. F. Henry and N. S. Heaps. Storm surges in the Southern Beaufort Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2362–2376, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-283.

Hamblin:1976:SCS

[2103] P. F. Hamblin. Seiches, circulation, and storm surges of an ice-free Lake Winnipeg. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(10):2377–2391, October 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-284.

Davies:1976:MMW

[2104] Ronald W. Davies and Gerald W. Thompson. Movements of mountain whitefish (Prosopium williamsoni) in the Sheep River Watershed, Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2395–2401, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-285.

Spangler:1976:PLT

[2105] G. R. Spangler and A. H. Berst. Performance of lake trout (Salvelinus namaycush) backcrosses, F_{1} splake (S. fontinalis × S. namaycush), and lake trout in Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2402–2407, November 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-286.

Yamamoto:1976:VAL

[2106] Tats Yamamoto, Richard D. Macdonald, Douglas C. Gillespie, and Russell K. Kelly. Viruses associated with lymphocystis disease and dermal sarcoma of walleye (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2408–2419, November 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-287.

Kincaid:1976:IRT

[2107] Harold L. Kincaid. Inbreeding in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2420–2426, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-288.

VanDenAvyle:1976:ASD

[2108] Michael J. Van Den Avyle. Analysis of seasonal distribution patterns of young largemouth bass (Micropterus salmoides) by use of frequencyof-capture data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11): 2427–2432, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-289.

Stockner:1976:EPE

[2109] John G. Stockner and David D. Cliff. Effects of pulpmill effluent on phytoplankton production in coastal marine waters of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2433–2442, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-290.

Schalles:1976:EDP

[2110] John F. Schalles and Thomas E. Wissing. Effects of dry pellet diets on the metabolic rates of bluegill (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2443–2449, November 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-291.

Stein:1976:CPC

[2111] Roy A. Stein and Michael L. Murphy. Changes in proximate composition of the crayfish Orconectes propinquus with size, sex, and life stage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2450–2458, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-292.

Johnson:1976:EAP

[2112] Lionel Johnson. Ecology of Arctic populations of lake trout, Salvelinus namaycush, lake whitefish, Coregonus clupeaformis, Arctic char, S. alpinus, and associated species in unexploited lakes of the Canadian Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11): 2459–2488, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-293.

Arthur:1976:PFA

[2113] J. R. Arthur, L. Margolis, and H. P. Arai. Parasites of fishes of Aishihik and Stevens Lakes, Yukon Territory, and potential consequences of their interlake transfer through a proposed water diversion for hydroelectrical purposes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2489–2499, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-294.

Jones:1976:TRS

[2114] Barry C. Jones and Glen H. Geen. Taxonomic reevaluation of the spiny dogfish (Squalus acanthias L.) in the Northeastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2500–2506, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-295.

Kabata:1976:ESS

[2115] Z. Kabata. Early stages of some copepods (Crustacea) parasitic on marine fishes of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2507–2525, November 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-296.

Schindler:1976:NWC

[2116] D. W. Schindler, R. W. Newbury, K. G. Beaty, and P. Campbell. Natural water and chemical budgets for a small Precambrian lake basin in Central Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2526–2543, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-297.

Swiss:1976:EDT

[2117] J. J. Swiss and M. G. Johnson. Energy dynamics of two benthic crustaceans in relation to diet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2544–2550, November 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-298.

Hinton:1976:GMF

[2118] David E. Hinton, Raymond T. Jones, and Roger Lee Herman. Glomerular mesangial fibrosis in hatchery-reared rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2551–2559, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-299.

Hiltz:1976:DCD

[2119] Doris Fraser Hiltz, Barbara Smith Lall, D. W. Lemon, and W. J. Dyer. Deteriorative changes during frozen storage in fillets and minced flesh of silver hake (Merluccius bilinearis) processed from round fish held in ice and refrigerated sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2560–2567, November 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-300.

Sameoto:1976:RRE

[2120] D. D. Sameoto. Respiration rates, energy budgets, and molting frequencies of three species of euphausiids found in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2568–2576, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-301.

Wellings:1976:FED

[2121] S. R. Wellings, C. E. Alpers, B. B. McCain, and B. S. Miller. Fin erosion disease of starry flounder (Platichthys stellatus) and English sole (Parophrys vetulus) in the estuary of the Duwamish River, Seattle, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2577–2586, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-302.

Zeitoun:1976:MMD

[2122] Ibrahim H. Zeitoun, Duane E. Ullrey, Werner G. Bergen, and William T. Magee. Mineral metabolism during the ontogenesis of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11): 2587–2591, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-303.

Haynes:1976:RNE

[2123] Evan Haynes, John F. Karinen, Jeffrey Watson, and Daniel J. Hopson. Relation of number of eggs and egg length to carapace width in the brachyuran crabs Chionoecetes bairdi and C. opilio from the Southeastern Bering Sea and C. opilio from the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2592–2595, November 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-304.

Stasko:1976:MCT

[2124] A. B. Stasko and R. M. Horrall. Method of counting tailbeats of freeswimming fish by ultrasonic telemetry techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2596–2598, November 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-305.

White:1976:GIC

[2125] A. W. White. Growth inhibition caused by turbulence in the toxic marine dinoflagellate Gonyaulax excavata. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2598–2602, November 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-306.

MacKinnon:1976:EIP

[2126] Colin N. MacKinnon and Edward M. Donaldson. Environmentally induced precocious sexual development in the male pink salmon (Oncorhynchus gorbuscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2602–2605, November 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-307.

Templeman:1976:TMS

[2127] Wilfred Templeman. Transatlantic migrations of spiny dogfish (Squalus acanthias). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2605–2609, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-308.

Mayer:1976:RDE

[2128] Foster L. Mayer. Residue dynamics of di-2-ethylhexyl phthalate in fathead minnows (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2610–2613, November 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-309.

Hendricks:1976:PBD

[2129] J. D. Hendricks, L. J. Hunter, and J. H. Wales. Postmortem bile damage to rainbow trout (Salmo gairdneri) livers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2613–2616, November 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-310.

Lee:1976:ASS

[2130] R. L. G. Lee and G. Power. Atlantic salmon (Salmo salar) of the Leaf River, Ungava Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11): 2616–2621, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-311.

Elazhary:1976:OIP

[2131] M. A. S. Y. Elazhary, A. Lagace, G. Cousineau, R. S. Roy, L. Berthiaume, P. Paulhus, and J. L. Frechette. Outbreak of infectious pancreatic necrosis in yearling brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2621–2625, November 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-312.

Rinne:1976:CSC

[2132] John N. Rinne. Coded spine clipping to identify individuals of the spinyrayed fish Tilapia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11): 2626–2629, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-313.

Nebeker:1976:EAS

[2133] Alan V. Nebeker, Donald G. Stevens, and Richard K. Stroud. Effects of air-supersaturated water on adult sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2629–2633, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-314.

Harvey:1976:AEQ

[2134] H. H. Harvey. Aquatic environmental quality: Problems and proposals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2634–2670, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-315.

Hamilton:1976:AEQ

[2135] R. D. Hamilton. Aquatic environmental quality: Toxicology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2671–2688, November 1976. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f76-316. See comments

[2464].

Anonymous:1976:BRCd

[2136] Anonymous. Book Reviews/ comptes rendus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(11):2689–2690, November 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-317.

Anonymous:1976:EE

[2137] Anonymous. Editorial/Editorial. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2697–2698, December 1976. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f76-318.

Anonymous:1976:AAb

[2138] Anonymous. Announcement/Avis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2698, December 1976. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f76-319.

Wedemeyer:1976:PRJ

[2139] Gary A. Wedemeyer. Physiological response of juvenile Coho salmon (Oncorhynchus kisutch) and rainbow trout (Salmo gairdneri) to handling and crowding stress in intensive fish culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2699–2702, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-320.

Peters:1976:ALW

[2140] G. B. Peters, H. J. Dawson, B. F. Hrutfiord, and R. R. Whitney. Aqueous leachate from western red cedar: Effects on some aquatic organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2703–2709, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-321.

Engel:1976:VHD

[2141] Sandy Engel and John J. Magnuson. Vertical and horizontal distributions of Coho salmon (Oncorhynchus kisutch), yellow perch (Perca flavescens), and cisco (Coregonus artedii) in Pallette Lake, Wisconsin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2710–2715, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-322.

Bams:1976:SPH

[2142] R. A. Bams. Survival and propensity for homing as affected by presence or absence of locally adapted paternal genes in two transplanted populations of pink salmon (Oncorhynchus gorbuscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2716–2725, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-323.

McKim:1976:LTE

[2143] J. M. McKim, G. F. Olson, G. W. Holcombe, and E. P. Hunt. Long-term effects of methylmercuric chloride on three generations of brook trout (Salvelinus fontinalis): Toxicity, accumulation, distribution, and elimination. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2726–2739, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-324.

Maki:1976:ETM

[2144] Alan W. Maki and Howard E. Johnson. Evaluation of a toxicant on the metabolism of model stream communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2740–2746, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-325.

Therriault:1976:NCI

[2145] Jean-Claude Therriault and Guy Lacroix. Nutrients, chlorophyll, and internal tides in the St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2747–2757, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-326.

Stockner:1976:MPG

[2146] John G. Stockner and Anne C. Costella. Marine phytoplankton growth in high concentrations of pulpmill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2758–2765, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-327.

Symons:1976:BGJ

[2147] P. E. K. Symons. Behavior and growth of juvenile Atlantic salmon (Salmo salar) and three competitors at two stream velocities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2766–2773, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-328.

Fong:1976:URK

[2148] Wang C. Fong. Uptake and retention of Kuwait crude oil and its effects on oxygen uptake by the soft-shell clam, Mya arenaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2774–2780, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-329.

Levy:1976:HSO

[2149] E. M. Levy and A. Walton. High Seas oil pollution: Particulate petroleum residues in the North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2781–2791, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-330.

McAllister:1976:VNN

[2150] Don E. McAllister and Jadwiga Aniskowicz. Vertebral number in North American sculpins of the Myoxocephalus quadricornis-complex. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2792–2799, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-331.

Lean:1976:PKL

[2151] D. R. S. Lean. Phosphorus kinetics in lake water: Influence of membrane filter pore size and low-pressure filtration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2800–2804, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-332.

Paerl:1976:VOP

[2152] H. W. Paerl and D. R. S. Lean. Visual observations of phosphorus movement between algae, bacteria, and abiotic particles in Lake Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2805–2813, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-333.

Stemberger:1976:NLM

[2153] Richard S. Stemberger. Notholca laurentiae and N. michiganensis, new rotifers from the Laurentian Great Lakes region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2814–2818, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-334.

Ackman:1976:VKA

[2154] R. G. Ackman. Volatile ketones and alcohols of codworms (Terranova decipiens) from axenic culture and fresh fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2819–2821, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-335.

Holub:1976:IFA

[2155] B. J. Holub, J. Piekarski, C. Y. Cho, and S. J. Slinger. Incorporation of fatty acids into phosphatidylcholine by Acyl-CoA: l-acyl-sn-glycero-3-phosphorylcholine acyltransferase in liver of rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2821–2826, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-336.

Sherburne:1976:PPG

[2156] Stuart W. Sherburne. Phosphoglucomutase polymorphism in the green crab, Carcinus maenas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2826–2829, December 1976. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f76-337.

Jones:1976:CTF

[2157] Dixon D. Jones and Carl J. Walters. Catastrophe theory and fisheries regulation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2829–2833, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-338.

Walton:1976:FDD

[2158] James M. Walton and Norman W. Bartoo. Flatfish densities determined with a diver-operated flounder sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2834–2836, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-339.

Bohn:1976:TMC

[2159] Arne Bohn and R. O. McElroy. Trace metals (As, Cd, Cu, Fe, and Zn) in Arctic cod, Boreogadus saida, and selected zooplankton from Strathcona Sound, Northern Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2836–2840, December 1976. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f76-340.

Crim:1976:GHT

[2160] L. W. Crim and D. M. Evans. Gonadotropic hormone treatment of rainbow trout (Salmo gairdneri): Plasma hormone profile following a single injection. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12): 2841–2844, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-341.

Vinyard:1976:ELT

[2161] Gary L. Vinyard and W. John O’Brien. Effects of light and turbidity on the reactive distance of bluegill (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2845–2849, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-342.

Fisher:1976:REF

[2162] William S. Fisher. Relationships of epibiotic fouling and mortalities of eggs of the Dungeness crab (Cancer magister). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2849–2853, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-343.

Buckley:1976:ATR

[2163] James A. Buckley. Acute toxicity of residual chlorine in wastewater to Coho salmon (Oncorhynchus kisutch) and some resultant hematologic changes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2854–2856, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-344.

Hansen:1976:IPA

[2164] C. B. Hansen and A. J. Lingg. Inert particle agglutination tests for detection of antibody to enteric redmouth bacterium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2857–2860, December 1976. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f76-345.

Brakke:1976:MWW

[2165] David F. Brakke. Modification of the whiteside–Williams pattern sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2861–2863, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f76-346.

Anonymous:1976:LLc

[2166] Anonymous. Letters/letters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 33(12):2864, December 1976. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f76-347.

Suzumoto:1977:RRT

[2167] Bruce K. Suzumoto, Carl B. Schreck, and John D. McIntyre. Relative resistances of three transferrin genotypes of Coho salmon (Oncorhynchus kisutch) and their hematological responses to bacterial kidney disease. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):1–8, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-001.

Buhler:1977:TAE

[2168] Donald R. Buhler, Robert M. Stokes, and Richard S. Caldwell. Tissue accumulation and enzymatic effects of hexavalent chromium in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):9–18, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-002.

Sutcliffe:1977:CFC

[2169] W. H. Sutcliffe, Jr., K. Drinkwater, and B. S. Muir. Correlations of fish catch and environmental factors in the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):19–30, January 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-003.

Peterson:1977:DAS

[2170] R. H. Peterson, H. C. E. Spinney, and A. Sreedharan. Development of Atlantic salmon (Salmo salar) eggs and alevins under varied temperature regimes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):31–43, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-004.

Jones:1977:ADE

[2171] Barry C. Jones and Glen H. Geen. Age determination of an elasmobranch (Squalus acanthias) by X-ray spectrometry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):44–48, January 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-005.

Keeney:1977:UFE

[2172] Ralph L. Keeney. A utility function for examining policy affecting salmon on the Skeena River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (1):49–63, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-006.

Hilborn:1977:DGS

[2173] Ray Hilborn and Carl J. Walters. Differing goals of salmon management on the Skeena River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (1):64–72, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-007.

Gerritsen:1977:EPC

[2174] Jeroen Gerritsen and J. Rudi Strickler. Encounter probabilities and community structure in zooplankton: a mathematical model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):73–82, January 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-008.

Zeitoun:1977:DRP

[2175] Ibrahim H. Zeitoun, Duane E. Ullrey, Werner G. Bergen, and William T. Magee. DNA, RNA, protein, and free amino acids during ontogenesis of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):83–88, January 1977. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f77-009.

Roberts:1977:ILP

[2176] J. R. Roberts, A. S. W. deFreitas, and M. A. J. Gidney. Influence of lipid pool size on bioaccumulation of the insecticide chlordane by northern redhorse suckers (Moxostoma macrolepidotum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):89–97, January 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-010.

Oliver:1977:BGG

[2177] D. R. Oliver. Bicinctus-group of the genus Cricotopus Van der Wulp (Diptera: Chironomidae) in the Nearctic with a description of a new species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):98–104, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-011.

Beverley-Burton:1977:MSO

[2178] Mary Beverley-Burton, O. L. Nyman, and John H. C. Pippy. The morphology, and some observations on the population genetics of Anisakis simplex larvae (Nematoda: Ascaridata) from fishes of the North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):105–112, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-012.

Carter:1977:ISV

[2179] John C. H. Carter and Judith K. Kwik. Instar succession, vertical distribution, and interspecific competition among four species of Chaoborus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):113–118, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-013.

Addison:1977:RCP

[2180] R. F. Addison and M. E. Zinck. Rate of conversion of ^{14}C-p,p-DDT to p,p’-DDE by brook trout (Salvelinus fontinalis): Absence of effect of pretreatment of fish with compounds related to p,p’-DDT. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):119–122, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-014.

Reisenbichler:1977:GDG

[2181] R. R. Reisenbichler and J. D. McIntyre. Genetic differences in growth and survival of juvenile hatchery and wild steelhead trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):123–128, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-015.

Alt:1977:ISL

[2182] Kenneth T. Alt. Inconnu, Stenodus leucichthys, migration studies in Alaska 1961–74. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1): 129–133, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-016.

Li:1977:NCL

[2183] M. F. Li and S. Clyburne. New cell line from the marine lumpfish, Cyclopterus lumpus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (1):134–139, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-017.

Zitko:1977:SDV

[2184] V. Zitko and W. G. Carson. Seasonal and developmental variation in the lethality of zinc to juvenile Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):139–141, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-018.

Robbins:1977:VSM

[2185] William H. Robbins and Hugh R. MacCrimmon. Vital statistics and migratory patterns of a potamodromous stock of smallmouth bass, Micropterus dolomieui. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (1):142–147, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-019.

Scott:1977:BCF

[2186] J. S. Scott. Back-calculated fish lengths and Hg and Zn levels from recent and 100-yr-old cleithrum bones from Atlantic cod (Gadus morhua). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):147–150, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-020.

Glova:1977:CSS

[2187] G. J. Glova and J. E. McInerney. Critical swimming speeds of Coho salmon (Oncorhynchus kisutch) fry to smolt stages in relation to salinity and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (1):151–154, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-021.

Flint:1977:CGL

[2188] R. Warren Flint and Charles R. Goldman. Crayfish growth in Lake Tahoe: Effects of habitat variation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):155–159, January 1977. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f77-022.

Anderson:1977:MDI

[2189] William C. Anderson and Patrick J. Manion. Morphological development of the integument of the sea lamprey, Petromyzon marinus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):159–163, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-023.

Vernberg:1977:RLM

[2190] Winona B. Vernberg, Bruce C. Coull, and Darwin D. Jorgensen. Reliability of laboratory metabolic measurements of meiofauna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):164–167, January 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-024.

Nakashima:1977:SDL

[2191] Brian S. Nakashima, Dominique Gascon, and William C. Leggett. Species diversity of littoral zone fishes along a phosphorus–production gradient in Lake Memphremagog, Quebec–Vermont. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(1):167–170, January 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-025.

Anonymous:1977:E

[2192] Anonymous. Editorial. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):173–176, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-026.

Harding:1977:USW

[2193] Gareth C. H. Harding and W. Peter Vass. Uptake from sea water and clearance of ^{14}C-p,p’-DDT by the marine copepod Calanus finmarchicus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):177–182, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-027.

Eberhardt:1977:PDM

[2194] L. L. Eberhardt and D. B. Siniff. Population dynamics and marine mammal management policies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):183–190, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-028.

Kabata:1977:HPR

[2195] Z. Kabata and B. Cousens. Host–parasite relationships between sockeye salmon, Oncorhynchus nerka, and Salmincola californiensis (Copepoda: Lernaeopodidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2): 191–202, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-029.

Antipa:1977:IPS

[2196] Ross Antipa and Donald F. Amend. Immunization of Pacific salmon: Comparison of intraperitoneal injection and hyperosmotic infiltration of Vibrio anguillarum and Aeromonas salmonicida bacterins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):203–208, February 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-030.

Shaw:1977:CSC

[2197] D. H. Shaw and J. R. Botta. Chemical and sensory changes in round inshore male capelin (Mallotus villosus) during prolonged storage for 24 months at −23 ^{o}C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2): 209–214, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-031.

Buckley:1977:HBH

[2198] James A. Buckley. Heinz body hemolytic anemia in Coho salmon (Oncorhynchus kisutch) exposed to chlorinated wastewater. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):215–224, February 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-032.

Harding:1977:SAE

[2199] Gareth C. H. Harding. Surface area of the euphausiid Thysanöessa raschii and its relation to body length, weight, and respiration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):225–231, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-033.

Turner:1977:CSS

[2200] J. L. Turner. Changes in the size structure of cichlid populations of Lake Malaŵi resulting from bottom trawling. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):232–238, February 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-034.

Cherry:1977:PAL

[2201] Donald S. Cherry, Kenneth L. Dickson, John Cairns, Jr., and Jay R. Stauffer. Preferred, avoided, and lethal temperatures of fish during rising temperature conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):239–246, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-035.

Rosenberg:1977:LHC

[2202] David M. Rosenberg, Allen P. Wiens, and Ole A. Sæther. Life histories of Cricotopus (Cricotopus) bicinctus and C. (C.) mackenziensis (Diptera: Chironomidae) in the Fort Simpson Area, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):247–253, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-036.

Rosenberg:1977:RCO

[2203] David M. Rosenberg, Allen P. Wiens, and Ole A. Sæther. Responses to crude oil contamination by Cricotopus (Cricotopus) bicinctus and C. (C.) mackenziensis (Diptera: Chironomidae) in the Fort Simpson Area, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):254–261, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-037.

Trojnar:1977:ELS

[2204] John R. Trojnar. Egg and larval survival of white suckers (Catostomus commersoni) at low pH. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):262–266, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-038.

Kligerman:1977:RCP

[2205] Andrew D. Kligerman and Stephen E. Bloom. Rapid chromosome preparations from solid tissues of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):266–269, February 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-039.

King:1977:NIS

[2206] Patricia A. King, Thomas E. Wissing, Edward J. De Villez, Michael J. Chimney, and William B. Randall, Jr. Nutritional implications of surface feeding by gizzard shad (Dorosoma cepedianum) on insect remains in an Ohio reservoir. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2): 269–273, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-040.

Arnold:1977:FBB

[2207] David C. Arnold. Fecundity of Balanus balanoides in Passamaquoddy Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):273–275, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-041.

Maki:1977:KLT

[2208] Alan W. Maki and Howard E. Johnson. Kinetics of lampricide (TFM, 3-trifluoromethyl-4-nitrophenol) residues in model stream communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):276–281, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-042.

Sherburne:1977:OPE

[2209] Stuart W. Sherburne. Occurrence of piscine erythrocytic necrosis (PEN) in the blood of the anadromous alewife, Alosa pseudoharengus, from Maine coastal streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):281–286, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-043.

Scavia:1977:CEM

[2210] Donald Scavia and Steven C. Chapra. Comparison of an ecological model of Lake Ontario and phosphorus loading models. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):286–290, February 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-044.

Eggers:1977:FID

[2211] Douglas M. Eggers. Factors in interpreting data obtained by diel sampling of fish stomachs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2): 290–294, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-045.

Dolan:1977:SRA

[2212] J. A. Dolan and G. Power. Sex ratio of American eels, Anguilla rostrata, from the Matamek River System, Quebec, with remarks on problems in sexual identification. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2): 294–299, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-046.

Reynolds:1977:FOB

[2213] William W. Reynolds. Fish orientation behavior: An electronic device for studying simultaneous responses to two variables. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):300–304, February 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-047.

Thomas:1977:DSP

[2214] Allan E. Thomas and Michael J. Donahoo. Differences in swimming performance among strains of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):304–307, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-048.

Povoledo:1977:PCC

[2215] D. Povoledo. Physical and chemical characteristics of aquatic humus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):308–309, February 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-049.

Penrose:1977:BRG

[2216] W. R. Penrose. Book review: Gas Bubble Disease. Proceedings of a Workshop held at Richland, Washington, October 8–9, 1974. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):309, February 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-050.

Gerking:1977:FM

[2217] S. D. Gerking. The fishes of Missouri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(2):310, February 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-051.

Kimura:1977:SAA

[2218] Daniel K. Kimura. Statistical assessment of the age–length key. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):317–324, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-052.

Amin:1977:HSA

[2219] Omar M. Amin and Janet M. Burrows. Host and seasonal associations of Echinorhynchus salmonis (Acanthocephala: Echinorhynchidae) in Lake Michigan fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (3):325–331, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-053.

Taubert:1977:DRO

[2220] Bruce D. Taubert and Daniel W. Coble. Daily rings in otoliths of three species of Lepomis and Tilapia mossambica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):332–340, March 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-054.

Dodson:1977:TPR

[2221] Julian J. Dodson and J. Clifton Young. Temperature and photoperiod regulation of rheotropic behavior in prespawning common shiners, Notropis cornutus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (3):341–346, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-055.

Keizer:1977:HEC

[2222] P. D. Keizer, D. C. Gordon, Jr., and J. Dale. Hydrocarbons in Eastern Canadian marine waters determined by fluorescence spectroscopy and gas–liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):347–353, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-056.

Winters:1977:ETE

[2223] G. H. Winters. Estimates of tag extrusion and initial tagging mortality in Atlantic herring (Clupea harengus harengus) released with abdominally inserted magnetic tags. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):354–359, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-057.

Werner:1977:HPF

[2224] Earl E. Werner, Donald J. Hall, Dennis R. Laughlin, Donald J. Wagner, Leni A. Wilsmann, and Frederick C. Funk. Habitat partitioning in a freshwater fish community. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):360–370, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-058.

Beamish:1977:VSD

[2225] R. J. Beamish and E. J. Crossman. Validity of the subspecies designation for the dwarf white sucker (Catostomus commersoni utawana). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):371–378, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-059.

Boudreault:1977:MLP

[2226] F.-Robert Boudreault, J.-Normand Dupont, and Claudine Sylvain. Modèle linéaires de prédiction des débarquements de homard aux Îles-de-la-Madeleine (Golfe du Saint-Laurent). (French) [Linear predictive models of lobster landings in the Magdalen Islands (Gulf of St. Lawrence)]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):379–383, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-060.

Newkirk:1977:GLT

[2227] G. F. Newkirk, D. L. Waugh, and L. E. Haley. Genetics of larval tolerance to reduced salinities in two populations of oysters, Crassostrea virginica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):384–387, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-061.

Bodaly:1977:PWE

[2228] R. A. Bodaly and C. C. Lindsey. Pleistocene watershed exchanges and the fish fauna of the Peel River Basin, Yukon Territory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):388–395, March 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-062.

Gunn:1977:FAF

[2229] J. M. Gunn, S. U. Qadri, and D. C. Mortimer. Filamentous algae as a food source for the brown bullhead (Ictalurus nebulosus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):396–401, March 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-063.

Udey:1977:ICA

[2230] Lanny R. Udey, Elaine Young, and Bennett Sallman. Isolation and characterization of an anaerobic bacterium, Eubacterium tarantellus sp. nov., associated with striped mullet (Mugil cephalus) mortality in biscayne Bay, Florida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34 (3):402–409, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-064.

Grice:1977:REP

[2231] George D. Grice and Victoria R. Gibson. Resting eggs in Pontella meadi (Copepoda: Calanoida). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):410–412, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-065.

White:1977:RTC

[2232] Alan W. White, Robert G. Sheath, and Johan A. Hellebust. A Red Tide caused by the marine ciliate Mesodinium rubrum in Passamaquoddy Bay, including pigment and ultrastructure studies of the endosymbiont. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):413–416, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-066.

Guerinot:1977:NFA

[2233] M. L. Guerinot, W. Fong, and D. G. Patriquin. Nitrogen fixation (acetylene reduction) associated with sea urchins (Strongylocentrotus droebachiensis) feeding on seaweeds and eelgrass. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):416–420, March 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-067.

Iredale:1977:ECA

[2234] D. G. Iredale and R. K. York. Effects of chemical additives on extending the shelf life of frozen minced sucker (Catostomidae) flesh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):420–425, March 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-068.

Eberhardt:1977:RBT

[2235] L. L. Eberhardt. Relationship between two stock–recruitment curves. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):425–428, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-069.

Wedemeyer:1977:STB

[2236] Gary A. Wedemeyer and Nancy C. Nelson. Survival of two bacterial fish pathogens (Aeromonas salmonicida and the enteric redmouth bacterium) in ozonated, chlorinated, and untreated waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):429–432, March 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-070.

Fisher:1977:TTE

[2237] William S. Fisher and Richard T. Nelson. Therapeutic treatment for epibiotic fouling on Dungeness crab (Cancer magister) larvae reared in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3): 432–436, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-071.

Luoma:1977:DTC

[2238] Samuel N. Luoma. Detection of trace contaminant effects in aquatic ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):436– 439, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-072.

Lemon:1977:HAM

[2239] D. W. Lemon and L. W. Regier. Holding of Atlantic mackerel (Scomber scombrus) in refrigerated sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):439–443, March 1977. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f77-073.

Berkes:1977:PEC

[2240] F. Berkes. Production of the euphausiid crustacean Thysanoëssa raschii in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):443–446, March 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-074.

Eales:1977:ECC

[2241] J. G. Eales. Eel capture, culture, processing and marketing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(3):447, March 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-075.

Forrester:1977:FSH

[2242] C. R. Forrester, H. Tsuyuki, S. Fuke, J. E. Smith, and J. Schnute. Flathead sole (Hippoglossoides) in the North Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):455–462, April 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-076.

Flossner:1977:VTP

[2243] Dietrich Flössner and Karl Kraus. On the variability and taxonomy of Pleuroxus denticulatus Birge (Cladocera: Chydoridae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):463–476, April 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-077.

McLeay:1977:DBS

[2244] Donald J. McLeay. Development of a blood sugar bioassay for rapidly measuring stressful levels of pulpmill effluent to salmonid fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):477–485, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-078.

Perrone:1977:PEC

[2245] Stephen J. Perrone and Thomas L. Meade. Protective effect of chloride on nitrite toxicity to Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):486–492, April 1977. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f77-079.

Sigmon:1977:RBD

[2246] Catherine F. Sigmon, Henry J. Kania, and R. J. Beyers. Reductions in biomass and diversity resulting from exposure to mercury in artificial streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4): 493–500, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-080.

Hodson:1977:EEA

[2247] Peter V. Hodson, Beverley R. Blunt, Douglas J. Spry, and Keith Austen. Evaluation of erythrocyte δ-amino levulinic acid dehydratase activity as a short-term indicator in fish of a harmful exposure to lead. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):501–508, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-081.

Adelman:1977:EBG

[2248] Ira R. Adelman. Effects of bovine growth hormone on growth of carp (Cyprinus carpio) and the influences of temperature and photoperiod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):509–515, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-082.

El-Sabh:1977:OFC

[2249] Mohammed I. El-Sabh. Oceanographic features, currents, and transport in Cabot Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4): 516–528, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-083.

Sherberger:1977:ETS

[2250] Fred F. Sherberger, Ernest F. Benfield, Kenneth L. Dickson, and John Cairns, Jr. Effects of thermal shocks on drifting aquatic insects: A laboratory simulation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4): 529–536, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-084.

Conway:1977:BUR

[2251] H. L. Conway, J. I. Parker, E. M. Yaguchi, and D. L. Mellinger. Biological utilization and regeneration of silicon in Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):537–544, April 1977. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f77-085.

Parker:1977:DDF

[2252] J. I. Parker, H. L. Conway, and E. M. Yaguchi. Dissolution of diatom frustules and recycling of amorphous silicon in Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):545–551, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-086.

Parker:1977:SPD

[2253] J. I. Parker, H. L. Conway, and E. M. Yaguchi. Seasonal periodicity of diatoms, and silicon limitation in offshore Lake Michigan, 1975. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):552–558, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-087.

Gershey:1977:CMD

[2254] R. M. Gershey, R. A. Nevé, D. L. Musgrave, and P. B. Reichardt. A colorimetric method for determination of saxitoxin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):559–563, April 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-088.

Tenore:1977:EMI

[2255] Kenneth R. Tenore, John H. Tietjen, and John J. Lee. Effect of meiofauna on incorporation of aged eelgrass, Zostera marina, detritus by the polychaete Nephthys incisa. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):563–567, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-089.

Lewis:1977:ABK

[2256] Frank G. Lewis III and Robert J. Livingston. Avoidance of bleached kraft pulpmill effluent by pinfish (Lagodon rhomboides) and Gulf killifish (Fundulus grandis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4): 568–570, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-090.

McGroddy:1977:ENN

[2257] P. M. McGroddy and R. L. Wyman. Efficiency of nets and a new device for sampling living fish larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):571–574, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-091.

Trojnar:1977:EHT

[2258] John R. Trojnar. Egg hatchability and tolerance of brook trout (Salvelinus fontinalis) fry at low pH. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):574–579, April 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-092.

Kerr:1977:JTA

[2259] S. R. Kerr. JPU theory and the apotheosis of simulacra. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(4):580–581, April 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-093.

Schnute:1977:IES

[2260] J. Schnute. Improved estimates from the Schaefer Production Model: theoretical considerations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):583–603, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-094.

Bernard:1977:FRC

[2261] F. R. Bernard. Fishery and reproductive cycle of the Red Sea urchin, Strongylocentrotus franciscanus, in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):604–610, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-095.

Alpers:1977:LDY

[2262] C. E. Alpers, B. B. McCain, M. S. Myers, and S. R. Wellings. Lymphocystis disease in yellowfin sole (Limanda aspera) in the Bering Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):611–616, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-096.

Franzin:1977:BGS

[2263] W. G. Franzin and J. W. Clayton. A biochemical genetic study of zoogeography of lake whitefish (Coregonus clupeaformis) in Western Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):617–625, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-097.

Lockhart:1977:MRS

[2264] W. L. Lockhart, D. A. Metner, and J. Solomon. Methoxychlor residue studies in caged and wild fish from the Athabasca River, Alberta, following a single application of blackfly larvicide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):626–632, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-098.

Heidinger:1977:ETF

[2265] Roy C. Heidinger and Steven D. Crawford. Effect of temperature and feeding rate on the liver–somatic index of the largemouth bass, Micropterus salmoides. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5): 633–638, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-099.

Hokanson:1977:ECT

[2266] Kenneth E. F. Hokanson, Charles F. Kleiner, and Todd W. Thorslund. Effects of constant temperatures and diel temperature fluctuations on specific growth and mortality rates and yield of juvenile rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5): 639–648, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-100.

Beamish:1977:PLA

[2267] R. J. Beamish and J. C. Van Loon. Precipitation loading of acid and heavy metals to a small acid lake near Sudbury, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):649–658, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-101.

Antia:1977:RCP

[2268] N. J. Antia and J. Y. Cheng. Reexamination of the carotenoid pigments of the unicellular blue-green alga Agmenellum quadruplicatum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):659–668, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-102.

Thomas:1977:RSH

[2269] Sylvan R. Thomas and Raymond A. Soltero. Recent sedimentary history of a eutrophic reservoir: Long Lake, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):669–676, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-103.

Leatherland:1977:EDA

[2270] J. F. Leatherland, C. Y. Cho, and S. J. Slinger. Effects of diet, ambient temperature, and holding conditions on plasma thyroxine levels in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):677–682, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-104.

Hilton:1977:FAS

[2271] J. W. Hilton, C. Y. Cho, and S. J. Slinger. Factors affecting the stability of supplemental ascorbic acid in practical trout diets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):683–687, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-105.

Cornick:1977:SAL

[2272] John W. Cornick and James E. Stewart. Survival of American lobsters (Homarus americanus) stored in a recirculating, refrigerated seawater system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):688– 692, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-106.

Rich:1977:PCA

[2273] Peter H. Rich and Barbara L. Pallotti. Potential contribution of atmospheric fallout to the phosphorus budget of Columbia Lake, Connecticut. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):692–697, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-107.

Thomson:1977:CJS

[2274] Richard E. Thomson. Currents in Johnstone Strait, British Columbia: Supplemental data on the Vancouver Island side. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):697–703, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-108.

Confer:1977:CIS

[2275] J. L. Confer and J. M. Cooley. Copepod instar survival and predation by zooplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5): 703–706, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-109.

Boyce:1977:EPE

[2276] Norbert P. Boyce and S. Behrens Yamada. Effects of a parasite, Eubothrium salvelini (Cestoda: Pseudophyllidea), on the resistance of juvenile sockeye salmon, Oncorhynchus nerka, to zinc. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):706–709, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-110.

Anonymous:1977:FAS

[2277] Anonymous. Fisheries and aquatic sciences in Canada: An overview. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):710–727, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-111.

Richards:1977:TPS

[2278] F. Paul Richards. Temperature preference studies in environmental impact assessments: An overview with procedural recommendations: Introduction. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):728– 729, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-112.

Crawshaw:1977:PBR

[2279] L. I. Crawshaw. Physiological and behavioral reactions of fishes to temperature change. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5): 730–734, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-113.

Reynolds:1977:TPF

[2280] William W. Reynolds. Temperature as a proximate factor in orientation behavior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5): 734–739, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-114.

Coutant:1977:CTP

[2281] Charles C. Coutant. Compilation of temperature preference data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):739–745, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-115.

Gift:1977:ATP

[2282] James J. Gift. Application of temperature preference studies to environmental impact assessment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):746–749, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-116.

McCauley:1977:LMD

[2283] Robert W. McCauley. Laboratory methods for determining temperature preference. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5): 749–752, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-117.

Anonymous:1977:PD

[2284] Anonymous. Panel discussion. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):752–754, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-118.

Anonymous:1977:Ra

[2285] Anonymous. Recommendations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):754–755, May 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-119.

Anonymous:1977:Rb

[2286] Anonymous. References. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):755–761, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-120.

Anonymous:1977:LLa

[2287] Anonymous. Letters/lettres. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):762–763, May 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-121.

Giovando:1977:CSB

[2288] L. F. Giovando. The coast and shelf of the Beaufort Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):764–766, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-122.

Greer:1977:STE

[2289] G. L. Greer. Survival in toxic environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(5):766, May 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-123.

Kelso:1977:PPM

[2290] J. R. M. Kelso, C. K. Minns, and R. J. P. Brouzes. Pulp and paper mill effluent in a freshwater environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):771–775, June 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-124.

Minns:1977:APP

[2291] Charles K. Minns. Analysis of a pulp and paper mill effluent plume. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):776–783, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-125.

Johnson:1977:CCA

[2292] M. G. Johnson. Caloric changes along pulp and paper mill effluent plumes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):784– 790, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-126.

Leslie:1977:CSP

[2293] J. K. Leslie. Characterization of suspended particles in some pulp and paper mill effluent plumes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):791–797, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-127.

Fox:1977:PDO

[2294] Michael E. Fox. Persistence of dissolved organic compounds in kraft pulp and paper mill effluent plumes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):798–804, June 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-128.

Polak:1977:CPC

[2295] J. Polak and M. D. Palmer. Concentration pattern of chemical constituents in a paper Mill’s effluent plume: Dynamics and model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):805–816, June 1977. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f77-129.

Sandilands:1977:EPM

[2296] R. G. Sandilands. Effect of pulp mill effluent on the surficial sediments of Western Nipigon Bay, Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):817–823, June 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-130.

VanderWal:1977:RBN

[2297] J. Vander Wal. Relation between Nipigon Bay benthic macroinvertebrates and selected aspects of their habitat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):824–829, June 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-131.

Brownlee:1977:DSO

[2298] B. Brownlee and W. M. J. Strachan. Distribution of some organic compounds in the receiving waters of a kraft pulp and paper mill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):830–837, June 1977. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f77-132.

Brownlee:1977:DDA

[2299] B. Brownlee, M. E. Fox, W. M. J. Strachan, and S. R. Joshi. Distribution of dehydroabietic acid in sediments adjacent to a kraft pulp and paper mill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6): 838–843, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-133.

Rokosh:1977:EEI

[2300] D. A. Rokosh, S. S. Rao, and A. A. Jurkovic. Extent of effluent influence on lake water determined by bacterial population distributions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):844–849, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-134.

Kaiser:1977:OCR

[2301] Klaus L. E. Kaiser. Organic contaminant residues in fishes from Nipigon Bay, Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6): 850–855, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-135.

Moore:1977:EPP

[2302] J. E. Moore and R. J. Love. Effect of a pulp and paper mill effluent on the productivity of periphyton and phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):856–862, June 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-136.

Cooley:1977:FRP

[2303] John M. Cooley. Filtering rate performance of Daphnia retrocurva in pulp mill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6): 863–868, June 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-137.

Whittle:1977:AAT

[2304] D. M. Whittle and K. W. Flood. Assessment of the acute toxicity, growth impairment, and flesh tainting potential of a bleached kraft mill effluent on rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):869–878, June 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-138.

Kelso:1977:DDM

[2305] J. R. M. Kelso. Density, distribution, and movement of Nipigon Bay fishes in relation to a pulp and paper mill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(6):879–885, June 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-139.

Margolis:1977:PHA

[2306] L. Margolis. Public health aspects of “Codworm” infection: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):887–898, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-140.

VanLoon:1977:HMC

[2307] J. C. Van Loon and R. J. Beamish. Heavy-metal contamination by atmospheric fallout of several flin flon area lakes and the relation to fish populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):899– 906, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-141.

Stockner:1977:PPD

[2308] John G. Stockner, David D. Cliff, and Douglas B. Buchanan. Phytoplankton production and distribution in Howe Sound, British Columbia: A coastal marine embayment–fjord under stress. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):907–917, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-142.

Peters:1977:AAO

[2309] Robert Henry Peters. Availability of atmospheric orthophosphate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):918–924, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-143.

Forney:1977:RYP

[2310] John L. Forney. Reconstruction of yellow perch (Perca flavescens) cohorts from examination of walleye (Stizostedion vitreum vitreum) stomachs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):925–932, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-144.

Zinn:1977:SSS

[2311] J. L. Zinn, K. A. Johnson, J. E. Sanders, and J. L. Fryer. Susceptibility of salmonid species and hatchery strains of Chinook salmon (Oncorhynchus tshawytscha) to infections by Ceratomyxa shasta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):933–936, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-145.

Addison:1977:ORM

[2312] R. F. Addison and P. F. Brodie. Organochlorine residues in maternal blubber, milk, and pup blubber from grey seals (Halichoerus grypus) from Sable Island, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):937–941, July 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-146.

Bell:1977:DEP

[2313] Graham Bell, Paul Handford, and Carl Dietz. Dynamics of an exploited population of lake whitefish (Coregonus clupeaformis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):942–953, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-147.

Handford:1977:GFC

[2314] Paul Handford, Graham Bell, and Tom Reimchen. A gillnet fishery considered as an experiment in artificial selection. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):954–961, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-148.

Sprules:1977:CZC

[2315] W. Gary Sprules. Crustacean zooplankton communities as indicators of limnological conditions an approach using principal component analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):962–975, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-149.

Stirling:1977:DAS

[2316] Ian Stirling, W. Ralph Archibald, and Douglas DeMaster. Distribution and abundance of seals in the Eastern Beaufort Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):976–988, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-150.

McCaughran:1977:GMA

[2317] Donald A. McCaughran and Guy C. Powell. Growth model for Alaska king crab (Paralithodes camtschatica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):989–995, July 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-151.

Tsuyuki:1977:SSD

[2318] H. Tsuyuki and S. N. Williscroft. Swimming stamina differences between genotypically distinct forms of rainbow (Salmo gairdneri) and steelhead trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7): 996–1003, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-152.

Fournier:1977:PDN

[2319] Robert O. Fournier, John Marra, Richard Bohrer, and Mary Van Det. Plankton dynamics and nutrient enrichment of the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1004–1018, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-153.

McCormick:1977:WSC

[2320] J. Howard McCormick, Bernard R. Jones, and Kenneth E. F. Hokanson. White sucker (Catostomus commersoni) embryo development, and early growth and survival at different temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1019–1025, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-154.

Schiewe:1977:SPI

[2321] Michael H. Schiewe and Harold O. Hodgins. Specificity of protection induced in Coho salmon (Oncorhynchus kisutch) by heat-treated components of two pathogenic vibrios. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1026–1028, July 1977. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f77-155.

Brown:1977:FSB

[2322] T. J. Brown and J. R. Sibert. Food of some benthic harpacticoid copepods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1028– 1031, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-156.

Tanner:1977:HOS

[2323] C. Tanner, M. W. Hawkes, P. A. Lebednik, and E. Duffield. A handoperated suction sampler for the collection of subtidal organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1031–1034, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-157.

Zeitoun:1977:ESE

[2324] Ibrahim H. Zeitoun, Leroy D. Hughes, and Duane E. Ullrey. Effect of shock exposures of chlorine on the plasma electrolyte concentrations of adult rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1034–1039, July 1977. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f77-158.

Busch:1977:EWP

[2325] Wolf-Dieter N. Busch, David H. Davies, and Stephen J. Nepszy. Establishment of white perch, Morone americana, in Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1039–1041, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-159.

Beamish:1977:CSG

[2326] R. J. Beamish and R. R. Miller. Cytotaxonomic study of Gila trout, Salmo gilae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1041– 1045, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-160.

Fletcher:1977:TRS

[2327] H. F. Fletcher. Toward a relevant science: Fisheries and aquatic scientific resource needs in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1046–1074, July 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-161.

Bell:1977:DMA

[2328] G. R. Bell. Diseases of marine aquarium fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1075, July 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-162.

Snieszko:1977:DTF

[2329] S. F. Snieszko. The dermatology of teleost fishes: a bibliography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(7):1075, July 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-163.

Bayley:1977:MFL

[2330] Peter B. Bayley. A method for finding the limits of application of the von Bertalanffy growth model and statistical estimates of the parameters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1079–1084, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-164.

Tilzey:1977:RHB

[2331] R. D. J. Tilzey. Repeat homing of brown trout (Salmo trutta) in Lake Eucumbene, New South Wales, Australia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1085–1094, August 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-165.

Steele:1977:CTS

[2332] J. H. Steele, D. M. Farmer, and E. W. Henderson. Circulation and temperature structure in large marine enclosures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1095–1104, August 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-166.

Gascon:1977:DAR

[2333] Dominique Gascon and William C. Leggett. Distribution, abundance, and resource utilization of littoral zone fishes in response to a nutrient/production gradient in Lake Memphremagog. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1105–1117, August 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-167.

Fuller:1977:LFI

[2334] D. W. Fuller, K. S. Pilcher, and J. L. Fryer. A leukocytolytic factor isolated from cultures of Aeromonas salmonicida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1118–1125, August 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-168.

Stirling:1977:APS

[2335] Ian Stirling and W. Ralph Archibald. Aspects of predation of seals by polar bears. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1126–1129, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-169.

Peterman:1977:SMC

[2336] Randall M. Peterman. A simple mechanism that causes collapsing stability regions in exploited salmonid populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1130–1142, August 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-170.

Engelhardt:1977:UCP

[2337] F. Rainer Engelhardt, Joseph R. Geraci, and Thomas G. Smith. Uptake and clearance of petroleum hydrocarbons in the ringed seal, Phoca hispida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1143–1147, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-171.

McKim:1977:ETE

[2338] James M. McKim. Evaluation of tests with early life stages of fish for predicting long-term toxicity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1148–1154, August 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-172.

Lett:1977:SMM

[2339] Patrick F. Lett and Terje Benjaminsen. A stochastic model for the management of the Northwestern Atlantic harp seal (Pagophilus groenlandicus) population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8): 1155–1187, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-173.

Walker:1977:PEN

[2340] Roland Walker and Stuart W. Sherburne. Piscine erythrocytic necrosis virus in Atlantic cod, Gadus morhua, and other fish: Ultrastructure and distribution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8): 1188–1195, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-174.

Gardner:1977:AZP

[2341] Grant A. Gardner. Analysis of zooplankton population fluctuations in the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1196–1206, August 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-175.

Druehl:1977:IKR

[2342] Louis D. Druehl and Stephen I. C. Hsiao. Intertidal kelp response to seasonal environmental changes in a British Columbia inlet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1207–1211, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-176.

Momot:1977:RCO

[2343] Walter T. Momot and Howard Gowing. Response of the crayfish Orconectes virilis to exploitation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1212–1219, August 1977. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f77-177.

Copeman:1977:PDR

[2344] D. G. Copeman. Population differences in rainbow smelt, Osmerus mordax: Multivariate analysis of mensural and meristic data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1220–1229, August 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-178.

Sameoto:1977:MOC

[2345] D. D. Sameoto, L. O. Jaroszynski, and W. B. Fraser. A multiple opening and closing plankton sampler based on the MOCNESS and N.I.O. nets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1230–1235, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-179.

Reinitz:1977:EDR

[2346] Gary L. Reinitz. Electrophoretic distinction of rainbow trout (Salmo gairdneri), West-slope cutthroat trout (S. clarki), and their hybrids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1236–1239, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-180.

Charlton:1977:CBL

[2347] Murray N. Charlton. Carbon budgets in Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1240–1241, August 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-181.

Sonstegard:1977:TAA

[2348] R. Sonstegard, J. F. Leatherland, and R. Moccia. Thyroid activity of alewives (Alosa pseudoharengus) and rainbow smelts (Osmerus mordax) in the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8): 1242–1244, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-182.

Bullock:1977:UTW

[2349] G. L. Bullock and H. M. Stuckey. Ultraviolet treatment of water for destruction of five Gram-negative bacteria pathogenic to fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1244–1249, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-183.

Anonymous:1977:LLb

[2350] Anonymous. Letters/lettres. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1250–1252, August 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-184.

Low:1977:WRW

[2351] C. J. Low. Water resources of the world. Selected statistics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1253, August 1977. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f77-185.

Farmer:1977:EDM

[2352] D. Farmer. Encyclopaedic dictionary of mathematics for engineers and applied scientists. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8): 1253–1254, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-186.

Low:1977:ZFP

[2353] C. J. Low. Zooplankton fixation and preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1254, August 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-187.

Low:1977:IKF

[2354] C. J. Low. An illustrated key to freshwater and soil amoebae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(8):1254–1255, August 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-188.

Stasko:1977:RUB

[2355] A. B. Stasko and D. G. Pincock. Review of underwater biotelemetry, with emphasis on ultrasonic techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1261–1285, September 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-189.

Jones:1977:RED

[2356] Barry C. Jones and Glen H. Geen. Reproduction and embryonic development of spiny dogfish (Squalus acanthias) in the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9): 1286–1292, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-190.

Strand:1977:SPI

[2357] J. A. Strand, M. P. Fujihara, R. D. Burdett, and T. M. Poston. Suppression of the primary immune response in rainbow trout, Salmo gairdneri, sublethally exposed to tritiated water during embryogenesis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1293–1304, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-191.

Beamish:1977:ADL

[2358] Richard J. Beamish and D. Chilton. Age determination of lingcod (Ophiodon elongatus) using dorsal fin rays and scales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1305–1313, September 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-192.

Nelson:1977:EGB

[2359] Joseph S. Nelson. Evidence of a genetic basis for absence of the pelvic skeleton in brook stickleback, Culaea inconstans, and notes on the geographical distribution and origin of the loss. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1314–1320, September 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-193.

Jackson:1977:NPP

[2360] George A. Jackson. Nearctic and Palaearctic Paracladopelma Harnisch and Saetheria n. gen. (Diptera: Chironomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1321–1359, September 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-194.

Pietsch:1977:SIH

[2361] John P. Pietsch, Donald F. Amend, and Christine M. Miller. Survival of infectious hematopoietic necrosis virus held under various environmental conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1360–1364, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-195.

Cherry:1977:SHA

[2362] Donald S. Cherry, Stephan R. Larrick, Kenneth L. Dickson, Robert C. Hoehn, and John Cairns, Jr. Significance of hypochlorous acid in free residual chlorine to the avoidance response of spotted bass (Micropterus punctulatus) and rosyface shiner (Notropis rubellus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1365–1372, September 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-196.

Farmer:1977:IWT

[2363] G. J. Farmer, F. W. H. Beamish, and P. F. Lett. Influence of water temperature on the growth rate of the landlocked sea lamprey (Petromyzon marinus) and the associated rate of host mortality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1373–1378, September 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-197.

Simms:1977:SDS

[2364] S. E. Simms and P. A. Larkin. Simulation of dispersal of sockeye salmon (Oncorhynchus nerka) underyearlings in Babine Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1379–1388, September 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-198.

Gordon:1977:SJB

[2365] Michael R. Gordon and Donald J. McLeay. Sealed-jar bioassays for pulpmill effluent toxicity: Effects of fish species and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1389–1396, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-199.

Snyder:1977:IGS

[2366] Darrel E. Snyder, Maryann B. Mulhall Snyder, and Susan C. Douglas. Identification of golden shiner, Notemigonus crysoleucas, spotfin shiner, Notropis spilopterus, and fathead minnow, Pimephales promelas, larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1397–1409, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-200.

Wissmar:1977:IAP

[2367] R. C. Wissmar, J. E. Richey, and D. E. Spyridakis. The importance of allochthonous particulate carbon pathways in a subalpine lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1410–1418, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-201.

Olver:1977:RPL

[2368] C. H. Olver and C. A. Lewis. Reproduction of planted lake trout, Salvelinus namaycush, in gamitagama, a small Precambrian lake in Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1419–1422, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-202.

Leggett:1977:OMR

[2369] William C. Leggett. Ocean migration rates of American shad (Alosa sapidissima). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1422–1426, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-203.

Rebach:1977:SAR

[2370] Steve Rebach. Simultaneous activity recording of multiple isolated marine organisms in an artificial saltwater recirculating system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1426–1430, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-204.

Bilinski:1977:TBF

[2371] E. Bilinski, R. E. E. Jonas, Y. C. Lau, and G. Gibbard. Treatments before frozen storage affecting thaw drip formation in Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1431–1435, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-205.

Perkins:1977:OTS

[2372] Judith Perkins and Hal Whitehead. Observations on three species of baleen whales off Northern Newfoundland and adjacent waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(9):1436–1440, September 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-206.

Anonymous:1977:EF

[2373] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1445–1446, October 1977. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f77-207.

Colby:1977:IPP

[2374] Peter J. Colby. Introduction to the Proceedings of the 1976 Percid International Symposium (PERCIS). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1447–1449, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-208.

Collette:1977:SZF

[2375] Bruce B. Collette and Petru Bănărescu. Systematics and zoogeography of the fishes of the family Percidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1450–1463, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-209.

Weatherley:1977:PFA

[2376] A. H. Weatherley. Perca fluviatilis in Australia: Zoogeographic expression of a life cycle in relation to an environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1464–1466, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-210.

Ali:1977:RSF

[2377] M. A. Ali and M. Anctil. Retinal structure and function in the walleye (Stizostedion vitreum vitreum) and sauger (S. canadense). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1467–1474, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-211.

Ali:1977:PVP

[2378] M. A. Ali, R. A. Ryder, and M. Anctil. Photoreceptors and visual pigments as related to behavioral responses and preferred habitats of perches (Perca spp.) and pikeperches (Stizostedion spp.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1475–1480, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-212.

Ryder:1977:EAL

[2379] R. A. Ryder. Effects of ambient light variations on behavior of yearling, subadult, and adult walleyes (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1481–1491, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-213.

Disler:1977:SOL

[2380] N. N. Disler and S. A. Smirnov. Sensory organs of the lateral-line canal system in two percids and their importance in behavior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1492–1503, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-214.

Thorpe:1977:MPB

[2381] J. E. Thorpe. Morphology, physiology, behavior, and ecology of Perca fluviatilis L. and P. flavescens Mitchill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1504–1514, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-215.

Marshall:1977:MPE

[2382] T. R. Marshall. Morphological, physiological, and ethological differences between walleye (Stizostedion vitreum vitreum) and pikeperch (S. lucioperca). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1515–1523, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-216.

Hokanson:1977:TRS

[2383] Kenneth E. F. Hokanson. Temperature requirements of some percids and adaptations to the seasonal temperature cycle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1524–1550, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-217.

Spanovskaya:1977:DFA

[2384] V. D. Spanovskaya and V. A. Grygorash. Development and food of age-0 Eurasian perch (Perca fluviatilis) in reservoirs near Moscow, USSR. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1551–1558, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-218.

Popova:1977:FFR

[2385] O. A. Popova and L. A. Sytina. Food and feeding relations of Eurasian perch (Perca fluviatilis) and pikeperch (Stizostedion lucioperca) in various waters of the USSR. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1559–1570, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-219.

Pivnicka:1977:FAS

[2386] K. Pivnička and M. Švátora. Factors affecting the shift in predominance from Eurasian perch (Perca fluviatilis) to roach (Rutilus rutilus) in the Kličava Reservoir, Czechoslovakia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1571–1575, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-220.

Bonar:1977:RBE

[2387] Alina Bonar. Relations between exploitation, yield, and community structure in Polish pikeperch (Stizostedion lucioperca) lakes, 1966–71. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1576–1580, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-221.

Nagiec:1977:PSL

[2388] Maria Nagieć. Pikeperch (Stizostedion lucioperca) in its natural habitats in Poland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1581–1585, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-222.

Entz:1977:ECP

[2389] B. Entz. Environmental conditions of percid waters in Central Europe. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1586–1591, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-223.

Johnson:1977:LCO

[2390] M. G. Johnson, J. H. Leach, C. K. Minns, and C. H. Olver. Limnological characteristics of Ontario lakes in relation to associations of walleye (Stizostedion vitreum vitreum), northern pike (Esox lucius), lake trout (Salvelinus namaycush), and smallmouth bass (Micropterus dolomieui). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1592– 1601, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f77-224.

Carlander:1977:BPY

[2391] Kenneth D. Carlander. Biomass, production, and yields of walleye (Stizostedion vitreum vitreum) and yellow perch (Perca flavescens) in North American lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1602–1612, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-225.

Adams:1977:YPS

[2392] G. F. Adams and C. H. Olver. Yield properties and structure of boreal percid communities in Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1613–1625, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-226.

Johnson:1977:IBW

[2393] Fritz H. Johnson and John G. Hale. Interrelations between walleye (Stizostedion vitreum vitreum) and smallmouth bass (Micropterus dolomieui) in four Northeastern Minnesota lakes, 1948–69. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1626–1632, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-227.

Johnson:1977:RWS

[2394] Fritz H. Johnson. Responses of walleye (Stizostedion vitreum vitreum) and yellow perch (Perca flavescens) populations to removal of white sucker (Catostomus commersoni) from a Minnesota lake, 1966. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1633–1642, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-228.

Swenson:1977:FCW

[2395] William A. Swenson. Food consumption of walleye (Stizostedion vitreum vitreum) and sauger (S. canadense) in relation to food availability and physical conditions in Lake of the Woods, Minnesota, Shagawa Lake, and Western Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1643–1654, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-229.

Kelso:1977:UPP

[2396] J. R. M. Kelso and F. J. Ward. Unexploited percid populations of West Blue Lake, Manitoba, and their interactions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1655–1669, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-230.

Hartmann:1977:PLC

[2397] J. Hartmann and W. Nümann. Percids of Lake Constance, a lake undergoing eutrophication. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1670–1677, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-231.

Biro:1977:EEI

[2398] Péter Biró. Effects of exploitation, introductions, and eutrophication on percids in Lake Balaton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1678–1683, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-232.

Lind:1977:RPS

[2399] Esko A. Lind. A review of pikeperch (Stizostedion lucioperca), Eurasian perch (Perca fluviatilis), and ruff (Gymnocephalus cernua) in Finland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1684–1695, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-233.

Chevalier:1977:CWS

[2400] J. R. Chevalier. Changes in walleye (Stizostedion vitreum vitreum) population in Rainy Lake and factors in abundance, 1924–75. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1696–1702, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-234.

Anthony:1977:FDC

[2401] D. D. Anthony and C. R. Jorgensen. Factors in the declining contributions of walleye (Stizostedion vitreum vitreum) to the fishery of Lake Nipissing, Ontario, 1960–76. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1703–1709, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-235.

Willemsen:1977:PDP

[2402] J. Willemsen. Population dynamics of percids in Lake IJssel and some smaller lakes in the Netherlands. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1710–1719, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-236.

Rundberg:1977:THP

[2403] Hugo Rundberg. Trends in harvests of pikeperch (Stizostedion lucioperca), Eurasian perch (Perca fluviatilis), and northern pike (Esox lucius) and associated environmental changes in Lakes Mälaren and Hjälmaren, 1914–74. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1720–1724, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-237.

Momot:1977:MWS

[2404] Walter T. Momot, Jack Erickson, and Frederick Stevenson. Maintenance of a walleye, Stizostedion vitreum vitreum, fishery in a eutrophic reservoir. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1725–1733, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-238.

Holcik:1977:CFC

[2405] Juraj Holčík. Changes in fish community of Klíčava Reservoir with particular reference to Eurasian perch (Perca fluviatilis), 1957–72. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1734–1747, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-239.

Nelson:1977:PDY

[2406] William R. Nelson and Charles H. Walburg. Population dynamics of yellow perch (Perca flavescens), sauger (Stizostedion canadense), and walleye (S. vitreum vitreum) in four main stem Missouri River reservoirs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1748–1763, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-240.

Bagenal:1977:EFE

[2407] T. B. Bagenal. Effects of fisheries on Eurasian perch (Perca fluviatilis) in Windermere. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1764–1768, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-241.

Elsey:1977:EWS

[2408] C. A. Elsey and R. T. Thomson. Exploitation of walleye (Stizostedion vitreum vitreum) in Lac des mille lacs, Northern Ontario, by commercial and sport fisheries, 1958–75. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1769–1773, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-242.

Smith:1977:WSV

[2409] Lloyd L. Smith, Jr. Walleye (Stizostedion vitreum vitreum) and yellow perch (Perca flavescens) populations and fisheries of the Red Lakes, Minnesota, 1930–75. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1774–1783, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-243.

Schupp:1977:TPY

[2410] D. H. Schupp and V. Macins. Trends in percid yields from Lake of the Woods, 1888–1973. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1784–1791, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-244.

Carlander:1977:YCA

[2411] Kenneth D. Carlander and Philip Michael Payne. Year-class abundance, population, and production of walleye (Stizostedion vitreum vitreum) in Clear Lake, Iowa, 1948–74, with varied fry stocking rates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1792–1799, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-245.

Kempinger:1977:DWS

[2412] James J. Kempinger and Robert F. Carline. Dynamics of the walleye (Stizostedion vitreum vitreum) population in Escanaba Lake, Wisconsin, 1955–72. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1800–1811, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-246.

Forney:1977:EII

[2413] John L. Forney. Evidence of inter- and IntraSpecific competition as factors regulating walleye (Stizostedion vitreum vitreum) biomass in Oneida Lake, New York. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1812–1820, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-247.

Wells:1977:CYP

[2414] LaRue Wells. Changes in yellow perch (Perca flavescens) populations of Lake Michigan, 1954–75. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1821–1829, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-248.

Eshenroder:1977:EIF

[2415] R. L. Eshenroder. Effects of intensified fishing, species changes, and spring water temperatures on yellow perch, Perca flavescens, in Saginaw Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1830–1838, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-249.

Spangler:1977:PCW

[2416] G. R. Spangler, N. R. Payne, and G. K. Winterton. Percids in the Canadian Waters of Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1839–1848, October 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-250.

Hurley:1977:DWF

[2417] D. A. Hurley and W. J. Christie. Depreciation of the warmwater fish community in the Bay of Quinte, Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1849–1860, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-251.

Nepszy:1977:CPP

[2418] Stephen J. Nepszy. Changes in percid populations and species interactions in Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1861–1868, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-252.

Johnston:1977:PDW

[2419] D. A. Johnston. Population dynamics of walleye (Stizostedion vitreum vitreum) and yellow perch (Perca flavescens) in Lake St. Clair, especially during 1970–76. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1869–1877, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-253.

Schneider:1977:WSV

[2420] J. C. Schneider and J. H. Leach. Walleye (Stizostedion vitreum vitreum) fluctuations in the Great Lakes and possible causes, 1800–1975. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1878–1889, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-254.

Collette:1977:BP

[2421] Bruce B. Collette, M. A. Ali, K. E. F. Hokanson, Maria Nagieć, S. A. Smirnov, J. E. Thorpe, A. H. Weatherley, and J. Willemsen. Biology of the percids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1890–1899, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-255.

Koonce:1977:FIY

[2422] J. F. Koonce, T. B. Bagenal, R. F. Carline, K. E. F. Hokanson, and M. Nagieć. Factors influencing year-class strength of percids: A summary and a model of temperature effects. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1900–1909, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-256.

Balon:1977:RGP

[2423] Eugene K. Balon, Walter T. Momot, and Henry A. Regier. Reproductive guilds of percids: Results of the paleogeographical history and ecological succession. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1910–1921, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-257.

Kitchell:1977:ABM

[2424] James F. Kitchell, Donald J. Stewart, and David Weininger. Applications of a bioenergetics model to yellow perch (Perca flavescens) and walleye (Stizostedion vitreum vitreum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1922–1935, October 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-258.

Kitchell:1977:PHR

[2425] James F. Kitchell, Murray G. Johnson, C. Kenneth Minns, Kenneth H. Loftus, Lorne Greig, and Charles H. Olver. Percid habitat: The river analogy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1936–1940, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-259.

MacLean:1977:SIP

[2426] James MacLean and John J. Magnuson. Species interactions in percid communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10): 1941–1951, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-260.

Kerr:1977:NTP

[2427] S. R. Kerr and R. A. Ryder. Niche theory and percid community structure. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1952–1958, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-261.

Kelso:1977:PUE

[2428] J. R. M. Kelso and T. B. Bagenal. Percids in unperturbed ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1959–1963, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-262.

Leach:1977:RPF

[2429] J. H. Leach, M. G. Johnson, J. R. M. Kelso, J. Hartmann, W. Nümann, and B. Entz. Responses of percid fishes and their habitats to eutrophication. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1964–1971, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-263.

Shuter:1977:DMW

[2430] B. J. Shuter and J. F. Koonce. A dynamic model of the Western Lake Erie walleye (Stizostedion vitreum vitreum) population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1972–1982, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-264.

Spangler:1977:RPE

[2431] G. R. Spangler, N. R. Payne, J. E. Thorpe, J. M. Byrne, H. A. Regier, and W. J. Christie. Responses of percids to exploitation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1983–1988, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-265.

Kerr:1977:STF

[2432] S. R. Kerr. Structure and transformation of fish production systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1989–1994, October 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-266.

Anonymous:1977:PIS

[2433] Anonymous. Percid International Symposium (PERCIS). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(10):1995–1999, October 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-267.

Dermott:1977:PCP

[2434] R. M. Dermott, J. Kalff, W. C. Leggett, and J. Spence. Production of Chironomus, Procladius, and Chaoborus at different levels of phytoplankton biomass in Lake Memphremagog, Quebec–Vermont. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2001–2007, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-268.

Bailey:1977:EPM

[2435] R. F. J. Bailey, K. W. Able, and W. C. Leggett. Evidence for the presence of a metamorphic check in capelin (Mallotus villosus) otoliths and implications for age determination. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2008–2014, November 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-269.

Jacquaz:1977:SDA

[2436] Bernadette Jacquaz, Kenneth W. Able, and William C. Leggett. Seasonal distribution, abundance, and growth of larval capelin (Mallotus villosus) in the St. Lawrence Estuary and Northwestern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2008–2014, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-270.

Bailey:1977:SVD

[2437] R. F. J. Bailey, K. W. Able, and W. C. Leggett. Seasonal and vertical distribution and growth of juvenile and adult capelin (Mallotus villosus) in the St. Lawrence Estuary and Western Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2015–2029, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-271.

Momot:1977:PPD

[2438] Walter T. Momot and Howard Gowing. Production and population dynamics of the crayfish Orconectes virilis in three Michigan lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2030–2040, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-272.

Momot:1977:REF

[2439] Walter T. Momot and Howard Gowing. Results of an experimental fishery on the crayfish Orconectes virilis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2041–2055, November 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-273.

Jones:1977:FFS

[2440] Barry C. Jones and Glen H. Geen. Food and feeding of spiny dogfish (Squalus acanthias) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2056–2066, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-274.

Ryan:1977:GRB

[2441] Patrick M. Ryan and Harold H. Harvey. Growth of rock bass, Ambloplites rupestris, in relation to the morphoedaphic index as an indicator of an environmental stress. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2067–2078, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-275.

Jarvinen:1977:LTT

[2442] A. W. Jarvinen, M. J. Hoffman, and T. W. Thorslund. Long-term toxic effects of DDT food and water exposure on fathead minnows (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2079–2088, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-276.

Thompson:1977:BCB

[2443] R. J. Thompson. Blood chemistry, biochemical composition, and the annual reproductive cycle in the giant scallop, Placopecten magellanicus, from Southeast Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2089–2103, November 1977. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f77-277.

Greisman:1977:NDS

[2444] Paul Greisman and Grant Ingram. Nutrient distribution in the St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2104–2116, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-278.

DeAngelis:1977:RFP

[2445] D. L. DeAngelis, S. W. Christensen, and A. G. Clark. Responses of a fish population model to young-of-the-year mortality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2117–2123, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-279.

Williams:1977:SAB

[2446] D. Dudley Williams, J. H. Mundie, and D. E. Mounce. Some aspects of benthic production in a salmonid rearing channel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2124–2132, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-280.

Schweigert:1977:EFF

[2447] Jacob F. Schweigert, F. J. Ward, and J. W. Clayton. Effects of fry and fingerling introductions on walleye (Stizostedion vitreum vitreum) production in West Blue Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2133–2141, November 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-281.

Campbell:1977:OMO

[2448] C. M. Campbell and D. R. Idler. Oocyte maturation and ovulation induced in hypophysectomized winter flounder (Pseudopleuronectes americanus) by preparations from pituitary glands of American plaice (Hippoglossoides platessoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2142–2150, November 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-282.

Paradis:1977:IIC

[2449] M. Paradis and R. G. Ackman. Influence of ice cover and man on the odd-chain hydrocarbons and fatty acids in the waters of Jeddore Harbour, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2151–2155, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-283.

McLeay:1977:LSH

[2450] Donald J. McLeay and Michael R. Gordon. Leucocrit: A simple hematological technique for measuring acute stress in salmonid fish, including stressful concentrations of pulpmill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2156–2163, November 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-284.

Smith:1977:CDF

[2451] E. A. H. Smith, C. I. Mayfield, and P. T. S. Wong. Colonization and decomposition of fish bone material in natural and synthetic aqueous solutions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2164–2175, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-285.

Ballintijn:1977:EPV

[2452] C. M. Ballintijn, D. D. Beatty, and R. L. Saunders. Effects of pseudobranchectomy on visual pigment density and ocular PO_{2} in Atlantic salmon, Salmo salar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2176–2184, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-286.

Khan:1977:BCA

[2453] R. A. Khan. Blood changes in Atlantic cod (Gadus morhua) infected with Trypanosoma murmanensis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2185–2192, November 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-287.

Geraci:1977:MSL

[2454] J. R. Geraci and D. J. St. Aubin. Mass stranding of the long-finned pilot whale, Globicephala melaena, on Sable Island, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2193–2196, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-288.

Prentice:1977:TRS

[2455] Earl F. Prentice and John E. Rensel. Tag retention of the spot prawn, Pandalus platyceros, injected with coded wire tags. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2196–2199, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-289.

Evans:1977:SPA

[2456] Patricia D. Evans and Kenneth H. Mann. Selection of prey by American lobsters (Homarus americanus) when offered a choice between sea urchins and crabs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2199–2203, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-290.

Hilton:1977:EAA

[2457] J. W. Hilton, C. Y. Cho, and S. J. Slinger. Evaluation of the ascorbic acid status of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2203–2207, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-291.

Barica:1977:EFM

[2458] J. Barica. Effect of freeze-up on major ion and nutrient content of a prairie winterkill lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2207–2210, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-292.

Johnson:1977:ALP

[2459] David W. Johnson and Dwight A. Webster. Avoidance of low pH in selection of spawning sites by brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2210–2215, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-293.

Levy:1977:TAD

[2460] E. M. Levy, C. C. Cunningham, C. D. W. Conrad, and J. D. Moffatt. A titration apparatus for the determination of dissolved oxygen in seawater. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2215–2218, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-294.

Anonymous:1977:CFC

[2461] Anonymous. Comments on “Fish catches and environment”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2218–2220, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-295.

Anonymous:1977:RCF

[2462] Anonymous. Reply to comments on “Fish catches and environment”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2221, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-296.

Anonymous:1977:CAE

[2463] Anonymous. Comments on “Aquatic environmental quality: problems and proposals”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11): 2221–2222, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f77-297.

Anonymous:1977:YCR

[2464] Anonymous. You can’t really get there from here. Reply to comments on “Aquatic environmental quality: problems and proposals”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2222–2225, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-298. See [2135].

Withler:1977:MRF

[2465] F. C. Withler. Marine recreational fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2225–2229, November 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-299.

Welch:1977:AAM

[2466] H. E. Welch. Assessment of the Arctic marine environment/selected topics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2230–2231, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-300.

Gillespie:1977:AAO

[2467] D. C. Gillespie. Aquatic applications of ozone. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2231–2232, November 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-301.

Stell:1977:BRR

[2468] W. K. Stell. Book review: Retinas of Fishes. An Atlas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2232, November 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/

10.1139/f77-302. See comment [2687] and reply [2688].

McAllister:1977:CZF

[2469] C. D. McAllister. Coastal zone: Framework for management in Atlantic Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(11):2232–2233, November 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-303.

Oglesby:1977:PSS

[2470] Ray T. Oglesby. Phytoplankton summer standing crop and annual productivity as functions of phosphorus loading and various physical factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2255–2270, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-304.

Oglesby:1977:RFY

[2471] Ray T. Oglesby. Relationships of fish yield to lake phytoplankton standing crop, production, and morphoedaphic factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2271–2279, December 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-305.

Brooks:1977:VSD

[2472] Arthur S. Brooks and Byron G. Torke. Vertical and seasonal distribution of chlorophyll a in Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2280–2287, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-306.

Thorne:1977:NDH

[2473] Richard E. Thorne. A new digital hydroacoustic data processor and some observations on herring in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2288–2294, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-307.

Sonntag:1977:IIM

[2474] Nicholas C. Sonntag and Wulf Greve. Investigation of the impact of mercury on enclosed water columns using a zooplankton simulation model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2295–2307, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-308.

Ware:1977:STE

[2475] D. M. Ware. Spawning time and egg size of Atlantic mackerel, Scomber scombrus, in relation to the plankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2308–2315, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-309.

Rulifson:1977:TWV

[2476] Roger A. Rulifson. Temperature and water velocity effects on the swimming performances of young-of-the-year striped mullet (Mugil cephalus), spot (Leiostomus xanthurus), and pinfish (Lagodon rhomboides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2316–2322, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-310.

Broderius:1977:RTF

[2477] Steven J. Broderius, Lloyd L. Smith, Jr., and David T. Lind. Relative toxicity of free cyanide and dissolved sulfide forms to the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2323–2332, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-311.

Reinitz:1977:TAT

[2478] Gary L. Reinitz. Tests for association of transferrin and lactate dehydrogenase phenotypes with weight gain in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2333–2337, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-312.

Crossman:1977:AHB

[2479] E. J. Crossman and James W. Meade. Artificial hybrids between Amur pike, Esox reicherti, and North American esocids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2338–2343, December 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-313.

Sheldon:1977:SPF

[2480] R. W. Sheldon, W. H. Sutcliffe, Jr., and M. A. Paranjape. Structure of pelagic food chain and relationship between plankton and fish production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2344–2353, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-314.

Saether:1977:HHG

[2481] Ole A. Sæther. Habrobaenus hudsoni n. gen., n. sp. and the immatures of Baeoctenus bicolor Sæther (Diptera: Chironomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2354–2361, December 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-315.

Wenner:1977:BMF

[2482] Charles A. Wenner and John A. Musick. Biology of the morid fish Antimora rostrata in the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2362–2368, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-316.

Hiltz:1977:SLR

[2483] Doris Fraser Hiltz, D. H. North, Barbara Smith Lall, and R. A. Keith. Storage life of refrozen silver hake (Merluccius bilinearis) processed as fillets am minced flesh from thawed, stored, round-frozen fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2369–2373, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-317.

Gales:1977:SCF

[2484] Lawrence E. Gales, James A. Buss, and L. J. Bledsoe. Simulation concepts for fishery systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2374–2380, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-318.

Poulet:1977:GMC

[2485] S. A. Poulet. Grazing of marine copepod developmental stages on naturally occurring particles. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2381–2387, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-319.

Peterson:1977:CSG

[2486] R. H. Peterson and J. L. Metcalfe. Changes in specific gravity of Atlantic salmon (Salmo salar) alevins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2388–2395, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-320.

Winters:1977:MAL

[2487] G. H. Winters. Migrations and activity levels of overwintering Atlantic herring (Clupea harengus harengus) along Southwest Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2396–2401, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-321.

Winters:1977:HWL

[2488] G. H. Winters. Healing of wounds and location of tags in Atlantic herring (Clupea harengus harengus) released with abdominally inserted magnetic tags. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2402–2404, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-322.

Smith:1977:EPA

[2489] E. A. Smith, C. I. Mayfield, and P. T. S. Wong. Effects of phosphorus from apatite on development of freshwater communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2405–2409, December 1977. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f77-323.

Busdosh:1977:SDR

[2490] M. Busdosh and A. E. Manning. A simple device for retrieving instruments or traps set under ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2409–2410, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-324.

Tyler:1977:EAI

[2491] A. V. Tyler. Evidence for the assumption of independence between gastric emptying rate and swimming activity using Atlantic cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2411–2413, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-325.

Lin:1977:RIR

[2492] C. Kwei Lin and John L. Blum. Recent invasion of a red alga (Bangia atropurpurea) in Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2413–2416, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-326.

Johnson:1977:YPR

[2493] Eric A. Johnson, Douglas E. Conklin, and Michael J. Lewis. The yeast Phaffia rhodozyma as a dietary pigment source for salmonids and crustaceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2417–2421, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-327.

White:1977:DTP

[2494] Alan W. White. Dinoflagellate toxins as probable cause of an Atlantic herring (Clupea harengus harengus) kill, and pteropods as apparent vector. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2421–2424, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-328.

Pickavance:1977:BRN

[2495] J. R. Pickavance. Book review: Nomenclature Oligochaetologica: A Catalogue of Names, Descriptions and Type Specimens of the Oligochaeta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2425, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f77-329.

Fee:1977:RRD

[2496] Everett J. Fee. Recovery and restoration of damaged ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2425–2426, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-330.

Webster:1977:HCD

[2497] G. R. Barrie Webster. Herbicides: Chemistry, degradation and mode of action. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2426, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-331.

Bardach:1977:FES

[2498] John E. Bardach. Farming the edge of the sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2427, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-332.

Cauvin:1977:EIE

[2499] Dennis M. Cauvin. Economic impacts of extended fisheries jurisdiction. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2427–2429, December 1977. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f77-333.

Yurkowski:1977:PFI

[2500] M. Yurkowski. Proceedings of the First International Conference on Aquaculture Nutrition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2429–2430, December 1977. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f77-334.

Resh:1977:IFC

[2501] Vincent H. Resh. Inland fishes of California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2430–2431, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-335.

Winterbottom:1977:FW

[2502] Richard Winterbottom. Fishes of the world. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 34(12):2431–2432, December 1977. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f77-336.

Groberg:1978:RWT

[2503] W. J. Groberg, Jr., R. H. McCoy, K. S. Pilcher, and J. L. Fryer. Relation of water temperature to infections of Coho salmon (Oncorhynchus kisutch), Chinook salmon (O. tshawytscha), and steelhead trout (Salmo gairdneri) with Aeromonas salmonicida and A. hydrophila. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):1–7, January 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-001.

Sanders:1978:RWT

[2504] J. E. Sanders, K. S. Pilcher, and J. L. Fryer. Relation of water temperature to bacterial kidney disease in Coho salmon (Oncorhynchus kisutch), sockeye salmon (O. nerka), and steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):8–11, January 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-002.

Keast:1978:FIB

[2505] Allen Keast. Feeding interrelations between age-groups of pumpkinseed (Lepomis gibbosus) and comparisons with bluegill (L. macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):12–27, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-003.

Stockner:1978:EAP

[2506] John G. Stockner and K. R. S. Shortreed. Enhancement of autotrophic production by nutrient addition in a coastal rainforest stream on Vancouver Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1): 28–34, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-004.

McCarty:1978:TCG

[2507] L. S. McCarty, J. A. C. Henry, and A. H. Houston. Toxicity of cadmium to goldfish, Carassius auratus, in hard and soft water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):35–42, January 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-005.

Lear:1978:CVS

[2508] W. H. Lear and R. K. Misra. Clinal variation in scale characters of Atlantic salmon (Salmo salar) based on discriminant function analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):43–47, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-006.

Craig:1978:MSR

[2509] P. C. Craig. Movements of stream-resident and anadromous Arctic char (Salvelinus alpinus) in a perennial spring on the Canning River, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):48–52, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-007.

Power:1978:FPS

[2510] G. Power. Fish population structure in Arctic lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):53–59, January 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-008.

Chadwick:1978:AGA

[2511] E. M. P. Chadwick, T. R. Porter, and P. Downton. Analysis of growth of Atlantic salmon (Salmo salar) in a small Newfoundland River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):60–68, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-009.

Alderdice:1978:RBT

[2512] D. F. Alderdice and F. P. J. Velsen. Relation between temperature and incubation time for eggs of Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):69–75, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-010.

Fried:1978:SMH

[2513] Stephen M. Fried, James D. McCleave, and George W. LaBar. Seaward migration of hatchery-reared Atlantic salmon, Salmo salar, smolts in the Penobscot River Estuary, Maine: Riverine movements. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):76–87, January 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-011.

Seegert:1978:DWF

[2514] Gregory L. Seegert and Arthur S. Brooks. Dechlorination of water for fish culture: Comparison of the activated carbon, sulfite reduction, and photochemical methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35 (1):88–92, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-012.

Farmer:1978:SAP

[2515] G. J. Farmer, J. A. Ritter, and D. Ashfield. Seawater adaptation and Parr–smolt transformation of juvenile Atlantic salmon, Salmo salar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):93–100, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-013.

Morrison:1978:DMR

[2516] C. M. Morrison and P. H. Odense. Distribution and morphology of the rodlet cell in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35 (1):101–116, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-014.

Cross:1978:GVA

[2517] T. F. Cross and R. H. Payne. Geographic variation in Atlantic cod, Gadus morhua, off Eastern North America: A biochemical systematics approach. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):117–123, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-015.

Sharp:1978:UMM

[2518] Joan C. Sharp, Kenneth W. Able, William C. Leggett, and James E. Carscadden. Utility of meristic and morphometric characters for identification of capelin (Mallotus villosus) stocks in Canadian Atlantic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):124–130, January 1978.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-016.

Keenleyside:1978:RIB

[2519] Miles H. A. Keenleyside. Reproductive isolation between pumpkinseed (Lepomis gibbosus) and longear sunfish (L. megalotis) (Centrarchidae) in the Thames River, Southwestern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):131–135, January 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-017.

Peterson:1978:EDS

[2520] Charles H. Peterson and Brian P. Bradley. Estimating the diet of a sluggish predator from field observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):136–141, January 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-018.

Barica:1978:URC

[2521] J. Barica. Unusual response of a cupric ion electrode in prairie Lake Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):141–143, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-019.

FitzGerald:1978:TTS

[2522] Gerard J. FitzGerald and Miles H. A. Keenleyside. Technique for tagging small fish with I^{131} for evaluation of predator–prey relationships. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):143–145, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-020.

Naevdal:1978:VAF

[2523] Gunnar Nævdal, Marianne Holm, Oscar Ingebrigtsen, and Dag Møller. Variation in age at first spawning in Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):145–147, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-021.

Reno:1978:USP

[2524] Paul W. Reno, Marie Philippon-Fried, Bruce L. Nicholson, and Stuart W. Sherburne. Ultrastructural studies of piscine erythrocytic necrosis (PEN) in Atlantic herring (Clupea harengus harengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):148–154, January 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-022.

Gordon:1978:BBD

[2525] M. J. Gordon, B. K. Swan, and C. G. Paterson. Baeoctenus bicolor (Diptera: Chironomidae) parasitic in unionid bivalve molluscs, and notes on other chironomid–bivalve associations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):154–157, January 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-023.

Kilambi:1978:ECC

[2526] Raj V. Kilambi, James C. Adams, and William A. Wickizer. Effects of cage culture on growth, abundance, and survival of resident largemouth bass (Micropterus salmoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):157–160, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-024.

Scherer:1978:FL

[2527] E. Scherer. Fishing with light. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):161, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-025.

Hamilton:1978:GN

[2528] A. L. Hamilton. Grassy narrows. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):161–162, January 1978. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f78-026.

Hamilton:1978:NSP

[2529] A. L. Hamilton. No safe place. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(1):161–162, January 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-027.

Leduc:1978:DEC

[2530] Gérard Leduc. Deleterious effects of cyanide on early life stages of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):166–174, February 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-028.

Symons:1978:SHB

[2531] P. E. K. Symons and M. Heland. Stream habitats and behavioral interactions of underyearling and yearling Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):175–183, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-029.

Westrheim:1978:BUA

[2532] S. J. Westrheim and W. E. Ricker. Bias in using an age–length key to estimate age-frequency distributions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):184–189, February 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-030.

Schindler:1978:PIC

[2533] D. W. Schindler, E. J. Fee, and T. Ruszczynski. Phosphorus input and its consequences for phytoplankton standing crop and production in the Experimental Lakes Area and in similar lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):190–196, February 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-031.

Marten:1978:CMR

[2534] Gerald G. Marten. Calculating mortality rates and optimum yields from length samples. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2): 197–201, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-032.

Whyte:1978:DPM

[2535] John N. C. Whyte and John R. Englar. Direct potentiometric measurement of chloride in fisheries products by ion-selective electrode analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):202–207, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-033.

Silvert:1978:PKF

[2536] William Silvert. The price of knowledge: Fisheries management as a research tool. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2): 208–212, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-034.

Medford:1978:PLC

[2537] B. A. Medford and W. C. Mackay. Protein and lipid content of gonads, liver, and muscle of northern pike (Esox lucius) in relation to gonad growth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):213–219, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-035.

Ware:1978:BPF

[2538] D. M. Ware. Bioenergetics of pelagic fish: Theoretical change in swimming speed and ration with body size. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):220–228, February 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-036.

Walter:1978:SYM

[2539] Gilbert G. Walter. A surplus yield model incorporating recruitment and applied to a stock of Atlantic mackerel (Scomber scombrus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):229–234, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-037.

Everitt:1978:MPM

[2540] Robert R. Everitt, Nicholas C. Sonntag, Martin L. Puterman, and Patrick Whalen. A mathematical programming model for the management of a renewable resource system: The kemano II development project. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):235–246, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-038.

Eidt:1978:MIL

[2541] D. C. Eidt and J. H. Meating. Method for incubating leaf tissue to study decomposition in streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):247–248, February 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-039.

Janssen:1978:FBR

[2542] John Janssen. Feeding-behavior repertoire of the alewife, Alosa pseudoharengus, and the ciscoes Coregonus hoyi and C. artedii. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):249–253, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-040.

Dawson:1978:SFS

[2543] L. E. Dawson, K. L. Uebersax, and M. A. Uebersax. Stability of freshwater sucker (Catostomus spp.) flesh during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):253–257, February 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-041.

Huntley:1978:MPP

[2544] M. E. Huntley and L. A. Hobson. Medusa predation and plankton dynamics in a temperate fjord, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):257–261, February 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-042.

Nebeker:1978:SST

[2545] Alan V. Nebeker, James D. Andros, Joel K. McCrady, and Donald G. Stevens. Survival of steelhead trout (Salmo gairdneri) eggs, embryos, and fry in air-supersaturated water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):261–264, February 1978. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f78-043.

Brunskill:1978:BRE

[2546] G. J. Brunskill. Book review: Environmental Biogeochemistry. Volume 1: Carbon, Nitrogen Phosphorus, Sulfur and Selenium Cycles. Volume 2: Metals Transfer and Ecological Mass Balances. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):265–266, February 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-044.

Sinclair:1978:EFM

[2547] William F. Sinclair. The economics of fisheries management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):266–267, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-045.

Pharo:1978:MCS

[2548] C. H. Pharo. Mineral cycling in southeastern ecosystems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):267, February 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-046.

Parsons:1978:PO

[2549] T. R. Parsons. Principles of oceanography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):267–268, February 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-047.

Giles:1978:BBP

[2550] M. A. Giles. Biochemical and biophysical perspectives in marine biology: Volume I. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):268–269, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-048.

Flannagan:1978:LNA

[2551] John F. Flannagan. Larvae of the North American caddisfly genera (Trichoptera). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):270, February 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-049.

Parsons:1978:LEO

[2552] T. R. Parsons. Laboratory exercises in oceanography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(2):270, February 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-050.

Loucks:1978:SFP

[2553] R. H. Loucks and W. H. Sutcliffe, Jr. A simple fish-population model including environmental influence, for two Western Atlantic shelf stocks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):279–285, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-051.

Conway:1978:SAC

[2554] H. L. Conway. Sorption of arsenic and cadmium and their effects on growth, micronutrient utilization, and photosynthetic pigment composition of Asterionella formosa. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):286–294, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-052.

Berezay:1978:BRC

[2555] Gordon Berezay and John H. Gee. Buoyancy response to changes in water velocity and its function in creek chub (Semotilus atromaculatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):295–299, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-053.

Scheider:1978:APB

[2556] W. A. Scheider. Applicability of phosphorus budget models to small Precambrian lakes, Algonquin Park, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):300–304, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-054.

Lewis:1978:CNC

[2557] A. G. Lewis. Concentrations of nutrients and chlorophyll on a crosschannel transect in Juan de Fuca Strait, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):305–314, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-055.

Peters:1978:CKP

[2558] Robert Henry Peters. Concentrations and kinetics of phosphorus fractions in water from streams entering Lake Memphremagog. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):315–328, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-056.

Rimmer:1978:FRA

[2559] D. M. Rimmer and G. Power. Feeding response of Atlantic salmon (Salmo salar) alevins in flowing and still water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):329–332, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-057.

Rabeni:1978:CTM

[2560] Charles F. Rabeni and K. Elizabeth Gibbs. Comparison of two methods used by divers for sampling benthic invertebrates in deep rivers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):332–336, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-058.

Hanic:1978:OMP

[2561] L. A. Hanic and J. D. Pringle. Outplant method for phenological studies of Chondrus crispus in mechanically harvested beds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):336–338, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-059.

Cossa:1978:STM

[2562] D. Cossa and S. A. Poulet. Survey of trace metal contents of suspended matter in the St. Lawrence Estuary and Saguenay Fjord. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):338–345, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-060.

Strange:1978:AHS

[2563] Richard J. Strange and Carl B. Schreck. Anesthetic and handling stress on survival and cortisol concentration in yearling Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):345–349, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-061.

Vallentyne:1978:FLT

[2564] J. R. Vallentyne. Facing the long term: An inquiry into opportunities to improve the climate for research with reference to limnology in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):350–369, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-062.

Anonymous:1978:O

[2565] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):370–371, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-063.

Cahn:1978:I

[2566] Phyllis H. Cahn. Ichthyology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):372, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-064.

Clarke:1978:BMW

[2567] R. McV. Clarke. Biological monitoring of water and effluent quality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):372–373, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-065.

Beamish:1978:BRI

[2568] R. J. Beamish. Book review: Impact of Acid Precipitation on Forest and Freshwater Ecosystems in Norway. Research Report 6/76. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):373–374, March 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-066.

Scott:1978:EN

[2569] D. P. Scott. Editor’s note. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(3):374, March 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-067.

Olson:1978:TUS

[2570] Kenneth R. Olson, Katherine S. Squibb, and Robert J. Cousins. Tissue uptake, subcellular distribution, and metabolism of ^{14}CH_{3}HgCl and CH_{3} ^{203}HgCl by rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):381–390, April 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-068.

Fraser:1978:CRP

[2571] J. M. Fraser. Comparative recoveries of planted yearling and fallfingerling brook trout (Salvelinus fontinalis) from Ontario lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):391–396, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-069.

White:1978:PTD

[2572] Alan W. White and Lucie Maranda. Paralytic toxins in the dinoflagellate Gonyaulax excavata and in shellfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):397–402, April 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-070.

Strout:1978:PVC

[2573] R. G. Strout, E. S. Sawyer, and B. A. Coutermarsh. Pathogenic vibrios in confinement-reared and feral fishes of the Maine–New Hampshire Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):403–408, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-071.

Chou:1978:PSN

[2574] C. L. Chou, J. F. Uthe, and E. G. Zook. Polarographic studies on the nature of cadmium in scallop, oyster, and lobster. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):409–413, April 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-072.

Houghton:1978:SMC

[2575] Robert W. Houghton, Peter C. Smith, and Robert O. Fournier. A simple model for cross-shelf mixing on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):414–421, April 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-073.

Weiler:1978:ICD

[2576] R. R. Weiler and J. O. Nriagu. Isotopic composition of dissolved inorganic carbon in the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):422–430, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-074.

Hilton:1978:EGL

[2577] J. W. Hilton, C. Y. Cho, and S. J. Slinger. Effect of graded levels of supplemental ascorbic acid in practical diets fed to rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):431– 436, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-075.

Hodson:1978:PIC

[2578] Peter V. Hodson, Beverly R. Blunt, and Douglas J. Spry. pH-induced changes in blood lead of lead-exposed rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):437–445, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-076.

Sullivan:1978:CPP

[2579] J. F. Sullivan, G. J. Atchison, D. J. Kolar, and A. W. McIntosh. Changes in the predator–prey behavior of fathead minnows (Pimephales promelas) and largemouth bass (Micropterus salmoides) caused by cadmium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):446–451, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-077.

Botta:1978:EDF

[2580] J. R. Botta and D. H. Shaw. Effect of double freezing and subsequent long-term refrozen storage at −23 ^{o}C on the quality of inshore male capelin (Mallotus villosus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):452–456, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-078.

OKeefe:1978:SSC

[2581] Timothy M. O’Keefe and Richard L. Noble. Storage stability of channel catfish (Ictalurus punctatus) in relation to dietary level of α-tocopherol. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):457–460, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-079.

Marshall:1978:PDD

[2582] J. S. Marshall. Population dynamics of Daphnia galeata mendotae as modified by chronic cadmium stress. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):461–469, April 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-080.

Law:1978:FCC

[2583] W. M. Law, R. D. Ellender, J. H. Wharton, and B. L. Middlebrooks. Fish cell culture: Properties of a cell line from the sheepshead, Archosargus probatocephalus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4): 470–473, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-081.

Bilinski:1978:CSD

[2584] E. Bilinski, R. E. E. Jonas, and Y. C. Lau. Chill stowage and development of rancidity in frozen Pacific herring, Clupea harengus pallasi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):473–477, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-082.

Payan:1978:CEB

[2585] P. Payan, P. Pic, and G. De Renzis. Comparaison des échanges branchiaux de Cl^{−} en eau douce chez Salmo gairdneri in vivo et in vitro sur la tête isolée perfusée. (French) [Comparison of branchial exchanges of Cl^{−} in freshwater in Salmo gairdneri in vivo and in vitro on the isolated perfused head]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4): 477–479, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-083.

Wong:1978:TMM

[2586] P. T. S. Wong, Y. K. Chau, and P. L. Luxon. Toxicity of a mixture of metals on freshwater algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):479–481, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-084.

Anonymous:1978:CDV

[2587] Anonymous. Comment on the distribution of vertebral number in North American sculpins of the Myoxocephalus quadricornis complex. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):482–483, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-085.

Anonymous:1978:EDR

[2588] Anonymous. Electrophoretic distinction in rainbow (Salmo gairdneri) and cutthroat (S. clarki) trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):483, April 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-086.

Schindler:1978:BRE

[2589] D. W. Schindler and G. J. Brunskill. Book review: Environmental Toxicity of Aquatic Radionuclides. Models and Mechanisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):484, April 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-087.

Healey:1978:BD

[2590] F. P. Healey. The biology of diatoms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):485–486, April 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-088.

Armstrong:1978:NSF

[2591] F. A. J. Armstrong. New strategic factors in the North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):486, April 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-089.

Armstrong:1978:PIN

[2592] F. A. J. Armstrong. The political implications of North Sea oil and gas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):486, April 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-090.

Armstrong:1978:NPF

[2593] F. A. J. Armstrong. Navigation primer for fishermen. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(4):486–488, April 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-091.

Anonymous:1978:EF

[2594] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):499–500, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-092.

Farrington:1978:MDM

[2595] J. W. Farrington. IN MEMORIAM: Dr max blumer organic geochemist and environmental scientist, 1923–77. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):501–504, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-093.

Vandermeulen:1978:ISR

[2596] J. H. Vandermeulen. Introduction to the Symposium on Recovery Potential of Oiled Marine Northern Environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):505–508, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-094.

Teal:1978:AAH

[2597] John M. Teal, Kathryn Burns, and John Farrington. Analyses of aromatic hydrocarbons in intertidal sediments resulting from two spills of no. 2 fuel oil in buzzards Bay, Massachusetts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):510–520, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-095.

Cretney:1978:LTF

[2598] W. J. Cretney, C. S. Wong, D. R. Green, and C. A. Bawden. Longterm fate of a heavy fuel oil in a spill-contaminated B.C. coastal bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):521–527, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-096.

Keizer:1978:RBC

[2599] P. D. Keizer, T. P. Ahern, J. Dale, and J. H. Vandermeulen. Residues of bunker C oil in Chedabucto Bay, Nova Scotia, 6 years after the Arrow spill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):528– 535, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-097.

VanVleet:1978:CCP

[2600] Edward S. Van Vleet and James G. Quinn. Contribution of chronic petroleum inputs to Narragansett Bay and Rhode Island Sound sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):536– 543, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-098.

Mackie:1978:PAP

[2601] P. R. Mackie, R. Hardy, and K. J. Whittle. Preliminary assessment of the presence of oil in the ecosystem at ekofisk after the blowout, April 22–30, 1977. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5): 544–551, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-099.

Mayo:1978:WCP

[2602] D. W. Mayo, D. S. Page, J. Cooley, E. Sorenson, F. Bradley, E. S. Gilfillan, and S. A. Hanson. Weathering characteristics of petroleum hydrocarbons deposited in fine clay marine sediments, Searsport, Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):552–562, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-100.

Owens:1978:MDO

[2603] E. H. Owens. Mechanical dispersal of oil stranded in the littoral zone. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):563–572, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-101.

Colwell:1978:MES

[2604] R. R. Colwell, A. L. Mills, J. D. Walker, P. Garcia-Tello, and V. Campos-P. Microbial ecology studies of the Metula spill in the Straits of Magellan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):573–580, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-102.

Stewart:1978:DAH

[2605] James E. Stewart and L. J. Marks. Distribution and abundance of hydrocarbon-utilizing bacteria in sediments of Chedabucto Bay, Nova Scotia, in 1976. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5): 581–584, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-103.

Atlas:1978:PCO

[2606] R. M. Atlas, A. Horowitz, and M. Busdosh. Prudhoe crude oil in Arctic marine ice, water, and sediment ecosystems: Degradation and interactions with microbial and benthic communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):585–590, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-104.

Gordon:1978:ISW

[2607] Donald C. Gordon, Jr., Jacqueline Dale, and Paul D. Keizer. Importance of sediment working by the deposit-feeding polychaete Arenicola marina on the weathering rate of sediment-bound oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):591–603, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-105.

Butler:1978:SSO

[2608] James N. Butler and Eric M. Levy. Session I. Summary and overview: Long-term fate of petroleum hydrocarbons after spills — compositional changes and microbial degradation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):604–605, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-106.

Roesijadi:1978:UHM

[2609] G. Roesijadi, J. W. Anderson, and J. W. Blaylock. Uptake of hydrocarbons from marine sediments contaminated with Prudhoe Bay crude oil: Influence of feeding type of test species and availability of polycyclic aromatic hydrocarbons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5): 608–614, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-107.

Grahl-Nielsen:1978:AHB

[2610] O. Grahl-Nielsen, J. T. Staveland, and S. Wilhelmsen. Aromatic hydrocarbons in benthic organisms from coastal areas polluted by Iranian crude oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):615– 623, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-108.

Cole:1978:PEG

[2611] Timothy J. Cole. Preliminary ecological–genetic comparison between unperturbed and oil-impacted Urosalpinx cinerea (Prosobranchia: Gastropoda) populations: Nobska Point (Woods Hole) and Wild Harbor (West Falmouth), Massachusetts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):624–629, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-109.

Gilfillan:1978:AGP

[2612] E. S. Gilfillan and J. H. Vandermeulen. Alterations in growth and physiology of soft-shell clams, Mya arenaria, chronically oiled with bunker C from Chedabucto Bay, Nova Scotia, 1970–76. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):630–636, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-110.

Stainken:1978:EUD

[2613] Dennis M. Stainken. Effects of uptake and discharge of petroleum hydrocarbons on the respiration of the soft-shell clam, Mya arenaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):637–642, May 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-111.

Vandermeulen:1978:AAH

[2614] J. H. Vandermeulen and W. R. Penrose. Absence of aryl hydrocarbon hydroxylase (AHH) in three marine bivalves. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):643–647, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-112.

Krebs:1978:LTE

[2615] C. T. Krebs and K. A. Burns. Long-term effects of an oil spill on populations of the salt-marsh crab Uca pugnax. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):648–649, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-113.

Percy:1978:ECE

[2616] J. A. Percy. Effects of chronic exposure to petroleum upon the growth and molting of juveniles of the Arctic marine isopod crustacean Mesidotea entomon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5): 650–656, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-114.

McCain:1978:BCO

[2617] B. B. McCain, H. O. Hodgins, W. D. Gronlund, J. W. Hawkes, D. W. Brown, M. S. Myers, and J. H. Vandermeulen. Bioavailability of crude oil from experimentally oiled sediments to English sole (Parophrys vetulus), and pathological consequences. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):657–664, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-115.

Payne:1978:PCM

[2618] J. F. Payne, J. W. Kiceniuk, W. R. Squires, and G. L. Fletcher. Pathological changes in a marine fish after a 6-month exposure to petroleum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):665–667, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-116.

Stegeman:1978:IEC

[2619] John J. Stegeman. Influence of environmental contamination on cytochrome P-450 mixed-function oxygenases in fish: Implications for recovery in the wild harbor marsh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):668–674, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-117.

Clark:1978:OSR

[2620] R. B. Clark. Oiled seabird rescue and conservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):675–678, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-118.

Anderson:1978:SIS

[2621] J. W. Anderson and D. C. Malins. Session II. Summary and overview: Physiological stresses and response in chronically oiled organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):679–680, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-119.

Southward:1978:RRS

[2622] A. J. Southward and Eve C. Southward. Recolonization of rocky shores in Cornwall after use of toxic dispersants to clean up the Torrey Canyon spill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):682–706, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-120.

Thomas:1978:COU

[2623] Martin L. H. Thomas. Comparison of oiled and unoiled intertidal communities in Chedabucto Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):707–716, May 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-121.

Sanders:1978:FOS

[2624] Howard L. Sanders. Florida oil spill impact on the Buzzards Bay benthic fauna: West Falmouth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):717–730, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-122.

Hampson:1978:NFO

[2625] George R. Hampson and Edwin T. Moul. No. 2 fuel oil spill in Bourne, Massachusetts: Immediate assessment of the effects on marine invertebrates and a 3-year study of growth and recovery of a salt marsh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):731–744, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-123.

Notini:1978:LTE

[2626] Mats Notini. Long-term effects of an oil spill on Fucus macrofauna in a small Baltic Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5): 745–753, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-124.

Clark:1978:OCW

[2627] Robert C. Clark, Jr., Benjamin G. Patten, and Edward E. DeNike. Observations of a cold-water intertidal community after 5 years of a lowlevel, persistent oil spill from the General M. C. Meigs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):754–765, May 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-125.

Leppakoski:1978:RBM

[2628] E. J. Leppäkoski and L. S. Lindström. Recovery of benthic macrofauna from chronic pollution in the sea area off a refinery plant, Southwest Finland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):766– 775, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-126.

Anderson:1978:RBA

[2629] J. W. Anderson, R. G. Riley, and R. M. Bean. Recruitment of benthic animals as a function of petroleum hydrocarbon concentrations in the sediment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5): 776–790, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-127.

Mann:1978:SIS

[2630] K. H. Mann and R. B. Clark. SESSION III. Summary and overview: Long-term effects of oil spills on marine intertidal communities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(5):791–795, May 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-128.

Peterman:1978:EFR

[2631] Randall M. Peterman and Marino Gatto. Estimation of functional responses of predators on juvenile salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):797–808, June 1978. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f78-129.

Dillon:1978:APS

[2632] P. J. Dillon, D. S. Jeffries, W. Snyder, R. Reid, N. D. Yan, D. Evans, J. Moss, and W. A. Scheider. Acidic precipitation in South-Central Ontario: Recent observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):809–815, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-130.

Brett:1978:MRE

[2633] J. R. Brett and J. M. Blackburn. Metabolic rate and energy expenditure of the spiny dogfish, Squalus acanthias. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):816–821, June 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-131.

Wedemeyer:1978:PTN

[2634] Gary A. Wedemeyer and W. T. Yasutake. Prevention and treatment of nitrite toxicity in juvenile steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):822–827, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-132.

Markiw:1978:MCF

[2635] Maria E. Markiw and Ken Wolf. Myxosoma cerebralis: Fluorescent antibody techniques for antigen recognition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):828–832, June 1978. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f78-133.

Botsford:1978:BAS

[2636] Louis W. Botsford and Daniel E. Wickham. Behavior of age-specific, density-dependent models and the Northern California Dungeness crab (Cancer magister) fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):833–843, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-134.

Power:1978:PFS

[2637] G. Power and Jean Gregoire. Predation by freshwater seals on the fish community of Lower Seal Lake, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):844–850, June 1978. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f78-135.

Robinson:1978:PES

[2638] D. G. Robinson and W. E. Barraclough. Population estimates of sockeye salmon (Oncorhynchus nerka) in a fertilized oligotrophic lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):851–860, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-136.

Eldridge:1978:FBE

[2639] Maxwell B. Eldridge, Tina Echeverria, and Sid Korn. Fate of^{14}C-benzene in eggs and larvae of Pacific herring (Clupea harengus pallasi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):861–865, June 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-137.

Wong:1978:WTF

[2640] S. L. Wong, B. Clark, M. Kirby, and R. F. Kosciuw. Water temperature fluctuations and seasonal periodicity of Cladophora and Potamogeton in shallow rivers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6): 866–870, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-138.

Fender:1978:HIF

[2641] D. C. Fender and D. F. Amend. Hyperosmotic infiltration: Factors influencing uptake of bovine serum albumin by rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):871– 874, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-139.

Wedemeyer:1978:SSV

[2642] Gary A. Wedemeyer, Nancy C. Nelson, and Cathy A. Smith. Survival of the salmonid viruses infectious hematopoietic necrosis (IHNV) and infectious pancreatic necrosis (IPNV) in ozonated, chlorinated, and untreated waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6): 875–879, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-140.

Brown:1978:RBC

[2643] D. A. Brown and T. R. Parsons. Relationship between cytoplasmic distribution of mercury and toxic effects to zooplankton and chum salmon (Oncorhynchus keta) exposed to mercury in a controlled ecosystem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):880–884, June 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-141.

Minns:1978:SDN

[2644] C. K. Minns, J. R. M. Kelso, and W. Hyatt. Spatial distribution of nearshore fish in the vicinity of two thermal generating stations, Nanticoke and Douglas Point, on the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):885–892, June 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-142.

Fagerlund:1978:DDR

[2645] U. H. M. Fagerlund and J. R. McBride. Distribution and disappearance of radioactivity in blood and tissues of Coho salmon (Oncorhynchus kisutch) after oral administration of^{3}H-testosterone. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):893–900, June 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-143.

Brodie:1978:ASD

[2646] P. F. Brodie. Alternative sampling device for aquatic organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):901–902, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-144.

Evelyn:1978:VEN

[2647] T. P. T. Evelyn and G. S. Traxler. Viral erythrocytic necrosis: Natural occurrence in Pacific salmon and experimental transmission. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):903–907, June 1978. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f78-145.

Symons:1978:LBJ

[2648] Philip E. K. Symons. Leaping behavior of juvenile Coho (Oncorhynchus kisutch) and Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):907–909, June 1978. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f78-146.

Fivizzani:1978:MSD

[2649] Albert J. Fivizzani and Richard E. Spieler. Modified staaland device with automatic recording techniques for determining salinity preference in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):910– 912, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-147.

DeMaster:1978:CAA

[2650] Douglas P. DeMaster. Calculation of the average age of sexual maturity in marine mammals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6): 912–915, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-148.

Loftus:1978:FPS

[2651] K. H. Loftus, M. G. Johnson, and H. A. Regier. Federal–provincial strategic planning for Ontario fisheries: Management strategy for the 1980s. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):916–927, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-149.

Mathias:1978:HPW

[2652] J. A. Mathias. Harvesting polluted waters: Waste heat and nutrientloaded effluents in the aquaculture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):928, June 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-150.

Ayles:1978:APT

[2653] G. Burton Ayles. Aquaculture practices in Taiwan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):929, June 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-151.

Iredale:1978:FSR

[2654] D. G. Iredale. Fishexpo ’76 Seminar, Report of Proceedings, Melbourne, 21–23 Sept. 1976. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6): 929–930, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-152.

Breen:1978:PSG

[2655] Paul A. Breen. Pacific seashores: A guide to intertidal ecology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):930–931, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-153.

Johnson:1978:EFB

[2656] Lionel Johnson. Energy flow — its biological dimension. A summary of I.B.P. in Canada, 1964–1974. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(6):931–933, June 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-154.

Healey:1978:FCE

[2657] M. C. Healey. Fecundity changes in exploited populations of lake whitefish (Coregonus clupeaformis) and lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):945–950, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-155.

Yurkowski:1978:ARP

[2658] M. Yurkowski, J. K. Bailey, R. E. Evans, Jo-Anne L. Tabachek, G. Burton Ayles, and J. G. Eales. Acceptability of rapeseed proteins in diets of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):951–962, July 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-156.

McFarlane:1978:EHM

[2659] G. A. McFarlane and W. G. Franzin. Elevated heavy metals: A stress on a population of white suckers, Catostomus commersoni, in Hamell Lake, Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7): 963–970, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-157.

Bisson:1978:DFS

[2660] Peter A. Bisson. Diel food selection by two sizes of rainbow trout (Salmo gairdneri) in an experimental stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):971–975, July 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-158.

Botta:1978:CSA

[2661] J. R. Botta, P. B. Noonan, and J. T. Lauder. Chemical and sensory analysis of ungutted offshore (nonspawning) capelin (Mallotus villosus) stored in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7): 976–980, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f78-159.

Moser:1978:LPJ

[2662] H. Geoffrey Moser and Elbert H. Ahlstrom. Larvae and pelagic juveniles of blackgill rockfish, Sebastes melanostomus, taken in midwater trawls off Southern California and Baja California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):981–996, July 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-160.

DeFoe:1978:EAF

[2663] D. L. DeFoe, G. D. Veith, and R. W. Carlson. Effects of Aroclor(R) 1248 and 1260 on the fathead minnow (Pimephales promelas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):997–1002, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-161.

Kutty:1978:AQS

[2664] M. N. Kutty. Ammonia quotient in sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):1003– 1005, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-162.

Hirtle:1978:DCV

[2665] R. W. M. Hirtle and K. H. Mann. Distance chemoreception and vision in the selection of prey by American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):1006–1008, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-163.

Kelly:1978:FCL

[2666] R. K. Kelly, B. W. Souter, and H. R. Miller. Fish cell lines: Comparisons of CHSE-214, FHM, and RTG-2 in assaying IHN and IPN viruses. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):1009–1011, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-164.

Bower:1978:IAS

[2667] Carol E. Bower and Joseph P. Bidwell. Ionization of ammonia in seawater: Effects of temperature, pH, and salinity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):1012–1016, July 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-165.

Sinclair:1978:MAS

[2668] William F. Sinclair. Management alternatives and strategic planning for Canada’s fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35 (7):1017–1030, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-166.

Andersen:1978:NHS

[2669] R. Andersen. The need for human sciences research in Atlantic Coast fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):1031– 1049, July 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-167.

McMahon:1978:TEI

[2670] J. W. McMahon. Thermal ecology II. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(7):1050, July 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-168.

DiMichele:1978:HPR

[2671] Leonard DiMichele and Malcolm H. Taylor. Histopathological and physiological responses of Fundulus heteroclitus to naphthalene exposure. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1060–1066, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-169.

Smith:1978:CVD

[2672] Peter C. Smith, Brian Petrie, and C. R. Mann. Circulation, variability, and dynamics of the Scotian Shelf and slope. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1067–1083, August 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-170.

Mauck:1978:EPB

[2673] Wilbur L. Mauck, Paul M. Mehrle, and Foster L. Mayer. Effects of the polychlorinated biphenyl Aroclor(R) 1254 on growth, survival, and bone development in brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1084–1088, August 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-171.

deMarch:1978:PSN

[2674] L. de March. Permanent sedimentation of nitrogen, phosphorus, and organic carbon in a High Arctic lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1089–1094, August 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-172.

Liao:1978:SCN

[2675] C. F.-H. Liao and D. R. S. Lean. Seasonal changes in nitrogen compartments of lakes under different loading conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1095–1101, August 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-173.

Liao:1978:NTW

[2676] C. F.-H. Liao and D. R. S. Lean. Nitrogen transformations within the trophogenic zone of lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1102–1108, August 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-174.

Borgmann:1978:RMG

[2677] U. Borgmann, O. Kramar, and C. Loveridge. Rates of mortality, growth, and biomass production of Lymnaea palustris during chronic exposure to lead. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1109–1115, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-175.

Petrie:1978:COB

[2678] Brian Petrie and Ken Drinkwater. Circulation in an open bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1116–1123, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-176.

Schreck:1978:SRC

[2679] Carl B. Schreck and Harold W. Lorz. Stress response of Coho salmon (Oncorhynchus kisutch) elicited by cadmium and copper and potential use of cortisol as an indicator of stress. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1124–1129, August 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-177.

Estes:1978:EAS

[2680] James A. Estes and James R. Gilbert. Evaluation of an aerial survey of Pacific walruses (Odobenus rosmarus divergens). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1130–1140, August 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-178.

Baross:1978:IME

[2681] John A. Baross, Patricia A. Tester, and Richard Y. Morita. Incidence, microscopy, and etiology of exoskeleton lesions in the Tanner crab, Chionoecetes tanneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8): 1141–1149, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-179.

Levenson:1978:DHW

[2682] Coleman Levenson and William T. Leapley. Distribution of humpback whales (Megaptera novaeangliae) in the Caribbean determined by a rapid acoustic method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8): 1150–1152, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-180.

Boyd:1978:DSV

[2683] Carl M. Boyd and John Marra. A device to simulate the variability in sunlight intensity for laboratory cultures of algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1152–1154, August 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-181.

Orr:1978:ADD

[2684] M. H. Orr, E. E. Hays, and F. R. Hess. Acoustic detection of demersal fish closer than 15 centimetres to the bottom in 80 metres of water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1155–1156, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-182.

Anonymous:1978:CFS

[2685] Anonymous. Comment on “Food of some benthic harpacticoid copepods”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1157, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-183.

Anonymous:1978:RCF

[2686] Anonymous. Reply to comment on “Food of some benthic harpacticoid copepods.”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1158, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-184.

Anonymous:1978:CBR

[2687] Anonymous. Comment on book review — “Retinas of fishes. An atlas”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1158–1159, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-185. See [2468, 2688].

Anonymous:1978:RCB

[2688] Anonymous. Reply to Comment on book review — “Retinas of fishes. An atlas”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(8):1159, August 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-186. See [2468, 2687].

Morgan:1978:LHA

[2689] Mark D. Morgan and Alfred M. Beeton. Life history and abundance of Mysis relicta in Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1165–1170, September 1978. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f78-187.

Sinclair:1978:SPV

[2690] Michael Sinclair. Summer phytoplankton variability in the Lower St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9): 1171–1185, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-188.

Griffin:1978:MCD

[2691] Billy R. Griffin and Eric M. Davis. Myxosoma cerebralis: Detection of circulating antibodies in infected rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1186–1190, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-189.

Teleki:1978:AEU

[2692] G. C. Teleki and A. J. Chamberlain. Acute effects of underwater construction blasting on fishes in long point Bay, Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1191–1198, September 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-190.

Bowering:1978:FWF

[2693] W. R. Bowering. Fecundity of witch flounder (Glyptocephalus cynoglossus) from St. Pierre Bank and the Grand Bank of Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1199–1206, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-191.

Sameoto:1978:ZSV

[2694] D. D. Sameoto. Zooplankton sample variation on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1207–1222, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-192.

Fleming:1978:RTG

[2695] Lesley C. Fleming and Michael D. B. Burt. Revision of the turbellarian genus Ectocotyla (Seriata, Monocelididae) associated with the crabs Chionoecetes opilio and Hyas araneus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1223–1233, September 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-193.

Young:1978:UAC

[2696] Charity L. Young and George B. Chapman. Ultrastructural aspects of the causative agent and renal histopathology of bacterial kidney disease in brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1234–1248, September 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-194.

Winters:1978:PMS

[2697] G. H. Winters. Production, mortality, and sustainable yield of Northwest Atlantic harp seals (Pagophilus groenlandicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1249–1261, September 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-195.

Dawson:1978:RMD

[2698] V. K. Dawson, P. D. Harman, D. P. Schultz, and J. L. Allen. Rapid method for determining concentrations of Bayer 73 in water during lampricide treatments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9): 1262–1265, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-196.

Niimi:1978:LAB

[2699] A. J. Niimi. Lag adjustment between estimated and actual physiological responses conducted in flow-through systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1265–1269, September 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-197.

Sutterlin:1978:PSM

[2700] A. M. Sutterlin, P. Harmon, and B. Young. Precocious sexual maturation in Atlantic salmon (Salmo salar) postsmolts reared in a seawater impoundment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9): 1269–1272, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-198.

Davis:1978:AEE

[2701] J. C. Davis. Adaptation to environment: Essays on the physiology of marine animals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35 (9):1273, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-199.

Vandermeulen:1978:CEB

[2702] John H. Vandermeulen. Coelenterate ecology and behavior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1273–1274, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-200.

Marshall:1978:IST

[2703] K. E. Marshall. Index to scientific and technical proceedings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(9):1274, September 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-201.

Leatherland:1978:LST

[2704] J. F. Leatherland and R. A. Sonstegard. Lowering of serum thyroxine and triiodothyronine levels in yearling Coho salmon, Oncorhynchus kisutch, by dietary Mirex and PCBs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1285–1289, October 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-202.

Colberg:1978:EOM

[2705] P. J. Colberg and A. J. Lingg. Effect of ozonation on microbial fish pathogens, ammonia, nitrate, nitrite, and BOD in simulated reuse hatchery water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1290–1296, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-203.

Curry:1978:EEG

[2706] Kevin D. Curry and Boyd Kynard. Effect of extended galvanonarcosis on behavior of rainbow trout, Salmo gairdneri, and channel catfish, Ictalurus punctatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10): 1297–1302, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-204.

Walters:1978:MEE

[2707] Carl J. Walters, Ray Hilborn, Randall M. Peterman, and Michael J. Staley. Model for examining early ocean limitation of Pacific salmon production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1303–1315, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-205.

OBrien:1978:NCM

[2708] R. N. O’Brien, S. Visaisouk, R. Raine, and D. F. Alderdice. Natural convection: A mechanism for transporting oxygen to incubating salmon eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1316–1321, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-206.

Cooley:1978:PED

[2709] John M. Cooley and Charles K. Minns. Prediction of egg development times of freshwater copepods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1322–1329, October 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-207.

McLaren:1978:GLS

[2710] I. A. McLaren. Generation lengths of some temperate marine copepods: Estimation, prediction, and implications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1330–1342, October 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-208.

Fisher:1978:AAC

[2711] William S. Fisher and Richard T. Nelson. Application of antibiotics in the cultivation of Dungeness crab, Cancer magister. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1343–1349, October 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-209.

Geraci:1978:PMA

[2712] J. R. Geraci, M. D. Dailey, and D. J. St. Aubin. Parasitic mastitis in the Atlantic white-sided dolphin, Lagenorhynchus acutus, as a probable factor in herd productivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1350–1355, October 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-210.

Beverley-Burton:1978:HAT

[2713] M. Beverley-Burton. Helminths of the alimentary tract from a stranded herd of the Atlantic white-sided dolphin, Lagenorhynchus acutus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1356–1359, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-211.

Meehan:1978:SFE

[2714] William R. Meehan and Richard A. Miller. Stomach flushing: Effectiveness and influence on survival and condition of juvenile salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1359–1363, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-212.

Seepersad:1978:UAB

[2715] B. Seepersad and R. W. Crippen. Use of aniline blue for distinguishing between live and dead freshwater zooplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1363–1366, October 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-213.

Sosnowski:1978:TCN

[2716] S. L. Sosnowski and J. H. Gentile. Toxicological comparison of natural and cultured populations of Acartia tonsa to cadmium, copper, and mercury. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1366–1369, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-214.

Drenner:1978:CPR

[2717] Ray W. Drenner, J. Rudi Strickler, and W. John O’Brien. Capture probability: The role of zooplankter escape in the selective feeding of planktivorous fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10): 1370–1373, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-215.

Kwain:1978:FEE

[2718] Wen hwa Kwain and J. A. Chappel. First evidence for even-year spawning pink salmon, Oncorhynchus gorbuscha, in Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1373–1376, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-216.

Li:1978:HHC

[2719] M. F. Li, G. S. Traxler, and S. Clyburne. Histopathological and hematological changes in mummichogs, Fundulus heteroclitus, infected by a pseudomonad. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10): 1376–1381, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-217.

Iverson:1978:SEI

[2720] Ernest W. Iverson. Status of Exosphaeroma inornata Dow and E. media George and Stromberg (Isopoda: Sphaeromatidae) with ecological notes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1381–1384, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-218.

Elner:1978:MBR

[2721] R. W. Elner and A. B. Stasko. Mating behavior of the rock crab, Cancer irroratus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10):1385–1388, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-219.

McLeay:1978:ESP

[2722] Donald J. McLeay and Michael R. Gordon. Effect of seasonal photoperiod on acute toxic responses of juvenile rainbow trout (Salmo gairdneri) to pulpmill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(10): 1388–1392, October 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-220.

Norstrom:1978:CSO

[2723] R. J. Norstrom, D. J. Hallett, and R. A. Sonstegard. Coho salmon (Oncorhynchus kisutch) and herring gulls (Larus argentatus) as indicators of organochlorine contamination in Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1401–1409, November 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-221.

Brezonik:1978:EOC

[2724] Patrick L. Brezonik. Effect of organic color and turbidity of secchi disk transparency. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11): 1410–1416, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-222.

Webb:1978:TEA

[2725] P. W. Webb. Temperature effects on acceleration of rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1417–1422, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-223.

Jalabert:1978:PIO

[2726] Bernard Jalabert, Frederick W. Goetz, Bernard Breton, Alexis Fostier, and Edward M. Donaldson. Precocious induction of oocyte maturation and ovulation in Coho salmon, Oncorhynchus kisutch. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1423–1429, November 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-224.

Kwain:1978:EAO

[2727] Wen-Hwa Kwain and Robert W. McCauley. Effects of age and overhead illumination on temperatures preferred by underyearling rainbow trout, Salmo gairdneri, in a vertical temperature gradient. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1430–1433, November 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-225.

Peterman:1978:TDD

[2728] Randall M. Peterman. Testing for density-dependent marine survival in Pacific salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11): 1434–1450, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-226.

Reno:1978:IPN

[2729] P. W. Reno, S. Darley, and M. Savan. Infectious pancreatic necrosis: Experimental induction of a carrier state in trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1451–1456, November 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-227.

Selgeby:1978:PRS

[2730] James H. Selgeby, Wayne R. MacCallum, and Donald V. Swedberg. Predation by rainbow smelt (Osmerus mordax) on lake herring (Coregonus artedii) in Western Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1457–1463, November 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-228.

Hall:1978:IGL

[2731] Lenwood W. Hall, Jr., Charles H. Hocutt, and Jay R. Stauffer, Jr. Implication of geographic location on temperature preference of white perch, Morone americana. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11): 1464–1468, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-229.

Leggett:1978:LVR

[2732] William C. Leggett and James E. Carscadden. Latitudinal variation in reproductive characteristics of American shad (Alosa sapidissima): Evidence for population specific life history strategies in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1469–1478, November 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-230.

Cloern:1978:BGM

[2733] James E. Cloern and Frederic H. Nichols. A von Bertalanffy growth model with a seasonally varying coefficient. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1479–1482, November 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-231.

Briand:1978:CBS

[2734] Frédéric Briand, Ramiro Trucco, and S. Ramamoorthy. Correlations between specific algae and heavy metal binding in lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1482–1485, November 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-232.

Bouck:1978:ECS

[2735] G. R. Bouck, M. A. Cairns, and A. R. Christian. Effect of capture stress on plasma enzyme activities in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1485–1488, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-233.

Eddy:1978:FMH

[2736] S. B. Eddy and M. W. Lankester. Feeding and migratory habits of Arctic char, Salvelinus alpinus, indicated by the presence of the swimbladder nematode Cystidicola cristivomeri White. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1488–1491, November 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-234.

Hartwick:1978:REO

[2737] E. B. Hartwick, P. A. Breen, and L. Tulloch. A removal experiment with Octopus dofleini (Wulker). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1492–1495, November 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-235.

Kettle:1978:VAZ

[2738] Dean Kettle and W. John O’Brien. Vulnerability of Arctic zooplankton species to predation by small lake trout (Salvelinus namaycush). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1495–1500, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-236.

Oliveira:1978:CSE

[2739] Luis Oliveira, Naval J. Antia, and Thana Bisalputra. Culture studies on the effects from fluoride pollution on the growth of marine phytoplankters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1500–1504, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-237.

Mori:1978:HBA

[2740] Katsuyoshi Mori and James E. Stewart. The hemolymph bactericidin of American lobster (Homarus americanus): Adsorption and activation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1504–1507, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-238.

Shaffer:1978:TEE

[2741] Edward H. Shaffer. Trading off environment, economics and energy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1508, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f78-239.

Boalch:1978:C

[2742] G. T. Boalch. The coastline. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(11):1508, November 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-240.

Able:1978:ISL

[2743] Kenneth W. Able. Ichthyoplankton of the St. Lawrence Estuary: Composition, distribution, and abundance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1518–1531, December 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-241.

Ostrofsky:1978:MPR

[2744] M. L. Ostrofsky. Modification of phosphorus retention models for use with lakes with low areal water loading. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1532–1536, December 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-242.

Rombough:1978:LHT

[2745] P. J. Rombough, S. E. Barbour, and J. J. Kerekes. Life history and taxonomic status of an isolated population of Arctic char, Salvelinus alpinus, from Gros Morne National Park, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1537–1541, December 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-243.

Saunders:1978:CGA

[2746] Richard L. Saunders and Eugene B. Henderson. Changes in gill ATPase activity and smolt status of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1542–1546, December 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-244.

Walton:1978:PIM

[2747] D. G. Walton, W. R. Penrose, and J. M. Green. The petroleuminducible mixed-function oxidase of cunner (Tautogolabrus adspersus Walbaum 1792): Some characteristics relevant to hydrocarbon monitoring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1547–1552, December 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-245.

Eggers:1978:LWE

[2748] D. M. Eggers, N. W. Bartoo, N. A. Rickard, R. E. Nelson, R. C. Wissmar, R. L. Burgner, and A. H. Devol. The Lake Washington ecosystem: The perspective from the fish community production and forage base. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1553–1571, December 1978. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f78-246.

Morgan:1978:IIK

[2749] Mark D. Morgan, Stephen T. Threlkeld, and Charles R. Goldman. Impact of the introduction of Kokanee (Oncorhynchus nerka) and opossum shrimp (Mysis relicta) on a subalpine lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1572–1579, December 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-247.

LeBrasseur:1978:ESS

[2750] R. J. LeBrasseur, C. D. McAllister, W. E. Barraclough, O. D. Kennedy, J. Manzer, D. Robinson, and K. Stephens. Enhancement of sockeye salmon (Oncorhynchus nerka) by lake fertilization in Great Central Lake: Summary report. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12): 1580–1596, December 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-248.

Nakashima:1978:DRY

[2751] Brian S. Nakashima and William C. Leggett. Daily ration of yellow perch (Perca flavescens) from Lake Memphremagog, Quebec–Vermont, with a comparison of methods for in situ determinations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1597–1603, December 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-249.

Hargrave:1978:ODM

[2752] B. T. Hargrave and S. Taguchi. Origin of deposited material sedimented in a marine bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12): 1604–1613, December 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-250.

Taguchi:1978:LRS

[2753] S. Taguchi and B. T. Hargrave. Loss rates of suspended material sedimented in a marine bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1614–1620, December 1978. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f78-251.

Hargrave:1978:SCO

[2754] B. T. Hargrave. Seasonal changes in oxygen uptake by settled particulate matter and sediments in a marine bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1621–1628, December 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-252.

Walters:1978:MOF

[2755] G. R. Walters and J. A. Plumb. Modified oxidation/fermentation medium for use in identification of bacterial fish pathogens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1629–1630, December 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-253.

Petrie:1978:NMC

[2756] Brian Petrie and Ken Drinkwater. Numerical model of the circulation in an open bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12): 1631–1635, December 1978. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f78-254.

Downes:1978:STD

[2757] Malcolm T. Downes and Hans W. Paerl. Separation of two dissolved reactive phosphorus fractions in lakewater. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1636–1639, December 1978. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f78-255.

Paerl:1978:BAL

[2758] Hans W. Paerl and Malcolm T. Downes. Biological availability of low versus high molecular weight reactive phosphorus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1639–1643, December 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-256.

Moffett:1978:APR

[2759] William L. Moffett and William S. Fisher. Ammonia production rates of Artemia salina under various culture conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 35(12):1643–1648, December 1978. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f78-257.

Stockner:1979:PEV

[2760] John G. Stockner and David D. Cliff. Phytoplankton ecology of Vancouver Harbor. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1): 1–10, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-001.

Cripe:1979:ADA

[2761] C. R. Cripe. An automated device (AGARS) for studying avoidance of pollutant gradients by aquatic organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):11–16, January 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-002.

Naiman:1979:CPO

[2762] R. J. Naiman and J. R. Sedell. Characterization of particulate organic matter transported by some Cascade Mountain streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):17–31, January 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-003.

Hsu:1979:PST

[2763] C. P. Hsu, A. Marchand, Y. Shimizu, and G. G. Sims. Paralytic shellfish toxins in the sea scallop, Placopecten magellanicus, in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):32–36, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-004.

Kelso:1979:ELF

[2764] John R. M. Kelso and John K. Leslie. Entrainment of larval fish by the Douglas Point Generating Station, Lake Huron, in relation to seasonal succession and distribution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):37–41, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-005.

Seibert:1979:DEP

[2765] G. H. Seibert, R. W. Trites, and S. J. Reid. Deepwater exchange processes in the Saguenay Fjord. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):42–53, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-006.

VanWinkle:1979:PAC

[2766] Webster Van Winkle, Bernadette L. Kirk, and Bert W. Rust. Periodicities in Atlantic Coast striped bass (Morone saxatilis) commercial fisheries data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):54–62, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-007.

Emanuel:1979:MRB

[2767] Mary E. Emanuel and Julian J. Dodson. Modification of the rheotropic behavior of male rainbow trout (Salmo gairdneri) by ovarian fluid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):63–68, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-008.

Peden:1979:MVL

[2768] Alex E. Peden. Meristic variation of Lycodapus mandibularis (Pisces: Zoarcidae) and oceanic upwelling on the West Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):69–76, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-009.

Phillips:1979:IDG

[2769] Glenn R. Phillips and Donald R. Buhler. Influences of dieldrin on the growth and body composition of fingerling rainbow trout (Salmo gairdneri) fed Oregon moist pellets or tubificid worms (Tubifex sp.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):77–80, January 1979. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f79-010.

Leithiser:1979:CHV

[2770] Ronald M. Leithiser, Karl F. Ehrlich, and Alan B. Thum. Comparison of a high volume pump and conventional plankton nets for collecting fish larvae entrained in power plant cooling systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):81–84, January 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-011.

Neville:1979:SEE

[2771] C. M. Neville. Sublethal effects of environmental acidification on rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):84–87, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-012.

Wolf:1979:MCM

[2772] Ken Wolf and Maria E. Markiw. Myxosoma cerebralis: a method for staining spores and other stages with silver nitrate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):88–89, January 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-013.

Schmidt:1979:PSD

[2773] Steven P. Schmidt and Edwin W. House. Precocious sexual development in hatchery-reared and laboratory-maintained male steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36 (1):90–93, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-014.

Trider:1979:SGJ

[2774] D. J. Trider, E. G. Mason, and J. D. Castell. Survival and growth of juvenile American lobsters (Homarus americanus) after eyestalk ablation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):93–97, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-015.

Larkin:1979:MYC

[2775] P. A. Larkin. Maybe you can’t get there from here: A foreshortened history of research in relation to management of Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):98–106, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-016.

Giles:1979:MPT

[2776] M. A. Giles. Marine pollutant transfer. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):107–108, January 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-017.

Stewart:1979:BRD

[2777] James E. Stewart. Book review: Disease Diagnosis and Control in North American Marine Aquaculture. Developments in Aquaculture and Fisheries Science, 6. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1): 108, January 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f79-018.

Clarke:1979:IDO

[2778] R. Allyn Clarke. Introductory dynamic oceanography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(1):108–109, January 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-019.

Wiegand:1979:FCR

[2779] R. C. Wiegand and S. Pond. Fluctuations of chlorophyll and related physical parameters in British Columbia coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):113–121, February 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-020.

Coote:1979:DNG

[2780] A. R. Coote and P. A. Yeats. Distribution of nutrients in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):

122–131, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-021.

Symons:1979:EEA

[2781] Philip E. K. Symons. Estimated escapement of Atlantic salmon (Salmo salar) for maximum smolt production in rivers of different productivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):132–140, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-022.

Beamish:1979:DAP

[2782] R. J. Beamish. Differences in the age of Pacific hake (Merluccius productus) using whole otoliths and sections of otoliths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):141–151, February 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-023.

Stauffer:1979:ECE

[2783] R. E. Stauffer, G. F. Lee, and David E. Armstrong. Estimating chlorophyll extraction biases. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):152–157, February 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-024.

Geraci:1979:TSD

[2784] J. R. Geraci and D. J. St. Aubin. Tissue sources and diagnostic value of circulating enzymes in cetaceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):158–163, February 1979. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f79-025.

Broderius:1979:LSE

[2785] Steven J. Broderius and Lloyd L. Smith, Jr. Lethal and sublethal effects of binary mixtures of cyanide and hexavalent chromium, zinc, or ammonia to the fathead minnow (Pimephales promelas) and rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2): 164–172, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-026.

Helfman:1979:TAY

[2786] Gene S. Helfman. Twilight activities of yellow perch, Perca flavescens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):173–179, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-027.

Sinclair:1979:AML

[2787] Michael Sinclair, Edric Keighan, and Jerry Jones. ATP as a measure of living phytoplankton carbon in estuaries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):180–186, February 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-028.

Sevigny:1979:SPD

[2788] J.-M. Sevigny, M. Sinclair, M. I. El-Sabh, Serge Poulet, and A. Coote. Summer plankton distributions associated with the physical and nutrient properties of the Northwestern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):187–203, February 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-029.

Beschta:1979:IFS

[2789] Robert L. Beschta and William L. Jackson. The intrusion of fine sediments into a stable gravel bed. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):204–210, February 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-030.

Mayzaud:1979:SBC

[2790] P. Mayzaud and S. Taguchi. Spectral and biochemical characteristics of the particulate matter in Bedford Basin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):211–218, February 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-031.

Bilinski:1979:CRF

[2791] E. Bilinski, R. E. E. Jonas, and Y. C. Lau. Control of rancidity in frozen Pacific herring, Clupea harengus pallasi: Use of sodium erythorbate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):219–222, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-032.

Gould:1979:IVS

[2792] R. W. Gould, R. Antipa, and D. F. Amend. Immersion vaccination of sockeye salmon (Oncorhynchus nerka) with two pathogenic strains of Vibrio anguillarum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2): 222–225, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-033.

Chapra:1979:EPL

[2793] Steven C. Chapra and Kenneth H. Reckhow. Expressing the phosphorus loading concept in probabilistic terms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):225–229, February 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-034.

Lingeman:1979:BRM

[2794] B. Lingeman. Book review: Multilingual Dictionary of Fish and Fish Products / Dictionnaire multilingue des poissons et produits de la pêche. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):230, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-035.

Lawrence:1979:BRM

[2795] S. G. Lawrence. Book review: Marine Ecology. Vol. III. Cultivation, part 2. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):230–231, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-036.

Flannagan:1979:IAI

[2796] John F. Flannagan. An introduction to the aquatic insects of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):231, February 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-037.

Kelly:1979:JFD

[2797] Russell K. Kelly. Journal of Fish Diseases, vol. 1, no. 1. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(2):231–232, February 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-038.

Williams:1979:LCD

[2798] J. D. H. Williams and A. E. Pashley. Lightweight corer designed for sampling very soft sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):241–246, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-039.

Harding:1979:USC

[2799] Gareth C. H. Harding and W. Peter Vass. Uptake from seawater and clearance of p,p’-DDT by marine planktonic crustacea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):247–254, March 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-040.

Spouge:1979:RP

[2800] John L. Spouge and P. A. Larkin. A reason for pleomerism. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):255–269, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-041.

Wiener:1979:CCC

[2801] James G. Wiener and John P. Giesy, Jr. Concentrations of Cd, Cu, Mn, Pb, and Zn in fishes in a highly organic softwater pond. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):270–279, March 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-042.

Sawyer:1979:CSA

[2802] Evelyn S. Sawyer, Richard G. Strout, and Bonita A. Coutermarsh. Comparative susceptibility of Atlantic (Salmo salar) and Coho (Oncorhynchus kisutch) salmon to three strains of Vibrio anguillarum from the Maine–New Hampshire Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):280–282, March 1979. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f79-043.

Hunt:1979:IBR

[2803] David W. C. Hunt and J. G. Eales. Iodine balance in rainbow trout (Salmo gairdneri) and effects of testosterone propionate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):282–285, March 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-044.

Schelske:1979:RPG

[2804] Claire L. Schelske. Role of phosphorus in Great Lakes eutrophication: Is there a controversy? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36 (3):286–288, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-045.

Rosenberg:1979:FBI

[2805] David M. Rosenberg. Freshwater benthic invertebrate life histories: Current research and future needs: Introduction. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):289–290, March 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-046.

Resh:1979:SVL

[2806] Vincent H. Resh. Sampling variability and life history features: Basic considerations in the design of aquatic insect studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):290–311, March 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-047.

Malley:1979:SSL

[2807] D. F. Malley and J. B. Reynolds. Sampling strategies and life history of non-insectan freshwater invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):311–318, March 1979. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f79-048.

Oliver:1979:CLH

[2808] D. R. Oliver. Contribution of life history information to taxonomy of aquatic insects. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3): 318–321, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-049.

Buikema:1979:UML

[2809] Arthur L. Buikema, Jr. and Ernest F. Benfield. Use of macroinvertebrate life history information in toxicity tests. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):321–328, March 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-050.

Lehmkuhl:1979:EDL

[2810] D. M. Lehmkuhl. Environmental disturbance and life histories: Principles and examples. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36 (3):329–334, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-051.

Anderson:1979:IDL

[2811] N. H. Anderson and Kenneth W. Cummins. Influences of diet on the life histories of aquatic insects. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):335–342, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-052.

Waters:1979:BLH

[2812] T. F. Waters. Benthic life histories: Summary and future needs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(3):342–345, March 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-053.

Cuff:1979:OSD

[2813] W. Cuff and N. Coleman. Optimal survey design: Lessons from a stratified random sample of macrobenthos. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):351–361, April 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-054.

Vanderploeg:1979:TEI

[2814] H. A. Vanderploeg and D. Scavia. Two electivity indices for feeding with special reference to zooplankton grazing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):362–365, April 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-055.

Spigarelli:1979:TSE

[2815] S. A. Spigarelli and M. M. Thommes. Temperature selection and estimated thermal acclimation by rainbow trout (Salmo gairdneri) in a thermal plume. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4): 366–376, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-056.

Kimura:1979:MVV

[2816] Daniel K. Kimura, Ruth R. Mandapat, and Sandra L. Oxford. Method, validity, and variability in the age determination of yellowtail rockfish (Sebastes flavidus), using otoliths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):377–383, April 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-057.

Hilborn:1979:ABC

[2817] Ray Hilborn and Max Ledbetter. Analysis of the British Columbia salmon purse-seine fleet: Dynamics of movement. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):384–391, April 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-058.

Ringler:1979:SPD

[2818] Neil H. Ringler. Selective predation by drift-feeding brown trout (Salmo trutta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):392– 403, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-059.

Mohn:1979:SAT

[2819] R. K. Mohn. Sensitivity analysis of two harp seal (Pagophilus groenlandicus) population models. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):404–410, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-060.

Maki:1979:CBD

[2820] A. W. Maki. Correlations between Daphnia magna and fathead minnow (Pimephales promelas) chronic toxicity values for several classes of test substances. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4): 411–421, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-061.

MacCrimmon:1979:WDA

[2821] Hugh R. MacCrimmon and Barra L. Gots. World distribution of Atlantic salmon, Salmo solar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4): 422–457, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-062.

Park:1979:CTU

[2822] Eun Ho Park and Sang Dai Park. A culture technique using marine fish kidney to obtain chromosomes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):458–461, April 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-063.

Brewster:1979:VMA

[2823] Frank Brewster and Bruce L. Nicholson. In vitro maintenance of amoebocytes from the American oyster (Crassostrea virginica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):461–467, April 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-064.

Bird:1979:PWP

[2824] Franklin H. Bird and Kenneth Roberson. Pygmy whitefish, Prosopium coulteri, in three lakes of the Copper River System in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):468–470, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-065.

Paloheimo:1979:IFT

[2825] J. E. Paloheimo. Indices of food type preference by a predator. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):470–473, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-066.

Cauvin:1979:CRM

[2826] Dennis M. Cauvin. Coastal resources management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):474, April 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-067.

Sutterlin:1979:AE

[2827] A. M. Sutterlin. Agricultural engineering. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):474–475, April 1979. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f79-068.

Addison:1979:PEM

[2828] R. F. Addison. Pollutant effects on marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):475, April 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-069.

Carey:1979:PAD

[2829] T. G. Carey. Planning of aquaculture development — an introductory guide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(4):476, April 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-070.

Diana:1979:TME

[2830] James S. Diana and W. C. Mackay. Timing and magnitude of energy deposition and loss in the body, liver, and gonads of northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):481– 487, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-071.

Healey:1979:DJS

[2831] M. C. Healey. Detritus and juvenile salmon production in the Nanaimo Estuary: I. Production and feeding rates of juvenile chum salmon (Oncorhynchus keta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5): 488–496, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-072.

Sibert:1979:DJS

[2832] J. R. Sibert. Detritus and juvenile salmon production in the Nanaimo Estuary: II. Meiofauna available as food to juvenile chum salmon (Oncorhynchus keta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5): 497–503, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-073.

Naiman:1979:DJS

[2833] R. J. Naiman and J. R. Sibert. Detritus and juvenile salmon production in the Nanaimo Estuary: III. Importance of detrital carbon to the estuarine ecosystem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5): 504–520, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-074.

Perkins:1979:NEB

[2834] J. S. Perkins and P. C. Beamish. Net entanglements of baleen whales in the inshore fishery of Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):521–528, May 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-075.

Hughes:1979:SSS

[2835] T. G. Hughes. Studies on the sediment of St. Margaret’s Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):529– 536, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-076.

Elner:1979:PSS

[2836] Robert W. Elner and Glen S. Jamieson. Predation of sea scallops, Placopecten magellanicus, by the rock crab, Cancer irroratus, and the American lobster, Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):537–543, May 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-077.

Redding:1979:PAS

[2837] J. M. Redding and C. B. Schreck. Possible adaptive significance of certain enzyme polymorphisms in steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):544–551, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-078.

Favro:1979:PGS

[2838] L. D. Favro, P. K. Kuo, and J. F. McDonald. Population-genetic study of the effects of selective fishing on the growth rate of trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):552–561, May 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-079.

Yamamoto:1979:IPN

[2839] T. Yamamoto and J. Kilistoff. Infectious pancreatic necrosis virus: Quantification of carriers in lake populations during a 6-year period. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):562–567, May 1979. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f79-080.

Demers:1979:STC

[2840] S. Demers, P. E. Lafleur, L. Legendre, and C. L. Trump. Short-term covariability of chlorophyll and temperature in the St. Lawrence Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):568–573, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-081.

Idler:1979:QVA

[2841] D. R. Idler, S. J. Hwang, and L. W. Crim. Quantification of vitellogenin in Atlantic salmon (Salmo salar) plasma by radioimmunoassay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):574–578, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-082.

Conway:1979:SCE

[2842] H. Lee Conway and Susan C. Williams. Sorption of cadmium and its effect on growth and the utilization of inorganic carbon and phosphorus of two freshwater diatoms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):579–586, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-083.

Swanson:1979:PIP

[2843] Robert N. Swanson and James H. Gillespie. Pathogenesis of infectious pancreatic necrosis in Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):587–591, May 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-084.

Woodward:1979:RDS

[2844] John J. Woodward, Edmundo Casillas, Lynwood S. Smith, and Brian G. D’Aoust. Rapid decompression stress accelerates fibrinolysis in fingerling salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5): 592–594, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-085.

Anonymous:1979:ONa

[2845] Anonymous. Obituary/nécrologie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):595–596, May 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-086.

Silvert:1979:CME

[2846] William Silvert. A coastal marine ecosystem. Simulation and analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(5):597–598, May 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-087.

Wedemeyer:1979:PBA

[2847] Gary A. Wedemeyer, Nancy C. Nelson, and William T. Yasutake. Physiological and biochemical aspects of ozone toxicity to rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):605– 614, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-088.

Levitan:1979:PSD

[2848] William M. Levitan and Malcolm H. Taylor. Physiology of salinitydependent naphthalene toxicity in Fundulus heteroclitus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):615–620, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-089.

Hillaby:1979:AAT

[2849] Betty A. Hillaby and David J. Randall. Acute ammonia toxicity and ammonia excretion in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):621–629, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-090.

Loudenslager:1979:KER

[2850] Eric J. Loudenslager and Gary H. Thorgaard. Karyotypic and evolutionary relationships of the Yellowstone (Salmo clarki bouvieri) and West-Slope (S. c. lewisi) cutthroat trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):630–635, June 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-091.

Anderson:1979:PFC

[2851] D. P. Anderson, B. S. Roberson, and O. W. Dixon. Plaque-forming cells and humoral antibody in rainbow trout (Salmo gairdneri) induced by immersion in a Yersinia ruckeri O-antigen preparation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):636–639, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-092.

Jeffries:1979:DPL

[2852] D. S. Jeffries, C. M. Cox, and P. J. Dillon. Depression of pH in lakes and streams in Central Ontario during snowmelt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):640–646, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-093.

Wood:1979:PDS

[2853] C. C. Wood, K. S. Ketchen, and R. J. Beamish. Population dynamics of spiny dogfish (Squalus acanthias) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):647–656, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-094.

Stockner:1979:PES

[2854] J. G. Stockner, D. D. Cliff, and K. R. S. Shortreed. Phytoplankton ecology of the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):657–666, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-095.

Yeo:1979:ICE

[2855] R. K. Yeo and M. J. Risk. Intertidal catastrophes: Effect of storms and hurricanes on intertidal benthos of the Minas Basin, Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):667–669, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-096.

Armstrong:1979:DMC

[2856] F. A. J. Armstrong and D. P. Scott. Decrease in mercury content of fishes in Ball Lake, Ontario, since imposition of controls on mercury discharges. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):670–672, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-097.

Chatarpaul:1979:RTW

[2857] L. Chatarpaul, J. B. Robinson, and N. K. Kaushik. Role of tubificid worms on nitrogen transformations in stream sediment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):673–678, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-098.

Tan:1979:CIR

[2858] F. C. Tan and P. M. Strain. Carbon isotope ratios of particulate organic matter in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):678–682, June 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-099.

Gillespie:1979:CGK

[2859] D. C. Gillespie and R. E. Evans. Composition of granules from kidneys of rainbow trout (Salmo gairdneri) with nephrocalcinosis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):683–685, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-100.

Weiler:1979:RLA

[2860] R. R. Weiler. Rate of loss of ammonia from water to the atmosphere. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):685–689, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-101.

Kaiser:1979:CET

[2861] Ivan I. Kaiser, Patricia A. Young, and Jerry D. Johnson. Chronic exposure of trout to waters with naturally high selenium levels: Effects on transfer RNA methylation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):689–694, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-102.

Grossnickle:1979:DEM

[2862] Nevin E. Grossnickle and Mark D. Morgan. Density estimates of Mysis relicta in Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):694–698, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-103.

Mitchell:1979:BCF

[2863] C. L. Mitchell. Bioeconomics of commercial fisheries management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):699–704, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-104.

Kelso:1979:PPE

[2864] John R. M. Kelso. Power plant entrainment. A biological assessment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):705, June 1979. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f79-105.

Sibert:1979:BBO

[2865] John R. Sibert. Biology of benthic organisms. 11th European Symposium on Marine Biology, Galway, October, 1976. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):706, June 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-106.

Faber:1979:DFM

[2866] Daniel J. Faber and Don E. McAllister. Development of fishes of the Mid-Atlantic Bight. An atlas of egg, larval and juvenile stages. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6):706–707, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-107.

McAllister:1979:BRA

[2867] Don E. McAllister. Book review: Anatomie de Latimeria chalumnae. Tome III. Appareil digestif — appareil respiratoire — appareil urogénital — glandes endocrines — appareil circulatoire — teguments — écailles — conclusions générales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(6): 707–708, June 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-108.

Pearse:1979:ISM

[2868] Peter H. Pearse. Introduction to the Symposium on Managing Fishing Effort. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):711– 714, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-109.

Needler:1979:ECF

[2869] A. W. H. Needler. Evolution of Canadian fisheries management towards economic rationalization. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):716–724, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-110.

Scott:1979:DET

[2870] Anthony Scott. Development of economic theory on fisheries regulation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):725–741, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-111.

Crutchfield:1979:ESI

[2871] J. A. Crutchfield. Economic and social implications of the main policy alternatives for controlling fishing effort. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):742–752, July 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-112.

Fraser:1979:LEE

[2872] G. Alex Fraser. Limited entry: Experience of the British Columbia salmon fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7): 754–763, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-113.

Pearse:1979:ICP

[2873] Peter H. Pearse and James E. Wilen. Impact of Canada’s Pacific salmon fleet control program. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7): 764–769, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-114.

Adasiak:1979:AEL

[2874] A. Adasiak. Alaska’s experience with limited entry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):770–782, July 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-115.

Rogers:1979:ALE

[2875] George W. Rogers. Alaska’s limited entry program: Another view. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):783–788, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-116.

Meany:1979:LEW

[2876] T. F. Meany. Limited entry in the Western Australian rock lobster and prawn fisheries: An economic evaluation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):789–798, July 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-117.

Kaczynski:1979:RAF

[2877] Wlodzimierz Kaczynski. Responses and adjustments of foreign fleets to controls imposed by coastal nations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):800–810, July 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-118.

MacKenzie:1979:RFM

[2878] W. C. MacKenzie. Rational fishery management in a depressed region: the Atlantic groundfishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):811–826, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-119.

Cauvin:1979:RAC

[2879] D. M. Cauvin. Regulating access in Canada’s inland fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):827–836, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-120.

Dickie:1979:PFB

[2880] L. M. Dickie. Perspectives on fisheries biology and implications for management. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):838– 844, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-121.

Huppert:1979:IMF

[2881] D. D. Huppert. Implications of multipurpose fleets and mixed stocks for control policies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7): 845–854, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-122.

Wilen:1979:FBD

[2882] James E. Wilen. Fisherman behavior and the design of efficient fisheries regulation programs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7): 855–858, July 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-123.

Moloney:1979:QRI

[2883] David G. Moloney and Peter H. Pearse. Quantitative rights as an instrument for regulating commercial fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(7):859–866, July 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-124.

Banse:1979:SPP

[2884] Karl Banse. Sabellidae (Polychaeta) principally from the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):869–882, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-125.

Persson:1979:SMO

[2885] Per-Edvin Persson. The source of muddy odor in bream (Abramis brama) from the Porvoo Sea area (Gulf of Finland). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):883–890, August 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-126.

Marten:1979:IFI

[2886] Gerald G. Marten. Impact of fishing on the inshore fishery of Lake Victoria (East Africa). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36 (8):891–900, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-127.

VandenBerg:1979:MCM

[2887] C. M. G. Van den Berg, P. T. S. Wong, and Y. K. Chau. Measurement of complexing materials excreted from algae and their ability to ameliorate copper toxicity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):901– 905, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f79-128.

Massalski:1979:SSC

[2888] A. Massalski and Gary G. Leppard. Survey of some Canadian lakes for the presence of ultrastructurally discrete particles in the colloidal size range. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):906–921, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-129.

Massalski:1979:MEF

[2889] A. Massalski and Gary G. Leppard. Morphological examination of fibrillar colloids associated with algae and bacteria in lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):922–938, August 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-130.

Mendelssohn:1979:DBT

[2890] Roy Mendelssohn. Determining the best trade-off between expected economic return and the risk of undesirable events when managing a randomly varying population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):939–947, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-131.

Sutterlin:1979:SBS

[2891] A. M. Sutterlin, K. J. Jokola, and B. Holte. Swimming behavior of salmonid fish in ocean pens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):948–954, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-132.

Thompson:1979:FRE

[2892] R. J. Thompson. Fecundity and reproductive effort in the blue mussel (Mytilus edulis), the sea urchin (Strongylocentrotus droebachiensis), and the snow crab (Chionoecetes opilio) from populations in Nova Scotia and Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36 (8):955–964, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-133.

Leavitt:1979:DIN

[2893] D. F. Leavitt, R. C. Bayer, M. L. Gallagher, and J. H. Rittenburg. Dietary intake and nutritional characteristics in wild American lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):965–969, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-134.

Peters:1979:CKP

[2894] Robert Henry Peters. Concentrations and kinetics of phosphorus fractions along the trophic gradient of Lake Memphremagog. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):970–979, August 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-135.

Barica:1979:ODW

[2895] Jan Barica and John A. Mathias. Oxygen depletion and winterkill risk in small prairie lakes under extended ice cover. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):980–986, August 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-136.

Macdonald:1979:AGS

[2896] P. D. M. Macdonald and T. J. Pitcher. Age-groups from size-frequency data: A versatile and efficient method of analyzing distribution mixtures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):987–1001, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-137.

Carignan:1979:QSP

[2897] R. Carignan and J. Kalff. Quantification of the sediment phosphorus available to aquatic macrophytes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):1002–1005, August 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-138.

Crecelius:1979:MOO

[2898] Eric A. Crecelius. Measurements of oxidants in ozonized seawater and some biological reactions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8):1006–1008, August 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-139.

Estes:1979:EMM

[2899] James A. Estes. Exploitation of marine mammals: r-selection of Kstrategists? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8): 1009–1017, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-140.

Anonymous:1979:CSRa

[2900] Anonymous. Comments on some recent methods for estimating food consumption by fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8): 1018, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-141.

Anonymous:1979:CSRb

[2901] Anonymous. Comments on some recent methods for estimating food consumption by fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8): 1018–1019, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-142.

Anonymous:1979:CSRc

[2902] Anonymous. Comments on some recent methods for estimating food consumption by fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(8): 1020, August 1979. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f79-143.

Bollingberg:1979:LSW

[2903] H. J. Bollingberg and P. Johansen. Lead in spotted wolffish, Anarhichas minor, near a zinc–lead mine in Greenland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1023–1028, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-144.

Hall:1979:DFB

[2904] Donald J. Hall, Earl E. Werner, James F. Gilliam, Gary G. Mittelbach, Donald Howard, Christopher G. Doner, Joyce A. Dickerman, and Arthur J. Stewart. Diel foraging behavior and prey selection in the golden shiner (Notemigonus crysoleucas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1029–1039, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-145.

Veith:1979:MEB

[2905] Gilman D. Veith, David L. DeFoe, and Barbara V. Bergstedt. Measuring and estimating the bioconcentration factor of chemicals in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1040–1048, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-146.

McLeay:1979:SCE

[2906] Donald J. McLeay and David A. Brown. Stress and chronic effects of untreated and treated bleached kraft pulpmill effluent on the biochemistry and stamina of juvenile Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1049–1059, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-147.

Gray:1979:SMA

[2907] Robert H. Gray and James M. Haynes. Spawning migration of adult Chinook salmon (Oncorhynchus tshawytscha) carrying external and internal radio transmitters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9): 1060–1064, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-148.

Traynor:1979:EDB

[2908] J. J. Traynor and J. E. Ehrenberg. Evaluation of the dual beam acoustic fish target strength measurement method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1065–1071, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-149.

Kendall:1979:STE

[2909] Michael W. Kendall and James E. Dale. Scanning and transmission electron microscopic observations of rainbow trout (Salmo gairdneri)

Gill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1072–1079, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-150.

Peebles:1979:MMD

[2910] John B. Peebles. Molting, movement, and dispersion in the freshwater prawn Macrobrachium rosenbergii. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1080–1088, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-151.

Cherry:1979:RIV

[2911] Donald S. Cherry, Stephan R. Larrick, Rufus K. Guthrie, Ernst M. Davis, and Fred F. Sherberger. Recovery of invertebrate and vertebrate populations in a coal ash stressed drainage system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1089–1096, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-152.

MacMillan:1979:ATS

[2912] John R. MacMillan and Dan Mulcahy. Artificial transmission to and susceptibility of Puget Sound fish to viral erythrocytic necrosis (VEN). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1097–1101, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-153.

Weber:1979:ASA

[2913] J. M. Weber and B. M. Zwicker. Aeromonas salmonicida in Atlantic salmon (Salmo salar): Occurrence of specific agglutinins to three bacterial pathogens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9): 1102–1107, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-154.

Burton:1979:ICT

[2914] Dennis T. Burton, Lenwood W. Hall, Jr., Stuart L. Margrey, and Robert D. Small. Interactions of chlorine, temperature change (Δ T), and exposure time on survival of striped bass (Morone saxatilis) eggs and prolarvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9): 1108–1113, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-155.

May:1979:JSB

[2915] Bernie May, James E. Wright, and Mark Stoneking. Joint segregation of biochemical loci in Salmonidae: Results from experiments with Salvelinus and review of the literature on other species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1114–1128, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-156.

Lee:1979:MRS

[2916] David Robert Lee and Arthur L. Buikema, Jr. Molt-related sensitivity of Daphnia pulex in toxicity testing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1129–1133, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-157.

Fong:1979:DMI

[2917] W. Fong and R. K. O’Dor. D-mannitol interferes with amino acid analysis of marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9): 1134–1137, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-158.

Peterson:1979:TPS

[2918] R. H. Peterson, A. M. Sutterlin, and J. L. Metcalfe. Temperature preference of several species of Salmo and Salvelinus and some of their hybrids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1137–1140, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-159.

Tomasso:1979:CIN

[2919] J. R. Tomasso, Bill A. Simco, and Kenneth B. Davis. Chloride inhibition of nitrite-induced methemoglobinemia in channel catfish (Ictalurus punctatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1141–1144, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-160.

Thorne:1979:HEA

[2920] Richard E. Thorne. Hydroacoustic estimates of adult sockeye salmon (Oncorhynchus nerka) in Lake Washington, 1972–75. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1145–1149, September 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-161.

Carmichael:1979:MMK

[2921] N. G. Carmichael, K. S. Squibb, and B. A. Fowler. Metals in the molluscan kidney: A comparison of two closely related bivalve species (Argopecten), using X-ray microanalysis and atomic absorption spectroscopy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1149–1155, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-162.

Patalas:1979:BRB

[2922] K. Patalas. Book review: Die Binnengewässer Band XXVI. Das Zooplankton der Binnengewässer. Teil 2. (German) [The Inland Waters. Volume XXVI. The inland zooplankton. Part 2]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1156, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-163.

Clarke:1979:ACO

[2923] R. McV. Clarke. The aqueous chlorination of organic compounds: Chemical reactivity and effects on environmental quality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1157, September 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-164.

Dyer:1979:FPC

[2924] W. J. Dyer. Fish protein concentrate: Panacea for protein malnutrition? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1157, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-165.

Anonymous:1979:CBA

[2925] Anonymous. Comment on “Bias in age–length keys”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1159–1160, September 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-166.

Anonymous:1979:RCB

[2926] Anonymous. Reply to comment on “Bias in age–length keys”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(9):1160–1161, September 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-167.

Makarewicz:1979:CAT

[2927] J. C. Makarewicz, R. I. Baybutt, and K. Damann. Changes in the apparent temperature optima of the plankton of Lake Michigan at Chicago, Illinois. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1169–1173, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-168.

Trust:1979:OAB

[2928] T. J. Trust, L. M. Bull, B. R. Currie, and J. T. Buckley. Obligate anaerobic bacteria in the gastrointestinal microflora of the grass carp (Ctenopharyngodon idella), goldfish (Carassius auratus), and rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1174–1179, October 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-169.

Collins:1979:REM

[2929] J. J. Collins. Relative efficiency of multifilament and monofilament nylon gill net towards lake whitefish (Coregonus clupeaformis) in Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1180–1185, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-170.

Hara:1979:OBE

[2930] Toshiaki J. Hara and S. B. Brown. Olfactory bulbar electrical responses of rainbow trout (Salmo gairdneri) exposed to morpholine during smoltification. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1186–1190, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-171.

Gowing:1979:IBT

[2931] H. Gowing and W. T. Momot. Impact of brook trout (Salvelinus fontinalis) predation on the crayfish Orconectes virilis in three Michigan lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1191–1196, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-172.

Wildish:1979:TES

[2932] D. J. Wildish and D. D. Kristmanson. Tidal energy and sublittoral macrobenthic animals in estuaries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1197–1206, October 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-173.

Jensen:1979:ECA

[2933] Arne Johan Jensen. Energy content analysis from weight and liver index measurements of immature Pollock (Pollachius virens). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1207–1213, October 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-174.

Bodaly:1979:MED

[2934] R. A. Bodaly. Morphological and ecological divergence within the lake whitefish (Coregonus clupeaformis) species complex in Yukon Territory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1214–1222, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-175.

Lemon:1979:SEP

[2935] D. D. Lemon, P. H. LeBlond, and T. R. Osborn. Seiche excitation in Port San Juan, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1223–1227, October 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-176.

Fournier:1979:ISB

[2936] R. O. Fournier, M. Van Det, J. S. Wilson, and N. B. Hargreaves. Influence of the shelf-break front off Nova Scotia on phytoplankton standing stock in winter. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1228–1237, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-177.

Freeman:1979:GMB

[2937] K. R. Freeman and L. M. Dickie. Growth and mortality of the blue mussel (Mytilus edulis) in relation to environmental indexing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1238–1249, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-178.

Pocklington:1979:TOM

[2938] R. Pocklington and J. D. Leonard. Terrigenous organic matter in sediments of the St. Lawrence Estuary and the Saguenay Fjord. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1250–1255, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-179.

Borgmann:1979:RME

[2939] U. Borgmann, C. Loveridge, and R. Cove. A rapid method for the estimation of, and some factors affecting, copepod production rates in the burlington canal. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10): 1256–1264, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-180.

Blackett:1979:ESO

[2940] R. F. Blackett. Establishment of sockeye (Oncorhynchus nerka) and Chinook (O. tshawytscha) salmon runs at Frazer Lake, Kodiak Island, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1265–1277, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-181.

Geraci:1979:IAC

[2941] J. R. Geraci, D. J. St. Aubin, and T. G. Smith. Influence of age, condition, sampling time, and method on plasma chemical constituents in free-ranging ringed seals, Phoca hispida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1278–1282, October 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-182.

Berg:1979:EMP

[2942] Richard E. Berg. External morphology of the pink salmon, Oncorhynchus gorbuscha, introduced into Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1283–1287, October 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-183.

Lau:1979:LSC

[2943] Y. L. Lau. Laboratory study of cylindrical sedimentation traps. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1288–1291, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-184.

Anonymous:1979:ONb

[2944] Anonymous. Obituary/nécrologie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1292–1293, October 1979. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f79-185.

Klaverkamp:1979:EPA

[2945] J. F. Klaverkamp. Effects of pollutants on aquatic organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(10):1294, October 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f79-186.

Roch:1979:RCI

[2946] M. Roch and E. J. Maly. Relationship of cadmium-induced hypocalcemia with mortality in rainbow trout (Salmo gairdneri) and the influence of temperature on toxicity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1297–1303, November 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-187.

Clark:1979:BMG

[2947] Colin W. Clark and Geoffrey P. Kirkwood. Bioeconomic model of the Gulf of Carpentaria prawn fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1304–1312, November 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-188.

Buroker:1979:OMP

[2948] N. E. Buroker. Overdominance of a muscle protein (Mp-1) locus in the Japanese oyster, Crassostrea gigas (Ostreidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1313–1318, November 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-189.

Wickham:1979:CEF

[2949] D. E. Wickham. Carcinonemertes errans and the fouling and mortality of eggs of the Dungeness crab, Cancer magister. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1319–1324, November 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-190.

Fortier:1979:CVC

[2950] L. Fortier and L. Legendre. Le contrôle de la variabilité à court terme du phytoplancton estuarien: stabilité verticale et profondeur critique. (French) [Controlling the short-term variability of estuarine phytoplankton: vertical stability and critical depth]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1325–1335, November 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-191.

Scavia:1979:EPC

[2951] Donald Scavia. Examination of phosphorus cycling and control of phytoplankton dynamics in Lake Ontario with an ecological model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1336–1346, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-192.

Ewing:1979:TPE

[2952] R. D. Ewing, S. L. Johnson, H. J. Pribble, and J. A. Lichatowich. Temperature and photoperiod effects on gill (Na + K)–ATPase activity in Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1347–1353, November 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-193.

Sprules:1979:BSF

[2953] W. Gary Sprules and L. Blair Holtby. Body size and feeding ecology as alternatives to taxonomy for the study of limnetic zooplankton Community Structure. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11): 1354–1363, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-194.

Healey:1979:IPN

[2954] F. P. Healey and L. L. Hendzel. Indicators of phosphorus and nitrogen deficiency in five algae in culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1364–1369, November 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-195.

Mitchum:1979:BKD

[2955] Douglas L. Mitchum, Loris E. Sherman, and George T. Baxter. Bacterial kidney disease in feral populations of brook trout (Salvelinus fontinalis), brown trout (Salmo trutta), and rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1370–1376, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-196.

Rich:1979:DCB

[2956] P. H. Rich. Differential CO_{2} and O_{2} benthic community metabolism in a soft-water lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11): 1377–1389, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-197.

Plumb:1979:FVD

[2957] J. A. Plumb, P. R. Bowser, J. M. Grizzle, and A. J. Mitchell. Fish viruses: A double-stranded RNA icosahedral virus from a North American cyprinid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11): 1390–1394, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-198.

Beamish:1979:NIL

[2958] Richard J. Beamish. New information on the longevity of Pacific Ocean perch (Sebastes alutus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1395–1400, November 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-199.

Stegeman:1979:TIB

[2959] John J. Stegeman. Temperature influence on basal activity and induction of mixed function oxygenase activity in Fundulus heteroclitus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1400–1405, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-200.

Grinnell:1979:RCI

[2960] Brian Grinnell and J. C. Leong. Recovery and concentration of infectious pancreatic necrosis (IPN) virus in water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1405–1408, November 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-201.

Sosiak:1979:FHR

[2961] A. J. Sosiak, R. G. Randall, and J. A. McKenzie. Feeding by hatcheryreared and wild Atlantic salmon (Salmo salar) Parr in streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1408–1412, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-202.

Liem:1979:DTB

[2962] A. J. Liem and V. A. Naish. Difference in taste between the right and left sides of tainted fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11): 1412–1413, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-203.

Soevik:1979:FAK

[2963] T. Soevik and O. R. Braekkan. Fluoride in Antarctic krill (Euphausia superba) and Atlantic krill (Meganyctiphanes norvegica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1414–1416, November 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-204.

Anonymous:1979:CMA

[2964] Anonymous. Comment on management alternatives and strategic planning for Canada’s fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1417, November 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-205.

Anonymous:1979:RCM

[2965] Anonymous. Reply to comment on management alternatives and strategic planning for Canada’s fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(11):1417–1418, November 1979. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f79-206.

Anonymous:1979:E

[2966] Anonymous. Editorial. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1423–1424, December 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-207.

Waters:1979:IBL

[2967] T. F. Waters. Influence of benthos life history upon the estimation of secondary production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12): 1425–1430, December 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-208.

Swift:1979:ZPD

[2968] Michael C. Swift and U. Theodore Hammer. Zooplankton population dynamics and Diaptomus production in Waldsea Lake, a saline meromictic lake in Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12): 1431–1438, December 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-209.

Moreau:1979:DPA

[2969] Guy Moreau and Claude Barbeau. Différenciation de populations anadromes et dulcicoles de Grands Corégones (Coregonus clupeaformis) par la composition minérale de leurs écailles. (French) [differentiation of anadromous and freshwater populations of large coregonids (Coregonus clupeaformis) by the mineral composition of their scales]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1439–1444, December 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-210.

Herman:1979:IVM

[2970] Alex W. Herman and Kenneth L. Denman. Intrusions and vertical mixing at the shelf/slope water front south of Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1445–1453, December 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-211.

Downing:1979:ATD

[2971] John A. Downing. Aggregation, transformation, and the design of benthos sampling programs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1454–1463, December 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-212.

Paterson:1979:DBK

[2972] W. D. Paterson, C. Gallant, D. Desautels, and L. Marshall. Detection of bacterial kidney disease in wild salmonids in the Margaree River System and adjacent waters using an indirect fluorescent antibody technique. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1464–1468, December 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-213.

Dunn:1979:PAH

[2973] Bruce P. Dunn and John Fee. Polycyclic aromatic hydrocarbon carcinogens in commercial seafoods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1469–1476, December 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-214.

Hilborn:1979:CFC

[2974] Ray Hilborn. Comparison of fisheries control systems that utilize catch and effort data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12): 1477–1489, December 1979. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f79-215.

Foote:1979:RLD

[2975] Kenneth G. Foote. On representing the length dependence of acoustic target strengths of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1490–1496, December 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-216.

Buckley:1979:RBR

[2976] L. J. Buckley. Relationships between RNA–DNA ratio, prey density, and growth rate in Atlantic cod (Gadus morhua) larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1497–1502, December 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-217.

Hughes:1979:MLF

[2977] T. G. Hughes. Mode of life and feeding in maldanid polychaetes from St. Margaret’s Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1503–1507, December 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-218.

Hawke:1979:BAD

[2978] John P. Hawke. A bacterium associated with disease of pond cultured channel catfish, Ictalurus punctatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1508–1512, December 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-219.

Levy:1979:INB

[2979] E. M. Levy. Intercomparison of niskin and blumer samplers for the study of dissolved and dispersed petroleum residues in seawater. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1513–1516, December 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-220.

Phillips:1979:AED

[2980] G. R. Phillips and R. W. Gregory. Assimilation efficiency of dietary methylmercury by northern pike (Esox lucius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1516–1519, December 1979. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f79-221.

Lee:1979:WEM

[2981] Richard F. Lee and Judd C. Nevenzel. Wax esters in the marine environment: Origin and composition of the wax from Bute Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1519–1523, December 1979. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f79-222.

Anonymous:1979:ONc

[2982] Anonymous. Obituary/nécrologie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1524–1526, December 1979. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f79-223.

Anonymous:1979:CPM

[2983] Anonymous. Comment on production, mortality, and sustainable yield of Northwest Atlantic harp seals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1527–1530, December 1979. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f79-224.

Anonymous:1979:RCP

[2984] Anonymous. Reply to comment on production, mortality, and sustainable yield of Northwest Atlantic harp seals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1530–1532, December 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-225.

Ayles:1979:STF

[2985] G. Burton Ayles. Salmon and trout farming in Norway. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1533, December 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-226.

Wegemann:1979:BRC

[2986] R. Wegemann. Book review: Chemical Modelling in Aqueous Systems. Speciation, Sorption, Solubility, and Kinetics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 36(12):1533–1534, December 1979. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f79-227.